A Complete Bibliography of Publications in Isis, 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 0.12

Title word cross-reference

ΣT OIXEIA [916].

-Tradition [1677].

1000 [1827, 1112]. 1100 [1345]. 116v [1799]. 1307 [179]. 1330 [1592]. 1471 [2586]. 1484 [1958]. 150 [1024]. 1500 [2214, 2126, 2818]. 1522 [437]. 1565 [37]. 1571 [2893]. 1599 [1280]. 15th [1676]. 15th-century [1676]. 16.-20 [916]. 1600 [918, 2894]. 1612 [122, 724]. 1640 [1087, 1474]. 1646 [578]. 1648 [44, 2129]. 1650 [1604, 612]. 1656 [1856]. 1664/5 [1536]. 1670 [2400]. 1672 [1733]. 1690s [2429]. 16th [510]. 17 [332]. 17.-20 [916]. 1700 [380, 2371, 910, 121]. 1700-talet [910]. 1702 [2904]. 1708 [1150]. 1722 [1136]. 1736 [1928]. 1750 [789, 1224]. 1753 [2081]. 1760 [2415]. 1776 [2908]. 1780 [2402, 2347]. 1782 [2284]. 1789 [71]. 1790 [2216]. 1791 [1369]. 1793 [446]. 1799 [2338]. 17th [510, 623, 1601]. 1800 [347, 468, 952, 2907, 149, 486, 2037, 2334]. 1800. [1091]. 1804 [1933]. 1805 [1413, 359]. 1809 [1188]. 1810 [2221, 2747]. 1814 [615]. 1815 [1465, 2825]. 1825 [843]. 1830 [2361, 1496]. 1831 [682]. 1832 [2641]. 1838 [1610]. 1840 [2887, 1482, 448, 1807]. 1842 [2406, 2105]. 1843 [1322, 2057]. 1848 [430]. 1850 [834, 861, 1017, 1312, 1579]. 1854 [336, 2207, 1913, 239]. 1855 [1667]. 1858 [163, 1497]. 1859 [1237, 1993, 927]. 1860 [2888, 1690, 1773, 2197, 2158]. 1860s [2385]. 1861 [1577]. 1863 [2289]. 1864 [959, 1498]. 1865 [175]. 1866 [2752, 928, 367]. 1867 [1988]. 1868 [1992]. 1869 [2702]. 1870 [728, 2455, 1765, 1840, 2643, 1374]. 1871 [1864]. 1872 [2702]. 1875 [1118]. 1876 [2653]. 1878 [1562, 1951, 1901]. 1880 [1841, 364, 2031]. 1882 [1628]. 1883 [1641, 1769]. 1884 [1504]. 1885 [2268, 2410]. 1886 [2744]. 1887 [2367]. 1888 [490]. 1890 [1321, 1650, 1835]. 1890s [2169]. 1894 [2703, 2132]. 1895 [2642, 2512]. 1899 [1051]. 18th [565, 1601]. 1900 [2834, 2096, 2248, 1164, 483, 2582, 2552, 2029, 160, 1476, 2434, 1569, 1821, 695]. 1902 [2654]. 1904 [2031, 2857]. 1905 [568, 1027]. 1909 [1921]. 1910 [2562, 2923, 296, 2444, 2142, 633, 630]. 1911 [1415]. 1912 [2251]. 1914 [477, 757, 1332, 1566, 226, 2830, 491, 2485, 2869, 1324, 584, 1900, 480, 1002]. 1915 [126, 294, 1738]. 1916 [1865]. 1918 [2566, 1247]. 1920 [1897, 1308, 148]. 1920s [2619]. 1921 [2777]. 1925 [989]. 1926 [708, 1923]. 1929 [451, 753]. 1930 [1946, 2048, 1272, 1245, 2511, 1261, 2424, 2770, 2116, 392, 1139]. 1930s [2619, 2294, 61, 323, 87]. 1932 [1771, 951]. 1933 [133, 55, 2932]. 1934 [1055]. 1935 [1753, 867, 517]. 1936 [1959, 440]. 1937 [2180, 2087, 933]. 1939 [2246, 2185, 2476, 817, 1912, 1866, 232]. 1940 [1819, 896, 1271, 2577, 2173, 1093, 1754, 349, 1086]. 1941 [934, 2507, 104]. 1944 [1427, 137, 2188]. 1945 [189, 515, 2927, 1529]. 1946 [1024, 2035, 658, 2377, 2759]. 1947 [1741]. 1948 [710]. 1950 [1554, 796, 2201, 2647, 1113, 1372, 2868, 1732]. 1951 [1500]. 1952 [210, 344]. 1953 [2853]. 1954 [2062, 2649, 2137]. 1955 [767, 2181]. 1956 [2386, 105]. 1957 [1096]. 1958 [2427, 2381]. 1960 [1167, 1898]. 1962 [1694]. 1963 [18]. 1965 [2088, 83]. 1966 [2113, 210]. 1967 [81]. 1968 [2843]. 1969 [1425]. 1970 [134, 2408]. 1972 [52]. 1975 [288, 108, 1544, 1573]. 1976 [184, 19, 46]. 1977 [221, 225, 345, 544, 985, 507, 496]. 1978 [474, 1693, 824, 513, 85, 913, 472, 410]. 1979 [64, 66, 9, 264, 289, 632, 614]. 1980 [465, 748, 464, 265, 332, 334, 2937, 697, 2299, 2849, 2861, 416, 307, 1210, 2783, 950]. 1981 [811, 670, 672, 671, 812, 669, 537, 1300, 1229, 1418, 768, 1011, 1228, 1251, 749]. 1982 [881, 882, 1001, 944, 1589, 1739, 1364, 1315]. 1983 [1135, 1000, 1644, 1133, 1879, 2281, 1990, 1760]. 1984 [2155, 1393, 1457, 1627, 1626, 1529, 2139]. 1985 [1804, 1709, 1878, 1942, 1801, 2661, 2017, 2558, 1822, 2707, 2428, 1825, 2504]. 1986 [2070, 2876, 2913, 2935, 2426, 2015]. 1987 [2716, 2474, 2387, 2718]. 1988 [2473, 2791]. 19th [281, 623, 285].

2 [2555, 29]. 20 [86]. 2000 [2724]. 20th [949]. 21-year [738]. 23 [2070]. 25-jährigen [2807]. 250jährigen [2489]. 27 [1133, 1457, 669, 2791]. 28 [881].

29 [2387].

31 [1801]. 360 [1429]. 365-300 [916].

4 [1120]. 400th [1315]. 41 [853]. 4th [283].

’500 [1442, 2160].

60 [691]. ’600 [2160, 1442, 626]. 68 [2225].

’700 [1749]. 72A [1786]. 79 [419].

80 [822]. 88 [2373].

Âge [1343].

A. [3]. A.11 [1439]. A.D [1827, 221, 1413, 1112, 1087, 1279, 239]. A.D. [2694, 635]. A/Part [1893]. Aachen [38]. Aachen-Agypten [38]. AAPT [1894]. Aaron [620, 384]. Aarsleff [1034]. Abakus [2897]. Abbacus [918]. Abbé [1602]. Abbott [2889]. Abdelhamid [1276]. Abell [972]. Abendland [1203]. Abhandlungen [1288]. Abi [1037]. Ability [1819]. Able [2646]. Able-Bodied [2646]. Abragam [2060]. Abraham [634, 1202, 1138]. Abrams [671, 2631]. Abramson [2507, 2765]. abrégé [2407, 2248, 483]. Abroad [2013, 458]. Abstammungs [836]. Abstammungs- [836]. Abu [2586]. Abu’l [1277]. Abu’l-Rayhan [1277]. Abulkasis [2586]. Abuse [2635]. Abusing [887]. Abyss [425, 1806, 1581]. Academic [2922, 814, 1620, 2484, 1473]. academica [1917]. académiciens [71]. Académie [456, 2429]. académies [2804, 71]. Academy [273, 2269, 2692, 478, 922, 18, 1400, 1042, 2067, 1131, 922]. Accademia [1130]. Acceptance [257]. according [976]. Account [246, 1807, 548, 549, 907, 2524, 1581]. Accounts [1940, 48, 1166, 2013]. Accuracy [597]. Achievements [1946, 1193]. Achinstein [1960]. achttiende [649]. Acid [2757]. acidez [442]. Ackermann [1728]. Acloque [726]. Acquisition [2137]. Acta [1882]. Actes [304, 496, 690]. Action [2787, 2324]. Activists [2294]. Activities [1166, 2237]. Activity [2883]. Actors [2537]. actually [1710]. Acupuncture [783]. Ad [1605, 604, 1882, 1681]. Ada [2144]. Adam [511, 2898, 2908, 1190]. Adams [1262, 2120, 199]. Addenda [1379]. Additions [2554]. Addresses [225]. Adel [2048]. Adelaide [2141]. Adelard [1508]. Adenauer [2758]. Adi [133]. Adjustment [563]. Adler [1424]. Administration [1596]. Administrators [2947]. Admirable [1585]. Ado [1140]. Adolph [1983]. Adoption [2005]. Adriaan [2576]. Adrian [1118]. Adrien [2460]. Adrienne [55]. Advancement [1910, 2288, 749]. Advances [1486, 485]. Advantage [2541]. Advent [1826, 2844]. Adventure [1025, 2765, 2774]. Adventures [2531, 704]. Adventurism [1124]. Advice [2756, 418]. Adviser [2949]. Advisory [2381, 1892]. AEG [1566]. Aegean [2123]. Aepinus [408]. Aerodynamics [930]. Aeronautical [2427]. Aeronautics [2381]. Aerospace [52]. Aerostation [1124]. Aesthetics [2335, 2293]. Aether [1522, 2500]. Affair [1784, 2022, 1211]. Affairs [2564]. affrontement [617]. Africa [1166, 1498, 2772]. African [32, 1498]. Africans [364]. After [1484, 425, 55, 1482, 2818, 2228]. Aftermath [6]. Again [1547]. against [1785, 2244, 2625, 887, 2021, 2402, 235]. Agassi [2311, 1155]. Agassiz [1263]. Agazzi [610, 609]. Age [1428, 1460, 2042, 2121, 2012, 1320, 842, 2301, 2117, 2565, 1087, 1641, 1910, 2037, 2435, 964, 752, 1515, 173, 2743, 2358, 1360, 132, 470, 1471, 2709, 2063, 1376, 2457, 2746, 2750, 2580, 1480]. Agency [1338]. Agent [147]. Ages [94, 2276, 1121, 2066, 877, 2213, 37, 2788, 2310, 2448, 1140, 1590, 236, 1307, 1849, 115, 1552, 1437, 600, 1510, 69]. Aggression [191]. Agnes [429, 2400]. Agnese [2338]. Agnosticism [2494]. Agony [2037]. Agricola [2901]. Agricultural [629, 838, 1835, 221, 1345, 1835]. Agriculture [2749, 1661, 1822, 2772, 1679]. Agriculturist [1524]. Agrimi [1675]. Aguilon [1783]. Agypten [38]. Ahead [822]. Ahmad [1277, 1979, 308, 2587, 238]. Ahsan [1279]. Aichelburg [133]. AIDS [2625, 2109, 2856]. Aim [1938]. Aims [431, 2665, 1560]. Ainsworth [765]. Air [2275, 1986, 2935, 2118]. Airplane [568]. Airs [1202]. Airships [2117]. Airway [2853]. Aitken [1771]. Akademie [2726, 891]. akademii [273]. Akademika [660]. akademischen [556]. Aksenov [2523]. Aktualität [475]. Al [2896, 1979, 308, 1844, 2595, 1682, 1439, 2586, 2587, 2814, 238, 1276, 1440, 1277, 504, 308]. Al-Biruni [2896, 1277]. Al-cilm [308]. al-Din [2814]. Al-Hasan [1276]. al-Hashimi [1439]. al-Hassan [2587, 238]. al-Hay [1440]. al-Hayawan [504]. al-Haytham [1844, 1276]. Al-Hiyal [308]. Al-Jamic [308]. Al-Jazari [308]. Al-Khayyam [1979]. al-Manazir [1276]. Al-Nafic [308]. al-Qasim [2586]. Al-Razzaz [308]. al-Tusi [2814]. al-Zahrawi [2586]. al-zijat [1439]. Alain [1857, 1819, 2352, 1319, 2170]. Alamos [515, 1293]. Alan [1827, 2929, 435, 469, 2092, 1447, 1787, 2828, 1835, 1163, 1733, 1869, 2493, 1291]. Alaska [1051]. Albers [2496]. Albert [216, 1559, 2738, 1757, 1192, 1193, 108, 1138, 2507, 2654, 1265, 356, 1317, 1193, 737, 133, 133, 319]. Albertini [566]. Alberto [2727, 1675, 2025, 1502]. Albertus [2695, 1780, 307]. Albrecht [1354, 1158, 2830]. Albritton [425]. Album [468, 1552]. Albury [1240, 381]. Alchemical [2788]. Alchemie [2210]. Alchemy [742, 2371, 1850, 1058, 435, 434, 109]. alchimique [1342]. alchimistes [845]. Alderotti [781]. Aldo [1253]. Aldras [824]. Aldrich [1163]. Aldrig [824]. Aldrovandi [644]. Aleksandr [2832]. Alex [1415, 2381]. Alexander [1249, 1837, 1184, 115, 385, 2488, 1530, 1400, 2033, 1982, 514, 432, 1527]. Alexanderson [2496]. Alexandria [2043, 114]. Alexandrovich [285]. Alfred [386, 757, 1906, 2029, 2444, 2668, 1309, 1831, 2620, 988, 427, 924, 2149, 2142]. Algebra [2004, 1744, 1980, 715, 1091, 281, 283]. Algebraic [878]. Algèbre [2814]. algebrique [1979, 280]. Alguns [2900]. Ali [1594, 1439]. Alic [2019]. Alice [2579, 2107, 2780, 2429, 393, 1614]. Alireza [1278]. Alison [122]. Alkaloide [2498]. alla [1009]. Allan [2564, 1540, 2109, 2585, 1841, 932, 1292, 635, 2666, 2212]. Allard [126, 848]. Alle [647]. Allemagne [2563]. Allen [2371, 2902, 1245, 1346, 1358, 12, 2086]. allgemeine [2483, 1026]. Alliance [2196, 1246]. Allianz [2490]. Allibone [1653]. Allied [962]. Allison [434, 865]. all’ottocento [488]. Almagest [2393, 1432]. Almanacs [493]. Almond [953]. Almost [190]. Aloft [1373]. Alonso [2593]. Alpers [1156]. Alphonse [2804]. Alphonsines [1591]. Alston [541]. Alten [1760]. Alter [1308]. alterblick [824]. Altschule [2885]. Alumni [2786]. Alun [85]. Alva [795]. Alvarez [1987, 2531, 2531]. Alvaro [2593]. Alwyne [768, 1582, 2103]. Am [1055, 985]. amanecer [860]. Amann [1030]. Amateur [2744]. Amateurs [256, 276]. Ambergris [666]. ambigua [2219]. Ambition [1834]. Ambitions [1072]. Ambroise [981, 2350]. ame [712]. Amedeo [1356]. Amelia [672]. America [387, 1809, 2239, 1273, 2326, 2201, 915, 2299, 290, 1647, 148, 2861, 122, 817, 2268, 1010, 2887, 2511, 160, 365, 2765, 1818, 1372, 232, 2302, 2435, 1163, 517, 284, 1086, 2367, 2106, 893, 1828, 49, 1939, 629, 1260, 930, 2175, 2312, 2307, 2176, 186, 2689, 1020, 2294, 1426, 1920, 2517, 1038, 1934, 360, 1311, 2179, 2823, 1270, 1373, 2111, 2889, 1768, 2273, 1562, 1892]. American [49, 294, 15, 34, 1371, 2406, 1427, 2631, 1692, 656, 2862, 225, 1897, 14, 366, 84, 2650, 2013, 620, 1271, 767, 2553, 96, 2510, 2774, 934, 914, 2922, 731, 2561, 2225, 774, 2835, 705, 1173, 127, 2582, 2215, 1978, 2682, 122, 1771, 1865, 2577, 1549, 416, 1923, 816, 2853, 1767, 2088, 2178, 2529, 2780, 2161, 2687, 1210, 2258, 2202, 2690, 542, 1921, 1964, 2475, 1054, 1641, 2034, 1582, 2291, 2478, 1818, 2721, 2781, 2547, 543, 869, 2457, 1223, 1504, 2158, 1649, 1666, 2328, 795, 2653, 2428, 1930, 630, 2718, 2506, 1057, 841, 2265, 2770, 1581, 2197, 297]. American [1197, 1889, 2414, 2509, 2039, 97, 351, 1529, 532, 59, 143, 741, 1221, 940, 1621, 527, 137, 834, 225, 1460, 1690, 2849, 2485, 2780, 1107, 1168, 1641, 2855, 633, 440, 2428, 1317, 898]. Americana [682]. Americans [568]. Americas [1760, 1039]. Amerikanische [1760]. amerikanischen [762]. Amerine [2079]. Amici [1181]. among [1226, 1880]. Amos [2310]. Ampere [1624, 996, 1256]. Amrine [2285]. Amsterdam [9]. Amsterdamse [649]. Amundsen [107]. Anaesthesia [2628]. analisi [693]. Analogies [1128]. Analogy [1126, 1797]. analyseos [1744]. Analyses [80]. Analysis [1408, 1186, 1231, 270, 2867, 1799, 1012, 269, 1645, 1079, 2295, 2250, 619, 136, 146]. Analysiskonzeptionen [1186]. Analyst [1297]. Analytic [2453, 676, 1744]. Analytical [684]. anastatica [1188]. Anatole [2060]. anatomía [444]. anatomie [315]. anatomiques [651]. Anatomized [645]. Anatomy [1652, 2644, 1842, 1957, 1921]. Anaximander [1303]. Ancestors [1118]. Ancien [1335]. Anciens [2810]. Ancient [1905, 361, 1775, 2368, 2694, 220, 2545, 1336, 778, 572, 1516, 907, 1534, 955, 1848, 1337, 1976, 571, 635, 2732, 851, 609, 2693, 2717, 1876]. Andean [2875]. Anders [219]. Anderson [2910, 163, 1236, 2640, 382, 1044, 2721, 1151].

Andersson [343]. Andre [848, 1745, 429, 2044, 1375, 2376, 2773, 2376]. Andrea [1818]. Andreas [2817, 621, 2533, 1938, 1029]. Andrés [2911]. Andrew [1291, 1345, 2422, 894, 2747, 919, 2334, 102, 1891]. Andrews [2834, 361, 1263]. Andy [1947]. Anesthesia [1768]. Anfange [1308, 174]. Anfangsgründe [2913]. Angeles [2508]. Angelo [355]. Anglican [62]. Anglicus [375]. Anglo [2225, 127, 713, 1329]. Anglo-Norman [713]. Anglo-Saxons [1329]. Animal [2362, 2750, 2735]. animale [2219]. animalibus [2695]. Animals [2445, 654, 974, 2214, 504, 559, 2240, 1580, 775, 1763]. Animated [1269]. Anita [2555, 548, 549, 2781, 754, 1024]. Anke [897]. anlasslich [1288, 1760]. Ann [76, 1262, 2916, 1974, 78, 2200, 1691, 2181, 1993, 2265, 2583]. Anna [2368, 868, 2654]. Annals [2671, 24]. Anne [2052, 2043, 1369, 1516, 1900, 2575, 114]. années [86]. Annegrit [1379]. Anneliese [2726, 852]. Annick [1175]. Anniversary [1111, 599, 2, 1315]. Annotated [1229, 1817, 2012, 1638, 948, 2140, 723, 373, 1635, 2161, 2195, 1737, 279, 1736, 2781, 2783, 1143, 1304]. Announced [10]. Annual [2070, 64, 1133, 881, 1457, 669, 85, 2791, 332, 2387, 1801]. Anonymous [374, 1977]. Anorexia [2886]. Anpassung [2330]. ans [1589]. Ansatze [637]. Anschluss [1285, 241]. Ansprachen [2807]. Answer [261]. Antarctic [1322, 2659]. Antarctica [2184, 1322]. ante [861]. Antebellum [2916, 2465, 833, 2749, 2111]. Antébi [2928]. Antecedents [1097]. anteriores [444]. Anthology [2879]. Anthony [2527, 2445, 971, 1173, 889, 1336, 2894, 572, 989, 2202, 775, 1209, 680, 1445, 926, 2772]. Anthropological [1840, 1535, 1171]. anthropologie [2942, 1495, 223]. Anthropologist [704, 741, 459]. Anthropologists [705]. Anthropology [331, 1108, 318, 2161, 630, 119, 295, 1664, 1492, 2322, 1970]. Anti [2062, 1164]. Anti-Darwinian [1164]. Anti-Submarine [2062]. Antibiotics [587]. antico [112]. Antike [1672, 36]. antiken [1906]. Antiquarians [2744]. Antique [2358, 613, 712]. antiquité [496]. Antiquities [756]. Antiquity [1307, 425, 1636, 808, 2140, 1552, 2019, 2731, 1434, 1545]. antiquos [2911]. Antivivisectionists [1390]. Antoine [2699, 922]. Antoine-Laurent [2699]. Anton [1336, 1052, 826, 1440, 2740, 638]. Antoni [2900, 647]. Antonino [1890]. Antonio [2727, 1189, 1543, 724, 1181]. Antony [2912]. antropologia [112]. anwa [1677]. Any [190]. aphoristica [41]. Apian [2822]. APL [1747]. Apocalyptic [1444]. apogée [1343]. Apollo [797, 248, 1528]. Apollo/Saturn [797]. Apollonios [309]. Apparate [2031]. Apparatus [349]. Appearances [2644]. Appel [2404]. Appendix [374]. Apple [1556, 2647]. Application [411]. Applications [409, 2448, 2415, 474]. Applied [1025, 950, 2234, 1012, 2128, 170]. Appointments [202]. Appraisal [1359]. Appraisals [409]. Apprentice [1618, 2190]. Apprenticeship [1536, 1474, 459]. Approach [2398, 1745, 1728, 1012, 2556, 2396]. Approaches [155, 1214, 2346, 431, 1510]. après [2259]. April [1135, 768, 2474, 2793, 265, 238]. Apulei [437]. Arab [2048]. Arabic [1439, 1678, 2391, 238, 504]. Arabische [2822]. arabischen [116, 641, 1595, 2393]. arabischer [117, 1433]. arabischmittelalterliche [2393]. Arabist [2390]. Arabists [1619]. Arabs [947]. arbeitende [1017]. Arbeitsmedizin [174]. Arber [2400]. Arbib [2343]. Arbussá [2627]. Arcana [1976]. Arch [2164, 1439]. Archaeoastronomy [1039]. Archaeologist [1493]. Archaeologists [2744]. Archaeology [377, 2161, 2428]. archaique [1431]. Archeology [1494]. Archer [791]. Archetypes [1118]. Archibald [1249]. Archimedean [1590]. Archimedes [37, 880, 499, 37, 600, 1590, 2729]. Architect [1783, 2820]. Architects [563]. Architecture [2120, 1502]. Archiv [2950]. Archival [55, 428]. Archive [1556]. Archives [202, 1196, 2555, 1083, 2339, 541, 754, 319, 1399]. Areas [819, 2768]. Aréjula [442]. Argentina [2727, 1868, 860]. Argument [2564, 491]. Arguments [2244]. Ariew [2124]. Ariotti [151]. Aris [1404]. Aristarchos [1708]. Aristarchus [1757]. Aristotele [372, 117]. Aristoteles [178, 504, 2126]. Aristotelian [1563, 2445, 2369]. Aristotelianism [2697, 1683]. Aristotelischen [178]. Aristotelismus [1285]. Aristotle [2585, 1547, 2894, 778, 775, 504, 2126, 1338, 2445, 1407, 1114, 2730, 974, 2213, 2212]. Arithmetic [529, 2815]. Arithmetica [2391]. Arithmetical [2905]. Arithmetik [283]. arithmétique [111]. arithmétiques [2391]. Arithmos [271]. Arithmos-Arrythmos [271]. Arizona [2224, 2018]. Ark [2834, 1165]. Armageddon [2839]. Armand [1394, 578, 2129]. Armed [1112]. Armenian [2125]. Armin [1866, 2760]. Armogathe [2050]. Arms [2839, 2225, 2606, 2934]. Armstrong [956, 2116]. Army [2768, 2207]. Arnald [849, 1846]. Arnaldi [849, 1846]. Arney [1665]. Arno [2854]. Arnold [386, 2288, 1361, 2396, 1464, 2379, 350, 1828]. Aronowitz [2798]. Aronson [1119]. Arpad [1845, 305]. Arranged [722]. Arrangement [386]. Arrhenius [1216]. Arrighi [506]. Arrow [2256]. Arrythmos [271]. Ars [1956, 910]. Arsenic [628]. Arsenij [2826]. Art [216, 1182, 2555, 508, 1156, 2176, 678, 1855, 180, 231, 1807, 85, 2037, 1651, 1956, 2550, 1744, 1808, 1769, 390, 2941, 1931, 2471, 1957, 2356]. Artefact [2471]. Arteries [1505]. arteriis [1505]. Arthur [2249, 2579, 2097, 204, 2688, 2881, 1644, 731, 904, 2057, 1211, 2517, 1168, 721, 2378, 1727, 43, 2641, 2523, 737]. Article [2883]. Articles [2180]. Artifact [2463]. Artificial [2244, 903]. artis [2171, 236]. Artisan [1493]. Artists [514, 951]. Arts [151, 1437, 2114, 308, 2666]. Arveiller [1423]. Arzneien [1760]. Arzneimittelehre [448]. Arzt [497]. Arzte [2199, 2483]. as-saniya [1440]. as-sunniya [1440]. as-Suyuti [1440]. Asa [2835]. Ascher [622, 622]. Ascribed [1508, 2621]. Ash [1237, 1915, 55]. Asher [2218]. Ashley [2146]. Ashmole [1644]. Ashmolean [1644]. Ashworth [2171]. Asia [1662, 378]. Asian [2449]. Asmis [1588]. Asot [1983]. aspectes [2900]. Aspectos [2025]. Aspects [2628, 634, 2336, 635, 2311]. Aspetti [703]. Aspray [2249, 2672]. Assault [1422]. Assessment [367, 350, 2667]. Association [225, 1641, 2288, 749, 97]. Associations [202]. assyrische [1201]. Astrolab [2897]. Astrolabes [1978, 2348]. Astrolabliteratur [1595]. Astrologer [1178]. Astrologie [641]. Astrologie-Meteorologie [641]. astrologue [185].

Astrology [1850, 493, 2516, 1346]. Astrometeorology [877]. Astronautics [2269, 1947, 2269]. Astronome [185]. Astronomer [942]. Astronomers [1219]. Astronomi [2338]. astronomia [2900]. Astronomical [330, 977, 1536, 2347, 25, 1858, 1439, 1466, 2250]. astronomicheskaia [1753]. Astronomie [626, 116, 761, 1753]. astronomiques [780, 1471]. astronomo [2595]. Astronomy [2570, 2723, 288, 1484, 1102, 938, 1574, 948, 1851, 998, 1732, 1261, 597, 1573, 901, 41, 405, 2569, 1316, 1018, 2734, 1848, 1947, 832, 1318]. Astrophysicist [1692, 1611]. Astrophysics [288, 1575, 257, 948, 406, 1692, 1732]. Asylum [1769]. Asylums [2633]. Atack [2749]. Atanasoff [2579]. Athanasius [1626]. Atheism [2505]. ati [1202]. Atil [1931]. Atlantic [1378]. Atlantis [190, 971]. Atlasband [2502]. Atmosphere [659]. Atom [1056, 2762, 1415, 586, 423, 618, 1694, 1529, 1653, 1633, 1469]. Atomen [1906]. Atomic [1056, 132, 1415, 1147, 1216, 322, 2519, 2457, 1257, 350]. Atomica [2758, 1868]. atomico [1868]. Atomism [1447, 2321]. Atomismus [1285]. atomteorin [219]. Atop [1281]. Atran [2942]. Atropin [823]. Attack [235]. attardés [2809]. Attfield [1420]. Atti [1685]. Attilio [1926]. Attitude [1045]. Attitudes [2265]. attribuite [1675]. Attributed [2391, 2126]. aube [1431]. Audience [359]. Audio [205]. Audio-Visual [205]. Audiovisual [1303]. Audrey [980, 1007]. Audubon [1050, 128]. Aufnahme [2726]. Aufsatze [1306, 2031]. Aufstieg [2199]. Augerius [1605]. August [474, 66, 2876, 9, 2558, 275, 1879, 1861, 2702, 1001, 1783]. Augustan [47, 1480]. Auguste [867]. Augustine [2020]. Augustus [878, 2004]. Auseinandersetzung [897]. Ausgabe [117]. Ausgewahlte [1288, 1776, 36]. Aussenseiter [1104]. Austellung [2031]. Austin [548]. Australasia [387, 2191]. Australia [76, 663]. Author [2284, 2661]. Authority [1863, 1453, 2503, 97]. Authors [1637]. Autobiographical [1798]. Autobiography [704, 1530, 1395, 2764, 1121, 172, 1449, 2286]. Autograph [1783]. Automata [612]. automatas [1154]. Automation [2939]. Automobile [2508, 2580]. Autonomous [35]. Auvergne [1341]. aux [2043, 690, 1274, 114, 2259, 223]. avait [2350]. Aveni [572]. aventura [1327]. Averch [1645]. Averley [1465]. Aviation [1182, 2853, 1372]. Avicenna [566, 2818]. Avogadro [1356]. Avoid [2267]. Award [598, 599, 2544]. Axel [946, 310]. Axelrod [16]. Axiom [1160]. Axiomatics [609]. Ayed [987]. Azbel [990]. Azophi [2822].

B [1859, 129, 632, 2573, 2451, 93, 2811, 1103, 1407, 471, 1580, 2176, 1254, 1563, 478, 2026, 2253, 1100, 439, 1114, 756, 1574, 2139, 101, 122, 2832, 2930, 2741, 2545, 829, 837, 868, 364, 2343, 176, 2247, 2366, 2869, 2213, 2938, 1010, 1836, 1751, 339, 2118, 962, 1601, 1683, 2765, 2036, 2426, 1504, 2607, 1661, 2344, 570, 900, 2323, 172, 1449, 839, 1007, 1962, 92, 2126, 2250, 2111, 619, 891, 787, 963, 2918, 1594, 376]. B-1 [2607]. B.C [2694, 1087]. Baader [1203]. Baark [2422]. Baatz [1822]. Babbage [2527, 333, 926, 1747]. Babies [2198]. Babini [1627]. Babylonia [2570]. babylonisch [1201]. babylonisch-assyrische [1201]. Bachelard [77]. Bacher [2180]. Back [57, 138, 192, 249, 253, 325, 396, 453, 518, 522, 525, 589, 661, 739, 799, 803, 874, 935, 992, 1059, 1063, 1122, 1212, 1294, 1382, 1386, 1451, 1531, 1615, 1695, 1699, 1792, 1870, 1936, 1998, 2002, 2064, 2150, 2229, 2313, 2317, 2382, 2461, 2535, 2608, 2612, 2614, 2710, 2784, 2871, 2952, 1547, 1581]. Background [1761, 417, 2397]. Backscheider [124]. Bacon [1177, 1786, 2525, 971, 2051, 1910, 2524, 2823, 198, 2232, 1127, 1126, 2844]. Bacterial [1454]. Bad [965]. Badash [515, 284, 1914]. Badham [2866]. Baer [2286, 2286]. Baeriana [130]. Báguena [2751]. baike [1102]. Bailes [104]. Bailey [2140, 1901]. Bain [308]. Baird [126]. Baja [963]. Baker [614, 1570, 1555]. Balaclava [2207]. Balan [315]. Balance [2296]. Baldini [1783]. Baldwin [1839, 835]. Balfour [3]. Ball [410]. Ballester [1846, 1781]. Balloon [1373]. Ballooning [1182, 1124]. Balme [2212]. Bammé [2854]. Ban [872, 2017]. Bancroft [594]. Band [38]. Band/Erste [38]. Banesh [319]. Bank [1924]. Banks [2103, 2635]. Bannister [127, 2577]. Banton [2437]. Baptista [1043]. Baptiste [963]. Baraca [355]. Barba [2593]. Barbara [2075, 2240, 84, 1413, 1614, 1958, 2949, 510, 1267, 1520, 1342, 1028, 2707, 349, 2662]. Barber [728, 2922]. Barchas [1815]. Bark’ [2125]. Barnabas [850]. Barnaby [2528]. Barnadas [2593]. Barnes [1174, 23, 2835, 1883, 751]. Baron [744, 860]. Barona [2751]. Barovier [724]. Barrett [2716, 2240, 2748, 729, 2767, 2362]. Barry [1187, 23, 2697, 972, 1883, 751]. Barrymore [296]. Barthelemy [1049]. Barthelemy-Madaule [1049]. Bartholomaeum [1681]. Bartholomaeus [375, 375]. Barton [1359, 2759]. Bartram [788]. Bas [1719]. Based [1146, 635]. Bashe [2038]. Basic [30, 2180]. básicas [2751]. Basis [929, 1391, 883]. Basler [2696, 174, 891]. basque [1335]. Bateman [2749]. Bateson [450, 1490, 1490]. Bath [1508]. Battail [2657]. Battista [2595, 2280]. Battle [1290, 1955]. Bauer [1587, 2351]. Baumgart [1008]. Baur [2626]. Baxby [769]. Bayan [1037]. Bayet [446]. Bayo [2555]. Bayor [2558]. Bazdar [2739]. Bazerman [2883]. Be [2440]. Beagle [48, 48]. Beardmore [369]. Beardsley [2648]. Bearman [428]. Beast [654, 1665]. Beasts [2695]. Beaulieu [1394, 578, 2129]. Beaumont [458]. Beauty [2335, 2707]. Bechler [1041]. Beck [2654, 289, 1104]. Becker [2489, 796]. Becquerel [555]. Bed [2201]. Beddoes [1687]. Bede [1774]. bedeutenden [426]. bedeutender [426]. Bedeutung [862, 727]. Bedingungen [1576]. Bedini [1445]. Bedside [2885]. Beeckman [1514]. Beeger [763]. Beek [2340]. Beekman [859]. Beer [1406]. Before [2054, 95, 1494, 995, 682, 2204, 21, 710, 549, 2404, 2434, 2126, 1831, 691, 2908]. Began [1320]. Beginn [1278, 553, 82]. Beginning [2701, 20, 1617, 1481]. Beginnings [1069, 1068, 305, 300, 1087, 1738, 1573, 554, 1694, 2290]. Begins [2205]. Begriff [2813, 475]. Begriffs [1519]. Behavior [559, 1795, 224, 2683]. Behavioral [1899]. Behaviorism [1897, 2196]. Behaviorist [172]. Behaviourism [1763]. Behind [2765, 2341, 1439]. Behram [2460]. Behrens [1566]. Beier [2638]. Beighton [2341, 2341]. Being [240, 1855, 1980, 2344].

Beings [2940]. Beispiel [1055, 2199]. Beitrag [706, 310, 82]. Beitrage [1275, 1990, 785, 637, 353, 2738]. Bel [1343]. Belhoste [1788]. Belief [1985, 1544, 211, 685, 2309, 685]. Believe [1481, 2899]. Bell [2239, 2534, 1923, 1921, 2675]. Bellarmine [1783]. Bellone [336, 960]. Belloni [1554, 2760]. Below [2848]. Bembridge [2426]. Ben [2876, 1851]. Ben-David [2876]. Benassi [1668]. Bendall [1264]. Benedict [1172, 2331]. Benfey [2052]. Benguigui [1602]. Benjamin [872, 1933, 2284, 1249, 636, 1819, 2032, 186, 892, 2221, 2334, 2284, 186, 183, 1521]. Bennett [384, 1759, 2723, 1157]. Benno [2927]. Bensaude [2761]. Bensaude-Vincent [2761]. Benson [2682, 2145, 1772]. Bentley [561]. Beobachtung [2808]. Berce [1496]. Berchem [1906]. Berdoulay [1149]. Berg [1592, 430]. Bergen [1665]. Berger [1446]. Berghage [2240]. Bergius [1296]. Bergkwerck [2901]. Bergman [2054]. Bergmann [2897]. Berichte [75]. Bering [125]. Berkel [1514]. Berkeley [788, 2146, 1297, 788, 2146]. Berkhofer [96]. Berkshire [2410]. Berlin [133, 2067, 2630]. Berliner [2726, 1954]. Berman [820]. Bern [2473]. Bernadette [2761]. Bernal [449]. Bernard [1353, 468, 1665, 501, 2452, 1924, 315, 24, 582, 1851, 1957, 1544, 2494, 2475, 983, 2068, 493, 1629, 734, 1848, 1808, 850, 367, 2008, 1711, 1391]. Bernat [2900]. Bernd [2854]. Bernhard [1906, 1909, 1379, 1309]. Bernier [2259]. Berninger [271]. Bernoulli [2904]. Bernstein [2420, 2024, 655, 2290]. Berrios [1584]. Berry [166]. Bert [1778, 2106]. Berthollet [624]. Bertier [111]. Bertrand [760, 371, 2859]. Berucksichtigung [1278, 830]. Berufliche [2854]. Beryl [640]. Berzelius [831, 219]. Besieged [618]. besoin [2350]. besonder [1278]. besonderer [830]. Best [2949, 2112]. Bestehen [2807]. Bestehens [985]. Beth [2299]. Bethe [2180, 655]. Betrachtungen [1026]. Betrayers [1099]. Bettina [2745]. Between [2816, 2924, 941, 276, 886, 1354, 1049, 550, 2409, 1520, 2154, 509, 2841, 2757, 2921, 2093, 2009, 2812, 1042, 1234]. Beverly [2407]. Bevilacqua [1749, 1255]. Beweisverfahren [310]. Beyer [2817]. Beyond [466, 2238, 1553, 2061, 84, 2294, 659, 2515, 840, 2503, 1587]. Bhardwaj [635]. Bhaskar [17, 413]. Bialas [718, 1686, 1485]. Bianchi [2801]. Bibby [2497]. Bibliografia [1904]. bibliografico [1682]. Bibliographia [916]. Bibliographie [2483, 1589]. Bibliographies [2553, 412]. Bibliography [1229, 435, 952, 2835, 1277, 2869, 1635, 2161, 1544, 252, 521, 802, 1062, 1385, 1698, 2001, 2316, 2611, 2613, 341, 1736, 2781, 1555, 2783, 412, 472, 1228, 340, 2554, 1091, 534, 810, 2140, 76, 1639, 1817, 1092, 2012, 1636, 1638, 948, 723, 2195, 339, 1737, 2300, 279, 925, 1558, 1554, 1143, 1304, 1231, 2079]. bibliometria [1782]. Bibliotheca [606]. bibliothèque [1348]. bien [2050]. Bierman [1184]. Biermann [385]. Bierstedt [1107]. Big [2238, 2714, 2226, 2668, 457, 2168]. Biggs [1669]. Bildung [2830]. Bildungspolitik [1008]. Bill [2443, 1303, 1633, 2222, 350]. Billington [231]. Billion [2030, 2226]. Bilson [1575]. Bilstein [797]. Bilyeu [2304]. Bina [917]. Binary [826]. Bio [952, 1682]. bio-bibliografico [1682]. Bio-Bibliography [952]. Biobibliographic [2298]. Biobibliography [2415]. Biochemical [952]. Biochemiker [134]. Biochemistry [428, 362, 700, 905, 588, 322]. biochimiques [1488]. Bioethics [337]. Biogeography [1165]. Biographical [1166, 2399, 879, 2140, 2780, 885, 2015, 1233, 15, 951]. Biographie [2838, 452]. Biographien [426]. Biographies [716, 607]. biographische [426]. Biography [1448, 532, 98, 87, 2475, 1190, 538, 2780, 1050, 1347, 2832, 128, 885, 1356, 733, 1839, 203]. Biohistory [2679]. Biologia [112]. Biological [2679, 1837, 550, 2588, 1679, 2029, 1843, 1718, 945, 759, 344]. biologie [2259]. Biologist [466, 929, 2262]. Biologists [59]. Biology [29, 2875, 428, 2445, 1967, 429, 1730, 2588, 1269, 5, 2864, 1679, 2404, 1005, 2169, 1569, 2926, 1231, 2585, 2682, 1489, 11, 896, 467, 2394, 1222]. Biomedical [481, 1840, 2087, 954, 905]. Biomedicine [1479]. Biosynthetic [362]. Biotechnology [2929, 2724, 2928, 2261]. Bird [1811]. Birds [2401, 1268, 782]. Birth [1785, 1192, 1852, 1414, 1245, 2932]. Birthday [2212]. Biruni [2896, 1277]. bis [91, 2726, 1055, 2696, 116, 641, 1990, 1021, 1908, 2504]. bisher [1677]. Bisio [1926]. Biyosa [444]. Björn [2401]. Black [1367, 1287, 1339, 43, 1528, 2251]. Black-Body [2251]. Blacker [1170]. Blackmore [1288]. Blacks [364, 2648]. Blackwelder [866]. Blackwell [2128, 1228]. Blaise [1743]. Blake [632, 101]. Blanchard [1369]. Blas [1904]. Blaug [2015, 2578]. Bleker [2504, 706]. Blinderman [2187]. Bliss [1109]. Blocher [867]. Blocker [2014]. Blondel [1256]. Blood [965]. Bloodletting [2518]. Blossoming [1975]. Blue [2607]. Blum [1612, 1893]. Blumenbach [5]. Blumenberg [2590]. Blundell [2445]. BN [1398]. Boakes [1763]. Boardman [2107, 1486]. Bob [689]. Bochner [1001]. Bodenstein [2726]. Bodied [2646]. Bodies [2499, 1281]. Bodleian [1439]. Body [2487, 2867, 375, 2251]. Boerhaave [1443, 1443]. Boewe [2407, 925]. Bogaert [227]. Bogen [1715]. Bohdan [1565]. Bohle [1566]. Böhme [2560, 209]. Bohr [2456, 1710, 1867, 1720, 451, 1866, 955, 1120, 392, 2456]. Bois [987, 2132]. Bois-Reymond [2132, 987]. Bologna [1503, 1009]. Bolter [1360]. Boltzmann [1185, 1288]. Bomb [2017, 586, 2457, 2583, 350, 2796]. Bomber [2607]. Bonasoni [1980]. Bonelli [1300]. Bones [2438]. Bongie [854]. Bonner [2485]. Bonnet [2402, 1044, 1354]. Bonnie [2304]. Bonnot [854]. Book [2570, 2456, 735, 2767, 1201, 981, 1475, 650, 2363, 2395, 276, 2725, 953, 216, 46, 91, 1981, 558, 1663, 222, 445, 854, 1353, 2439, 49, 834, 728, 946, 1759, 2260, 2062, 2075, 2249, 344, 2166, 2890, 289, 514, 2113, 29, 387, 857, 1863, 2198, 1785, 2737, 2824, 2836, 1827, 1969, 47, 569, 1291, 2567, 152, 2914, 1252, 1481, 1493, 1196, 1286, 1809, 2239, 135, 872, 1056, 1811, 2839, 274, 888, 431, 381, 1282, 294, 1408, 1646, 2238, 1859, 1905, 2726, 1148, 2477, 2798, 15, 129, 2799, 2035, 76, 2455, 2497, 34, 79, 1371, 2649, 2771, 2848, 11, 1972]. Book [2076, 2635, 474, 1548, 1398, 1360, 629, 1559, 1327, 1987, 347, 468, 786, 1154, 1253, 1933, 2406, 2723, 81, 1756, 2920, 726, 1662, 1710, 2762, 555, 660, 2077, 2677, 632, 2284, 2350, 2880, 1208, 2447, 499, 848, 975, 1278, 1672, 1844, 2891, 271, 1186, 281, 132, 154, 886, 1160, 1289, 1583, 1055, 677, 1403, 1950, 1187, 2309, 1195, 492, 2499, 246, 2563, 1423, 2289, 2574, 1050, 776, 1428, 2244, 321, 1780, 2605, 169, 766, 1015, 1118, 1448, 1608, 2187, 2322, 2437, 2047, 2275, 1249, 185, 2398, 684, 1199, 2597, 2625, 2752, 2227, 2854, 2329, 337, 965]. Book [1111, 1427, 1665, 703, 136, 736, 1032, 2703, 22, 2630, 903, 219, 831, 2016, 74, 273, 1255, 1415, 1612, 1656, 1893, 1954, 2246, 2631, 386, 1231, 1605, 2147, 2834, 2203, 2626, 2267, 2634, 2888, 288, 336, 610, 1575, 1692, 2263, 1162, 2073, 656, 221, 481, 2223, 2937, 1260, 2862, 2929, 501, 1343, 51, 94, 611, 887, 2556, 444, 861, 48, 291, 559, 1048, 1105, 2240, 2822, 225, 819, 1273, 1460, 1897, 2915, 2194, 1652, 227, 2110, 2727, 2276, 2695, 2100, 636, 1285, 1745, 2738, 2826, 2913, 1765, 604, 2579, 603, 2676, 2021, 1657, 2564, 2326, 557]. Book [843, 929, 1362, 1928, 2660, 2942, 2810, 80, 1262, 1200, 653, 2097, 1166, 477, 1147, 2431, 1424, 1314, 697, 1429, 1746, 2685, 1039, 860, 377, 2138, 969, 2453, 1237, 511, 212, 757, 836, 2286, 2769, 2663, 2192, 779, 1508, 1673, 614, 2573, 1495, 240, 1838, 1915, 184, 408, 688, 1488, 1604, 1949, 31, 14, 719, 1631, 1329, 23, 155, 495, 2269, 2354, 730, 2041, 370, 796, 930, 1367, 2443, 980, 2638, 2532, 380, 639, 845, 2531, 366, 112, 2748, 2359, 237, 84, 896, 1944, 2042, 2121, 1243, 2667, 2650, 1229, 119, 2679, 56, 676, 1788, 2739]. Book [2910, 567, 2208, 899, 1567, 1607, 1946, 2569, 2842, 2311, 2593, 2736, 2175, 2534, 2270, 683, 1511, 272, 1185, 1722, 2013, 2312, 1177, 1307, 1847, 631, 1491, 1639, 1840, 1328, 204, 958, 1302, 2096, 2467, 2815, 583, 1864, 1959, 2248, 2530, 2566, 1182, 1023, 1487, 1658, 2661, 1025, 1425, 2452, 2688, 1837, 113, 304, 777, 1505, 2327, 627, 1484, 1947, 2335, 2916, 1345, 2766, 1102, 1441, 2245, 2451, 2700, 88, 1786, 1043, 1782, 1854, 2373, 16, 466, 911, 1916, 2706, 2399, 2555, 1194, 1290, 1203, 1779, 1853, 30, 224, 1159, 190, 395, 797, 1413, 1924, 2082, 2224, 2360, 2568]. Book [2017, 435, 470, 720, 820, 1041, 1603, 1984, 2884, 391, 1594, 1857, 2115, 2201, 2440, 93, 163, 540, 1236, 2216, 1517, 2595, 2820, 2811, 489, 1103, 1354, 2402, 2307, 1614, 44, 123, 508, 2298, 2692, 2775, 2266, 2644, 210, 586, 2304, 379, 516, 762, 864, 1104, 1149, 1321, 2027, 2186, 1402, 620, 1095, 469, 1096, 2091, 1918, 1271, 434, 643, 1275, 1407, 1551, 1817, 1958, 2696, 2801, 424, 1156, 1303, 1435, 1541, 1774, 2733, 883, 1540, 2881, 738, 952, 1115, 471, 702, 767, 1165, 1266, 1580, 1833, 2362, 378, 1205, 217, 2374, 2817, 2950, 428, 607, 950, 1092, 2131]. Book [2176, 2553, 2763, 2949, 1597, 1116, 1401, 298, 1254, 2527, 1242, 1974, 96, 704, 1121, 361, 859, 2145, 2376, 2353, 1819, 161, 315, 906, 1049, 32, 1164, 2259, 425, 1418, 2375, 550, 1333, 1556, 2112, 2904, 510, 1563, 1644, 2525, 2657, 78, 2892, 1971, 1518, 1144, 243, 915, 77, 270, 355, 409, 575, 608, 1189, 1543, 1685, 1784, 2219, 2281, 2598, 2664, 2109, 2299, 2487, 2645, 2159, 2724, 1985, 203, 1973, 2185, 2510, 2020, 2774, 2012, 2123, 1184, 2866, 2167, 322, 2180, 1943, 2637, 125, 478, 1636, 2445, 482, 483, 1094, 1775, 2782, 1331]. Book [934, 1925, 2099, 914, 1170, 2458, 168, 2773, 134, 362, 700, 2498, 2368, 2209, 500, 1906, 98, 2922, 2026, 2253, 1334, 1100, 1634, 1721, 2468, 50, 423, 732, 2050, 1613, 1155, 982, 1117, 1752, 266, 1004, 1206, 1335, 1668, 1320, 1849, 2819, 53, 768, 476, 1014, 1396, 2032, 2412, 191, 245, 2585, 1240, 1764, 2184, 302, 1355, 2257, 2699, 290, 1358, 626, 718, 1344, 1686, 1757, 1909, 2409, 658, 12, 354, 467, 1267, 1829, 2183, 166, 505, 814, 1332, 1192, 1503, 1553, 1781, 1860, 1882, 2048, 2592, 2751, 1588, 2590, 2898, 186, 731, 986, 2105, 394]. Book [1193, 2092, 2658, 2647, 2642, 116, 439, 641, 681, 708, 1432, 24, 922, 2371, 2476, 2883, 2182, 2833, 153, 1522, 1653, 1887, 2218, 793, 1648, 760, 2200, 45, 553, 2394, 189, 692, 1411, 1691, 2496, 654, 115, 780, 852, 1114, 1778, 282, 507, 623, 693, 1485, 2672, 2777, 825, 1890, 2694, 2882, 1479, 487, 988, 1660, 2028, 2148, 75, 247, 385, 1222, 1633, 1789, 1990, 2055, 2285, 2355, 2704, 1272, 1842, 389, 2349, 165, 414, 679, 2125, 1442, 2480, 2821, 427, 2561, 2768, 1146, 1713, 1960, 313, 649, 961, 1259, 1514, 2340]. Book [2030, 2135, 2680, 2691, 71, 813, 2283, 207, 278, 580, 678, 971, 1089, 1158, 1235, 1377, 1394, 1735, 1886, 2045, 2120, 2160, 2372, 52, 1022, 2222, 1108, 1172, 1970, 601, 1983, 2446, 2730, 131, 369, 1139, 67, 1152, 1667, 2009, 2901, 714, 756, 1471, 1675, 2558, 156, 1967, 2225, 213, 1019, 1638, 318, 2137, 2689, 1520, 1997, 2656, 150, 1098, 1681, 1726, 2237, 2492, 698, 546, 2606, 1550, 515, 1161, 1867, 2293, 241, 117, 309, 376, 774, 1935, 1347, 2698, 2907, 2835, 2683, 2521, 342, 1724, 832, 1316, 1574, 2709, 1469, 823, 2838, 2488, 2867, 218, 345, 2678, 2640]. Book [1766, 120, 2025, 2095, 2816, 2196, 2279, 1690, 183, 682, 1020, 578, 705, 1173, 1693, 1791, 2442, 2840, 1647, 2925, 2339, 485, 2143, 420, 957, 2857, 248, 127, 148, 2294, 2655, 582, 931, 2059, 770, 541, 948, 429, 277, 382, 1288, 1720, 2254, 966, 2632, 2849, 2776, 577, 655, 1018, 889, 970, 1112, 1566, 2063, 208, 821, 1730, 1834, 2500, 2599, 2701, 2917, 782, 585, 2902, 897, 432, 2861, 556, 1363, 1530, 1728, 1888, 1241, 2449, 1011, 2040, 2582, 149, 509, 725, 1248, 1444, 1855, 1986, 2052, 226, 2936, 220, 1659, 2051, 1500, 918, 2552, 842, 502, 108, 2546]. Book [1021, 2830, 2863, 146, 491, 1167, 1250, 1406, 1426, 1521, 1571, 1841, 2043, 2136, 2485, 2684, 2697, 2845, 1911, 602, 2846, 1684, 2084, 375, 1547, 1097, 1812, 2927, 151, 1810, 2189, 640, 1438, 233, 2274, 1277, 1851, 2896, 2928, 1965, 106, 393, 1245, 2139, 1585, 1610, 2215, 2562, 858, 2588, 223, 838, 1995, 101, 547, 1595, 1978, 2627, 2813, 1359, 1761, 2682, 122, 574, 904, 1040, 1417, 1577, 89, 324, 357, 1716, 1138, 1602, 484, 1771, 2507, 2330, 215, 1447, 1465, 2479, 2600, 1379, 178, 234, 305, 352, 2140, 2832, 2930, 2665, 817, 2087, 2181, 18, 214, 390, 1399]. Book [1562, 1865, 2741, 2545, 696, 829, 2080, 2827, 1717, 837, 924, 17, 503, 2459, 371, 968, 1369, 1671, 2753, 616, 2859, 1523, 1884, 2057, 180, 1814, 1012, 868, 990, 295, 364, 928, 1301, 1951, 2089, 2486, 2577, 2812, 2410, 1322, 2104, 1497, 2629, 2923, 2297, 2803, 1436, 27, 717, 1512, 2756, 104, 1421, 1378, 1674, 1549, 2348, 771, 1336, 1907, 100, 1106, 1047, 1169, 1395, 1750, 1370, 2860, 824, 910, 984, 1044, 1246, 1284, 2262, 2308, 2401, 2576, 1477, 2343, 167, 764, 2948, 2303, 2471, 2643, 2843, 2885, 35, 176, 231, 368, 416, 956, 1923, 624]. Book [816, 2853, 709, 1767, 2301, 2190, 544, 974, 1269, 475, 706, 862, 892, 987, 1306, 2132, 2199, 1305, 2418, 2557, 723, 1920, 2247, 2366, 1034, 70, 2869, 2841, 2903, 1820, 2022, 162, 1381, 1670, 2088, 2117, 2172, 118, 174, 587, 908, 1966, 2403, 1957, 307, 373, 978, 1552, 2242, 979, 1437, 2213, 2584, 2894, 2565, 638, 2417, 707, 2292, 2107, 778, 847, 1405, 1013, 2808, 1807, 2134, 2226, 2337, 2601, 2938, 1366, 1400, 1528, 2268, 1499, 1560, 2164, 2054, 2178, 2529, 2780, 1420, 1734, 2083, 2306, 1490, 87, 564, 2652, 1712, 1719, 750, 1741, 659, 2708, 1599, 2206]. Book [159, 269, 2681, 37, 283, 372, 1136, 1635, 1743, 1979, 2038, 2128, 2515, 346, 701, 2864, 1679, 316, 545, 1832, 2163, 95, 833, 1010, 1330, 1524, 840, 2033, 2161, 2204, 2648, 1211, 2029, 2149, 2687, 2940, 1107, 2195, 228, 565, 1207, 1975, 2365, 2441, 2728, 494, 972, 1463, 2517, 729, 763, 1026, 1268, 1836, 1210, 513, 1714, 1502, 21, 312, 615, 1042, 1150, 1632, 2282, 2429, 2484, 572, 977, 1038, 2258, 710, 2935, 384, 2346, 1365, 1934, 2749, 835, 1527, 2228, 2837, 1430, 1472, 103, 1227, 1501, 1006, 1137, 1349, 2085, 1903, 1037, 1232, 1850]. Book [1611, 452, 300, 621, 1168, 1225, 2177, 2720, 737, 358, 625, 989, 1751, 1830, 1881, 2202, 2056, 2173, 2646, 2887, 2133, 2220, 128, 1516, 1813, 687, 548, 613, 865, 1009, 2205, 605, 1016, 1409, 1473, 1593, 1642, 1731, 1988, 2019, 2086, 2405, 2528, 2744, 55, 1467, 2511, 2690, 542, 1919, 2174, 716, 69, 179, 1087, 1461, 2277, 885, 1052, 1324, 1578, 2074, 2191, 2502, 339, 617, 1544, 1628, 2094, 2494, 275, 612, 761, 985, 1280, 1380, 1600, 1908, 2210, 2516, 2856, 2886, 1557, 2755, 2393, 121, 2221, 1310, 1669, 411, 822, 2014, 2241, 2341, 2413]. Book [1093, 1926, 2490, 2705, 2118, 280, 1325, 2945, 1921, 54, 205, 319, 657, 1293, 1826, 775, 1175, 962, 479, 157, 109, 473, 549, 606, 724, 818, 884, 1088, 1815, 2101, 2370, 2719, 2058, 2745, 1513, 1601, 2347, 2251, 2380, 1058, 83, 2602, 846, 721, 1964, 785, 2392, 2427, 1725, 2878, 1645, 1773, 33, 1483, 2475, 1468, 2419, 600, 1732, 2897, 1264, 160, 365, 685, 1054, 1539, 1641, 1737, 2034, 2855, 317, 584, 690, 1003, 1397, 1476, 1900, 1912, 2352, 2551, 2761, 2411, 170, 1238, 1519, 2310, 2300, 1805, 2919, 686, 1564, 2345, 341, 827, 1017, 1119, 1650, 229]. Book [648, 1179, 417, 279, 1683, 1787, 2519, 2805, 338, 983, 1180, 1606, 1910, 1982, 2053, 2130, 360, 426, 1494, 1576, 2168, 1326, 1823, 2404, 177, 1582, 85, 947, 1510, 1637, 2214, 2899, 2272, 1682, 211, 789, 1283, 1598, 2589, 2291, 2603, 2852, 1883, 758, 2754, 2850, 912, 2505, 2575, 781, 1030, 2433, 2673, 1209, 1843, 86, 895, 1005, 1090, 1261, 1410, 1754, 2416, 2503, 1742, 2169, 2478, 2765, 25, 438, 486, 1664, 1474, 296, 2628, 1287, 1131, 1376, 2604, 1342, 2124, 2369, 175, 2514, 618, 1470, 2591, 2893, 72, 1654, 2157, 2290, 2470, 1736, 2036, 2426]. Book [2207, 1589, 973, 2423, 310, 784, 917, 1592, 2734, 137, 2559, 2103, 2560, 2804, 2325, 441, 447, 488, 579, 788, 925, 951, 1770, 1818, 1858, 1862, 1901, 1913, 1993, 2015, 2037, 2146, 2288, 2324, 2361, 2364, 2434, 2721, 2781, 2829, 2947, 815, 561, 2377, 2858, 1274, 1506, 932, 2671, 2828, 2900, 418, 913, 1113, 1361, 826, 959, 1558, 1885, 20, 133, 451, 652, 711, 867, 1609, 1868, 1917, 2333, 2422, 2548, 2654, 512, 1053, 1265, 2659, 2264, 285, 870, 2924, 105, 1439, 2408, 2501, 299, 1496, 493, 689, 1368, 1565, 1555, 1866, 1561, 1896, 2444, 2547, 2932]. Book [437, 754, 907, 1109, 1348, 1651, 1760, 1956, 2010, 2102, 2400, 2550, 2524, 1188, 1643, 1790, 1263, 2668, 543, 869, 1311, 1629, 1738, 2457, 356, 622, 694, 2078, 1223, 1374, 1586, 1929, 1233, 2378, 1718, 99, 552, 562, 727, 926, 954, 2142, 2879, 374, 1678, 2586, 1744, 2060, 2472, 1002, 2757, 566, 1029, 1584, 1902, 2179, 2495, 2921, 734, 1504, 2158, 40, 642, 1279, 287, 1153, 1356, 1689, 1755, 1861, 2702, 363, 560, 651, 945, 130, 2931, 1292, 1372, 1649, 1927, 2607, 2851, 2847, 2823, 232, 449, 68, 633, 1270, 1666, 1739, 2328, 2746, 2271, 2939, 267, 634]. Book [2332, 230, 894, 2114, 2302, 268, 765, 1191, 1661, 1835, 320, 1142, 2243, 2435, 2764, 2944, 124, 158, 1350, 1443, 1568, 2305, 2430, 2636, 2783, 1824, 2357, 2432, 2489, 2674, 1431, 1190, 2747, 2750, 413, 674, 759, 927, 2396, 964, 755, 955, 1715, 308, 752, 1318, 1776, 1848, 2391, 2587, 2814, 2895, 1422, 1763, 1554, 1323, 2448, 412, 2594, 39, 238, 504, 976, 1204, 1276, 1440, 2865, 2344, 36, 110, 303, 497, 712, 1028, 1202, 1337, 1433, 1507, 1852, 1904, 1976, 2049, 2122, 2518, 2193, 2578, 855, 902, 1806]. Book [1822, 1024, 1729, 2481, 2778, 675, 472, 1143, 440, 570, 2831, 1312, 2520, 900, 2707, 551, 890, 1046, 795, 1163, 1486, 1723, 2420, 2797, 82, 187, 539, 791, 1414, 2460, 1478, 2796, 2670, 2735, 2508, 2580, 1313, 188, 517, 1309, 2338, 164, 905, 284, 2759, 1226, 1777, 1228, 13, 430, 646, 919, 1183, 1515, 2156, 2334, 1480, 2217, 2802, 2079, 42, 242, 311, 388, 506, 573, 856, 960, 1181, 1340, 1727, 1856, 2129, 2280, 2454, 2526, 1596, 415, 480, 1151, 1464, 2424, 1992, 844, 1346, 2108, 2323, 1086, 422, 1145, 1412, 2941, 1687, 773, 644]. Book [849, 1178, 1339, 1846, 181, 571, 783, 1816, 2450, 1101, 1569, 2119, 292, 863, 2572, 1570, 340, 790, 102, 301, 314, 635, 733, 769, 1033, 1258, 1364, 1587, 1630, 1749, 1922, 1931, 2024, 2666, 1352, 2686, 172, 293, 563, 1449, 2090, 2732, 2806, 26, 1808, 2675, 1948, 792, 1051, 2653, 433, 2428, 1930, 2581, 2772, 920, 1482, 2800, 209, 673, 306, 1140, 2011, 126, 866, 923, 1579, 873, 1110, 2367, 2669, 1952, 630, 839, 1492, 1839, 1898, 2295, 2596, 2491, 2758, 2379, 2513, 2342, 446, 1007, 2081, 2825, 2407, 798, 1257, 1120, 1416, 1962, 2252, 2571, 2844]. Book [43, 92, 442, 2718, 2641, 2061, 2506, 244, 350, 419, 1057, 1317, 1772, 1319, 1466, 1733, 1573, 38, 235, 436, 850, 1977, 2126, 2171, 1341, 2044, 1968, 828, 1304, 367, 841, 2018, 2250, 359, 645, 1045, 1419, 2046, 2256, 2436, 1980, 2144, 1762, 1869, 680, 1230, 1740, 1899, 2554, 898, 1251, 1821, 1036, 2265, 2770, 1831, 851, 1035, 1375, 1590, 1769, 2162, 1581, 353, 637, 1308, 2116, 2512, 2633, 496, 1174, 1591, 2729, 967, 2809, 1758, 323, 830, 949, 1450, 1825, 1996, 568, 1373, 2381, 2106, 2504, 554, 349, 691, 749, 1008, 1446, 1546, 1880, 1894, 1994, 2093]. Book [2421, 2740, 2943, 1655, 236, 916, 1176, 1572, 1680, 1845, 2722, 2212, 2731, 715, 1091, 73, 206, 921, 1157, 1953, 1932, 1677, 443, 2908, 2111, 2127, 1694, 2197, 111, 909, 713, 421, 619, 891, 1027, 1224, 1462, 1753, 1914, 2523, 2549, 2909, 1247, 2889, 450, 1434, 107, 609, 787, 1542, 1783, 2742, 1338, 1315, 297, 348, 1197, 1445, 1768, 1889, 2098, 2356, 2583, 498, 2818, 1244, 1357, 2278, 2662, 2946, 772, 2639, 19, 871, 1234, 2023, 2493, 2926, 1525, 1509, 2483, 2651, 1239, 2170, 1545, 1640, 1351, 1711, 2906, 147, 2211, 2255, 2351, 576, 90]. Book [392, 1891, 1963, 2141, 2533, 2760, 1085, 2482, 538, 2415, 286, 173, 628, 893, 1198, 2414, 2273, 2509, 695, 1031, 1141, 1256, 1747, 1748, 1989, 2743, 2868, 114, 581, 901, 963, 1961, 1991, 2522, 647, 28, 1489, 383, 722, 1281, 1688, 991, 1404, 1828, 2039, 2358, 2918, 2934, 699, 2912, 490, 1945, 2031, 2287, 97, 1099, 2296, 351, 2779, 2425, 2397, 2933, 1529, 343, 2165, 2336, 1171, 2905, 239, 794, 1955, 588, 2261, 410, 41, 182, 751, 1526, 1892, 2469, 1676, 753, 171, 1498, 2188, 2331, 2438, 448, 2807, 1895, 933, 2693, 2717, 2911, 1241]. Book [2401]. Books [2695, 1508, 668, 1947, 2373, 471, 950, 918, 101, 1280, 2014, 109, 884, 504, 1276, 1816, 1932, 2391, 1310]. Boole [1911, 959]. Boole-De [959]. Borax [1756]. borghese [2601]. Bork [1103]. Borkin [2490]. Born [131]. Borodin [2832]. Bos [575, 827]. Boscovich [506]. Bosk [31]. Bosković [2909]. Boss [1381]. Boston [1501, 1142, 2646]. Botanic [907, 415]. Botanical [1166, 2555, 2086, 1325, 754, 907, 1651, 2550, 2081, 1166, 1760]. Botanicus [2405]. Botaniques [2555]. Botanist [2146]. Botanists [2555, 2086, 951]. Botany [2873, 702, 907, 2102, 2400, 2588, 2576]. Bottazzini [693]. Bottled [970]. Bottles [2198, 2508]. Bough [1892]. Boulle [1207]. Boulliau [1398]. Boulton [843]. Boundary [941, 2364]. Bounds [1128]. Bourgeois [2368, 1298, 2303, 70]. Boussingault [1524]. Bowditch [1142]. Bowen [158, 2603]. Bower [2491]. Bowers [2209, 707]. Bowl [261]. Bowler [1548, 1164, 2188]. Bowman [2200, 1533, 516]. Box [1408, 1367]. Boy [1743]. Boyd [1566]. Boyde [700]. Boyden [2679]. Boyer [2457]. Boyle [1493, 1986, 1982, 1257, 807, 580]. Bra [2583]. Bracegirdle [346]. Brackman [386]. Bradshaw [2656]. Bragg [341, 2141]. Brahe [2409, 1938]. Brain [2575, 2518, 318, 2163]. Brains [2858]. Bramall [2115]. Branch [658]. Brandt [2109, 1841]. Brannigan [2020]. Brasch [2719]. Brashear [2764, 2764]. Brass [261]. Braun [736, 350, 33]. Braunschweig [2830]. Bray [2449, 1679, 512]. Brazier [2573, 1601]. Brazil [2705]. Breakdown [2153]. Breaking [2922]. Breakthrough [16]. Breakthroughs [1946]. Breasts [2198]. Breckinridge [1249]. Breeder [871]. Breeding [327, 2735]. Breger [1017]. Breit [2193]. Brendan [1885]. Brent [730]. Brera [2338]. Breton [2513]. Brewer [2399]. Brewster [1992, 1657]. Brian [2203, 2525, 166, 2777, 2678, 1610, 346, 121, 789, 1654, 356, 374, 734, 107]. Briand [2170]. Bridge [233, 1671, 2753, 1671]. Bridges [2274, 231]. Bridson [2555, 2102, 680]. Brief [2258, 1357]. Briefe [1847]. Briefen [1861, 2702]. Briefwechsel [2904, 2132, 451, 1824, 1184, 385]. Brieven [647]. Brigaglia [1253]. Bright [1497, 2168, 52, 1497]. Brigitte [823, 86]. Brindle [2550]. Bringing [930]. Brings [2290]. Brinkgreve [1443]. Britain [1015, 221, 1401, 2442, 1167, 2486, 1741, 710, 2887, 160, 177, 1482, 1821, 1124, 461, 937, 144, 1219, 197, 2848, 1032, 1194, 1717, 1209, 1862, 1689, 1033, 2225]. britische [2512]. British [569, 768, 189, 1272, 756, 2088, 87, 2708, 710, 2086, 2405, 1093, 1664, 2207, 2103, 2288, 1312, 415, 680, 749, 2415, 1489, 351, 1498, 939, 471]. Britta [1495]. Brittain [2558]. Brittany [177]. Britz [754]. Broad [1099]. Broberg [719]. Broc [379]. Broca [318]. Brock [791, 1740, 2634, 1206, 1479, 1861, 1825]. Brockliss [2451]. Brockman [1787]. Brockway [415]. Broda [1185]. Brodhead [923]. Broglie [2604]. Broida [515]. Bromberg [991]. Bronisław [459]. Bronowski [151, 151]. Bronze [1087, 2012]. Brook [1556]. Brooke [1286, 1671, 1376, 773, 792]. Brooklyn [1671]. Brookman [417]. Brooks [248, 1831]. Brosse [1348]. Brothers [1182]. Brotons [1782, 1681, 1232]. Brought [1581, 2201]. Brouw [627]. Brown [656, 221, 2405, 1414, 1646, 347, 2762, 557, 2660, 2663, 2091, 183, 2308, 2130]. Browne [1088, 1165]. Brownell [1809]. Brozek [2836, 431, 1491]. Bruce [2175, 2510, 1730, 837, 180, 2348, 416, 1366, 2687, 205, 1554, 976, 1416, 949]. Bruchey [2862]. Bruck [1573]. Bruderschaft [570]. Brumberg [2886]. Brundell [2697]. Bruni [1904]. Bruno [2354, 1788, 156, 2337, 1706, 2006]. Brunot [1369]. Brunschwig [1174]. Brunsdon [782]. Brush [1554, 2831, 1412, 2676, 447, 1257, 2868]. Bruxellensis [1676]. Bruzzaniti [555]. Bryan [1194, 2765]. Bryant [2701]. Bryce [511]. Bryder [2848]. Buber [16]. Bucciarelli [692]. Buch [309, 309]. Buchborn [2808]. Buchdruck [2696]. Buchdruckerstadt [2696]. Bücher [2901]. Buches [117]. Buck [2420, 440]. Buck-Morss [2420]. Buckland [1419]. Buckle [1081]. Bucknell [1101]. Bud [269, 1689, 1828]. Budd [1971]. Buddensieg [1566]. Buffetaut [2501]. Buffon [857]. Buge [2080]. Bugliarello [176]. Buican [1323]. Building [81, 1182, 1562, 2559, 2510, 2274, 2204]. Builds [2796]. Built [569]. Bujosa [1782]. Bullet [2109, 1841]. Bulletin [540]. Bullock [656]. Bullough [2780]. Bulmer [2486]. Bun [944]. Bun-Ichi [944]. Bunge [90]. Burbank [1995]. Burch [2091]. Burdet [2555]. Bureaucratic [2175]. Bureaucrats [971]. Burghard [1640]. Buridanus [2813]. Buried [1581]. Burke [2159, 1515, 2800, 2255]. Burks [2579, 2249, 2579]. Burndy [1073]. Burnett [1774]. Burnham [1666, 2803, 2547]. Burning [808]. Burning-Mirrors [808]. Burnyeat [1174]. Bursey [828]. Burt [293]. Burton [2484]. Busard [1508]. Busch [1661]. Business [2709, 149, 1923, 2366, 1921, 2668]. Butterfield [264]. Butterflies [1268]. Buttner [762, 1198]. Butts [2913, 1716]. Butzer [900]. Buxton [2527, 2329]. Byatt [710]. Bylebyl [173]. Bynum [2439, 1842, 1088]. Byrne [2133]. Byzantine [2047, 1852]. byzantinischen [1593].

C [49, 2113, 569, 2567, 1809, 294, 1559, 1933, 2762, 766, 386, 1231, 2267, 511, 1177, 1864, 1959, 257, 627, 1290, 93, 2644, 1435, 767, 1205, 425, 1971, 2185, 914, 1206, 1014, 1829, 2092, 2182, 2349, 2125, 1713, 971, 2558, 2698, 1724, 183, 485, 127, 2632, 2917, 1571, 2845, 838, 18, 837, 2459, 2577, 2803, 368, 2213, 2337, 2938, 1719, 316, 95, 710, 1751, 2221, 109, 1054, 686, 1119, 1494, 912, 2765, 25, 1287, 2591, 959, 1558, 133, 2654, 2547, 1929, 552, 1755, 1666, 765, 2122, 2778, 2831, 2759, 2802, 1412, 635, 2581, 2011, 446, 2407]. C [244, 680, 2729, 2093, 715, 1244, 1525, 1509, 2273, 647, 1529, 2453, 106, 2037, 1113, 374]. C. [1972]. ca [1779, 116, 641, 1087, 1544, 1980]. Cabanis [650]. Cabinet [2467]. Cabinets [2465]. Cable [2753, 367]. cachées [2880]. Caddes [2134]. cadre [906]. Caine [700]. Cajal [444]. Calaby [1839]. Calcination [1941]. Calcium [2101]. calcolo [693]. Calcul [1592, 1743]. Calculating [1746, 1254, 1020, 1746]. Calculations [2790]. Calculators [1254]. Calculus [623, 693, 484, 1768]. calcun [848]. Caldani [1179]. Caldwell [331, 2649]. Calendar [387, 1628, 1315, 1047]. cAli [376]. California [2080, 2079, 963]. Calinger [1411]. Calley [2176]. camal [308]. Cambridge [702, 2843, 2894, 13, 1989, 1068, 2787, 240, 2147, 181, 691, 1526]. Came [1401, 1038, 1494]. Cameron [350, 697]. Camille [29]. Campaign [2771, 2267]. Campbell [1571, 919, 2943, 2298, 2281, 2191, 2919]. Campo [2592]. Campus [2465]. Can [809]. Canada [344, 210, 658, 1500, 1609, 2757, 20, 1446, 628]. Canadian [2113, 2869, 2118, 1464, 2116, 967, 410, 2713]. Cancer [481, 229, 1534]. Candid [1418]. Candolle [2804]. Canguilhem [228]. Cannon [812]. Canon [2818]. canons [1591]. Cantelon [1290, 1529]. Cantor [2566, 162, 2334, 695, 1482]. Canziani [2160]. Cape [1166, 1610]. Capitalism [2815]. Capitalist [2329]. capitulo [1781]. Caplan [204, 2881]. Capp [493]. Cappellari [1525]. Captain [2559, 1808]. Captured [1974]. Car [34]. Card [1746]. Cardano [1178]. Cardenas [963]. Cardinal [2281]. Care [2634, 1974, 1556, 2172, 2648, 1737, 2036]. Careers [1940, 1217]. Cargill [2269]. Caribbean [1843]. caricature [1332]. Carl [969, 1838, 864, 984, 987, 300, 1309, 1446, 289, 2605, 762, 864, 1558, 1046]. Carla [2131, 2149]. Carleton [704, 771]. Carlisle [1293]. Carlo [2567, 2673]. Carlos [2696]. Carlson [946, 137, 1498]. Carmelia [1676]. Carmen [1037]. Carnot [218, 585, 2220]. Caroe [2167]. Carol [2299, 1320, 2092, 2052, 2080, 1820]. Carolina [2375, 828]. Carolinas [2146]. Caroline [2140]. Carolyn [1050, 583, 2530, 608, 2105, 2134, 228, 338, 2344, 1346, 2686]. Carozzi [1559, 1045]. Carpenter [684, 1972, 1310]. Carraud [2050]. carriére [625]. Carrillo [442]. Carroll [915, 1828]. Carsten [2946]. Carteggio [44, 1179, 1856]. Carter [1809, 1208, 290, 2837]. Cartesian [1042]. cartésien [311]. cartésienne [1117]. cartesische [2245]. Cartography [2545, 486, 2356, 904, 2545, 2257]. Cartwright [1144]. Carver [1191]. cas [2829]. Case [386, 611, 887, 1362, 796, 2881, 32, 270, 2319, 1235, 1269, 269, 665, 1921, 1683, 72, 2472, 350, 871]. Cash [909]. Caspard [2170]. Caspard-Karydis [2170]. Caspian [1378]. Cassebaum [1046]. Cassedy [1690, 2197]. Cassidy [2654, 1555]. Castillejo [720]. Castle [1562]. Catalan [1192]. Catalana [1192]. Catalog [1559, 2835, 918, 2258, 2469, 818, 349]. Catalogo [2338]. Catalogs [2475]. Catalogue [1196, 2373, 2399, 1556, 1644, 756, 101, 1465, 2348, 1207, 605, 109, 2102, 1816, 2555, 133, 2555, 2413, 2103]. Catalogues [2284, 541]. Catalogus [373]. Catalunya [2900]. Catedra [1613]. Catherine [2487, 213, 1490, 1207, 2520, 573]. Cathode [736]. Cathode-Ray [736]. Catholic [317]. Catlett [2414]. Catt [2917]. Cattell [490]. Cattermole [2843]. Cauchy [693, 1788]. Causa [2053]. Causal [1724]. Causality [1547]. Causation [1948]. Cause [2324, 1281]. Causes [1971]. causis [1505, 1778]. Caveing [1274]. Cavendish [2320]. Cavini [2122]. Cayleff [2440]. CBI [333]. Cecelia [2307]. Cecil [360]. Cecilia [1395]. Celebrate [768, 1193]. Celebration [1111, 2459, 1671, 2428, 2141]. Celesti [1000, 1685]. Celestial [2066, 486, 783]. celestis [1774]. Cell [2441]. celle [1162]. Celli [1904]. Celluloid [1670]. Censer [2289]. Censure [1855]. Census [2551, 2721, 1466, 642]. Centenari [1192]. Centenary [768, 1192, 1867, 2459, 513, 133, 632]. Centennial [2138, 1193, 1671, 416, 1578]. Center [1804, 1556, 2368]. Central [969, 1613, 1751]. centrale [2350]. Centrality [1541]. Centre [46, 184, 906, 541]. cents [1589, 2900]. Centuries [2451, 1601, 1466, 1373, 2549, 15, 510]. Century [2439, 1481, 1282, 2848, 281, 2275, 2267, 2100, 80, 2638, 1639, 2467, 1023, 2733, 1115, 378, 1401, 1985, 1355, 439, 2833, 1222, 2120, 601, 2009, 1520, 2816, 725, 1444, 2052, 1406, 2846, 106, 101, 1359, 1465, 2140, 1717, 2753, 368, 956, 1269, 987, 1920, 2584, 1832, 2648, 2441, 2517, 312, 2646, 2205, 1473, 2511, 1732, 1003, 2919, 827, 648, 758, 2575, 2157, 1862, 2324, 1113, 285, 560, 232, 1739, 124, 1727, 1145, 102, 1898, 92, 2718, 1977, 645, 2046, 2633, 1768, 576, 1082, 1961, 1489, 97, 2188, 1676, 529, 2539, 1704, 2321, 1921, 1455, 1130, 942, 2464, 1217]. Century [1535, 2634, 2573, 2815, 1854, 2644, 1156, 1541, 702, 2131, 425, 2487, 168, 2922, 623, 2222, 1447, 968, 180, 565, 2365, 2019, 2310, 1823, 2754, 447, 1858, 652, 711, 299, 2495, 1350, 2430, 1486, 1816, 2686, 1317, 1341, 1590, 949, 1546, 90, 1989, 2743, 2574, 1472, 1731, 1952]. Ceramics [2358]. CERN [2760]. Certain [1349]. Certainty [1520]. Cerullo [1033]. Cervellera [2751]. cervello [112]. cete [2050]. Cetina [1409]. Chabàs [2900]. Chain [932, 2061, 1419]. Chair [2544]. Challenge [1973]. Challenger [2766]. Chamberlain [1038]. Chambers [269, 616, 269]. Champaign [1559]. Champier [107]. Chan [2453, 571]. Chance [1814]. Chancellor [128]. Chandler [2348, 2668, 1025, 180]. Chandrasekhar [2335, 1611]. Change [1362, 80, 2312, 166, 277, 1363, 2410, 1934, 2646, 1919, 1280, 317, 2158, 882, 1930, 1317, 1821, 609, 1338, 147, 2203, 2688, 2020, 433, 2878]. Changed [1405, 2159]. Changing [1023, 425, 1248, 2797, 2265]. Chant [685]. Chantilly [46, 906]. Chapel [828]. Chapman [1578, 2775]. Chapoulie [2170]. Chappe [963]. Chapple [2754]. Chapter [2109, 2485, 2400, 1977]. Chapters [1851]. Character [2367]. Characteristics [80]. Charakter [1340]. Charge [2427]. Chariots [248]. Charlene [2832]. Charles [1809, 1187, 2187, 2752, 1427, 736, 386, 1273, 477, 31, 2748, 896, 2312, 583, 1864, 2530, 1182, 2661, 16, 1354, 1407, 1774, 2527, 1563, 478, 2883, 1114, 333, 52, 714, 393, 390, 2741, 928, 1951, 2038, 513, 1714, 2191, 1628, 2475, 1683, 1598, 925, 1913, 1885, 299, 1651, 230, 2114, 327, 863, 2666, 2407, 1772, 2126, 787, 297, 1351, 196, 331, 260, 1175, 1218, 926, 593, 2008, 49, 1286, 129, 730, 2748, 1170, 1044, 316, 1052, 1578, 2364]. Charles-Saget [1175]. Charleston [349]. Charleton [1855]. Charlotte [2406]. Charlton [968]. Charmasson [2170, 639]. Chartrand [2725]. Charts [1577]. Chase [1931]. Chassagne [2170]. Chatelet [2282]. Chatterjee [917]. chaudron [1496]. Chaussées [1369]. Chautemps [2555]. Checklist [682]. Checklists [1263]. Chemical [2166, 831, 403, 302, 2339, 970, 2136, 1013, 1456, 665, 2370, 2589, 194, 1922, 1304, 1447, 2572]. Chemici [649]. Chemie [82, 358, 1830]. Chemielehrbücher [2745]. Chemiker [2726, 2726]. Chemikern [2726]. chemische [287]. chemischen [823, 174]. Chemist [2223, 2832, 1524, 1687, 1530]. Chemistry [81, 1216, 163, 322, 2009, 2339, 1620, 2181, 1456, 1881, 818, 2370, 2589, 404, 285, 1689, 905, 43, 92, 2926, 805, 1751, 2677, 2359, 1751, 2551, 1755, 164, 1758, 1198, 2455, 2371, 165, 207, 594, 1828]. Chemists [143, 2005, 828].

Chemotherapy [1534]. Chen [2450]. Cheng [2026]. chercher [2050]. chercheur [2170]. Cherniack [2302]. Chernow [2949]. Chesterton [2306]. Chevalley [573]. chez [1343, 414, 1175]. Chi [378, 571]. Chiara [608]. Chiarugi [2639]. Chicago [2195]. Chief [1334]. Child [777, 2107]. Childbearing [2201]. Childbed [1208]. chimique [624, 624]. chimiste [625]. China [2363, 377, 119, 1102, 2588, 1679, 2370, 2589, 441, 2422, 440, 1537, 2386, 743, 2622, 118, 1598, 2796, 1085]. Chinese [1205, 1597, 269, 181, 571, 2450, 1218, 261, 1876, 405]. chinoises [1441]. Chirurgica [910]. chirurgie [1343]. chirurgien [1343]. Chirurgische [1475]. chisel [281]. Chodorow [953]. Choice [1160]. Choices [2419]. Cholera [2752, 2923, 2641, 1210, 175]. Chomsky [2420]. Choppin [2170]. Chouet [1042]. chrétien [237]. Chris [1642]. Chrisman [1280]. Christa [1523, 2132, 133]. Christen [617]. Christentum [2355]. Christiaan [575, 2128]. Christian [1864, 2374, 385, 1951, 27, 356, 2807, 997, 1220, 2542]. Christianity [2093]. Christianson [909, 1603]. Christie [2334, 1992]. Christine [1343, 542, 1052, 1256]. Christoffel [900]. Christoph [2210, 82, 1340]. Christophe [717]. Christopher [2910, 1786, 2399, 1152, 2887, 2628, 2116, 1157]. Chronological [1864, 928, 1951]. Chronology [1946, 2552, 1827]. chuang [378]. Chubin [481, 1229]. Chung [571]. Chung-kuo [571]. Chuquet [1958, 280]. Church [2780, 997, 1221, 1859]. Churchland [1719, 2163]. Chymical [2373]. Chyrurgery [1585]. cielos [2025]. ciencia [861, 860, 1781, 1860, 1232, 1917, 1917, 443]. ciencias [2751]. Ciències [1192]. Científica [2914]. cientificas [1917]. cientifico [861]. cientificos [1782]. cilm [308]. Cimino [703, 544]. Cinema [2637]. Cinq [1589]. Cioc [2758]. Circle [2723, 186, 2670]. circolo [1253]. Circulation [173]. Circumnavigation [1512]. Circumstances [2538]. Citation [2071, 2388, 753, 411]. cite [1375]. Citizen [2380]. Citizens [687]. City [670, 2642, 1110, 841, 4, 2508]. Civil [2289, 2842, 2375]. Civilisation [2588, 1679, 2370, 2589, 1085, 1431, 301]. Civilization [2679, 2445, 118, 542, 2641, 2693]. Civilizations [892]. Civilized [2581]. Clagett [600, 37, 436, 1590, 2448]. Claims [2717]. Claire [1946]. Clanchy [179]. Clare [1346, 2147]. Clark [15, 342, 1322, 606, 1718, 2396, 2469, 2484, 1564, 2377, 2061, 2755]. Clarke [2574, 1370, 982, 2183, 328]. Clash [2916]. Clásicos [444]. Classical [2499, 700, 1235, 1638, 285, 1145, 673, 2652]. Classics [2767, 2362, 1632, 2550, 1411, 92]. Classification [222, 257, 2445, 2639]. classique [496, 1431]. Claude [1488, 425, 2699, 582, 624, 979, 2068, 551, 2809, 1391]. Claude-Louis [624]. Claudia [1817, 2199]. Claudine [2629]. Claudio [1919]. Claudius [2393]. Claus [556, 1153]. Clavis [2235]. Clay [1430]. Clayton [1057]. Clelia [1609]. Clemencon [1819]. Clercq [1952]. Clergyman [1473]. Clerk [1522, 2599, 733]. Cleveland [1578]. Clifford [336, 1719]. Climate [312]. Cline [1020]. Clinical [1668, 1198]. Clinton [1693]. Clipped [2291]. Clock [2128]. Clocks [1377, 2348, 612, 2347, 2348, 2128, 25, 278]. Clockwork [612]. Close [1816]. Closely [2081]. Closs [2215]. Closure [2881]. Cloud [2136]. Clouds [2768]. Clubs [2562]. Clutius [1605]. Co. [895]. Coan [986]. Coast [2833, 1378]. Coastal [1577]. Coates [367]. Cocain [823]. Cochrane [18]. Cockroft [1653]. Cocks [1756, 1916]. Codd [1166]. Code [622]. Codell [1208, 2837]. Codex [977, 1676]. Coe [1227]. coelestium [1686]. Cognition [2664, 1012, 2865]. Cognitive [2556, 1391, 1012, 1714, 1896]. Cohen [474, 544, 1712, 1467, 468, 1604, 2452, 1020, 364, 228, 2221, 983, 1629, 2344, 1711]. Cohn [2176]. Cole [246, 246, 2768]. Coleman [29, 2793, 390, 512]. Coleridge [758]. Colette [657]. Coley [685]. Colin [2223, 118, 685, 1190, 1226, 1821]. Collaboration [423]. Colleagues [2173, 2531]. Collected [1612, 1893, 2654, 647, 1120, 1959, 1159, 2293]. Collecting [2176, 2260]. Collection [1398, 1559, 2555, 163, 681, 2080, 2348, 771, 2401, 2258, 2241, 109, 606, 818, 1815, 374, 1976, 2469, 2348]. Collections [2870, 2951, 819, 2013, 2399, 2555, 2339, 2337, 2014, 951, 1818, 2414, 1644, 2550, 374]. Collective [607, 532, 87]. Collectors [1166, 2399]. Colleen [2339]. College [1196, 474, 240, 1097, 1812, 87, 349, 656]. Colleges [1229, 1835]. Collegio [1688]. Collingridge [350]. Collini [491]. Collins [1241]. Colloque [304, 906, 690, 496, 507]. Colloques [551]. Colloquium [2480]. Colomb [717]. Colonial [2576, 415, 909]. Colonialism [2425]. colonialismo [1609]. Colonies [2266]. Colonization [2666]. Color [60, 2915]. Colours [1321]. Colton [2862]. Columbus [96]. Combridge [2347]. comburentibus [1177]. Combustion [1941]. Come [160, 80]. Comes [1428, 1641]. Comet [2714, 1752]. cometa [1681, 1681]. Cometary [1961]. comete [1705]. Comets [1827]. Coming [752]. Commemorate [1641, 1315]. Commemorating [1512]. Commemoration [657]. Commemorative [307, 1287, 1578]. Commentaar [1780]. Commentaire [979, 2043, 114]. Commentaries [1779, 2293, 373]. commentariorum [373]. Commentary [2531, 1505, 1910, 917, 1202, 1507, 777, 1786, 1778, 1851, 1440, 242]. Commentum [1846]. Commercial [2511]. Commercium [1983]. Commission [2289, 1093, 1056, 126]. Commissioned [2103]. Committee [2017, 1083, 419, 2381]. Common [2185, 67, 95]. Commonly [1508]. Commonwealth [2631]. Communauté [86]. Communication [243]. Communications [2686, 147]. Communities [2424]. Community [257, 1266, 1013, 743, 2889, 1298, 268]. Companion [129]. Company [347, 1166, 2843, 1818, 1922]. Comparative [1243, 1173, 2486, 839, 2918]. comparée [315]. Comparison [1158, 2425]. Comparisons [1129, 1919]. Compass [1445]. Compendium [308, 2171]. Compilation [85]. Compiled [541]. complement [2594]. Complementarity [1720]. Complete [2857, 1981]. Completed [1958]. Completion [1844]. Complex [2261]. complexionis [1846]. Composition [722]. Compounds [829]. Comprised [1544]. Compton [1772, 1533]. Compulsory [1199]. Computation [1746]. Computational [2656]. Computer [2249, 1360, 1428, 2579, 2685, 1935, 1739, 1748, 569, 1428, 926]. Computers [569, 2567, 2358, 2038]. Computing [2249, 1084, 2272, 24]. Computistical [2621]. Comune [1009]. Con [2911, 1188]. Concept [1208, 51, 1631, 153, 1761, 528, 2033, 1601, 566, 1584, 350, 2397, 2165]. Conception [1718, 1570]. Conceptions [425, 933, 695]. Concepts [2635, 380, 487, 220, 2845, 1919, 404, 2574]. Conceptual [2445, 417, 1145, 1317]. Concern [1420]. concerning [886, 2128]. Concise [2674, 885]. conclusione [1188]. Concordance [2240, 729]. Concorde [2442]. Condillac [445, 854, 381]. Condit [969, 300]. Condition [1956]. conduire [2050]. Conference [474, 1415, 610, 1000, 768, 200, 65, 2622, 909, 609, 1315, 410]. Confessing [1701]. Confidence [1769]. Confirmation [2481]. Conflict [2916, 61, 2419, 2291]. congetture [1845]. Congress [1415, 2756, 1760, 2100, 2337]. Congressi [1643]. Conic [600]. Conics [1844]. conjurateurs [1673]. Conkin [543]. connaissance [2127]. Connelly [754]. conoscenza [2801]. Conrad [126, 891]. Conry [1966]. Conscience [892]. Consciousness [2821, 667]. Consent [1862]. Consequences [995, 2538, 1449]. Conservatoire [2555]. Conspiracy [2491]. Constance [1161, 2034, 2772]. Constant [567]. Constantin [227]. Constants [2024]. Constituting [664]. Constitution [62]. constitutione [1774]. Constraint [2557]. Constructing [1887, 1891]. Construction [1015, 212, 2354, 1243, 156, 2343, 2204, 548, 549, 2274, 2557]. Constructive [2775, 2776]. Consultations [1668]. consummatio [236]. Consumption [2649]. Contained [1290, 1439]. Contemporaries [592]. Contemporary [224, 541, 2346, 413, 2767, 1041, 2362]. Contention [2438]. contesti [355]. Contet [1207]. Context [1840, 971, 2225, 791, 173, 1424, 394]. Contexts [481]. Conti [506]. contineatur [1505]. Continent [1322, 790]. Continental [2149]. Contingency [317]. continu [89]. Continuation [541]. continue [1274]. Continuing [1199]. Continuity [198, 1919, 851]. Continuous [1771]. Continuum [1307]. contra [2915]. Contribution [2576, 2392, 1762, 2127]. Contributions [1565, 1658, 1582]. contributo [1255]. Contributors [2792]. Control [1195, 2222, 2225, 2546, 1245, 616, 35, 2932]. Controlling [1694]. controverse [1819]. controversia [2219]. Controversial [929]. Controversies [2360, 2881, 160, 2438, 821]. Controversy [256, 1898, 1880, 28, 196, 269, 1244, 1058, 1587, 1895, 1539]. Convegno [1685]. Convergence [1691]. Converging [1214]. Conway [1516, 2052]. Cook [1808, 2110]. Coon [704, 704, 2360]. Cooper [2373, 986, 395, 2863, 1030]. Coordination [829, 92]. Cooter [1862]. Copans [223]. Copenhaver [107]. Copernic [2824]. Copernican [270, 2590]. copernicana [2025]. Copernicanism [1541]. Copernicus [1981, 1909, 2006, 462, 259, 2570, 716, 1600]. Copper [2874, 1599]. Copy [668, 1783]. Coquand [1369]. Corbin [2352]. Corgan [833]. Corn [1372, 2039]. Cornelia [2489]. Cornelis [313, 1394, 578, 2129]. Cornelius [830]. Corner [1121]. Corporate [2937]. Corps [1369, 1343, 712, 1343].

Corpus [2445, 1379]. Corpuscles [1982]. Correlation [2789]. Correspondance [1983, 1602, 1743, 1394, 578, 2129]. Correspondence [1759, 1933, 511, 1354, 2402, 1047, 987, 1628, 2221, 2288, 959, 1866, 1666, 2156, 511, 2650, 1052, 356]. Corridor [1371]. Corrigendum [853]. corrispondenti [1181]. corrispondenza [2338]. Corsi [1230]. Cortada [1635]. Cosa [2368]. Coser [1647]. Cosman [180]. Cosmic [1757, 1997, 2430, 1953, 1018, 2255]. Cosmogonical [2790]. cosmographicum [1518]. Cosmological [2024]. Cosmologists [2519]. Cosmology [806, 1038, 2124, 976, 1440, 1597, 1258, 2024, 2522, 1575]. Cosmos [720, 278, 306, 880, 1567, 1814]. Cosslett [1546]. Costabel [2904, 1983, 2594]. Costs [2172]. Coudert [434]. Coues [923]. Couloubaritsis [776]. Coulston [1182]. Council [344, 80, 210, 658, 1500]. Count [183]. Counted [529]. Counterthemes [1832]. Country [2284, 2890]. cours [2350, 86]. Course [2242]. Courtenay [2584]. Courtney [248]. Cousin [1857]. Covers [2297]. Coveted [738]. Cowan [1365]. Cowdrey [1265]. Coyne [2884, 1784, 1783, 1315]. Cracow [1466]. Cradle [1415]. Craft [298]. Craig [886, 2842, 93]. Cranefield [987]. Cranz [373]. Crater [2360]. Craters [196]. Crawford [2551, 1738, 1309]. Creating [1727, 2685]. Creation [11, 1710, 1307, 1540, 1985, 2030, 2118, 1539, 2157, 907, 1244, 1125, 1875, 1256, 316, 2094]. Creation-Evolution [1539]. Creationism [887, 1244]. Creative [216, 929, 380, 2656]. Creativity [1634, 1217, 1404, 1881]. Creator [1603]. Creatures [2750]. Creed [2179]. Crime [1631]. crise [2773]. Crisi [1000, 1685, 1790]. Crisis [2175, 1502, 97, 1290]. Crispin [324, 1570]. Criteria [352]. Critic [2295]. Critical [445, 2453, 1177, 217, 1092, 1778, 1022, 2655, 1851, 2948, 1107, 2220, 252, 521, 802, 1062, 1385, 1698, 2001, 2316, 2611, 1637, 913, 335, 412, 1440, 2295, 2023, 1464]. Critically [2678]. Criticism [2293]. Critics [857, 79, 2519]. Critique [726, 2244, 500, 1970, 145, 413, 2043, 624, 1274, 1943]. Croatia [1462]. Crombie [2011]. Cronica [1868]. Cronin [1567]. Crook [2636, 1521]. Crops [1345]. Crosby [2029]. Cross [2422, 2425]. Cross-Cultural [2422, 2425]. Crossroads [1730]. Crouch [568, 1373]. Crow [1273]. Crowd [1003]. Crowe [2082]. Croyances [1496]. Crusade [2649]. Crusaders [1054]. Crystallization [1874]. Crystallography [1260]. cuiuslibet [236]. Cullen [43]. Cult [1056]. cultura [604, 1517, 1189, 1790]. Cultural [1403, 2451, 1193, 1970, 2698, 2867, 269, 252, 521, 802, 1062, 1385, 1698, 2001, 2316, 2611, 2613, 1862, 2422, 694, 147, 893, 2425]. culturali [1326]. Culture [1360, 1231, 2307, 1479, 213, 2492, 2308, 2529, 750, 2033, 1087, 1280, 1376, 447, 1361, 1392, 2547, 2457, 1279, 1312, 2800, 839, 1247, 2493, 239, 2331, 2521, 2178, 2515, 2687, 621, 622, 1649, 1994, 2331, 1403, 68, 839]. Culturele [2825]. culturelles [2880]. Cultures [2618, 2482, 1226, 2433]. Cumulative [524, 2541, 412, 534]. Cunningham [1662]. Cuore [112]. Cup [2469]. Cure [2440, 2111]. Cures [2931]. Curiosities [2467, 2465]. Curiosity [2408, 730]. Current [614, 2613, 2858, 2390]. Currents [2326, 562]. Currie [152]. Curry [2516]. Cursory [2453]. Curti [816]. Curtis [2146, 630]. Cusanus [1777]. Cushing [1713]. Cutcliffe [542]. Cutright [923]. Cuvier [387, 1863, 2404]. Cuvier-Geoffroy [2404]. Cybernetics [2293, 2940]. Cycle [767, 2256, 2276]. Cyclones [168]. Cynthia [1958]. Cypher [1747]. Cyril [678, 1599, 293]. cytologie [2259]. Cytoplasmic [2503]. Czartoryski [1981, 73].

D [2737, 1827, 2937, 1262, 2453, 511, 1864, 113, 304, 2555, 93, 1236, 123, 2692, 298, 1556, 2647, 852, 1479, 52, 2196, 2840, 2339, 582, 2063, 1521, 1965, 357, 771, 2885, 778, 2938, 1136, 1207, 763, 1714, 2405, 1264, 685, 1510, 1843, 2765, 1287, 2659, 2102, 2378, 363, 449, 1431, 752, 2122, 2707, 594, 1992, 1258, 2428, 2011, 1839, 1772, 680, 967, 568, 1373, 1680, 73, 891, 2763]. D.D [2146]. D.O.’s [1270]. d’acque [1503]. Dadic [1462, 2909]. Dagut [2731]. Dahms [2489]. Daiches [2216]. Daintith [885]. Dal [1009, 604, 1682, 488, 1979]. d’al-Khayyam [1979]. d’Albert [1192]. Dale [1971]. Dales [1435]. Dalesman [1190]. d’Alexandre [677]. d’Alexandrie [2043, 114]. Dalla [608]. dalle [1749]. Dalrymple [2941]. Dam [2175, 2774]. Dame [672]. Damerow [1572]. Dan [1260]. Danaher [2677]. Dance [1621, 2260]. Danforth [2469]. Daniel [1056, 71, 2088, 2887, 2133, 711, 99, 699]. Danielle [1506]. Daniels [1031]. dans [975, 677, 639, 237, 906, 500, 2050, 2392, 1506, 1431, 712, 496]. d’après [2804]. d’aquisition [505]. D’Archimède [2880]. d’Aristote [776, 2691]. Dark [2028, 2063, 1806]. Darkness [2568]. Darnton [149]. Darrel [107]. Darstellung [309, 764]. Darstellungen [426]. d’Arts [2349]. Darwin [1759, 129, 1187, 2398, 386, 48, 2240, 636, 757, 730, 2748, 2013, 2091, 2689, 1406, 1547, 2843, 1966, 316, 729, 1052, 1578, 2191, 617, 1628, 2221, 160, 1539, 317, 1005, 2305, 927, 691, 2873, 1701, 1872, 1218, 593, 327, 2233, 1048, 2565, 685, 2404, 1598, 863, 1327, 2920, 2683, 1406, 685, 1326, 895, 2921, 187, 1994]. Darwinian [1164, 389, 2680, 2074, 160, 794, 794]. Darwinism [757, 1164, 550, 2048, 1967, 127, 1832, 1564, 1763, 2309, 821]. darwinisme [1966]. Darwinist [2542, 1521]. Daryl [481, 1229]. Dash [20]. D’Assigny [1855]. Daston [2652]. d’astronomie [2895]. Data [80, 1635, 531, 1176, 1728]. Date [809]. datis [850]. Dauben [1884, 1636, 825, 162]. Daughter [1490]. Daumas [1393, 301]. d’Auteroche [963]. Davenport [2365]. Davey [76]. David [2890, 2113, 1360, 610, 2937, 1949, 2748, 1177, 1505, 16, 720, 2216, 428, 32, 322, 2876, 1240, 1344, 1377, 948, 2863, 1250, 2928, 838, 122, 2545, 616, 231, 2853, 2022, 1437, 2164, 2529, 564, 269, 2161, 865, 1293, 606, 2519, 2168, 2591, 2654, 1555, 2921, 40, 642, 2939, 2396, 164, 1992, 2669, 350, 1772, 1869, 949, 1450, 2093, 2212, 450, 609, 2356, 1509, 1748, 2743, 2918, 239, 794, 350, 1493, 2912]. Davies [2762, 513, 1321]. Davis [2399, 1885, 1404, 1007]. Davy [359, 793]. Davydd [1764]. Dawn [2457, 947]. Dawson [1959]. Day [11, 2299, 2159, 1529, 2349, 907]. Days [1166, 2247, 2603]. Dazzi [703]. DDT [628, 687]. Dea [1368]. Dealers [2077]. Dealing [374]. Dean [1634, 2181, 176, 126, 1899, 2705]. Dear [1268]. Death [2022, 338, 1339, 420, 2645, 1004, 2923]. debat [651]. Debate [899, 204, 2082, 2713, 2281, 939, 398, 2788, 2404, 1261, 940, 2758, 2188, 350, 1199, 1560, 648, 1586]. Debates [789]. Debility [365]. Deborah [214, 486]. Debra [2372]. Debris [2255]. Debru [1488]. Debus [2371, 2902]. debut [1375, 1743]. debuts [708, 1342, 2350, 305]. Decades [1164, 2404]. Decay [284]. Deceit [1099]. December [653, 748, 732, 1709, 1457, 671, 2057, 669, 2791, 1626, 1300, 64]. Decision [2419]. Decisions [1429, 1194, 350, 2112, 2419]. Declaration [1783]. Decline [2110, 1597, 2885, 2202, 2690, 1299, 2931, 306, 1631]. Decoding [977]. découverte [2571]. Dederick [1559]. Dedication [1804]. Deductive [846]. Dee [41]. Deep [1024]. Deeper [1080]. Defence [1785, 2185, 95]. Defended [821]. Defenders [2921]. Defending [2289]. Defense [1541, 2131, 807, 2758]. Defied [1087]. Defining [2169]. Degenhardt [1379]. degli [1643, 2338]. DeGregori [1919]. Deified [1087]. DeKosky [306]. Del [1038, 1765, 1000, 1543, 1685, 1781, 2592, 2751, 2160, 1681, 2627, 579, 2900, 1868, 1188, 1643, 1749]. Delaney [1713]. Delaporte [2641, 291]. Delatte [1050]. Delbruck [1949]. d’Elée [1274]. Delicate [386]. Dell [1492, 693, 2601]. Della [42, 2597, 1255, 77, 488, 579, 1790, 2338, 1845]. Dellian [2906]. Delmonte [1925]. Demaitre [501]. Democedes [2892]. Democracy [2353]. Democrat [1687]. Democratic [2623]. Democritea [1906]. Democrito [372]. Democritus [899, 778]. Demon [1522]. Demonstrated [1855]. Demonstration [2128]. Dempster [1355]. Dendrochronology [1800]. Denis [2782, 1906, 1323]. Dennis [2915, 2431, 1100, 1498]. d’Entrecasteaux [2403]. Department [81, 1290, 756, 2715]. Depew [1730]. Depression [2264, 1621, 1533]. depuis [2804]. Derek [2238, 2040, 1632, 2636, 2217, 348, 1171, 1538]. Derivations [2098]. Derral [2168]. Derrick [769]. Descarte [982]. Descartes [2050, 1180, 2594, 28, 1215, 1129, 1126, 531, 2084, 1283]. Descent [757]. Describing [1156]. Desert [2766]. Desiderio [1327]. Design [2466, 1938, 1168, 232, 1168, 1927]. Desire [2730]. Desmond [1048, 982, 1911, 2305, 1118]. deSolla [1538]. Despair [2684]. Despalatović [2826]. Desperate [2931]. Destitute [2645]. Determinantenkalkul [553]. Determinantentheorie [553]. Determines [1214]. d’Etudes [906]. d’Euclide [2594]. deutsche [2512, 715, 2330, 19]. Deutschen [1789, 1969, 1915, 2533]. Deutscher [1011, 2483]. Deutschland [1650, 2946, 706]. Deutschsprachige [2745]. Deutung [864]. deux [2773, 2804]. Develop [2934]. Development [2035, 1160, 1692, 2360, 2775, 1597, 302, 658, 1550, 2449, 2846, 2682, 2479, 352, 1421, 1549, 95, 710, 989, 404, 2422, 285, 907, 2472, 267, 927, 2518, 1226, 1145, 571, 306, 630, 1740, 619, 1889, 498, 137, 2253, 825, 2107, 2133, 1919, 679]. Developments [1746, 1755]. développement [1753]. deviations [651]. Device [439]. Devices [1674, 1932]. Devil [1539, 398]. Devonian [1880]. DeVorkin [948]. Dexter [599]. d’Histoire [1423, 2248, 483]. Dhombres [89, 690, 1397]. Dialectical [1728, 2262]. dialectique [414]. Dialogue [979, 1281, 1669, 1710]. Dialogues [2078]. Diamond [2077]. Diana [2930, 2670]. Diane [1956]. Diaries [1610]. Diary [2455]. Diaz [1917]. Dibner [2473, 2241]. Diccionario [1232]. Dick [2117, 867, 899]. Dickson [2669]. Dictionary [15, 203, 2125, 2140, 21, 885, 951, 2015, 815, 1233, 2780, 2780, 2186, 1088]. Dictionnaire [2782, 1369]. Did [1131, 1539]. Diderot [161]. Diego [1154, 1932, 2592]. Diesel [2227]. Diet [365]. Dieter [1574, 108, 891, 448]. Diethard [36]. Dietrich [1847, 547, 2483, 744, 1760]. Dieudonne [2248, 483]. Differential [2357]. Difficult [2296]. Diffraction [2844]. Diffusion [1345]. DiFilippo [2202]. Digby [1900]. Digital [1748]. Dijksterhuis [2729]. Dikshoorn [2729]. Dilemma [2504]. Dilemmas [2333]. Dilg [785]. Dilg-Frank [785]. Dilke [2891]. dilo [1352]. Dilworth [886]. Dilwyn [2126]. Dimension [2941]. Dimensions [1757, 880, 2880, 1832]. Din [2814]. Dinosaurs [2022]. Diophante [2391]. Diophantu [2391]. Diophantus [2391]. Dioptrique [2050]. Dioscorides [1678, 2010]. Diplomacy [2137]. Dirac [2460]. Direct [1276]. Direction [1243, 2884]. Directions [614, 543, 2879]. Directory [205]. Direktor [2740]. Dirichlet [1184]. Dirk [2674]. Disaster [2021, 1339, 2302]. discepoli [44]. Discipleship [59]. Disciplinary [1396, 2484, 2153, 594]. disciplinas [1917]. Discipline [1391, 1579, 298, 905]. Discontents [953]. Discontinuity [2251, 1006]. Discours [2050]. Discourse [2798, 2600, 1796, 1079, 1408, 2779]. Discovered [1322]. Discoverers [1461, 1581]. Discoveries [704, 1750, 43, 2662, 996, 2499, 422, 883]. Discovering [2531]. Discovery [1940, 2246, 2834, 611, 2319, 476, 678, 2656, 2069, 1018, 544, 932, 1109, 1263, 2787, 2256, 348, 6, 173, 1991, 1529, 735, 1484, 835, 2765, 2435, 1899]. Discrete [2673]. Discussion [2762, 1977]. Disease [2109, 1004, 1841, 2930, 2267, 2886, 1195, 2246, 2888, 32, 2856, 2641]. Diseases [777, 1596]. Disintegration [2894]. Diskin [1430]. Diskussion [836]. Dispute [941, 2757]. Disputes [2881, 328, 556]. Dissecting [2543]. Dissection [2645]. Dissent [2910, 2442, 618]. Dissertation [721]. Dissertations [1458]. Distance [1921]. Distancing [2539]. Distinguished [1611]. Distorting [1003]. Distributors [2475]. Disvoery [2703]. Disziplin [1576, 908]. Disziplinen [1650, 2720]. Disziplingeschichte [764]. d’Italia [604]. diverse [1846]. Diversification [1239]. Diversity [945]. Divided [2723]. divin [1175]. divina [528, 984]. divinatoire [639]. Divinatory [439]. Divine [528, 2458]. Divinity [2309]. Divis [1352]. Divise [1352]. Division [2722]. Djafari [1278]. Djebbar [975, 1979]. Djen [2588, 2347, 783]. DNA [1058]. Do [1539]. Dobzhansky [561, 837]. Dock [2365]. Doctor [1334, 501, 2365]. Doctoral [1458]. Doctorates [80]. Doctors [1425, 2632, 2485, 2301, 2851, 894, 1853]. Doctrina [1846]. Doctrines [110, 43, 1274]. Documentary [514, 122, 817, 2243, 1529]. Documentation [2703]. documenti [1543]. Doggett [2821]. Dohrn [1052]. Dokumentation [1660]. Dokumente [2132]. Dollar [2226]. Domain [1250]. Domin [2739].

Dominique [1754, 2760]. Don [1864, 1473]. Donahue [963]. Donald [1015, 2227, 1692, 2240, 2564, 614, 1639, 1333, 1752, 2496, 729, 962, 2102, 2587, 563, 2428, 2943, 1545]. Donegan [2111]. Doner [176]. Donley [2092]. Donna [754]. Donnini [2122]. Donors [1459, 1803, 2072, 2389]. Donovan [721, 43]. Dorinda [387, 1863]. Dorion [2030]. Doris [1584]. Dorman [2057]. Dorothy [380, 1158, 2297, 618, 788, 2146, 1687, 2144, 1244]. Dott [359]. Double [1716, 269, 2575]. Douglas [1746, 2667, 1571, 2117, 339, 1480, 2118, 2423]. Douglass [2874]. Dov [927, 464]. Dovring [2933]. Down [668, 1038]. Doyce [963]. Doz [963]. Dr [1652, 2286]. Dr. [47, 1497]. Drachman [1425]. Drago [2739]. Dragon [2759]. Drake [1512, 1964, 1787, 722, 1281]. Drake-Brockman [1787]. Drakenstein [2576]. Dransfeld [2808]. Draught [2214]. Drawings [636, 2103, 93]. Drawn [1651]. Drazen [544]. Dream [266, 568]. Dreams [1178]. drei [287]. Dresden [2924]. Dresdensis [715]. Drewnowski [120]. Dreyer [1995]. Drift [427, 2149]. Dritten [154, 2490, 353]. Drive [1459, 1803, 2072, 2389]. Dronke [2733]. Drugs [1272, 1760, 1955]. dualism [2954, 1416]. Duane [2861]. Dublin [513, 1150]. Dubos [2263, 2263]. Duchesne [2725, 83]. Duchesneau [920]. Dudich [1658]. Dudley [2916, 1561]. due [1592, 1035]. Duel [738]. Duerr [1094]. Duffy [1507]. Dufour [2753]. Dugald [511]. Duhem [77, 2416, 2124, 1734, 317]. Dukas [319]. Dumez [2829]. Dumont [1990, 2691]. d’un [624, 625]. Dunbar [1417, 1736]. Duncan [1518]. Dunes [1266]. Dunlap [687]. Dunn [951]. Dunwoody [1820]. durant [625, 2309]. Durbin [2949, 1499, 2944, 1364]. durch [2696]. Durham [1258]. during [2634, 1984, 2429, 2758]. Durkheim [2324]. Durling [1846]. Durr [1612, 1893]. Dutch [1156, 2040, 2576, 417, 2340]. Dutton [2930]. Duval [1578]. Dux [1094]. Dvorak [1468]. Dworsky [692]. Dwyer [2633]. Dyeing [818]. Dynamic [50, 2183]. Dynamics [2849, 352, 61]. Dynamis [1882]. Dynamos [277]. Dynasty [2190]. Dyson [2135].

E. [257, 900]. Eads [233]. Eagle [2406, 1373]. Earliest [1072, 1378, 1106, 2519, 1816, 996]. Early [344, 2239, 1605, 1692, 2223, 1166, 1307, 1484, 1345, 1453, 1853, 2644, 2131, 1644, 1004, 245, 302, 938, 2618, 2200, 809, 2821, 961, 971, 2446, 2689, 1098, 2543, 1020, 705, 2902, 2753, 771, 997, 778, 659, 2038, 2321, 737, 1380, 758, 743, 2288, 652, 2654, 2457, 2332, 2520, 2735, 1183, 242, 1962, 1341, 2018, 645, 2265, 1091, 147, 392, 1489, 2570, 569, 756, 1997, 904]. Early-Modern [147]. Earman [1305]. Earth [697, 1567, 93, 1103, 1541, 425, 53, 1022, 167, 1038, 2101, 1806, 1143, 1486, 645, 2030, 1622, 63, 2183, 2028, 1512, 645, 1103]. Earthquakes [698]. Easlea [789]. East [1166, 2399, 2772]. Eastern [2811]. Eastwood [205, 976]. Eatough [1511]. eau [1375]. Ebbinghaus [2031]. Eberhard [859, 45, 553, 2808]. Eberhardt [1379]. Eccles [244, 980]. Eckel [1010]. Eckert [1746]. Eclipse [1164, 1592]. Eclipses [1827]. Ecliptic [999]. École [2350, 958, 480, 1149]. Écoles [2349, 237]. Ecological [1359, 1421, 2029]. Ecologist [53]. Ecologists [2192]. Ecology [767, 1761, 338, 1298, 2397, 14, 1833]. Economic [2840, 2546, 710, 2578]. Economics [2015, 2620, 1367, 2578]. Economies [2449]. Economists [2922, 2015, 2193]. Economo [227]. Economy [2862, 2724, 1004, 530, 430, 2908, 2933]. ecrits [772]. Ed [2906]. Eddington [1611]. Edelstein [818, 1804]. Eden [907]. Edgar [1286, 1468]. Edge [317, 350]. Edición [442]. Edinburgh [163, 1573, 1859, 2036]. Edinye [2846]. Edison [2475, 795, 2475, 2475]. Edited [916]. Edith [1293]. Editing [2243, 1089]. Edition [1177, 1864, 1505, 1668, 1778, 1851, 928, 1951, 624, 283, 2220, 1440, 2043, 2475, 2373, 1152, 2080, 729, 2475, 1142]. Editionen [916]. Editor [262, 2794, 745, 2877]. Editorial [1794, 455, 1388, 994, 663, 2617, 2231, 2537, 140, 2795, 2463, 255, 527, 1065, 2296]. edizioni [2160]. Edmund [2500, 788, 2146]. Edna [2480]. EDSAC [1748]. Edsall [428, 588]. Eduard [207, 1983, 1758]. Education [614, 1067, 366, 2451, 1819, 1812, 2465, 956, 1238, 142, 1596, 691, 490, 1401, 844, 1889, 22, 1236, 2349]. Educational [564]. Educator [792]. Edward [1981, 1809, 2275, 2223, 1147, 779, 567, 1864, 2568, 1541, 2606, 1693, 2274, 373, 2648, 21, 312, 1600, 1325, 2448, 2420, 2572, 1140, 73, 713, 1526, 171, 1709]. Edwardian [2640]. Edwards [382, 1637, 2905]. Edwin [2574, 1108, 2582, 352, 1010, 1513, 1948]. Eeuw [649]. Effect [2541]. Effects [2114, 1487]. Efficiency [2847]. Effluvial [1259]. Effort [1975]. Egerton [1325, 14]. Eggleston [210]. Egmond [918]. Egypt [1331]. egzaktnih [1462]. Eheschliessungsvorgange [1908]. Ehr [1908]. Ehrman [2282]. eigentlich [1710]. Eighteenth [2439, 1282, 2267, 2451, 601, 2009, 725, 1444, 101, 1717, 648, 1823, 560, 124, 645, 576, 529, 1704]. Eighteenth-Century [2439, 1282, 2267, 601, 2009, 725, 1444, 1717, 648, 560, 124, 576, 1704]. Eighth [11, 1062]. Eilhard [2726]. Einde [649]. einem [1285]. Einfluss [1186, 1340]. Einfuhrung [1660, 2897]. Eingebettet [287]. Einheit [1468]. Einladungs [1760]. Einleitung [1306]. Einsatzgewichte [2469]. Einstein [2880, 2097, 1613, 1192, 394, 1193, 1887, 2254, 737, 319, 1826, 685, 133, 451, 2654, 1866, 1707, 1192, 1193, 142, 2715, 1303, 1138, 54, 657, 1483, 133, 2654, 955, 960, 2073, 2431, 2092, 2658]. Einstein. [133, 133]. Eisele [583, 2530, 583]. Eiseley [187]. Eisen [1544]. Eisenreich [1495]. Eisenstein [1380]. Elaine [2368]. Elasticity [692]. Elborg [2629]. Elder [2446]. Electric [1923, 1921, 2475, 400, 805]. Electrical [970, 939, 1920, 710, 1739, 2746, 2205]. Electricity [1032, 408, 710, 710, 1259, 2500]. Electrification [969, 1139]. électriques [1602]. Electrochemicals [774]. Electrochemistry [696]. Electrodynamics [996]. electrodynamique [1256]. Electromagnetic [2703, 2599]. Electron [485]. Electronic [2579, 1748]. Electronics [485, 33]. Electrons [989]. Elementarmathematik [283]. Elemente [916]. Elementen [1780]. elementis [2594]. Elements [1508, 976, 1680, 2594, 846, 961, 234]. Elena [1917, 2025]. Elevation [196]. Eleventh [2001]. Elias [1860, 1644]. Elimination [2927]. Eliot [1481, 1406, 2089]. Elisabeth [2551, 1738, 1309, 1525]. Elites [2916]. Elizabeth [76, 2246, 1575, 2276, 1487, 2555, 2376, 2368, 1588, 1681, 2928, 885, 1380, 2364, 2328, 158, 2831, 1412, 1197]. Elizabethan [470]. Elkana [1193, 133]. Ellen [445, 620, 2694, 2633, 2469]. Ellenberger [2810]. Elli [718]. Elliot [2442, 1750, 2931, 1580, 1997]. Elliott [15, 2203, 923]. Ellis [822, 1163, 2662, 2119]. Elliston [2086]. Elman [1898]. Elof [946]. Eloge [2716, 1538, 1135, 2155, 748, 2876, 1393, 1709, 1878, 2474, 2793, 2473, 672, 671, 1942, 2624, 1879, 1627, 537, 1001, 1626, 944, 2008, 1300, 465, 66, 264, 464, 334]. Eloges [811, 478]. Elogue [265]. Elsasser [132]. Else [2300]. Elshout [1443]. Elster [1362]. Elucidated [2678]. Elusive [495]. Embryo [53]. Embryology [2769, 648]. Emergence [831, 2680, 1667, 2835, 2683, 1316, 1796, 2515, 737, 2886, 1002, 1579, 2908, 97]. Emerging [216, 1559]. Emerson [1429, 2038]. Emigration [55]. Emil [1983, 987, 2132, 479, 2702]. Emilie [439]. Emilio [2499, 422]. Emily [2867]. Emmanuel [780, 1471, 438, 1591]. Emmy [867]. Emory [2075]. Emotions [2240]. Empedocles [847]. Emphasis [1544]. Empire [2446, 1865, 1912, 711, 1902, 1852, 1400]. Empiric [298]. Empirical [1714, 2878, 531, 1416, 2954]. Empiricism [530, 1719]. Empirisme [920]. Employment [407]. Emrich [2950]. Emulation [773]. Enchanted [912, 1127]. Encounter [364, 2921, 2093]. Encyclopedia [1102, 885, 337]. Encyclopédie [149]. End [290, 813, 1617, 2470]. Endangered [1209]. Ende [292]. Endeavor [159]. Endeavour [2103]. Endless [419]. Ends [697]. Enduring [1632]. energia [1868]. Énergie [301]. Energiebegriffs [1017]. Energy [1056, 1290, 970, 2133, 991, 2933, 1145, 51, 655, 2377, 1031]. Engel [2287, 1266, 1473]. Engelbert [1185]. Engelhardt [2881, 1659, 547, 209, 2483]. Engend’ring [2735]. Engeser [2586]. Engfer [1186]. Engine [843, 489, 1370, 368]. Engineer [2820, 1927, 1170, 2581, 2427]. Engineering [636, 2208, 2257, 970, 1500, 2753, 1920, 2205, 822, 1739, 194, 844, 2349, 231, 2558]. Engineers [2510, 416, 1032]. Engines [1428, 2158]. England [2888, 1425, 2399, 1520, 226, 491, 1245, 956, 2744, 179, 1474, 137, 2742, 1239, 490, 1096, 2539, 2464, 1703, 47, 2275, 980, 2638, 2736, 2640, 725, 2584, 1903, 1016, 1238, 1683, 2735, 646, 102]. English [1329, 1177, 1505, 435, 1402, 1819, 550, 2907, 1444, 640, 1438, 2277, 1637, 2214, 175, 1901, 2654, 2823, 2750, 1419, 1762, 1281, 145, 41, 510, 2525, 279, 493]. English-Language [435]. English-Speaking [2823]. Englishman [374]. Engstrand [1282, 963]. Enid [1676]. Enigma [2351, 1291]. enjeux [1410]. Enlarged [730]. Enlightened [577]. Enlightenment [650, 1859, 2216, 1116, 149, 1042, 2036, 43, 2652, 1402, 2283, 2282, 964, 919, 1183, 2684]. Enough [1299]. Enquiries [2748]. Enquiring [832]. Enquiry [2345]. Enrico [336, 1668, 960]. Enseignement [690, 237, 2170]. Enseignment [975]. Enstehung [108]. Entdeckung [897, 1830]. Enterprise [914, 1921, 13, 1930]. Entfremdete [209]. Enthusiast [1947]. Entities [1403]. Entomological [2835]. Entomology [2835, 628]. Entrance [2173]. entre [861, 2773, 1602]. Entstehung [282, 2720, 1017, 1650, 1777, 1572].

Entstehungsgeschichte [310]. Entwicklung [1186, 2854, 727, 82, 1655]. Entwicklungsdenken [603]. Entwicklunsgeschichte [836]. Enumeration [1205]. Environment [2458, 2861, 774]. Environmental [2362, 167, 1420, 2705, 1265, 2707]. Enzymes [700]. Ephemeriden [718]. Ephemerides [1686]. Ephemeris [977]. Ephemeroptera [1605]. Ephesos [1433]. Ephesus [775]. Epic [2589]. Epicuru [1588]. Epicurus [1430]. epidemias [2911]. Epidemics [1596, 2200]. Epidemiology [2619, 634, 2332]. Episodes [2569, 1833]. Episodic [1618]. Epistemologia [77]. Epistemological [652]. epistemologicos [2025]. Epistémologie [89]. epistemologique [574]. epistemologiques [677]. Epistemology [198, 2327, 2321, 2943, 1949, 2598, 1948]. epistolicum [1983]. Equality [2681]. Equations [2357]. Equatoires [438]. Era [2771, 1087, 2490, 1873, 126, 1027, 1954, 2375, 894, 2758]. Erasmus [1048, 2873, 2305]. Eratosthenes [999, 597]. Erdgestalt [1485]. erdwissenschaftliche [1660]. Erebus [1322]. eretico [1542]. Erfahrung [1306, 1945]. Erfahrungsraum [1019]. Erhaltung [1523]. Erhard [241, 1021]. Eric [832, 577, 137, 2501, 909, 2287, 2624]. Erice [544]. Erik [2422]. Eriksson [1284]. Erinnerungen [2533]. Erkennens [1019]. Erkenntnis [1468]. Erlen [1737]. Erlenmeyer [2702]. Erna [1981, 716, 73]. Ernan [153, 2557, 2094]. Ernest [225, 2326, 208, 1528, 33, 350, 341, 126]. Ernst [271, 1055, 836, 2286, 1949, 247, 2678, 2864, 2502, 1470, 2061, 154]. Error [2488]. Erste [38, 715]. ersten [241, 117]. Erster [2904, 479, 38]. Eruption [1487]. Erwin [384]. escribió [2592]. Esin [1931]. España [861, 1765, 1613, 1781, 2914, 442, 1232]. español [861]. española [443]. españoles [444, 1782]. Espionage [2379]. espiritu [1327]. Esq [2527]. Essai [776, 291, 2594, 624]. Essais [2880, 2050]. Essay [408, 1949, 692, 285, 1948]. Essays [1785, 2239, 888, 2309, 1575, 377, 511, 676, 583, 1115, 1644, 1775, 354, 731, 825, 1842, 678, 1886, 2655, 208, 151, 2486, 1512, 167, 1034, 307, 778, 2178, 1719, 1330, 1137, 1751, 1988, 2419, 1641, 686, 2673, 1664, 2668, 1629, 99, 230, 2448, 2865, 2011, 2093, 297, 2023, 2493, 1404, 2331, 2076, 366, 1417, 2516, 1318, 2090, 2356, 2076, 2359, 2209, 2181, 1336, 2885, 2837, 613, 1557, 1492]. Essential [1246, 2668, 2450]. Establishment [1545]. Estate [2750]. Estes [909]. estestvoznaniia [74]. Esther [2282]. Estimating [620]. Estudios [1860]. eta [1503]. Etats [2563]. Etats-Unis [2563]. Ether [695]. Ethic [1168, 1421]. Ethical [204]. Ethics [119, 2293, 1420, 226, 842]. Ethlie [2583]. Ethnographic [1330]. Ethnological [1898]. Ethnology [1173]. Ethology [559, 1105]. Etienne [854, 1547, 381, 311]. Etiology [1208]. étoiles [581]. étrangères [500]. Ettling [770]. Etude [975, 712, 1274]. Euclid [1508, 234, 846, 1176]. Euclide [1780]. Euclidean [2941]. Euclideana [916]. euclidienne [2809]. Eudoxischen [178]. Eudoxos [178]. Eugene [2076, 1289, 1985, 986, 2515, 1557, 411, 2460, 2358]. eugenesia [1987]. Eugenics [1759, 1805, 2847, 1702]. Eugenio [1782, 1232, 1346]. Euklid [916]. Euler [693, 1983]. europaisch [762]. europaisch-amerikanischen [762]. europaischen [1595, 2210, 2696]. Europe [387, 2811, 2055, 2545, 2029, 1821, 22, 679, 2821, 2902, 2277, 1380, 1339, 2046, 1961, 2467, 147]. European [2013, 756, 417, 711, 1279, 238, 2760, 2619, 1808]. Eustaquio [1846]. Evald [682]. Evaluating [2195]. Evaluation [1299]. Evan [831, 890]. Evandro [610, 609]. Evangelical [2921]. Evangelista [2934]. Evans [2168, 734, 2923]. Evelyn [53, 1267]. Everett [1528, 1629]. Everitt [1412, 2831]. Everyday [2779]. Everywhere [2853]. evich [2832]. Evidence [1187, 342, 2211]. Evolution [2395, 2309, 2563, 1243, 395, 2775, 2091, 1164, 1764, 2680, 970, 821, 1562, 2033, 1544, 1921, 1773, 1539, 25, 2400, 945, 2396, 2188, 2396, 929, 2030, 1547, 1053, 1831, 1548, 1187, 766, 1730, 2189, 2681, 2094, 1264, 1968]. Evolutionary [1949, 467, 2394, 2683, 1250, 1406, 1832, 686, 2921, 839, 2233, 1836]. evolutionisme [1323]. Evolutionismus [2355]. Evolutionist [2864]. Evolutionists [187]. Evolving [220, 60]. Ewan [951, 951]. Ewen [2018]. ex [1368]. Exact [1344, 852, 642, 1462]. exakten [1777, 1572]. Examinations [2789]. Examining [2211]. Excavations [1570]. Excellence [2667]. Exceptional [1289]. Exchange [2123]. Exercise [778]. Exercises [2594]. Exhibition [1644, 133, 133, 2102]. Exile [1525]. Existence [972]. Expanding [720, 1261]. Expansion [2029, 415, 1239]. Expedición [2914]. Expedition [2105, 1051]. Expeditions [1913, 1282]. Experience [2638, 2606, 2088, 1210, 498, 446, 2862, 2840, 1930, 2482]. Experiences [1647]. Experiment [1960, 838, 1390, 2808, 2321, 1835, 2464, 2463, 1572, 1281, 965, 1601, 2468]. Experimental [2769, 495, 2883, 2466, 1970, 1667, 1986, 807, 2031]. Experimental-psychologische [2031]. experimentale [726]. Experimentalvorlesung [1153]. experimentelle [2031]. Experiments [1078, 809, 2543, 1405, 2470]. Expert [269]. Experten [2199]. Expertise [2947]. Explaining [2442, 2556, 1362]. Explanation [1147, 1146, 1724, 2481, 888, 2778]. Explanations [683]. Exploration [1166, 1693, 1881, 211, 1163]. Explorations [2656, 1491]. Explored [1012, 486]. Explorer [318, 704]. Exploring [2105]. Explosion [1529]. Explosives [1504]. Exposition [1580]. Expositions [1865]. Expository [1742]. Express [105]. Expression [2240]. Extant [847]. Extensive [1958]. External [1982]. extraordinaria [1613]. Extraparliamentary [618]. Extraterrestrial [899, 2082]. Extreme [2722]. Eye [151, 1490, 865, 2191]. Eyes [1357]. Eyler [314]. Ezell [1693, 1693].

F [2439, 1809, 660, 1583, 860, 1025, 2327, 2700, 395, 1407, 96, 1418, 2657, 270, 2180, 1613, 731, 1842, 1146, 1713, 971, 756, 1316, 2917, 2136, 2479, 2600, 2753, 2057, 1370, 2843, 2853, 987, 373, 2213, 159, 2515, 1524, 572, 1232, 885, 1088, 1773, 33, 932, 451, 2408, 1866, 1689, 2939, 2831, 900, 388, 1412, 2666, 172, 1449, 2106, 2905]. F.R.S [548]. F.R.S. [2527]. fa [571]. fa-chan [571]. Fabiani [2751]. Fabio [2597, 1255, 1749]. Fabric [1918]. Faces [2265, 2880]. Fachausdrucke [1595]. Fachprosa [1275]. Fachprosa-Studien [1275]. faciles [2043, 114]. facsímil [442]. Facsimile [2401, 1439, 722]. Fact [267, 2481]. Factors [2395]. Facts [2654, 2354, 156, 2662]. Faculty [2079]. Faherty [2528]. Failed [594]. Failure [566, 1927, 31]. Failures [2442]. Fair [1865, 268, 1955]. Fairs [1621]. Faith [1200, 1784, 1910, 2847]. Fakultat [1475]. Falk [1290]. Fall [2534, 2543, 1173, 2671]. Falling [2499]. Fallstudie [1830]. Fallstudien [764, 1572]. Falschung [2254]. False [2488]. Falsely [2126]. Families [971]. Family [2408, 2141, 2741]. Famous [1512]. Fan [1621]. Fancher [224, 1895]. Fancy [327]. Fang [2450, 378]. Far [1051]. Faraday [1988, 960, 530, 1887, 1988]. Farben [2490]. Farber [1886, 1579]. Farewells [2617]. Farley [1324, 28]. Farm [2273]. farmacia [1332, 2627]. Farmall [2273]. Farmer [1355]. Farmers [1835]. Farming [1345, 2214, 2933]. Farms [2509]. Farquhar [771]. Farr [314]. Farrall [270, 616]. Farthest [697]. Fast [2271]. Fasting [2886]. Fate [37]. Father [2149]. Fattori [2051]. Faul [1320, 1320]. Fausz [1809]. Fauvel [685, 2432]. Favoro [1181]. Faw [812]. Fawcett [228, 2344, 608]. Faye [812]. Fears [1968]. Feather [2077]. Febrer [2751]. February [537, 334]. Federal [2237, 2087, 95, 2377, 140]. Federale [2563]. Federov [2523]. Federspiel [305]. Fee [2246]. Feeding [2647, 2198]. Feeling [1267, 2303]. Feferman [1959]. Feher [900]. Feier [479]. Fein [2172]. Feinberg [2024]. Feingold [1474]. Feldman [2442]. Felice [723]. Felix [870, 1824, 1204, 1873]. Fellmann [1983]. Fellows [1098]. Feminism [2417, 840]. Feminist [2282, 1368]. Fennoscandia [2811]. Ferdinand [736]. Ferenc [1529]. Fermat [1743]. Ferment [601]. Ferments [403]. Fern [1501]. Fernand [2824]. Ferretti [1188]. Ferrone [1517]. Festschrift [2095, 2728]. Festvortrag [479]. feu [218]. Feuer [2823]. Fever [175, 2332, 1498, 1208]. Fevers [1971]. Feyerabend [1095]. FF [1398]. Fi [308, 1439, 1440]. Fichman [924]. Fiction [2920, 1406]. Fictions [621]. Field [1418, 2846, 771, 402, 2430, 2522, 2599]. Field-Notes [771]. Fields [1455, 216]. Fieldwork [1330]. Fiery [2206]. Fierz [1178]. Fifteen [1032]. Fifteenth [2815, 2816, 1961]. Fifteenth-Century [2816, 1961]. Fifth [252, 2622]. Fiftieth [1111]. Fifty [1484, 2032, 2086, 2250, 1899, 2844]. Figala [271]. Fighters [835]. Figural [2334]. Figure [63]. Figured [1263]. Figures [224, 2433]. Figurovskii [2832, 285]. Fildes [2198]. Filopono [372]. filosofia [1189, 1890, 1442]. filosofici [2160]. Filosofo [2595]. Fin [911, 237, 1335, 979, 1375]. Fin-de-Siècle [911]. Final [1083, 1547]. finde [1640]. Fine [2097]. Fink [576]. Finn [2475, 367]. Finocchiaro [508]. Finucane [1885]. Fire [2220, 2348, 893, 893]. Fire-Clocks [2348]. Firearms [2120]. Firms [2938]. First [1032, 1415, 2579, 1508, 1604, 1701, 2136, 2846, 640, 1040, 18, 390, 729, 1292, 2823, 2328, 238, 2718, 2, 1955, 410, 2724]. Fisch [1864, 2374, 928]. Fischer [1949, 1970]. Fischler [2722]. Fish [126]. Fisher [1759, 2112, 125]. Fishery [2368]. Fishes [1263]. Fishlock [2928]. fisica [1255, 372, 2053, 1255]. fisico [1188]. fisico-matematico [1188]. fisiki [2846]. Físiques [1192]. Fiske [1487]. Fission [1290, 1529]. Fitch [2835]. Fitchen [2204]. Fitness [1054]. Fitzpatrick [925]. Five [2231, 1151]. Fizikalna [2739]. fiziki [660]. Flamm [2685]. Fleck [1306]. Fleetwood [1583]. Flegg [1958]. Fleming [1527]. Flesche [741]. Fletcher [1827]. Fleuves [500]. Flew [2912]. Flexner [1954, 2632]. Fliess [2857]. Flight [2270, 2936]. Flink [2580]. Flora [1166, 378, 2550]. Florence [1000, 1494, 581, 1853, 2056]. Florey [321, 2228]. Floris [1604]. Florkin [66, 362]. Florman [2581]. Floyd [2559]. Flu [1211]. Fludd [240, 240]. Flyer [2208]. Flying [1372]. Focus [2836]. Folch [2627]. Folgen [2254]. Folia [130]. Folke [2933]. Folkerts [2095, 916]. Folklore [1337]. Follette [1244, 1014]. Folse [1720]. Fond [824]. fondements [573, 2942]. Fontana [723]. Food [1955]. Forbes [2624, 577, 1585, 990, 1901]. Forbidden [2865]. Force [720, 1112, 2935, 2118, 1145, 1684, 2118, 898]. Forces [2939]. Ford [1056]. Fordson [2273]. Forecasting [2935]. Foreign [2890]. Forensic [1901]. Foreseen [971]. Forest [1237]. Forester [2514]. Forgan [793]. Forg’d [2742]. Forgive [31]. Forgotten [2921, 1980, 2583]. Form [1480, 1630]. Formal [2089]. Forman [133]. Formation [273, 1013, 2519, 2506, 1149]. Formative [1096]. Formed [412]. Forms [1742]. formulario [1037]. Formwalt [2221, 1933]. Forrester [492]. Forschung [2854]. Forster [2629]. Fort [1206, 1993]. Fortifications [2120]. Fortschritt [2512]. fortuna [2051]. Fortwirken [178]. Forward [2271]. Fossil [361, 2377, 1263]. Foster [2558, 13, 841]. Foucault [726, 2799]. Foul [2352]. Foundation [2786, 656, 1096, 1644, 1669, 1077, 700]. Foundations [1954, 508, 1943, 546, 62, 324, 2361, 2004]. Founder [318]. Founders [1404, 2732]. Founding [767, 2247, 1641, 2018, 294]. Four [796, 2030, 778, 705, 359]. Fourcy [2350]. Fourteenth [180, 2584, 1977]. Fourteenth-Century [2584, 1977]. Fourth [378, 2941]. Fowler [2428, 2692]. Fox [1267, 2887, 584, 2680, 2088, 2220, 218]. Fraassen [1719]. Fracastoro [1511, 1511]. fragments [845, 1274, 847]. Fragrant [2352]. Frame [1258]. Framework [1720]. francais [625]. francaise [1149, 585, 223, 280]. francaises [500]. France [46, 2563, 614, 1631, 2736, 2257, 708, 813, 1667, 714, 312, 584, 1912, 1754, 618, 1002, 2424, 1375, 853, 2236, 1124, 1704, 1535, 1863, 302, 2699, 2442, 1564, 1323, 107, 1004, 2009, 1003, 921]. Frances [470, 672]. Francesc [444, 1782]. Francesca [2449, 1679]. Francis [1987, 1786, 911, 2186, 2553, 2285, 318, 1512, 1910, 296, 879, 741, 1740, 1786, 2176, 897, 2051, 2524]. Francisco [2592]. Francissen [1605]. Francois [958, 1090, 1744, 920, 1783, 291, 2810, 2376, 388, 2641]. Frangsmyr [1353]. Franjo [2739]. Frank [2455, 14, 2443, 2815, 2091, 2446, 1248, 2080, 1988, 1325, 932, 1258, 73, 510, 785, 1662]. Franke [1204]. Frankland [2223]. Franks [870]. Franksen [1747]. Frans [1605, 2155]. Franz [1201, 1660, 1990, 1470]. Fraser [2927]. Fraud [2488, 1099]. Fred [269, 2749]. Frederic [1983, 744]. Frederick [1809, 2289, 1227, 2542, 2719, 1962]. Frederique [1673]. Fredson [2209]. Free [1095, 1855]. Freeland [2411]. Freeman [1103, 2135, 339, 129, 471, 2225, 1171]. Freemasons [1116]. Freer [1931]. Frege [356, 1570, 357, 1570]. Frei [1824]. Freiberg [1847]. Freidrich [1340]. Freidson [2089]. Freiesleben [27]. Freitag [2205]. Fremont [1913]. French [650, 1854, 2349, 2446, 318, 2606, 1867, 725, 1417, 364, 317, 1476, 2352, 2404, 133, 844, 1676, 2451, 21]. Frenchmen [614]. Frerichs [2441]. Fresquet [2751]. Freud [1424, 16, 2857, 2505, 1666, 1422, 191, 466, 2706, 245, 1108, 1991]. Freudenthal [2476, 1469]. Freundschaft [2132]. Fridolf [497]. Friedel [2475, 1670]. Frieden [1027]. Friedman [2092, 1820, 2780]. Friedrich [1055, 684, 736, 1838, 1499, 985, 1558, 356, 2287, 1296]. Friend [1855]. Friendly [2757]. Frijhoff [615]. Frisch [188]. Frischtak [1919]. Frits [417]. Fritz [2726, 22, 643, 75, 1019, 1566, 1777, 856, 798]. Front [1, 58, 139, 193, 251, 254, 326, 397, 454, 520, 523, 526, 590, 662, 740, 801, 804, 875, 936, 993, 1061, 1064, 1123, 1213, 1295, 1384, 1387, 1452, 1532, 1616, 1697, 1700, 1793, 1871, 1937, 2000, 2003, 2065, 2151, 2230, 2315, 2318, 2383, 2462, 2536, 2610, 2615, 2616, 2711, 2785, 2872, 664]. Frontier [796, 2224, 923, 419, 2247]. frontiere [1326]. Frontiers [1163]. Frontiersmen [2936]. FRS [791]. Fruchtman [1444]. Fruhdrucke [916]. fruhen [1186, 2822, 1789]. Fruhzeit [823]. Fruitful [932]. Fruits [53]. Fruton [588, 952]. Fuad [1439]. Fuat [116, 641]. Fuel [2377, 1296, 2933]. Fulke [182]. Full [1413]. Fuller [2101]. Fullerton [126]. Functionalism [1664]. Functions [1742, 1396]. Fund [1459, 1803, 2072, 2389, 2631]. Fundamental [2846]. Fundamentalist [2542]. fundamental’nykh [2846]. Funding [2429]. Funghi [2122]. Fungus [835]. Funkenstein [2310]. Furley [1505, 2519]. Furnace [2206]. Furumoto [2328]. Fusion [991]. Future [151, 1671, 2428, 619, 2039].

G [46, 2737, 1408, 1187, 1111, 2703, 2676, 1915, 2359, 2910, 1425, 2884, 93, 163, 2402, 510, 1543, 1784, 53, 2585, 186, 1432, 2371, 2833, 45, 2496, 1158, 2225, 248, 577, 2917, 2902, 858, 574, 1771, 1322, 2843, 2117, 494, 1836, 513, 2056, 1578, 2014, 2490, 657, 685, 2855, 1637, 2603, 781, 1589, 2377, 932, 959, 907, 560, 1666, 765, 1554, 2344, 1337, 2049, 2831, 2420, 1515, 388, 1412, 1570, 2772, 1482, 2800, 1257, 2844, 949, 2106, 206, 891, 2523, 1315, 498, 2818, 2255, 695, 2868, 2358, 1498]. Gabbard [2637, 2637]. Gabler [2582]. Gabriel [953, 356, 2436]. Gabriele [1566, 637]. Gad [2476]. Gago [442]. Gailuzzi [383]. Gajdusek [771, 771]. Galambos [2534]. Galaxy [2515]. Gale [1603, 1187]. Galen [1505, 1779, 110, 303, 1202, 2125, 1505, 1202, 2518]. Galenic [976]. Galianosi [2125]. Galieni [849, 1846]. Galilean [1000, 1281, 1128]. Galilei [44, 2281, 1543]. galileiani [1685, 383]. Galileo [44, 1115, 1543, 1784, 2598, 1750, 609, 1783, 1281, 1688, 1799, 2543, 2790, 195, 1126, 2712, 1622, 1389, 1125, 1875, 1181, 2526, 508, 2281, 242, 2454, 1542, 722, 1688]. Galishoff [2642]. Galison [2470]. Gallagher [2487]. Galleons [2131]. Galler [24]. Gallery [1931]. Gallerye [182]. Galluzi [1181]. Galluzzi [44, 1856, 2820, 1685]. Galton [1987, 879, 2789]. Galvani [2219]. Gambit [2084]. Game [2097]. Gametes [1324]. Gandt [388]. Gaposchkin [1395]. Garber [2831, 1412]. Garcia [1781, 1846, 1154, 1932, 2592]. Garcia-Diego [1154, 1932]. Garden [907]. Gardener [1995]. Gardeners [1644]. Gardens [415]. Gardner [1896]. Garfield [2076, 1557, 411]. Garin [1346]. Garland [12, 691]. Garma [861]. Garnet [1921]. Garry [902]. Gärten [2896]. Garvin [1101]. Gary [409, 1713, 2656, 1417, 87, 1736, 2654]. Gas [620, 571]. Gascoigne [2552, 1816]. Gases [2831]. Gassendi [2697, 1285, 2053]. Gaston [155, 351]. Gathering [2137]. Gattopardi [2595]. Gaukroger [1283]. Gaunse [2823]. Gauss [1558]. Gautrey [2748]. Gavin [2555, 2102, 680]. Gay [1185, 2706, 2505, 70, 625]. Gay-Lussac [70, 625]. Gaylord [2408, 1581]. Gazda [2368]. GE [1923]. Geared [25]. Gears [2307]. Geburtstages [479]. Gedachtniss [2031]. Geddes [2628]. Gedenkbibliothek [473]. Gegensatze [2287]. Gegenwart [2031]. Geheimnis [309]. Geiger [2838, 963]. Geison [1476, 1311, 13]. Geistesgeschichte [2696, 1275]. Geisteslinien [916]. Gelbsucht [1433]. Gelehrte [1104]. Gelehrten [1990, 426, 2199]. Gelfand [565]. Geltung [864]. Gender [1944, 2301, 1797]. Gene [2503]. Geneeskunde [206]. General [2185, 1923, 1381, 95, 285, 359, 2926, 805, 2627, 500, 1707]. générale [301]. Generalis [241]. Generated [645]. Generation [777, 2385, 974, 648, 876, 940, 2328, 28, 2259]. Generations [781]. Generous [1769]. Genes [946]. Genese [1469, 772]. Genesis [766, 2590, 229, 1576, 394, 267]. Genetic [2327, 1058]. Geneticist [929]. Genetics [2363, 2154, 144, 837, 1805, 2386, 2503]. genetique [1323, 2259]. Geneva [143, 1042]. Genéve [2555]. génèvois [1602]. Genie [2935]. Genius [289, 2216, 793, 1995, 2190, 1734, 2746, 1450, 1634]. Genocide [2851]. genootschappen [2825]. Genre [2883]. Genres [1637]. Gentle [1053]. Gentleman [1355]. Gentleman-Farmer [1355]. Gentlemanly [1880]. Gentlemen [2873, 2288]. Geodasie [1485]. Geodesy [597]. Geoffrey [1511, 1916, 516, 2684, 2511, 1238, 2334]. Geoffroy [2404]. Geographen [1104]. Geographic [2765]. Geographical [697, 166]. Geographie [762, 379, 1149]. Geography [289, 814, 1965, 1417, 1736, 1869, 2186, 557, 2893]. Geologen [763]. Geological [834, 1658, 95, 833, 1010, 1163, 359, 2256]. geologiche [488]. Geologie [1576, 2810, 2436]. Geologiegeschichte [426]. Geologist [1418, 1190]. Geologists [1418, 951, 1964, 472, 2554]. Geology [558, 1559, 2748, 2185, 1320, 95, 360, 2361, 472, 359, 1419, 2554, 49, 487, 1964, 1045, 1559, 1023]. Geomagnetic [549]. Geomagnetism [2258]. geomancie [639]. Geomancy [439]. Geometrica [1980]. Geometrical [2128, 2522]. Geometrie [2050, 1035, 1743, 280, 2814, 2809, 1780, 236]. Geometry [808, 1744, 2941, 1590, 236]. Geomorphology [2027]. Geophysics [1829, 2182]. Geophysik [764]. Georg [859, 2901, 1976, 1319, 2566, 162]. George [2920, 1448, 1427, 2949, 1121, 1970, 2521, 2339, 1406, 2927, 1978, 2832, 829, 176, 1923, 159, 972, 1751, 1921, 584, 1664, 137, 2671, 2408, 230, 792, 1962, 92, 1581, 1694, 1783, 2639, 2493, 1031, 2331, 1066, 1619, 1618, 2542, 1481, 1911, 1191]. Georges [387, 234, 228, 2604, 2407, 1863]. Georgia [2558]. Geowissenschaft [1659]. Geowissenschaften [426]. Gerald [1493, 2650, 1103, 1271, 32, 2253, 1193, 2059, 220, 602, 33, 1476, 72, 1311, 13, 2024].

Geraldine [1809]. Gerard [677, 1819, 1472, 613, 343, 185]. Gerasa [111]. Geraud [1117]. Gerd [2652]. Gereon [2254]. Gerhard [154, 2700, 1203, 161, 891]. Gerhardt [2813]. Gerhart [1540]. Gericke [1672, 283, 2095]. Gerlof [1840]. Germ [2768, 770]. Germain [692]. Germaine [2558]. German [1840, 1222, 2485, 1417, 1013, 55, 612, 2333, 871, 1707, 59, 5, 1796, 665, 2153, 1296, 1702, 2623, 322]. German-Jewish [322]. German-Speaking [55]. Germania [2601]. Germany [477, 757, 2927, 2932, 2758, 1239, 490, 2154, 1620, 140, 2227, 577, 618]. Gernot [2560, 209]. Gerrit [2058]. Gerrylynn [1689]. Gersh [2447]. Gerson [1851, 1775]. Gerstein [1899]. Gertrude [2655]. Gesammelte [1306, 1612, 1893, 1686]. Gesamtausgabe. [1909, 1288]. Geschichte [2950, 1485, 1011, 706, 358, 1830, 830, 1655, 19, 1640, 2807, 1094, 2031, 862, 2489, 91, 1278, 1915, 424, 116, 641, 1021, 27, 908, 283, 637, 891, 1141]. Geschiedenis [206]. Gesellschaft [46, 1011, 108, 19, 2483, 2950, 275, 497]. gesellschaftlichen [1576]. Gesetz [2245]. Gestirne [1340]. Gesundheit [1340]. Gesundheitswesen [1593]. Getting [457]. Geuter [1969, 1915]. Gevitz [1270]. gewesen [1710]. Gezelschap [649]. Ghada [238]. Ghalioungui [1331]. Ghent [1072, 2544, 1907]. Ghinst [1071]. Ghislaine [2628]. Ghost [2762]. Ghostwriting [2076]. Giacconi [336, 1791]. Giambattista [1668, 42]. Gian [2567, 2673]. Gian-Carlo [2567, 2673]. Giancarlo [1442]. Gianni [604, 2160]. Giant [369]. Giants [1735]. Gibbs [1978]. Gibian [2808]. Gibson [2247, 244]. Gideon [1469]. Giebelhaus [2558]. Gieben [1435]. Giere [2556]. Gifford [137]. Gift [67]. Gigerenzer [2652]. Gigon [1906]. Gilbert [1408, 2634, 958, 489, 806]. Gilded [2301]. Gilded-Age [2301]. Gill [1748]. Gille [371, 2859]. Gillespie [316]. Gillian [1819, 1406]. Gillispie [1182]. Gillmor [1829, 2182]. Gilly [2696]. Gilman [165, 2856]. Gilson [1547, 311]. Gingerich [1413, 288]. Gingras [2725]. Gino [506]. Giordano [1706]. Giorgio [1188]. Giovan [1188, 506, 2280]. Giovanni [2595]. Girls [2886]. Girnus [2726]. Girolamo [1511, 1819, 933, 1130, 1178]. Gisela [2245, 168]. Gittins [1638]. Giuliano [1326, 1643]. Giulio [372]. Giuseppe [1905, 555, 1179, 552, 506]. Given [1146, 109, 359]. Gjertsen [1632, 2217]. Gladys [1973]. Glanvill [398]. Glas [207]. Glaser [826]. Glass [1127, 561]. Glasses [2348]. Glazer [2173]. Gleanings [2390]. Glebov [2532]. Glen [2637, 1022]. Glenn [872, 150]. Glick [1232, 1613]. Glimpses [319]. Glimpsing [1316]. Global [1290, 2148, 890]. globe [574]. Globig [2950]. Gloelampenfabrieken [2040]. Glomar [2766]. Gloria [2301]. Glossar [1595]. Glusker [1260]. Glymour [342]. Glyndwr [964]. Goal [1243]. Gobar [1237]. Goblot [505]. God [2179, 1446, 1875, 88, 2093]. Goddess [216]. Goddu [2044]. Godel [1959]. Godet [2469]. Godless [2505]. Gods [328, 1248]. Godwin [2147]. Godwins [897]. Goethe [161, 2055, 2915, 2285]. Goethezeit [1990, 763]. Goetzmann [1051, 2435]. Goggin [2353]. Gohau [2436]. Gold [1575]. Goldberg [2312, 1648]. Golden [2117, 1892, 2756, 418]. Goldenberg [2020]. Goldhaber [2420]. Goldhammer [2641]. Goldkind [2244]. Goldman [1522]. Goldschmidt [929]. Goldsmith [449, 2295]. Goldstein [1851, 947, 1848]. Goldstine [623]. Golley [2443]. Golomb [1949]. Gomez [1502]. Gomis [1819]. Good [1610, 866, 2258]. Goodfield [735, 688]. Gooding [1988]. Goodly [182]. Goodman [2591]. Goodstein [2080]. Gorden [2666]. Gordon [501, 1321, 2304]. Goring [1916]. Gorn [2935]. Gornick [1242]. Gospel [1372]. Gosse [339]. Got [2358]. Gottfried [2738, 356, 1339]. Gotthard [2896]. Gotthelf [2585, 2212]. Göttingen [1873, 2489]. Gottleber [729]. Gottlob [356]. Gould [2681, 2256]. Goupil [624]. Gourevitch [1506]. Governing [2353]. Government [2685, 2688, 2237, 1716, 2486, 2690, 1455, 2591, 2377, 1478, 2947]. Gowing [350]. Grabiner [693]. Grace [990]. Graceanna [214]. Graceful [1775]. graduate [1401]. gradués [615]. Graf [2117, 1158]. Graham [1786, 1958, 1396, 302, 986, 2860, 365, 196, 2366, 1234]. Grand [969, 2043, 848, 617]. Grande [73]. Grand’Henry [1594, 376]. Granollachs [2900]. Grant [2448, 2693, 779, 1541, 1140]. Grapes [2079]. Grassblade [1872]. Grattan [484, 973]. Grattan-Guinness [484, 973]. Grau [891]. Graumann [1838]. Graves [2360, 320]. Gray [2176, 2624, 2357, 2432]. Grayson [1545]. Grazia [1749]. Grazyna [2816, 1466]. Grdenić [2739]. Great [321, 1665, 1946, 224, 489, 1254, 934, 1167, 868, 987, 2117, 1741, 160, 1261, 2435, 964, 1621, 2460, 1419, 1880, 1533, 144, 1405, 1501, 2931, 2802]. Greater [1102]. Greatest [2821, 2379]. Greatrakes [592, 591]. grec [1431]. greca [1845]. greci [2122]. greco [1506]. greco-romain [1506]. grecque [2392]. grecques [305]. grecs [845, 371]. Greece [2693, 1337]. Greek [2012, 2445, 2125, 305, 1336, 1269, 778, 2519, 1976, 2520, 498, 2717, 938, 1674]. Greek-Armenian [2125]. Greeks [2731]. Green [2106, 1028]. Greenaway [2446]. Greene [1325, 1023, 686]. Greenwich [348]. Greenwood [1764]. Greg [2583]. Gregoras [1592]. Gregorian [1315]. Gregorio [1608]. Gregory [152, 1160, 1959, 1490, 977, 2258, 2478, 683, 450]. Greiner [2726]. Grelon [2773]. Grene [2372, 1832]. Greppin [2125]. Greta [550, 2341]. Grete [1762]. Gridgeman [344]. Griechen [36]. griechische [497]. Griffin [2379, 213, 54]. Griffith [513]. Griggs [1028]. Grigorijan [1983]. Griminger [2654]. Grimwood [248]. Grinstein [2298]. Grmek [544, 1207, 304, 582, 363, 1431]. Grob [2650, 1271]. Grocer [347]. Groeben [1052]. Groneman [2299]. Grossbritannien [1308]. Grosse [556, 1104]. Grosseteste [1435, 2277, 1341, 2621, 978]. Grounds [211]. Group [725, 2357]. Grove [489]. Growing [2560]. Growth [80, 23, 2303, 2938, 2277, 1773, 945, 284, 350, 841, 2197, 2482, 701]. Gruender [610, 609]. Grund [475]. Grundlagen [1029, 2906]. Gründung [2763]. Gruppe [2498]. Gudehus [86]. Guenard [958]. Guenter [2036]. Guericke [856]. Guerlac [1878, 922, 2852, 790, 235, 676]. guerres [2773]. Guerrini [2781]. Guerrino [2727]. Guest [1794, 455, 1388, 994, 663, 140]. Guide [819, 1947, 428, 2553, 2949, 2236, 2030, 2339, 821, 640, 2080, 750, 2195, 542, 1637, 973, 951, 1142, 2243, 2126, 2162, 2098, 2662, 2170, 810, 2414, 2868, 2555, 1399, 55, 754]. Guides [373]. Guido [1253, 703, 2160, 544, 2338].

Guilford [2052]. Guillaume [2753]. Guillerme [1375]. Guillermo [2627]. Guinness [484, 973]. Guiton [2636]. Guldin [1132]. Gulyga [2826]. Gun [2119]. Gundolf [1203, 1275, 1340]. Gunn [1166]. Gunnar [719, 1779, 1284, 343]. Gunpowder [2589]. Gunston [2443]. Guntau [2720, 426, 1576]. Gunter [1094, 1990]. Gunther [1949, 988, 358, 1824, 759, 502]. Gupta [784]. Gursch [836]. Gustaf [1309]. Gustafsson [984]. Gustav [1184]. Gustavianska [1284]. Gutenberg [2515]. Guthorn [1577]. Gutting [409, 1713]. Guy [1653, 2037, 1348]. Gvishiani [619]. Gwei [2588, 2347, 783]. Gwei-Djen [2588, 2347, 783]. Gwen [1465]. Gwendy [2167]. Gwyn [321, 1527]. Gymnasia [1455]. Gynaecology [980]. Gynecological [640]. Gypsies [2927].

H [953, 1759, 2716, 1809, 1282, 34, 271, 1160, 1608, 2597, 965, 1612, 1893, 2834, 2634, 559, 636, 557, 1508, 1604, 495, 2748, 1639, 1864, 1959, 2692, 2881, 1205, 2374, 298, 2527, 1049, 575, 125, 934, 1170, 1206, 768, 1332, 623, 2148, 165, 649, 1089, 2558, 2689, 2907, 1690, 948, 970, 1112, 1730, 2701, 556, 1241, 904, 817, 2080, 2057, 928, 100, 2869, 2117, 2038, 2648, 729, 763, 2935, 1751, 2646, 55, 542, 1921, 265, 2101, 2058, 2347, 1494, 2899, 211, 758, 2893, 2659, 105, 869, 1586, 562, 2435, 755, 791, 919, 415, 635, 1587, 2666, 1051]. H [1492, 1839, 2596, 359, 2116, 323, 1825, 2106, 2197, 909, 1239, 2351, 2482, 1031, 1404, 946, 116, 641]. H. [556, 2924]. H.-P [1612, 1893]. H.M.S [48]. ha-Me [1202]. Haas [1922]. Haber [2136]. Hacker [1570, 2759]. Hacking [886, 1721]. Hackmann [2062]. Haddad [1439]. Haeckel [2502]. Haeckels [836]. Hafertepe [1562]. Haffter [2728]. Hagen [1760, 2107]. Hagerstrand [2083]. Hagley [2162]. Hahn [1656, 1047, 340, 1035]. Haines [2081]. Hainsworth [2523]. Hale [170]. Half [2846, 1486]. Half-Century [1486]. Halina [514]. Hall [2347, 2426, 2106, 2737, 2269, 509, 1097, 685, 1224]. Haller [1354, 648, 633, 1158]. Haller-Wolff [648]. Halleux [845]. Halley [1752, 1705, 2714, 1757]. Hallion [2936]. Halls [1099]. Hallyn [2824]. Halpern [2849]. Hamann [502]. Hamburg [2923, 2807]. Hamburger [2769]. Hamilton [1607, 1614, 564, 539]. Hämmerle [174]. Hammond [1151]. Hammurapi [1745]. Hamon [1237]. Han [378]. Hanbury [2308]. Handbook [640, 2217, 2553]. Handlist [471, 884]. Handmaiden [2232]. Handschrift [715]. Handschriften [916]. handschriftlichen [2819]. Hanen [494]. Haneveld [206]. Hankins [2283]. Hanle [133, 930]. Hannah [1032, 710]. Hannaway [1960]. Hanno [289, 1104]. Hannover [46, 108]. Hans [91, 1186, 2769, 2566, 2180, 1094, 134, 2590, 988, 931, 2274, 1379, 178, 27, 1034, 2292, 2808, 1830, 133, 652, 356, 2489, 1141, 655]. Hans-Christian [27]. Hans-Gunther [988]. Hans-Joachim [1379, 178, 2489]. Hans-Jurgen [1186, 133]. Hans-Werner [1830]. Hansen [1778, 2427]. Hansjakob [1949]. Happened [595]. Hard [2419]. Harden [2087]. Harding [2417]. Hardware [1428]. Hardwick [1786]. Hardy [2326]. Harem [533]. Harford [586]. Harley [2545, 2268]. Harlow [2205]. Harman [1145, 1145, 1989]. Harmon [80]. Harmony [380]. Harnessing [2935]. Harold [1950, 2110, 2663, 50, 2117]. Harre [1405, 2418]. Harrie [2708]. Harriet [2750]. Harrigan [614]. Harriman [1051]. Harrington [2575]. Harriot [1347]. Harris [2032, 766, 84]. Harrison [2568]. Harrower [1328]. Harry [2786, 2147, 676, 1997, 1751, 317, 1912, 2673, 1101, 1193]. Hart [2741]. Hartcup [1653]. Hartenstein [2740]. Hartner [811, 1776]. Hartree [1746]. Hartwig [32]. Harvard [1067]. Harvesting [2275]. Harvey [2631, 2892, 2285, 1645, 173, 680, 675]. Harwit [1018]. Hasan [1276]. hasard [1743]. Hashimi [1439]. Haskell [97]. Hasler [862]. Hassan [2587, 238, 308]. Hassel [1470]. Hatch [1398, 2561]. Haubelt [1352]. Haupt [2745]. hauptsachlich [1026]. Havas [2654]. Hawk [2302]. Hawkes [756]. Hawkins [2780, 1293]. Hawrylchak [55]. Hawthorn [2684]. Hay [1958, 1440]. Hayawan [504]. Hayes [2490]. Haygood [2375]. Hayles [2430]. Hays [2707, 2707]. Haytham [1844, 1276]. Hayyan [2392]. Hazen [834, 834]. Headrick [711]. Heal [1889]. Healed [2440]. Healers [2638]. Healing [2037, 2049, 230]. Health [2113, 2771, 1199, 2246, 1556, 1004, 770, 1426, 640, 2803, 2088, 2172, 2648, 2646, 365, 1054, 1737, 1209, 296, 2114, 2707, 1596, 1110, 2111, 2440, 1200, 2088, 299]. Health/Medicine [1200]. Healthiest [1110]. Hearnshaw [2325, 293, 2250]. Heart [1425, 2303, 566]. Hearth [1365]. Heat [1182, 922, 2678, 2320]. Heath [1568, 1146, 1101]. Heath-Stubbs [1568]. Heavenly [2347]. Heavens [855]. Heaviside [2917, 2746]. Heavy [302]. Hebe [1917]. Hebert [2056]. Hebrew [1202, 330]. Hechler [419]. Hedwig [503]. Heelan [2945]. Heerding [2040]. Heers [717]. Hegel [414, 388, 2704, 1467]. Heidecker [1566]. Heidelberger [2652, 1945]. Heidi [867]. Heijenoort [1959]. Heil [2825]. Heilberufe [2504]. Heilbron [2551, 949, 1984, 961, 1554]. Heilkunde [2819]. Heilkunst [36]. Heims [420]. Hein [1760]. Heinekamp [108, 2738]. Heinen [1440]. Heinrich [385, 547, 2489, 2701]. Heinz [1011, 2808, 1046, 2504]. Heinz-Peter [2504]. Heisenberg [1612, 1893, 451, 1555, 1866, 1525, 1525]. Heitmann [2444]. Held [2103, 909]. Helden [1757]. Hele [2737, 1048, 2305]. Helen [1344, 319, 2559]. Hélène [2476, 2403]. Helfand [1332]. Helga [1178]. Helge [2548]. Helix [269]. Hellenismus [497]. Hellenistic [1174, 597]. Heller [2884, 1784]. Helmholtz [1523, 2132]. Helmont [1043, 2278]. Helmreich [2137]. Helmuth [1672, 2095, 2830, 283]. Hematology [1975]. Hemerobio [1605]. Henderson [1066, 1597, 2941]. Hendrik [2576]. Hendry [1526, 2599]. Heniger [2576]. Henk [627, 827, 2472]. Henning [1566]. Henri [1343, 505, 507, 2753, 1319]. Henry [1933, 2634, 51, 636, 653, 676, 395, 361, 2376, 732, 1320, 731, 922, 1720, 2057, 1907, 339, 2221, 1815, 2719, 341, 417, 2852, 25, 1927, 2747, 2156, 790, 1587, 126, 1996, 2351, 1878, 2320, 731, 2375, 2528, 1183]. Her [84, 1322, 213]. Heranziehung [764]. Herausbildung [1576]. Herbal [1028, 1676]. Herball [1762]. Herbals [2400, 2081]. Herbarium [437]. Herbert [2455, 2748, 282, 1146, 264, 2656, 515, 1926, 606, 827, 1017, 2420, 353, 863]. Herbolario [437]. Hercules [1415]. heredite [1819]. Heredity [1759, 2152, 2735, 1805]. Heretic [2454]. Heretics [2802]. Herf [2492].

Heribert [1909]. Heritage [2769, 1018, 2019, 1688, 1766, 2074]. Herman [2105, 623, 932, 2420, 2079]. Hermann [1523, 2132, 1026, 451, 1866, 1443, 1573, 2760, 2031]. Hermes [356]. Hermeticism [2902]. Heroes [971]. Herries [1321, 513]. Herrmann [1574]. Herschel [1574, 1610, 2600]. Hertz [2701, 2703]. Hertzsprung [1574]. Herve [280, 2555, 2829]. Herz [2722]. Herz-Fischler [2722]. Herzenberg [2140]. Herzlich [2629]. Hesse [474, 2343, 545]. Hetherington [2569]. Heuer [187]. Heuser [2702, 2702]. heute [2807]. Hew [2207]. Hexaemeron [1435]. Heyck [1016]. Heyd [1042]. Heyday [728]. Heyer [1005]. Heyningen [1210]. Heytesbury [1977]. Heyward [1479]. Hibner [1691]. Hideyo [1390]. Hiebert [384]. Hiertas [824]. Higgins [2780, 1571]. High [2685, 796, 2042, 2121, 2709, 2442, 2938, 597, 236, 2253, 2514]. High-Accuracy [597]. High-Speed [796]. High-Tech [2709, 2514]. Higham [543]. Higher [2451, 2922, 1239]. Highlights [322]. Highway [2510, 105, 2111]. Hilary [2643, 421]. Hilbert [1824]. Hilbert-Felix [1824]. Hilfstein [73, 1981, 716]. Hilgard [2326, 225]. Hill [1237, 2927, 858, 2587, 828]. Hillel [725]. Hilliard [2757]. Hillier [122]. Himmelfarb [2655]. Himself [1461]. Himsworth [1950]. Hinde [1105]. Hindenburg [2117]. Hindle [834, 1286, 1671, 1376, 773]. Hine [2453, 445]. Hinkle [294]. Hinsley [630]. Hintergrund [287]. Hintergrunde [556]. Hintikka [610, 609]. Hinton [350]. Hipocratica [1904]. Hippocrates [110, 2911, 1507, 1849]. Hippocratic [113, 777, 1849, 2264, 1202]. Hippocratica [304]. hippocratique [304, 1589, 1506]. Hippolytus [2520]. Hirendranath [784]. Hirsch [1196]. Hirschfelder [515]. Hirsh [1316]. Hirst [791]. Hispanica [1882]. Histoire [1488, 2043, 2513, 2942, 315, 2476, 2392, 690, 1090, 853, 89, 2725, 2350, 2810, 2266, 714, 2804, 1323, 301, 2513, 2436]. Histoires [1495]. Historia [1917, 861, 2900, 1860, 2627]. Historiam [1882]. Historian [2236, 1078, 884, 923]. Historians [1623, 2744, 591]. Historic [99]. Historical [222, 1940, 2016, 225, 1273, 2326, 366, 119, 631, 1556, 245, 1193, 825, 207, 1235, 1550, 966, 2479, 1269, 1137, 1087, 2516, 1093, 685, 2345, 417, 2628, 2325, 1818, 2858, 2668, 68, 940, 242, 1816, 1948, 1762, 967, 2093, 2212, 2356, 1031, 901, 1828, 2530, 2478, 1584]. Historically [2678]. Historicized [1664]. historico [1232]. Histories [796, 1078]. Historiography [601, 2548, 2549, 431, 1948]. Historiometric [1634]. historique [726, 1274]. Historisch [1152]. Historisch-kritische [1152]. Historische [637, 1572, 287]. Historischen [1019, 764, 908]. History [1794, 941, 558, 728, 2260, 2075, 2166, 29, 857, 2198, 1785, 47, 569, 1084, 1252, 1905, 2799, 2035, 2497, 2848, 1972, 1548, 629, 1559, 2406, 2723, 1756, 1583, 1428, 169, 455, 1111, 2246, 1231, 336, 610, 819, 1745, 2326, 1166, 697, 969, 614, 2573, 1329, 2269, 380, 2359, 676, 2569, 683, 1491, 2530, 1658, 1025, 30, 1081, 797, 2224, 2017, 2115, 2811, 1103, 2027, 1303, 2733, 952, 471, 702, 1165, 1580, 1833, 1388, 428, 1092, 2553, 1974, 859, 32, 2892, 243, 915, 2109, 746, 994, 1636, 934, 2099, 168, 362, 423, 1320, 768, 191, 670, 1553, 2647, 681]. History [24, 692, 623, 2694, 2148, 1842, 2480, 333, 1089, 2222, 1108, 369, 67, 1638, 1520, 1997, 150, 2237, 2234, 1574, 1076, 535, 2339, 248, 2655, 663, 948, 208, 2500, 2902, 2040, 2582, 149, 725, 220, 2552, 146, 1841, 2684, 1097, 2274, 106, 122, 904, 1417, 2507, 817, 2181, 2465, 2545, 696, 968, 2859, 295, 1674, 1169, 2576, 167, 2843, 176, 368, 416, 2853, 2869, 587, 2242, 2894, 2164, 2178, 564, 1712, 2206, 1635, 346, 701, 95, 1010, 2648, 1463, 2258, 103, 1227, 1501, 300, 1168, 613, 2019, 2086, 1461, 2014, 2241, 2118, 205, 962, 818, 1088]. History [2101, 2719, 1601, 1058, 252, 521, 802, 1062, 1385, 1698, 2001, 2316, 2611, 2613, 1964, 2427, 33, 1054, 1221, 1737, 1238, 686, 407, 827, 229, 2168, 1582, 2272, 1537, 2068, 2893, 1736, 2426, 973, 2103, 447, 1818, 2434, 2781, 913, 1113, 20, 285, 2501, 401, 1896, 2400, 1929, 2378, 99, 562, 2004, 1755, 2537, 2939, 634, 230, 894, 765, 1077, 2432, 2674, 335, 927, 1715, 1318, 1554, 412, 238, 1028, 2578, 855, 1806, 1822, 1024, 472, 140, 1226, 882, 1079, 2334, 258, 1464, 2323, 181, 571, 783, 2622, 2119, 1258, 1587, 1922, 563, 2011, 596, 1492, 1007]. History [419, 1057, 1317, 1772, 1304, 1419, 255, 680, 1230, 1065, 2554, 1758, 949, 1894, 2722, 73, 713, 1462, 2549, 609, 2742, 297, 1197, 406, 1357, 1509, 1082, 1891, 286, 810, 628, 893, 2273, 695, 2868, 2039, 1529, 410, 405, 2792, 2070, 64, 1133, 881, 1457, 2652, 669, 2791, 332, 2387, 1801, 1526, 2041, 84, 2451, 1333, 2167, 1396, 2218, 2561, 678, 277, 1965, 1978, 484, 816, 1034, 1107, 1268, 2258, 2755, 1310, 2705, 1325, 2855, 1582, 1843, 2721, 1265, 543, 734, 2587, 340, 2255, 2676, 14, 2269, 2026, 768, 1829, 2182, 2672, 487, 1884, 2107, 2708, 2033]. History [826, 907, 363, 873, 2718, 1963, 2760, 1198]. Hitherto [359]. Hiyal [308]. Hobbes [1986, 198, 2279]. Hobbs [645]. Hobby [1973]. Hobhouse [491]. Hochheiser [1922]. Hochschule [2830]. Hochschulfach [82]. Hochschulreform [154]. Hocker [2081]. Hoddeson [1414]. Hodge [695]. Hodges [1291]. Hodierna [2595, 2280]. Hodson [1861]. Hoffmann [319]. Hofmann [1861, 2702]. Hogendijk [1844]. Holbrook [2396]. Holbrow [2666]. Holland [2747]. Hollandsche [649]. Holley [432]. Hollis [1477]. Hollywood [2271]. Holton [1193, 602, 72]. Homar [1782]. Home [1638, 458, 1259, 1463]. Homer [659]. Homes [2633]. Homme [223]. hommes [1819, 1819]. Honest [1684]. Honor [2239, 1575, 676, 1886, 1629, 2448, 73]. Honoured [1855]. Hooke [579, 1806]. Hooker [1719]. Hoover [606, 2774]. Hope [1610, 537]. Hopes [1968]. Hopkins [1111, 2246, 2106, 1333, 2565]. Hoppe [823]. Hoppen [1150]. Horace [2365, 2411, 2843]. Horizontal [1921]. horlogerie [438]. Hornberger [182]. Horne [1809, 1933, 2221]. Horowitz [2218]. Horses [2214]. Hortus [2576]. Horwitch [2291]. Hoskin [1315, 901]. Hoskovec [2836]. Hospital [2035, 1425, 1852, 1425, 2036]. hospitales [1037]. Hotbed [2216]. Houghton [221, 2874]. Hounshell [2937]. House [2464]. Household [1365]. Howard [1954, 1273, 2208, 2664, 1967, 233, 2765, 1896, 1649, 1334, 1348, 321, 2228]. Howe [2075]. Howell [2780]. Howlett [1428]. Howse [348]. Howson [1238]. Hoyer [1120]. Hoyrup [2300]. Hoyt [2360, 320]. Hrvata [1462]. Hsia [377].

Hsing [2588]. Hsing-Tsung [2588]. Hsü [2766]. Hsuin [2370]. Hu [1205]. Huang [2588]. Hubbel [844]. Hubert [49, 2720, 552]. Huddersfield [2643]. Hudson [1195, 2882]. Huemer [2850]. Huemul [1868]. Huerkamp [2199]. Hufbauer [1013]. Huff [1663, 892]. Hugh [1329, 1580, 1771, 2177, 1448]. Hughes [349, 850]. Hugo [2854, 627, 170]. Hugonnard [1319]. Hugonnard-Roche [1319]. Hui [378]. Hui-Lin [378]. Huisman [2782]. Hull [416]. Hulse [2106]. Hulst [627]. Human [2564, 1918, 1970, 1834, 1855, 8, 159, 2681, 701, 2940, 1919, 1795, 319, 1805, 1005, 1956, 1927, 413, 1729, 1339, 1545, 2482, 2188, 2438, 816, 170]. Humanism [121, 161]. humanisme [2127]. Humanismus [643]. Humanist [2396, 235]. Humanitarian [214]. Humanities [2671, 2396, 2431, 411]. Humanity [654]. Humans [1764]. Humboldt [514, 1184, 385]. Hume [369, 2912]. Humphry [793, 359, 359]. Hunayn [976]. Hundred [2138, 18, 1169, 548, 1642, 2086, 252, 521, 802, 1062, 1385, 1698, 2001, 2316, 2611, 360, 2765, 1151, 2250, 2106]. Hünemörder [2807]. Hungarian [1040]. Hunt [2555, 12, 1106, 754, 2550, 2491, 962, 902]. Hunter [2439, 1662, 1644, 1206, 1098, 368, 1929, 646]. Hunters [361, 2252]. Hunting [329, 789]. Husain [1726]. Hutchinson [53]. Hutchison [2908]. Hutton [93, 721]. Huw [2628]. Huxley [1608]. Huygens [9, 2128, 575, 921]. Huyssen [621]. Hyde [586]. Hydrographic [2833]. Hydropathic [2111]. Hygiene [2246, 2847, 2651]. Hyman [926, 2527]. Hymes [1492]. Hypatia [2019]. Hypochondria [2626]. Hypotheses [531]. Hypothesis [1216, 1960, 1137]. Hysteria [1991].

I-II-III [1276]. I-XLIII [837]. I. [609]. I.G [2490]. Ia [130]. iadernoi [660]. Iain [777, 1566, 350]. Ian [886, 697, 2281, 1721, 1250, 846, 1312]. Ianniello [1749]. IBM [1429, 2038]. Ibn [1037, 308, 1594, 2392, 1439, 2391, 2895, 976, 1276, 1844]. Icarus [2270]. Ice [2184, 94]. Iceland [2747]. ich [1640]. Ichi [944]. Iconography [514]. Ida [817]. Idea [2082, 2598, 1167, 2889, 1548, 395]. Ideal [691]. Ideas [2406, 2652, 312, 1324, 657, 1964, 675, 314, 1341, 2537, 915, 166, 686, 229, 983, 1313, 2583, 312, 1982, 955]. Ideengeschichte [441]. idées [585, 2392]. Identification [2083]. Identitat [908]. Identity [2268, 744]. Ideological [136]. ideologici [703]. Ideologie [353, 1966]. Ideology [2798, 2490, 686, 1311, 1337, 1049, 421]. iectigatione [849]. Ignas [2209]. Ignaz [1208]. Ihde [384]. ihre [2254, 1026]. ihren [1861, 2702]. ihrer [727, 2489, 82, 1655]. ihres [985]. II [1285, 1864, 2259, 2904, 2458, 1909, 1860, 2254, 2247, 2652, 2221, 2118, 2591, 1866, 1929, 238, 1276, 2415, 106, 2541, 610, 2182, 2480, 1983, 1336, 2852, 1776, 2379]. III [1981, 604, 653, 511, 2185, 375, 1951, 37, 1913, 1276, 92, 2350, 1484]. IIIe [1375]. IIIe-debut [1375]. ilal [1439]. Ilgauds [2566]. Ilkka [2085]. Illinois [1559, 81]. Illness [2638, 1271, 2856, 2931, 2629]. Illustrandum [1882]. Illustrated [1166, 1807, 1599, 549, 2587, 1556]. Illustrating [1559, 678]. Illustrationen [836]. Illustrations [983]. illustré [482]. im [1969, 1278, 1186, 154, 2822, 836, 1838, 1915, 2700, 1203, 1094, 1789, 1019, 1469, 2838, 547, 503, 2199, 174, 1593, 985, 2490, 2897, 1204, 497, 798]. imagerie [1342]. Imagery [1727]. Images [96, 1556, 2887, 2856, 350, 133, 167, 1719]. imaginaire [1343]. Imagination [1128, 621, 2310, 1126, 72, 2671, 1569, 2352]. Imagined [735]. Imagining [2039]. Imaginings [2626]. Imbrie [94, 94]. Immanence [937]. Immanuel [855, 1747, 2826]. Immobility [1541]. Immortality [478, 1855]. Impact [1746, 370, 2360, 2048, 1648, 1647, 2206, 1903, 657, 33, 2560, 2273, 1596]. Impacts [794]. Impartial [1855]. Imperial [119, 2131, 2409, 1097, 1620]. Imperialism [2029, 711]. Impetustheorie [424]. Implications [1840, 2606]. Importance [1182]. Important [312]. Impresa [1326]. impresos [1782]. Impressions [2134]. Impure [1211]. Imre [152]. Inc. [2378]. Incandescent [2040]. Incident [2302]. Including [1759]. Incorporating [1281]. Index [524, 632, 2661, 607, 753, 2140, 311]. Indexes [2453, 2949]. Indexing [411, 412]. India [1166, 2422, 1596, 2732]. Indian [96, 1173, 1279, 635, 741, 1392]. Indiana [1266]. Indians [1038]. Indicators [1828]. Indiens [848]. Indies [859]. Individual [1080]. Individuals [915]. Individuum [1469]. indivisibilibus [2898]. Induction [888, 217, 2797]. Inductive [474]. industria [1009]. Industrial [2035, 2227, 2175, 2866, 1004, 2349, 971, 369, 2689, 368, 1923, 621, 1168, 2202, 2602, 1929, 2158, 232, 2939, 2302, 194, 2806, 2261, 1773, 1030]. Industrialization [274, 665]. Industrie [174]. Industriekultur [1566]. industrielle [301]. industriellen [2854]. Industries [1370]. Industry [2021, 1429, 2685, 934, 302, 52, 774, 2861, 1097, 710, 665, 2490, 1773, 2444, 1504, 571, 1296, 2506, 805, 1921, 2040, 2860, 2202, 2506]. Ines [1584]. Infancy [1415]. Infant [2198, 2647]. Inference [807]. inferieurs [1819]. infini [2433]. Infinite [162]. Infinity [2124, 851]. Infirmary [2036]. Influence [186, 133, 900, 1160, 2882, 2446, 2596]. Influences [252, 521, 802, 1062, 1385, 1698, 2001, 2316, 2611, 2613, 1407]. Informal [2167]. Information [250, 519, 800, 1060, 1383, 1696, 1999, 2314, 2609, 2953, 2076, 333, 2546, 621, 542, 1557, 2514, 2799, 1230]. informatique [2513]. Infrared [2005]. Ingenieur [499]. Ingenieure [1011, 2854]. ingénieurs [2773, 1369]. Ingold [2189]. Ingolstadt [2822]. Ingrid [2902]. Inheritance [2503, 945]. inizi [1845]. inizio [1749]. Inkster [1312]. Inkunabeln [916]. Inner [2650, 1525]. inneren [547]. Innovation [2203, 1345, 971, 1454, 1921, 2214, 1030, 932, 2005, 2860, 2934]. Innovationen [2897]. Innovator [2821]. Inorganic [829]. Inquest [2187]. Inquirers [1953]. Inquiries [1634]. Inquiry [1200, 1194, 602]. Insane [2265]. Insanity [709, 102, 2639]. insegnamento [1255]. Insight [2677]. Inspiration [1975]. Institut [1475]. Institute [1111, 2555, 333, 2080, 1795, 754, 2550, 392, 2623, 2555, 1025, 1916]. Institution [2167, 1562, 2241, 1738, 630, 1098]. Institutional [2634, 2484, 417]. Institutionalization [2089]. Institutions [1084, 80, 565, 2423, 1535, 412]. Instituts [2763, 2807]. Instrument [347, 2284, 2843]. Instrument-Maker [2284]. Instrument-Makers [347]. Instrumentation [2166, 1007]. Instrumente [1470]. Instruments [2723, 1746, 780, 756, 1471, 1728, 2348, 549, 2347, 1952, 438, 756, 613, 756, 1472, 1024]. Insufficiency [2321]. Insulin [1109]. Insurance [2113, 1199, 2172]. Insuring [2021]. Integration [593]. Intellect [564]. Intellectual [16, 2902, 2485, 1034, 840, 2541, 1813, 1016, 543, 99, 2108, 2046, 2073]. intellectuelle [2657]. Intelligence [2244, 189, 1895, 2244]. Intelligentsia [2385, 104]. Intelligibilität [503]. intensionalen [2245]. Interaction [610, 550, 675]. Interdependence [2511]. interdisciplinaire [507]. interdisciplines [906]. Interessengemeinschaft [2490]. Interest [2112]. Interests [1313, 23]. Interférences [500]. Interioribus [1846]. International [2929, 2269, 1639, 1025, 2811, 1000, 746, 150, 2606, 2840, 2485, 1865, 1625, 1134, 1554, 238, 472, 2622, 2554, 2623, 906, 507, 9, 1919]. Internationale [1423, 1288, 785, 86]. internazionale [1685]. Interpretation [737]. Interpretations [2276, 592, 591, 2800]. Interpreter [1178]. Interrelationships [810]. Interstate [105]. Intertwist [1892]. Intervening [1721]. Interviews [406, 2496]. introduccion [1765, 442]. Introduction [1928, 1177, 1370, 1042, 103, 789, 2548, 1007, 350, 2421, 2926, 2336, 794, 376, 2075, 2412, 1888, 2325, 2691, 765, 111]. Introductions [2617]. Introductory [32, 1721, 484, 2164]. Introduzione [2595]. Inventario [1782]. Inventing [2087]. Invention [1182, 2789, 665, 1921, 2583, 991, 2880, 2475, 773]. Inventions [2889, 2662]. Inventor [736]. Inventors [2660]. Inventory [1554, 1082, 1398]. Inversion [403]. Investigaciones [2593]. Investigations [1561]. Invisible [1229, 1129, 1316]. Invitational [1760]. Ioffe [660]. Ionists [805]. IQ [1895, 734]. Ira [199]. Ireland [1321, 1482, 1583]. Irene [76]. Irenee [1071]. Iris [1282, 560, 963]. Irish [1885]. Iron [1773, 2162, 752]. Ironies [1365]. Irrationale [1094]. Irrationality [2828]. Irving [2077]. Isaac [1603, 1602, 1136, 2419, 1733, 2906, 538, 1514, 2596]. Isabel [1662, 391]. Isaiah [516, 1533]. Isham [2845]. Ishaq [976]. ISHTCP [1082]. Isidore [63]. Isis [2231, 412, 1877, 667, 412, 534]. Islam [2392, 1204]. Islamic [1345, 1344, 2390, 439, 2734, 2587, 39, 1440, 1931]. Island [1290, 196]. Isles [2086, 680]. Ismael [1398]. Isomerism [2319]. Isomorphismus [1830]. Israel [2480, 2475, 1388]. Issadore [50]. Issue [558]. Issues [2032, 2585, 1464, 1898, 204, 2032, 1031]. istorii [285]. Istoriia [2549]. istoriografiia [2549]. Italia [703, 1326]. Italian [918, 781, 2818, 1130, 608]. italiana [1517]. italiane [1332]. Italiani [1643]. italienischen [1379]. Italy [2818, 1794, 1957]. itinéraire [2604]. Ito [1176]. Iu [273]. Iushkevich [281, 74]. IV [2289, 1847, 777, 732, 1983, 2293, 600, 2391, 301, 1676, 375, 373, 642]. Ivan [733]. Ivo [499, 827]. Ivor [2917]. IX [1188].

J [946, 1759, 2075, 1809, 2239, 431, 2238, 2455, 1972, 1548, 1360, 2723, 1756, 2762, 2677, 1844, 1423, 1428, 1780, 2275, 2597, 1665, 2703, 903, 2834, 221, 2240, 2110, 2695, 1765, 1657, 1746, 2138, 1237, 511, 614, 2748, 84, 2667, 899, 2175, 2815, 1864, 1505, 2766, 2373, 2555, 1159, 1984, 2884, 489, 2298, 516, 2027, 469, 1096, 1066, 1266, 1580, 2374, 2376, 1164, 1418, 2657, 575, 1784, 2866, 2099, 2922, 423, 1764, 2184, 1355, 2183, 166, 2105, 2092, 1432, 1887, 2496, 2055, 2285, 679, 961, 1886, 1970, 2225, 1638, 1520, 1867, 2683, 2442, 420, 2294, 1720, 1730, 3, 2845]. J [2084, 151, 1577, 1771, 1447, 234, 2545, 2080, 868, 295, 1951, 1322, 2923, 1512, 2348, 1370, 2576, 167, 2843, 2584, 2938, 2054, 2652, 2038, 2128, 701, 1524, 2441, 494, 729, 384, 2346, 1349, 1168, 1473, 1988, 2405, 2014, 657, 1088, 33, 690, 2551, 1805, 229, 360, 2168, 2754, 486, 1592, 1818, 2721, 2550, 2378, 449, 2746, 158, 2674, 964, 1554, 1486, 2420, 1478, 2580, 1228, 1992, 1846, 1033, 2675, 2428, 2011, 923, 1839, 1120, 2171, 2250, 1740, 2729, 949, 1157, 1694, 909, 713, 619, 297, 576, 2415, 173, 893, 1198, 2414, 695, 1748, 963, 1991]. J [2522, 2039, 1892, 2188]. J. [2836]. Jaakko [610, 609]. Jabir [2392, 2392]. Jack [221, 2724, 1444, 2288, 1312]. Jackie [712]. Jackson [2667, 1346, 2264]. Jacksonian [365]. Jacob [2673, 1116, 2698, 1183, 831]. Jacobs [2886, 270, 150]. Jacobson [2632]. Jacques [1594, 376, 429, 717, 137, 2858, 2391, 1174, 2127, 772]. Jacyna [2574]. Jaggi [1596]. Jagtenberg [1243]. Jahnke [652]. Jahr [985]. Jahre [626, 452, 715]. Jahrestages [1288]. Jahrhundert [1915, 2199, 916, 91, 1186, 2822, 2700, 762, 2817, 547, 1908, 1576]. Jahrhunderten [287]. Jahrhunderts [2897, 82, 916, 292, 2287]. jährigen [2807]. Jaki [88, 215, 1734, 2306, 2416, 855, 286]. Jakob [151, 727]. Jallabert [1602]. James [2767, 1646, 81, 1289, 965, 2267, 2695, 2359, 2555, 2186, 2362, 2159, 2209, 768, 2585, 1358, 1522, 654, 1713, 698, 1690, 957, 248, 1684, 375, 1812, 1322, 1750, 416, 307, 978, 564, 1635, 833, 1751, 548, 885, 2427, 160, 1054, 2130, 1823, 1598, 2102, 2550, 2179, 2271, 1142, 1715, 752, 2580, 1183, 1101, 635, 2367, 2197, 1357, 576, 963, 490, 2455, 2916, 1988, 1626, 1692, 51, 93, 986, 2599, 721, 2559, 733]. Jamic [308]. Jamie [2453]. Jamin [223]. Jamison [2422]. Jan [2656, 910, 2894, 2808, 2058, 2503, 1279, 571]. Jane [2289, 2286, 2682, 1814, 2169, 2111, 1961]. Janet [2453, 1165]. Janine [2629, 1207, 111]. Janis [981, 2350]. Jannetta [2200]. Jansky [1484]. Janssen [897]. January [1538, 653, 811, 2876, 500, 732, 2057, 252, 521, 802]. Japan [2422, 239, 2200, 707]. Japanese [2271]. Japon [2563]. Jaramillo [1022]. Jarausch [1239]. Jarcho [1668, 566]. Jardin [2555, 1348]. Jardine [1785]. Jarrell [1464, 410]. Jarrett [1722, 2270]. Jaskot [2821]. Jay [2681, 1902, 2851, 2256]. Jayawardene [884]. Jazari [308]. JCAST [1083]. Jean [2263, 1162, 227, 1959, 2248, 1857, 2115, 2216, 483, 2050, 2699, 708, 507, 2691, 756, 223, 89, 1602, 214, 2728, 1042, 1397, 1319, 1036, 1591, 2809, 2926, 2170, 963]. Jean-Baptiste [963]. Jean-Claude [2699, 2809]. Jean-Henri [507]. Jean-Michel [2170]. Jean-Paul [2691]. Jean-Robert [2050, 1042]. Jeanne [2904, 432]. Jeannine [541]. Jefferson [925, 1460]. Jeffrey [2041, 2492, 2835, 2857, 2221, 2475, 232, 1422, 1828]. Jelliffe [1666]. Jenkin [2141]. Jenkins [956, 1869]. Jenner [769]. Jennifer [843, 1787]. Jenny [1260]. Jeremy [655, 2749, 2290, 2357, 2432, 2024, 2428]. Jerome [1434, 173]. Jeronimo [1681]. Jerry [155, 1910, 351]. Jerusalem [1804]. Jervis [1961]. Jerzy [120]. Jet [52, 1057]. Jett [1931]. Jevons [270, 269, 2620]. Jew [2505, 2823]. Jewish [681, 1873, 322, 1100]. Jews [2927, 2823]. Jill [2213, 548]. Jim [1290]. Jimenez [2265]. Jirí [2836]. Jo [1415, 2243, 2424]. Joachim [2566, 247, 1152, 1379, 178, 2823, 2489, 1319, 465]. Joan [2115, 731, 705, 2057, 2886, 133, 1368, 839, 350, 991, 1043]. Joanne [2635, 2368]. Joaquin [963]. Jodrell [1924]. Joel [1428, 2862]. Joellen [2780]. Johann [2904, 1468, 573]. Johanna [706, 2504, 2740]. Johannes [1518, 718, 1686, 2813, 974, 283, 2533, 2441]. John [2249, 2166, 857, 152, 1809, 2035, 2497, 34, 79, 1371, 1538, 1933, 632, 2284, 1583, 2309, 492, 2047, 2937, 51, 94, 2021, 2564, 1314, 2041, 796, 2443, 1959, 1918, 428, 1597, 1644, 2281, 2012, 1925, 2368, 1206, 302, 2704, 369, 1667, 756, 770, 1288, 2599, 2701, 1728, 1571, 1547, 858, 101, 2600, 1717, 837, 2104, 2803, 1674, 1336, 1106, 1305, 587, 1552, 638, 707, 2107, 2268, 2038, 2204, 2441, 1210, 1714, 312, 1006, 1751, 2133, 128, 55, 885, 1324, 2191, 2221, 2347, 2380, 685, 686, 2345, 2214, 211, 25, 2893, 788, 1913, 1993, 2377, 2654, 2444]. John [2547, 907, 2550, 543, 1374, 1902, 633, 1666, 2764, 1568, 2432, 2448, 412, 1515, 2334, 1151, 844, 2119, 314, 1033, 2800, 126, 244, 1977, 1831, 2255, 2141, 2760, 28, 2336, 588, 671, 876, 788, 420, 1683, 1606, 792, 41]. Johnny [2273]. Johns [1111, 2246, 2106]. Johnson [2458, 2038, 1996, 2021, 2783]. Joint [1083]. Jole [1675]. Jon [1362, 2162]. Jonathan [1148, 474, 2730, 2137, 382, 2783, 2506, 1174]. Jonathon [1737]. Jones [2216, 1322, 1956, 965, 2216, 550, 189, 85]. Joppien [1808]. Jordan [558, 2040, 1964]. Jordanova [167, 2023]. Jordanus [850]. Jorg [1659]. Jorland [677]. Jose [1154, 1782, 1613, 1743, 1037, 1232, 1584, 1932, 1627, 444, 2592, 2751, 2627, 443]. Josef [2836, 431, 1491, 2257, 1470, 2740]. Josep [2593, 2900]. Joseph [222, 47, 2477, 1403, 653, 2910, 2311, 952, 2774, 1636, 934, 732, 1155, 1240, 731, 825, 515, 1444, 2043, 2588, 1884, 2057, 1421, 162, 118, 1679, 2490, 2370, 2347, 721, 2589, 2103, 951, 1651, 1372, 2778, 2420, 783, 43, 2039, 588, 2876, 334, 384, 2602, 1287, 512, 1053]. Josip [2739]. Jou [2627]. Joule [51]. Journal [1103, 213, 10, 1641, 2747, 490, 2032, 150, 594]. Journalists [1820]. Journals [2661, 791, 1894]. Journey [2184, 1547, 2292]. Joyce [347]. Józef [2598]. JPL [1057]. Jr [2474, 2014, 1142]. Jr. [248, 1444, 792, 2239]. Juan [442, 235]. Juanelo [1154, 1932]. Judgment [2664]. Judiciary [2756]. Judith [2201, 693, 1172, 1426, 2080, 771, 542, 1110]. Judson [2411]. juillet [1592]. juin [184]. Jules [2571, 49]. Julian [2658, 1030]. Julien [2827]. July [465, 2155, 264, 2793, 671, 1641, 1626, 944, 1529]. June [735, 1804, 688, 922, 1878, 464, 265, 1346]. Jung [1666]. junge [1285, 241]. Jungius [1152]. Juni [985]. Jupiter [2405]. Jurgen [1186, 1026, 761, 133, 2757, 2654, 2498]. Juskevic [1983]. Just [2654, 1528, 1831]. Justin [1827]. Justus [2058, 2702, 1861, 2702]. Jutta [727, 36]. Juyces [2303].

Kaal [356]. Kabinettes [2740]. Kac [2673]. Kahn [2378]. Kaiser [2763]. Kaiser-Wilhelm-Instituts [2763]. Kaiserreichs [1008]. Kaiserzeit [1470]. Kambartel [356]. Kamen [2063]. Kan [824]. Kanel [1673]. Kangro [91]. Kanigel [2190]. Kant [2826, 2082, 1716, 475, 1718, 855, 5, 899, 2913, 1716]. kantiano [1845]. kantiano-popperiane [1845]. K’aokuhsue [377]. Kapitel [2489, 2946]. Kapitza [1914]. Karen [1953]. Karg [2555, 754]. Kargon [869]. Karin [271, 1566, 283, 1409]. Karl [2286, 489, 864, 2776, 1011, 1013, 2728, 1119, 1866, 1739, 576, 1533, 889]. Karl-Heinz [1011]. Karlsruhe [1415]. Karmi [238]. Karp [1956]. Kartsev [2532]. Karydis [2170]. Kassler [2453, 2453]. Katalog [1379]. Kate [350]. Katharine [1853]. Katherine [94, 2100, 2430, 517]. Kathleen [731, 2057, 621, 2130]. Katz [2017]. Kauffman [1751, 2832, 829, 1962, 92]. Kaufman [2780]. Kay [1051]. Kayas [234]. Kaye [1967]. Kealey [2275, 713]. Kearns [312]. Kechishih [377]. Keck [871]. Keele [1680]. Keeler [1692]. Keeney [1197]. Keeping [1769]. Kegel [1443]. Kegel-Brinkgreve [1443]. Keift [384]. Keil [1340, 1203, 1275]. Keilschrifttexte [1201]. Keir [2811]. Keith [1252, 2027, 2682]. Kekules [1153]. Keller [1415, 1267, 830]. Kelly [757]. Kelvin [50, 2918]. kemiska [219]. Kemp [2075, 786]. Kempski [1026]. Ken [419]. Kendall [104]. Kenly [2937]. Kennard [1103]. Kennedy [1344, 1344, 1867, 1439, 552, 872]. Kenner [2177]. Kenneth [1972, 288, 481, 2660, 1200, 2685, 2766, 2374, 245, 2882, 1562, 1310, 1843, 2049, 187, 1680, 1889]. Kenney [2261]. Kenny [2894]. Kent [569]. Kenzer [2605]. Kepler [1785, 1857, 1518, 1686, 2409, 573, 2824, 718, 185, 2522]. Kern [1247]. Kerr [1273, 1750]. Ketner [2374]. Kettering [1381]. Keuren [2161]. Kevin [1574]. Kevles [1805]. Key [240, 1078, 2235]. Keynes [48]. Kh [273]. Khan [1277]. Khayyam [1979, 1979]. khimii [285]. Kholodilin [657]. Khrushchev [872]. Kibre [1849, 1850]. Kidd [1521]. Killan [1493]. Killian [1812]. Killing [2851]. Kilmister [2459]. Kimball [393]. Kindly [53]. King [2737, 2131, 1344, 25, 1048, 966, 1641, 2734, 2305]. King-Hele [2737, 1048, 2305]. Kings [328]. Kingsland [1833]. Kingston [410]. Kinnon [657]. Kiple [1843]. Kirchheimer [1660]. Kirkbride [1769]. Kirsanov [2903]. Kirsh [2252]. Kirshenbaum [2253]. Kirsten [133, 1523, 2132]. Kirurgi [910]. Kitab [1439, 504, 1276]. Kitchener [2327]. Kitcher [887, 2672, 1834]. Kite [865]. Kitt [1357]. Kitzinger [2693]. Klaaren [1985]. Klaas [1514]. klassicheskoi [285]. klassischen [424]. Klaus [503, 2808, 612]. Klausner [2194]. Kleene [1959]. Klein [1824, 2239, 1100, 1873, 1204]. Klein-Franke [1204]. Kleine [358]. Kleinert [2533]. Kline [2243]. Klingemann [2946]. klinischen [706]. Kloeppel [1357]. Kloesel [1864, 2374, 1951]. Kloppenburg [2724]. Klosterman [2138]. Klostertypika [1593]. Klotz [2077]. Kluttz [2274]. Knatz [1789]. Kneller [159]. Knight [164, 2743]. Knobloch [45, 553]. Knoefel [723]. Knorr [1409, 2729]. Knorr-Cetina [1409]. Know [1461]. Knowing [2420]. Knowledge [1950, 2327, 270, 2883, 601, 148, 1241, 2665, 616, 1907, 1400, 2268, 2164, 1795, 2419, 1912, 85, 211, 2560, 1561, 2865, 791, 1880, 1015, 212, 23, 352, 1012, 2089, 2494, 157, 2420, 2477, 1302, 890, 306]. Knowles [658, 1500]. Known [1980, 1044]. Knox [2126]. Koblitz [1691]. Koch [2837, 2837]. Kochan [2352]. Kocher [1201]. Kodo [1599]. Koelsch [2755]. Koenigsberger [380]. Koffka [1328]. kognitiven [908, 1576]. Kohlenforschung [2763]. Kohler [905]. Kohlstedt [2478]. Kohn [2748, 2488]. Kollesch [36]. Kolmogorov [281]. Kolodziej [2606]. Kombinatorik [45]. Komkov [891]. Kommentar [2586]. Kommunikation [247]. Kompendien [2819]. Kong [1205]. Kongress [2031]. Kongresses [1760]. Konig [1011]. Konopka [428]. Konrad [1285, 241, 1239]. Konstitutionserforschung [2498]. Konstitutionsforschung [823]. Kontext [503]. Kontinente [427]. Kontinuum [178]. Konvitz [2257]. Kopelevich [273]. Kopernik [120]. Koppes [1057]. Korber [988]. Korean [2347]. korespondencji [120]. koristi [2739]. Korrespondenz [247]. Koslowski [2355]. Kosmologie [1485]. Kosmos [785]. Kossbiel [2854]. Kotz [2607]. Kovalevskaia [1691]. Koyre [677]. Krafft [643, 75, 1019, 856, 798]. Kraft [1519, 1523]. Kragh [2548]. Krakatau [1487]. Kramer [2509]. Kramers [2924]. Kramish [2379]. Kranzberg [842, 1671]. Kratz [1153, 287]. Kraus [2052, 2392]. Kraye [2213]. Kreatur [785]. Krebs [134, 931, 1080]. Kremlin [1914]. Kren [1817]. Kretzman [2894]. Kretzmann [851]. Krieck [154]. Krieg [2504]. Krieger [2813]. Krige [2760, 1006]. Krimsky [1058]. Krin [2637]. Krisciunas [1574]. Kristeller [373]. kritische [1152]. Kritischer [1909]. Krohn [352]. Kruger [2652]. Kruk [504]. Kryptologie [2700]. Krzhizhanovsky [2532]. Ku [571]. Ku-tai [571]. Kudlien [497]. Kuehn [469]. Kuhn [409, 751, 479, 2251]. Kuhn-Schnyder [479]. Kuhner [606]. Kula [2805]. Kultur [2210]. Kultur- [2210]. kultureller [292]. Kulturgeschichte [1485, 1776, 761]. kung [571]. kung-yeh [571]. Kunitzsch [2098, 2822, 1595, 2393, 1677]. Kunst [2817]. Kunsttriebe [1026]. Kunzl [1470, 1470]. kuo [571]. Kupsch [487]. Kursunoglu [2460]. Kurt [1959, 1184, 626, 385, 283, 715, 1942, 1328]. Kurt-R [1184]. Kurtz [2195]. Kuru [771]. Kurylo [736]. Kutschmann [1519]. Kutzbach [168]. Kuznetsov [2344]. Kuznick [2294].

L [981, 854, 2075, 2455, 474, 2574, 1199, 2631, 1231, 1273, 2915, 1200, 1166, 1508, 31, 366, 1946, 2736, 683, 204, 2451, 88, 2555, 1290, 1984, 1321, 469, 2881, 2353, 78, 1973, 1335, 166, 2751, 692, 2496, 961, 2283, 278, 1983, 1967, 2656, 698, 485, 491, 2136, 574, 215, 2140, 167, 1820, 1437, 2780, 1734, 2306, 2441, 513, 1501, 1988, 2856, 1879, 2551, 1564, 229, 2416, 2325, 754, 1311, 694, 232, 1739, 1554, 855, 570, 900, 1163, 2508, 13, 293, 1579, 873, 2596, 1120, 43, 442, 1035, 949, 2093, 891, 2549, 2760, 286, 1961, 647, 1828, 97, 1529]. L [239, 1440, 485, 1037]. l-Bayan [1037]. l-hay [1440]. Labor [2521, 2690]. laboratoire [1348]. Laboratories [1455]. laboratorio [2601]. Laboratory [1938, 2009, 2294, 2471, 2427, 2018, 1526, 2471, 865, 2806, 1057, 2354, 156]. laboris [2661]. Labour [689]. Labours [2527]. l’Académie [446]. L’accueil [585]. Lacy [110, 1661]. Ladd [1681]. Laennec [1207]. Laidi [2450]. Laine [2374]. Laitko [2720]. Lajos [1040]. Lakatos [152]. Lake [788, 2874]. Lalla [917]. Lamarck [1580, 906, 906, 1049]. Lamb [1445]. Lambert [2597, 507]. Lambright [1996]. Lambros [776]. Lammot [1504]. Lamp [1237, 2040, 1740]. Lancastrian [2223]. Lancets [783]. lancette [1496]. Land [1265]. Landau [1257, 868]. Landes [1377]. landessprachigen [2819]. Landmarks [2337, 1325]. Lands [2185, 95, 2435]. Lane [1480]. Lanfranco [1554]. Lang [288]. Langdon [2042, 2121, 35, 1831, 2214]. Langenbruch [684]. Langevin [2761]. Langham [1250]. Langins [2350]. Langley [2656, 2427]. Language [2564, 435, 2009, 1034, 2334, 576, 492]. Languages [1350, 238, 873, 2023]. langue [280]. Lanning [1902]. Lapérouse [2403]. Laplace [922, 1047]. Laporte [2408]. Laqueur [2487]. L’architecture [1175]. Lardner [2271]. lardomshistoria [1284]. Large [2099]. Larry [1185, 1560, 2014]. Lars [824, 984]. Larsson [910]. L’arte [724]. LaSalle [866]. Last [2649, 1665]. l’astrolabe [1036]. L’Astronomie [496]. Laszlo [473]. Late [123, 2527, 852, 1444, 256, 2788, 1003, 2899, 13, 2686, 1317, 2731]. lateinischen [2897]. lateinischer [117, 1433]. Later [756, 2894]. Latimer [1671]. Latin [2447, 1849, 1438, 373, 600, 242, 2126, 1176, 41, 2066]. Latinus [1849, 504, 2126]. Latour [2354, 156]. Latrobe [1933, 2221, 636, 1809]. Laudan [1363, 1560, 2361]. Laughing [2628]. Laughlin [1714]. Launch [797]. Launching [2653, 2760]. Laura [1142]. Laureate [131]. Laureates [2193]. Laurel [2328]. Laurence [2363, 854, 2196, 109]. Laurent [2699, 922, 2191]. Lauretis [621]. Laurie [1414]. L’avenement [776]. l’avenir [1410]. L’aventure [2809]. Lavington [569]. Lavoisier [468, 2699, 922, 442, 1080, 1131, 1941, 1881]. Law [2289, 1520, 2543, 1246, 1244, 2165]. Lawn [374]. Lawrence [1886, 2558, 515, 341, 1510, 2628, 1661, 284, 1914, 2910, 1066, 2887]. Laws [1144, 142]. Lay [1280]. Layman [2949]. Layton [352]. Lazarus [1040]. Laziness [770]. Lazzaro [1179]. Lead [1434, 1770, 1434]. Leadership [1634, 2424]. Leading [1166, 1429]. Leaf [295]. Leahey [2326, 562]. Lear [2730]. Learned [1010, 1280]. Learning [2922, 2365, 1740, 1239, 2083, 1910, 781, 1889]. l’eau [505]. Leavitt [2201, 1426, 1110]. Leben [2838, 798]. Lebendigen [475]. L’économiste [2829]. LeConte [1053]. Lectiones [1783]. Lectures [731, 1146, 544, 2169, 1703, 1733, 359, 1783, 1289]. Lee [1933, 2221, 1325, 1913]. Leendert [117]. Leenhardt [1301]. Leeuwenhoek [647, 647]. Lefevre [1572]. Legacies [2528]. Legacy [2658, 1770, 450, 2737, 2144]. Legend [466]. légende [585]. Legendre [1745]. Legitimacy [1835]. Legitimation [1243]. legitimiste [1788]. Lehre [1906, 178]. Leibniz [46, 1285, 2738, 45, 241, 108, 668, 328, 2801, 509, 2130, 46].

Leibniz-Gesellschaft [46, 108]. Leibnizens [553]. Leicestershire [2634, 2634]. Leinwoll [243]. Leisibach [1475]. Leisure [276]. Leith [1644]. Leith-Ross [1644]. Lejeune [1184]. Lelievre [958]. Lemaine [1819]. Lenhardt [1340]. Lenin [104]. Lennox [2585]. Lens [2765]. Lenz [864]. Leo [2138, 2340, 2408]. Leon [766, 963]. Leona [135]. Leonard [1759, 186, 2768, 1923, 2337]. Leonardo [468, 1513, 947, 786, 2820, 1680]. Leonhard [1983]. Leonie [929]. Leopold [839]. Leopoldo [1179]. Lepenies [1396, 984, 2812, 908, 292]. Leplin [1722]. Lerner [1319]. Lesch [1667]. Lesley [1858]. Leslie [1032, 710, 962, 1381]. L’esprit [1488]. Lesser [2033]. Lester [966, 2195, 384, 1641, 1053]. Letter [1855]. Lettere [506]. Letters [387, 1048, 1073, 2857, 393, 1610, 858, 771, 2794, 2058, 1993, 561, 2408, 2877, 1554, 647, 490, 1614]. Lettre [651]. Leufsta [2401]. Levere [1089, 758]. Levi [1851, 2419, 2419]. Leviathan [1986]. Levidow [689, 2174, 2090]. Levin [245]. Levine [2493, 47, 2920, 2744]. Levins [2262]. Leviton [1163]. Levsin [891]. Lévy [2433]. Lewin [2438]. Lewis [1647, 214, 2853, 1826, 2823, 274]. Lewontin [837, 2262]. Lexikon [38]. Leyde [845]. l’histoire [585, 2170]. l’Huillier [280]. Li [378, 571]. Li-sheng [571]. Libavius [1938]. Libby [135]. Liber [2586]. Liberal [1437, 2666, 691, 1765]. Liberalism [491]. Liberation [616]. Librarians [1669]. Libraries [2241]. Library [1196, 1559, 668, 681, 2009, 101, 2337, 2161, 822, 2241, 1818, 2781, 1142, 258, 2414, 1748, 1073, 1111, 2401, 109, 2162, 2413]. libri [375, 849]. Libro [1681]. Lichte [174]. Lichter [987]. Lichtspekulation [503]. Lick [2559]. Lidz [2194]. Lie [1144]. Liebenau [2506]. Liebig [2058, 1861, 2702]. Liebigs [1153]. Liedman [2287]. Lieferung [38]. Life [2767, 946, 169, 736, 2276, 2826, 2354, 899, 2706, 395, 2082, 1614, 516, 767, 2527, 1267, 731, 2777, 1222, 2030, 2680, 278, 131, 156, 2143, 420, 2059, 937, 1911, 1995, 1138, 1528, 1734, 1010, 1881, 548, 1016, 1988, 2528, 648, 2290, 2036, 788, 925, 869, 552, 363, 2823, 449, 2746, 1337, 2712, 1352, 866, 2596, 2407, 2061, 244, 2144, 1581, 2633, 2523, 1445, 1525, 2135, 1172, 1161, 1986, 2189, 2928, 1346]. Lifetime [1752]. Lifting [1674]. Lifton [2851]. Light [260, 1855, 399, 2475, 2457, 187, 60, 1482, 2011]. Lightman [1544, 2494]. like [2735, 2735]. Lilienfeld [136, 634]. Lillian [1286, 2741, 1414]. Lily [1644]. Limited [232]. Limits [2042, 2121, 2494, 1723, 350, 602]. Limoges [29]. Lin [378]. Linda [2848, 1693, 1438, 1564, 1191, 2941]. Lindberg [2093, 1177, 1509]. Linden [2373]. Lindgren [2095, 1739]. Lindqvist [1928, 913]. Lindroth [748, 1284]. Lindsey [80]. Linear [932, 2357]. Lingua [2700]. Linguistic [1492]. Links [1106]. Linnaean [719]. Linnaeus [528, 1353, 719]. Linne [984]. Linnean [2555]. L’institution [446]. l’intelligence [1819]. Lintner [1809]. Linus [2850]. Lipscombe [350]. Lipset [450]. Lisa [991]. Lise [103]. List [629, 2662]. Lister [1494, 1494]. Lists [373]. Lit [1642]. Literal [2334]. Literary [2635, 2293, 957, 2430, 1480, 1569]. Literati [1859]. Literatur [1640]. Literature [834, 1084, 1231, 1166, 2307, 2949, 2907, 151, 2812, 2195, 2754, 402, 2783, 412, 898, 1251, 1101, 1520, 124, 2023, 2493, 949]. Littera [1681]. litteraire [853, 714]. Little [2378, 188, 2238, 290]. Littleton [2178]. Littlewood [1159, 1159]. Littmann [1752]. Liu [2450]. Livanova [868]. Livelong [2299]. Lives [1254, 1691, 1209, 2328, 2193]. Living [1665, 876, 2212, 1772]. Livingston [2180]. Livingstone [2529, 2921]. Livre [1594, 2043, 1676, 376]. Lloyd [1111, 1775, 1337, 498]. Lobato [2592]. Local [21]. Loch [2351]. Lochak [2604]. Lock [2296]. Locke [1034, 1606, 1982]. Loeb [872]. Loeck [2245]. L’oeuvre [1979]. Logic [474, 2530, 508, 476, 2899, 1702, 381, 26, 235]. logica [2122]. Logicae [2171]. Logical [2196, 1570]. Logik [1094]. logika [281]. logique [381]. logique/Logic [381]. Lois [169]. Lomonosov [2523]. London [2455, 1118, 2555, 1159, 593, 2110]. Long [1921, 20, 1581, 1357, 2045, 375]. Long-Distance [1921]. Longevity [2114]. Longeway [1977]. Longitude [348]. Longley [1934]. Lonie [777, 1854]. Look [532]. Looking [1051]. Looks [1567]. Lopez [1765, 1782, 1232, 1584, 444, 2751, 443, 2627]. Loptson [1516]. Lord [1138, 1268, 50]. L’ordre [651, 315]. Loren [1396, 187, 1234]. Lorenz [2838, 2652, 479]. Lorenzen [2776]. Loretta [2780]. L’orientation [1117]. Lorin [1044]. Loris [1847]. Lorna [350]. Lorraine [2652, 2628]. Losee [2345]. Losing [2709]. Lost [93, 2803, 744, 1581]. Lothar [1789, 1306]. L’ottica [1749]. Louis [2100, 2534, 1857, 692, 368, 624, 2604, 1929, 651, 2604]. Louis-Paul [1857]. Louise [2298]. Louisiana [2444]. Louvain [1783]. lovanienses [1783]. Love [2149]. Loved [2764]. Lovell [1924]. Loves [2873]. Low [1992, 2355, 475]. Lowe [2142]. Lowell [1289, 2082]. Löwendahl [2401]. Lowenthal [2890]. Lowood [341, 1815, 2719, 341]. Lowrance [1729]. Loyd [2138, 394, 248]. Lozoya [2914]. Lu [2588, 2347, 783]. Luc [2725]. Lucchetta [372]. luces [2914]. Luciani [1543]. Lucile [415]. Lucinda [2638]. lucis [503]. Luck [1976]. Luckin [2222]. Lucrèce [500]. Lucretius [1430]. Lucy [1050]. Ludmerer [1889]. Ludmilla [2023]. Ludo [227]. Ludwig [431, 2777, 1288, 1011, 862, 1655, 987, 1185]. Ludwik [1306]. Luetin [1390]. Luigi [2801, 1188]. Luis [2531, 1781, 2751, 2627, 1037, 1846, 235]. Luisa [608, 42, 1300]. Luke [501]. Lukes [1477]. lumière [1162, 184]. lumières [71, 920]. Lunar [1413, 248, 2018]. Lunari [2900]. Lunatics [2634]. Lundgren [219]. Lunds [910]. l’Université [496]. Lupovici [1207]. Luqa [2391]. Luria [246]. Lusis [1947]. Lussac [70, 625]. Lüst [2808]. Luther [1995]. Luyendijk [1443]. Luyendijk-Elshout [1443]. Luz [1782, 2751, 444]. Lwoff [429]. Lydie [1207]. Lyding [2368]. Lyell [49, 196]. Lyle [2038]. Lynch [2471]. Lyndsay [270, 616]. Lynn [728, 1864, 2030, 2474]. Lyon [857, 1547]. Lysenkoism [2363].

M [2570, 558, 834, 47, 569, 2406, 2077, 132, 1780, 831, 1605, 221, 2929, 1897, 2194, 1765, 1147, 2286, 2041, 2593, 2096, 583, 1959, 2688, 777, 1345, 1782, 1854, 2555, 2884, 1103, 2186, 1918, 702, 859, 161, 1049, 1518, 575, 1543, 1784, 2598, 2109, 1985, 2180, 1613, 982, 1909, 2751, 2590, 1272, 1713, 649, 580, 1970, 1983, 2558, 2656, 832, 248, 582, 1241, 1571, 1841, 1874, 122, 305, 968, 1322, 2843, 2841, 1820, 2022, 2088, 847, 1712, 1719, 2708, 701, 1211, 1975, 615, 977, 1232, 2887, 55, 542, 179, 657, 818, 2058, 1513, 685, 2034, 2919, 1119]. M [1179, 279, 137, 2434, 2858, 20, 1917, 2924, 754, 2102, 1263, 1678, 1584, 363, 634, 1350, 1443, 2396, 1440, 1507, 2122, 890, 791, 1414, 164, 2323, 1145, 1101, 2572, 1570, 314, 2666, 792, 630, 828, 1740, 1899, 2770, 2162, 1994, 2212, 619, 1315, 1889, 695, 1989, 490]. M. [1765]. M.D [1687]. M.D. [331]. M.N.B. [1855]. Mabberley [2405]. Macchina [488]. macchine [1188, 1009]. Maccio [2122]. MacDonald [2869]. MacDougall [1329]. Macey [278]. Macfarlane [321, 1527]. MacGregor [1644]. Mach [1055, 2678, 2254, 2254]. MacHale [1911]. Machaut [180]. Machina [1368]. Machine [2410, 2242, 2202, 2119]. Machinery [430]. Machines [1746, 1746, 1932, 2244]. Machs [247]. MacKenzie [1015, 2041]. MacKinnon [1147]. Maclear [1610]. MacLeod [2947, 791]. Mad [894, 2633]. Mad-Doctors [894]. Madame [2282]. Madaule [1049]. Madeleine [1049, 180]. Madhouses [894]. Madmen [894]. Madness [1631, 1842, 2856, 2034, 102, 2265, 2742, 1900]. Madrid [1613, 698, 442]. Maeyama [1776]. Maghreb [975]. magia [2801, 108]. Magic [2848, 2109, 1841, 789, 1976, 1351, 498]. Magical [279]. Magner [169]. magnete [806]. Magnetic [2472]. Magnetism [408]. Magnetospheric [1358]. Magnificent [2105]. Magnus [1780, 2695, 307]. mago [42]. Mahmoud [1678]. Mai [1990]. Maienschein [2682, 2169]. Maier [852]. Maillart [231]. Main [2326, 562]. Maine [838]. Mainz [1990]. Maize [2772]. Majesty [1322]. Major [1637, 2015, 2347]. Make [1481, 1539]. Make-Believe [1481]. Makepeace [1642]. Maker [2284]. Makers [347, 52, 432, 2157, 1895, 2510, 1952, 1953]. Making [246, 321, 614, 2487, 1377, 2925, 2410, 1923, 2190, 1920, 1670, 1815, 2419, 417, 2505, 2508, 905, 430, 844, 2806, 839, 255, 1351, 2351, 2509, 1171, 2606, 2632, 2670]. Makromolekule [556]. Malabaricus [2576]. malade [1506]. maladie [1506, 712]. maladies [1335, 1431]. Malaria [2619]. Malcolm [2353, 20, 1174]. malicia [1846]. Malina [2080]. Malinowski [459, 2331]. Mallikarjuna [917]. Malone [2226]. Maloney [1589]. Malthusian [2233]. Mamluks [998]. Mammalogy [2811]. Mammals [1581]. Man [1353, 1360, 2263, 2240, 227, 2276, 730, 380, 1567, 1185, 96, 12, 1173, 897, 2054, 1527, 1461, 2341, 2333, 2142, 2764, 844, 126, 786, 2917, 1106, 1680, 1252, 2695, 220, 295, 674]. Manacles [2742]. Management [1665, 2688, 2690, 1996]. Managers [1032]. Managing [31]. Manazir [1276]. Manchester [2035, 1642]. Mandelker [1027]. Mandrino [2338]. Mandrou [121]. Manetti [2122]. Manfred [762, 2763, 39, 1433]. Manhattan [2759]. Mankind [2631, 2161, 964, 2206]. Mann [1990]. Manned [248]. Mannigfaltigkeitsbegriffs [1021]. Manning [2833, 1528]. Manuel [442, 1248]. Manufactures [914]. Manufacturing [1549, 1876]. Manuli [112]. Manuscript [387, 819, 240, 1786, 1958, 2339, 2161, 1093, 1466, 680, 6]. Manuscripts [1196, 681, 918, 109, 1119, 1093, 2836, 782, 2220, 2014, 2241]. Manuscrito [2592]. manuscrits [1705, 1207]. Many [1321, 2889]. Map [964, 512]. Mapmaker [1040]. Mapmakers [815]. Mapping [1321, 904, 95, 1658, 1814]. Maranjian [2780]. Marburg [345, 82]. Marc [1092, 731, 2057, 1012, 1024, 2425]. Marcel [362, 1014, 1244, 66]. Marcello [2219]. March [1000, 1709, 2474, 537, 2839]. Marchand [576]. Marcia [622]. Marco [1782, 2053]. Marcos [1232]. Marcou [49]. Marcucci [1845]. Marcus [1970, 2420, 1835]. Margalit [2480]. Margaret [834, 629, 428, 1116, 1886, 2698, 2479, 1671, 2366, 1490, 494, 2019, 2191, 2478, 175, 350, 1171]. Margin [276]. marginadas [1781]. Margins [212]. Margo [2721]. Margolis [1403, 2664, 2105]. Margot [2555]. Marguerite [2368, 1045]. Margulis [2030]. Maria [1847, 1782, 608, 1188, 1749, 2067, 1300, 444, 444, 2751, 2627, 443]. Mariá-Luz [444]. Maria-Rita [1847]. Marianne [2586, 2278]. Mariano [1379]. Marie [1343, 2555, 1496]. Marie-Christine [1343]. Marijan [2826]. Marilyn [2140]. Marin [1394, 578, 2129]. Marine [2266, 1370, 756]. Mario [1608, 112, 2595, 1868, 1356, 2280, 90]. Marion [439]. Mariscotti [1868]. Maritime [2893]. Marius [1855]. Marjorie [2176, 2372, 1832, 537]. Mark [2567, 2799, 2649, 1249, 1752, 1638, 556, 2885, 2673, 2015, 105, 2578, 2758, 841, 1031, 705, 2292, 932]. Markell [1767]. Markers [11]. Marketing [2606]. Marketplace [1955]. Markie [2084]. Markova [2549]. Markus [174, 1178]. Marland [2643]. Marquis [922]. Marrone [1907, 1341]. Mars [395, 1693]. Marsha [494]. Marshall [135, 37, 600, 964, 2448, 436, 1590, 2620]. Marshes [1421]. Marta [2051]. Martens [2854]. Martha [774, 277, 2646, 512, 1555, 691, 2145]. Martin [650, 786, 2284, 2605, 1200, 1522, 1018, 2063, 924, 2486, 1477, 2780, 2149, 1349, 2720, 426, 1576, 2302, 1768, 2261, 583, 516, 2867, 2562, 898]. Martine [2170]. Marton [485, 485]. Marty [1200]. Martyr [1992]. Maruani [2926]. Marvellous [786]. Marvelous [6]. Marvels [1778, 981]. Marvin [953, 2686]. Marx [272, 1467, 1119, 1005, 414, 1119, 617]. Marxism [2799, 401]. Mary [2166, 474, 1415, 656, 1166, 2453, 2573, 2916, 1540, 2185, 1344, 2694, 2343, 1490, 2652, 545, 95, 1669, 109, 1601, 2423, 1913, 2243, 158, 2424, 1968, 2265, 699]. Mary-Jo [2243]. Masani [2293]. Masayoshi [239]. Maschine [2245, 1017]. Masotto [1253]. Mass [1549]. Massachusetts [909, 909]. Massey [2708]. Massimo [2801, 1436]. Masson [2857, 1422]. Mastai [1188]. Masters [2034]. Mastery [2073]. Masuda [1599]. Matalon [1819]. matematica [693, 372, 1845]. Matematicheskaia [281]. matematico [2595, 1188]. Matematika [281]. Matemátiques [1192]. Materia [1205, 1678]. Material [929, 2331, 1376]. Materialien [718]. Materialism [1717, 5]. Materialismus [2245]. Materialistische [603]. Materials [387, 1925, 55, 2781, 2358]. matériaux [301]. Math [2815]. Mathemata [2095]. mathematicae [1744]. Mathematical [1746, 1302, 583, 1159, 1958, 1704, 1136, 1119, 2734, 356, 2722, 1157, 1989, 347, 2496, 825]. Mathematician [1958, 1474, 1178, 1980]. Mathematicians [2409]. mathematicien [1788]. mathematico [1253]. Mathematics [281, 1946, 2298, 2692, 1636, 825, 420, 918, 162, 1850, 846, 1238, 1283, 2673, 142, 622, 2004, 2879, 1873, 2432, 900, 1313, 1939, 2415, 2567, 1411, 2672, 2215, 324, 357, 305, 827, 2674, 436, 2891, 1571, 694, 2448]. Mathematik [2807, 1672]. Mathematiker [499, 1278]. Mathematique [1175]. mathematiques [975, 958, 2248, 483, 305, 2433, 2814]. Mathematische [2906]. mathematischen [45]. mathematischer [1186]. Mathematization [2620]. Mathematized [1235]. Matrix [995]. Mats [1762]. Matter [57, 138, 192, 249, 253, 1, 58, 139, 193, 251, 325, 396, 453, 518, 522, 254, 326, 397, 454, 520, 525, 589, 661, 739, 799, 803, 523, 526, 590, 662, 740, 801, 874, 935, 992, 1059, 1063, 804, 875, 936, 993, 1061, 1122, 1212, 1294, 1382, 1386, 1064, 1123, 1213, 1295, 1384, 1451, 1531, 1615, 1695, 1699, 1387, 1452, 1532, 1616, 1697, 1792, 1870, 1936, 1998, 2002, 1700, 1793, 1871, 1937, 2000, 2064, 2150, 2229, 2313, 2317, 2003, 2065, 2151, 2230, 2315, 2382, 2461, 2535, 2608, 2612, 2318]. Matter [2383, 2462, 2536, 2610, 2614, 2710, 2784, 2871, 2952, 2615, 2616, 2711, 2785, 2872, 1949, 742, 806, 153, 1717, 2628, 955, 1145, 1449, 1257, 1825, 648]. Matthew [428, 1355, 170, 2934, 2541]. Maulitz [2644]. Maupertuis [1983, 651]. Maurer [885]. Maurice [508, 1935, 1301, 612, 1274, 449, 2460, 301, 2295, 828, 1748, 1393]. Maurizio [44, 1856]. Mauskopf [495, 211]. Max [2726, 1949, 1864, 2374, 131, 928, 2333, 916, 2950, 1663, 341]. Max-Planck-Gesellschaft [2950]. Maxima [1977]. Maxime [848]. Maximo [860]. Maxine [2145, 430]. Maxwell [2145, 2599, 1975, 733, 2665, 1522, 2831, 1412]. Maxwellians [2703]. May [811, 66, 2155, 1878, 535, 1627, 949, 1001, 909]. Mayer [577]. Maynard [2079, 963]. Mayr [914, 2864, 612]. Mazene [1308]. Mazuzan [1694]. Mazzeo [158]. Mazzolini [2402, 560]. McCann [2368, 165]. McClellan [1823]. McClelland [477]. McClintock [1267]. McCloskey [2564]. McConnell [1024, 548, 549]. McCosh [1524]. McCraw [2668]. McCray [2638]. McCusker [1654]. McCutcheon [350]. McDonald [962, 2168, 2175]. McEvoy [978]. McGehee [2631]. McGovern [2034]. McGrath [1991]. McGucken [1741]. McGuinness [356, 2777, 2678]. McGuire [1349, 1666, 1715]. McHugh [432]. McIntosh [1325, 1761, 2397]. McIntyre [93]. McKeen [490]. McLachlan [1260]. McLaughlin [888]. McLeod [2455]. McMackin [691]. McMahon [1920]. McManus [1714]. McMath [2558]. McMillan [932]. McMullin [153, 2557, 2094]. McMurry [1191]. McNeill [1112]. McNiven [445]. McVaugh [495, 1438, 1325, 849]. Me [1202]. Mead [1490, 1171]. Meadows [679, 1740]. Mean [2722]. Meaning [1887, 1967, 2698, 1862]. Measurement [1819, 2434, 2891]. Measures [21, 1919, 2469, 2805]. Measuring [2348, 1757]. Meat [2509]. mécaniciens [371]. Mechanic [2177]. Mechanical [1674, 416, 308, 1500, 1934]. Mechanics [1710, 1222, 546, 2675]. Mechanik [424, 1655]. Mechanisering [1514]. mechanistischen [1285, 1469]. Mechanization [2410, 2204, 2772]. Medal [2071, 2388, 598]. Medals [1802]. médecin [376, 1506]. medecine [1496]. medecines [1676]. medecins [2266]. Media [622, 688]. Medica [1956, 1678, 849, 1846, 1205]. Medical [650, 2439, 1111, 1427, 656, 2110, 31, 1631, 1854, 1974, 1121, 1556, 2112, 2632, 2930, 2172, 2268, 565, 2413, 721, 1641, 2855, 781, 137, 1956, 1902, 2851, 905, 1596, 1007, 2506, 1889, 2818, 119, 1766, 966, 2172, 2506]. medicamento [2627]. médicas [2751]. medicin [1594]. medicina [1781, 1682, 2122, 1332, 1442]. Medicinae [1882]. Medicine [2113, 1583, 455, 1111, 1954, 2888, 896, 2910, 2013, 1779, 1853, 1556, 266, 2647, 1667, 1690, 2632, 709, 1269, 707, 565, 2365, 2887, 2014, 205, 1641, 2855, 1900, 2300, 2850, 175, 1901, 299, 2547, 1956, 374, 633, 1270, 402, 2390, 1596, 314, 400, 2278, 2651, 1335, 2635, 2047, 2631, 1200, 1043, 2949, 2892, 1089, 2488, 101, 2885, 1767, 1850, 1606, 39, 1028, 1464, 2450, 1230, 713, 297, 2278, 2296, 2035, 1665, 1200, 2736, 2643, 750, 2575, 2010, 1223, 1007, 967, 2197, 909]. medicinsk [910]. Medico [2081, 712]. Medico-Botanical [2081]. medico-philosophique [712]. Medicus [632, 713]. Medieval [2276, 779, 1115, 1775, 852, 2045, 756, 2521, 2816, 782, 2545, 373, 2894, 37, 1850, 2277, 600, 2899, 1848, 1339, 306, 235, 436, 851, 1176, 1509, 8, 1534, 2390, 639, 1817, 640, 2124, 1956, 1035, 713]. Medina [2592]. Medioevo [1436]. Mediterranean [2766, 2893, 2693, 2545]. Medium [897]. Medizin [1201, 547, 706, 1029, 1204, 19, 1906, 2808, 1029, 2630, 1203, 441, 2504]. Mediziner [134]. Medizinisch [1475]. Medizinisch-Chirurgische [1475]. Medizinische [1470]. Medizinischen [1475, 36]. Medusen [2502]. Meet [707]. Meeting [2070, 64, 1784, 1973, 1133, 881, 1457, 669, 2791, 332, 2387, 1801]. Megalithic [177]. Megill [2564]. Mehrtens [827, 353, 282]. Meikle [232]. Meinel [1152, 2210, 82]. Mel [2291]. Melancholia [2264]. Melanesian [1301]. Melba [1894, 1963]. Melcher [1487]. Melhado [831]. Mellon [109]. Melnick [349]. Melody [2636]. Meltzer [2428]. Melvin [1540, 842, 1671, 2779]. Memoir [2063, 1490, 1206, 1812, 2527, 922]. mémoires [223]. Memoirs [132, 1418, 1935]. Memories [1429]. Memory [1855, 2178, 179]. Men [1169, 2435, 1895, 1264, 2805]. Menard [2148]. Mendel [955, 2919]. Mendelsohn [1629]. Menschdom [2825]. Menschen [1340]. Menschheit [1485]. Menso [2095, 916]. Mensuration [1035]. Mental [1289, 1254, 1819, 734, 2931, 1271]. Mentalities [1551]. Mentasi [724]. Mentor [59]. Meo [1119]. Merchant [338]. Merchants [2077]. Mercury [1483]. Merit [1819]. Meritocratic [564]. Merkel [2902]. Merle [816]. Merriman [699]. Merritt [1930]. Merry [2192]. Mersenne [1394, 578, 2129]. Mersits [2760]. Merton [1096, 2108, 2731, 941, 1735, 2544, 2540]. Merzbach [1558]. Mesmerism [400]. Messeri [2053]. Messianism [2907]. Messrs [922]. Messtechnik [761]. mestiere [1009]. mesure [89]. metafisica [2597]. Metal [635]. metallica [606]. Metallurgy [606]. Metals [2206, 962]. Metalwork [1931]. Metamorphose [1519]. Metamorphosis [1934]. Metaphor [2256, 2233, 1994, 2863]. Metaphoric [1540]. Metaphysical [2748, 62]. Metaphysician [384]. Metaphysics [1145]. Metaphysische [2913]. Metastudies [2671]. Meteor [2360]. Meteores [2050]. Meteorites [2224, 2255]. Meteorological [168]. Meteorologie [641, 117]. Meteors [182]. Method [1249, 2016, 1215, 1716, 2481, 2802, 508, 217, 1588, 1726]. methode [1594, 2050, 1090, 2476, 2691, 376, 446, 2050]. Methoden [2031]. Methodenmodelle [1186]. Methodology [1663, 670, 1716, 2164, 2324, 609, 1718, 1137, 2943]. Methods [1746, 1254, 2665, 2332, 314, 635]. Métiers [2349]. metodo [2160, 644]. Metropolis [1312]. Metropolitan [1371]. Mettler [108]. Mettrie [2827]. Metzger [2476]. Mexico [1860, 670, 2914, 572]. Mey [1012]. Meyenn [451, 1866]. Meyer [1662, 556]. Michael [1353, 1408, 2771, 246, 2437, 819, 495, 2175, 204, 2555, 2082, 2775, 424, 1303, 2950, 1644, 2704, 1842, 389, 1970, 369, 1098, 821, 1438, 2215, 357, 2343, 2471, 1269, 2226, 2652, 1042, 2935, 1409, 1988, 1908, 775, 157, 2899, 2272, 618, 652, 754, 1109, 954, 2420, 646, 849, 209, 923, 2770, 901, 490, 1945, 2287, 2693, 13]. Michaels [310]. Michal [1920]. Michaux [2376]. Michel [1441, 500, 414, 1743, 1319, 2170]. Michel-Pierre [1319]. Michela [1790]. Michele [1163]. Micheli [604]. Michelle [624]. Micro [976]. Micro-Cosmology [976]. Microbial [2030]. Microcomputers [2515]. Microcosmos [2030]. Microfiche [2080, 181]. Microfilm [2475, 1142]. Micronesia [270]. Microscope [613]. Microtechnique [346]. Microwave [1586, 1365]. Microwaves [2701]. Mid [2455]. Mid-Victorian [2455]. Middle [1307, 1849, 115, 2066, 877, 1438, 1552, 1437, 2213, 37, 69, 2788, 600, 2310, 1510, 2448, 1140, 1590, 236, 2447, 1637]. Middleton [210, 658, 1500]. Midgley [1968]. Midwives [2115]. Miedzy [2816]. Migdal [2173]. Mijnhardt [2825]. Mikhail [2523]. Mikolaj [120]. Mikulinskii [74]. Mikulinsky [619]. Mild [1370]. Mile [1290]. Milestones [2662]. milieu [237]. militaires [1369]. Militarization [2139]. Military [2120, 2589, 2934, 2202, 1930]. Milk [2509]. Mill [2648]. Millburn [2284, 2284, 25]. Millenarian [725]. Millenarianism [2907]. Millennialism [1444]. Miller [1448, 428, 2741, 737, 1180, 1286, 233, 1180, 1852, 2481, 1727]. Millikan [869]. Milton [2017]. Milwaukee [1110]. Mind [246, 1428, 466, 2055, 2683, 382, 295, 2277, 2575, 1896, 99, 654, 832, 2163, 1949, 683, 2742]. Mind-Brain [2163]. Mind-Forg’d [2742]. Minds [2858, 1763, 2482]. Miner [2375]. Mineral [95]. Mineralogie [1830]. Mineralogy [2361]. Minerals [2185, 95]. Miniature [33]. Miniatures [1956]. Minima [1977]. minime [578, 2129, 1394]. Mining [606, 2823, 2861]. Minne [824]. Minor [1981]. minora [916]. Mintz [2033]. Minutiae [2619]. Mira [784]. mirabilium [1778]. Mirage [266]. Mirail [496]. Mirandole [2127]. Mireille [1819]. Mirella [336]. Miriam [1268, 2173, 1280, 2352]. Mirko [304, 544, 1207, 1431]. miroir [2433]. Mirrors [808, 1003]. Miscellaneous [1933, 2221]. Misconception [596]. Mismeasure [674]. Missing [2714, 1106]. Mission [481]. Mistichelli [542]. Mistrust [620]. Misunderstanding [941]. Mitchell [1915, 885, 1955]. Mitscherlich [2726]. Mitschrift [1153]. Mittelalter [1278, 2897]. mittelalterlichen [1203, 1275, 1595, 503, 1945]. Mittelalters [38]. Mittelitalien [1379]. Mittelstaedt [26]. Mobilizing [2625]. Mobius [911, 911]. Mode [2799].

Model [2381]. modeles [920]. Modeling [1833]. Modell [1172]. Models [532, 1363, 2430, 272, 2472]. Moderate [1859]. Modern [1252, 431, 2798, 1195, 22, 1166, 1631, 2487, 1985, 153, 2200, 2672, 2821, 1960, 961, 1235, 1377, 1967, 2816, 148, 948, 2902, 62, 840, 565, 1502, 1751, 1516, 1087, 1380, 2886, 657, 1815, 1539, 1910, 947, 1537, 1736, 2325, 743, 2264, 1002, 2823, 1554, 440, 2735, 2508, 422, 2941, 1033, 2024, 2653, 350, 1091, 1953, 2098, 406, 1351, 147, 2868, 1404, 2358, 588, 1729]. moderna [1232, 1917, 1503]. moderne [573]. modernen [108, 2720, 1650, 1029]. moderner [1308]. Modernism [2492]. Modernist [2307]. Modernity [892]. Modernization [2444]. modernos [2911]. Mogenet [2043, 1592, 334]. moitié [2810, 625]. Mol [1605]. Moldow [2301]. Molecular [428, 429, 1798, 2850, 2926]. Molecules [1264, 2831, 2099, 1162, 2926]. Molella [731, 2057]. Moll [1285, 241]. Mollan [1885, 2661, 513]. Molly [1328]. Moment [1970, 2116]. Momento [383]. monades [854]. Monck [2938]. monde [2824, 979, 1506, 1431, 1857]. Mondeville [1343]. Mondo [1517]. Money [2607, 2509]. Monod [429]. Monopoly [2116]. Monsters [981]. monstres [651]. Montgolfier [1182]. Montgomery [586, 1142]. Monti [1142]. Montmorillonite [2101]. Monument [2559]. Moone [897]. Moons [1413]. Moore [1864, 1959, 774, 277, 837, 1151, 1160, 160]. Mora [2639]. Moral [2912]. Morality [1900, 759]. Morals [2547]. Morantz [1767]. Morantz-Sanchez [1767]. Morbid [2644]. Mordechai [1474]. Mordier [708]. Moreau [1428, 651]. Morello [488]. Morelon [2895]. Morgagni [1668]. Morgan [12, 2652, 959, 878, 2004]. moriscos [1781]. Morison [602]. Moritz [1475]. Morphology [1098, 2153]. Morrell [2288, 1312]. Morris [820, 297, 2099, 1525]. Morrison [1992]. Morrison-Low [1992]. Morselli [1356]. Morsink [974]. Morss [2420]. Mortality [2200, 299]. mortice [218]. Mortimer [2552, 1816]. Morton [907, 1529, 1569]. Moruzzi [2555]. Moruzzi-Bayo [2555]. Mosaic [1874]. Moses [2146]. Moshe [818]. Moss [1958, 369]. Mossotti [860]. Most [738, 2026, 312, 1611, 1516, 2410]. Mostly [2555]. Mother [1365]. Mothers [2647, 2583]. Motion [2543, 2128, 195, 2475, 1622, 722]. Motivations [2335]. Motive [154, 2220]. Motors [1381]. Mott [1023, 2143, 554]. motuum [1686]. Mould [2922]. Mount [2662, 822]. Mountain [2848, 2360, 2145, 951]. Moussaieff [2857, 1422]. Move [1281]. Movement [2771, 2440, 1741, 775, 1622, 1100]. Moxa [783]. Moyal [76]. Moyen [1343, 1471]. Moyer [1317]. Mozley [76]. Mr [187]. Mr. [390, 1747]. MS [1439]. mu [378]. Much [1855, 1839, 1140]. Mueller [846]. Muhle [891]. Muhly [752]. Muir [2028]. Mulder [2058]. Mülheim [2763]. Mulholland [2168]. Mulhouse [507]. Mulkay [1408, 1409, 157]. Muller [2927, 946, 2441, 154, 2498]. Muller-Hill [2927]. Multhauf [2258, 1304]. multiplicatione [1177]. Mulvaney [1839]. Mundane [2779]. mundi [1774, 1976]. Munoz [1681, 1037]. Münsterberg [170]. Muraro [42]. Murder [1701]. Murderous [2927]. Murdin [1858]. Murdoch [1552, 2448]. Murray [295, 1956, 115]. Muse [2092, 2177]. Museet [370]. Museum [1644, 768, 1978, 390, 2348, 2258, 2103, 2162, 756, 2405]. Museums [2467, 2465, 2331, 2882, 102]. Music [2453, 1604]. Musical [123]. Musicke [2453]. Musschenbroek [313]. Musser [2475]. Musso [2808]. mutation [574]. My [131]. Myers [2059]. Myles [1174]. Mysteries [2762]. Mysterious [187]. Mysterium [1518]. Mystery [94, 2351]. Mysticism [1313]. Myth [1329, 2092, 127, 1301, 2256, 6, 1527, 795, 1171]. Mythical [1049]. Myths [621].

N [1785, 1756, 1662, 281, 169, 2556, 2240, 2564, 614, 14, 2667, 2650, 627, 361, 2040, 226, 2846, 2529, 384, 989, 1325, 657, 1483, 99, 2921, 2430, 2460, 2334, 1482, 1753, 2651, 695]. NACA [796]. Nacci [1790]. Nach [715, 2813]. nachalo [74]. Nachdruck [2586]. nachgelassene [553]. Nachmansohn [322]. Nachtrage [2702]. Nafic [308]. Nagel [2740, 208, 1777]. Nahin [2746]. Nai [377]. Nails [2372]. Naini [1278]. naissance [500, 371]. naixenca [1192]. Name [1805, 1762]. Names [2835, 1762, 2098]. Namias [2104]. Namnun [238]. Nan [1035, 378]. Nan-fang [378]. Nancy [1144, 1887, 692, 1167, 1610, 2930, 2346, 781, 1769, 1027, 2818]. Napoli [563]. naravi [2739]. Narcotic [1272]. Narratio [1319]. Narrative [1406, 2755]. Nathan [653, 1367, 732, 731, 2561, 817, 2057, 1919, 2425]. Nathaniel [1142, 2529]. Nation [2329, 2642, 2192]. Nation-State [2329]. National [46, 344, 80, 184, 1631, 210, 469, 1096, 658, 2154, 1500, 101, 1978, 18, 2484, 2258, 2118, 1494, 2765, 2847, 1077, 2381, 1357, 1526, 2765, 210]. Nationalisation [710]. Nationalised [1032]. Nationalization [2194]. Nationalsozialismus [2489, 1969, 1838]. nationalsozialistische [154]. Nations [1806]. Native [2874, 2215]. Natur [2817, 309, 1469, 862, 1017, 1945]. Natur-und [1469]. natura [1505, 1517, 63]. Natural [728, 857, 1252, 2406, 1608, 903, 559, 1237, 779, 1786, 2399, 471, 1580, 859, 1721, 768, 1355, 852, 1235, 1520, 1173, 382, 2697, 151, 2600, 390, 2465, 837, 2418, 2130, 1582, 2103, 927, 2448, 855, 2481, 328, 258, 1145, 1178, 1339, 571, 421, 1759, 1818, 2781, 68, 791, 680]. naturalis [108]. Naturalism [413]. Naturalist [1355, 561, 258, 923, 2145, 986]. Naturalistic [2398]. Naturalists [2192, 2555, 951, 1489]. Nature [857, 786, 886, 2073, 1105, 1177, 1302, 777, 1833, 166, 2218, 1778, 678, 213, 1363, 816, 399, 338, 1510, 758, 1005, 142, 1651, 919, 2011, 126, 1825, 1446, 1994, 2093, 2212, 2023, 291, 1855, 327, 1834, 1235, 1269, 1005, 1090, 1338]. naturelle [1162, 2942]. Naturforscher [2483]. Naturforschers [2533]. Naturgeschichte [292]. naturhistorische [706]. Naturonderzoeker [313]. Naturphilosophie [448, 2906]. Naturselbstdruckes [1660]. Naturtheorie [603]. Naturwissenschaft [2913, 1906, 2808, 275, 19, 2245, 1945, 353]. Naturwissenschaften [1847, 345, 108, 1640, 2807, 643, 2704, 1141]. Naturwissenschaftler [499, 2854]. naturwissenschaftliche [1576]. Natuurwetenschappen [206]. Nauchnaya [2903]. nauk [273]. nauka [2816, 2549]. nautischen [1655]. Naval [2833, 865, 1024, 792]. Navarro [1782, 1681, 1232]. Navigation [2723, 756, 27]. Navigators [270]. Navy [2062, 865]. Naylor [2113]. Nazi [2490, 2851, 2491]. nazioni [579]. Nazis [2651]. Neal [316]. Near [1080]. Nearly [1980]. Neckers [81]. Nederland [2825]. Needell [1292]. Needham [2402, 2588, 118, 1679, 2370, 2347, 2589, 876, 783]. Ne’eman [2252]. néerlandaise [615]. Neglect [2468, 2648]. Neighbors [2757]. Neil [1934, 1921, 1899]. nel [112, 1517, 1890, 2051, 1436, 1609, 644]. Nelken [514]. Nelkin [618, 2297, 1244]. nella [604, 1189, 1332, 2601]. Nelles [2116]. nell’eta [1643]. nello [1682]. Nelly [1029]. Nelson [2564, 892]. Nemesis [266, 2022, 528, 984]. Nemore [850]. Neoplatonism [2447, 2278]. Neptune [2787, 67]. Neri [724]. Nersessian [1887, 2346]. Nervosa [2886]. Nervous [2303]. Ness [2351]. Nesson [1501]. Nest [2406, 2302]. Nesta [951]. Nested [2469]. Nesting [2469]. Network [2787]. Networks [1139]. Neu [412]. Neuen [1760]. neuere [637]. Neugebauer [1318]. Neuman [1693]. Neumann [2249, 420]. Neumeyer [1566]. Neuorientierung [2838]. Neural [350]. Neuroanatomy [1513]. Neurophilosophy [2163]. Neurophysiology [2573, 1601]. Neuropsychiatry [911]. Neuroscientific [2574]. Neuroses [245]. Neurosis [1584]. Neutron [2844, 2571]. Neuzeit [1278]. neuzeitlichen [1945]. Never [538]. Nevill [2143, 554]. Neville [2485]. Nevskaia [1753]. Newark [2642]. Newell [1894, 659]. Newland [76]. Newly [1012]. Newman [2673]. News [263, 536, 943, 463, 747, 1820]. Newton [2915, 720, 1603, 1984, 509, 1136, 1349, 1858, 2217, 1482, 1257, 1733, 1351, 2906, 668, 1872, 742, 809, 1536, 2235, 755, 60, 6, 2525, 1136, 1515, 1733, 2737, 2884, 790, 538]. Newton-Smith [755]. Newtoniaans [313]. Newtonian [2452, 1041, 983, 1703, 1684]. newtoniano [1517]. Newtons [1469]. newtonsche [1519]. Neyman [1161]. Nicephore [1592]. Nicephorus [1592]. Nicholas [1981, 1657, 738, 217, 354, 2120, 2665, 2948, 1586, 2865, 1723, 2735, 1099]. Nichomachus [111]. Nichtwissenschaftlichen [836]. Nick [2607]. Nickel [36]. Nickles [611, 476]. Nico [2560]. Nicolaidis [1753]. Nicolas [280, 1419, 2592, 1958]. Nicolaus [1909, 1777]. Nicole [1778, 399, 1778]. Nicoletta [488]. Nicollet [512]. Nicolson [537]. Niels [2456, 1720, 392, 2456, 1120, 1867]. Nielson [1120, 1487]. Nigel [1408, 756]. Night [2568, 673]. Nightingale [2056]. NIH [2087]. Niiniluoto [2085]. Nikolai [2832, 285]. Nilo [1739]. Niman [1178]. Nine [1744]. Nineteenth [1481, 15, 2634, 2573, 1023, 2644, 2487, 168, 2922, 1355, 2833, 1222, 2222, 1406, 1447, 2140, 968, 2753, 1269, 987, 2441, 2517, 2646, 1473, 1731, 2019, 1003, 827, 758, 2754, 2575, 447, 1862, 2324, 652, 711, 1145, 1816, 102, 2686, 1317, 2633, 1546, 2549, 1768, 1989, 2743, 1489, 97, 1455, 1217, 1535, 2574, 1465, 1472, 1952]. Nineteenth-Century [1481, 2833, 1222, 1406, 2753, 1269, 2441, 2517, 2646, 1473, 1003, 827, 758, 2575, 1862, 2324, 1145, 102, 2633, 1768, 1489, 97, 1455, 1217, 1535, 2574, 1472, 1952]. Ninive [1201]. Nino [703]. Ninth [1385]. Nissenbaum [365]. Nitschke [275]. Nizar [238]. No [500, 2109, 1841, 2628, 1024]. Noam [2420]. Nobel [736, 738, 131, 2551, 1738, 2193, 1216, 1309]. Nobis [1909]. Noble [2618, 2939]. Noel [685]. Noether [867]. Noguchi [1390, 391]. Noll [1249]. Nollet [1602]. Nombre [89]. nomenclatura [442]. Nomenclature [1758]. Nominators [2551]. Nominees [2551]. Non [341, 2941, 2809]. Non-Technical [341]. Nondecimal [826]. Nonevolutionary [1764]. Nonsense [204]. Nora [2636]. Norbert [2293, 420, 2940]. nord [1335, 1375]. Nordström [1074]. Normal [1454, 228]. normale [958]. Normalität [2330]. Norman [344, 1512, 2894, 605, 1504, 1270, 851, 713, 2165, 410]. Norms [352]. Norriss [2569]. North [49, 1260, 2453, 2399, 122, 1038, 1964, 1582, 2142, 2823, 2332, 1163, 1051, 2011, 2749]. North-East [2399]. Northeast [558]. Northeastern [558]. Norton [2299]. Nostri [2661]. Notebook [942, 863, 1349]. Notebooks [2748, 242]. Notes [2453, 1177, 2399, 1877, 771, 1536, 722, 2497, 582]. Nothing [1140]. notre [906]. nova [1744, 54]. novae [1686]. Novelists [2920]. Novels [2303]. November [2716, 1135, 672, 108, 812, 1879, 882, 2387, 1801, 1300]. Novembre [46]. Novita [1000, 1685]. nowozytna [2816]. NRC [210]. Nriagu [1434]. NS [353]. NS-Ideologie [353]. Nuclear [2839, 1656, 2021, 1194, 1290, 2017, 2180, 2225, 2840, 2063, 2472, 2758, 1694, 871, 2760, 1529, 323]. Nueva [2914, 442]. nuevo [1681]. Nuffield [956]. Numa [379]. Number [558, 2032, 2052, 460, 2671, 1101, 1745]. Numbers [80, 1570, 1199, 366, 2736, 2093]. Numeracy [1020]. Numeration [826]. numeris [850]. Nunis [963]. Nurnberg [626]. Nursing [2780]. Nutton [2892, 303]. Nye [1415, 2424].

O [2891, 2605, 2186, 2637, 760, 487, 2561, 2045, 515, 2052, 1814, 2104, 2708, 972, 1191, 1318, 1596, 1434, 1315, 2739]. O.P [1626]. Oakley [1974]. Obituaries [2293]. Objectivity [2569, 2577, 2663]. Objects [678, 1570, 2331]. Obliquity [999]. obra [2593]. O’Brien [778]. Obrist [1342]. obschei [285]. Obscure [260]. observaciones [2911]. Observateurs [223]. Observation [530, 1848, 1708, 2663, 1960]. Observational [942]. Observations [2864, 963]. Observatoire [1705]. Observatory [2916, 2168, 2559, 1357]. Observed [1322, 1330, 1409]. Observers [1330]. Obsidian [2123]. Obstetrics [980]. Occidens [1776]. Occident [237, 639]. occidentale [1431]. Occult [1273, 470, 595, 2902, 109, 1976, 1551]. Occultist [107]. Ocean [2148, 1421]. Oceanographic [190, 1024]. Oceanography [699]. Ocherk [285]. Ockham [2044]. Octavio [860]. October [2070, 670, 1133, 264, 881, 1942, 812, 332, 2387, 1801, 944]. Octobre [496]. Odelberg [824]. O’Donnell [1897, 1607]. Odor [2352]. oeuvre [625]. Oeuvres [2814]. of/in [421]. oferta [1613]. Off [2247]. Office [2833, 1024]. Official [2118]. Offray [2827]. oggi [604]. Ogilvie [2140]. O’Hara [2703]. O’Hear [889]. Ohm [2073]. Oil [620]. Oken [479, 1877]. Okruhlik [2130]. Okun [1955]. Old [2110, 813, 1997, 2655, 1760, 2795, 2802, 2686]. Oldenburg [448, 2156]. Oldroyd [1250, 2164, 794]. Ole [1747]. Oleg [2532]. Oleson [2368, 2012, 1674]. Olga [2780, 2654]. Olive [2640]. Oliver [2917, 2746]. Olmsted [2289]. Olof [1906, 2401]. Olson [1087]. omnia [1847, 1983, 849, 1846]. Once [1752]. One [252, 521, 802, 1062, 1385, 1698, 2001, 2316, 2611, 2765, 2290, 1151, 2250, 1932, 2138, 2106, 2493]. O’Neil [2570]. Ongaro [1179]. Only [1980]. Onsager [1257]. Onslaught [2271]. Ontario [788, 410]. Ontologie [2245]. Oort [627, 627]. op [1780]. Oparin [28]. Open [1365]. Opening [1239, 1408]. Opera [916, 849, 1846, 1847, 1983]. operaia [1009]. opere [44, 2280]. Ophir [133]. opinion [2804]. Opinions [2518]. Oppenheimer [393, 2286]. Opportunity [1236]. Opsomer [1676]. Optical [2319, 809, 1733]. Optics [2816, 457, 2384, 1509, 1482]. optique [573]. Optyka [2816]. Opus [1744]. Orations [1443]. orbites [388]. Orbs [2066]. Order [1448, 3, 1044, 170, 1455, 68, 2797, 722, 2828]. Ores [2137]. Oresme [1778, 1675, 399]. Organic [1872, 2359, 78, 2181, 1758]. Organischen [475]. Organised [513]. Organism [1267]. Organisms [876]. Organization [746, 148, 937, 1813, 1773, 584, 1862, 102, 2116, 2760]. Organizer [2769]. Organizing [1795, 2806]. Oriens [1776]. Oriens-Occidens [1776]. Orient [1278, 1672]. Orientalism [171]. origen [1868]. Origin [1187, 666, 2253, 1648, 2040, 7, 215, 2479, 729, 1220, 865, 2400, 1831, 498, 215, 769]. Original [1893, 48, 2401, 2407]. Originalarbeiten [1893]. origine [651]. origines [223]. origini [1749]. Origins [1160, 492, 2574, 559, 1897, 80, 495, 2667, 567, 899, 2467, 2775, 879, 1985, 2445, 1100, 302, 814, 693, 2466, 1173, 62, 2546, 104, 1923, 2441, 2494, 1921, 2034, 170, 137, 1584, 2179, 2823, 2518, 472, 919, 844, 2554, 1800, 286, 588, 2030, 2438, 1925, 1358, 2135, 429]. Orme [2634]. Orna [2166]. Ornithology [1579]. Orreries [25]. Orthodoxy [1042]. Osborne [2520]. Oscillations [726]. Oscilloscope [736]. Osier [1639]. Oskar [373]. Osler [1886, 494, 1334]. Osnovanie [273]. Ospovat [464, 927]. Osteopathic [1270]. Osterbrock [1692]. Osterreich [1055]. Ostwald [696]. Other [216, 2076, 1105, 2553, 2055, 208, 1395, 2213, 826, 954, 2931, 2750, 2865, 1596]. Otte [652]. Otthein [2330]. ottiche [1188]. Otto [1354, 914, 1158, 612, 287, 1153, 188, 871, 1656, 134, 856]. ottocentesco [1609]. Ottocento [2601]. Ottosson [1779]. ouab [1673]. Our [2706, 2562, 1405]. Outer [902]. Outrage [1448]. Outram [387, 1863]. ouvrages [1705]. Overman [2943]. Overview [932]. Ovitt [2521]. Owen [288, 1413, 298, 1960]. Own [2562, 2749, 1131]. Oxen [2214]. Oxford [474, 1644, 1146, 675, 1069, 1706, 1473]. Oxygen [165]. Ozminski [2240].

P [1286, 2762, 1844, 281, 1195, 1780, 74, 1612, 1893, 1605, 1260, 2431, 511, 123, 1205, 575, 1775, 731, 2055, 1983, 2225, 515, 1867, 2293, 578, 2936, 1521, 2846, 2215, 1761, 2057, 1336, 1907, 231, 416, 638, 2938, 2780, 494, 2258, 103, 1325, 1180, 2850, 1770, 2324, 133, 543, 1649, 964, 900, 2707, 2460, 2129, 1596, 1145, 1570, 1352, 1952, 1341, 680, 1036, 619, 107, 2415, 1989, 1828, 2397, 588, 1394]. P.-L [1983]. Pacey [1361]. Pachella [109]. Pacific [1808]. Padova [1890, 1682]. padre [1987]. Padua [1000]. Paganini [2160]. Pagano [1543]. Pagden [1173]. Pagel [1135, 2278, 1043]. Pages [668]. Pagnoni [1847].

Pagnoni-Sturlese [1847]. Pain [654, 1768]. Pais [1138]. paix [86]. Palaeontology [1118, 2501]. Paleographical [242]. Palermo [1253]. Pallister [981]. Palm [647]. Palmer [94, 1334, 1585, 2038]. Pamela [2045, 1497]. Pancaldi [1326, 1643]. Pandora [1408]. Pantheists [1116]. Paola [112]. Paolo [2820, 44, 1685, 579, 1856, 1980, 383]. Papadimitriou [718]. Paper [336, 403, 2410, 2370]. Paperclip [2491]. Papers [1056, 1933, 2289, 2100, 843, 653, 2555, 1159, 700, 732, 768, 1146, 580, 382, 858, 2741, 2057, 1136, 513, 1150, 2221, 1093, 2475, 2654, 754, 1142, 238, 2024, 92, 1733, 152, 1893, 720, 382, 1884, 2033, 2943, 2249]. papiri [2122]. Papon [1410]. Papp [1327]. Pappus [1132]. Papyrus [845]. Paracelso [2801]. Paracelsus [2278, 1351]. Paracelsusforschung [785]. Paradigm [1012]. Paradigmatic [165]. Paradigms [409]. Paradis [957]. Paradise [907]. Paradox [1652, 1425]. Paralipomenes [573]. parallèles [2809]. Paranormal [972]. Parascandola [1751, 1197, 587]. Pardo [1782, 2751]. Pare [981]. Paris [46, 184, 304, 478, 565, 977, 2641, 1705, 1348]. Parisian [2429]. Park [98, 1494, 1853]. Parker [1855]. Parkinson [1946]. Parks [2938]. Parliament [749]. Parodi [1436]. Parrott [1366]. Parsons [2329]. Parssinen [1272]. Part [1893, 2555, 362, 1686, 106, 2588, 1379, 2348, 1679, 2370, 2589, 2103, 754, 164, 172, 1449, 92, 2415]. partes [1035]. Partial [1240]. Participant [1078]. Particle [1414, 1416, 2252, 1891, 2954]. partielle [975]. Partnership [843]. Parts [504]. Pascal [1743, 2905]. Paskoff [1773]. Pasquale [2338]. Passage [1811]. Passenger [34, 2117]. Passions [1272, 2880]. Passmore [79]. Past [2890, 1254, 1080, 1486, 2428, 699]. Pasteur [403, 2319]. Pat [2656]. Patenting [2786]. Pater [313]. Patey [1480]. Pathological [228]. pathologique [1431]. Pathology [2644, 765]. Paths [2547]. Pathways [362]. Patient [2112]. patológica [444]. Paton [2183]. Patrice [819]. Patricia [1020, 2163, 2441]. Patrick [614, 546, 2516, 2945, 54, 1606, 651, 1355]. Patristic [2232]. Patron [1096]. Patronage [1703, 1389]. Patrons [391]. Pattaro [644]. Patterns [2920, 2679, 1172, 2664]. Patterson [32]. Paul [726, 2240, 2822, 1424, 930, 2748, 119, 911, 1857, 2298, 1095, 1205, 607, 2553, 2949, 2281, 1331, 191, 290, 2691, 1886, 2139, 1595, 2057, 987, 373, 1499, 1719, 729, 1903, 2690, 2393, 109, 2392, 1773, 2475, 441, 133, 2457, 1153, 2746, 2944, 2460, 1364, 1931, 923, 1579, 1230, 1677, 2098, 628, 2716, 478, 317, 1912, 318]. Paula [124]. Pauli [1710, 452, 451, 1866]. Pauling [2850]. Paulo [1181]. Pausek [2739]. Pausek-Bazdar [2739]. Paved [224]. Pavlova [2523]. Pavlovic [2776]. Pavone [2595, 2280]. Pawel [1981, 73]. Pawnee [1038]. Pax [2758]. Payment [2113]. Payne [1395]. Payne-Gaposchkin [1395]. Pays [1335]. Peace [258, 1869]. Peacocke [2495]. Peak [1357]. Peale [1286, 390, 2741]. Peano [552, 552]. Pearce [953, 121]. Pearl [1849, 1850]. Peasants [1333]. Pedersen [1315]. Pediatrics [2849]. Peer [2296]. Peierls [1811]. Peiffer [2904]. Peirce [583, 1864, 2530, 928, 1951, 217, 2374]. Peking [656]. Pelaez [1987]. Pellegrin [2445]. Pelletier [607]. Pelling [175]. Pelner [180]. Pena [1037]. Pendula [2128]. Pendulum [2128]. Penelope [2170]. Penicillin [2228, 2061, 912, 1973]. Penick [698]. Penina [2173]. Pennsylvania [300, 1773, 1818]. Penny [2562]. Penrose [2845]. Penser [958]. pensiero [112, 2595, 644]. Penury [2591]. People [569, 2567, 1428, 529, 2496, 1020, 1571, 296, 1885, 1537, 135]. Pepper [2120, 1869]. Per-Gunnar [1779]. Pera [2219]. Perception [2384, 2945]. Perceptions [964]. Percy [2270]. Perez [1502]. Perez-Gomez [1502]. Perfection [2521]. Performance [2253]. Pergamon [110]. periferica [1917]. Perihelion [1483]. Periodic [1810]. Periodicals [1947, 2479]. Periphery [2875]. Permanence [2707]. Pernick [1768]. Perrot [2699]. Persian [635]. persischer [1278]. Persistence [1764]. Person [627, 1301]. Personal [246, 746, 2148, 1997, 1359, 2292, 2646]. Personalities [412]. Personality [2331]. Persons [634]. perspectivas [1917]. Perspective [225, 2679, 2268, 1227, 839, 1031, 2929, 2208, 245, 417, 306, 1397]. Perspectives [2530, 467, 825, 2052, 2687, 1409, 1539, 2478, 1368, 1930, 1273, 631, 1193, 957, 2858, 2422, 967, 2183]. Perspectivist [457]. Perspektiven [2854]. Peru [2875]. Pervasive [842]. pervoi [2846]. Pesenti [1682]. Pesticide [269]. Pestre [2760, 1754]. Peter [2260, 1548, 1949, 730, 2531, 2748, 2534, 2706, 2399, 2216, 2733, 1164, 2012, 1184, 1094, 2099, 2368, 1396, 2355, 1960, 2294, 1566, 2084, 1995, 1577, 352, 1674, 723, 1516, 2341, 1982, 2505, 2470, 2654, 1956, 2524, 1738, 2518, 440, 1163, 2323, 1569, 26, 1308, 2504, 1008, 1572, 392, 2188, 2822]. Peterburgskoi [273]. Peters [2753]. Petersburg [273]. Petersburgskaia [1753]. Petit [114, 482]. Petroleum [571]. Petroski [1927]. Petrus [313]. Petry [2704]. Pettijohn [1418]. Peyrère [2596]. pflanzliche [1760]. Pfund [2930]. pH [700]. Ph.D [1401]. Phaff [2079]. Pharaonic [1331]. Pharmaceutical [2506]. pharmaciens [2266]. Pharmacology [1197]. Pharmacy [2010, 1028]. Pharmazie [2838]. Phase [1465]. Phases [447]. PhD [80]. Phenomenon [759]. Phil. [1642]. Philadelphia [1822, 1196, 1818]. Philanthropic [1954]. Philanthropy [770, 2426]. Philebus [460]. Philip [887, 2270, 1290, 2672, 1834, 2040, 2591, 909, 2889, 1489, 1529, 339]. Philipp [2838]. Philippa [2744]. Philippe [1343, 2667, 2513, 772, 744]. Philipps [345]. Phillip [857, 110]. Phillips [1568, 680, 2131, 1894, 1963]. Philological [1796]. Philopon [1036]. Philoponus [2369]. Philosopher [434, 2282, 1507, 1178, 2912]. Philosophers [2694, 61, 509]. philosophes [2782]. Philosophic [1950, 1480]. Philosophical [2456, 511, 2311, 2013, 1840, 270, 1943, 207, 580, 1235, 382, 974, 1269, 2161, 1150, 2346, 1349, 2419, 1818, 2781, 413, 2421, 2212, 1338, 2414, 1489, 2336, 152, 2585, 356, 43]. Philosophicall [240]. Philosophie [1855, 547, 862, 2287, 677, 1488, 958, 1175, 1186, 475]. philosophique [2476, 712, 2127]. philosophiques [2827, 2416]. philosophischen [1029]. philosophischer [1186]. Philosophy [650, 1785, 2477, 995, 684, 610, 2913, 1362, 1237, 1177, 583, 1786, 1779, 470, 1580, 217, 78, 409, 608, 1721, 982, 2371, 2672, 2480, 1713, 1235, 382, 1720, 208, 1730, 357, 1716, 1884, 1336, 176, 162, 978, 2894, 2164, 2864, 545, 1463, 1714, 2945, 2130, 789, 1283, 133, 1561, 2524, 2879, 363, 1077, 1715, 2448, 2865, 2520, 1228, 328, 2334, 1145, 1364, 2011, 609, 787, 2398, 779, 2733, 1775, 153, 852, 2776, 2697, 151, 2600, 778, 1516, 1180, 2673, 1148, 2412, 1888, 1499, 846, 2345]. Phlebotomy [1438]. Phlogiston [165]. Phosphorus [2618]. Photographers [951]. Photographic [1556]. Photographie [1660]. Photographing [2887]. Photographs [1556, 1997]. Phrenology [1862]. Phylosopher [1185]. Physic [2426]. Physicae [1152]. Physical [2913, 2453, 1960, 404, 900, 594, 1939, 242, 1258, 2731, 805, 2165, 1526, 1712]. Physician [1666, 1687, 1178]. Physicians [1196, 366, 1331, 1767, 1027]. physicien [1602]. physiciens [1754]. Physicist [132, 1185, 868, 2290, 2460, 459, 1811, 2531, 2060, 673]. Physicists [2499, 1554, 422, 61, 1989, 2925]. Physick [1585]. Physics [1415, 2073, 2676, 1746, 1216, 1144, 322, 2180, 530, 1704, 1638, 485, 2846, 1456, 2551, 133, 1554, 1313, 517, 1257, 1317, 436, 2044, 323, 949, 2731, 1338, 1800, 2926, 90, 1082, 1989, 1526, 210, 1145, 1148, 2762, 1656, 1147, 683, 1358, 961, 1712, 1283, 1414, 1727, 1145, 554, 1891, 1963, 2918, 2100, 1984, 2157, 1894, 753]. Physik [91, 1019, 761, 1017, 1650]. physikalische [287]. physikalischen [1019, 2740]. Physiological [2718]. Physiologie [920, 582]. Physiologists [675]. Physiology [1391, 207, 1667, 1454, 13, 1217, 560, 222]. Physique [651, 776, 500, 1754]. Piaget [2775, 2327]. Pic [2127]. Pickering [257, 533, 1891]. Pickstone [2035]. Pictorial [48, 220, 1024, 2141]. Picture [2475, 457]. Pictures [2162]. Pie [1188]. Pier [1188]. Pierangelo [2601]. Pierce [990]. Piero [151]. Pierre [2880, 1285, 237, 2266, 2904, 2445, 482, 1335, 708, 922, 1983, 1047, 1734, 1743, 2053, 895, 1410, 2416, 2124, 651, 2594, 1319, 2170, 2697]. Pierre-Jose [1743]. Pierret [2629]. Pieter [117, 1758]. Pietro [77, 585, 2454, 1230, 1542]. Pigeaud [712]. Pigeons [327]. Piggott [1378]. Pighetti [1609]. Pilcher [2270]. Piles [2469]. Pilots [2936]. Piltdown [2187]. Pilz [626]. Pinborg [2894]. Pinel [772]. Pinero [1782, 1232, 1584, 1765, 444, 2751, 443]. Pingree [40, 642]. Pinsel [1024]. Pio [1188]. Pioneer [569, 1692, 986, 577, 926, 1962, 1935, 2602, 1670]. Pioneers [2275, 322, 2099, 2340, 2134, 224]. Pioniere [1104]. Piotr [2108]. Piperidin [2498]. Piquer [2911]. Pisa [610, 1000, 609]. Piternick [929]. Pitfalls [2930]. Pitt [2778]. Pizzamiglio [1188]. Place [1608, 922, 1463, 2124, 1994]. Places [1885, 634, 1202]. Placides [979]. Plague [2263, 1903]. Plamondon [78]. Planck [2950, 341, 2333]. Planet [1567, 1693]. planétaire [438]. planétaires [1441]. Planetarians [286]. Planetariums [25]. Planetary [2018, 286, 1413]. planètes [438, 388]. Planets [902, 320, 286]. Planispheric [1978]. Planners [841]. Plant [1662, 1166, 702, 767, 2362, 2724, 765, 1762]. Plant-Name [1762]. Plantae [2364]. Plantas [2914]. Plantes [1348]. Plants [361, 2873]. Planude [848]. Plastic [1670]. Plate [1320, 574]. Platinum [962]. Plato [2692, 778, 460, 638, 110]. Platonick [1855]. Platonism [2447]. Play [1955]. Playfair [163]. Playing [1455]. plena [2661]. Plesset [391]. Pliny [2446]. Plotin [1175]. Plots [1406]. Plough [1822]. Pluchon [2266]. Plurality [2082, 2124, 899]. plus [2050]. Pluto [320]. Poems [279, 1568]. Poetic [1158]. poétique [2824]. Poetry [2304, 758]. Poets [2305]. Poincaré [1021, 2357]. Point [2067]. Poison [859, 1770]. Poisoning [1434]. Poisonous [2136]. Polar [664]. polarisation [1162]. Polemics [2915]. Policies [2237, 2088, 2377, 1529]. Policy [2017, 2619, 2510, 2225, 2139, 2087, 95, 417, 2757, 1031, 469, 2172, 687, 1645]. Policymaking [871]. poliia [2846]. Polish [1565]. politica [2601]. Political [2477, 2329, 529, 2724, 1004, 2294, 491, 35, 430, 2908, 2933]. politiche [1332]. Politicians [1032]. Politics [2113, 2016, 1631, 620, 2112, 1240, 2492, 2063, 1444, 1533, 2923, 2088, 2714, 1211, 2133, 1910, 105, 2607, 1661, 2707, 328, 2772, 1110, 2669, 1991, 991, 1498, 133, 2534, 2833, 2174, 1366]. politique [2563, 2563]. Polity [813]. Pollak [618]. Pollner [2779]. Pollution [2222]. polovine [2846]. Polyalphabetic [7]. Polymath [2459]. Polymer [932, 2099]. Polymers [2253, 404, 2026, 932]. Polyolefins [2026]. Polytechnique [2350, 480, 2350]. Polywater [870]. Pongratz [431]. Pont [2937, 1504, 2809]. Ponts [1369]. Pope [2281]. Popkin [2907, 2596]. Popper [889]. popperiane [1845]. Poppi [1890]. Poppin [2273]. populaires [1496]. Popular [2099, 390, 1862, 493, 2295]. Popularization [757, 1742]. Popularizing [2803]. Populated [2768]. Population [1833, 1245, 2551]. Populations [837]. Porfir [2832]. Port [969, 2368, 300]. Porta [42]. Portela [1782, 1232]. Porter [1842, 2938, 2439, 2406, 2888, 2096, 469, 167, 1088, 1143, 645, 2742]. Portrait [1607, 1521, 456, 869, 1328, 2108]. Portraits [514, 2555, 1242, 605, 2134, 2342]. Portrayed [2550]. posible [1613]. Positions [1413]. Positivism [2196]. positivismo [1643]. Positivismus [1055]. positivistica [1189]. Possibility [2244, 413]. Post [1863, 2175, 1401, 2515, 621, 160, 914]. Post-Darwinian [160]. Post-graduate [1401]. Post-Industrial [2175, 621]. Post-Revolutionary [1863]. Post-Typographic [2515]. postanku [2739]. Postel [772]. Poster [2799]. Postlethwaite [2670]. Postlewait [957]. Postscript [1798, 1735]. Pot [1770]. Pottery [2358]. Pouchelle [1343]. Poulle [780, 1471, 438, 1591]. Poupard [2281]. Pour [2942, 2050, 582, 2594, 1410]. pouvoir [83, 2829, 480]. Povijest [1462]. Powaski [2839]. Powell [756]. Power [2477, 2798, 1928, 1290, 1139, 2840, 1112, 368, 2887, 2220, 1912, 1374, 1929, 890, 616, 56, 1929, 2591]. Powers [2089, 1148]. Practical [918, 236]. prácticas [2911]. Practice [2113, 1427, 530, 2268, 2855, 1689, 308, 1480, 2717, 939, 2346, 1174, 2797]. Practices [1858, 973]. Practitioner [501]. praecognitione [303]. Praelectiones [1152]. Pragmatism [2374]. Prairies [767]. Prakash [2732]. praktischen [2813]. Pratike [236]. Praxis [2830]. Precambrian [487]. Précis [2407]. Précurseurs [2809]. Precursor [1049]. Predication [1715]. Preface [966]. Prefaced [548]. Pregnant [1974]. Préhistoire [2809]. Prehistoric [2123, 2545]. préhistorique [1431]. Preliminary [2600]. Prelude [1115]. Premices [2416]. premiere [625, 2810, 280]. premières [1441]. premiers [772]. Prentiss [721]. Preparation [1748]. Prescher [763]. Prescott [51]. Presence [1603]. Present [702, 1636, 2140, 907, 1529, 2839, 969, 1254, 96]. Presentation [2295]. Presented [2100, 513, 2011, 2212, 722]. Preservation [2166]. Preserved [387]. Presidency [1984, 1996]. President [1812, 2756, 418]. Presidential [1996]. Presidents [225]. Presocratics [2520]. Press [1707, 2297, 147, 493]. presso [1188]. Prest [907]. Prestele [1651]. Presti [1949]. Preston [2930]. Presuppositions [1840, 759]. pretres [1673]. pretres-ouab [1673]. Pretty [866]. Preus [2445, 1336, 775]. Preussen [1008]. Preussens [2199]. preventive [1496]. Previously [403]. Price [1538, 1444, 2238, 768]. Prices [2534]. Pricha [2703]. Priesner [556, 1153]. Priestley [2910, 1444, 384]. prima [1319]. Primary [2869, 685]. primi [2595]. Primitive [839]. primo [1517]. Princes [1333]. Princeton [1249, 653, 732, 2057]. principales [2804, 1035]. Principia [2737]. Principles [2678, 2846, 1516, 973, 1180]. Printed [387, 682, 918, 101, 1818, 2781]. Printing [1380, 147, 2370]. Printmaking [2102]. printsipov [2846]. Prinz [2817]. Prism [2480]. Pritchard [435]. Private [593, 1703, 2113]. Priyadaranjan [784]. Prize [1738, 738]. Prizes [1216, 2551, 1738]. Prizewinner [736]. Probabilistic [2652, 610]. probabilités [1743]. Probability [2652, 2324, 1520, 124, 1480]. Probable [722]. Probes [220]. Probing [972]. Problem [2152, 1872, 316, 2324, 2378, 1145]. Problemi [2160, 1503]. Problems [995, 155, 670, 1571, 652, 954, 335, 2334, 1655, 619]. Problemstellung [241]. Proceedings [2363, 474, 610, 2269, 323, 609, 1315, 238]. procès [581]. Process [216, 1540, 2319, 1714, 2346, 1125, 2656, 689]. Processes [1876, 1925]. Processing [333, 1635]. processo [1543]. Proclus [1175]. Proctor [2651]. Prodigies [1254]. Prodigy [1607]. Producers [2475]. Production [2799, 1549, 2939, 2123]. Profession [1920, 2484, 1473, 1902, 563, 2773, 2034]. Professional [2744, 1311, 173, 97, 2089]. Professionalisierung [1969]. Professionalism [2849, 1768]. professionalita [1009]. Professionalization [1070, 1239, 2301]. Professionalizing [565]. Professionals [256, 2947]. professionellen [2199]. Professions [84, 2561, 2173, 2114, 1476, 1311]. Professor [501]. Professori [1682]. Profile [1418]. Profiles [2496]. Prognosis [303]. prognostico [1681]. Prognosticon [1507]. Program [1077, 594, 1057, 871]. Programming [1624, 873]. Programs [201, 796, 2007, 1748]. progres [582, 651]. Progress [886, 719, 1240, 2083, 354, 2930, 953, 343]. Progression [775]. Progressive [2771, 2085, 126]. Progymnasmata [2594]. Project [2915, 1926, 2759, 569, 1293]. Projectile [2543, 195]. projet [624]. Prokopa [1352]. Prolégomènes [1274]. Prominent [607]. Promoting [1822]. promotori [1682]. Proof [1622]. Propaedeumata [41]. Proper [2161]. Properties [375]. Property [2541]. Prophet [655, 1053]. Prophets [2488, 725]. Prophylaxis [1208]. Proportion [1193]. Proportionentheorie [178]. proposto [1188]. proprietatibus [375]. Propulsion [1057]. Prose [1637, 374]. Proseminar [535]. Prospective [1746, 1410]. Prospects [719, 619, 2478]. prospettive [2160]. proteines [1488]. Proteins [588]. Protestant [160]. Protestantism [1183]. Protocol [143]. Protomedicato [1902]. Proton [1825]. Protyle [1825]. Prout [1825]. Provenance [1785]. Provenzo [2515]. Province [1312, 71]. Provinces [2424]. Provincial [21, 2424]. provinciaux [71]. Provine [837]. Prozessen [1576]. Prucha [2553]. Prudence [1644]. Pryor [2893]. Pseudo [309, 494, 1932, 1407, 1774, 2213, 2517, 2126]. Pseudo-Apollonios [309]. Pseudo-Aristoteles [2126]. Pseudo-Aristotle [1407, 2213]. Pseudo-Bede [1774]. Pseudo-Juanelo [1932]. Pseudo-Science [494, 2517]. Pseudodialecticians [235]. psicologia [703]. psiquiatría [2727]. psychanalyse [708]. Psychiatric [2034]. psychiatrie [772, 1423]. Psychiatry [2650, 1840, 1269, 894, 1905, 1842, 1840, 2637]. Psychical [495, 2209, 1033, 973]. Psychoanalysis [1943, 137, 1948, 1991, 492, 1269, 2505, 137]. Psychoanalyst [1666]. Psychoanalytic [466, 1100, 1948]. psychogenetische [1908]. Psychological [225, 2326, 2453, 1220, 562, 563, 2106, 2770]. Psychologie [1969, 1915, 2031, 1838]. Psychologiegeschichte [2031]. psychologische [2031]. Psychologist [293]. Psychologists [2328]. Psychology [2836, 431, 225, 819, 1897, 2326, 683, 1491, 1639, 224, 1254, 879, 2032, 245, 2325, 2547, 99, 562, 2851, 1763, 2384, 490, 1081, 246, 30, 170, 99, 2326, 631]. Psychosurgery [2931]. Psychotherapy [137, 1916]. Ptacek [2583]. Ptolemaüs [2393]. Ptolemee [2043, 438, 114]. Ptolemy [1432, 2384, 1432]. Public [2113, 2771, 2246, 2564, 1004, 2618, 461, 770, 95, 687, 1209, 1142, 593, 1596, 1587, 1703, 197, 2075]. Publication [2538]. Publications [1639, 1959, 682, 2079, 500]. Publick [1855]. publics [2350]. Published [1544, 1554, 412]. Publishers [1669]. Publishing [679, 149, 258, 1151, 1740]. Pugh [1429, 2038]. puissance [218]. Pulos [1168]. pulsuum [1505]. Pump [1986]. Punched [1746]. Pupils [781]. Pure [1096, 1455, 1955, 950, 1211]. Puritans [145]. Purkert [2566]. Purkyne [2836]. Purrington [1258]. Pursell [915]. Pursuit [1112]. Pusey [1598]. Puzzles [270]. Pyatt [1526]. Pyenson [1826]. Pyne [489, 2184, 893]. Pyridin [2498]. Pyridin-Piperidin-Gruppe [2498]. Pythagoreer [570].

Qaisar [1279]. qanats [505]. Qasim [2586]. Qeuvres [2827, 2895]. Qin [2363]. Qingdao [2363, 2386]. Quadrans [1035]. Quadrant [1445]. Quadreni [355]. Quadricentennial [1512]. Quadrivium [2897]. Quaestiones [1675]. Quaker [2602]. Qualities [595, 1982]. Quality [1014]. quanshu [1102]. Quantifying [1604]. quantiques [933]. Quantitative [1299, 340]. Quantum [1710, 2762, 2097, 546, 1550, 2251, 2675, 2845, 517, 26]. Quark [329]. Quarks [422, 1891, 1654]. Quarrel [509]. Quasi [1590]. Quasi-Archimedean [1590]. quatorzième [853, 714]. quatre [2900]. quatre-cents [2900]. Quattrocento [1890]. que [2592]. Québec [83, 2725]. Queen [474, 1644]. Quen [137, 2858]. Queron [2127]. Querschnitt [2696]. Quest [2270, 1834, 2577, 1654, 1835, 576]. Question [1415, 1245, 2417, 430].

Questions [1349, 374, 242, 291]. qui [2050]. quimeras [2025]. química [442]. Químiques [1192]. Quinn [122]. Quinta [233]. Quintas [2938]. Quipu [622]. Quotations [2186]. Qurra [2895]. Qusta [2391].

R [1759, 2737, 381, 129, 1371, 2762, 2284, 246, 684, 2016, 74, 1656, 2834, 288, 2937, 225, 2100, 2579, 2326, 80, 1746, 495, 2269, 56, 2910, 2312, 631, 583, 2661, 1854, 2555, 2224, 391, 163, 1236, 1333, 270, 575, 608, 1184, 700, 1334, 2183, 189, 385, 2349, 1108, 369, 213, 1638, 2845, 375, 1812, 1438, 1851, 1965, 2682, 2080, 837, 1369, 1814, 100, 167, 1923, 2869, 847, 2938, 1528, 1712, 2038, 2195, 228, 2441, 1463, 1210, 513, 2484, 710, 384, 687, 1919, 2277, 339, 2427, 160, 2272, 25, 1589, 1592, 133, 711, 1956, 2102, 2495, 124, 1848]. R [2587, 1554, 2344, 1337, 188, 919, 1183, 2334, 1992, 1948, 1952, 1898, 1416, 1057, 680, 1899, 2729, 949, 498, 1234, 90, 1963, 2415, 1489, 794, 588, 410]. Rabb [1553]. Rabbitt [2185, 95]. Rabi [2380]. Rabkin [2841]. Race [2839, 738, 934, 1167, 2301, 1797, 2934, 2847]. Rachel [2821, 1363, 2361, 2693]. Racial [2437, 1329, 2651]. Radar [865, 658, 2603, 2852]. Radiant [2063]. Radiation [946, 2703, 2759]. Radical [1116, 2931, 1183, 2090]. Radio [2499, 1484, 243, 658, 1771]. Radioactivity [284]. Radioaktivitat [830]. radioattività [555]. Radiochemistry [1962]. Radnitzky [343]. Rafail [819]. Rafinesque [2407, 925]. Rahn [2131]. Rail [969, 300]. Railroad [34]. Railroads [1371, 1182, 2367]. Railway [2041]. Rain [2757]. Rainger [2682]. Raised [667]. raison [2050]. Rajkumari [2925]. Ralph [221, 2724, 1293, 349]. Rammstedt [2330]. Ramo [2709]. Ramon [442]. Ramsay [760]. Ramunni [1819, 933]. rana [2219]. Rand [2821]. Randier [756]. Randles [574]. Randolph [2239]. Randomization [2466]. Rank [1655]. Raphael [511]. rapide [574]. Rapp [684, 1499]. rapport [582]. Raquel [1987]. Rare [2137]. Rarities [1644]. Rasch [2763]. Rashed [1979, 2391, 2814]. Rashi [2172]. Raskin [2420]. rasprava [2739]. Rassenmythos [2946]. Rather [2441, 229]. Rating [1220]. Ratio [999, 2722]. Rationale [2418, 783]. Rationality [1646, 755, 343, 476, 1194, 1477]. Rauch [2493]. Raup [2022]. Raushenbush [98]. Ravenel [2375]. Ravenswaay [1651]. Ray [1665, 736, 1039, 702, 1997, 1316, 1791, 784]. Rayhan [1277]. Raymond [2725, 224, 125, 2026, 2253, 2056, 83, 694, 2454, 1895, 2206]. Rays [422]. Razvitie [285, 74, 2846]. Razzaz [308]. RCA [2366]. RDS [2661]. Re [907, 606]. Re-Creation [907]. Reaction [1541, 317]. Reactionary [2492]. Reactor [2042, 2121, 871]. Read [922]. Reader [2949, 1106, 2432]. Readings [857, 204]. Ready [2662]. Real [2914]. Realism [2948, 2418, 1719, 1722]. Realist [17]. realité [1431]. Realities [2840]. Reality [2097, 1713, 1798, 2481, 2779, 2343, 2557]. Realized [758]. Realm [1357]. Reappraisal [2285, 1108, 2948, 2524]. Reappraisals [2655]. Reason [1775, 2191, 498, 2779, 2327, 115, 1561, 2344]. Reasoning [508]. Reasons [1439]. Reassessment [2196]. Rebecca [2551]. Rebel [2056]. Reception [458, 211, 107]. recettes [845]. rechazo [442]. Rechenberg [1612, 1893].

Rechenstein [1572]. Recherche [46, 184, 1410, 975, 2563, 2403]. Recherches [906, 677, 2407, 639]. rechter [1908]. reciente [1917]. Reckoning [654]. Recollected [1070]. Recollections [1811, 53, 131, 1618, 1525, 393, 1395]. Recombinant [1058]. Reconstruction [778, 2692]. Reconstructions [545]. Reconstructive [2420]. Record [263, 536, 943, 48, 463, 179, 747]. Recorded [2823]. Records [221, 48, 2555, 2347, 754]. Red [1693, 863, 269]. Rediscovered [1988]. Rediscovery [2894]. Redmond [569]. Redondi [2454, 77, 585, 1542]. Reduced [722]. Redux [1217]. Reed [2558]. Reede [2576]. Reenchantment [820]. Rees [1786]. Reese [1180]. Reeve [1598]. Reference [428, 2553, 2662, 1748, 884]. References [629]. Reflection [399]. Reflections [1484, 1079, 931, 688, 1944, 2060]. Reflexiones [442]. Reflexions [2220, 218]. Reform [1027, 1315, 365, 1110]. Reformer [1043, 2056]. Reformers [1054]. Refraction [1215]. Refugee [1647]. Refusenik [990]. Rega [2898]. Regard [1580]. Regency [2742]. Regime [2110, 1335, 813]. Regina [1767]. Regine [2702]. Regiomontanus [502]. Regiomontanus-Studien [502]. Region [2035]. Régis [2895]. règne [291]. Regulation [1902, 2116, 1694, 551]. Regulatory [2840]. Rehbock [1489]. Rehn [1847]. Reich [1916, 1923, 283, 2490, 337, 154, 2492]. Reiches [353]. Reid [1161, 1836]. Reimar [2808]. Reimarus [1026]. Reingold [731, 817, 2057, 2425, 653, 732, 817]. Reinhard [836, 2355, 475]. Reinterpretation [1630]. Reinterpreting [2526]. Reisacherum [1681]. Rejane [2259]. Rejected [212]. Rejection [2369]. Rektor [479]. Related [819, 2348, 374, 2079, 2081]. Relation [198, 886, 712]. Relations [1105, 1049, 2225, 2485, 1741, 1795, 2783, 1313, 194, 222, 207, 2560, 1251]. Relationships [1520]. Relativism [1477]. Relativita [960]. Relativitatstheorie [2254]. Relativity [394, 1648, 737, 1707, 1826]. Relevance [2308]. Relevant [1363]. Reliable [211]. religieux [1394, 578, 2129]. Religion [1200, 2539, 1520, 1544, 1221, 1968, 1546, 2308, 1886, 2278]. religione [1517]. Religiose [570]. Religious [2309, 2091, 1540, 1985, 2049]. relojes [1154]. Remains [177]. Remarks [668]. Rembrandt [1652]. Remedies [1272]. Remember [31, 1530, 188]. Reminiscences [1997, 2460, 515, 931]. Remke [504]. Remsen [199]. Rena [226]. renacimiento [2592]. Renaissance [91, 1854, 1853, 379, 1958, 1563, 2902, 918, 1957, 373, 1683, 1346, 1980, 107, 787, 2818, 2278, 2259, 123, 1114, 380, 1551]. renaissant [2127]. Renate [1401]. Renato [2402, 560]. Rene [1428, 2263, 2050, 1180, 2594]. Renee [2078]. Renn [2654]. Reorganized [1823]. Repertorio [1682]. Repertorium [1849]. Replication [2797]. Report [1083, 1131, 428]. Reporting [1820]. Reports [200, 65]. Representation [272, 2856]. Representing [1721]. Reprints [2676]. Reproduction [1324, 1439, 2867]. reproduzione [1188]. Republic [460, 1537, 140, 2623]. Republica [1613]. Republican [1444]. Republicans [1116]. République [2350]. rerum [2801, 375, 63]. Rescher [217, 354, 2948, 2865, 1723]. Research [344, 2786, 1624, 80, 1243, 1025, 627, 190, 210, 738, 2553, 658, 1500, 2087, 2486, 2471, 1923, 2107, 2133, 865, 85, 973, 1596, 635, 1033, 2806, 2381, 2106, 810, 1923, 481, 719, 155, 495, 1041, 489, 2209, 2366, 564, 1661, 759, 2323, 2760, 85, 1364]. Researches [1431]. Resistance [1940]. Resnik [357]. Resolution [2881, 2281]. Resonance [2472]. Resources [431, 95, 680, 620]. Respiration [1505]. respirationis [1505]. Response [364, 1279]. Responses [2020]. Responsibility [709, 2944]. Rest [538]. restitutae [1744]. Restivo [1313]. Restoration [2521, 398, 646, 2742]. Results [1571]. Retailer [2284]. Retaliation [528]. Rethinking [2520]. Retreat [1900]. Retrospect [323, 588]. Retrospective [367]. Returns [710]. Revelations [1701, 766]. Reverend [2146]. revidierte [1285]. Review [2570, 2456, 735, 2767, 1201, 981, 1475, 650, 2363, 2395, 276, 2725, 953, 216, 46, 91, 1981, 558, 1663, 222, 445, 854, 1353, 2439, 49, 834, 728, 946, 1759, 2260, 2062, 2075, 2249, 344, 2166, 2890, 289, 514, 2113, 29, 387, 857, 1863, 2198, 1785, 2737, 2824, 2836, 1827, 1969, 47, 569, 1291, 2567, 152, 2914, 1252, 1481, 1493, 1196, 1286, 1809, 2239, 135, 872, 1056, 1811, 2839, 274, 888, 431, 381, 1282, 294, 1408, 1646, 2238, 1859, 1905, 2726, 1148, 2477, 2798, 15, 129, 2799, 2035, 76, 2455, 2497, 34, 79, 1371, 2649, 2771, 2848, 11, 1972]. Review [2076, 2635, 474, 1548, 1398, 1360, 629, 1559, 1327, 1987, 347, 468, 786, 1154, 1253, 1933, 2406, 2723, 81, 1756, 2920, 726, 1662, 1710, 2762, 555, 660, 2077, 2677, 632, 2284, 2350, 2880, 1208, 2447, 499, 848, 975, 1278, 1672, 1844, 2891, 271, 1186, 281, 132, 154, 886, 1160, 1289, 1583, 1055, 677, 1403, 1950, 1187, 2309, 1195, 492, 2499, 246, 2563, 1423, 2289, 2574, 1050, 776, 1428, 2244, 321, 1780, 2605, 169, 766, 1015, 1118, 1448, 1608, 2187, 2322, 2437, 2047, 2275, 1249, 185, 2398, 684, 1199, 2597, 2625, 2752, 2227, 2854, 2329, 337, 965]. Review [1111, 1427, 1665, 703, 136, 736, 1032, 2703, 22, 2630, 903, 219, 831, 2016, 74, 273, 1255, 1415, 1612, 1656, 1893, 1954, 2246, 2631, 386, 1231, 1605, 2147, 2834, 2203, 2626, 2267, 2634, 2888, 288, 336, 610, 1575, 1692, 2263, 1162, 2073, 656, 221, 481, 2223, 2937, 1260, 2862, 2929, 501, 1343, 51, 94, 611, 887, 2556, 444, 861, 48, 291, 559, 1048, 1105, 2240, 2822, 225, 819, 1273, 1460, 1897, 2915, 2194, 1652, 227, 2110, 2727, 2276, 2695, 2100, 636, 1285, 1745, 2738, 2826, 2913, 1765, 604, 2579, 603, 2676, 2021, 1657, 2564, 2326, 557]. Review [843, 929, 1362, 1928, 2660, 2942, 2810, 80, 1262, 1200, 653, 2097, 1166, 477, 1147, 2431, 1424, 1314, 697, 1429, 1746, 2685, 1039, 860, 377, 2138, 969, 2453, 1237, 511, 212, 757, 836, 2286, 2769, 2663, 2192, 779, 1508, 1673, 614, 2573, 1495, 240, 1838, 1915, 184, 408, 688, 1488, 1604, 1949, 31, 14, 719, 1631, 1329, 23, 155, 495, 2269, 2354, 730, 2041, 370, 796, 930, 1367, 2443, 980, 2638, 2532, 380, 639, 845, 2531, 366, 112, 2748, 2359, 237, 84, 896, 1944, 2042, 2121, 1243, 2667, 2650, 1229, 119, 2679, 56, 676, 1788, 2739]. Review [2910, 567, 2208, 899, 1567, 1607, 1946, 2569, 2842, 2311, 2593, 2736, 2175, 2534, 2270, 683, 1511, 272, 1185, 1722, 2013, 2312, 1177, 1307, 1847, 631, 1491, 1639, 1840, 1328, 204, 958, 1302, 2096, 2467, 2815, 583, 1864, 1959, 2248, 2530, 2566, 1182, 1023, 1487, 1658, 2661, 1025, 1425, 2452, 2688, 1837, 113, 304, 777, 1505, 2327, 627, 1484, 1947, 2335, 2916, 1345, 2766, 1102, 1441, 2245, 2451, 2700, 88, 1786, 1043, 1782, 1854, 2373, 16, 466, 911, 1916, 2706, 2399, 2555, 1194, 1290, 1203, 1779, 1853, 30, 224, 1159, 190, 395, 797, 1413, 1924, 2082, 2224, 2360, 2568]. Review [2017, 435, 470, 720, 820, 1041, 1603, 1984, 2884, 391, 1594, 1857, 2115, 2201, 2440, 93, 163, 540, 1236, 2216, 1517, 2595, 2820, 2811, 489, 1103, 1354, 2402, 2307, 1614, 44, 123, 508, 2298, 2692, 2775, 2266, 2644, 210, 586, 2304, 379, 516, 762, 864, 1104, 1149, 1321, 2027, 2186, 1402, 620, 1095, 469, 1096, 2091, 1918, 1271, 434, 643, 1275, 1407, 1551, 1817, 1958, 2696, 2801, 424, 1156, 1303, 1435, 1541, 1774, 2733, 883, 1540, 2881, 738, 952, 1115, 471, 702, 767, 1165, 1266, 1580, 1833, 2362, 378, 1205, 217, 2374, 2817, 2950, 428, 607, 950, 1092, 2131]. Review [2176, 2553, 2763, 2949, 1597, 1116, 1401, 298, 1254, 2527, 1242, 1974, 96, 704, 1121, 361, 859, 2145, 2376, 2353, 1819, 161, 315, 906, 1049, 32, 1164, 2259, 425, 1418, 2375, 550, 1333, 1556, 2112, 2904, 510, 1563, 1644, 2525, 2657, 78, 2892, 1971, 1518, 1144, 243, 915, 77, 270, 355, 409, 575, 608, 1189, 1543, 1685, 1784, 2219, 2281, 2598, 2664, 2109, 2299, 2487, 2645, 2159, 2724, 1985, 203, 1973, 2185, 2510, 2020, 2774, 2012, 2123, 1184, 2866, 2167, 322, 2180, 1943, 2637, 125, 478, 1636, 2445, 482, 483, 1094, 1775, 2782, 1331]. Review [934, 1925, 2099, 914, 1170, 2458, 168, 2773, 134, 362, 700, 2498, 2368, 2209, 500, 1906, 98, 2922, 2026, 2253, 1334, 1100, 1634, 1721, 2468, 50, 423, 732, 2050, 1613, 1155, 982, 1117, 1752, 266, 1004, 1206, 1335, 1668, 1320, 1849, 2819, 53, 768, 476, 1014, 1396, 2032, 2412, 191, 245, 2585, 1240, 1764, 2184, 302, 1355, 2257, 2699, 290, 1358, 626, 718, 1344, 1686, 1757, 1909, 2409, 658, 12, 354, 467, 1267, 1829, 2183, 166, 505, 814, 1332, 1192, 1503, 1553, 1781, 1860, 1882, 2048, 2592, 2751, 1588, 2590, 2898, 186, 731, 986, 2105, 394]. Review [1193, 2092, 2658, 2647, 2642, 116, 439, 641, 681, 708, 1432, 24, 922, 2371, 2476, 2883, 2182, 2833, 153, 1522, 1653, 1887, 2218, 793, 1648, 760, 2200, 45, 553, 2394, 189, 692, 1411, 1691, 2496, 654, 115, 780, 852, 1114, 1778, 282, 507, 623, 693, 1485, 2672, 2777, 825, 1890, 2694, 2882, 1479, 487, 988, 1660, 2028, 2148, 75, 247, 385, 1222, 1633, 1789, 1990, 2055, 2285, 2355, 2704, 1272, 1842, 389, 2349, 165, 414, 679, 2125, 1442, 2480, 2821, 427, 2561, 2768, 1146, 1713, 1960, 313, 649, 961, 1259, 1514, 2340]. Review [2030, 2135, 2680, 2691, 71, 813, 2283, 207, 278, 580, 678, 971, 1089, 1158, 1235, 1377, 1394, 1735, 1886, 2045, 2120, 2160, 2372, 52, 1022, 2222, 1108, 1172, 1970, 601, 1983, 2446, 2730, 131, 369, 1139, 67, 1152, 1667, 2009, 2901, 714, 756, 1471, 1675, 2558, 156, 1967, 2225, 213, 1019, 1638, 318, 2137, 2689, 1520, 1997, 2656, 150, 1098, 1681, 1726, 2237, 2492, 698, 546, 2606, 1550, 515, 1161, 1867, 2293, 241, 117, 309, 376, 774, 1935, 1347, 2698, 2907, 2835, 2683, 2521, 342, 1724, 832, 1316, 1574, 2709, 1469, 823, 2838, 2488, 2867, 218, 345, 2678, 2640]. Review [1766, 120, 2025, 2095, 2816, 2196, 2279, 1690, 183, 682, 1020, 578, 705, 1173, 1693, 1791, 2442, 2840, 1647, 2925, 2339, 485, 2143, 420, 957, 2857, 248, 127, 148, 2294, 2655, 582, 931, 2059, 770, 541, 948, 429, 277, 382, 1288, 1720, 2254, 966, 2632, 2849, 2776, 577, 655, 1018, 889, 970, 1112, 1566, 2063, 208, 821, 1730, 1834, 2500, 2599, 2701, 2917, 782, 585, 2902, 897, 432, 2861, 556, 1363, 1530, 1728, 1888, 1241, 2449, 1011, 2040, 2582, 149, 509, 725, 1248, 1444, 1855, 1986, 2052, 226, 2936, 220, 1659, 2051, 1500, 918, 2552, 842, 502, 108, 2546]. Review [1021, 2830, 2863, 146, 491, 1167, 1250, 1406, 1426, 1521, 1571, 1841, 2043, 2136, 2485, 2684, 2697, 2845, 1911, 602, 2846, 1684, 2084, 375, 1547, 1097, 1812, 2927, 151, 1810, 2189, 640, 1438, 233, 2274, 1277, 1851, 2896, 2928, 1965, 106, 144, 393, 1245, 2139, 1585, 1610, 2215, 2562, 858, 2588, 223, 838, 1995, 101, 547, 1595, 1978, 2627, 2813, 1359, 1761, 2682, 122, 574, 904, 1040, 1417, 1577, 89, 324, 357, 1716, 1138, 1602, 484, 1771, 2507, 2330, 215, 1447, 1465, 2479, 2600, 1379, 178, 234, 305, 352, 2140, 2832, 2930, 2665, 817, 2087, 2181, 18, 214, 390]. Review [1399, 1562, 1865, 2741, 2545, 696, 829, 2080, 2827, 1717, 837, 924, 17, 503, 2459, 371, 968, 1369, 1671, 2753, 616, 2859, 1523, 1884, 2057, 180, 1814, 1012, 868, 990, 295, 364, 928, 1301, 1951, 2089, 2486, 2577, 2812, 2410, 1322, 2104, 1497, 2629, 2923, 2297, 2803, 1436, 27, 717, 1512, 2756, 104, 1421, 1378, 1674, 1549, 2348, 771, 1336, 1907, 100, 1106, 1047, 1169, 1395, 1750, 1370, 2860, 824, 910, 984, 1044, 1246, 1284, 2262, 2308, 2401, 2576, 1477, 2343, 167, 764, 2948, 2303, 2471, 2643, 2843, 2885, 35, 176, 231, 368, 416, 956, 1923]. Review [624, 816, 2853, 709, 1767, 2301, 2190, 544, 974, 1269, 475, 706, 862, 892, 987, 1306, 2132, 2199, 1305, 2418, 2557, 723, 1920, 2247, 2366, 1034, 70, 2869, 2841, 2903, 1820, 2022, 162, 1381, 1670, 2088, 2117, 2172, 118, 174, 587, 908, 1966, 2403, 1957, 307, 373, 978, 1552, 2242, 979, 1437, 2213, 2584, 2894, 2565, 638, 2417, 707, 2292, 2107, 778, 847, 1405, 1013, 2808, 1807, 2134, 2226, 2337, 2601, 2938, 1366, 1400, 1528, 2268, 1499, 1560, 2164, 2054, 2178, 2529, 2780, 1420, 1734, 2083, 2306, 1490, 87, 564, 2652, 1712, 1719, 750, 1741, 659, 2708, 1599]. Review [2206, 159, 269, 2681, 37, 283, 372, 1136, 1635, 1743, 1979, 2038, 2128, 2515, 346, 701, 2864, 1679, 316, 545, 1832, 2163, 95, 833, 1010, 1330, 1524, 840, 2033, 2161, 2204, 2648, 1211, 2029, 2149, 2687, 2940, 1107, 2195, 228, 565, 1207, 1975, 2365, 2441, 2728, 494, 972, 1463, 2517, 729, 763, 1026, 1268, 1836, 1210, 513, 1714, 1502, 21, 312, 615, 1042, 1150, 1632, 2282, 2429, 2484, 572, 977, 1038, 2258, 710, 2935, 384, 2346, 1365, 1934, 2749, 835, 1527, 2228, 2837, 1430, 1472, 103, 1227, 1501, 1006, 1137, 1349, 2085, 1903, 1037, 1232]. Review [1850, 1611, 452, 300, 621, 1168, 1225, 2177, 2720, 737, 358, 625, 989, 1751, 1830, 1881, 2202, 2056, 2173, 2646, 2887, 2133, 2220, 128, 1516, 1813, 687, 548, 613, 865, 1009, 2205, 605, 1016, 1409, 1473, 1593, 1642, 1731, 1988, 2019, 2086, 2405, 2528, 2744, 55, 1467, 2511, 2690, 542, 1919, 2174, 716, 69, 179, 1087, 1461, 2277, 885, 1052, 1324, 1578, 2074, 2191, 2502, 339, 617, 1544, 1628, 2094, 2494, 275, 612, 761, 985, 1280, 1380, 1600, 1908, 2210, 2516, 2856, 2886, 1557, 2755, 2393, 121, 2221, 1310, 1669, 411, 822, 2014, 2241, 2341]. Review [2413, 1093, 1926, 2490, 2705, 2118, 280, 1325, 2945, 1921, 54, 205, 319, 657, 1293, 1826, 775, 1175, 962, 479, 157, 109, 473, 549, 606, 724, 818, 884, 1088, 1815, 2101, 2370, 2719, 2058, 2745, 1513, 1601, 2347, 2251, 2380, 1058, 83, 2602, 846, 721, 1964, 785, 2392, 2427, 1725, 2878, 1645, 1773, 33, 1483, 2475, 1468, 2419, 600, 1732, 2897, 1264, 160, 365, 685, 1054, 1539, 1641, 1737, 2034, 2855, 317, 584, 690, 1003, 1397, 1476, 1900, 1912, 2352, 2551, 2761, 2411, 170, 1238, 1519, 2310, 2300, 1805, 2919, 686, 1564, 2345, 341, 827, 1017, 1119, 1650, 229]. Review [648, 1179, 417, 279, 1683, 1787, 2519, 2805, 338, 983, 1180, 1606, 1910, 1982, 2053, 2130, 360, 426, 1494, 1576, 2168, 1326, 1823, 2404, 177, 1582, 85, 947, 1510, 1637, 2214, 2899, 2272, 1682, 211, 789, 1283, 1598, 2589, 2291, 2603, 2852, 1883, 758, 2754, 2850, 912, 2505, 2575, 781, 1030, 2433, 2673, 1209, 1843, 86, 895, 1005, 1090, 1261, 1410, 1754, 2416, 2503, 1742, 2169, 2478, 2765, 25, 438, 486, 1664, 1474, 296, 2628, 1287, 1376, 2604, 1342, 2124, 2369, 175, 2514, 618, 1470, 2591, 2893, 72, 1654, 2157, 2290, 2470, 1736, 2036, 2426, 2207]. Review [1589, 973, 2423, 310, 784, 917, 1592, 2734, 137, 2559, 2103, 2560, 2804, 2325, 441, 447, 488, 579, 788, 925, 951, 1770, 1818, 1858, 1862, 1901, 1913, 1993, 2015, 2037, 2146, 2288, 2324, 2361, 2364, 2434, 2721, 2781, 2829, 2947, 815, 561, 2377, 2858, 1274, 1506, 932, 2671, 2828, 2900, 418, 913, 1113, 1361, 826, 959, 1558, 1885, 20, 133, 451, 652, 711, 867, 1609, 1868, 1917, 2333, 2422, 2548, 2654, 512, 1053, 1265, 2659, 2264, 285, 870, 2924, 105, 1439, 2408, 2501, 299, 1496, 493, 689, 1368, 1565, 1555, 1866, 1561, 1896, 2444, 2547, 2932, 437]. Review [754, 907, 1109, 1348, 1651, 1760, 1956, 2010, 2102, 2400, 2550, 2524, 1188, 1643, 1790, 1263, 2668, 543, 869, 1311, 1629, 1738, 2457, 356, 622, 694, 2078, 1223, 1374, 1586, 1929, 1233, 2378, 1718, 99, 552, 562, 727, 926, 954, 2142, 2879, 374, 1678, 2586, 1744, 2060, 2472, 1002, 2757, 566, 1029, 1584, 1902, 2179, 2495, 2921, 734, 1504, 2158, 40, 642, 1279, 287, 1153, 1356, 1689, 1755, 1861, 2702, 363, 560, 651, 945, 130, 2931, 1292, 1372, 1649, 1927, 2607, 2851, 2847, 2823, 232, 449, 68, 633, 1270, 1666, 1739, 2328, 2746, 2271, 2939, 267, 634]. Review [2332, 230, 894, 2114, 2302, 268, 765, 1191, 1661, 1835, 320, 1142, 2243, 2435, 2764, 2944, 124, 158, 1350, 1443, 1568, 2305, 2430, 2636, 2783, 1824, 2357, 2432, 2489, 2674, 1431, 1190, 2747, 2750, 413, 674, 759, 927, 2396, 964, 755, 955, 1715, 308, 752, 1318, 1776, 1848, 2391, 2587, 2814, 2895, 1422, 1763, 1554, 1323, 2448, 412, 2594, 39, 238, 504, 976, 1204, 1276, 1440, 2865, 2344, 36, 110, 303, 497, 712, 1028, 1202, 1337, 1433, 1507, 1852, 1904, 1976, 2049, 2122, 2518, 2193, 2578, 855, 902, 1806]. Review [1822, 1024, 1729, 2481, 2778, 675, 472, 1143, 440, 570, 2831, 1312, 2520, 900, 2707, 551, 890, 1046, 795, 1163, 1486, 1723, 2420, 2797, 82, 187, 539, 791, 1414, 2460, 1478, 2796, 2670, 2735, 2508, 2580, 1313, 188, 517, 1309, 2338, 164, 905, 284, 2759, 1226, 1777, 1228, 13, 430, 646, 919, 1183, 1515, 2156, 2334, 1480, 2217, 2802, 2079, 42, 242, 311, 388, 506, 573, 856, 960, 1181, 1340, 1727, 1856, 2129, 2280, 2454, 2526, 1596, 415, 480, 1151, 1464, 2424, 1992, 844, 1346, 2108, 2323, 1086, 422, 1145, 1412, 2941, 1687, 773, 644]. Review [849, 1178, 1339, 1846, 181, 571, 783, 1816, 2450, 1101, 1569, 2119, 292, 863, 2572, 1570, 340, 790, 102, 301, 314, 635, 733, 769, 1033, 1258, 1364, 1587, 1630, 1749, 1922, 1931, 2024, 2666, 1352, 2686, 172, 293, 563, 1449, 2090, 2732, 2806, 26, 1808, 2675, 1948, 792, 1051, 2653, 433, 2428, 1930, 2581, 2772, 920, 1482, 2800, 209, 673, 306, 1140, 2011, 126, 866, 923, 1579, 873, 1110, 2367, 2669, 1952, 630, 839, 1492, 1839, 1898, 2295, 2596, 2491, 2758, 2379, 2513, 2342, 446, 1007, 2081, 2825, 2407, 798, 1257, 1120, 1416, 1962, 2252, 2571, 2844]. Review [43, 92, 442, 2718, 2641, 2061, 2506, 244, 350, 419, 1057, 1317, 1772, 1319, 1466, 1733, 1573, 38, 235, 436, 850, 1977, 2126, 2171, 1341, 2044, 1968, 828, 1304, 367, 841, 2018, 2250, 359, 645, 1045, 1419, 2046, 2256, 2436, 1980, 2144, 1762, 1869, 680, 1230, 1740, 1899, 2554, 898, 1251, 1821, 1036, 2265, 2770, 1831, 851, 1035, 1375, 1590, 1769, 2162, 1581, 353, 637, 1308, 2116, 2512, 2633, 496, 1174, 1591, 2729, 967, 2809, 1758, 323, 830, 949, 1450, 1825, 1996, 568, 1373, 2381, 2106, 2504, 554, 349, 691, 749, 1008, 1446, 1546, 1880, 1894, 1994, 2093]. Review [2421, 2740, 2943, 1655, 236, 916, 1176, 1572, 1680, 1845, 2722, 2212, 2731, 715, 1091, 73, 206, 921, 1157, 1953, 1932, 1677, 443, 2908, 2111, 2127, 1694, 2197, 111, 909, 713, 421, 619, 891, 1027, 1224, 1462, 1753, 1914, 2523, 2549, 2909, 1247, 2889, 450, 1434, 107, 609, 787, 1542, 1783, 2742, 1338, 1315, 297, 348, 1197, 1445, 1768, 1889, 2098, 2356, 2583, 498, 2818, 1244, 1357, 2278, 2662, 2946, 772, 2639, 19, 871, 1234, 2023, 2493, 2926, 1525, 1509, 2483, 2651, 1239, 2170, 1545, 1640, 1351, 1711, 2906, 147, 2211, 2255, 2351, 576, 90]. Review [392, 1891, 1963, 2141, 2533, 2760, 1085, 2482, 538, 2415, 286, 173, 628, 893, 1198, 2414, 2273, 2509, 695, 1031, 1141, 1256, 1747, 1748, 1989, 2743, 2868, 114, 581, 901, 963, 1961, 1991, 2522, 647, 28, 1489, 383, 722, 1281, 1688, 991, 1404, 1828, 2039, 2358, 2918, 2934, 699, 2912, 490, 1945, 2031, 2287, 97, 1099, 2296, 351, 2779, 2425, 2397, 2933, 1529, 343, 2165, 2336, 1171, 2905, 239, 794, 1955, 588, 2261, 410, 41, 182, 751, 1526, 1892, 2469, 1676, 753, 171, 1498, 2188, 2331, 2438, 448, 2807, 1895, 933, 2693, 2717, 2911, 150]. Review [1101]. Reviews [2293]. Revisions [2419]. Revisited [1297, 277, 208, 2364]. revolucion [2025]. Revolution [11, 2929, 1604, 971, 1022, 595, 226, 2546, 2665, 2652, 21, 1006, 1456, 1380, 983, 789, 2157, 2158, 2046, 1027, 1224, 297, 2868, 270, 269, 33, 650, 336, 567, 2452, 1648, 389, 2698, 2652, 1600, 338, 2514, 2924, 1896, 1140, 794, 1377, 1486, 1711]. Revolutionary [1863, 1934, 1691]. Revolutionen [275]. revolutionibus [1909]. révolutionnaires [2350]. Revolutions [886, 270, 409, 1623, 2717, 545]. revolutsia [2903]. Revue [500, 1423, 624]. Reward [351]. Reweaving [1918]. Rexmond [18]. Reymond [2132, 987]. Reynaud [2407]. Reynolds [2567, 1169, 2242]. Rezeption [1906, 2819]. Rezeptionsgeschichte [2738]. Rheticus [1319]. Rhetoric [2016, 2564, 1453]. Rhetorical [508, 1215]. Rhinoceros [1231]. Rhinoceroses [1231]. Rhoda [2032]. Riccardo [336, 1791]. Rice [1025, 2449]. ricerca [2160]. Rich [2133]. Richard [1859, 2016, 48, 929, 683, 1487, 2327, 2224, 1435, 471, 2866, 1886, 2907, 1316, 2776, 970, 1444, 2936, 1245, 1497, 2923, 100, 1750, 2262, 2128, 835, 1813, 1087, 109, 2805, 2850, 1742, 2369, 2559, 1770, 1744, 2481, 1228, 1464, 1846, 866, 873, 2596, 2523, 410, 2403, 265, 929, 513]. Richards [2041, 2412, 2683, 1888]. Richardson [1286, 2869, 2645, 657]. Riché [237]. richesse [2699]. Richter [452, 353]. Riddle [2568, 769]. Rider [949, 1225, 1091]. Ridge [596]. Ridwan [1594, 376]. Rieber [631]. Riegert [628]. Riemann [1021, 2357].

Riforma [2597]. Rigden [2380]. Righini [1300]. Rights [2178]. rigore [849]. Rigorous [693]. Riley [2267]. Rima [1556, 2647]. Rinascimento [604]. Ring [912]. Ringer [22]. Rings [1412, 1750]. Rink [682]. Rio [1348]. Rise [136, 1656, 559, 2096, 298, 914, 2349, 774, 2812, 2860, 2885, 707, 1473, 1912, 869, 2931, 1296]. Risse [2036]. Rita [1847, 151, 235]. Ritter [289, 762, 864, 864]. Ritual [1194]. Ritvo [2750]. Rivalry [2833]. River [788]. Rivers [2331, 100]. Rizzo [606]. Road [88, 1022]. Roads [892, 2104]. Roaring [213]. Robert [276, 834, 274, 888, 1646, 1398, 1195, 136, 1105, 2913, 240, 845, 1639, 1959, 1025, 1290, 30, 1435, 96, 510, 2180, 914, 2458, 2050, 1553, 2590, 2028, 2355, 2821, 2030, 580, 1735, 2558, 2683, 218, 127, 1728, 149, 2052, 2552, 602, 1761, 1716, 1865, 2577, 2190, 544, 1670, 978, 1107, 228, 494, 1836, 1042, 2258, 384, 2837, 2220, 1593, 2405, 1467, 121, 1325, 1921, 2101, 2475, 584, 279, 2130, 1494, 85, 2654, 869, 622, 1689, 2851, 1763, 2344, 905, 1339, 1816, 306, 2342, 350, 2171, 1341, 1304, 359, 1980, 1994, 1783, 2651, 2211, 2273, 1828, 2031]. Robert [2397, 1529, 2621, 393, 807, 2544, 240, 98, 231, 2277, 2108]. Roberts [1864, 1676, 1510, 1689]. Robertson [2101, 1346, 392]. Robin [2123, 1420, 1225, 949, 1091]. Robins [2708]. Robinson [99, 904, 2117, 605, 99, 949]. Roca [2900]. Roccatagliata [1905]. Rocchietta [1332]. Roche [1319, 71, 2011]. Rochow [2358]. Rocke [1447]. Rockefeller [656, 770]. Rocketry [2269]. Rocks [1321]. Rocky [2145, 951]. Rodda [1163]. Roderick [1259, 2757]. Rodger [1180]. Roe [1930, 2402, 648]. Roemer [184]. Roger [2453, 797, 2446, 1047, 709, 1578, 2124, 1862, 2193, 340, 199, 323, 2421, 2722, 1404, 2438, 2453, 2232, 1262, 2453, 1177]. Rogers [1585, 1325, 1901, 832, 1820]. Rogge [1566]. Rohm [1922]. Roland [2381]. Role [266, 461, 1560, 404, 1927, 415, 1797, 619]. Rolf [2630, 19]. Rolle [2819]. Roller [1158]. Rom [2418, 1405]. romain [1506]. Roman [2012, 2445, 2368, 2446, 1288, 1674, 1976]. Romance [1372]. Romanell [1606]. Romano [1688]. Romans [897]. Romantic [1170, 2305]. Romantische [1029, 448]. Rome [2520, 2693]. Römer [36]. romischen [1470]. Romm [1766]. Ron [1613, 818]. Ronald [2839, 1199, 2556, 366, 2736, 767, 1266, 1411, 2558, 2682, 21, 1030, 2783, 2061, 898, 2093]. Ronan [118, 1226]. Ronayne [1478]. Rookmaaker [1231]. Roos [1464]. Roots [840, 1910, 2850, 1416, 2954]. Rosa [724]. Rosalind [840]. Roschlaub [1029]. Roscoe [294]. Rose [1644, 421, 1529, 2138, 105, 421, 1031]. Rosemarie [785]. Rosen [1981, 1427, 1709, 1600, 230, 73]. Rosenberg [1919, 2752, 1427, 1367, 1837, 840, 230, 297]. Rosenfeld [1120]. Rosenthal [2454]. Roseveare [1483]. Roshdi [1979, 2391, 2814]. Rosie [1448]. Rosińska [2816, 1466]. Rosmorduc [1162]. Ross [1644, 2032, 1322, 2602, 511, 1322]. Rossi [579]. Rossii [74]. Rossiter [629, 2478]. Rota [2673, 2567]. rotatoire [1162]. Rotberg [1553]. Rothenberg [731, 2057, 216, 1092, 2425]. Rothschild [1268, 1268, 1368]. Rothstein [2855]. Rouse [2477]. Rousseau [858, 2671, 2218]. Roving [561].

Rowan [1607, 539]. Rowland [640]. Roy [953, 2439, 2888, 212, 1842, 17, 1088, 2947, 413, 1143, 791, 645, 2742, 540, 784]. Royal [2062, 697, 1984, 2167, 922, 1098, 513, 605, 1731, 1093, 2118, 2036, 1902, 415, 1453, 2618, 461, 2429]. Royale [456, 446]. Royaume [2563, 2699]. Royaume-Uni [2563]. Royez [1592]. Roysdon [542]. Rubber [1926, 2705]. Rubinstein [1714]. Rud [1120]. Rudbeck [2401]. Rud– der [2909]. Rudiger [1808]. Rudolf [1196, 1811, 1847, 643, 345, 2808, 1650]. Rudrum [1787]. Rudwick [575]. Rufus [1433]. Ruggiero [506]. Ruhlen [2240]. Ruhr [2763]. Rules [1977]. Rumford [183]. Rumpf [859]. Rumphius [859]. Run [866]. Rupert [509, 1097, 2426, 1224]. Rupke [1419]. Rupp [1495]. Rupp-Eisenreich [1495]. Ruprecht [984]. Rural [893]. Ruscelli [1130]. Ruse [2398, 204, 389, 821, 1269, 954]. Rushton [2667]. Russel [386, 924, 1831]. Russell [2644, 517, 2666, 923, 2223, 2929, 1540, 1755, 2735, 1821]. Russia [1239, 1298]. Russian [2385, 2233, 1027]. russischen [426]. Russo [1090]. Rustum [540]. Ruth [2645, 1365, 1172]. Rutherford [423, 341, 1914, 1404, 555, 1450, 90]. Rutherford-Soddy [423]. Rutkoff [2323]. Ryan [1407, 2213, 2431]. Rydell [1865]. Ryden [1762].

S [650, 2260, 2239, 872, 2574, 2605, 74, 1954, 2564, 80, 2138, 511, 1949, 2208, 2569, 583, 1864, 2530, 1487, 1505, 2335, 395, 2298, 952, 702, 704, 575, 2253, 1344, 394, 1411, 487, 2055, 1377, 1886, 369, 2840, 248, 966, 2040, 602, 233, 858, 1417, 928, 1951, 1169, 167, 1923, 544, 2903, 2242, 2226, 2938, 2652, 1712, 228, 835, 1611, 1467, 205, 1293, 884, 2101, 2347, 2251, 2380, 2475, 1264, 1641, 2168, 177, 1637, 1209, 2765, 1736, 2325, 2671, 1439, 1263, 2931, 2823, 633, 759, 1852, 2122, 2420, 919, 1339, 1570, 293, 563, 2675, 126, 2407, 841]. S [2554, 967, 2731, 619, 2523, 1768, 1525, 695, 588, 751]. S.J [2884, 1783]. sa [2050, 1506]. Sabie [987]. Sabine [1566]. Sabloff [2428]. Sabra [1276]. Sachs [955]. Sacred [1266]. Sadek [1678]. Sadi [218, 585, 2220]. Sadoun [624]. Sadoun-Goupil [624]. Saeed [1277]. Safety [2759]. Saga [190, 2559, 33]. Sagan [2030]. Sage [2746, 449]. Sages [2078]. Saget [1175]. Saggio [2597, 1188]. Said [171]. Saint [2056]. Sakralgeometrie [310]. Sal [1313]. Salavert [2751]. Salernitan [374]. Saliba [1978]. Sally [1481, 2478, 2081]. Salman [1568]. Salmon [1724]. Salomon [446, 1001]. Salomon-Bayet [446]. Salor [1846]. Salt [67]. Salzinger [631]. Samaran [714]. Sambursky [2731, 2731]. Same [922]. Samhita [784]. Samion [1207]. Samion-Contet [1207]. Samoa [1171]. Sample [2242]. Samuel [2194, 1206, 278, 1855, 1026, 758, 2707, 2581, 2943, 1694, 1990]. Sanborn [183]. Sanchez [1613, 1846, 1767]. Sanchez-Ron [1613]. Sand [2348]. Sand-Glasses [2348]. Sander [2856]. Sanderson [2171]. Sandra [2748, 2417, 55, 644, 863]. Sands [1266]. SANE [2017]. sangue [112]. sanguis [1505]. sanitario [1765]. Sanitary [2289]. saniya [1440]. Sanskrit [642]. Santiago [861]. Santucci [1189]. Sapere [1000, 2597, 1685]. Sapp [2503]. Sarah [1487, 298, 885, 1338]. Sarasvati [2732]. Sargent [852].

Sarjeant [487, 472, 2554]. Sarkar [1726]. Sarton [1802, 2071, 1072, 1073, 1066, 1074, 1619, 2388, 598, 2544, 1618, 1075, 1065, 1071]. Sassonia [1675]. Satellite [243]. Satire [47]. Saturn [797, 1412]. Satya [2732]. Sauer [2605]. Saul [1668, 566]. Saussure [1034]. Savage [2376, 439, 2376]. Savage-Smith [439]. Savages [630]. savant [482]. savante [2657]. savants [2350, 2804]. Saving [767]. Saviours [2192]. Savoie [1589]. Savoir [1343, 2880, 480]. Saxe [1591]. Saxons [1329]. Scales [1220]. Scanlan [2695]. Scarborough [2328, 2047]. Scattering [2844]. Scene [137, 1371]. Scepticism [1725]. Schachter [1172]. Schaffer [1402, 1986]. Schallenberg [265, 970]. Schallmayer [1702, 2847]. Schatten [798]. Schatzberg [2783, 1251]. Scheele [1046]. Scheikundigen [649]. Schelling [862]. Schemnitzer [473]. Schickard [985]. Schiera [2601]. Schiller [222, 911, 318]. Schipperges [547]. Schlee [190]. Schmandt [2757]. Schmidt [727, 1980]. Schmiedebach [2504]. Schmitt [1379, 2126, 2008, 1407, 1563, 1114, 2213, 1683, 787, 866]. Schmitz [643, 345]. Schneer [360]. Schneider [2363, 499, 827]. Schnelle [1306]. Schnyder [479]. Schofield [384, 1174]. Scholarly [2479]. Scholars [1647, 2584]. Scholarship [2564, 2671, 2783]. Scholastic [1541, 1779]. Scholasticism [2894]. Scholz [1021]. Schönburg [2740]. School [2246, 767, 13, 2323, 1419]. Schools [1236, 2855, 2584]. Schoonheim [117]. Schöpfung [309]. Schove [1827]. Schrift [1433]. Schriften [1847, 1776]. Schrifttum [36, 836]. Schrifttums [116, 641]. Schroder [764]. Schrödinger [995, 2459]. Schroeder [86]. Schroeder-Gudehus [86]. Schroter [1908]. Schufle [2054]. Schule [706, 570]. Schuler [279]. Schulmann [2654]. Schulz [1957]. Schumacher [385]. Schupbach [1652]. Schuster [2016]. Schutt [1830]. Schwartz [1365, 2673, 1240, 725]. Schwarzbach [2149, 427]. Schwinger [2658]. Science [1794, 216, 2792, 1863, 1785, 47, 1481, 2238, 198, 15, 629, 468, 2920, 1249, 455, 610, 887, 2556, 559, 1460, 2915, 2913, 377, 2138, 212, 779, 688, 1604, 2538, 155, 495, 1243, 676, 1946, 683, 2013, 2070, 2530, 1182, 1025, 88, 1043, 2399, 224, 64, 540, 1103, 123, 2304, 1096, 1817, 1303, 2881, 217, 1388, 950, 1092, 1083, 2949, 1242, 2353, 161, 1049, 1418, 1644, 2525, 1985, 746, 322, 994, 2099, 2412, 2539, 1078, 1240, 2257, 670, 2048, 186, 681, 2618, 2833, 852, 1133, 679, 2480, 461, 678, 1089, 1886, 881, 1938, 1520, 2154, 2234, 2293, 2709, 2488]. Science [2816, 1076, 705, 957, 127, 663, 2632, 577, 2063, 1888, 62, 3, 2552, 398, 1167, 1097, 1533, 2927, 1457, 1138, 215, 352, 2140, 2385, 1454, 390, 1884, 180, 1012, 295, 2486, 2812, 2297, 2803, 256, 2308, 2471, 956, 1923, 1767, 807, 892, 1820, 118, 1966, 997, 1552, 2232, 2417, 2714, 1807, 2337, 1528, 1560, 2178, 1719, 750, 1741, 269, 2163, 95, 1211, 565, 1975, 494, 1042, 1632, 2935, 2346, 2085, 1850, 1751, 2541, 2019, 1087, 1544, 2516, 121, 411, 2241, 205, 145, 884, 252, 521, 802, 1062, 1385, 1698, 2001, 2316, 2611, 2613, 2878, 1645, 669, 685, 1221]. Science [584, 1912, 2761, 170, 2300, 2345, 407, 417, 947, 2673, 1742, 2478, 1474, 1537, 2068, 2591, 447, 1862, 2288, 2324, 2361, 932, 1885, 2005, 1392, 401, 1565, 1896, 907, 2524, 1738, 1718, 954, 2791, 374, 734, 2823, 940, 1739, 2537, 141, 268, 1077, 1835, 2783, 335, 593, 2396, 1621, 2448, 412, 238, 2390, 2067, 1729, 440, 1312, 140, 194, 1228, 882, 1515, 2334, 1464, 1992, 181, 2622, 1101, 1569, 4, 332, 340, 2387, 2090, 2806, 2653, 2011, 1797, 2295, 2506, 350, 419, 527, 1230, 1065, 1740, 1899, 1251, 1996, 197, 749, 2421, 2943, 1680, 1157, 1801]. Science [619, 1224, 2549, 609, 1244, 2662, 1298, 1234, 2023, 2493, 1389, 576, 810, 2743, 2868, 1281, 991, 97, 2336, 410, 1892, 1498, 677, 2563, 776, 1117, 83, 2392, 1410, 2829, 446, 270, 2236, 151, 2652, 269, 211, 2278, 2880, 2238, 2477, 1403, 1608, 1362, 2842, 2530, 1837, 2335, 2884, 78, 409, 608, 1784, 2458, 1721, 1155, 982, 1014, 12, 2883, 1713, 1960, 2340, 1022, 2143, 208, 1730, 1728, 1716, 2665, 17, 2022, 2164, 2529, 2306, 659, 2708, 545, 1463, 1714, 1502, 1227, 1225, 1409, 2945, 1088, 1815, 2719, 211, 1883, 758, 2369, 2288, 2548]. Science [2659, 1565, 1561, 869, 158, 1568, 755, 1715, 1723, 284, 2669, 2295, 2093, 787, 498, 1351, 1711, 576, 1688, 1404, 2912, 1099, 351, 343, 751, 2942, 2567, 2239, 274, 2798, 76, 79, 2937, 1944, 2910, 2569, 2402, 1402, 1095, 469, 478, 1155, 793, 414, 1713, 813, 2283, 2045, 601, 2446, 1667, 2237, 1097, 2588, 817, 2756, 1246, 2869, 2841, 638, 2938, 1560, 159, 1679, 494, 972, 1463, 2517, 1714, 1006, 1137, 1409, 1642, 2174, 69, 1087, 157, 2370, 1725, 686, 1910, 1823, 2589, 2754, 2671, 418, 2333, 689, 1689, 1661, 1337, 1478, 1309, 1226, 646, 415, 2424]. Science [2800, 1821, 1174, 1446, 1546, 73, 619, 1085, 239, 410, 2717]. Sciences [344, 2284, 273, 2564, 1229, 1103, 952, 478, 922, 2285, 226, 18, 307, 638, 1400, 1712, 833, 456, 2429, 1737, 85, 2671, 652, 1629, 2495, 642, 1337, 1143, 1486, 2420, 1939, 2732, 2513, 446, 1462, 2725, 2050, 2476, 690, 2804, 1131, 2767, 1663, 2077, 169, 1105, 2194, 2362, 1344, 2372, 1970, 167, 2418, 1813, 1467, 1005, 954, 363, 413, 2481, 900, 2421, 421, 2926, 1090, 651, 1397]. Scientia [241]. Scientiarumque [1882]. Scientific [735, 1940, 2284, 886, 1015, 2016, 1893, 1954, 336, 1746, 1604, 2354, 2875, 272, 204, 583, 2661, 2327, 508, 1551, 1215, 883, 1540, 217, 298, 510, 203, 1588, 1887, 1648, 189, 1158, 756, 156, 1098, 2698, 2069, 541, 382, 595, 1363, 1241, 602, 1623, 2927, 352, 1012, 1907, 2843, 2190, 8, 544, 1305, 2557, 1405, 1560, 87, 1210, 1632, 2346, 1472, 1137, 1881, 613, 885, 1600, 657, 2878, 317, 1299, 2310, 279, 338, 983, 1823, 789, 72, 784, 2560, 743, 2472, 1356, 2823, 68, 267, 268, 1350, 2430, 1130, 675, 2797, 13, 2802, 1727, 2424, 1816, 635]. Scientific [1630, 1140, 1579, 1952, 349, 1880, 2743, 2868, 1125, 1404, 1646, 886, 1950, 611, 1147, 2667, 1722, 2881, 476, 1146, 2656, 1724, 2948, 1866, 350, 1466, 2425]. scientifica [1326]. scientifici [703, 2160]. scientifico [2597]. Scientifique [46, 86, 480, 184, 2880, 2604]. Scientifiques [86, 2392, 1207, 690, 2804]. Scientism [2577, 350]. Scientist [1408, 2076, 1262, 738, 532, 1121, 2375, 731, 1691, 1080, 214, 70, 2282, 384, 2380, 1191, 792, 619, 1783, 321, 227, 2055, 1557, 450]. Scientists [1282, 607, 2294, 541, 2140, 987, 835, 687, 1731, 1093, 2078, 2802, 1086, 1816, 630, 885, 1314, 1233, 2491, 56, 1820, 1741]. Scienza [604, 2601, 77, 1517, 1189, 1890, 1609]. scienze [355, 644]. Scienziati [1643]. scienziato [42]. scolastico [311]. scolastico-cartésien [311]. Scope [1018]. Score [2226]. Scotland [1236, 2036, 1263, 1374]. Scott [2942, 233, 1879, 2508, 2323]. Scottish [1859, 2216, 369, 919, 43]. Scripture [1249]. Scrutinizing [2878]. Scrutiny [1463]. Scull [894, 102]. scuola [1009]. Scuole [1188]. Scurvy [1972]. Sea [1024, 1378]. Seaborg [872]. Seal [1210]. Sean [1607]. Search [2042, 2121, 666, 395, 361, 678, 1018, 2172, 1461, 2411, 1455, 2324, 1561, 2078, 2438]. Searle [2917]. Sears [699, 2276]. Seas [1024]. Seaton [2407]. Seck [985]. secolo [2051]. Second [336, 2205, 2157, 2435, 2868, 291]. seconde [624]. Secrecy [2768]. Secret [1056, 1407, 1518, 1272, 1747, 1857, 1272]. Secretarius [1040]. secreto [1868]. Secrets [1407, 1585]. Sections [600]. Secular [1165]. Secularization [1033]. Sedgwick [1190]. Seduction [1422]. Seed [2724]. Seek [2062]. Seeking [2889]. Seely [2510]. Seer [2306]. Segal [1649]. Segelns [1655]. Segni [2801, 579]. Segonds [1857, 1319, 1036]. Segré [2499, 422]. Segreta [1130]. Segunda [1613]. seicento [488]. Seienden [503]. Seiki [1227]. Seiler [1949]. sein [475]. Seine [862, 497, 475]. seines [241, 1288, 479]. seit [82]. Sekhmet [1673]. Sekido [1997]. Selbstreflexion [764]. Selbstverstandlichkeiten [292]. Selbstverstandnis [1019]. Seld [1439]. Select [2012]. Selected [1759, 1448, 2676, 843, 2531, 2650, 1817, 859, 852, 678, 1158, 1638, 948, 2741, 1884, 2033, 1736, 1318, 2943, 580]. Selection [1759, 1355, 2927, 927, 2395]. Selections [1809, 48]. Selective [1092, 1636, 2475, 1737]. Self [423, 2629, 1328]. Self-Portrait [1328]. Self-Splitting [423]. Sellers [390]. Selling [1670, 2297]. selon [726, 848, 438]. Semeiotic [2374]. Semenov [891]. Semiconductor [33]. Seminaire [958]. Seminal [2180]. Seminary [681]. Semitico [504]. Semitico-Latinus [504]. Semmelweis [1208]. Sensation [798]. Sense [204, 1128, 151, 2351]. Sensibility [2303, 1252]. sensu [1675]. senza [372]. Seoane [1765]. Separate [840]. Separatism [1425]. Sepper [2915]. September [1538, 465, 748, 464, 672, 1942, 513]. septembre [304, 507]. Sepulkralfunden [1470]. Serge [2170]. Sergio [1543, 1332]. Serial [1639]. Series [1893, 1983, 1465, 642]. Seriously [2398]. Serket [1673]. Serres [500]. Service [768, 2102, 1024, 1898]. servitoris [2586]. Servus [1435]. ses [615]. Sesiano [2391]. Set [2453, 484]. Seton [126]. settecento [1517]. settembre [1188]. Settings [2484]. seu [1744]. Seven [1121, 1437, 2193]. Seventeenth [15, 2638, 2467, 2451, 1156, 1541, 2131, 1985, 1520, 2052, 312, 2310, 1858, 1350, 2539, 2321, 2464]. Seventeenth-Century [2638, 2467, 1985, 1520, 2052, 312, 2539, 2321, 2464]. Seventh [802]. Seventieth [2212]. Seventy [678, 2231, 1151]. Seventy-Five [2231]. Seventy-Two [678]. Severin [977]. Sewall [2394]. Sex [1214, 365]. Sexist [954]. Sexl [1288, 133]. Sexton [1884]. Sexual [1324]. Sexuality [2487, 8]. Seymour [495, 2026, 2253, 211]. Sezgin [116, 641]. Shadow [1424, 1858]. Shafer [1306]. Shaky [2097]. Shaler [2529]. Shall [2899]. Shandean [1735]. Shank [2899]. Shape [2670]. Shaped [1429]. Shapere [1561]. Shapin [1986]. Shaping [1195, 702, 2557, 2325, 172, 1880, 2883]. Shapiro [510, 1733, 2030, 1520]. Sharaf [2814]. Shark [1416, 2954]. Sharlin [50, 50]. Sharon [1833, 1766, 1978, 1820]. Sharp [1819]. Shaw [2528, 2364, 1374]. Shea [1656, 1235, 90]. Sheaffer [2211]. Sheehan [912]. Sheets [1321]. Sheffield [2511]. Sheila [246, 2847]. Shelby [1486]. Sheldon [2020, 1058, 1922]. Shell [2260]. Sheller [1809]. Shelley [2636]. Shen [571]. sheng [571]. Shepherd [1842]. Sher [1859]. Sherrington [244]. Shia [1227]. Shield [47, 2713]. Shift [165]. Shifting [2307]. shih [571]. shih-yu [571]. shihyu [571]. Shinn [1742, 480]. Ship [2688]. Shipbuilding [1370]. Shipley [1662]. Ships [1322]. Shirakawa [1599]. Shirley [2402, 648, 1347]. Shiu [1205]. Shiu-ying [1205]. Shiv [2806]. Shizhen [2363]. shkola [1753]. Shop [2471, 2843]. Shore [1421]. Short [2027, 702, 382, 1097, 101, 1465, 1113, 2501, 2098]. Shorter [118]. Shortt [967]. Shoulders [1735]. Show [1225]. Showman [2284]. Shown [720]. Shrank [1049]. Shumaker [41]. Shuntaro [1176]. Shurkin [1428]. Shuttleworth [1481]. Sibley [2186]. Sica [2828]. Sicherman [1614]. Sicht [2696]. Side [2028, 319]. Sidney [2680, 2741, 2033, 818]. Siècle [911, 853, 237, 71, 1602, 979, 651, 291, 2810, 237, 1788, 315, 714, 438, 2814, 1375, 625]. siècles [975, 1342, 2804, 2170, 1495]. Siege [2120]. Siegfried [359]. Siena [2120]. Siglo [1860, 1781, 2592, 2751]. siglos [1782, 443]. Sigma [2138]. Sigmund [2857, 1666]. Sign [1445]. Signatura [2801]. Significance [1087, 1785]. Signs [213]. Sigrun [1945]. Silicon [2134, 2226, 2342, 2358]. Silicones [2358]. Silliman [186]. Silver [599]. Silverstein [1211]. Silvestro [1845]. Silvio [1445]. Simcock [1644]. Simkin [1487]. Simmel [16]. Simmons [1206]. Simon [569, 1402, 922, 2120, 2656, 2709, 1986, 1047, 1822, 185, 358]. Simonton [1634]. Simple [1450, 2408]. simples [1676]. Simpson [2408, 1401, 1287, 1581]. Simulation [2245]. Sinac [308]. since [952, 2109, 2632, 1112, 1841, 2887, 2919]. Sinclair [416, 2842, 2687]. Singer [972]. Sinnstrukturen [1094]. Sir [2223, 858, 1992, 359, 264, 1987, 1512, 539]. Siraisi [781, 2818]. Siry [1421]. sistema [1765, 644]. Sisyadhivrddhida [917]. Site [1529]. Situation [2854]. Six [548, 2356, 2571, 2131]. Sixteenth [2467, 1854, 1115, 425, 2120, 299, 1130]. Sixteenth- [2467]. Sixteenth-Century [1115, 2120, 1130]. Sixth [2269, 521]. Size [256, 778]. Skeptic [2030]. Sketch [627, 925]. Sketches [828, 1809]. Skidi [1038]. Skinner [49, 511, 49, 919, 172, 1449]. Skizzen [271, 452]. Skrupskelis [2209]. Sky [486]. Skylab [1772]. Skywatchers [572]. Slack [1903]. Slater [2173, 2342]. Slaughter [1350]. Slave [1843]. Slide [1556]. Sloan [857, 1051]. Slobodin [100]. Smallpox [1333, 769]. Smart [2098]. Smelser [1899]. Smelting [1599]. Smith [511, 439, 678, 393, 564, 1599, 2163, 1921, 788, 2146, 2654, 1808, 2908, 569, 2937, 2748, 113, 2555, 1254, 1971, 302, 439, 678, 2196, 2861, 838, 709, 1293, 1261, 296, 959, 755, 1808, 1930]. Smithsonian [1399, 1562, 2241, 630, 2224]. Smollett [2303]. Smuts [2107]. Snelders [575, 649, 2058]. Snjezana [2739]. sobre [2593, 442]. Social [1663, 2848, 1015, 481, 2194, 212, 2041, 1243, 2311, 883, 550, 2109, 1004, 2032, 2647, 971, 2222, 156, 1967, 127, 2849, 2582, 226, 3, 1521, 1841, 2189, 352, 2410, 1741, 269, 2646, 1813, 1409, 1280, 1058, 2034, 2352, 170, 1664, 2560, 2324, 2721, 2671, 2828, 652, 1223, 734, 2939, 894, 2481, 675, 791, 1313, 1596, 844, 2323, 2119, 102, 314, 433, 350, 1899, 1821, 2421, 2943, 713, 297, 1239, 173, 97, 2336, 751, 480, 550, 2020, 127, 2486, 1564, 827]. sociale [551]. Socialist [1298]. Socialists [87]. sociedad [1917, 443]. società [604, 355]. Societat [1192]. société [615, 223]. sociétés [2804, 551]. Societies [822, 1823, 2449]. Society [946, 2239, 2035, 76, 2227, 22, 2888, 2862, 477, 697, 2538, 896, 2555, 1984, 1271, 2487, 1155, 115, 1272, 1139, 1098, 2293, 1112, 226, 2923, 104, 2643, 416, 1741, 1010, 2517, 513, 1150, 605, 1731, 2086, 2516, 1310, 657, 1221, 2899, 1005, 2765, 1474, 2514, 2560, 1818, 940, 1130, 1822, 2420, 1309, 646, 2428, 2718, 2770, 967, 909, 2792, 2070, 1453, 1159, 64, 2618, 1133, 461, 881, 1457, 1751, 669, 2791, 332, 2387, 1801, 2798, 2263, 2013, 2115, 540, 1103, 1095, 2866, 2045, 213, 2629, 1010, 2161, 2940, 494, 2781, 890, 13, 619, 2414, 370, 2546]. Sociobiological [759]. Sociobiology [204, 1834, 204, 1967]. Sociological [1408, 1646, 146, 1107, 1891, 2779]. Sociologist [98]. Sociology [941, 294, 2329, 155, 1155, 2412, 2480, 150, 1888, 491, 2195, 157, 85, 2684, 2812, 1229, 1241, 2577]. Soddy [423, 1962]. Soderqvist [2192]. Soemmering [1990]. Soffer [226]. Sofia [1691]. Software [1428]. Soil [2749, 2509]. Sokal [819, 2770, 490]. Solar [1413, 1800]. soleil [1592]. Solid [554]. solides [2594]. solidorum [2594]. Solitude [2746]. Solla [2238]. Solomon [1949, 1959, 1202]. Solovay [1959]. Solov’ev [2832]. Soloway [1245]. Solstice [1708]. Solutions [2411]. Solvay [1415]. Solvers [2378]. Solving [94, 1977]. Som [824]. Some [668, 1418, 1417, 8, 1581, 1193, 1885]. Sommer [2031]. son [906, 2050, 1506]. Sonar [2062]. Songs [2304]. Sonnet [2170]. Sonntag [1354]. Sons [793, 1651]. Sophie [793, 692]. Sophismata [1977]. Sopka [2100, 517]. Sorabji [1307, 2369]. Sorbom [1309, 433]. Sorell [2279]. Soul [1303, 1855, 375, 1033, 1774]. Sound [2562, 1455]. Source [288, 1786, 1241, 991]. Sourcebook [2298]. Sources [1407, 1115, 428, 330, 2446, 2869, 2161, 205, 685, 1218, 2800, 1466, 1230, 406, 1082, 1688, 2933, 2259, 780, 542, 1197]. South [2375, 2558, 1265, 1808, 1581, 770, 833, 2648]. Southeast [378]. Southern [81, 2834, 1166, 2277]. Southgate [2529]. Southwestern [1494]. Soviet [246, 2842, 990, 104, 1366, 1030, 619, 2934, 1623, 10, 61]. sowjetischen [426]. Sowjetunion [426]. sozialen [908]. Sozialismus [1789]. Sozialphilosophie [1469]. Sozialwissenschaften [2946]. sozialwissenschaftlicher [2946]. Sozio [1908]. Sozio- [1908]. Soziologie [2330, 908]. Space [888, 2658, 535, 220, 2845, 2139, 2292, 659, 2708, 1292, 124, 1140, 1057, 1772, 1247, 2358, 2945, 2666]. Space-Perception [2945]. Spacecraft [248]. Spaemann [2355]. Spain [2736, 2131, 2591]. Spalek [55]. Spallanzani [1179]. Spanien [2696]. spanische [2696]. Spanish [1902, 1619, 1282]. Spannungsfeld [2838]. Spark [243]. spasmo [849]. Spatmittelalters [2819]. spazio [1436]. Speak [1855]. Speakable [2675]. Speaking [55, 2823]. Speakman [1190]. Special [558, 737, 1748]. Specialists [2947, 1880]. specierum [1177]. Species [2748, 1547, 729, 1187]. Specimens [2835, 1263]. Spectacle [2618]. Spectral [257]. Spectroscopy [2005, 2250]. Speculation [1174]. Speculations [144]. speculis [1177]. Speed [796, 399]. Spemann [2769]. Spence [1913]. Spencer [56, 1146, 1839, 1963, 2193, 126]. Spending [2202]. Sphaera [503]. Sphere [84]. Spheres [840]. Sphujidhvaja [40]. Spider [382]. spiegazione [2053]. Spiegel [1593]. Spiel [2287]. Spiers [2572]. Spies [586]. Spinoza [2372]. Spirit [676]. Split [2858, 2858]. Splitting [423]. Spokesman [2333]. Spontaneous [28]. Spores [1324]. Spouting [261]. Sprache [1572]. sprachen [2700]. Spread [1020]. Spring [558, 2052]. Springs [1404]. Squires [2688]. ’sredniowieczna [2816]. SRI [2247]. Srigley [1353]. Srinivasan [588]. SSSR [660]. SST [2291]. St [2100, 788]. Staat [497, 1308]. stabilité [726]. Stabilizing [2395]. Stachel [2654]. Stadler [1055]. Staff [2079]. Stage [1604]. Stages [878, 797]. Stagner [2032]. Stahl [932]. Stahlindustrie [2512]. Stalin [104, 1298]. Stand [2199]. Stanislaw [2567]. Stanley [2798, 684, 1262, 88, 243, 2180, 1648, 678, 215, 2181, 1734, 2306, 1599, 2416, 2264, 2620, 855, 2106, 286, 1748, 1874]. stanovlenie [660]. Stansfield [976, 1687]. Stanton [2373]. Stapleton [2221, 636, 2013, 2689]. Star [2098]. Stares [2139]. Stark [2533]. Starkey [2235]. Starlight [2250]. Starobinski [2728]. Starowolski [716]. Stars [8, 1038, 25, 2764]. State [2329, 1741, 1476, 402, 554, 477]. Statecraft [2257]. States [387, 2839, 1289, 2752, 1092, 2109, 1841, 1577, 2181, 368, 1796, 95, 55, 1929, 2707, 2934, 2154, 144, 1491, 2486, 2803, 95, 2757, 563, 1239, 2933, 814, 1549]. Statesman [1262]. Station [2041, 2292, 838, 300]. Stations [1835]. statique [624]. Statistic [2846]. Statistical [2096, 1690, 456, 1257]. statisticheskoi [2846]. Statistics [2434, 2497, 1015]. Status [2842]. Staudenmaier [1918]. Staudinger [556]. Staudt [1884]. Staum [650]. Stay [1281]. Steam [1928, 368, 1929]. Stearn [1676, 2400]. Stebbins [276]. Steck [916]. Steel [1370, 2511, 2690, 2162, 432]. Steenbock [2786]. Stefan [491]. Stefano [506]. Steffen [452, 353]. Steffens [51]. Stegena [1040]. Stehr [2560]. Steig [2479]. Stein [2780, 2144]. Steinberg [2832]. Steiner [731, 2057]. Stellar [901]. Stellung [497]. Sten [1284, 748]. Steneck [1586]. Stent [1949, 759]. Step [2789]. Stepan [1167]. Stepansky [1424, 191]. Stephanus [1507]. Stephen [1809, 2447, 2240, 2676, 2679, 1959, 489, 1819, 2184, 2681, 542, 365, 1283, 447, 2324, 2434, 1554, 2831, 1412, 1257, 2256, 236, 1247, 893, 2414, 2868, 2296]. Stephens [1053]. Stereochemistry [760]. Stereoregular [932]. Sterling [2811]. Sternbilder [2822]. Sternkatalog [2393]. Sternnamen [1677]. Steve [2354, 156, 420]. Steven [899, 2312, 1254, 852, 1986, 1907, 1477, 1341, 421]. Stevens [2774]. Stevenson [1111]. Stewart [511, 2412, 1829, 2182, 1888, 2106, 580]. Stichweh [1650]. Sticker [1909]. Stieber [2555, 754]. Stiefel [2046]. Stiftelsen [824]. Stigler [2434]. Stilgoe [1371]. Stillman [722, 1281]. Stimpson [213]. Stock [2166]. Stockholm [845]. Stocking [2331, 1330, 1664, 2331]. Stoddart [1965]. Stoics [2731]. Stokes [2918]. stoletti [285]. Stone [2123, 1320, 2358, 434]. Stone-Age [2358]. Stonebarger [1303, 1633, 1303, 1633]. Stoneburner [2052]. Stones [700]. Stony [1556]. Storage [970]. Storey [2938]. Storia [693, 579, 77, 355, 1255, 604]. Storr [1370]. Story [735, 569, 2625, 1522, 2022, 2226, 1975, 912, 2514, 1372, 2379, 1573, 1769, 1529, 2579, 2443, 1293]. Storytellers [1918]. Strachan [2207]. Straighteners [932]. Strain [2312]. Strange [386, 1968]. Strangeness [1193]. Stranger [1968]. Stranges [989]. Strasbourg [1280]. Strasser [2700]. Strassmann [798]. Strategic [2329, 1645]. Strategies [2928, 2430, 1086]. Strategy [2937, 1222, 1921]. Stratton [221]. Streetcar [841]. Streeter [2469]. Strike [2062]. Strip [911]. Strohmaier [2896]. Stroker [592, 591]. Stronger [1169]. Stroup [2429]. Struan [270]. Structure [1427, 1837, 678, 1724, 1623, 1921, 846, 1773, 404, 1313, 60, 2824, 1162]. Structures [968, 2482]. Struggle [1266, 1401, 2705, 160, 2503]. Struggles [1086]. Struik [2674]. Struktur [556]. Stuart [2244, 2862, 2110, 980, 2642, 1378, 1381, 1903]. Stubbe [1183]. Stubbs [1568]. Stuckelberger [1906]. Students [2531, 2079, 199]. Studer [481]. Studi [1685, 383, 703, 1188, 2122]. Studie [2245, 1576]. Studien [1186, 45, 553, 503, 908, 2720, 1908, 2897, 1275, 502]. Studies [336, 224, 2881, 1165, 217, 1000, 2892, 608, 1155, 186, 793, 2480, 1235, 601, 1997, 774, 956, 1269, 564, 2878, 685, 2893, 1744, 68, 1715, 1337, 1313, 340, 1174, 2212, 73, 498, 2278, 695, 1989, 2336, 405, 2717, 350, 2016, 611, 32, 2242, 72, 901, 29, 779, 583, 16, 123, 575, 1344, 546, 1850, 436, 787, 1509]. Studio [1682]. Studium [2817]. Study [445, 886, 2703, 1362, 1243, 119, 2534, 1779, 2007, 1958, 161, 1049, 270, 1764, 1778, 1520, 1444, 2562, 2600, 1012, 2089, 2471, 1034, 2164, 269, 701, 1214, 2161, 1632, 710, 2202, 665, 1409, 2705, 160, 1900, 229, 417, 2575, 622, 1440, 2390, 2518, 675, 350, 1762, 2918, 410, 1498, 974, 1338]. Stuewer [323, 1404]. Sturchio [1828]. Sturlese [1847]. Styles [2875, 2154]. Subjects [810, 511, 2079]. sublime [693]. Submarine [2062, 367]. Subroutines [1748]. Substance [1807, 1145]. Substitution [7]. Subtle [1138]. Succeed [1926]. Success [2134]. Successful [1927]. Successive [886]. Such [1855]. Sudund [1379]. Sue [2112, 2445]. Suède [2657]. Sufferers [2638]. Suffering [1768]. Sugar [403, 2444]. Sugimoto [239]. Suicide [2640]. Suite [853, 714]. suivi [651]. Sulayman [1439]. Sulfur [1876]. sull’elettricita [2219]. Sullivan [1484]. Summa [1681]. Summary [387, 1083]. Summer [2623, 1708]. Summerfield [2654]. Sun [1529]. Suñé [2627]. sunniya [1440]. Sunspots [1622]. sunt [1035]. suoi [1188]. Superhighway [841]. superieure [958]. superieurs [1819]. Superior [2874]. Supernatural [972]. Superpowers [2841]. Supersensibility [1716]. Superstition [2803]. Suppes [546]. Supplement [2284, 600, 2554]. Supply [1032, 710, 1501]. Supporters [1459, 1803, 2072, 2389]. Suppression [1422]. supra [1846]. Surgeons [565, 1901]. Surgery [298]. Suri [917]. Survey [32, 2833, 95, 913, 2326]. Surveying [2723]. Surviving [1644, 2220]. Susan [2631, 2626, 190, 2440, 2821, 213, 122, 2303, 2420, 812]. Susanna [1470]. Suspension [2753]. Susruta [784]. Susskind [736]. Sutherland [1819]. Sutt [130]. Sutton [2666]. Suyuti [1440]. Svami [2732]. Svante [1928, 913]. Svedberg [1798]. Sven [2287]. Sven-Eric [2287]. Svensk [1284]. Svetlana [1156]. svolte [77]. Swain [239]. Swartz [2165]. Sweden [1928, 882, 994]. ’sweietle [120]. swej [120]. Swenson [2138, 248, 394]. Swetz [2815]. Swine [1211]. Sydney [2748, 2849, 1544]. Sykes [868]. Syllabi [2242]. Sylvester [365]. Sylvia [2138]. Symbol [1191]. Symbolism [2541]. Sympathy [1767]. Symphorien [107]. Symposia [2269, 370, 2007, 1625, 1134]. Symposion [46, 108, 985]. Symposium [1804, 670, 1193, 9, 513, 2386, 1760, 238, 882, 323, 2363, 558, 2104, 587, 1287, 1292, 1163, 2047]. Symposiums [1990]. Syndrome [2341]. Synopsis [784]. Synthesis [467, 1836]. Synthetic [932, 1296, 1926]. Syphilis [965, 1511]. System [49, 1954, 969, 2534, 257, 2510, 914, 1549, 300, 2502, 1921, 351, 831, 1429, 2606, 1494, 694]. System/360 [1429]. Systematics [768]. Systementwurfs [241]. systèmes [445]. Systems [136, 1650, 286]. Szabó [305, 1845]. Szasz [1529]. Szmant [1751]. Szreter [2805]. Sztompka [2108].

T [2767, 344, 2166, 872, 337, 903, 1023, 1484, 2555, 2362, 428, 2949, 2099, 700, 2183, 1653, 1713, 1983, 491, 352, 968, 2756, 1499, 1136, 729, 615, 1016, 179, 1578, 1964, 1483, 2673, 137, 418, 754, 2400, 1744, 1372, 2944, 102, 1931, 367, 2943, 206, 588, 1676]. T. [1608, 751]. Tabarroni [1188]. Table [1810]. Tables [780, 2043, 1439, 1466, 1591, 114]. Tabukasch [2287]. Taccola [1379]. Taconic [49]. Tactics [2207]. Taddeo [781]. Tagliaferri [2338]. Tagung [1288]. Tai [2026, 571]. Tail [2759]. Take [2247, 269]. Take-Off [2247]. Taking [2398]. Talcott [2329]. Tale [1747]. Tales [2077]. talet [910]. Talk [2471]. Talking [1079]. Tamamushi [944]. Tamara [2375]. Taming [1665, 1764]. Tamny [1349]. Tania [606]. Tann [843]. Tanner [701]. Tantra [917]. Tapia [2592, 551]. Tarbell [1262, 2181, 1262, 2181]. Tate [1902]. Tatsache [1306]. Tauber [220]. Taxonomic [2835]. Taxonomy [1350]. Taylor [2178, 758, 1740, 1289]. te [571]. Teacher [868]. Teaching [2497, 201, 163, 2365, 2036, 913, 2818]. Teboul [1441]. Tech [2558, 2709, 2514]. Technical [1362, 682, 1438, 352, 104, 341, 682]. Technics [35]. Technics-out-of-Control [35]. Techniek [206]. Technik [353, 19, 1640, 2807, 1011, 637]. Technik-Geschichte [637]. technique [639, 505]. Techniques [1345, 1397, 301, 1375, 2859]. Technische [2830, 2830]. technischer [2512]. Technological [2688, 797, 2020, 354, 774, 277, 1363, 2546, 1934, 621, 2511, 2214, 2005, 844, 1930, 350, 2312, 2422, 1649]. technologie [2563, 585, 1410, 371]. Technologies [2689, 420, 2686, 2934]. Technologist [2823]. Technology [76, 2227, 377, 2138, 2538, 370, 1367, 2013, 540, 2307, 469, 1817, 1092, 1083, 2949, 2353, 2012, 2099, 670, 2045, 2492, 2521, 2442, 1112, 2449, 1097, 2274, 106, 2588, 1771, 352, 2080, 616, 2756, 1549, 35, 2869, 2242, 2938, 1366, 750, 1679, 2687, 2935, 1365, 621, 1642, 1919, 411, 2241, 2370, 2788, 1645, 584, 2300, 2589, 1537, 2514, 2591, 2893, 2207, 913, 1113, 711, 1392, 689, 1279, 2587, 1309, 1464, 635, 1364, 2666, 433, 2772, 1007, 350, 419, 1304, 367, 841, 1996, 619, 2662, 2039, 410, 1498, 684, 2685, 2042, 2121, 2881, 1240, 2339, 842, 2297, 1674]. Technology [176, 1499, 710, 205, 818, 2272, 1361, 1885, 1368, 2336, 2203, 2862, 1928, 2311, 1918, 915, 104, 542, 2944]. tecnica [604, 2592, 443, 1790]. Tectonics [1320, 2148]. Ted [1404]. Tegni [1779]. Teichmann [761]. Teil [2904, 1485]. Tejon [1993]. Tekniska [370]. Telegrapher [2582]. Telegraphy [367]. Teleological [807]. Teleology [1872, 530, 1222, 208]. Telepathy [2466]. Telephone [1921]. Telescope [256, 1389]. Telescopes [1953, 1924]. Television [934, 2507]. Teller [2771]. Tellico [2175]. Temin [2534]. témoinages [1274]. Temperature [447]. tempo [579, 1436]. temps [315, 906, 1375]. Ten [374]. Tendenzen [154]. Tender [2688]. Tenenbaum [620]. Tenth [1698]. tenu [496]. teori [910]. teoría [442]. teoriche [77]. teorici [703]. Teorie [488]. teorii [2846]. teoriia [281]. TePaske [1902]. Tercentenary [1644]. Terence [2908]. Teresa [621]. Terminal [969, 300]. Terminologia [2051]. Terms [160]. terra [488, 579]. Terrada [1782, 2751, 444]. terre [574]. terrestre [574]. terrestrisque [1774]. territoriale [2699]. Terror [1322]. Terry [1272, 1169, 2242, 1742, 480]. Terzian [1575]. Tesing [2808]. Tess [1546]. Test [872, 531, 2936]. testi [2160, 355]. Testing [1819, 2874, 734, 2770, 1305]. Tests [2768]. Teutons [1329]. Texas [2138]. Texera [1917]. Text [1439, 1909]. Textband [2502, 553]. Texte [36, 500, 2043]. Texten [1201]. Textes [2476, 780]. Textfassung [2586]. Textkritische [117, 916]. textos [1860]. Texts [1089, 181, 99, 2213, 1976]. Thabit [2895]. Thackrah [2114]. Thackray [1828, 2288]. Thaddeus [423, 1892]. Thames [2222]. Thauer [2808]. Theerman [2057]. Their [514, 2499, 80, 1166, 1746, 2224, 1418, 2253, 207, 59, 1647, 1560, 2749, 835, 549, 341, 2400, 2670, 422, 1952, 350, 1953, 619, 2098, 2583, 6]. Them [569]. Thematic [904]. Theme [35]. Themes [1832, 1489, 1190]. Theodicee [668]. Theodor [684, 2052]. Theodora [2562]. Theodore [2096, 1553, 1150, 752, 182]. Théodoridès [227]. Theodosius [561]. Theologian [384]. Theological [1249, 681]. Theology [2455, 1249, 2213, 2495, 2921, 927, 2310]. Theon [2043, 114, 114]. Theorem [1132]. Theoretical [2073, 631]. Theoretical-Historical [631]. Theorie [2700, 2808, 1655, 438, 1659]. Theoriebildung [1789]. Theories [1307, 1164, 2683, 2500, 937, 2846, 621, 2124, 1482, 1140, 286, 695, 1441, 2259, 886, 2437, 1887, 1305, 920, 2866, 778, 2778, 2188, 1602]. Theory [2395, 294, 995, 136, 1745, 2152, 408, 2327, 742, 93, 217, 2664, 806, 168, 692, 1259, 1726, 2678, 939, 2599, 1728, 484, 17, 968, 1107, 1836, 2320, 2321, 737, 989, 1919, 460, 411, 195, 2251, 2519, 2668, 2357, 927, 308, 976, 855, 60, 1480, 1257, 645, 1831, 1174, 2482, 1961, 2249, 2097, 2453, 2360, 550, 1761, 696, 1832, 2441, 1422, 2397, 1550, 342, 1848, 609]. Therapeutic [2268, 297]. Thérapeutique [376]. There [597]. Therese [639, 2170]. Thermal [168]. Thermodynamics [336, 610]. Thermometer [548]. thesaurus [1782]. Thesen [556]. Thesis [941, 2540]. Thévenin [2563]. Thiel [356]. Thiele [465, 247]. Thiere [1026]. Thiessen [1945]. Thillaud [1335]. Things [1715, 2212]. Think [2423]. Thinkers [312]. Thinking [610, 380, 2096, 2664, 1690, 966, 2686, 1717]. Third [2269, 1916, 2492, 2846].

Thirteenth [439, 2611, 1341, 1590, 942]. Thirteenth-Century [439, 942]. Thirties [1526]. Thom [177, 177]. Thomas [569, 1575, 611, 2326, 2192, 469, 409, 2487, 1613, 476, 2833, 1990, 2283, 957, 2485, 1585, 2930, 1306, 1232, 687, 1919, 2251, 2475, 1787, 947, 925, 1901, 2668, 562, 791, 1769, 1828, 97, 1252, 2227, 2838, 2840, 1246, 2475, 12, 1347, 2475, 795, 1687]. Thomasius [1220]. Thomasset [979]. Thompson [183, 126]. Thomson [20]. Thorpe [559]. Those [1581, 1494]. Thought [516, 550, 168, 2907, 3, 1907, 816, 2457, 945, 1727, 2233, 576, 2326, 1950, 1195, 2826, 2775, 1115, 1985, 2445, 278, 127, 2059, 1250, 35, 2575, 562, 2921, 244, 851]. Thoughts [2673, 1941, 673]. Three [1290, 878, 2518, 1449, 2409, 2509]. Threshold [852]. Thrower [1512]. Thuillier [2880, 2880, 482, 895]. Tianwenxue [1102]. Tiby [50]. Tichi [2307]. tiden [1284]. tiempo [2627]. t’ien [571]. t’ien-jan-ch’i [571]. Tiffany [2690]. Tiger [1416, 2954]. Tihon [1592, 2043, 114]. Tijdschrift [206]. Tilmann [1566]. Tim [2189, 2098]. Time [888, 702, 425, 2821, 278, 1377, 1530, 2348, 822, 2124, 124, 1806, 1309, 2256, 1247, 348, 2706, 2658, 2845, 2054, 1611, 2256, 1581, 1307, 2821, 2256]. Timekeeping [2821, 20]. Timeo [979]. Times [220, 2264, 907, 2746, 955, 571, 1816, 1445, 1166, 1603, 2264, 634]. Timothy [2632, 729, 1852]. Tina [1809, 2555, 2046]. Tincal [1756]. Tingyu [2450]. Tinkler [2027]. Title [101, 1465]. Titre [2773]. Tiziana [1682]. Tobacco [1874]. Tobey [767]. Tobias [2303, 577]. Toby [1663, 892, 565, 2404]. Today [220, 882]. Todd [1530, 1513]. Todes [1288]. Todesco [2597]. Toil [2299]. Tolstoy [733]. Tom [1243, 1487, 2279, 2753, 2565, 2514, 2491, 568, 1373]. Tomas [1782, 2751]. Tomes [1769]. Tomorrow [2039, 557]. Tony [2433]. Too [1024]. Tool [2202]. Tooley [815]. Tools [2123, 711, 2358]. Toomer [1432]. Tootil [885]. Topics [1721]. Torbern [2054]. Tore [1353]. Torrence [2123]. Torrini [44, 1856]. Torsten [2083]. Tort [651]. tot [2576, 2825]. Totius [1453]. Touched [1995]. Toulouse [496]. Toulouse-Le [496]. Tournadre [1117]. Tovmas [130]. Towler [2115]. Town [290]. tra [2219, 1442]. Trabulse [1860]. tractatum [1846]. Tractatus [2898]. Tractor [2273]. Tracy [1262, 2181]. Trade [2368]. Trades [2114]. Tradescant [1644]. Tradescants [1644]. Tradition [1563, 330, 2874, 2232, 2337, 2393, 2924, 1629, 882, 2800, 916, 1677, 107, 712, 2447, 113, 2106]. Traditional [2450, 239]. Traditions [1200, 600]. traduction [2043]. Tragicomical [336]. Trail [1756]. Traité [445, 1036]. Trajectories [2543]. Tranquility [1070]. Transatlantic [59, 367]. Transcendence [937]. Transcendent [1718]. Transcendental [164]. Transcription [240, 1786]. Transfer [2689, 1919, 2772]. Transformation [2312, 1016, 983, 1223, 1629]. Transformations [1239, 147]. Transformed [165]. Transit [963]. Transition [633, 1317, 1242]. Transitions [2753]. Translating [259, 462]. Translatio [849]. Translation [1508, 1177, 1505, 1786, 1958, 1778, 1851, 2391, 1440, 1202, 2518, 242, 1176, 1281, 2654, 1438]. Translations [1417, 373]. translationum [373]. Transmission [2152, 1921]. Transmutation [1701, 2748, 903]. Transplant [656]. Transport [1378, 1031, 433]. Transportation [841]. Transurane [830]. Transylvania [1237, 2413]. Trapped [2409, 990]. Trautmann [2635]. travaux [2350, 2407]. Travel [1807, 561]. Travelers [2013]. Travels [788, 1913, 2407]. travers [1705]. través [2627]. Travis [1756]. Traxel [2031]. Traynham [2359]. Treatise [1774, 1855, 1202]. Treatises [777]. Treatment [2265]. Treatments [2931]. Treder [133]. Tree [2567, 859]. Trees [2537]. tremore [849]. Trends [407, 2390]. Trenn [903, 423, 1892]. Trescott [774, 277]. treti [2846]. Trevor [1089, 106, 2228, 758, 1113, 1233]. Trial [1928, 2021, 709]. Trials [709]. triangle [1506, 2905]. Tribute [429, 2850, 932, 2605]. Triebe [1026]. Trigg [2421]. Trigonometry [1744]. Trimble [2853]. Trinity [240, 1349, 1529]. Tristram [2881]. Troitzsch [637]. Trojans [1329]. Tropanalkaloide [823]. Tropfke [283]. Trower [2531]. True [2443]. Truesdell [336, 2802]. Truman [2014]. Truslow [1293]. Trust [2426]. Truth [2148, 1422, 1341, 1099, 2335, 1907]. Trystram [581]. ts’ao [378]. ts’ao-mu [378]. Tselos [2339]. Tsien [2370]. Tsouyopoulos [1029]. Tsuen [2370]. Tsuen-Hsuin [2370]. Tsuna [1599]. Tsung [2588]. Tube [33]. Tuberculosis [2771, 2848, 2263]. Tübingen [985]. Tucci [2338]. Tucher [1818]. Tucker [1791, 1953, 1953]. Tucson [2224]. Tudor [980, 1903]. Tugnoli [644]. Tulp. [1652]. Tummers [1780]. Tumor [2441]. Tunnelling [2860]. Turberville [876]. Turbojet [567]. turbulences [500, 500]. Turing [1291, 1360]. Turn [2365, 1921, 90, 1646]. Turned [668]. Turned-Down [668]. Turner [2289, 613, 2114, 654, 2348, 1472, 613, 2324, 2179, 2420]. Turning [2067]. Turriano [1932, 1154]. Tusi [2814]. Tuskegee [965]. TVA [2175]. Twain [707]. Tweedale [2511]. Twelfth [2275, 2733, 2316, 2046]. Twelfth-Century [2275, 2733, 2046]. Twelve [1632, 1190]. Twentieth [2848, 2846, 106, 956, 1832, 2648, 1732, 2157, 1113, 2495, 2430, 1727, 92, 2549, 1082, 232]. Twentieth-Century [1113]. Twenty [1405, 1932]. Twenty-One [1932]. Twice [1529]. Two [738, 678, 59, 835, 548, 1642, 781, 172, 2633, 1373, 987, 360]. Tyana [309]. Tycho [1785, 2409, 1938]. Tylor [839]. Tyne [33]. Type [2835]. Typographic [2515].

U [80, 119, 1288, 441, 133, 1163, 126, 1462]. u.a [451]. U.R.S.S [1753]. U.S [2289, 2105, 2139, 1923, 1024]. U.S. [2833]. Uberblick [1915]. Ubergang [1285]. Uberoi [2055]. Übersetzung [2586, 1433]. Ubersetzungen [2393]. Übersicht [2483]. Udelson [934]. UdSSR [891]. Uebersetzung [117]. Uexkulls [727]. UFO [2211]. Ugo [1783]. uit [649]. Ulam [2567]. Ulfried [1969, 1915]. Ulisse [644]. Ullmann [2480, 429, 39, 1433]. Ullmann-Margalit [2480]. Ullrich [2551]. Ulrich [1120, 637, 2512]. Ulrike [2950, 2838, 2760]. Umberto [693]. umjetnog [2739]. Umweltlehre [727]. Umweltschutz [174]. unbekannten [1677]. Unbelief [1544, 2494, 2179]. Unbelievers [2402]. Uncertainty [2434]. Unconventional [211]. Understand [2899, 2730]. Understanding [1728, 2346, 2540, 272, 1540, 1648, 2396, 2421]. Uneasy [1734]. Unequal [2173]. Unfinished [2625, 1836]. Unfolded [99]. Unforgettable [2583]. Ungar [305]. Unger [2032]. Unhealthiest [2642]. unheilige [2490]. Uni [2563]. Unification [2598, 467]. Unified [2846, 2163, 2482]. Union [2289, 656, 2934, 990, 1366, 1030]. Unique [1439]. Unis [2563]. United [387, 2839, 2752, 1491, 1092, 2109, 814, 1841, 2181, 2486, 2803, 1549, 368, 95, 55, 1929, 2757, 2707, 563, 1239, 2934, 2933, 2154, 144, 1577]. Units [21]. Unity [1403, 2445, 2658, 2078]. Universal [2700, 1350, 855]. Universal-sprachen [2700]. universalis [2700]. Universe [627, 1774, 1518, 2159, 1757, 1316, 220, 612, 1261, 1258, 898, 2568, 1791]. Universidad [1613]. Universita [1890]. Universitat [1475, 985, 82, 345, 479, 2489, 2807]. Universitet [910]. Universities [2485, 1474, 1002, 2818, 1236, 1563]. University [1859, 1559, 81, 477, 1237, 163, 2401, 2242, 2755, 2413, 109, 2899, 2079, 2018, 2261, 2786, 1072, 1076, 2544, 1146]. University-Industrial [2261]. Unless [2899]. Unmaking [1171]. Unnatural [734]. Unpopular [354]. Unpublished [2836, 720, 2789, 403, 359]. Unraveled [597]. Unravelling [362]. Unrepentant [1418]. Unresolved [2419]. Unruly [1618]. Unschuld [119, 441]. Unspeakable [2675]. Unsung [2304]. Unternehmensstrategien [2512]. Untersuchungen [1201, 424]. until [1166, 1573]. Untold [912, 2379, 2328]. Unwitting [1328]. Unwritten [270]. Uppsala [2401, 882]. Upright [2333]. Upton [1786]. Uranium [135, 2137]. Uranus [902]. Urbach [2524]. Urban [841]. Urbana [1559]. Urbana-Champaign [1559]. Urmanuskript [2031]. Ursachen [309, 556]. Ursprung [424, 2720]. Ursula [309]. Ursus [1785, 2409]. USA [913]. Usage [1762]. Use [1081, 1370, 2940, 548, 2214, 531, 1748, 549, 2635]. Used [2026]. Useful [1855]. Uses [1396, 616, 1751, 1805, 1515, 350]. Usher [1280]. USSR [1400, 274]. usu [1505]. Uta [2095, 1558]. Utilities [2116]. Utopia [2823]. Utopian [1910]. Utopianism [1649]. Utopie [1789].

V [1201, 2035, 1559, 2563, 796, 1639, 2884, 2949, 1644, 1784, 362, 189, 693, 2040, 2057, 1421, 2903, 564, 1544, 2370, 2589, 2754, 451, 2550, 1142, 1352, 1590, 2116, 891, 1783, 1315, 2415, 1748, 2522, 660, 74, 2846, 285]. v. [1753]. Vaccine [769]. Vaccinia [769]. Vaclava [1352]. Vacuum [33, 1140]. Vadée [414]. Valence [989]. Valencia [2751]. Valenstein [2931]. Valentine [1180]. Valerie [2198, 680]. Validity [2518]. Vallasi [2949]. Valley [2134, 2226]. Value [2419]. Values [957, 352, 1560, 542, 1729, 1234, 1314]. Valverde [2627, 1037]. Vance [685]. Vanderbilt [2632]. Vanderburg [2482]. van’t [2825]. Vare [2583]. Variations [623]. Varieties [2418]. Varignon [2904]. Vary [367]. Vasco [161]. Vasilievich [2523]. Vatican [1315]. Vaughan [1787, 1787]. Vaulting [1834]. Vaux [1200]. VCR [2271]. Vdovichenko [2846]. Ve [237]. vecino [2592]. Veda [2049]. vedischen [310]. végétalité [291]. Vegetti [112]. Vehicles [797]. veka [281, 74, 2846, 2903]. Velazquez [963]. Velikhov [619]. Velikovsky [1587]. Venerate [1822]. Venereal [2109, 1841]. Venezuela [1917, 1917]. Venice [1000]. Venomed [2636].

Venus [963, 651]. Veraart [356]. Verba [1453]. Verbandstheorie [282]. Verbindungen [547]. Verbrugge [2646]. Verdet [1319]. Verdict [2211]. Verdoorn [2155]. verdrängte [2489]. verdrängtes [2946]. Vereins [1011]. Vererbung [2152]. vergleichenden [727]. Verhaltenesforschung [727]. Verified [2373]. verite [2050]. Verkade [1758]. Vern [2780]. Vernadsky [1075]. Vernon [1926]. Vernunft [2813]. veroiatnostei [281]. Verrier [1483]. Versammlungen [2483]. Version [504, 2407]. Versuche [287]. versus [2799, 1215, 403, 1938, 256, 1689, 6]. Vertebrate [2501]. Vertical [1169]. Verwandtes [641]. Verwey [1840]. Vessel [190]. Vessuri [1917, 1917]. Vestigia [1906]. vetraria [724]. vetus [1035]. VI [1201, 2949, 116, 1983, 382, 106, 2588, 1679]. Vicente [2751, 963]. Vickers [1809, 2525]. Vico [198, 579, 1806]. Victor [2194, 1782, 1681, 1232, 1866, 2142, 236]. Victoria [2087]. Victorian [2455, 2920, 1118, 1236, 654, 2640, 957, 709, 1016, 2744, 2494, 1209, 1689, 2746, 894, 2750, 2670, 13, 1446, 1994, 2918, 256, 50, 2322, 957, 1544, 314]. vid [910]. vida [2593, 2592]. VideoDisc [2366]. Vienna [2899]. Viete [1744]. Vietnam [2458]. View [1567, 746, 938, 220, 1405, 686, 2743, 1886]. Views [1809, 1023, 1764, 8, 2091]. vigilance [2761]. Vignoles [1170, 1170]. VII [641, 106, 460, 2589, 2391]. VIII [2076, 1508, 1288, 1136, 223]. Viktor [2769]. Vilar [2751]. Villanova [849, 1846]. Ville [2555]. Vincent [1567, 1149, 2050, 2761]. Vincenzo [1517, 2639]. Vinci [786, 2820, 1680, 1513, 947]. Vintage [569]. Viola [2105]. Virginia [2166, 1425, 1884]. Virgins [277]. Virology [103]. Virtuosi [510]. Virus [1874, 769, 2625]. Viscountess [2052]. Visible [1129, 87]. Vision [976, 1276, 1808, 979, 1657]. Visionary [151]. Visions [1865, 1003]. Visser [575]. Visual [205, 976, 2384]. Visvanathan [2806]. vita [2280]. Vital [403, 5, 1569]. Vitamin [1972]. Vitellion [573]. vitesse [184]. vivant [446]. vivants [315]. Vives [235]. Vivian [1242, 2892, 303]. Viviani [2712]. Vizgin [2846]. Vklad [660]. Vladimir [2532]. Vocation [1863, 2329]. Vogel [283, 1942, 297, 715]. Voices [2562]. Void [2124]. Voights [1438]. Vol [500, 451, 849]. Volcanic [1487]. volgare [437]. Volk [1593]. Volker [2819, 718, 1686, 1485]. Volta [2219]. Voltaire [1045]. Volterra [1188]. Volume [1981, 558, 250, 519, 800, 1060, 1383, 1696, 1999, 2314, 2609, 2953, 29, 2076, 1933, 2289, 610, 604, 843, 2810, 653, 511, 845, 1847, 1864, 1959, 2811, 44, 1096, 2949, 2259, 2904, 2185, 2458, 732, 2032, 1686, 1909, 1829, 1860, 116, 641, 2182, 2777, 2694, 2480, 1235, 1983, 714, 2293, 241, 376, 578, 485, 382, 1288, 2040, 2052, 106, 2588, 2627, 1379, 2545, 2057, 928, 1951, 2348, 1336, 368, 2247, 373, 778, 2652, 37, 283, 1136, 1679, 95, 2393, 2221, 2118, 2370, 600, 2519, 85, 2589, 1913, 2671, 2654, 1866, 1929, 2142, 642, 1776, 412]. Volume [238, 1024, 2129, 1846, 571, 301, 1364, 1733, 1590, 609, 2926, 2760, 853, 133, 2654, 1101, 2138, 1867, 2247]. Volumes [1201, 1120]. Volz [984]. Vom [2199, 2504, 1285, 762, 2817, 1990, 178, 1908, 1055, 2901, 1340]. voor [206]. Vorgeschichte [2763]. Vorlaufer [1475]. Vorlesungen [1019, 1204]. Vorlesungsbuch [1153]. Vortrage [2031, 985, 2483, 2807]. Vowles [1353].

Voyage [48, 2766, 1512, 2103, 2403, 1807]. Voyager [1750]. Voyagers [2105]. Voyages [125, 1322, 2407, 1808]. vs [2833, 939]. Vucinich [1400]. vv [2549].

W [2570, 46, 2439, 381, 1282, 629, 1933, 81, 2762, 2891, 2703, 1612, 1893, 1605, 559, 2579, 2021, 1429, 969, 511, 1949, 2531, 2910, 567, 272, 631, 1864, 1959, 1484, 2451, 163, 210, 1407, 2374, 2527, 1121, 32, 915, 575, 1636, 1355, 658, 45, 825, 487, 2148, 1842, 2558, 1638, 1347, 1500, 574, 2600, 1865, 1717, 2459, 671, 1884, 2057, 1951, 1512, 956, 2853, 2366, 1381, 2213, 2107, 1712, 1635, 2038, 1330, 1524, 2033, 2029, 2365, 1463, 1210, 300, 1016, 1467, 2277, 2221, 2118, 1921, 1088, 2602, 317, 1912, 1261, 2478, 1664, 2659, 2264, 1142, 755, 2122, 2193, 1729]. W [2481, 2831, 388, 1412, 635, 1931, 2825, 2061, 1772, 1740, 1825, 1996, 576, 628, 2905, 171, 2331, 120, 2816]. W. [1874, 2620, 100]. Waard [1394, 578, 2129]. Wachhorst [795]. Wade [616, 269, 1657, 738, 1099]. Wadland [126]. Waerden [570]. Wagner [1437]. Wahlvorschläge [2726]. Waite [2783]. Waites [734]. Waith [2694]. Wakefield [2643]. wal [308]. wal-camal [308]. Waldenfels [731]. Waldo [866]. Walentynowicz [1565]. Wales [956]. Walker [2555, 123, 1694]. Wallace [1746, 1791, 382, 1831, 1953, 1847, 1115, 2590, 971, 1108, 837, 242, 2526, 1948, 1688, 386, 924]. Wallis [2415, 212, 2415]. Walras [2829]. Walter [132, 2566, 1043, 1855, 812, 2783, 1251, 2278, 1135]. Walters [702]. Walton [2659]. Walzer [2201, 1426, 1110]. Wandel [1094, 1019, 292, 761]. Wangensteen [298, 298]. War [2649, 2289, 2375, 189, 509, 2852, 2893, 2715, 258, 2379, 2136, 1869]. Warburg [134]. Ward [1559, 756, 2628, 2367]. Warfare [2062, 2768, 2120, 2136, 2572]. Warfield [1249]. Warner [1610, 1610, 214, 2268, 486]. Warnow [133]. Warren [337, 918, 162, 2705]. Wars [2154]. Wartofsky [272, 1712, 1467]. Was [2766, 597, 1710]. Waschkies [178]. Wash [2440]. Washington [2301, 1760, 1191]. Wasserman [1921]. Wasserstein [1202]. Water [2440, 1674, 2348, 1169, 1501, 261, 2111, 1281, 1374]. Water-Clocks [2348]. Water-Cure [2440, 2111]. Water-Lifting [1674]. Water-Spouting [261]. Watercolors [1809, 2401]. Waterloo [2207]. Waterlow [1338]. Waterpower [368]. Waters [1501, 1202]. Waterson [103]. Waterston [93]. Watkins [1725]. Watson [2780, 1345, 30]. Watt [843]. Wave [2703, 1771, 1416, 2954]. Wave-Particle [1416, 2954]. Waves [2499]. Wawn [2747]. Way [224, 2358]. Wayne [1151, 41]. Ways [88, 2022, 2420]. Wear [1854]. Weart [1963, 56]. Web [2430]. Weber [1663, 2324, 1730, 2078, 2828]. Webster [896, 398, 145, 299, 1351]. Wedeen [1770]. Wegeler [2489]. Wegener [427, 2149, 988]. Weierstrass [693]. Weigels [241]. Weight [778]. Weights [21, 2469]. Weil [1745]. Weimar [2492]. Weinberger [1910]. Weindling [1230]. Weiner [393]. Weingart [1396, 352]. Weinshank [2240, 729]. Weir [1259]. Weisheipl [1626, 307]. Weiss [2847, 844, 1640]. Weisser [309, 1340]. Weisskopf [451, 1866]. Weisz [584]. Weitling [1789].

Welch [1111]. Weldon [2247]. Welfare [2631, 2185, 95]. Wellcome [2426]. Wells [1814]. Welsh [2726]. Welt [1760]. Weltanschauung [1485]. Weltauffassung [1055]. Weltbilder [1094]. Weltbildes [761]. Welther [1413]. Wende [762, 1576]. Wendel [2102]. Wengenroth [2512]. Went [2690]. Were [2686, 996]. Wereldbeeld [1514]. Werk [2838]. Werke [1612, 1893, 1686]. Werner [1612, 1893, 1830, 2897, 1519, 1525, 2031, 1555]. Werskey [87]. Wertheimer [339]. Wertime [752]. Wesley [113, 1724]. West [417, 2758, 871, 1498, 2860]. West-European [417]. Western [1360, 2733, 986, 278, 1139, 707, 2337, 1087, 1921, 1163, 1596, 2800, 628, 2679, 2048, 1596]. Westfall [1886]. Wexelblat [873]. Weyant [494]. Whale [2042, 2121]. Wharton [2602]. Wheaton [1554, 949, 1416]. Wheatstone [1657]. Wheel [1169]. Wheeled [1378]. Wheeler [2175, 2186, 768, 1748, 2285, 1582, 2103]. Where [888]. Which [80, 312]. Whig [62]. Whirlwind [569]. Whiston [1684]. Whitaker [2018]. White [2263, 2555, 96, 2474, 364, 60, 34, 1314, 2102, 2550]. Whitehead [2142, 78]. Whiteside [1136]. Whitfield [2239]. Whither [125, 616]. Whitley [1813, 1742]. Whitlock [221]. Whitney [1923]. Whittaker [2500]. Whittle [2443, 2443]. Who [569, 529, 224, 1322, 1494, 2015, 2764, 1581, 2015]. Whole [702, 2670]. Whole-Plant [702]. Whorton [1751, 1054]. Whyte [1566]. Wickey [2339]. Widely [2026]. Wider [1250]. Wie [1710, 1640]. wieku [2816]. Wiener [2293, 420, 2940]. Wightman [511]. Wigner [2460]. Wilbur [2729]. Wild [2607]. Wildavsky [620]. Wilder [694, 594]. Wildes [1739]. Wildland [893]. Wiley [1151]. Wilfrid [210]. Wilfried [764]. Wilhelm [2738, 2763, 1789, 2857, 1566, 696, 824, 985, 1861, 2702, 2847, 1046, 1738, 1702]. Wilhelmus [615]. Wilkes [1748, 1935]. Wilkie [1505]. Wilkinson [103, 1504]. Will [2368]. Willeford [384]. Willem [2062, 627, 2482]. Willey [2224]. William [558, 2249, 29, 2677, 2329, 1665, 1656, 2634, 1652, 2175, 1847, 2360, 2375, 1971, 2209, 2922, 1334, 1332, 2672, 1235, 1022, 1112, 837, 364, 2756, 1246, 2584, 1741, 2528, 2755, 1964, 2855, 341, 418, 1885, 2400, 1861, 1666, 1661, 320, 2435, 2193, 1729, 472, 890, 539, 791, 2323, 314, 43, 244, 1977, 2044, 645, 1419, 2554, 1825, 90, 1991, 1688, 1099, 182, 1676, 806, 2793, 458, 341, 2439, 1289, 1607, 2373, 1206, 2059, 1684, 1341, 173]. Williams [1290, 106, 350, 1529, 2228, 2272, 2037, 1113, 1233, 964, 2162, 2273]. Williamson [1039, 2925]. Willis [1923]. Willson [1286, 390, 2741]. Willy [1776, 811]. Wilson [186, 1879, 1450, 2918]. Win [738]. Winau [2630, 19]. Wind [190]. Winder [2278, 2278]. Windmill [2275]. Windscale [1194]. Wine [2795]. Wines [2079]. Wineskins [2795]. Winged [1372, 1372]. Wings [2291, 568, 1372]. Winifred [98]. Winkelmann [2067]. Winner [2042, 2121, 35, 269]. Winning [2709]. Winter [1983]. Wintrobe [1975]. Wirken [2838, 798]. Wirkungs [2738]. Wirkungs- [2738]. Wirkungsgeschichte [1055, 2946]. Wirtschaft [823]. Wisconsin [2786, 1076]. Wisdom [2665, 2308, 2717]. Wise [1923]. Wiskunde [206]. Wissenschaft [1789, 823, 2830, 1576, 2896, 310, 570, 209, 1308]. Wissenschaften [2720, 292, 891, 2287, 1777, 1572]. Wissenschaftler [1094].

Wissenschaftliche [1893, 764, 2807, 247]. Wissenschaftlichen [836, 1055, 452, 1519]. Wissenschaftlicher [451, 2720, 1650]. Wissenschafts [1275, 1776]. Wissenschafts- [1275, 1776]. Wissenschaftsbegriff [1789]. Wissenschaftsforschung [764]. Wissenschaftsgeschichte [603, 764, 353, 271, 75, 2210, 985, 2483, 2807]. Wissenschaftslehre [154]. Wissenschaftspolitik [1308]. Wissenschaftsreform [154]. wissenschaftstheoretische [241]. Wiswall [1158]. Witch [789]. Witch-Hunting [789]. Witchcraft [398]. within [213]. Without [2179, 1187, 2179]. Witmer [1744]. Witold [2805]. Witten [109]. Wittfoht [2274]. Wittgenstein [2777, 324]. Wizard [189, 1381]. Wo [1908]. Wodeham [2898]. Woeful [2626]. Woerden [627]. Wohl [1209]. Wohlauf [637]. Wohltatigkeit [1593]. Wolf [1396, 1659, 2812, 984, 908, 292]. Wolfe [2843]. Wolff [648, 424]. Wolffian [596]. Wolfgang [2726, 1011, 352, 451, 1866, 1760, 1572, 452, 451]. Wolfgang-Hagen [1760]. Wolfram [2817]. Wolko [2208]. Wolters [2254]. Woman [2867, 640, 213]. Womanhood [2646]. Womb [1974]. Women [84, 1425, 2440, 2112, 2299, 2694, 213, 277, 1219, 2562, 2385, 1767, 835, 2173, 2019, 338, 2328, 268, 2067, 2111, 2583, 2140, 141, 1974, 2298, 1242, 1426, 2140, 2301, 2300, 1086]. Won [2803]. Wonderful [2410]. Wonders [1372]. Wood [2898, 2028]. Wooden [1376]. Woods [2537]. Woodward [47, 2545, 621, 621, 2356]. Wooldridge [1372]. Woolf [1193, 676]. Woolgar [2354, 156]. Word [1182]. Work [276, 946, 1605, 2299, 1267, 1713, 1799, 2471, 1734, 1365, 1988, 878, 2142, 2823, 2746, 141, 900, 1352, 2596, 1980, 1353, 2660, 1911, 2523]. Worked [312]. Workers [2648]. Working [1772]. Works [1981, 786, 1893, 2531, 2621, 1959, 1158, 2293, 382, 1277, 548, 1544, 1787, 1818, 552, 2081, 2126, 511, 1120, 1981, 1612, 2553, 1632, 1736, 2518, 511]. Workshop [2390]. World [735, 1252, 1282, 336, 2650, 2736, 1345, 738, 1242, 2026, 2048, 2349, 1886, 2052, 2136, 122, 1865, 180, 778, 1461, 686, 2852, 2781, 1226, 2379, 350, 1581, 2731, 2743, 2154, 2765, 1621, 2439, 1286, 1023, 820, 290, 2590, 1377, 1724, 1405, 1982, 2670, 1301]. World-View [2743, 1886]. Worlds [899, 2082, 2304, 1760, 964, 1976, 2124]. Worm [269]. Worrall [152]. Worse [2930]. Worst [2302]. Worth [909]. Worthington [1359]. Woycke [2932]. Wozniak [1639]. Wranglers [1989]. Wren [1157]. Wright [2208, 2394, 2364, 324, 847, 2542, 2559, 1570]. Wrightianae [2364]. Writer [1691, 1355]. Writing [2069, 2478, 158]. Writings [2456, 2453, 30, 859, 1849, 852, 382, 951, 1466, 341, 1448, 435, 341, 1555, 1864, 928, 1951]. Written [48, 2013, 2883, 1855, 179, 374]. Wrobel [2517]. Wrong [2690]. Wundt [1081]. Wussing [1141]. Wygant [2014]. Wyn [795]. Wynarczyk [2938]. Wynne [1194]. Wyss [1906].

X [1316, 1791, 833, 320, 187, 422]. X-Ray [1316, 1791]. X-Rays [422]. X.36 [1508]. Xantus [1993]. Xavier [2914]. XI [1686, 504, 2138]. XI-XIV [504]. XIe [237]. XII [2901]. XIIe [2814]. XIIIe [975, 979, 438]. XIIIe-XIVe [975]. XIV [578, 504]. XIVe [975, 1342]. XIVe-XVe [1342]. XIVth [1466].

XIVth-XVIth [1466]. XIX [281, 285, 2549, 2751]. XIX-XX [2549]. XIXe [1495, 1788, 315, 625, 1375, 2170]. XIXe-XXe [2170]. XLIII [837]. XV [1782, 2816, 1846]. XV.2 [1508]. XVe [1342]. XVI [1782, 2592, 2129, 849, 443, 1781]. XVIe [1495, 438]. XVIe-XIXe [1495]. XVII [2051, 2903, 1860, 443]. XVIIe [2810]. XVIII [74, 1753]. XVIIIe [291, 1602, 651]. XVIth [1466]. XX [74, 2846, 2549]. XXe [2170].

Ya [990]. Yakov [2841]. Yale [109, 1795]. Yanchi [2450]. Yankee [2146, 914]. Yapp [782]. Yataro [1997]. Yates [672, 470, 2602]. Yavanajataka [40]. year [738]. Years [344, 2752, 1032, 2223, 2138, 1484, 1096, 2281, 2458, 2032, 2030, 1022, 18, 2057, 2923, 2247, 659, 1642, 360, 2765, 2654, 1292, 1024, 1151, 2844, 2250, 1899, 2106, 6, 392, 653, 732, 705, 548, 2086, 2018]. yeh [571]. Yehuda [1193, 133]. Yellow [2332]. Yeo [2016]. Yeomans [1752]. Yervant [1575]. Yesterday [557]. ying [1205]. Yochelson [1163]. Yolanda [1917]. Yolton [1717]. Yonder [2607]. Yoram [2252]. York [969, 300, 1900, 1934, 4]. Young [2777, 1826, 689, 459, 1994]. Younger [2401]. yu [377, 571]. Yurii [2832]. Yuval [2252]. Yuzhe [1102]. Yves [2725, 617, 1496]. Yves-Marie [1496]. Yvette [1966].

Z [2194, 707]. Zahlentheorie [1278]. Zahrawi [2586]. Zak [2469]. Zandvoort [2472]. Zaniello [2565]. Zanier [1442]. Zarko [1462, 2909]. Zealand [2834]. Zeichnungen [1379]. Zeigler [1864]. Zeit [1019, 1008]. Zeitalter [1469, 497]. Zeitproblemen [897]. Zellphysiologer [134]. Zenon [1274]. Zenryu [1599]. Zeppelin [2117]. Zermelo [1160]. Zest [866]. Zev [1041]. Zhang [1102]. Zhongguo [1102]. Ziadat [2048]. Ziggelaar [1783]. zijat [1439]. Zilsel [1468]. Ziman [211, 2336]. Zimmermann [2819, 1659]. Zivot [1352]. Zloczower [1217]. znanosti [1462]. Zodiac [1346]. Zoological [2834, 1580]. Zoologists [951]. Zoology [2103]. Zott [2702]. zraka [2739]. Zsamboki [473]. Zucker [2285]. zum [91, 762, 424, 2817, 553, 1019, 2199, 2807]. Zumbach [1718]. Zupko [21]. Zur [762, 1469, 1017, 2726, 1186, 1055, 2738, 836, 1847, 1275, 2950, 45, 75, 556, 503, 706, 908, 2720, 1830, 479, 785, 2897, 310, 1776, 353, 1008, 1572, 1640, 1945, 2031, 1650]. Zurich [1475, 479, 1475]. Zuroku [1599]. zusammenkommen [1908]. zwei [1908]. Zwinger [1993]. zwischen [1184, 385, 2830, 547, 2132]. ’Zyciński [2598, 1784, 2884]. Zysk [2049]. Zytkow [2656].

References

Anonymous:1980:FMa

[1] Anonymous. Front matter. Isis, 71(1):1–6, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

      /www.jstor.org/stable/230310.

Thackray:1980:FA

[2] Arnold Thackray. First anniversary. Isis, 71(1):7–10, March 1980. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230311.

Jacyna:1980:SSO

[3] L. S. Jacyna. Science and social order in the thought of A. J. Balfour. Isis, 71(1):11–34, March 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230312.

Sloan:1980:SNY

[4] Douglas Sloan. Science in New York City, 1867–1907. Isis, 71(1):35– 76, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230313.

Lenoir:1980:KBV

[5] Timothy Lenoir. Kant, Blumenbach, and vital materialism in German biology. Isis, 71(1):77–108, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230314.

Westfall:1980:NMY

[6] Richard S. Westfall. Newton’s marvelous years of discovery and their aftermath: Myth versus manuscript. Isis, 71(1):109–121, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230315.

Kahn:1980:OPS

[7] David Kahn. On the origin of polyalphabetic substitution. Isis, 71(1): 122–127, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230316.

Lemay:1980:SHS

[8] Helen Lemay. The stars and human sexuality: Some Medieval scientific views. Isis, 71(1):127–137, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230317.

Hall:1980:IHS

[9] A. Rupert Hall and Albert Van Helden. International Huygens Symposium, Amsterdam, August 22–25, 1979. Isis, 71(1):138–139, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230318.

Mikulinsky:1980:NSJ

[10] S. R. Mikulinsky. New Soviet journal announced. Isis, 71(1):139, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230319.

Bearman:1980:BRB

[11] David Bearman. Book review: The Eighth Day of Creation: The Markers of the Revolution in Biology. Isis, 71(1):140–142, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230320.

Glass:1980:BRB

[12] Bentley Glass. Book review: Thomas Hunt Morgan: The Man and His Science by Garland E. Allen. Isis, 71(1):143–146, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230321.

Shapin:1980:BRB

[13] Steven Shapin. Book review: Michael Foster and the Cambridge School of Physiology. The Scientific Enterprise in Late Victorian Society by Gerald L. Geison. Isis, 71(1):146–149, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230322.

Coleman:1980:BRB

[14] William Coleman. Book review: History of American Ecology by Frank N. Egerton. Isis, 71(1):150–152, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230323.

Barr:1980:BRB

[15] E. Scott Barr. Book review: Biographical Dictionary of American Science: The Seventeenth through the Nineteenth Centuries by Clark A. Elliott. Isis, 71(1):153–154, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230324.

Decker:1980:BRB

[16] Hannah S. Decker. Book review: Studies in Intellectual Breakthrough: Freud, Simmel, Buber by Charles David Axelrod. Isis, 71(1):154, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230325.

Kourany:1980:BRB

[17] Janet A. Kourany. Book review: A Realist Theory of Science by Roy Bhaskar. Isis, 71(1):154–155, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230326.

Kohlstedt:1980:BRBa

[18] Sally Gregory Kohlstedt. Book review: The National Academy of Sciences: The First Hundred Years, 1863–1963 by Rexmond C. Cochrane. Isis, 71(1):155–157, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230327.

Weingart:1980:BRB

[19] Peter Weingart. Book review: Deutsche Gesellschaft für Geschichte der Medizin, Naturwissenschaft und Technik 1901–1976 by Rolf Winau. Isis, 71(1):157, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230328.

Pyenson:1980:BRBa

[20] Lewis Pyenson. Book review: The Beginning of the Long Dash: A History of Timekeeping in Canada by Malcolm M. Thomson. Isis, 71(1): 157–158, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230329.

McClellan:1980:BRB

[21] James E. McClellan. Book review: French Weights and Measures before the Revolution: A Dictionary of Provincial and Local Units by Ronald Edward Zupko. Isis, 71(1):158–159, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230330.

Brock:1980:BRB

[22] William H. Brock. Book review: Education and Society in Modern Europe by Fritz K. Ringer. Isis, 71(1):159, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230331.

Collins:1980:BRBa

[23] H. M. Collins. Book review: Interests and the Growth of Knowledge by Barry Barnes. Isis, 71(1):159–160, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230332.

Goldstine:1980:BRB

[24] Herman H. Goldstine. Book review: Annals of the History of Computing by Bernard A. Galler. Isis, 71(1):160, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230333.

Pedersen:1980:BRB

[25] Olaf Pedersen. Book review: Geared to the Stars: The Evolution of Planetariums, Orreries and Astronomical Clocks by Henry C. King; John R. Millburn. Isis, 71(1):160–161, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230334.

Stachel:1980:BRB

[26] John Stachel. Book review: Quantum Logic by Peter Mittelstaedt. Isis, 71(1):162, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230335.

Lamb:1980:BRB

[27] Ursula Lamb. Book review: Geschichte der Navigation by Hans-Christian Freiesleben. Isis, 71(1):162–163, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230336.

Winsor:1980:BRB

[28] Mary P. Winsor. Book review: The Spontaneous Generation Controversy from Descartes to Oparin by John Farley. Isis, 71(1):163–164, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230337.

Appel:1980:BRB

[29] Toby A. Appel. Book review: Studies in History of Biology. Volume 2 by William Coleman; Camille Limoges. Isis, 71(1):164–165, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230338.

Diamond:1980:BRBa

[30] Solomon Diamond. Book review: Basic Writings in the History of Psychology by Robert I. Watson. Isis, 71(1):165, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230339.

Cole:1980:BRB

[31] Jack W. Cole. Book review: Forgive and Remember: Managing Medical Failure by Charles L. Bosk. Isis, 71(1):165–166, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230340.

Farley:1980:BRB

[32] John Farley. Book review: Disease in African History: An Introductory Survey and Case Studies by Gerald W. Hartwig; K. David Patterson. Isis, 71(1):166–167, March 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230341.

Norberg:1980:BRB

[33] Arthur L. Norberg. Book review: Saga of the Vacuum Tube by Gerald F. J. Tyne; Revolution in Miniature: The History and Impact of Semiconductor Electronics by Ernest Braun. Isis, 71(1):167–168, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230342.

Basalla:1980:BRBa

[34] George Basalla. Book review: The American Railroad Passenger Car by John H. White. Isis, 71(1):168–169, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230343.

Layton:1980:BRBa

[35] Edwin T. Layton, Jr. Book review: Autonomous Technology: Technicsout-of-Control as a Theme in Political Thought by Langdon Winner. Isis, 71(1):169, March 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230344.

Scarborough:1980:BRBa

[36] John Scarborough. Book review: Antike Heilkunst. Ausgewählte Texte aus dem medizinischen Schrifttum der Griechen und Römer by Jutta Kollesch; Diethard Nickel. Isis, 71(1):169–170, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230345.

Mahoney:1980:BRB

[37] Michael S. Mahoney. Book review: Archimedes in the Middle Ages. Volume III: The Fate of the Medieval Archimedes, 1300 to 1565 by Marshall Clagett. Isis, 71(1):170–171, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230346.

Sylla:1980:BRBa

[38] Edith Sylla. Book review: Lexikon des Mittelalters. Erster Band/Erste Lieferung. Aachen-Agypten. Isis, 71(1):171–172, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230347.

Savage-Smith:1980:BRBa

[39] Emilie Savage-Smith. Book review: Islamic Medicine by Manfred Ullmann. Isis, 71(1):172–173, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230348.

Rocher:1980:BRB

[40] Ludo Rocher. Book review: The Yavanajataka of Sphujidhvaja by David Pingree. Isis, 71(1):173–174, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230349.

Zetterberg:1980:BRBa

[41] J. Peter Zetterberg. Book review: John Dee on Astronomy: Propaedeumata aphoristica (1558 and 1568), Latin and English by Wayne Shumaker. Isis, 71(1):175, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230350.

Shea:1980:BRBa

[42] William Shea. Book review: Giambattista Della Porta mago e scienziato by Luisa Muraro. Isis, 71(1):175–176, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230351.

Sturchio:1980:BRBa

[43] Jeffrey L. Sturchio. Book review: Philosophical Chemistry in the Scottish Enlightenment: The Doctrines and Discoveries of William Cullen and Joseph Black by Arthur L. Donovan. Isis, 71(1):176–177, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230352.

Drake:1980:BRBa

[44] Stillman Drake. Book review: Le opere dei discepoli di Galileo Galilei. Volume I: Carteggio 1642–1648 by Paolo Galluzzi; Maurizio Torrini. Isis, 71(1):177–178, March 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230353.

Gould:1980:BRB

[45] Henry W. Gould. Book review: Die mathematischen Studien von G. W. Leibniz zur Kombinatorik by Eberhard Knobloch. Isis, 71(1):178, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230354.

Aiton:1980:BRBa

[46] E. J. Aiton. Book review: Leibniz à Paris (1672–1676). Symposion de la G. W. Leibniz-Gesellschaft (Hannover) et du Centre National de la Recherche Scientifique (Paris) à Chantilly (France) du 14 au 18 Novembre 1976. Isis, 71(1):178–179, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230355.

Ashworth:1980:BRB

[47] William B. Ashworth, Jr. Book review: Dr. Woodward’s Shield: History, Science, and Satire in Augustan England by Joseph M. Levine. Isis, 71 (1):179–180, March 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230356.

Burkhardt:1980:BRB

[48] Frederick Burkhardt. Book review: The Beagle Record: Selections from the Original Pictorial Records and Written Accounts of the Voyage of H.M.S. Beagle by Richard Darwin Keynes. Isis, 71(1):180–181, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230357.

Aldrich:1980:BRB

[49] Michele L. Aldrich. Book review: Charles Lyell on North American Geology by Hubert C. Skinner; Jules Marcou on the Taconic System of North America by Hubert C. Skinner. Isis, 71(1):181, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230358.

Garber:1980:BRB

[50] Elizabeth Garber. Book review: Lord Kelvin: The Dynamic Victorian by Harold Issadore Sharlin; Tiby Sharlin. Isis, 71(1):182–183, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230359.

Burchfield:1980:BRBa

[51] Joe D. Burchfield. Book review: James Prescott Joule and the Concept of Energy by Henry John Steffens. Isis, 71(1):183, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230360.

Hallion:1980:BRB

[52] Richard P. Hallion. Book review: The Jet Makers: The Aerospace Industry from 1945 to 1972 by Charles D. Bright. Isis, 71(1):183–184, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230361.

Ghiselin:1980:BRB

[53] Michael T. Ghiselin. Book review: The Kindly Fruits of the Earth: Recollections of an Embryo Ecologist by G. Evelyn Hutchinson. Isis, 71(1):184–185, March 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230362.

Moyer:1980:BRBa

[54] Albert E. Moyer. Book review: NOVA: Einstein by Patrick Griffin. Isis, 71(1):185, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230363.

Milton:1980:BRB

[55] Sybil Milton. Book review: Guide to the Archival Materials of the German-Speaking Emigration to the United States after 1933 by John M. Spalek; Adrienne Ash; Sandra H. Hawrylchak. Isis, 71(1):186, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230364.

Crosland:1980:BRB

[56] Maurice Crosland. Book review: Scientists in Power by Spencer R. Weart. Isis, 71(1):186–187, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230365.

Anonymous:1980:BMa

[57] Anonymous. Back matter. Isis, 71(1):188–190, March 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230366.

Anonymous:1980:FMb

[58] Anonymous. Front matter. Isis, 71(2):191–195, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230170.

Hawkins:1980:TDT

[59] Hugh Hawkins. Transatlantic discipleship: Two American biologists and their German mentor. Isis, 71(2):196–210, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230171.

Shapiro:1980:ESN

[60] Alan E. Shapiro. The evolving structure of Newton’s theory of white light and color. Isis, 71(2):211–235, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230172.

Vucinich:1980:SPP

[61] Alexander Vucinich. Soviet physicists and philosophers in the 1930s: Dynamics of a conflict. Isis, 71(2):236–250, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230173.

Jacob:1980:AOM

[62] James R. Jacob and Margaret C. Jacob. The Anglican origins of modern science: The metaphysical foundations of the Whig Constitution. Isis, 71(2):251–267, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230174.

Stevens:1980:FEI

[63] Wesley M. Stevens. The figure of the Earth in Isidore’s “de natura rerum”. Isis, 71(2):268–277, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230175.

Donovan:1980:AMH

[64] Arthur Donovan and Sally Gregory Kohlstedt. Annual meeting of the History of Science Society, December 27–31, 1979. Isis, 71(2):278–284, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230176.

Shapin:1980:CR

[65] Steven Shapin and N. Sivin. Conference reports. Isis, 71(2):284–285, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230177.

Edsall:1980:EMF

[66] John T. Edsall. Éloge: Marcel Florkin, August 15, 1900–May 3, 1979. Isis, 71(2):286–288, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230178.

Hannaway:1980:BRB

[67] Owen Hannaway. Book review: Neptune’s Gift: A History of Common Salt. Isis, 71(2):289–291, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230179.

Rosenberg:1980:BRB

[68] Charles E. Rosenberg. Book review: Natural Order: Historical Studies of Scientific Culture. Isis, 71(2):291–295, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230180.

Molland:1980:BRBa

[69] A. G. Molland. Book review: Science in the Middle Ages. Isis, 71(2): 295–298, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230181.

Levere:1980:BRB

[70] Trevor H. Levere. Book review: Gay-Lussac: Scientist and Bourgeois. Isis, 71(2):298–300, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230182.

Hahn:1980:BRB

[71] Roger Hahn. Book review: Le siècle des lumières en province. Académies et académiciens provinciaux, 1680–1789 by Daniel Roche. Isis, 71(2): 301–302, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230183.

Pickering:1980:BRB

[72] Andrew Pickering. Book review: The Scientific Imagination: Case Studies by Gerald Holton. Isis, 71(2):302–303, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230184.

VanHelden:1980:BRBa

[73] Albert Van Helden. Book review: Science and History: Studies in Honor of Edward Rosen by Erna Hilfstein; Pawel Czartoryski; Frank D. Grande. Isis, 71(2):303–304, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230185.

Brooks:1980:BRB

[74] Nathan M. Brooks. Book review: Razvitie estestvoznaniia v Rossii (XVIII–nachalo XX veka) by S. R. Mikulinskii; A. P. Iushkevich. Isis, 71(2):304–305, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230186.

Gregory:1980:BRBa

[75] Frederick Gregory. Book review: Berichte zur Wissenschaftsgeschichte by Fritz Krafft. Isis, 71(2):305–306, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230187.

Barton:1980:BRB

[76] Ruth Barton. Book review: Science, Technology, and Society in Australia: A Bibliography by Ann Mozley Moyal; Elizabeth Newland; Irene Davey. Isis, 71(2):306, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230188.

Finocchiaro:1980:BRB

[77] Maurice A. Finocchiaro. Book review: Epistemologia e storia della scienza: Le svolte teoriche da Duhem a Bachelard by Pietro Redondi. Isis, 71(2):306–307, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230189.

Felt:1980:BRB

[78] James W. Felt. Book review: Whitehead’s Organic Philosophy of Science by Ann L. Plamondon. Isis, 71(2):307, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230190.

Basalla:1980:BRBb

[79] George Basalla. Book review: Science and Its Critics by John Passmore. Isis, 71(2):308, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230191.

Carroll:1980:BRB

[80] P. Thomas Carroll. Book review: A Century of Doctorates: Data Analyses of Growth and Change: U. S. PhD’s — Their Numbers, Origins, Characteristics, and the Institutions from Which They Come by National Research Council; Lindsey R. Harmon. Isis, 71(2):308–309, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230192.

Beer:1980:BRB

[81] John J. Beer. Book review: The Building of a Department: Chemistry at Southern Illinois University, 1927–1967 by James W. Neckers. Isis, 71 (2):309, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230193.

Schutt:1980:BRB

[82] Hans-Werner Schutt. Book review: Die Chemie an der Universitat Marburg seit Beginn des 19. Jahrhunderts. Ein Beitrag zu ihrer Entwicklung als Hochschulfach by Christoph Meinel. Isis, 71(2):309–311, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230194.

Nelles:1980:BRB

[83] H. V. Nelles. Book review: La science et le pouvoir au Québec, 1920– 1965 by Raymond Duchesne. Isis, 71(2):311–312, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230195.

Cowan:1980:BRB

[84] Ruth Schwartz Cowan. Book review: Beyond Her Sphere: Women and the Professions in American History by Barbara J. Harris. Isis, 71(2): 312, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230196.

Overington:1980:BRB

[85] Michael A. Overington. Book review: Research in Sociology of Knowledge, Sciences and Art: An Annual Compilation of Research. Volume I, 1978 by Robert Alun Jones. Isis, 71(2):312–313, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230197.

Paul:1980:BRB

[86] H. W. Paul. Book review: Les Scientifiques et la paix. La Communauté scientifique internationale au cours des années 20 by Brigitte Schroeder-Gudehus. Isis, 71(2):313–314, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230198.

MacKenzie:1980:BRB

[87] Donald MacKenzie. Book review: The Visible College: The Collective Biography of British Scientific Socialists in the 1930s by Gary Werskey. Isis, 71(2):314–315, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230199.

Deason:1980:BRB

[88] Gary Bruce Deason. Book review: The Road of Science and the Ways to God by Stanley L. Jaki. Isis, 71(2):315–316, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230200.

Kitcher:1980:BRB

[89] Philip Kitcher. Book review: Nombre, mesure et continu: Epistémologie et histoire by Jean Dhombres. Isis, 71(2):316–317, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230201.

Wheaton:1980:BRB

[90] Bruce R. Wheaton. Book review: Rutherford and Physics at the Turn of the Century by Mario Bunge; William R. Shea. Isis, 71(2):317–318, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230202.

Aiton:1980:BRBb

[91] E. J. Aiton. Book review: Geschichte der Physik: Renaissance bis zum 18 Jahrhundert by Hans Kangro. Isis, 71(2):318–319, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230203.

Sturchio:1980:BRBb

[92] Jeffrey L. Sturchio. Book review: Classics in Coordination Chemistry. Part III: Twentieth–Century Papers (1904–1935) by George B. Kauffman. Isis, 71(2):319–320, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230204.

Donovan:1980:BRBa

[93] Arthur Donovan. Book review: James Hutton’s Theory of the Earth: The Lost Drawings by G. Y. Craig; D. B. McIntyre; C. D. Waterston.

Isis, 71(2):320–321, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230205.

Burchfield:1980:BRBb

[94] Joe D. Burchfield. Book review: Ice Ages: Solving the Mystery by John Imbrie; Katherine Palmer Imbrie. Isis, 71(2):321–322, June 1980. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230206.

Manning:1980:BRB

[95] Thomas G. Manning. Book review: Minerals, Lands, and Geology for the Common Defence and General Welfare. Volume I: Before 1879. United States Geological Survey. A History of Public Lands, Federal Science and Mapping Policy, and Development of Mineral Resources in the United States by Mary C. Rabbitt. Isis, 71(2):322–323, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230207.

Erickson:1980:BRB

[96] Paul A. Erickson. Book review: The White Man’s Indian: Images of the American Indian from Columbus to the Present by Robert F. Berkhofer. Isis, 71(2):323–324, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230208.

Woolf:1980:BRB

[97] Patricia Woolf. Book review: The Emergence of Professional Social Science: The American Social Science Association and the Nineteenth-Century Crisis of Authority by Thomas L. Haskell. Isis, 71(2):324, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230209.

Furner:1980:BRB

[98] Mary O. Furner. Book review: Robert E. Park: Biography of a Sociologist by Winifred Raushenbush. Isis, 71(2):325, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230210.

Richards:1980:BRB

[99] Robert J. Richards. Book review: An Intellectual History of Psychology by Daniel N. Robinson; The Mind Unfolded: Essays on Psychology’s Historic Texts by Daniel N. Robinson. Isis, 71(2):325–326, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230211.

Langham:1980:BRB

[100] Ian G. Langham. Book review: W. H. R. Rivers by Richard Slobodin. Isis, 71(2):327, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230212.

King:1980:BRB

[101] Lester S. King. Book review: A Short Title Catalogue of Eighteenth Century Printed Books in the National Library of Medicine by John B. Blake. Isis, 71(2):327–328, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230213.

Smith:1980:BRB

[102] Roger Smith. Book review: Museums of Madness: The Social Organization of Insanity in Nineteenth-Century England by Andrew T. Scull. Isis, 71(2):328, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230214.

McLean:1980:BRB

[103] Donald M. McLean. Book review: An Introduction to the History of Virology by A. P. Waterson; Lise Wilkinson. Isis, 71(2):328–329, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230215.

Lampert:1980:BRB

[104] N. Lampert. Book review: Technology and Society under Lenin and Stalin: Origins of the Soviet Technical Intelligentsia, 1917–1941 by Kendall E. Bailes. Isis, 71(2):329–330, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230216.

Rae:1980:BRB

[105] John B. Rae. Book review: Interstate: Express Highway Politics, 1941– 1956 by Mark H. Rose. Isis, 71(2):330–331, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230217.

Kevles:1980:BRB

[106] Daniel J. Kevles. Book review: A History of Technology. Volume VI: The Twentieth Century, c. 1900 to c. 1950, Part I by Trevor I. Williams; A History of Technology. Volume VII: The Twentieth Century, c. 1900 to c. 1950, Part II. Isis, 71(2):331, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230218.

Walker:1980:BRB

[107] D. P. Walker. Book review: Symphorien Champier and the Reception of the Occultist Tradition in Renaissance France by Brian P. Copenhaver; Darrel Amundsen. Isis, 71(2):331–332, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230219.

Jobe:1980:BRB

[108] Thomas H. Jobe. Book review: Magia naturalis und die Enstehung der modernen Naturwissenschaften: Symposion der Leibniz-Gesellschaft Hannover, 14. und 15. November 1975 by Albert Heinekamp; Dieter Mettler. Isis, 71(2):333, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230220.

Multhauf:1980:BRB

[109] Robert P. Multhauf. Book review: Alchemy and the Occult: A Catalogue of Books and Manuscripts from the Collection of Paul and Mary Mellon Given to the Yale University Library by Laurence C. Witten; Richard Pachella. Isis, 71(2):333–334, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230221.

Scarborough:1980:BRBb

[110] John Scarborough. Book review: On the Doctrines of Hippocrates and Plato by Phillip De Lacy; Galen of Pergamon. Isis, 71(2):334–335, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230222.

Vogel:1980:BRB

[111] Kurt Vogel. Book review: Introduction arithmétique by Nichomachus of Gerasa; Janine Bertier. Isis, 71(2):336, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230223.

Corsi:1980:BRB

[112] Pietro Corsi. Book review: Cuore, sangue e cervello: Biologia e antropologia nel pensiero antico by Paola Manuli; Mario Vegetti. Isis, 71(2):336–337, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230224.

DeLacy:1980:BRB

[113] Phillip De Lacy. Book review: The Hippocratic Tradition by Wesley D. Smith. Isis, 71(2):337–338, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230225.

Wilson:1980:BRB

[114] N. G. Wilson. Book review: Le “Petit commentaire” de Théon d’Alexandrie aux tables faciles de Ptolémée by Theon of Alexandria; Anne Tihon. Isis, 71(2):339–340, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230226.

Grant:1980:BRB

[115] Edward Grant. Book review: Reason and Society in the Middle Ages by Alexander Murray. Isis, 71(2):340–341, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230227.

Goldstein:1980:BRB

[116] Bernard R. Goldstein. Book review: Geschichte des arabischen Schrifttums. Volume VI: Astronomie bis ca. 430 H by Fuat Sezgin. Isis, 71(2): 341–342, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230228.

Heinen:1980:BRB

[117] Anton Heinen. Book review: Aristoteles’ Meteorologie in arabischer und lateinischer Uebersetzung. Textkritische Ausgabe des ersten Buches by Pieter Leendert Schoonheim. Isis, 71(2):342–343, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230229.

Libbrecht:1980:BRB

[118] Ulrich Libbrecht. Book review: The Shorter Science and Civilization in China by Joseph Needham; Colin A. Ronan. Isis, 71(2):343, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230230.

Croizier:1980:BRB

[119] Ralph Croizier. Book review: Medical Ethics in Imperial China: A Study in Historical Anthropology by Paul U. Unschuld. Isis, 71(2):343–344, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230231.

Hilfstein:1980:BRB

[120] Erna Hilfstein. Book review: Mikolaj Kopernik w ’sweietle swej korespondencji by Jerzy Drewnowski. Isis, 71(2):344–345, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230232.

Morgan:1980:BRB

[121] Bruce T. Morgan. Book review: From Humanism to Science 1480–1700 by Robert Mandrou; Brian Pearce. Isis, 71(2):345–346, June 1980. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230233.

Kish:1980:BRB

[122] George Kish. Book review: New American World: A Documentary History of North America to 1612 by David B. Quinn; Alison M. Quinn; Susan Hillier. Isis, 71(2):346–347, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230234.

Drake:1980:BRBb

[123] Stillman Drake. Book review: Studies in Musical Science in the Late Renaissance by D. P. Walker. Isis, 71(2):347–348, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230235.

Rousseau:1980:BRBa

[124] G. S. Rousseau. Book review: Probability, Time, and Space in Eighteenth-Century Literature by Paula R. Backscheider. Isis, 71(2): 348–349, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230236.

Foust:1980:BRB

[125] Clifford M. Foust. Book review: Bering’s Voyages: Whither and Why by Raymond H. Fisher. Isis, 71(2):349–350, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230237.

Sterling:1980:BRB

[126] Keir B. Sterling. Book review: Spencer Fullerton Baird and the U. S. Fish Commission by Dean Conrad Allard; Ernest Thompson Seton: Man in Nature and the Progressive Era 1880–1915 by John Henry Wadland.

Isis, 71(2):350–351, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230238.

Hollinger:1980:BRBa

[127] David A. Hollinger. Book review: Social Darwinism: Science and Myth in Anglo–American Social Thought by Robert C. Bannister. Isis, 71(2): 351–352, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230239.

Mengel:1980:BRB

[128] Robert M. Mengel. Book review: Audubon: A Biography by John Chancellor. Isis, 71(2):352–353, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230240.

Barrett:1980:BRB

[129] Paul Howard Barrett. Book review: Charles Darwin: A Companion by R. B. Freeman. Isis, 71(2):353, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230241.

Rogers:1980:BRB

[130] James Allen Rogers. Book review: Folia Baeriana by Tovmas Ia. Sutt. Isis, 71(2):353–354, June 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230242.

Hanle:1980:BRB

[131] Paul A. Hanle. Book review: My Life. Recollections of a Nobel Laureate, by Max Born. Isis, 71(2):354–355, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230243;   http://www.jstor.org/stable/
pdfplus/230243.pdf.

Beyerchen:1980:BRBa

[132] Alan Beyerchen. Book review: Memoirs of a Physicist in the Atomic Age by Walter M. Elsasser. Isis, 71(2):355–356, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230244.

Pyenson:1980:BRBb

[133] Lewis Pyenson. Book review: Einstein. A Centenary Exhibition by Paul Forman; Paul Hanle; Albert Einstein in Berlin 1913–1933 by Christa Kirsten; Hans-Jurgen Treder; Images of Einstein: A Catalogue by Joan Warnow; Einstein 1897–1979. Exhibition by Yehuda Elkana; Adi Ophir; Einstein. A Centenary Volume by A. P. French; Albert Einstein. His Influence on Physics, Philosophy and Politics by P. C. Aichelburg; R. U. Sexl. Isis, 71(2):356–359, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230245.

Fruton:1980:BRB

[134] Joseph S. Fruton. Book review: Otto Warburg: Zellphysiologer, Biochemiker, Mediziner, 1883–1970 by Hans Krebs. Isis, 71(2):359–360, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230246.

Badash:1980:BRB

[135] Lawrence Badash. Book review: The Uranium People by Leona Marshall Libby. Isis, 71(2):360–361, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230247.

Brittain:1980:BRB

[136] James E. Brittain. Book review: The Rise of Systems Theory: An Ideological Analysis by Robert Lilienfeld. Isis, 71(2):361–362, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230248.

Pitts:1980:BRB

[137] John A. S. Pitts. Book review: Psychoanalysis, Psychotherapy, and the New England Medical Scene 1894–1944 by George E. Gifford; American Psychoanalysis: Origins and Development by Jacques M. Quen; Eric T. Carlson. Isis, 71(2):362–363, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230249.

Anonymous:1980:BMb

[138] Anonymous. Back matter. Isis, 71(2):364–368, June 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230250.

Anonymous:1980:FMc

[139] Anonymous. Front matter. Isis, 71(3):369–373, September 1980. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230116.

Schutt:1980:GEH

[140] Hans-Werner Schütt. Guest editorial: History of science in the Federal Republic of Germany. Isis, 71(3):374–380, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230117.

Rossiter:1980:WWS

[141] Margaret W. Rossiter. “Women’s work” in science, 1880–1910. Isis, 71 (3):381–398, September 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230118.

Pyenson:1980:EEM

[142] Lewis Pyenson. Einstein’s education: Mathematics and the laws of nature. Isis, 71(3):399–425, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230119.

Jones:1980:ACG

[143] Daniel P. Jones. American chemists and the Geneva protocol. Isis, 71 (3):426–440, September 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230120.

Kevles:1980:GUS

[144] Daniel J. Kevles. Genetics in the United States and Great Britain, 1890– 1930: A review with speculations. Isis, 71(3):441–455, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230121.

Mulligan:1980:PES

[145] Lotte Mulligan. Puritans and English science: A critique of Webster. Isis, 71(3):456–469, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230122.

Jones:1980:BRB

[146] Robert Alun Jones. Book review: A History of Sociological Analysis. Isis, 71(3):470–473, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230123.

Westman:1980:BRB

[147] Robert S. Westman. Book review: The Printing Press as an Agent of Change: Communications and Cultural Transformations in Early-Modern Europe. Isis, 71(3):474–477, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230124.

Hollinger:1980:BRBb

[148] David A. Hollinger. Book review: The Organization of Knowledge in Modern America, 1860–1920. Isis, 71(3):478–480, September 1980. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230125.

Jacob:1980:BRB

[149] Margaret C. Jacob. Book review: The Business of Enlightenment: A Publishing History of the Encyclopédie, 1775–1800 by Robert Darnton. Isis, 71(3):481–482, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230126.

Haskell:1980:BRB

[150] Thomas L. Haskell. Book review: Journal of the History of Sociology: An International Review by Glenn Jacobs. Isis, 71(3):482–483, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230127.

Kauffman:1980:BRB

[151] George B. Kauffman. Book review: A Sense of the Future: Essays in Natural Philosophy by J. Bronowski; Piero E. Ariotti; Rita Bronowski; The Visionary Eye: Essays in the Arts, Literature, and Science by Jakob Bronowski; Piero E. Ariotti; Rita Bronowski. Isis, 71(3):483–484, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230128.

Asquith:1980:BRB

[152] Peter D. Asquith. Book review: Philosophical Papers by Imre Lakatos; John Worrall; Gregory Currie. Isis, 71(3):484–486, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230129.

Gooding:1980:BRB

[153] David C. Gooding. Book review: The Concept of Matter in Modern Philosophy by Ernan McMullin. Isis, 71(3):486, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230130.

Beyerchen:1980:BRBb

[154] Alan Beyerchen. Book review: Ernst Krieck und die nationalsozialistische Wissenschaftsreform: Motive und Tendenzen einer Wissenschaftslehre und Hochschulreform im Dritten Reich by Gerhard Müller. Isis, 71 (3):486–487, September 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230131.

Collins:1980:BRBb

[155] H. M. Collins. Book review: The Sociology of Science: Problems, Approaches, and Research by Jerry Gaston. Isis, 71(3):487–488, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230132.

Haraway:1980:BRB

[156] Donna Haraway. Book review: Laboratory Life: The Social Construction of Scientific Facts by Bruno Latour; Steve Woolgar. Isis, 71(3):488–489, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230133.

Mullins:1980:BRB

[157] Nicholas Mullins. Book review: Science and the Sociology of Knowledge by Michael Mulkay. Isis, 71(3):489–490, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230134.

Rousseau:1980:BRBb

[158] G. S. Rousseau. Book review: Writing about Science by Mary Elizabeth Bowen; J. A. Mazzeo. Isis, 71(3):490, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230135.

Magner:1980:BRB

[159] Lois N. Magner. Book review: Science as a Human Endeavor by George F. Kneller. Isis, 71(3):490–491, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230136.

Numbers:1980:BRB

[160] Ronald L. Numbers. Book review: The Post-Darwinian Controversies: A Study of the Protestant Struggle to Come to Terms with Darwin in Great Britain and America, 1870–1900 by James R. Moore. Isis, 71(3):

491–492, September 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230137.

Farber:1980:BRB

[161] Paul Lawrence Farber. Book review: Diderot and Goethe: A Study in Science and Humanism by Gerhard M. Vasco. Isis, 71(3):492–493, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230138.

Lewis:1980:BRB

[162] Albert C. Lewis. Book review: Georg Cantor: His Mathematics and Philosophy of the Infinite by Joseph Warren Dauben. Isis, 71(3):493–494, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230139.

Donovan:1980:BRBb

[163] Arthur Donovan. Book review: The Playfair Collection and the Teaching of Chemistry at the University of Edinburgh 1713–1858 by R. G. W. Anderson. Isis, 71(3):494, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230140.

Servos:1980:BRB

[164] John W. Servos. Book review: The Transcendental Part of Chemistry by David M. Knight. Isis, 71(3):494–495, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230141.

Guedon:1980:BRB

[165] Jean-Claude Guédon. Book review: Chemistry Transformed: The Paradigmatic Shift from Phlogiston to Oxygen by H. Gilman McCann. Isis, 71(3):495–497, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230142.

Glick:1980:BRB

[166] Thomas F. Glick and Jonathan Olsson. Book review: The Nature of Change in Geographical Ideas by Brian J. L. Berry. Isis, 71(3):497–498, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230143.

Laudan:1980:BRB

[167] Rachel Laudan. Book review: Images of the Earth: Essays in the History of the Environmental Sciences by L. J. Jordanova; R. S. Porter. Isis, 71 (3):498–499, September 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230144.

Frisinger:1980:BRB

[168] H. Howard Frisinger. Book review: The Thermal Theory of Cyclones: A History of Meteorological Thought in the Nineteenth Century by Gisela Kutzbach. Isis, 71(3):499–500, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230145.

Bowler:1980:BRB

[169] Peter J. Bowler. Book review: A History of the Life Sciences by Lois N. Magner. Isis, 71(3):500–501, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230146.

ODonnell:1980:BRB

[170] John M. O’Donnell. Book review: Human Science and Social Order: Hugo Münsterberg and the Origins of Applied Psychology by Matthew Hale. Isis, 71(3):501, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230147.

vanKeuren:1980:BRB

[171] D. K. van Keuren. Book review: Orientalism by Edward W. Said. Isis, 71(3):502, September 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230148.

Sokal:1980:BRB

[172] Michael M. Sokal. Book review: The Shaping of a Behaviorist: Part Two of an Autobiography by B. F. Skinner. Isis, 71(3):502–503, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230149.

Whitteridge:1980:BRB

[173] Gweneth Whitteridge. Book review: William Harvey and His Age: The Professional and Social Context of the Discovery of the Circulation by Jerome J. Bylebyl. Isis, 71(3):503–504, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230150.

Liebenau:1980:BRB

[174] Jonathan Liebenau. Book review: Die Anfänge der Basler chemischen Industrie im Lichte von Arbeitsmedizin und Umweltschutz by Markus Hämmerle. Isis, 71(3):504, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230151.

Peterson:1980:BRB

[175] M. Jeanne Peterson. Book review: Cholera, Fever and English Medicine 1825–1865 by Margaret Pelling. Isis, 71(3):504–505, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230152.

Layton:1980:BRBb

[176] Edwin T. Layton, Jr. Book review: The History and Philosophy of Technology by George Bugliarello; Dean B. Doner. Isis, 71(3):505, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230153.

Ovenden:1980:BRB

[177] Michael W. Ovenden. Book review: Megalithic Remains in Britain and Brittany by A. Thom; A. S. Thom. Isis, 71(3):506, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230154.

Knorr:1980:BRBa

[178] Wilbur R. Knorr. Book review: Von Eudoxos zu Aristoteles: Das Fortwirken der Eudoxischen Proportionentheorie in der Aristotelischen Lehre vom Kontinuum by Hans-Joachim Waschkies. Isis, 71(3):506–508, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230155.

Molland:1980:BRBb

[179] A. G. Molland. Book review: From Memory to Written Record: England 1066–1307 by M. T. Clanchy. Isis, 71(3):508, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230156.

Kren:1980:BRB

[180] Claudia Kren. Book review: Machaut’s World: Science and Art in the Fourteenth Century by Madeleine Pelner Cosman; Bruce Chandler. Isis, 71(3):508–509, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230157.

Sivin:1980:BRB

[181] N. Sivin. Book review: Cambridge Texts in the History of Chinese Science on Microfiche. Isis, 71(3):509–511, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230158.

Zetterberg:1980:BRBb

[182] J. Peter Zetterberg. Book review: A Goodly Gallerye: William Fulke’s Book of Meteors by Theodore Hornberger. Isis, 71(3):511–512, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230159.

Hindle:1980:BRB

[183] Brooke Hindle. Book review: Benjamin Thompson, Count Rumford by Sanborn C. Brown. Isis, 71(3):512–513, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230160.

Cohen:1980:BRB

[184] I. Bernard Cohen. Book review: Roemer et la vitesse de la lumière: Paris 16 et 17 juin 1976 by Centre National de la Recherche Scientifique. Isis, 71(3):513–514, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230161.

Brackenridge:1980:BRB

[185] J. Bruce Brackenridge. Book review: Kepler: Astronome, astrologue by Gérard Simon. Isis, 71(3):514–515, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230162.

Goetzmann:1980:BRB

[186] William H. Goetzmann. Book review: Benjamin Silliman and His Circle: Studies on the Influence of Benjamin Silliman on Science in America by Leonard G. Wilson. Isis, 71(3):515–517, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230163.

Schwartz:1980:BRB

[187] Joel S. Schwartz. Book review: Darwin and the Mysterious Mr. X: New Light on the Evolutionists by Loren Eiseley; Kenneth Heuer. Isis, 71(3):

517, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230164.

Seidel:1980:BRB

[188] Robert W. Seidel. Book review: What Little I Remember by Otto R. Frisch. Isis, 71(3):517–518, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230165.

Gowing:1980:BRB

[189] Margaret Gowing. Book review: The Wizard War: British Scientific Intelligence 1939–1945 by R. V. Jones. Isis, 71(3):518–519, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230166.

Dick:1980:BRB

[190] Myvanwy M. Dick. Book review: On Almost Any Wind: The Saga of the Oceanographic Research Vessel “Atlantis” by Susan Schlee. Isis, 71(3): 519, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230167.

Gifford:1980:BRBa

[191] Sanford Gifford. Book review: A History of Aggression in Freud by Paul E. Stepansky. Isis, 71(3):519–520, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230168.

Anonymous:1980:BMc

[192] Anonymous. Back matter. Isis, 71(3):521–524, September 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230169.

Anonymous:1980:FMd

[193] Anonymous. Front matter. Isis, 71(4):525–529, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230498.

Servos:1980:IRS

[194] John W. Servos. The industrial relations of science: Chemical engineering at MIT, 1900–1939. Isis, 71(4):530–549, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230499.

Naylor:1980:GTP

[195] R. H. Naylor. Galileo’s theory of projectile motion. Isis, 71(4):550–570, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230500.

Dean:1980:GIC

[196] Dennis R. Dean. Graham Island, Charles Lyell, and the Craters of Elevation Controversy. Isis, 71(4):571–588, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230501.

Turner:1980:PSB

[197] Frank M. Turner. Public science in Britain, 1880–1919. Isis, 71(4):589– 608, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230502.

Barnouw:1980:VCS

[198] Jeffrey Barnouw. Vico and the continuity of science: The relation of his epistemology to Bacon and Hobbes. Isis, 71(4):609–620, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230503.

Tarbell:1980:SIR

[199] D. S. Tarbell, Ann T. Tarbell, and R. M. Joyce. The students of ira remsen and roger Adams. Isis, 71(4):620–626, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230504.

Hoddeson:1980:CR

[200] Lillian Hoddeson, Nobuo Kawamiya, and Martin Rudwick. Conference reports. Isis, 71(4):627–629, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230505.

Buck:1980:TP

[201] Peter Buck. Teaching programs. Isis, 71(4):629–631, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230506.

Anonymous:1980:AAA

[202] Anonymous. Appointments, archives, associations. Isis, 71(4):631–632, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230507.

Fleming:1980:BRB

[203] Donald Fleming, Joseph Needham, Edward Grant, and Jacques Roger. Book review: The Dictionary of Scientific Biography. Isis, 71(4):633–652, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230508.

Darden:1980:BRB

[204] Lindley Darden. Book review: The Sociobiology Debate: Readings on Ethical and Scientific Issues by Arthur L. Caplan; Sociobiology: Sense or Nonsense? by Michael Ruse. Isis, 71(4):653–654, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230509.

Moyer:1980:BRBb

[205] Albert E. Moyer. Book review: Directory of Audio-Visual Sources: History of Science, Medicine, and Technology by Bruce S. Eastwood. Isis, 71(4):654, December 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230510.

VanHelden:1980:BRBb

[206] A. Van Helden. Book review: Tijdschrift voor de Geschiedenis der Geneeskunde, Natuurwetenschappen, Wiskunde en Techniek by G. T. Haneveld. Isis, 71(4):654–655, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230511.

Hall:1980:BRB

[207] Thomas S. Hall. Book review: Chemistry and Physiology in Their Historical and Philosophical Relations by Eduard Glas. Isis, 71(4):655–656, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230512.

Hull:1980:BRB

[208] David L. Hull. Book review: Teleology Revisited and Other Essays in the Philosophy and History of Science by Ernest Nagel. Isis, 71(4):656–657, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230513.

Stehr:1980:BRB

[209] Nico Stehr. Book review: Entfremdete Wissenschaft by Gernot Böhme; Michael von Engelhardt. Isis, 71(4):657–658, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230514.

Dufour:1980:BRB

[210] Paul Dufour. Book review: National Research in Canada: The NRC, 1916–1966 by Wilfrid Eggleston; Physics at the National Research Council of Canada, 1929–1952 by W. E. K. Middleton. Isis, 71(4):658–659, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230515.

Palter:1980:BRB

[211] Robert Palter. Book review: The Reception of Unconventional Science by Seymour H. Mauskopf; Reliable Knowledge: An Exploration of the Grounds for Belief in Science by John Ziman. Isis, 71(4):660–661, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230516.

Chubin:1980:BRB

[212] Daryl E. Chubin. Book review: On the Margins of Science: The Social Construction of Rejected Knowledge by Roy Wallis. Isis, 71(4):661–662, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230517.

Harding:1980:BRB

[213] Sandra Harding. Book review: Signs: Journal of Women in Culture and Society by Catherine R. Stimpson; Woman and Nature: The Roaring within Her by Susan Griffin. Isis, 71(4):662–664, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230518.

Kohlstedt:1980:BRBb

[214] Sally Gregory Kohlstedt. Book review: Graceanna Lewis: Scientist and Humanitarian by Deborah Jean Warner. Isis, 71(4):664–665, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230519.

Klosterman:1980:BRB

[215] Leo J. Klosterman. Book review: The Origin of Science and the Science of Its Origin by Stanley L. Jaki. Isis, 71(4):665, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230520.

Aitken:1980:BRB

[216] Hugh G. J. Aitken. Book review: The Emerging Goddess: The Creative Process in Art, Science, and Other Fields by Albert Rothenberg. Isis, 71 (4):665–666, December 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230521.

Eisele:1980:BRB

[217] Carolyn Eisele. Book review: Peirce’s Philosophy of Science: Critical Studies in His Theory of Induction and Scientific Method by Nicholas Rescher. Isis, 71(4):666–667, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230522.

Hiebert:1980:BRB

[218] Erwin N. Hiebert. Book review: Réflexions sur la puissance mortice du feu by Sadi Carnot; Robert Fox. Isis, 71(4):667, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230523.

Brooke:1980:BRB

[219] John Hedley Brooke. Book review: Berzelius och den kemiska atomteorin by Anders Lundgren. Isis, 71(4):668, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230524.

Jaki:1980:BRB

[220] Stanley L. Jaki. Book review: Man’s View of the Universe, a Pictorial History: Evolving Concepts of the Universe from Ancient Times of Today’s Space Probes by Gerald E. Tauber. Isis, 71(4):668, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230525.

Bud-Frierman:1980:BRB

[221] Lisa Bud-Frierman. Book review: Agricultural Records in Britain A.D. 220–1977 by J. M. Stratton; Jack Houghton Brown; Ralph Whitlock. Isis, 71(4):669, December 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230526.

Albury:1980:BRB

[222] William Randall Albury. Book review: Physiology and Classification: Historical Relations by Joseph Schiller. Isis, 71(4):669–670, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230527.

Kilborne:1980:BRB

[223] Benjamin Kilborne. Book review: Aux origines de l’anthropologie française: Les mémoires de la Société des Observateurs de l’Homme en l’an VIII by Jean Copans; Jean Jamin. Isis, 71(4):670, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230528.

Diamond:1980:BRBb

[224] Solomon Diamond. Book review: Pioneers of Psychology: Studies of the Great Figures Who Paved the Way for the Contemporary Science of Behavior by Raymond E. Fancher. Isis, 71(4):671–672, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230529.

Burnham:1980:BRB

[225] John C. Burnham. Book review: American Psychology in Historical Perspective: Addresses of the Presidents of the American Psychological Association, 1892–1977 by Ernest R. Hilgard. Isis, 71(4):672, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230530.

Jacyna:1980:BRB

[226] Stephen Jacyna. Book review: Ethics and Society in England: The Revolution in the Social Sciences 1870–1914 by Rena N. Soffer. Isis, 71 (4):672–673, December 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230531.

Bynum:1980:BRB

[227] William F. Bynum. Book review: Constantin von Economo (1876–1931): The Man and the Scientist by Ludo van Bogaert; Jean Théodoridès. Isis, 71(4):673–674, December 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230532.

Maulitz:1980:BRB

[228] Russell Maulitz. Book review: On the Normal and the Pathological by Georges Canguilhem; Carolyn R. Fawcett; Robert S. Cohen. Isis, 71(4): 674, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230533.

Oppenheimer:1980:BRB

[229] Jane M. Oppenheimer. Book review: The Genesis of Cancer: A Study in the History of Ideas by L. J. Rather. Isis, 71(4):675, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230534.

Rosner:1980:BRB

[230] David Rosner. Book review: Healing and History: Essays for George Rosen by Charles Rosenberg. Isis, 71(4):675–676, December 1980. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230535.

Layton:1980:BRBc

[231] Edwin T. Layton, Jr. Book review: Robert Maillart’s Bridges: The Art of Engineering by David P. Billington. Isis, 71(4):677, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230536.

Rose:1980:BRB

[232] Mark H. Rose. Book review: Twentieth Century Limited: Industrial Design in America, 1925–1939 by Jeffrey L. Meikle. Isis, 71(4):677–678, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230537.

Kemp:1980:BRB

[233] Emory Kemp. Book review: The Eads Bridge by Quinta Scott; Howard S. Miller. Isis, 71(4):678, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230538.

Knorr:1980:BRBb

[234] Wilbur R. Knorr. Book review: Les Éléments by Euclid; Georges J. Kayas. Isis, 71(4):678–679, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230539.

Sylla:1980:BRBb

[235] Edith D. Sylla. Book review: Juan Luis Vives against the Pseudodialecticians. A Humanist Attack on Medieval Logic by Rita Guerlac. Isis, 71 (4):679–680, December 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230540.

Unguru:1980:BRB

[236] Sabetai Unguru. Book review: Practical Geometry in the High Middle Ages: Artis cuiuslibet consummatio and the Pratike de geometrie by Stephen K. Victor. Isis, 71(4):680–681, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230541.

Courtenay:1980:BRB

[237] William J. Courtenay. Book review: Les écoles et l’enseignement dans l’Occident chrétien de la fin du Ve siècle au milieu du XIe siècle by Pierre Riché. Isis, 71(4):681–682, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230542.

Savage-Smith:1980:BRBb

[238] Emilie Savage-Smith. Book review: Proceedings of the First International Symposium for the History of Arabic Science, April 5–12, 1976. Volume II: Papers in European Languages by Ahmad Y. al-Hassan; Ghada Karmi; Nizar Namnun. Isis, 71(4):682–683, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230543.

Yoshida:1980:BRB

[239] Tadashi Yoshida. Book review: Science and Culture in Traditional Japan, A.D. 600–1854 by Masayoshi Sugimoto; David L. Swain. Isis, 71 (4):684–685, December 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230544.

Clulee:1980:BRB

[240] Nicholas H. Clulee. Book review: Robert Fludd and His Philosophicall Key: Being a Transcription of the Manuscript at Trinity College, Cambridge by Robert Fludd. Isis, 71(4):685, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230545.

Heinekamp:1980:BRB

[241] Albert Heinekamp. Book review: Der junge Leibniz. Volume I: Die wissenschaftstheoretische Problemstellung seines ersten Systementwurfs. Der Anschluss an Erhard Weigels Scientia Generalis by Konrad Moll. Isis, 71(4):686, December 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230546.

Shea:1980:BRBb

[242] William Shea. Book review: Galileo’s Early Notebooks: The Physical Questions. A Translation from the Latin, with Historical and Paleographical Commentary by William A. Wallace. Isis, 71(4):686, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230547.

Finn:1980:BRB

[243] Bernard S. Finn. Book review: From Spark to Satellite: A History of Radio Communication by Stanley Leinwoll. Isis, 71(4):687, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230549.

Swazey:1980:BRB

[244] Judith P. Swazey. Book review: Sherrington: His Life and Thought by John C. Eccles; William C. Gibson. Isis, 71(4):687, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230548.

Gifford:1980:BRBb

[245] Sanford Gifford. Book review: Freud’s Early Psychology of the Neuroses: A Historical Perspective by Kenneth Levin. Isis, 71(4):687–688, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230550.

Bloor:1980:BRB

[246] David Bloor. Book review: The Making of Mind: A Personal Account of Soviet Psychology by A. R. Luria; Michael Cole; Sheila Cole. Isis, 71 (4):688–689, December 1980. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230551.

Gregory:1980:BRBb

[247] Frederick Gregory. Book review: Wissenschaftliche Kommunikation: Die Korrespondenz Ernst Machs by Joachim Thiele. Isis, 71(4):689, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230552.

Holley:1980:BRB

[248] I. B. Holley. Book review: Chariots for Apollo: A History of Manned Lunar Spacecraft by Courtney G. Brooks; James M. Grimwood; Loyd S. Swenson, Jr. Isis, 71(4):689–690, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230553.

Anonymous:1980:BMd

[249] Anonymous. Back matter. Isis, 71(4):691–694, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230554.

Anonymous:1980:VI

[250] Anonymous. Volume information. Isis, 71(4):695–710, December 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230497.

Anonymous:1980:FMe

[251] Anonymous. Front matter. Isis, 71(5):1–4, ???? 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230491.

Neu:1980:OHF

[252] John Neu. One hundred fifth critical bibliography of the history of science and its cultural influences (to January 1980). Isis, 71(5):5–225+227– 295, ???? 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230492.

Anonymous:1980:BMe

[253] Anonymous. Back matter. Isis, 71(5):296, ???? 1980. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230493.

Anonymous:1981:FMa

[254] Anonymous. Front matter. Isis, 72(1):1–5, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231173.

Thackray:1981:EMH

[255] Arnold Thackray. Editorial: Making history. Isis, 72(1):6–10, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231174.

Lankford:1981:AVP

[256] John Lankford. Amateurs versus professionals: The controversy over telescope size in late Victorian science. Isis, 72(1):11–28, March 1981.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231175.

DeVorkin:1981:CSC

[257] David H. DeVorkin. Community and spectral classification in astrophysics: The acceptance of E. C. Pickering’s system in 1910. Isis, 72(1): 29–49, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231176.

Sheets-Pyenson:1981:WPN

[258] Susan Sheets-Pyenson. War and peace in natural history publishing: The naturalist’s library, 1833–1843. Isis, 72(1):50–72, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231177.

Swerdlow:1981:TC

[259] N. M. Swerdlow and Nicholas Copernicus. Translating Copernicus. Isis, 72(1):73–82, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231178.

See comments [462].

Hahn:1981:MLC

[260] Roger Hahn. More light on Charles the Obscure. Isis, 72(1):83–86, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231179.

Schufle:1981:ACB

[261] J. A. Schufle. Answer: Chinese brass water-spouting bowl. Isis, 72(1): 86–88, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231180.

Gillispie:1981:E

[262] Charles C. Gillispie. To the Editor. Isis, 72(1):88, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231181.

Anonymous:1981:NR

[263] Anonymous. News of record. Isis, 72(1):89, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231182.

Hall:1981:ESH

[264] A. Rupert Hall. Éloge: Sir Herbert Butterfield, 7 October 1900–20 July 1979. Isis, 72(1):90–91, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231183.

Moyer:1981:ERH

[265] Albert E. Moyer. Elogue: Richard H. Schallenberg, 3 June 1943–18 April 1980. Isis, 72(1):91–92, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231184.

Geison:1981:BRB

[266] Gerald L. Geison. Book review: The Role of Medicine: Dream, Mirage, or Nemesis? Isis, 72(1):93–95, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231185.

Rosenkrantz:1981:BRB

[267] Barbara Gutmann Rosenkrantz. Book review: Genesis and Development of a Scientific Fact. Isis, 72(1):96–99, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231186.

Rossiter:1981:BRB

[268] Margaret W. Rossiter. Book review: Fair Science: Women in the Scientific Community. Isis, 72(1):99–103, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231187.

Magner:1981:BRB

[269] Lois N. Magner. Book review: Winner Take All: Case Study of the Double Helix by Fred Jevons; A Worm in the Bud: Case Study of the Pesticide Controversy by David Wade Chambers; On the Social Analysis of Science by David Wade Chambers; Red and Expert: A Case Study of Chinese Science in the Cultural Revolution by David Wade Chambers. Isis, 72(1):104, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231188.

Finocchiaro:1981:BRBa

[270] Maurice A. Finocchiaro. Book review: Unwritten Knowledge: Case Study of the Navigators of Micronesia by Lyndsay Farrall; Puzzles and Revolutions: Case Study of the Copernican Revolution by F. R. Jevons; On the Philosophical Analysis of Science by Struan Jacobs. Isis, 72(1):104– 105, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231189.

Bernstein:1981:BRB

[271] Howard Bernstein. Book review: Arithmos-Arrythmos: Skizzen aus der Wissenschaftsgeschichte by Karin Figala; Ernst H. Berninger. Isis, 72 (1):105–106, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231190.

Curd:1981:BRB

[272] Martin V. Curd. Book review: Models: Representation and the Scientific Understanding by Marx W. Wartofsky. Isis, 72(1):106–107, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231191.

Brooks:1981:BRB

[273] Nathan M. Brooks. Book review: Osnovanie Peterburgskoi akademii nauk (The Formation of the Petersburg Academy of Sciences) by Iu. Kh. Kopelevich. Isis, 72(1):107–108, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231192.

Bailes:1981:BRB

[274] K. E. Bailes. Book review: Science and Industrialization in the USSR by Robert Lewis. Isis, 72(1):108–109, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231193.

Moran:1981:BRB

[275] Bruce T. Moran. Book review: Revolutionen in Naturwissenschaft und Gesellschaft by August Nitschke. Isis, 72(1):109, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231194.

Ainley:1981:BRB

[276] Marianne Gosztonyi Ainley. Book review: Amateurs: On the Margin between Work and Leisure by Robert A. Stebbins. Isis, 72(1):109–110, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231195.

Hounshell:1981:BRB

[277] David A. Hounshell. Book review: Dynamos and Virgins Revisited: Women and Technological Change in History by Martha Moore Trescott. Isis, 72(1):110, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231196.

Hall:1981:BRB

[278] A. Rupert Hall. Book review: Clocks and the Cosmos: Time in Western Life and Thought by Samuel L. Macey. Isis, 72(1):110–111, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231197.

Ormsby-Lennon:1981:BRB

[279] Hugh Ormsby-Lennon. Book review: English Magical and Scientific Poems to 1700: An Annotated Bibliography by Robert M. Schuler. Isis, 72(1):111–112, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231198.

Morse:1981:BRB

[280] Joann Morse. Book review: La géométrie: Première géométrie algébrique en langue française (1484) by Nicolas Chuquet; Herve l’Huillier. Isis, 72 (1):112–113, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231199.

Bers:1981:BRB

[281] Lipman Bers. Book review: Matematika XIX veka: Matematicheskaia logika, algebra, teoriia chisel, teoriia veroiatnostei (Mathematics in the 19th Century.) by A. N. Kolmogorov; A. P. Iushkevich. Isis, 72(1): 113, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231200.

Grattan-Guiness:1981:BRB

[282] I. Grattan-Guiness. Book review: Die Entstehung der Verbandstheorie by Herbert Mehrtens. Isis, 72(1):113–114, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231201.

Mahoney:1981:BRBa

[283] Michael S. Mahoney. Book review: Geschichte der Elementarmathematik. 4th Edition. Volume I: Arithmetik und Algebra by Johannes Tropfke; Kurt Vogel; Karin Reich; Helmuth Gericke. Isis, 72(1):115– 116, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231202.

Serwer:1981:BRB

[284] Daniel Serwer. Book review: Radioactivity in America: Growth and Decay of a Science by Lawrence Badash. Isis, 72(1):116–117, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231203.

Rabkin:1981:BRB

[285] Yakov M. Rabkin. Book review: Ocherk obschei istorii khimii: Razvitie klassicheskoi khimii v XIX stoletti (An Essay on the General History of Chemistry: The Development of Classical Chemistry in the 19th Century) by Nikolai Alexandrovich Figurovskii. Isis, 72(1):117, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231204.

Whitrow:1981:BRB

[286] G. J. Whitrow. Book review: Planets and Planetarians: A History of Theories of the Origins of Planetary Systems by Stanley L. Jaki. Isis, 72(1):118, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231205.

Rocke:1981:BRB

[287] Alan J. Rocke. Book review: Historische chemische und physikalische Versuche: Eingebettet in den Hintergrund von drei Jahrhunderten by Otto Krätz. Isis, 72(1):119, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231206.

Brush:1981:BRBa

[288] Stephen G. Brush. Book review: A Source Book in Astronomy and Astrophysics, 1900–1975 by Kenneth R. Lang; Owen Gingerich. Isis, 72(1):119–120, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231207.

Andrews:1981:BRBa

[289] Ilse Andrews. Book review: Carl Ritter, Genius of Geography 1779–1979 by Hanno Beck. Isis, 72(1):120–121, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231208.

Gillmor:1981:BRB

[290] C. Stewart Gillmor. Book review: Little America: Town at the End of the World by Paul A. Carter. Isis, 72(1):121, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231209.

Burkhardt:1981:BRB

[291] Richard W. Burkhardt, Jr. Book review: Le second règne de la nature: Essai sur les questions de la végétalité au XVIIIe siècle by François Delaporte. Isis, 72(1):121–123, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231210.

Sloan:1981:BRB

[292] Phillip R. Sloan. Book review: Das Ende der Naturgeschichte: Wandel kultureller Selbstverstandlichkeiten in den Wissenschaften des 18. und 19. Jahrhunderts by Wolf Lepenies. Isis, 72(1):123–124, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231211.

Sokal:1981:BRB

[293] Michael M. Sokal. Book review: Cyril Burt, Psychologist by L. S. Hearnshaw. Isis, 72(1):124–125, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231212.

Bannister:1981:BRB

[294] Robert C. Bannister. Book review: Founding Theory of American Sociology, 1881–1915 by Roscoe C. Hinkle. Isis, 72(1):125–126, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231213.

Kuklick:1981:BRBa

[295] Henrika Kuklick. Book review: Man, Mind, and Science: A History of Anthropology by Murray J. Leaf. Isis, 72(1):126–127, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231214.

Pelling:1981:BRB

[296] Margaret Pelling. Book review: The People’s Health 1830–1910 by Francis Barrymore Smith. Isis, 72(1):127–128, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231215.

Warner:1981:BRBa

[297] John Harley Warner. Book review: The Therapeutic Revolution: Essays in the Social History of American Medicine by Morris J. Vogel; Charles E. Rosenberg. Isis, 72(1):128–129, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231216.

English:1981:BRB

[298] Peter C. English. Book review: The Rise of Surgery: From Empiric Craft to Scientific Discipline by Owen H. Wangensteen; Sarah D. Wangensteen. Isis, 72(1):129–130, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231217.

Ramsey:1981:BRB

[299] Matthew Ramsey. Book review: Health, Medicine and Mortality in the Sixteenth Century by Charles Webster. Isis, 72(1):130–131, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231218.

Meikle:1981:BRB

[300] Jeffrey L. Meikle. Book review: The Port of New York: A History of the Rail and Terminal System from the Beginnings to Pennsylvania Station by Carl W. Condit. Isis, 72(1):131–132, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231219.

Smith:1981:BRBa

[301] Cecil O. Smith, Jr. Book review: Histoire générale des techniques. Volume IV: Les techniques de la civilisation industrielle: Énergie et matériaux by Maurice Daumas. Isis, 72(1):132–133, March 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231220.

Gillispie:1981:BRB

[302] Charles C. Gillispie. Book review: The Origins and Early Development of the Heavy Chemical Industry in France by John Graham Smith. Isis, 72(1):133, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231221.

Scarborough:1981:BRBa

[303] John Scarborough. Book review: De praecognitione. (On Prognosis.) by Galen; Vivian Nutton. Isis, 72(1):134, March 1981. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231222.

DeLacy:1981:BRB

[304] Phillip De Lacy. Book review: Hippocratica: Actes du Colloque hippocratique de Paris (4–9 septembre 1978) by Mirko D. Grmek. Isis, 72 (1):134–135, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231223.

Knorr:1981:BRBa

[305] Wilbur Knorr. Book review: The Beginnings of Greek Mathematics by Árpád Szabó; A. M. Ungar; Les débuts des mathématiques grecques by Árpád Szabó; M. Federspiel. Isis, 72(1):135–136, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231224.

Steneck:1981:BRB

[306] Nicholas H. Steneck. Book review: Knowledge and Cosmos: Development and Decline of the Medieval Perspective by Robert K. DeKosky. Isis, 72 (1):136–137, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231225.

Lindberg:1981:BRBa

[307] David C. Lindberg. Book review: Albertus Magnus and the Sciences: Commemorative Essays 1980 by James A. Weisheipl. Isis, 72(1):137– 138, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231226.

Saliba:1981:BRB

[308] George Saliba. Book review: Al-Jamic Bain Al-cilm wal-camal Al-Nafic Fi Sinac at Al-Hiyal (A Compendium on the Theory and Practice of the Mechanical Arts.) by Ibn Al-Razzaz Al-Jazari; Ahmad Y. Hassan. Isis, 72(1):138–139, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231227.

Heinen:1981:BRBa

[309] Anton Heinen. Book review: Buch über das Geheimnis der Schöpfung und die Darstellung der Natur (Buch der Ursachen) von Pseudo-Apollonios von Tyana by Ursula Weisser. Isis, 72(1):140, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231228.

Pingree:1981:BRB

[310] David Pingree. Book review: Beweisverfahren in der vedischen Sakralgeometrie: Ein Beitrag zur Entstehungsgeschichte von Wissenschaft by Axel Michaels. Isis, 72(1):140–141, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231229.

Shea:1981:BRBa

[311] William Shea. Book review: Index scolastico-cartésien by Étienne Gilson. Isis, 72(1):141, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231230.

McClellan:1981:BRB

[312] James E. McClellan. Book review: Ideas in Seventeenth-Century France: The Most Important Thinkers and the Climate of Ideas in Which They Worked by Edward John Kearns. Isis, 72(1):141–142, March 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231231.

Hackman:1981:BRB

[313] Willem D. Hackman. Book review: Petrus van Musschenbroek (1692– 1761), een Newtoniaans Naturonderzoeker by Cornelis de Pater. Isis, 72(1):142–143, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231232.

Smith:1981:BRBb

[314] F. B. Smith. Book review: Victorian Social Medicine: The Ideas and Methods of William Farr by John M. Eyler. Isis, 72(1):143, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231233.

Farber:1981:BRB

[315] Paul Lawrence Farber. Book review: L’ordre et le temps: L’anatomie comparée et l’histoire des vivants au XIXe siècle by Bernard Balan. Isis, 72(1):143–144, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231234.

Manier:1981:BRB

[316] Edward Manier. Book review: Charles Darwin and the Problem of Creation by Neal C. Gillespie. Isis, 72(1):144–145, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231235.

Nye:1981:BRB

[317] Mary Jo Nye. Book review: The Edge of Contingency: French Catholic Reaction to Scientific Change from Darwin to Duhem by Harry W. Paul. Isis, 72(1):145–146, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231236.

Harris:1981:BRB

[318] Ruth Harris. Book review: Paul Broca: Founder of French Anthropology, Explorer of the Brain by Francis Schiller. Isis, 72(1):146–147, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231237.

Moyer:1981:BRB

[319] Albert E. Moyer. Book review: Albert Einstein, the Human Side: New Glimpses from His Archives by Helen Dukas; Banesh Hoffmann. Isis, 72(1):147–148, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231238.

Rothenberg:1981:BRB

[320] Marc Rothenberg. Book review: Planets X and Pluto by William Graves Hoyt. Isis, 72(1):148, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231239.

Borell:1981:BRB

[321] Merriley Borell. Book review: Howard Florey: The Making of a Great Scientist by Gwyn Macfarlane. Isis, 72(1):148–149, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231240.

Forman:1981:BRB

[322] Paul Forman. Book review: German-Jewish Pioneers in Science 1900– 1933: Highlights in Atomic Physics, Chemistry and Biochemistry by David Nachmansohn. Isis, 72(1):150, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231241.

Trenn:1981:BRB

[323] Thaddeus J. Trenn. Book review: Nuclear Physics in Retrospect: Proceedings of a Symposium on the 1930’s by Roger H. Stuewer. Isis, 72 (1):150–151, March 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231242.

Kitcher:1981:BRBa

[324] Philip Kitcher. Book review: Wittgenstein on the Foundations of Mathematics by Crispin Wright. Isis, 72(1):151–152, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231243.

Anonymous:1981:BMa

[325] Anonymous. Back matter. Isis, 72(1):153–156, March 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231244.

Anonymous:1981:FMb

[326] Anonymous. Front matter. Isis, 72(2):157–161, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230967.

Secord:1981:NFC

[327] James A. Secord. Nature’s fancy: Charles Darwin and the breeding of pigeons. Isis, 72(2):162–186, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230968.

Shapin:1981:GKN

[328] Steven Shapin. Of gods and kings: Natural philosophy and politics in the Leibniz–Clarke disputes. Isis, 72(2):187–215, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230969.

Pickering:1981:HQ

[329] Andrew Pickering. The hunting of the quark. Isis, 72(2):216–236, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230970.

Goldstein:1981:HAT

[330] Bernard R. Goldstein. The Hebrew astronomical tradition: New sources. Isis, 72(2):237–251, June 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230971.

Erickson:1981:ACC

[331] Paul Erickson. The anthropology of Charles Caldwell, M.D. Isis, 72(2): 252–256, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230972.

Sloan:1981:AMH

[332] Phillip R. Sloan and Sally Gregory Kohlstedt. Annual meeting of the History of Science Society. 17–19 October 1980. Isis, 72(2):257–262, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230973.

Gullard:1981:CBI

[333] Pamela Gullard. The Charles Babbage Institute for the History of Information Processing (CBI). Isis, 72(2):262–264, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230974.

Tihon:1981:EJM

[334] Anne Tihon. Éloge: Joseph Mogenet, 26 February 1913–18 February 1980. Isis, 72(2):265–266, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230975.

Rudwick:1981:CPH

[335] Martin Rudwick, William Coleman, Edith Sylla, and Lorraine Daston. Critical problems in the history of science. Isis, 72(2):267–283, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230976.

Brush:1981:BRBb

[336] Stephen G. Brush. Book review: A World on Paper: Studies on the Second Scientific Revolution by Mirella Giacconi; Enrico Bellone; Riccardo Giacconi; The Tragicomical History of Thermodynamics, 1822– 1854 by Clifford A. Truesdell. Isis, 72(2):284–286, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230977.

Brieger:1981:BRB

[337] Gert H. Brieger. Book review: Encyclopedia of Bioethics by Warren T. Reich. Isis, 72(2):286–287, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230978.

Osler:1981:BRB

[338] Margaret J. Osler. Book review: The Death of Nature: Women, Ecology, and the Scientific Revolution by Carolyn Merchant. Isis, 72(2):287–288, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/230979.

Moore:1981:BRB

[339] James R. Moore. Book review: Philip Henry Gosse: A Bibliography by R. B. Freeman; Douglas Wertheimer. Isis, 72(2):288–289, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230980.

Small:1981:BRB

[340] Henry Small. Book review: A Bibliography of Quantitative Studies on Science and Its History by Roger Hahn. Isis, 72(2):289, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230981.

Olesko:1981:BRB

[341] Kathryn M. Olesko. Book review: Max Planck. A Bibliography of His Non-Technical Writings by Henry Lowood; William Henry Bragg and William Lawrence Bragg. A Bibliography of Their Non-Technical Writings by Henry Lowood; Ernest Rutherford. A Bibliography of His Non-Technical Writings by Henry Lowood. Isis, 72(2):289–290, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230982.

Hesse:1981:BRB

[342] Mary Hesse. Book review: Theory and Evidence by Clark Glymour. Isis, 72(2):290–291, June 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230983.

Wykstra:1981:BRB

[343] Stephen Wykstra. Book review: Progress and Rationality in Science by Gerard Radnitzky; Gunnar Andersson. Isis, 72(2):291–292, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230984.

Anderson:1981:BRB

[344] Frances Anderson and Raymond Duchesne. Book review: Biological Sciences at the National Research Council of Canada: The Early Years to 1952 by Norman T. Gridgeman. Isis, 72(2):292–293, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230985.

Hiebert:1981:BRB

[345] Erwin Hiebert. Book review: Die Naturwissenschaften an der Philipps– Universität Marburg, 1527–1977 by Rudolf Schmitz. Isis, 72(2):293– 294, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230986.

Maienschein:1981:BRB

[346] Jane Maienschein. Book review: A History of Microtechnique by Brian Bracegirdle. Isis, 72(2):294, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230987.

Bedin:1981:BRB

[347] Silvio A. Bedin. Book review: Mathematical Instrument-Makers in the Grocers’ Company 1688–1800 by Joyce Brown. Isis, 72(2):294–295, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230988.

Warner:1981:BRBb

[348] Deborah Jean Warner. Book review: Greenwich Time and the Discovery of the Longitude by Derek Howse. Isis, 72(2):295, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230989.

Turner:1981:BRB

[349] G. L’E. Turner. Book review: Catalog of the Scientific Apparatus at the College of Charleston, 1800–1940 by Barbara Hughes; Ralph Melnick. Isis, 72(2):295–296, June 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230990.

Swenson:1981:BRBa

[350] Loyd S. Swenson, Jr. Book review: Scientific Images and Their Social Uses: An Introduction to the Concept of Scientism by Iain Cameron; David Edge; Limits of a Modern World: A Study of the “Limits to Growth” Debate by Robert McCutcheon; The Atomic Bomb by Margaret Gowing; Lorna Arnold; Are Science and Technology Neural? by Joan Lipscombe; Bill Williams; Assessment of Technological Decisions– Case Studies by Ernest Braun; David Collingridge; Kate Hinton. Isis, 72(2):296–297, June 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230991.

Woolgar:1981:BRB

[351] Steve Woolgar. Book review: The Reward System in British and American Science by Jerry Gaston. Isis, 72(2):297–298, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230992.

Knorr:1981:BRBb

[352] Karin Knorr. Book review: The Dynamics of Science and Technology: Social Values, Technical Norms and Scientific Criteria in the Development of Knowledge by Wolfgang Krohn; Edwin T. Layton,; Peter Weingart. Isis, 72(2):298–299, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230993.

Todd:1981:BRB

[353] Edmund Todd. Book review: Naturwissenschaft, Technik und NS-Ideologie: Beiträge zur Wissenschaftsgeschichte des Dritten Reiches by Herbert Mehrtens; Steffen Richter. Isis, 72(2):299–300, June 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230994.

Glass:1981:BRBa

[354] Bentley Glass. Book review: Unpopular Essays on Technological Progress by Nicholas Rescher. Isis, 72(2):300, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230995.

Finocchiaro:1981:BRBb

[355] Maurice A. Finocchiaro. Book review: Testi e contesti: Quadreni di scienze, storia e società by Angelo Baraca. Isis, 72(2):300–301, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230996.

Resnik:1981:BRB

[356] Michael D. Resnik. Book review: Philosophical and Mathematical Correspondence by Gottlob Frege; Gottfried Gabriel; Hans Hermes; Friedrich Kambartel; Christian Thiel; Albert Veraart; Brian McGuinness; Hans Kaal. Isis, 72(2):301–302, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230997.

Kitcher:1981:BRBb

[357] Philip Kitcher. Book review: Frege and the Philosophy of Mathematics by Michael D. Resnik. Isis, 72(2):302–303, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230998.

Melhado:1981:BRB

[358] Evan M. Melhado. Book review: Kleine Geschichte der Chemie by Gunther Simon. Isis, 72(2):303, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230999.

Taylor:1981:BRB

[359] Kenneth L. Taylor. Book review: Four Hitherto Unpublished Geological Lectures Given by Sir Humphry Davy in 1805 by Humphry Davy; Humphry Davy on Geology: The 1805 Lectures for the General Audience by Humphry Davy; Robert Siegfried; Robert H. Dott. Isis, 72(2): 303–304, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231000.

Ospovat:1981:BRBa

[360] Alexander M. Ospovat. Book review: Two Hundred Years of Geology in America by Cecil J. Schneer. Isis, 72(2):304–305, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231001.

Ewan:1981:BRB

[361] Joseph Ewan. Book review: The Fossil Hunters: In Search of Ancient Plants by Henry N. Andrews. Isis, 72(2):305–306, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231002.

Fruton:1981:BRB

[362] Joseph S. Fruton. Book review: A History of Biochemistry. Part V. The Unravelling of Biosynthetic Pathways by Marcel Florkin. Isis, 72 (2):306–307, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231003.

Roe:1981:BRB

[363] Shirley A. Roe. Book review: History and Philosophy of the Life Sciences by M. D. Grmek. Isis, 72(2):307–308, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231004.

Kuklick:1981:BRBb

[364] Henrika Kuklick. Book review: The French Encounter with Africans: White Response to Blacks, 1530–1880 by William B. Cohen. Isis, 72(2): 308–309, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231005.

Numbers:1981:BRB

[365] Ronald L. Numbers. Book review: Sex, Diet, and Debility in Jacksonian America: Sylvester Graham and Health Reform by Stephen Nissenbaum. Isis, 72(2):309, June 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231006.

Corner:1981:BRB

[366] George W. Corner and Senior. Book review: The Education of American Physicians: Historical Essays by Ronald L. Numbers. Isis, 72(2):309– 310, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231007.

Tarr:1981:BRB

[367] Joel A. Tarr. Book review: A Retrospective Technology Assessment: Submarine Telegraphy–The Transatlantic Cable of 1866 by Vary T. Coates; Bernard Finn. Isis, 72(2):310–311, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231008.

Layton:1981:BRBa

[368] Edwin T. Layton, Jr. Book review: A History of Industrial Power in the United States, 1780–1930. Volume I: Waterpower in the Century of the Steam Engine by Louis C. Hunter. Isis, 72(2):311–312, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231009.

Hannah:1981:BRB

[369] Leslie Hannah. Book review: Beardmore: The History of a Scottish Industrial Giant by John R. Hume; Michael S. Moss. Isis, 72(2):312, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231010.

Constant:1981:BRB

[370] Edward W. Constant, II. Book review: Tekniska Museet Symposia: Technology and Its Impact on Society. Isis, 72(2):312–313, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231011.

Kranakis:1981:BRB

[371] Eda Fowlks Kranakis. Book review: Les mécaniciens grecs: La naissance de la technologie by Bertrand Gille. Isis, 72(2):313–314, June 1981.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231012.

Mahoney:1981:BRBb

[372] Michael S. Mahoney. Book review: Una fisica senza matematica: Democrito, Aristotele, Filopono by Giulio A. Lucchetta. Isis, 72(2):314, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231013.

Lindberg:1981:BRBb

[373] David C. Lindberg. Book review: Catalogus translationum et commentariorum: Medieval and Renaissance Latin Translations and Commentaries: Annotated Lists and Guides. Volume IV by F. Edward Cranz; Paul Oskar Kristeller. Isis, 72(2):314–315, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231014.

Riddle:1981:BRB

[374] John M. Riddle. Book review: The Prose Salernitan Questions: An Anonymous Collection Dealing with Science and Medicine Written by an Englishman c. 1200, with an Appendix of Ten Related Collections by Brian Lawn. Isis, 72(2):315–316, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231015.

Kaplan:1981:BRB

[375] Barbara Beigun Kaplan. Book review: Bartholomaeus Anglicus on the Properties of Soul and Body: “De proprietatibus rerum libri III et IV” by Bartholomaeus Anglicus; R. James Long. Isis, 72(2):316–317, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231016.

Heinen:1981:BRBb

[376] Anton Heinen. Book review: Le livre de la méthode du médecin de cAli b. Rid.wan (998–1067). Volume I. Introduction — Thérapeutique by Jacques Grand’henry. Isis, 72(2):317, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231017.

Chang:1981:BRB

[377] K. C. Chang. Book review: K’aokuhsue yu Kechishih (Essays on Archaeology of Science and Technology in China) by Hsia Nai. Isis, 72(2):

317–318, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231018.

Eide:1981:BRB

[378] Elling O. Eide. Book review: Nan-fang ts’ao-mu chuang: A Fourth Century Flora of Southeast Asia by Chi Han; Hui-Lin Li. Isis, 72(2): 318, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231019.

Dunbar:1981:BRBa

[379] G. S. Dunbar. Book review: La géographie de la Renaissance (1420– 1620) by Numa Broc. Isis, 72(2):319, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231020.

Copenhaver:1981:BRB

[380] Brian Copenhaver. Book review: Renaissance Man and Creative Thinking: A History of Concepts of Harmony, 1400–1700 by Dorothy Koenigsberger. Isis, 72(2):319–320, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231021.

Baker:1981:BRB

[381] Keith Michael Baker. Book review: La logique/Logic by Étienne de Condillac; W. R. Albury. Isis, 72(2):320–321, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231022.

Hovenkamp:1981:BRB

[382] Herbert Hovenkamp. Book review: The Works of Jonathan Edwards, Volume VI: Scientific and Philosophical Writings: The “Spider” Papers, “Natural Philosophy,” “The Mind,” Short Scientific and Philosophical Papers by Jonathan Edwards; Wallace E. Anderson. Isis, 72(2):321, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231023.

Wisan:1981:BRB

[383] Winifred Lovell Wisan. Book review: Momento: Studi galileiani by Paolo Gailuzzi. Isis, 72(2):321–322, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231024.

McEvoy:1981:BRB

[384] John McEvoy. Book review: Joseph Priestley: Scientist, Theologian, and Metaphysician by Erwin N. Hiebert; Aaron J. Ihde; Robert E. Schofield; Lester Keift; Bennett R. Willeford. Isis, 72(2):322–323, June 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231025.

Gregory:1981:BRB

[385] Frederick Gregory. Book review: Briefwechsel zwischen Alexander von Humboldt und Heinrich Christian Schumacher by Kurt R. Biermann. Isis, 72(2):323–324, June 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231026.

Browne:1981:BRB

[386] Janet Browne. Book review: A Delicate Arrangement: The Strange Case of Charles Darwin and Alfred Russel Wallace by Arnold C. Brackman. Isis, 72(2):324, June 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231027.

Appel:1981:BRB

[387] Toby A. Appel. Book review: The Letters of Georges Cuvier: A Summary Calendar of Manuscript and Printed Materials Preserved in Europe, the United States of America, and Australasia by Dorinda Outram. Isis, 72(2):324–325, June 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231028.

Shea:1981:BRBb

[388] William Shea. Book review: Les orbites des planètes by G. W. F. Hegel; François De Gandt. Isis, 72(2):325, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231029.

Gruber:1981:BRB

[389] Howard E. Gruber. Book review: The Darwinian Revolution by Michael Ruse. Isis, 72(2):325–327, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231030.

Kohlstedt:1981:BRB

[390] Sally Gregory Kohlstedt. Book review: Mr. Peale’s Museum: Charles Willson Peale and the First Popular Museum of Natural Science and Art by Charles Coleman Sellers. Isis, 72(2):327–328, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231031.

Dolman:1981:BRB

[391] Claude Dolman. Book review: Noguchi and His Patrons by Isabel R. Plesset. Isis, 72(2):328–329, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231032.

Wheaton:1981:BRB

[392] Bruce R. Wheaton. Book review: The Early Years: The Niels Bohr Institute, 1921–1930 by Peter Robertson. Isis, 72(2):329–330, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231033.

Kevles:1981:BRA

[393] Daniel J. Kevles. Book review: Alice Kimball Smith and Charles Weiner: Robert Oppenheimer: Letters and Recollections. Isis, 72(2):330, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231034.

Goldberg:1981:BRB

[394] Stanley Goldberg. Book review: Genesis of Relativity: Einstein in Context by Loyd S. Swenson. Isis, 72(2):330–331, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231035.

Dick:1981:BRB

[395] Steven J. Dick. Book review: The Search for Life on Mars: Evolution of an Idea by Henry S. F. Cooper. Isis, 72(2):331, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231036.

Anonymous:1981:BMb

[396] Anonymous. Back matter. Isis, 72(2):332–336, June 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231037.

Anonymous:1981:FMc

[397] Anonymous. Front matter. Isis, 72(3):337–341, September 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230253.

Jobe:1981:DRS

[398] Thomas Harmon Jobe. The Devil in restoration science: The Glanvill– Webster witchcraft debate. Isis, 72(3):342–356, September 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230254.

Marshall:1981:NON

[399] Peter Marshall. Nicole Oresme on the nature, reflection, and speed of light. Isis, 72(3):357–374, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230255.

Sutton:1981:EMM

[400] Geoffrey Sutton. Electric medicine and mesmerism. Isis, 72(3):375–392, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230256.

Ravetz:1981:MHS

[401] Jerome Ravetz and Richard S. Westfall. Marxism and the history of science. Isis, 72(3):393–405, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230257.

Rousseau:1981:LMS

[402] G. S. Rousseau. Literature and medicine: The state of the field. Isis, 72 (3):406–424, September 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230258.

Geison:1981:PVV

[403] Gerald L. Geison and Louis Pasteur. Pasteur on vital versus chemical ferments: A previously unpublished paper on the inversion of sugar. Isis, 72(3):425–445, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230259.

Pritykin:1981:RCS

[404] L. M. Pritykin. The role of concepts of structure in the development of the physical chemistry of polymers. Isis, 72(3):446–456, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230260.

Zezong:1981:CSH

[405] Xi Zezong. Chinese studies in the history of astronomy, 1949–1979. Isis, 72(3):456–470, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230261.

Weart:1981:ISH

[406] Spencer R. Weart and David H. DeVorkin. Interviews as sources for history of modern astrophysics. Isis, 72(3):471–477, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230262.

Olesko:1981:ETH

[407] Kathryn M. Olesko. Employment trends in history of science. Isis, 72 (3):477–479, September 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230263.

Cohen:1981:BRB

[408] I. Bernard Cohen. Book review: Aepinus’s Essay on the Theory of Electricity and Magnetism. Isis, 72(3):480–489, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230264.

Finocchiaro:1981:BRBc

[409] Maurice A. Finocchiaro. Book review: Paradigms and Revolutions: Appraisals and Applications of Thomas Kuhn’s Philosophy of Science by Gary Gutting. Isis, 72(3):490–491, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230265.

Zaslow:1981:BRB

[410] Morris Zaslow. Book review: Science, Technology, and Canadian History: The First Conference on the Study of the History of Canadian Science and Technology, Kingston, Ontario, 1978 by Richard A. Jarrell; Norman R. Ball. Isis, 72(3):491, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230266.

Morman:1981:BRB

[411] Edward T. Morman. Book review: Citation Indexing: Its Theory and Application in Science, Technology, and Humanities by Eugene Garfield. Isis, 72(3):491–492, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230267.

Sarjeant:1981:BRB

[412] William A. S. Sarjeant. Book review: Isis Cumulative Bibliography 1966– 1975. A Bibliography of the History of Science, Formed from Isis Critical Bibliographies 91–100 Indexing Literature Published from 1965 to 1974. Volume I: Personalities and Institutions by John Neu. Isis, 72(3):492– 493, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230268.

Ruse:1981:BRB

[413] Michael Ruse. Book review: The Possibility of Naturalism: A Philosophical Critique of the Contemporary Human Sciences by Roy Bhaskar. Isis, 72(3):493–495, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230269.

Guedon:1981:BRB

[414] Jean-Claude Guedon. Book review: Science et dialectique chez Hegel et Marx by Michel Vadée. Isis, 72(3):495, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230270.

Sheets-Pyenson:1981:BRBa

[415] Susan Sheets-Pyenson. Book review: Science and Colonial Expansion: The Role of the British Royal Botanic Gardens by Lucile H. Brockway. Isis, 72(3):495–496, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230271.

Layton:1981:BRBb

[416] Edwin T. Layton, Jr. Book review: A Centennial History of the American Society of Mechanical Engineers, 1880–1980 by Bruce Sinclair; James P. Hull. Isis, 72(3):497, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230272.

Orlans:1981:BRB

[417] Harold Orlans. Book review: The Making of a Science Policy: A Historical Study of the Institutional and Conceptual Background to Dutch Science Policy in a West-European Perspective by Frits Henry Brookman. Isis, 72(3):497–498, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230273.

Pursell:1981:BRB

[418] Carroll W. Pursell, Jr. Book review: Science Advice to the President by William T. Golden. Isis, 72(3):498–499, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230274.

Swenson:1981:BRBb

[419] Loyd S. Swenson, Jr. Book review: Toward the Endless Frontier: History of the Committee on Science and Technology, 1959–79 by Ken Hechler. Isis, 72(3):499–500, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230275.

Hodes:1981:BRB

[420] Elizabeth Hodes. Book review: John von Neumann and Norbert Wiener: From Mathematics to the Technologies of Life and Death by Steve J. Heims. Isis, 72(3):500–501, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230276.

Vucinich:1981:BRB

[421] Alexander Vucinich. Book review: Ideology of/in the Natural Sciences by Hilary Rose; Steven Rose. Isis, 72(3):501–502, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230277.

Siegel:1981:BRB

[422] Daniel M. Siegel. Book review: From X-Rays to Quarks: Modern Physicists and Their Discoveries by Emilio Segrè. Isis, 72(3):502, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230278.

Garber:1981:BRB

[423] Elizabeth Garber. Book review: The Self-Splitting Atom: The History of the Rutherford-Soddy Collaboration by Thaddeus J. Trenn. Isis, 72(3): 503, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230279.

Eastwood:1981:BRB

[424] Bruce Stansfield Eastwood. Book review: Geschichte der Impetustheorie: Untersuchungen zum Ursprung der klassischen Mechanik by Michael Wolff. Isis, 72(3):503–504, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230280.

Faul:1981:BRB

[425] Henry Faul. Book review: The Abyss of Time: Changing Conceptions of the Earth’s Antiquity after the Sixteenth Century by Claude C. Albritton. Isis, 72(3):504–505, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230281.

Ospovat:1981:BRBb

[426] Alexander M. Ospovat. Book review: Biographien bedeutender Geowissenschaften der Sowjetunion: 19 biographische Darstellungen zu bedeutenden Gelehrten der russischen und sowjetischen Geologiegeschichte by Martin Guntau. Isis, 72(3):505, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230282.

Guntau:1981:BRB

[427] Martin Guntau. Book review: Alfred Wegener und die Drift der Kontinente by Martin Schwarzbach. Isis, 72(3):505–506, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230283.

Elliott:1981:BRB

[428] Clark A. Elliott. Book review: Archival Sources for the History of Biochemistry and Molecular Biology: A Reference Guide and Report by David Bearman; John T. Edsall; Margaret Miller; Matthew Konopka. Isis, 72(3):506–507, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230284.

Hotchkiss:1981:BRB

[429] Rollin D. Hotchkiss. Book review: Origins of Molecular Biology: A Tribute to Jacques Monod by Andre Lwoff; Agnes Ullmann. Isis, 72(3): 507–508, September 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230285.

Shapin:1981:BRB

[430] Steven Shapin. Book review: The Machinery Question and the Making of Political Economy, 1815–1848 by Maxine Berg. Isis, 72(3):508–509, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230286.

Bakan:1981:BRB

[431] David Bakan. Book review: Historiography of Modern Psychology: Aims, Resources and Approaches by Josef Brožek; Ludwig J. Pongratz. Isis, 72(3):509, September 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230287.

Hyde:1981:BRB

[432] Charles K. Hyde. Book review: Alexander Holley and the Makers of Steel by Jeanne McHugh. Isis, 72(3):509–511, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230288.

Stapleton:1981:BRB

[433] Darwin H. Stapleton. Book review: Transport Technology and Social Change by Per Sörbom. Isis, 72(3):511, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230289.

Eamon:1981:BRB

[434] William Eamon. Book review: Alchemy: The Philosopher’s Stone by Allison Coudert. Isis, 72(3):511–512, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230290.

Dobbs:1981:BRBa

[435] B. J. T. Dobbs. Book review: Alchemy: A Bibliography of English-Language Writings by Alan Pritchard. Isis, 72(3):512, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230291.

Sylla:1981:BRB

[436] Edith Sylla. Book review: Studies in Medieval Physics and Mathematics by Marshall Clagett. Isis, 72(3):512–513, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230292.

Reeds:1981:BRB

[437] Karen Reeds. Book review: Herbarium Apulei, 1481; Herbolario volgare, 1522. Isis, 72(3):513, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230293.

Pedersen:1981:BRBa

[438] Olaf Pedersen. Book review: Les instruments de la théorie des planètes selon Ptolémée: Equatoires et horlogerie planétaire du XIIIe au XVIe siècle by Emmanuel Poulle. Isis, 72(3):513–514, September 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230294.

Goldstein:1981:BRB

[439] Bernard R. Goldstein. Book review: Islamic Geomancy and a Thirteenth-Century Divinatory Device by Emilie Savage-Smith; Marion B. Smith. Isis, 72(3):514–515, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

230295.

Schneider:1981:BRB

[440] Laurence A. Schneider. Book review: American Science and Modern China, 1876–1936 by Peter Buck. Isis, 72(3):515–516, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230296.

Porkert:1981:BRB

[441] Manfred Porkert. Book review: Medizin in China: Eine Ideengeschichte by Paul U. Unschuld. Isis, 72(3):516–517, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230297.

Sturchio:1981:BRB

[442] Jeffrey L. Sturchio. Book review: La introducción de la nueva nomenclatura química y el rechazo de la teoría de la acidez de Lavoisier en España: Edición facsímil de las Reflexiones sobre la nueva nomenclatura química (Madrid, 1788) de Juan Manuel de Aréjula by Ramon Gago; Juan L. Carrillo. Isis, 72(3):517, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230298.

Vassberg:1981:BRB

[443] David E. Vassberg. Book review: Ciencia y técnica en la sociedad española de los siglos XVI y XVII by José María López Piñero. Isis, 72 (3):517–518, September 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230299.

Burke:1981:BRB

[444] Michael Burke. Book review: Clásicos españoles de la anatomía patológica anteriores a Cajal by José María López Piñero; Francesc Biyosa; Mariá-Luz Terrada. Isis, 72(3):518, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230300.

Albury:1981:BRB

[445] W. R. Albury. Book review: A Critical Study of Condillac’s Traité des systèmes by Ellen McNiven Hine. Isis, 72(3):519, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230301.

Stroup:1981:BRB

[446] Alice Stroup. Book review: L’institution de la science et l’expérience du vivant: Méthode et expérience à l’Académie Royale des Sciences, 1666–1793 by C. Salomon-Bayet. Isis, 72(3):519–520, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/230302.

Porter:1981:BRBa

[447] Roy Porter. Book review: The Temperature of History: Phases of Science and Culture in the Nineteenth Century by Stephen Brush. Isis, 72(3): 520–521, September 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230303.

vonEngelhardt:1981:BRB

[448] Dietrich von Engelhardt. Book review: Romantische Naturphilosophie und Arzneimittelehre 1800–1840 by Dieter Oldenburg. Isis, 72(3):521– 522, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/230304.

Rose:1981:BRB

[449] Hilary Rose. Book review: Sage: A Life of J. D. Bernal by Maurice Goldsmith. Isis, 72(3):522–523, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/230305.

Wagner:1981:BRB

[450] Roy Wagner. Book review: Gregory Bateson: The Legacy of a Scientist by David Lipset. Isis, 72(3):523–524, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230306.

Pyenson:1981:BRB

[451] Lewis Pyenson. Book review: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg, u.a. Vol. I: 1919–1929 by Wolfgang Pauli; A. Hermann; K. V. Meyenn; V. F. Weisskopf. Isis, 72 (3):524–525, September 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230307.

Mehra:1981:BRB

[452] Jagdish Mehra. Book review: Wolfgang Pauli: Die Jahre 1918–1930. Skizzen zu einer wissenschaftlichen Biographie by Steffen Richter. Isis, 72 (3):525–526, September 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230308.

Anonymous:1981:BMc

[453] Anonymous. Back matter. Isis, 72(3):527–530, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230309.

Anonymous:1981:FMd

[454] Anonymous. Front matter. Isis, 72(4):531–535, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231246.

Brieger:1981:GEH

[455] Gert H. Brieger. Guest editorial: The history of medicine and the history of science. Isis, 72(4):536–540, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231247.

McClellan:1981:ARS

[456] James E. McClellan, III. The Académie royale des sciences, 1699–1793: A statistical portrait. Isis, 72(4):541–567, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231248.

Smith:1981:GBP

[457] A. Mark Smith. Getting the big picture in perspectivist optics. Isis, 72 (4):568–589, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231249.

Numbers:1981:WBR

[458] Ronald L. Numbers and William J. Orr, Jr. William Beaumont’s reception at home and abroad. Isis, 72(4):590–612, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231250.

Sredniawa:1981:AYP

[459] Bronislaw ’Sredniawa. The anthropologist as a Young physicist: Bronisław Malinowski’s apprenticeship. Isis, 72(4):613–620, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231251.

Mohr:1981:NTP

[460] Richard D. Mohr. The number theory in Plato’s Republic vii and philebus. Isis, 72(4):620–627, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231252.

Hall:1981:PSB

[461] Marie Boas Hall. Public science in Britain: The role of the Royal Society. Isis, 72(4):627–629, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231253.

Rosen:1981:TC

[462] Edward Rosen, Erna Hilfstein, and N. M. Swerdlow. “Translating Copernicus”. Isis, 72(4):629–631, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231254. See [259].

Lindberg:1981:NR

[463] David C. Lindberg and Ronald L. Numbers. News of record. Isis, 72 (4):632–634, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231255.

Holmes:1981:EDO

[464] Frederic L. Holmes. Éloge: Dov Ospovat, 14 June 1947–28 September 1980. Isis, 72(4):635–636, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231256.

Blackmore:1981:EJT

[465] John T. Blackmore. Éloge: Joachim Thiele, 13 September 1931–19 July 1980. Isis, 72(4):636–637, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231257.

Decker:1981:BRB

[466] Hannah S. Decker. Book review: Freud, Biologist of the Mind: Beyond the Psychoanalytic Legend. Isis, 72(4):638–642, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231258.

Glass:1981:BRBb

[467] Bentley Glass. Book review: The Evolutionary Synthesis: Perspectives on the Unification of Biology. Isis, 72(4):642–647, December 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231259.

Bedini:1981:BRB

[468] Silvio A. Bedini. Book review: Album of Science: From Leonardo to Lavoisier, 1450–1800 by I. Bernard Cohen. Isis, 72(4):648–649, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231260.

Dupree:1981:BRB

[469] A. Hunter Dupree. Book review: Science, Technology, and National Policy by Thomas J. Kuehn; Alan L. Porter. Isis, 72(4):649–650, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231261.

Dobbs:1981:BRBb

[470] B. J. T. Dobbs. Book review: The Occult Philosophy in the Elizabethan Age by Frances A. Yates. Isis, 72(4):650, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231262.

Egerton:1981:BRB

[471] Frank N. Egerton, III. Book review: British Natural History Books, 1495–1900: A Handlist by Richard B. Freeman. Isis, 72(4):651, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231263.

Schneer:1981:BRB

[472] Cecil J. Schneer. Book review: Geologists and the History of Geology: An International Bibliography from the Origins to 1978 by William Sarjeant. Isis, 72(4):651–652, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231264.

Multhauf:1981:BRB

[473] Robert P. Multhauf. Book review: Die Schemnitzer Gedenkbibliothek by Laszlo Zsamboki. Isis, 72(4):652, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231265.

Beatty:1981:BRB

[474] John Beatty. Book review: Applications of Inductive Logic: Proceedings of a Conference at the Queen’s College, Oxford, 21–24 August 1978 by L. Jonathan Cohen; Mary Hesse. Isis, 72(4):653–654, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231266.

Lenoir:1981:BRB

[475] Timothy Lenoir. Book review: Philosophie des Lebendigen: Der Begriff des Organischen bei Kant, sein Grund und seine Aktualität by Reinhard Löw. Isis, 72(4):654–655, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231267.

Giere:1981:BRB

[476] Ronald N. Giere. Book review: Scientific Discovery, Logic, and Rationality by Thomas Nickles. Isis, 72(4):655–656, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231268.

Cassidy:1981:BRB

[477] David C. Cassidy. Book review: State, Society, and University in Germany, 1700–1914 by Charles E. McClelland. Isis, 72(4):656–657, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231269.

Fox:1981:BRB

[478] Robert Fox. Book review: Science and Immortality: The Éloges of the Paris Academy of Sciences (1699–1791) by Charles B. Paul. Isis, 72 (4):657–658, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231270.

Mullen:1981:BRB

[479] Pierce C. Mullen. Book review: Lorenz Oken (1779–1851), Erster Rektor der Universität Zürich: Festvortrag zur Feier seines 200. Geburtstages by Emil Kuhn-Schnyder. Isis, 72(4):658, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231271.

Sheets-Pyenson:1981:BRBb

[480] Susan Sheets-Pyenson. Book review: Savoir scientifique et pouvoir social: L’École polytechnique 1794–1914 by Terry Shinn. Isis, 72(4):658–659, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231272.

Bud:1981:BRB

[481] Robert Bud. Book review: The Cancer Mission: Social Contexts of Biomedical Research by Kenneth E. Studer; Daryl E. Chubin. Isis, 72 (4):659–660, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231273.

Freudenthal:1981:BRBa

[482] Gad Freudenthal. Book review: Le petit savant illustré by Pierre Thuillier. Isis, 72(4):660, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231274.

Freudenthal:1981:BRBb

[483] Hans Freudenthal. Book review: Abrégé d’histoire des mathématiques, 1700–1900 by Jean Dieudonné. Isis, 72(4):660–661, December 1981. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231275.

Kline:1981:BRB

[484] Morris Kline. Book review: From the Calculus to Set Theory, 1630–1910: An Introductory History by I. Grattan-Guinness. Isis, 72(4):661–662, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231276.

Hoddeson:1981:BRB

[485] Lillian Hoddeson. Book review: Advances in Electronics and Electron Physics. Volume L by L. Marton; C. Marton. Isis, 72(4):662–663, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231277.

Pedersen:1981:BRBb

[486] Olaf Pedersen. Book review: The Sky Explored: Celestial Cartography, 1500–1800 by Deborah J. Warner. Isis, 72(4):663, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231278.

Greene:1981:BRB

[487] Mott Greene. Book review: History of Concepts in Precambrian Geology by W. O. Kupsch; W. A. S. Sarjeant. Isis, 72(4):663–664, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231279.

Porter:1981:BRBb

[488] Roy Porter. Book review: La Macchina della terra: Teorie geologiche dal seicento all’ottocento by Nicoletta Morello. Isis, 72(4):664–665, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231280.

Dott:1981:BRB

[489] R. H. Dott, Jr. Book review: Grove Karl Gilbert: A Great Engine of Research by Stephen J. Pyne. Isis, 72(4):665–666, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231281.

Woodward:1981:BRB

[490] William Woodward. Book review: An Education in Psychology: James McKeen Cattell’s Journal and Letters from Germany and England, 1880– 1888 by James McKeen Cattell; Michael M. Sokal. Isis, 72(4):666–667, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231282.

Jones:1981:BRB

[491] Robert Jones. Book review: Liberalism and Sociology: L. T. Hobhouse and Political Argument in England, 1880–1914 by Stefan Collini. Isis, 72 (4):667–668, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231283.

Blight:1981:BRB

[492] James G. Blight. Book review: Language and the Origins of Psychoanalysis by John Forrester. Isis, 72(4):668, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231284.

Rattansi:1981:BRB

[493] P. M. Rattansi. Book review: English Almanacs, 1500–1800: Astrology and the Popular Press by Bernard Capp. Isis, 72(4):668–669, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231285.

Mauskopf:1981:BRB

[494] Seymour H. Mauskopf. Book review: Science, Pseudo-Science, and Society by Marsha P. Hanen; Margaret J. Osler; Robert G. Weyant. Isis, 72 (4):669–670, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231286.

Collins:1981:BRB

[495] H. M. Collins. Book review: The Elusive Science: Origins of Experimental Psychical Research by Seymour H. Mauskopf; Michael R. McVaugh. Isis, 72(4):670–671, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231287.

Toomer:1981:BRB

[496] G. J. Toomer. Book review: L’Astronomie dans l’antiquité classique: Actes du colloque tenu à l’Université de Toulouse-Le Mirail, 21–23 Octobre 1977. Isis, 72(4):671–672, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231288.

Scarborough:1981:BRBb

[497] John Scarborough. Book review: Der griechische Arzt im Zeitalter des Hellenismus: Seine Stellung in Staat und Gesellschaft by Fridolf Kudlien. Isis, 72(4):672–673, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231289.

Wasserstein:1981:BRB

[498] A. Wasserstein. Book review: Magic, Reason, and Experience: Studies in the Origin and Development of Greek Science by G. E. R. Lloyd. Isis, 72 (4):673–674, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231290.

Berggren:1981:BRB

[499] J. L. Berggren. Book review: Archimedes: Ingenieur, Naturwissenschaftler und Mathematiker by Ivo Schneider. Isis, 72(4):674–675, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231291.

Furley:1981:BRB

[500] David J. Furley. Book review: La naissance de la physique dans le texte de Lucrèce: Fleuves et turbulences by Michel Serres; Critique: Revue général des publications françaises et étrangères. Vol. 35, No. 380, January 1979: Michel Serres: Interférences et turbulences. Isis, 72(4):675– 676, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231292.

Bullough:1981:BRB

[501] Vern L. Bullough. Book review: Doctor Bernard de Gordon: Professor and Practitioner by Luke E. Demaitre. Isis, 72(4):677, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231293.

Jervis:1981:BRB

[502] Jane Jervis. Book review: Regiomontanus-Studien by Günther Hamann. Isis, 72(4):677–678, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231294.

Kraft:1981:BRB

[503] Kent Kraft. Book review: Sphaera lucis: Studien zur Intelligibilität des Seienden im Kontext der mittelalterlichen Lichtspekulation by Klaus Hedwig. Isis, 72(4):678–679, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231295.

Savage-Smith:1981:BRB

[504] Emilie Savage-Smith. Book review: Aristoteles Semitico-Latinus. The Arabic Version of Aristotle’s Parts of Animals: Books XI-XIV of the Kitab al-H .ayawan by Remke Kruk. Isis, 72(4):679–680, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231296.

Glick:1981:BRB

[505] Thomas F. Glick. Book review: Les qanats: Une technique d’aquisition de l’eau by Henri Goblot. Isis, 72(4):680–681, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231297.

Shea:1981:BRBc

[506] William Shea. Book review: Lettere a Giovan Stefano Conti by Ruggiero Giuseppe Boscovich; Gino Arrighi. Isis, 72(4):681–682, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231298.

Grattan-Guinness:1981:BRB

[507] I. Grattan-Guinness. Book review: Colloque international et interdisciplinaire Jean-Henri Lambert, Mulhouse, 26–30 septembre 1977. Isis, 72(4):682, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231299.

Drake:1981:BRB

[508] Stillman Drake. Book review: Galileo and the Art of Reasoning: Rhetorical Foundations of Logic and Scientific Method by Maurice A. Finocchiaro. Isis, 72(4):682–683, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231300.

Jacob:1981:BRB

[509] James R. Jacob. Book review: Philosophers at War: The Quarrel between Newton and Leibniz by A. Rupert Hall. Isis, 72(4):683–684, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231301.

Feingold:1981:BRB

[510] Mordechai Feingold. Book review: English Scientific Virtuosi in the 16th and 17th Centuries by Barbara Shapiro; Robert G. Frank. Isis, 72 (4):684–685, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231302.

Christie:1981:BRB

[511] J. R. R. Christie. Book review: Works and Correspondence by Adam Smith; D. D. Raphael; A. S. Skinner; Works and Correspondence. Volume III: Essays on Philosophical Subjects by W. P. D. Wightman; J. C. Bryce; Dugald Stewart; I. S. Ross. Isis, 72(4):685–686, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231303.

Pyne:1981:BRB

[512] Stephen Pyne. Book review: Joseph Nicollet and His Map by Martha Coleman Bray. Isis, 72(4):686–687, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231304.

McCartney:1981:BRB

[513] Paul J. McCartney. Book review: Richard Griffith, 1784–1878: Papers Presented at the Centenary Symposium Organised by the Royal Dublin Society, 21 and 22 September 1978 by G. L. Herries Davies; R. Charles Mollan. Isis, 72(4):687, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231305.

Andrews:1981:BRBb

[514] Ilse Andrews. Book review: Alexander von Humboldt: His Portraits and Their Artists; A Documentary Iconography by Halina Nelken. Isis, 72 (4):687–688, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231306.

Heims:1981:BRB

[515] Stephen J. Heims. Book review: Reminiscences of Los Alamos, 1943– 1945 by Lawrence Badash; Joseph O. Hirschfelder; Herbert P. Broida. Isis, 72(4):688, December 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231307.

Dunbar:1981:BRBb

[516] G. S. Dunbar. Book review: The Life and Thought of Isaiah Bowman by Geoffrey J. Martin. Isis, 72(4):688–689, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231308.

Seidel:1981:BRB

[517] Robert Seidel. Book review: Quantum Physics in America, 1920–1935 by Katherine Russell Sopka. Isis, 72(4):689, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231309.

Anonymous:1981:BMd

[518] Anonymous. Back matter. Isis, 72(4):690–693, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231310.

Anonymous:1981:VI

[519] Anonymous. Volume information. Isis, 72(4):694–706, December 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231245.

Anonymous:1981:FMe

[520] Anonymous. Front matter. Isis, 72(5):1–20, ???? 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/230494.

Neu:1981:OHS

[521] John Neu. One hundred sixth critical bibliography of the history of science and its cultural influences (to January 1981). Isis, 72(5):5–19+21– 248, ???? 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230495.

Anonymous:1981:BMe

[522] Anonymous. Back matter. Isis, 72(5):??, ???? 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/230496.

Anonymous:1982:FMa

[523] Anonymous. Front matter. Isis, 73(??):1–3, ???? 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231432.

Anonymous:1982:CI

[524] Anonymous. Cumulative index. Isis, 73(??):4–167, ???? 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231433.

Anonymous:1982:BMa

[525] Anonymous. Back matter. Isis, 73(??):168, ???? 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231434.

Anonymous:1982:FMb

[526] Anonymous. Front matter. Isis, 73(1):1–5, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232077.

Thackray:1982:EAS

[527] Arnold Thackray. Editorial: On American science. Isis, 73(1):6–10, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232078.

Lepenies:1982:LND

[528] Wolf Lepenies. Linnaeus’s Nemesis divina and the concept of divine retaliation. Isis, 73(1):11–27, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232079.

Buck:1982:PWC

[529] Peter Buck. People who counted: Political arithmetic in the Eighteenth Century. Isis, 73(1):28–45, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232080.

Gooding:1982:EPT

[530] David Gooding. Empiricism in practice: Teleology, economy, and observation in Faraday’s physics. Isis, 73(1):46–67, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232081.

Sakellariadis:1982:DUE

[531] Spyros Sakellariadis. Descartes’s use of empirical data to test hypotheses. Isis, 73(1):68–76, March 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232082.

Elliott:1982:MAS

[532] Clark A. Elliott. Models of the American scientist: A look at collective biography. Isis, 73(1):77–93, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232083.

Welther:1982:PH

[533] Barbara L. Welther. “Pickering’s harem”. Isis, 73(1):94, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232084.

Whitrow:1982:BCB

[534] Magda Whitrow. Isis cumulative bibliography. Isis, 73(1):95, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232085.

Hirsh:1982:PSH

[535] Richard F. Hirsh. Proseminar on space history, 22 may 1981. Isis, 73(1): 96–97, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232086.

Anonymous:1982:NR

[536] Anonymous. News of record. Isis, 73(1):97, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232087.

Rousseau:1982:EMH

[537] G. S. Rousseau. Eloge: Marjorie Hope Nicolson, 18 February 1894– 9 March 1981. Isis, 73(1):98–99, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232088.

Whiteside:1982:BRB

[538] D. T. Whiteside. Book review: Never at Rest: A Biography of Isaac Newton. Isis, 73(1):100–107, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232089.

Schweber:1982:BRBa

[539] Silvan S. Schweber. Book review: Sir William Rowan Hamilton. Isis, 73(1):107–109, March 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232090.

Donovan:1982:BRB

[540] Arthur Donovan. Book review: Bulletin of Science, Technology and Society by Rustum Roy. Isis, 73(1):110, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232091.

Hoskin:1982:BRB

[541] Michael Hoskin. Book review: The Catalogues of Archives of Scientists, Compiled by the Contemporary Scientific Archives Centre. 1973– 1978, and Continuation by Jeannine Alston. Isis, 73(1):111, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232092.

Mitcham:1982:BRB

[542] Carl Mitcham. Book review: Technology and Values in American Civilization: A Guide to Information Sources by Stephen H. Cutcliffe; Judith A. Mistichelli; Christine M. Roysdon. Isis, 73(1):111–112, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232093.

Reingold:1982:BRB

[543] Nathan Reingold. Book review: New Directions in American Intellectual History by John Higham; P. K. Conkin. Isis, 73(1):112, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232094.

Lennox:1982:BRB

[544] James Lennox. Book review: On Scientific Discovery: The Erice Lectures, 1977 by Mirko Drazen Grmek; Robert S. Cohen; Guido Cimino. Isis, 73(1):112–113, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232095.

Manier:1982:BRB

[545] Edward Manier. Book review: Revolutions and Reconstructions in the Philosophy of Science by Mary Hesse. Isis, 73(1):113–114, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232096.

Healey:1982:BRB

[546] Richard A. Healey. Book review: Studies in the Foundations of Quantum Mechanics by Patrick Suppes. Isis, 73(1):114–115, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232097.

King:1982:BRB

[547] Lester S. King. Book review: Die inneren Verbindungen zwischen Philosophie und Medizin im 20. Jahrhundert by Dietrich von Engelhardt; Heinrich Schipperges. Isis, 73(1):115–116, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232098.

Middleton:1982:BRBa

[548] W. E. K. Middleton. Book review: The Construction and Use of a Thermometer, by James Six, F.R.S., Prefaced by an Account of His Life and Works and the Use of His Thermometer over Two Hundred Years by Jill Austin; Anita McConnell. Isis, 73(1):116, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232099.

Multhauf:1982:BRBa

[549] Robert P. Multhauf. Book review: Geomagnetic Instruments before 1900: An Illustrated Account of Their Construction and Use by Anita McConnell. Isis, 73(1):116–117, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232100.

Fee:1982:BRB

[550] Elizabeth Fee. Book review: Social Darwinism and English Thought: The Interaction between Biological and Social Theory by Greta Jones. Isis, 73(1):117, March 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232101.

Schroeder-Gudehus:1982:BRB

[551] Brigitte Schroeder-Gudehus. Book review: Colloques et sociétés: La régulation sociale by Claude Tapia. Isis, 73(1):117–118, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232102.

Richards:1982:BRBa

[552] Joan L. Richards. Book review: Peano: Life and Works of Giuseppe Peano by Hubert C. Kennedy. Isis, 73(1):118–119, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232103.

Gould:1982:BRB

[553] Henry W. Gould. Book review: Der Beginn der Determinantentheorie: Leibnizens nachgelassene Studien zum Determinantenkalkul. Textband by Eberhard Knobloch. Isis, 73(1):119, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232104.

Turnbull:1982:BRB

[554] David Turnbull. Book review: The Beginnings of Solid State Physics by Nevill Mott. Isis, 73(1):120, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232105.

Belloni:1982:BRBa

[555] Lanfranco Belloni. Book review: La radioattività da Becquerel a Rutherford by Giuseppe Bruzzaniti. Isis, 73(1):120–121, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232106.

Ihde:1982:BRB

[556] Aaron J. Ihde. Book review: H. Staudinger, H. Mark und K. H. Meyer: Thesen zur Grosse und Struktur der Makromolekule: Ursachen und Hintergrunde eines akademischen Disputes by Claus Priesner. Isis, 73(1): 121–122, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232107.

Carbonell:1982:BRB

[557] Armando Carbonell. Book review: Geography Yesterday and Tomorrow by E. H. Brown. Isis, 73(1):122, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232108.

Albritton:1982:BRB

[558] Claude Albritton. Book review: Northeastern Geology. Special Issue, Volume 3, Number 1 (Spring 1981): History of Geology in the Northeast: A Symposium by William M. Jordan. Isis, 73(1):123, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232109.

Burkhardt:1982:BRB

[559] Richard W. Burkhardt, Jr. Book review: The Origins and Rise of Ethology: The Science of the Natural Behavior of Animals by W. H. Thorpe. Isis, 73(1):123–124, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232110.

Roe:1982:BRBa

[560] Shirley A. Roe. Book review: The Iris in Eighteenth-Century Physiology by Renato G. Mazzolini. Isis, 73(1):124–125, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232111.

Powell:1982:BRB

[561] Jeffrey R. Powell. Book review: The Roving Naturalist: Travel Letters of Theodosius Dobzhansky by Bentley Glass. Isis, 73(1):125, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232112.

Richards:1982:BRBb

[562] Robert J. Richards. Book review: A History of Psychology: Main Currents in Psychological Thought by Thomas H. Leahey. Isis, 73(1):125– 127, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232113.

Sokal:1982:BRB

[563] Michael M. Sokal. Book review: The Architects of Adjustment: The History of the Psychological Profession in the United States by Donald S. Napoli. Isis, 73(1):127–128, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232114.

MacKenzie:1982:BRB

[564] Donald MacKenzie. Book review: The Meritocratic Intellect: Studies in the History of Educational Research by James V. Smith; David Hamilton. Isis, 73(1):128, March 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232115.

Maulitz:1982:BRB

[565] Russell C. Maulitz. Book review: Professionalizing Modern Medicine: Paris Surgeons and Medical Science and Institutions in the 18th Century by Toby Gelfand. Isis, 73(1):128–129, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232116.

Risse:1982:BRB

[566] Guenter B. Risse. Book review: The Concept of Heart Failure from Avicenna to Albertini by Saul Jarcho. Isis, 73(1):129–130, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232117.

Crouch:1982:BRB

[567] Tom D. Crouch. Book review: The Origins of the Turbojet Revolution by Edward W. Constant. Isis, 73(1):130–131, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232118.

Trimble:1982:BRB

[568] William F. Trimble. Book review: A Dream of Wings: Americans and the Airplane, 1875–1905 by Tom D. Crouch. Isis, 73(1):131–132, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232119.

Aspray:1982:BRB

[569] William Aspray. Book review: Early British Computers: The Story of Vintage Computers and the People Who Built Them by Simon Lavington; Project Whirlwind: The History of a Pioneer Computer by Kent C. Redmond; Thomas M. Smith. Isis, 73(1):132–133, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232120.

Schneider:1982:BRB

[570] Ivo Schneider. Book review: Die Pythagoreer: Religiose Bruderschaft und Schule der Wissenschaft by B. L. van der Waerden. Isis, 73(1):133– 135, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232121.

Sivin:1982:BRBa

[571] N. Sivin. Book review: Chung-kuo shihyu kung-yeh fa-chan shih. Volume I: Ku-tai te shih-yu yu t’ien-jan-ch’i. (A History of the Development of the Chinese Petroleum Industry. Volume I: Petroleum and Natural Gas in Ancient Times.) by Shen Li-sheng. Isis, 73(1):135, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232122.

McCluskey:1982:BRB

[572] Stephen C. McCluskey. Book review: Skywatchers of Ancient Mexico by Anthony F. Aveni. Isis, 73(1):135–136, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232123.

Shea:1982:BRB

[573] William R. Shea. Book review: Les fondements de l’optique moderne: Paralipomenes a Vitellion (1604) by Johann Kepler; Catherine Chevalley. Isis, 73(1):136, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232124.

Kish:1982:BRB

[574] George Kish. Book review: De la terre plate au globe terrestre: Une mutation epistemologique rapide (1480–1520) by W. G. L. Randles. Isis, 73(1):136–137, March 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232125.

Finocchiaro:1982:BRBa

[575] Maurice A. Finocchiaro. Book review: Studies on Christiaan Huygens by H. J. M. Bos; M. J. S. Rudwick; H. A. M. Snelders; R. P. W. Visser. Isis, 73(1):137–138, March 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232126.

Wetzels:1982:BRB

[576] Walter D. Wetzels. Book review: The Quest for the New Science: Language and Thought in Eighteenth-Century Science by Karl J. Fink; James W. Marchand. Isis, 73(1):138–139, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232127.

Hufbauer:1982:BRBa

[577] Karl Hufbauer. Book review: Tobias Mayer (1723–62): Pioneer of Enlightened Science in Germany by Eric G. Forbes. Isis, 73(1):139–140, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232128.

Hine:1982:BRB

[578] William L. Hine. Book review: Correspondance du P. Marin Mersenne, religieux minime. Volume XIV: 1646 by Marin Mersenne; Cornelis de Waard; Armand Beaulieu. Isis, 73(1):140, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232129.

Porter:1982:BRBa

[579] Roy Porter. Book review: I segni del tempo: Storia della terra e storia delle nazioni da Hooke a Vico by Paolo Rossi. Isis, 73(1):140–141, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232130.

Hall:1982:BRBa

[580] Marie Boas Hall. Book review: Selected Philosophical Papers of Robert Boyle by M. A. Stewart. Isis, 73(1):141, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232131.

Wilson:1982:BRB

[581] Curtis Wilson. Book review: Le procès des étoiles by Florence Trystram. Isis, 73(1):141–142, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232132.

Holmes:1982:BRB

[582] Frederic L. Holmes. Book review: Notes pour le rapport sur les progres de la physiologie by Claude Bernard; M. D. Grmek. Isis, 73(1):142– 143, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232133.

Dauben:1982:BRB

[583] Joseph W. Dauben. Book review: Studies in the Scientific and Mathematical Philosophy of Charles S. Peirce: Essays by Carolyn Eisele by Carolyn Eisele; R. M. Martin. Isis, 73(1):143–144, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232134.

Nye:1982:BRBa

[584] Mary Jo Nye. Book review: The Organization of Science and Technology in France 1808–1914 by Robert Fox; George Weisz. Isis, 73(1):144– 145, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232135.

Hutchison:1982:BRB

[585] Keith Hutchison. Book review: L’accueil des idées de Sadi Carnot et la technologie française de 1820 à 1860: De la légende à l’histoire by Pietro Redondi. Isis, 73(1):145–146, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232136.

Dufour:1982:BRB

[586] Paul Dufour. Book review: The Atom Bomb Spies by Harford Montgomery Hyde. Isis, 73(1):146–147, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232137.

Liebenau:1982:BRB

[587] Jonathan Liebenau. Book review: The History of Antibiotics: A Symposium by John Parascandola. Isis, 73(1):148, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232138.

Yoxen:1982:BRB

[588] Edward Yoxen. Book review: The Origins of Modern Biochemistry: A Retrospect on Proteins by P. R. Srinivasan; Joseph S. Fruton; John T. Edsall. Isis, 73(1):148–149, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232139.

Anonymous:1982:BMb

[589] Anonymous. Back matter. Isis, 73(1):150–154, March 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232140.

Anonymous:1982:FMc

[590] Anonymous. Front matter. Isis, 73(2):155–159, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231671.

Steneck:1982:GSI

[591] Nicholas H. Steneck. Greatrakes the stroker: The interpretations of historians. Isis, 73(2):160–177, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231672.

Kaplan:1982:GSI

[592] Barbara Beigun Kaplan. Greatrakes the stroker: The interpretations of his contemporaries. Isis, 73(2):178–185, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231673.

Rudwick:1982:CDL

[593] Martin J. S. Rudwick. Charles Darwin in London: The integration of public and private science. Isis, 73(2):186–206, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231674.

Servos:1982:DPF

[594] John W. Servos. A disciplinary program that failed: Wilder D. Bancroft and the Journal of Physical Chemistry, 1896–1933. Isis, 73(2):207–232, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231675.

Hutchison:1982:WHO

[595] Keith Hutchison. What happened to occult qualities in the scientific revolution? Isis, 73(2):233–253, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231676.

Stephens:1982:WRH

[596] Trent D. Stephens. The Wolffian Ridge: History of a misconception. Isis, 73(2):254–259, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231677.

Rawlins:1982:EGU

[597] Dennis Rawlins. Eratosthenes’ geodesy unraveled: Was there a highaccuracy Hellenistic astronomy? Isis, 73(2):259–265, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231678.

Holmes:1982:ASM

[598] Frederic L. Holmes. Award of the Sarton Medal for 1981. Isis, 73(2): 266–268, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231679.

Ihde:1982:DAS

[599] Aaron J. Ihde. Dexter award: Silver anniversary. Isis, 73(2):268–270, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231680.

North:1982:BRB

[600] J. D. North. Book review: Archimedes in the Middle Ages, by Marshall Clagett. Volume IV: A Supplement on the Medieval Latin Traditions of Conic Sections (1150–1566). Isis, 73(2):271–274, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231681.

Hankins:1982:BRB

[601] Thomas L. Hankins. Book review: The Ferment of Knowledge: Studies in the Historiography of Eighteenth-Century Science. Isis, 73(2):274– 279, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231682.

Joravsky:1982:BRB

[602] David Joravsky. Book review: Limits of Scientific Inquiry by Gerald Holton; Robert S. Morison. Isis, 73(2):280–281, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231683.

Caneva:1982:BRB

[603] Kenneth L. Caneva. Book review: Materialistische Wissenschaftsgeschichte: Naturtheorie und Entwicklungsdenken. Isis, 73(2):281, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231684.

Cambrosio:1982:BRB

[604] Alberto Cambrosio. Book review: Storia d’Italia. Volume III: Scienza e tecnica nella cultura e nella società dal Rinascimento ad oggi by Gianni Micheli. Isis, 73(2):281–282, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231685.

Miller:1982:BRB

[605] David Philip Miller. Book review: The Royal Society Catalogue of Portraits by Norman Robinson. Isis, 73(2):282, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231686.

Multhauf:1982:BRBb

[606] Robert P. Multhauf. Book review: Bibliotheca de re metallica: The Herbert Clark Hoover Collection of Mining and Metallurgy by David Kuhner; Tania Rizzo. Isis, 73(2):283, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231687.

Elliott:1982:BRB

[607] Clark A. Elliott. Book review: Prominent Scientists: An Index to Collective Biographies by Paul A. Pelletier. Isis, 73(2):283–284, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231688.

Finocchiaro:1982:BRBb

[608] Maurice A. Finocchiaro. Book review: Italian Studies in the Philosophy of Science by Maria Luisa Dalla Chiara; Carolyn R. Fawcett. Isis, 73(2): 284–285, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231689.

Wallace:1982:BRBa

[609] William A. Wallace. Book review: Theory Change, Ancient Axiomatics, and Galileo’s Methodology. Proceedings of the 1978 Pisa Conference on the History and Philosophy of Science, Volume I. by Jaakko Hintikka; David Gruender; Evandro Agazzi. Isis, 73(2):285–286, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231690.

Brush:1982:BRB

[610] Stephen G. Brush. Book review: Probabilistic Thinking, Thermodynamics, and the Interaction of the History and Philosophy of Science: Proceedings of the 1978 Pisa Conference on the History and Philosophy of Science, Volume II by Jaakko Hintikka; David Gruender; Evandro Agazzi. Isis, 73(2):286–287, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231691.

Burian:1982:BRB

[611] Richard M. Burian. Book review: Scientific Discovery: Case Studies by Thomas Nickles. Isis, 73(2):287–288, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231692.

Moran:1982:BRBa

[612] Bruce T. Moran. Book review: The Clockwork Universe: German Clocks and Automata, 1550–1650 by Klaus Maurice; Otto Mayr. Isis, 73(2):288– 289, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231693.

Middleton:1982:BRBb

[613] W. E. Knowles Middleton. Book review: Essays on the History of the Microscope by Gerard L’E. Turner; Antique Scientific Instruments by Gerard L’E. Turner. Isis, 73(2):289–290, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231694.

Clark:1982:BRB

[614] Linda Loeb Clark. Book review: The Making of Frenchmen: Current Directions in the History of Education in France, 1679–1979 by Donald N. Baker; Patrick J. Harrigan. Isis, 73(2):290–291, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231695.

McClellan:1982:BRB

[615] James E. McClellan. Book review: La société néerlandaise et ses gradués, 1575–1814 by Wilhelmus T. M. Frijhoff. Isis, 73(2):291, June 1982.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231696.

Kranzberg:1982:BRB

[616] Melvin Kranzberg. Book review: Whither Technology? by Lyndsay Farrall; Liberation and Control: The Uses of Knowledge and Power by David Wade Chambers. Isis, 73(2):291–292, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231697.

Moore:1982:BRB

[617] James R. Moore. Book review: Marx et Darwin: Le grand affrontement by Yves Christen. Isis, 73(2):292–293, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231698.

Pfetsch:1982:BRB

[618] Frank R. Pfetsch. Book review: The Atom Besieged: Extraparliamentary Dissent in France and Germany by Dorothy Nelkin; Michael Pollak. Isis, 73(2):293–294, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231699.

Vucinich:1982:BRB

[619] Alexander Vucinich. Book review: Science, Technology and the Future: Soviet Scientists’ Analysis of the Problems of and Prospects for the Development of Science and Technology and Their Role in Society by E. P. Velikhov; J. M. Gvishiani; S. B. Mikulinsky. Isis, 73(2):294–295, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231700.

Dunlap:1982:BRB

[620] Thomas R. Dunlap. Book review: The Politics of Mistrust: Estimating American Oil and Gas Resources by Aaron Wildavsky; Ellen Tenenbaum. Isis, 73(2):295, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231701.

Meikle:1982:BRB

[621] Jeffrey L. Meikle. Book review: The Myths of Information: Technology and Post-Industrial Culture by Kathleen Woodward; The Technological Imagination: Theories and Fictions by Teresa de Lauretis; Andreas Huyssen; Kathleen Woodward. Isis, 73(2):295–296, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231702.

Restivo:1982:BRBa

[622] Sal Restivo. Book review: Code of the Quipu: A Study in Media, Mathematics, and Culture by Marcia Ascher; Robert Ascher. Isis, 73(2):296– 297, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231703.

Grattan-Guinness:1982:BRBa

[623] I. Grattan-Guinness. Book review: A History of the Calculus of Variations: From the 17th through the 19th Century by Herman H. Goldstine. Isis, 73(2):297, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231704.

LeGrand:1982:BRB

[624] H. E. Le Grand. Book review: Revue de l’essai de statique chimique: Edition critique d’un projet de seconde edition de L’essai de statique chimique by Claude-Louis Berthollet; Michelle Sadoun-Goupil. Isis, 73 (2):298, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231705.

Melhado:1982:BRB

[625] Evan M. Melhado. Book review: Gay-Lussac: La carriére et l’oeuvre d’un chimiste français durant la premiere moitié du XIXe siècle. Isis, 73(2):298–299, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231706.

Gingerich:1982:BRBa

[626] Owen Gingerich. Book review: 600 Jahre Astronomie in Nurnberg by Kurt Pilz. Isis, 73(2):299, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231707.

DeVorkin:1982:BRB

[627] David H. DeVorkin. Book review: Oort and the Universe: A Sketch of Oort’s Research and Person by Hugo van Woerden; Willem N. Brouw; Henk C. van de Hulst. Isis, 73(2):300, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231708.

Whorton:1982:BRB

[628] James Whorton. Book review: From Arsenic to DDT: A History of Entomology in Western Canada by Paul W. Riegert. Isis, 73(2):300– 301, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231709.

Becker:1982:BRB

[629] Stanley L. Becker. Book review: A List of References for the History of Agricultural Science in America by Margaret W. Rossiter. Isis, 73(2): 301, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231710.

Stocking:1982:BRB

[630] George W. Stocking, Jr. Book review: Savages and Scientists: The Smithsonian Institution and the Development of American Anthropology, 1846–1910 by Curtis M. Hinsley. Isis, 73(2):301–302, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231711.

Danziger:1982:BRB

[631] Kurt Danziger. Book review: Psychology: Theoretical-Historical Perspectives by R. W. Rieber; Kurt Salzinger. Isis, 73(2):302–303, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231712.

Benison:1982:BRB

[632] Saul Benison. Book review: Centenary of Index Medicus, 1879–1979 by John B. Blake. Isis, 73(2):303–304, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231713.

Rosenberg:1982:BRB

[633] Charles E. Rosenberg. Book review: American Medicine in Transition, 1840–1910 by John S. Haller. Isis, 73(2):304, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231714.

Rosenkrantz:1982:BRB

[634] Barbara Gutmann Rosenkrantz. Book review: Times, Places, and Persons: Aspects of the History of Epidemiology by Abraham M. Lilienfeld. Isis, 73(2):304–305, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231715.

Smith:1982:BRBa

[635] Cyril Stanley Smith. Book review: Persian Metal Technology, 790 to 1300 A.D. by James W. Allan; Aspects of Ancient Indian Technology: A Research Based on Scientific Methods by H. C. Bhardwaj. Isis, 73(2):

305–306, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231716.

Calhoun:1982:BRB

[636] Daniel Calhoun. Book review: The Engineering Drawings of Benjamin Henry Latrobe by Darwin H. Stapleton. Isis, 73(2):306–307, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231717.

Todd:1982:BRB

[637] Edmund N. Todd. Book review: Technik-Geschichte: Historische Beitrage und neuere Ansatze by Ulrich Troitzsch; Gabriele Wohlauf. Isis, 73(2):307–308, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231718.

Lloyd:1982:BRB

[638] G. E. R. Lloyd. Book review: Science and the Sciences in Plato by John P. Anton. Isis, 73(2):308–309, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231719.

Copenhaver:1982:BRB

[639] Brian P. Copenhaver. Book review: Recherches sur une technique divinatoire: La geomancie dans l’occident medieval by Therese Charmasson. Isis, 73(2):309, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231720.

Kealey:1982:BRB

[640] Edward J. Kealey. Book review: Medieval Woman’s Guide to Health: The First English Gynecological Handbook by Beryl Rowland. Isis, 73 (2):310, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231721.

Goldstein:1982:BRBa

[641] Bernard R. Goldstein. Book review: Geschichte des arabischen Schrifttums. Volume VII: Astrologie-Meteorologie und Verwandtes, bis ca. 430 H by Fuat Sezgin. Isis, 73(2):310–311, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231722.

Rocher:1982:BRB

[642] Ludo Rocher. Book review: Census of the Exact Sciences in Sanskrit. Series A, Volume IV by David Pingree. Isis, 73(2):311–312, June 1982.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231723.

Eamon:1982:BRB

[643] William Eamon. Book review: Humanismus und Naturwissenschaften by Rudolf Schmitz; Fritz Krafft. Isis, 73(2):312, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231724.

Siraisi:1982:BRB

[644] Nancy G. Siraisi. Book review: Metodo e sistema delle scienze nel pensiero di Ulisse Aldrovandi by Sandra Tugnoli Pattaro. Isis, 73(2):313, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231725.

Taylor:1982:BRB

[645] Kenneth L. Taylor. Book review: The Earth Generated and Anatomized: An Early Eighteenth Century Theory of the Earth by William Hobbs; Roy Porter. Isis, 73(2):313–314, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231726.

Shapin:1982:BRB

[646] Steven Shapin. Book review: Science and Society in Restoration England by Michael Hunter. Isis, 73(2):314–315, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231727.

Winder:1982:BRB

[647] Marianne Winder. Book review: Alle de Brieven van Antoni van Leeuwenhoek: The Collected Letters of Antoni van Leeuwenhoek by Antoni van Leeuwenhoek; L. C. Palm. Isis, 73(2):315, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231728.

Oppenheimer:1982:BRB

[648] Jane Oppenheimer. Book review: Matter, Life, and Generation: Eighteenth-Century Embryology and the Haller-Wolff Debate by Shirley A. Roe. Isis, 73(2):315–316, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231729.

Hackmann:1982:BRB

[649] Willem D. Hackmann. Book review: Het Gezelschap der Hollandsche Scheikundigen: Amsterdamse Chemici uit het Einde van de achttiende Eeuw by H. A. M. Snelders. Isis, 73(2):316–317, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231730.

Adams:1982:BRB

[650] Thomas M. Adams. Book review: Cabanis: Enlightenment and Medical Philosophy in the French Revolution by Martin S. Staum. Isis, 73(2): 317–318, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231731.

Roe:1982:BRBb

[651] Shirley A. Roe. Book review: Venus Physique, suivi de Lettre sur le progres des sciences by Pierre Louis Moreau de Maupertuis; L’ordre et les monstres: Le debat sur l’origine des deviations anatomiques au XVI-IIe siècle by Patrick Tort. Isis, 73(2):318–319, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231732.

Pyenson:1982:BRBa

[652] Lewis Pyenson. Book review: Epistemological and Social Problems of the Sciences in the Early Nineteenth Century by Hans Jahnke; Michael Otte. Isis, 73(2):319–320, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231733.

Carter:1982:BRB

[653] Edward C. Carter, II. Book review: The Papers of Joseph Henry. Volume III: January 1836–December 1837: The Princeton Years by Nathan Reingold. Isis, 73(2):320–321, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231734.

Granshaw:1982:BRB

[654] Lindsay Granshaw. Book review: Reckoning with the Beast: Animals, Pain, and Humanity in the Victorian Mind by James Turner. Isis, 73(2): 321–322, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231735.

Hufbauer:1982:BRBb

[655] Karl Hufbauer. Book review: Hans Bethe: Prophet of Energy by Jeremy Bernstein. Isis, 73(2):322, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231736.

Buck:1982:BRB

[656] Peter Buck. Book review: An American Transplant: The Rockefeller Foundation and Peking Union Medical College by Mary Brown Bullock. Isis, 73(2):322–323, June 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231737.

Moyer:1982:BRB

[657] Albert E. Moyer. Book review: The Impact of Modern Scientific Ideas on Society: In Commemoration of Einstein by Colette M. Kinnon; A. N. Kholodilin; J. G. Richardson. Isis, 73(2):323–324, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231738.

Gingras:1982:BRB

[658] Yves Gingras. Book review: Radar Development in Canada: The Radio Branch of the National Research Council of Canada, 1939–1946 by W. E. Knowles Middleton. Isis, 73(2):324, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231739.

Mack:1982:BRB

[659] Pamela E. Mack. Book review: Beyond the Atmosphere: Early Years of Space Science by Homer E. Newell. Isis, 73(2):324–325, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231740.

Belloni:1982:BRBb

[660] Lanfranco Belloni. Book review: Vklad Akademika A. F. Ioffe v stanovlenie iadernoi fiziki v SSSR. Isis, 73(2):326, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231741.

Anonymous:1982:BMc

[661] Anonymous. Back matter. Isis, 73(2):327–330, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231742.

Anonymous:1982:FMd

[662] Anonymous. Front matter. Isis, 73(3):331–335, September 1982. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231438.

Home:1982:GEH

[663] R. W. Home. Guest editorial: History of science in Australia. Isis, 73 (3):336–342, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231439.

Friedman:1982:CPF

[664] Robert Marc Friedman. Constituting the polar front, 1919–1920. Isis, 73 (3):343–362, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231440.

Meyer-Thurow:1982:IIC

[665] Georg Meyer-Thurow. The industrialization of invention: A case study from the German chemical industry. Isis, 73(3):363–381, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231441.

Dannenfeldt:1982:ASO

[666] Karl H. Dannenfeldt. Ambergris: The search for its origin. Isis, 73(3): 382–397, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231442.

Merchant:1982:BCR

[667] Carolyn Merchant. Isis’ consciousness raised. Isis, 73(3):398–409, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231443.

Cohen:1982:NCL

[668] I. Bernard Cohen. Newton’s copy of Leibniz’s Theodicee: With some remarks on the turned-down pages of books in Newton’s library. Isis, 73 (3):410–414, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231444.

Numbers:1982:AMH

[669] Ronald L. Numbers, David C. Lindberg, and Sally Gregory Kohlstedt. Annual meeting of the History of Science Society 27–30 December 1981. Isis, 73(3):415–421, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231445.

Glick:1982:SMP

[670] Thomas F. Glick. Symposium on methodology and problems in the history of science and technology, Mexico City, 26–28 October 1981. Isis, 73(3):421, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231446.

Kranzberg:1982:EJW

[671] Melvin Kranzberg. Eloge: John W. Abrams, 15 December 1913–11 July 1981. Isis, 73(3):422–424, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231447.

Jacob:1982:EDF

[672] Margaret Jacob and Edward Gosselin. Eloge: Dame Frances Amelia Yates, 28 November 1899–29 September 1981. Isis, 73(3):424–426, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231448.

Steiner:1982:BRB

[673] George Steiner. Book review: Night Thoughts of a Classical Physicist. Isis, 73(3):427–429, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231449.

Ruse:1982:BRBa

[674] Michael Ruse. Book review: The Mismeasure of Man. Isis, 73(3):430– 431, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231450.

Schmitt:1982:BRB

[675] Charles B. Schmitt. Book review: Harvey and the Oxford Physiologists: A Study of Scientific Ideas and Social Interaction. Isis, 73(3):432–433, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231451.

Crosland:1982:BRB

[676] Maurice Crosland. Book review: The Analytic Spirit: Essays on the History of Science in Honor of Henry Guerlac by Harry Woolf. Isis, 73 (3):434–435, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231452.

Blackwell:1982:BRB

[677] Richard J. Blackwell. Book review: La science dans la philosophie: Les recherches epistemologiques d’Alexandre Koyre by Gerard Jorland. Isis, 73(3):435, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231453.

Hall:1982:BRBb

[678] A. Rupert Hall. Book review: From Art to Science: Seventy-Two Objects Illustrating the Nature of Discovery by Cyril Stanley Smith; A Search for Structure: Selected Essays on Science, Art, and History by Cyril Stanley Smith. Isis, 73(3):435–437, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231454.

Guedon:1982:BRB

[679] Jean-Claude Guedon. Book review: Development of Science Publishing in Europe by A. J. Meadows. Isis, 73(3):437–438, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231455.

Thackray:1982:BRB

[680] Arnold Thackray. Book review: Natural History Manuscript Resources in the British Isles by Gavin D. R. Bridson; Valerie C. Phillips; Anthony P. Harvey. Isis, 73(3):438–439, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231456.

Goldstein:1982:BRBb

[681] Bernard R. Goldstein. Book review: The History of Science: A Collection of Manuscripts from the Library of the Jewish Theological Seminary. Isis, 73(3):439–440, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231457.

Hindle:1982:BRB

[682] Brooke Hindle. Book review: Technical Americana: A Checklist of Technical Publications Printed before 1831 by Evald Rink. Isis, 73(3):440–441, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231458.

Cummins:1982:BRB

[683] Robert Cummins. Book review: Mind in Science: A History of Explanations in Psychology and Physics by Richard L. Gregory. Isis, 73(3):441, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231459.

Brandom:1982:BRB

[684] Robert B. Brandom. Book review: Analytical Philosophy of Technology by Friedrich Rapp; Stanley R. Carpenter; Theodor Langenbruch. Isis, 73 (3):441–442, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231460.

Numbers:1982:BRB

[685] Ronald L. Numbers. Book review: Darwin to Einstein: Historical Studies on Science and Belief by Colin Chant; John Fauvel; Darwin to Einstein: Primary Sources on Science and Belief by Noel G. Coley; Vance M. D. Hall. Isis, 73(3):442–443, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231461.

Oldroyd:1982:BRB

[686] David Oldroyd. Book review: Science, Ideology, and World View: Essays in the History of Evolutionary Ideas by John C. Greene. Isis, 73(3): 443–444, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231462.

Metcalf:1982:BRB

[687] Robert L. Metcalf. Book review: DDT: Scientists, Citizens, and Public Policy by Thomas R. Dunlap. Isis, 73(3):444–445, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231463.

Cohen:1982:BRB

[688] Cheryl Sandra Cohen. Book review: Reflections on Science and the Media by June Goodfield. Isis, 73(3):445–446, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231464.

Ravetz:1982:BRB

[689] Jerome R. Ravetz. Book review: Science, Technology and the Labour Process by Les Levidow; Bob Young. Isis, 73(3):446, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231465.

Nye:1982:BRBb

[690] Mary Jo Nye. Book review: Actes du colloque: Enseignement de l’histoire des sciences aux scientifiques by J. Dhombres. Isis, 73(3):446–447, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231466.

Turner:1982:BRBa

[691] Frank M. Turner. Book review: Cambridge before Darwin: The Ideal of a Liberal Education, 1800–60 by Martha McMackin Garland. Isis, 73 (3):447–448, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231467.

Grabiner:1982:BRB

[692] Judith V. Grabiner. Book review: Sophie Germain: An Essay in the History of the Theory of Elasticity by Louis L. Bucciarelli; Nancy Dworsky. Isis, 73(3):448–449, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231468.

Grattan-Guinness:1982:BRBb

[693] I. Grattan-Guinness. Book review: Il calcolo sublime: Storia dell’ analisi matematica da Euler a Weierstrass by Umberto Bottazzini; The Origins of Cauchy’s Rigorous Calculus by Judith V. Grabiner. Isis, 73(3):449– 450, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231469.

Restivo:1982:BRBb

[694] Sal Restivo. Book review: Mathematics as a Cultural System by Raymond L. Wilder. Isis, 73(3):450–451, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231470.

Williams:1982:BRB

[695] L. Pearce Williams. Book review: Conceptions of Ether: Studies in the History of Ether Theories, 1740–1900 by G. N. Cantor; M. J. S. Hodge. Isis, 73(3):451–452, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231471.

Kopperl:1982:BRB

[696] Sheldon J. Kopperl. Book review: Electrochemistry: History and Theory by Wilhelm Ostwald. Isis, 73(3):452–453, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231472.

Cawood:1982:BRB

[697] J. A. Cawood. Book review: To the Farthest Ends of the Earth: The History of the Royal Geographical Society 1830–1980 by Ian Cameron. Isis, 73(3):453, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231473.

Hazen:1982:BRB

[698] Robert M. Hazen. Book review: The New Madrid Earthquakes by James L. Penick. Isis, 73(3):454, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231474.

Wolff:1982:BRB

[699] Torben Wolff. Book review: Oceanography: The Past by Mary Sears; Daniel Merriman. Isis, 73(3):454–455, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231475.

Fruton:1982:BRB

[700] Joseph S. Fruton. Book review: Foundation Stones of Biochemistry: Classical Papers on Enzymes and pH by T. R. Caine Boyde. Isis, 73(3): 455, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231476.

Maienschein:1982:BRB

[701] Jane Maienschein. Book review: A History of the Study of Human Growth by J. M. Tanner. Isis, 73(3):455–456, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231477.

Egerton:1982:BRBa

[702] Frank N. Egerton. Book review: The Shaping of Cambridge Botany: A Short History of Whole-Plant Botany in Cambridge from the Time of Ray into the Present Century by S. M. Walters. Isis, 73(3):456–457, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231478.

Bringmann:1982:BRB

[703] Wolfgang G. Bringmann and Gustav A. Ungerer. Book review: Gli studi di psicologia in Italia: Aspetti teorici scientifici e ideologici by Guido Cimino; Nino Dazzi. Isis, 73(3):457, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231479.

Erickson:1982:BRB

[704] Paul Erickson. Book review: Adventures and Discoveries: The Autobiography of Carleton S. Coon, Anthropologist and Explorer by Carleton S. Coon. Isis, 73(3):457–458, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231480.

Hinsley:1982:BRB

[705] C. M. Hinsley, Jr. Book review: Four Anthropologists: An American Science in Its Early Years by Joan Mark. Isis, 73(3):458–459, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231481.

Lenoir:1982:BRB

[706] Timothy Lenoir. Book review: Die naturhistorische Schule 1825–1845: Ein Beitrag zur Geschichte der klinischen Medizin in Deutschland by Johanna Bleker. Isis, 73(3):459, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231482.

Lock:1982:BRB

[707] Margaret Lock. Book review: When the Twain Meet: The Rise of Western Medicine in Japan by John Z. Bowers. Isis, 73(3):459–460, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231483.

Goldstein:1982:BRBc

[708] Jan Goldstein. Book review: Les debuts de la psychanalyse en France, 1895–1926 by Jean–Pierre Mordier. Isis, 73(3):460, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231484.

Leavitt:1982:BRB

[709] Judith Walzer Leavitt. Book review: Trial by Medicine: Insanity and Responsibility in Victorian Trials by Roger Smith. Isis, 73(3):460–461, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231485.

McCraw:1982:BRB

[710] Thomas K. McCraw. Book review: The British Electrical Industry 1875– 1914: The Economic Returns of a New Technology by I. C. R. Byatt; Electricity before Nationalisation: A Study of the Development of the Electricity Supply Industry in Britain to 1948 by Leslie Hannah. Isis, 73 (3):461–462, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231486.

Pyenson:1982:BRBb

[711] Lewis Pyenson. Book review: The Tools of Empire: Technology and European Imperialism in the Nineteenth Century by Daniel R. Headrick. Isis, 73(3):462–463, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231487.

Scarborough:1982:BRB

[712] John Scarborough. Book review: La maladie de l’ame: Etude sur la relation de l’ame et du corps dans la tradition medico-philosophique antique by Jackie Pigeaud. Isis, 73(3):463–464, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231488.

Voigts:1982:BRB

[713] Linda Ehrsam Voigts. Book review: Medieval Medicus: A Social History of Anglo-Norman Medicine by Edward J. Kealey. Isis, 73(3):464–465, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231489.

Hansen:1982:BRBa

[714] Bert Hansen. Book review: Histoire litteraire de la France. Volume 41: Suite du quatorzième siècle by Charles Samaran. Isis, 73(3):465–466, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231490. See corri-

gendum [853].

VanEgmond:1982:BRB

[715] Warren Van Egmond. Book review: Die erste deutsche Algebra aus dem Jahre 1481: Nach einer Handschrift aus C 80 Dresdensis by Kurt Vogel. Isis, 73(3):466, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231491.

Moesgaard:1982:BRB

[716] Kristian Peder Moesgaard. Book review: Starowolski’s Biographies of Copernicus by Erna Hilfstein. Isis, 73(3):466–467, September 1982. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231492.

Lamb:1982:BRB

[717] Ursula Lamb. Book review: Christophe Colomb by Jacques Heers. Isis, 73 (3):467–468, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231493.

Gingerich:1982:BRBb

[718] Owen Gingerich. Book review: Materialien zu den Ephemeriden von Johannes Kepler by Volker Bialas; Elli Papadimitriou. Isis, 73(3):468– 469, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231494.

Coleman:1982:BRB

[719] William Coleman. Book review: Linnaeus: Progress and Prospects in Linnaean Research by Gunnar Broberg. Isis, 73(3):469–470, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231495.

Dobbs:1982:BRB

[720] B. J. T. Dobbs. Book review: The Expanding Force in Newton’s Cosmos, as Shown in His Unpublished Papers by David Castillejo. Isis, 73(3):470, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231496.

Neve:1982:BRB

[721] Michael Neve. Book review: James Hutton’s Medical Dissertation by Arthur Donovan; Joseph Prentiss. Isis, 73(3):470–471, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231497.

Wisan:1982:BRB

[722] Winifred Lovell Wisan. Book review: Galileo’s Notes on Motion: Arranged in Probable Order of Composition and Presented in Reduced Facsimile by Stillman Drake. Isis, 73(3):471–472, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231498.

Lesch:1982:BRB

[723] John E. Lesch. Book review: Felice Fontana 1730–1805: An Annotated Bibliography by Peter K. Knoefel. Isis, 73(3):472, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231499.

Multhauf:1982:BRBc

[724] Robert P. Multhauf. Book review: L’arte vetraria: 1612 by Antonio Neri; Rosa Barovier Mentasi. Isis, 73(3):472–473, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231500.

Jacob:1982:BRB

[725] Margaret Jacob. Book review: The French Prophets: The History of a Millenarian Group in Eighteenth-Century England by Hillel Schwartz. Isis, 73(3):473–474, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231501.

Bell:1982:BRB

[726] James F. Bell. Book review: Oscillations et stabilité selon Foucault: Critique historique et experimentale by Paul Acloque. Isis, 73(3):474, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231502.

Richards:1982:BRBc

[727] Robert J. Richards. Book review: Die Umweltlehre Jakob von Uexkulls in ihrer Bedeutung fur die Entwicklung der vergleichenden Verhaltenesforschung by Jutta Schmidt. Isis, 73(3):474–475, September 1982. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231503.

Allen:1982:BRB

[728] D. E. Allen. Book review: The Heyday of Natural History, 1820– 1870 by Lynn Barber. Isis, 73(3):475–476, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231504.

Mayr:1982:BRBa

[729] Ernst Mayr. Book review: Concordance to Darwin’s Origin of Species, First Edition by Paul H. Barrett; Donald J. Weinshank; Timothy T. Gottleber. Isis, 73(3):476–477, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231505.

Colp:1982:BRB

[730] Ralph Colp, Jr. Book review: Charles Darwin: A Man of Enlarged Curiosity by Peter Brent. Isis, 73(3):477–478, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231506.

Goetzmann:1982:BRB

[731] William H. Goetzmann. Book review: A Scientist in American Life: The Essays and Lectures of Joseph Henry by Joseph Henry; Arthur P. Molella; Nathan Reingold; Marc Rothenberg; Joan F. Steiner; Kathleen Waldenfels. Isis, 73(3):478–479, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231507.

Garber:1982:BRB

[732] Elizabeth Garber. Book review: The Papers of Joseph Henry. Volume IV: January 1838–December 1840: The Princeton Years by Nathan Reingold. Isis, 73(3):479–480, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231508.

Smith:1982:BRBb

[733] Crosbie Smith. Book review: James Clerk Maxwell: A Biography by Ivan Tolstoy. Isis, 73(3):480, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231509.

Robinson:1982:BRB

[734] Daniel N. Robinson. Book review: IQ and Mental Testing: An Unnatural Science and Its Social History by Brian Evans; Bernard Waites. Isis, 73 (3):480–481, September 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231510.

Abir-Am:1982:BRB

[735] Pnina Abir-Am. Book review: An Imagined World: A Story of Scientific Discovery by June Goodfield. Isis, 73(3):481–482, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231511.

Brittain:1982:BRB

[736] James E. Brittain. Book review: Ferdinand Braun: A Life of the Nobel Prizewinner and Inventor of the Cathode-Ray Oscilloscope by Friedrich Kurylo; Charles Susskind. Isis, 73(3):482–483, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231512.

Melcher:1982:BRA

[737] Horst Melcher. Book review: Arthur I. Miller: Albert Einstein’s Special Theory of Relativity: Emergence (1905) and Early Interpretation (1905–11). Isis, 73(3):483–484, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231513.

Edsall:1982:BRB

[738] John T. Edsall. Book review: The Nobel Duel: Two Scientists’ 21-year Race to Win the World’s Most Coveted Research Prize by Nicholas Wade. Isis, 73(3):484–485, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231514.

Anonymous:1982:BMd

[739] Anonymous. Back matter. Isis, 73(3):486–490, September 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231515.

Anonymous:1982:FMe

[740] Anonymous. Front matter. Isis, 73(4):491–495, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232142.

Mark:1982:FFA

[741] Joan Mark. Francis La flesche: The American Indian as anthropologist. Isis, 73(4):496–510, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232143.

Dobbs:1982:NAH

[742] B. J. T. Dobbs. Newton’s alchemy and his theory of matter. Isis, 73 (4):511–528, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232144.

Porter:1982:SCE

[743] Jonathan Porter. The scientific community in early modern China. Isis, 73(4):529–544, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232145.

Perrin:1982:LIP

[744] C. E. Perrin. A lost identity: Philippe Frédéric, Baron de Dietrich (1748– 1793). Isis, 73(4):545–551, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232146.

Reingold:1982:E

[745] Nathan Reingold. To the Editor. Isis, 73(4):551, December 1982. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232147.

Forbes:1982:IOH

[746] Eric G. Forbes. The international organization of the history of science: A personal view. Isis, 73(4):552–557, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232148.

Nye:1982:NR

[747] Mary Jo Nye. News of record. Isis, 73(4):557–560, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232149.

Frangsmyr:1982:ESL

[748] Tore Frangsmyr. Eloge: Sten Lindroth, 28 December 1914–1 September 1980. Isis, 73(4):561–562, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232150.

Turner:1982:BRBb

[749] Frank M. Turner. Book review: The Parliament of Science: The British Association for the Advancement of Science 1831–1981. Isis, 73(4):563– 566, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232151.

MacLeod:1982:BRB

[750] Roy MacLeod. Book review: A Guide to the Culture of Science, Technology, and Medicine. Isis, 73(4):567–571, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232152.

Ziman:1982:BRB

[751] J. M. Ziman. Book review: T. S. Kuhn and Social Science by Barry Barnes. Isis, 73(4):572, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232153.

Saliba:1982:BRB

[752] George Saliba. Book review: The Coming of the Age of Iron by Theodore A. Wertime; James D. Muhly. Isis, 73(4):572–573, December 1982. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232154.

deSollaPrice:1982:BRB

[753] Derek deSolla Price. Book review: Physics Citation Index, 1920–1929. Isis, 73(4):573–574, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232155.

Reeds:1982:BRB

[754] Karen Reeds. Book review: Guide to the Botanical Records and Papers in the Archives of the Hunt Institute. Part I by Michael T. Stieber; Anita L. Karg; Donna M. Connelly; Karen K. Britz. Isis, 73(4):574, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232156.

Rynasiewicz:1982:BRB

[755] Robert Rynasiewicz. Book review: The Rationality of Science by W. H. Newton-Smith. Isis, 73(4):574–575, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232157.

Hansen:1982:BRBb

[756] Bert Hansen. Book review: Early Scientific Instruments by Nigel Hawkes; Marine Navigation Instruments by Jean Randier; John E. Powell; A Catalogue of European Scientific Instruments in the Department of Medieval and Later Antiquities of the British Museum by F. A. B. Ward. Isis, 73(4):575–576, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232158.

Churchill:1982:BRB

[757] Frederick B. Churchill. Book review: The Descent of Darwin: The Popularization of Darwinism in Germany, 1860–1914 by Alfred Kelly. Isis, 73 (4):576–578, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232159.

Paradis:1982:BRB

[758] James Paradis. Book review: Poetry Realized in Nature: Samuel Taylor Coleridge and Early Nineteenth-Century Science by Trevor H. Levere.

Isis, 73(4):578, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232160.

Ruse:1982:BRBb

[759] Michael Ruse. Book review: Morality as a Biological Phenomenon: The Presuppositions of Sociobiological Research by Gunther S. Stent. Isis, 73(4):579, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232161.

Gortler:1982:BRB

[760] Leon Gortler. Book review: Stereochemistry by O. Bertrand Ramsay. Isis, 73(4):579–580, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232162.

Moran:1982:BRBb

[761] Bruce T. Moran. Book review: Wandel des Weltbildes: Astronomie, Physik und Messtechnik in der Kulturgeschichte by Jurgen Teichmann. Isis, 73(4):580–581, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232163.

Dunbar:1982:BRB

[762] G. S. Dunbar. Book review: Carl Ritter: Zur europaisch-amerikanischen Geographie an der Wende vom 18. zum 19. Jahrhundert by Manfred Buttner. Isis, 73(4):581–582, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232164.

Mayr:1982:BRBb

[763] Ernst Mayr. Book review: Geologen der Goethezeit by H. Prescher; D. Beeger. Isis, 73(4):582, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232165.

Laudan:1982:BRB

[764] Rachel Laudan. Book review: Disziplingeschichte als wissenschaftliche Selbstreflexion der historischen Wissenschaftsforschung: Eine Darstellung unter Heranziehung von Fallstudien der Wissenschaftsgeschichte der Geophysik by Wilfried Schroder. Isis, 73(4):582–583, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232166.

Rossiter:1982:BRB

[765] Margaret W. Rossiter. Book review: Introduction to the History of Plant Pathology by G. C. Ainsworth. Isis, 73(4):583–584, December 1982. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232167.

Bowler:1982:BRB

[766] Peter J. Bowler. Book review: Evolution: Genesis and Revelations by C. Leon Harris. Isis, 73(4):584, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232168.

Egerton:1982:BRBb

[767] Frank N. Egerton. Book review: Saving the Prairies: The Life Cycle of the Founding School of American Plant Ecology, 1895–1955 by Ronald C. Tobey. Isis, 73(4):584–586, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232169.

Ghiselin:1982:BRB

[768] Michael T. Ghiselin. Book review: History in the Service of Systematics: Papers from the Conference to Celebrate the Centenary of the British Museum (Natural History) 13–16 April, 1981 by Alwyne Wheeler; James H. Price. Isis, 73(4):586, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232170.

Smith:1982:BRBc

[769] Ellen Smith. Book review: Jenner’s Smallpox Vaccine: The Riddle of Vaccinia Virus and Its Origin by Derrick Baxby. Isis, 73(4):586, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232171.

Horn:1982:BRB

[770] Margo Horn. Book review: The Germ of Laziness: Rockefeller Philanthropy and Public Health in the New South by John Ettling. Isis, 73(4): 587, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232172.

Lang:1982:BRB

[771] David J. Lang. Book review: Kuru: Early Letters and Field-Notes from the Collection of D. Carleton Gajdusek by Judith Farquhar; D. Carleton Gajdusek. Isis, 73(4):587–588, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232173.

Weiner:1982:BRB

[772] Dora B. Weiner. Book review: Genese de la psychiatrie: Les premiers ecrits de Philippe Pinel by Jacques Postel. Isis, 73(4):588–589, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232174.

Sinclair:1982:BRB

[773] Bruce Sinclair. Book review: Emulation and Invention by Brooke Hindle. Isis, 73(4):589–590, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232175.

Heitmann:1982:BRB

[774] John A. Heitmann. Book review: The Rise of the American Electrochemicals Industry, 1880–1910: Studies in the American Technological Environment by Martha Moore Trescott. Isis, 73(4):590–591, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232176.

Mueller:1982:BRB

[775] Ian Mueller. Book review: On the Movement and Progression of Animals by Aristotle; Michael of Ephesus; Anthony Preus. Isis, 73(4):591–592, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232177.

Bogaard:1982:BRB

[776] Paul A. Bogaard. Book review: L’avenement de la science physique: Essai sur la Physique d’Aristote by Lambros Couloubaritsis. Isis, 73 (4):592–593, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232178.

DeLacy:1982:BRB

[777] Phillip De Lacy. Book review: The Hippocratic Treatises “On Generation,” “On the Nature of the Child,” “Diseases IV”: A Commentary by Iain M. Lonie. Isis, 73(4):593–594, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232179.

Longrigg:1982:BRB

[778] James Longrigg. Book review: Theories of Weight in the Ancient World: Four Essays on Democritus, Plato and Aristotle. Volume I: Democritus, Weight and Size: An Exercise in the Reconstruction of Early Greek Philosophy by D. O’Brien. Isis, 73(4):594–595, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232180.

Clagett:1982:BRB

[779] Marshall Clagett. Book review: Studies in Medieval Science and Natural Philosophy by Edward Grant. Isis, 73(4):595–596, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232181.

Grant:1982:BRB

[780] Edward Grant. Book review: Les sources astronomiques (Textes, Tables, Instruments) by Emmanuel Poulle. Isis, 73(4):596–597, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232182.

Park:1982:BRB

[781] Katharine Park. Book review: Taddeo Alderotti and His Pupils: Two Generations of Italian Medical Learning by Nancy G. Siraisi. Isis, 73 (4):597–598, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232183.

Hutchinson:1982:BRB

[782] G. E. Hutchinson. Book review: Birds in Medieval Manuscripts by Brunsdon Yapp. Isis, 73(4):598, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232184.

Sivin:1982:BRBb

[783] N. Sivin. Book review: Celestial Lancets: A History and Rationale of Acupuncture and Moxa by Lu Gwei-Djen; Joseph Needham. Isis, 73 (4):598–599, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232185.

Pingree:1982:BRB

[784] David Pingree. Book review: Susruta Samhita: A Scientific Synopsis by Priyadaranjan Ray; Hirendranath Gupta; Mira Roy. Isis, 73(4):600, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232186.

Newman:1982:BRB

[785] William Newman. Book review: Kreatur und Kosmos: Internationale Beitrage zur Paracelsusforschung by Rosemarie Dilg-Frank. Isis, 73(4): 600–601, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232187.

Bedini:1982:BRB

[786] Silvio A. Bedini. Book review: Leonardo da Vinci: The Marvellous Works of Nature and Man by Martin Kemp. Isis, 73(4):601, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232188.

Wallace:1982:BRBb

[787] William A. Wallace. Book review: Studies in Renaissance Philosophy and Science by Charles B. Schmitt. Isis, 73(4):601–602, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232189.

Porter:1982:BRBb

[788] Charlotte M. Porter. Book review: The Life and Travels of John Bartram: From Lake Ontario to the River St. John by Edmund Berkeley; Dorothy Smith Berkeley. Isis, 73(4):602, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232190.

Palter:1982:BRB

[789] Robert Palter. Book review: Witch-Hunting, Magic and the New Philosophy: An Introduction to Debates of the Scientific Revolution, 1450– 1750 by Brian Easlea. Isis, 73(4):602–603, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232191.

Smeaton:1982:BRB

[790] W. A. Smeaton. Book review: Newton on the Continent by Henry Guerlac. Isis, 73(4):603–604, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232192.

Schweber:1982:BRBb

[791] Silvan S. Schweber. Book review: Natural Knowledge in Social Context: The Journals of Thomas Archer Hirst, FRS by William H. Brock; Roy M. MacLeod. Isis, 73(4):604–605, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232193.

Stanton:1982:BRB

[792] William Stanton. Book review: John M. Brooke: Naval Scientist and Educator by George M. Brooke, Jr. Isis, 73(4):605–606, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232194.

Goodstein:1982:BRB

[793] Judith R. Goodstein. Book review: Science and the Sons of Genius: Studies on Humphry Davy by Sophie Forgan. Isis, 73(4):606–607, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232195.

Young:1982:BRB

[794] Robert M. Young. Book review: Darwinian Impacts: An Introduction to the Darwinian Revolution by David R. Oldroyd. Isis, 73(4):607–608, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232196.

Schultz:1982:BRB

[795] Susan Schultz. Book review: Thomas Alva Edison: An American Myth by Wyn Wachhorst. Isis, 73(4):608–609, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232197.

Constant:1982:BRB

[796] Edward W. Constant. Book review: The High-Speed Frontier: Case Histories of Four NACA Programs, 1920–1950 by John V. Becker. Isis, 73(4):609–610, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232198.

Dick:1982:BRB

[797] Steven J. Dick. Book review: Stages to Saturn: A Technological History of the Apollo/Saturn Launch Vehicles by Roger E. Bilstein. Isis, 73(4): 610–611, December 1982. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232199.

Stuewer:1982:BRB

[798] Roger H. Stuewer. Book review: Im Schatten der Sensation: Leben und Wirken von Fritz Strassmann by Fritz Krafft. Isis, 73(4):611, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232200.

Anonymous:1982:BMe

[799] Anonymous. Back matter. Isis, 73(4):612–615, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232201.

Anonymous:1982:VI

[800] Anonymous. Volume information. Isis, 73(4):616–626, December 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232141.

Anonymous:1982:FMf

[801] Anonymous. Front matter. Isis, 73(5):1–18, ???? 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231516.

Neu:1982:OHS

[802] John Neu. One hundred seventh critical bibliography of the history of science and its cultural influences (to January 1982). Isis, 73(5):5–17+19– 198, ???? 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231517.

Anonymous:1982:BMf

[803] Anonymous. Back matter. Isis, 73(5):??, ???? 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231518.

Anonymous:1983:FMa

[804] Anonymous. Front matter. Isis, 74(1):1–5, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232276.

Wise:1983:IIP

[805] George Wise. Ionists in industry: Physical chemistry at general electric, 1900–1915. Isis, 74(1):6–21, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232277.

Freudenthal:1983:TMC

[806] Gad Freudenthal. Theory of matter and cosmology in William Gilbert’s de magnete. Isis, 74(1):22–37, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232278.

Lennox:1983:RBD

[807] James G. Lennox. Robert Boyle’s defense of teleological inference in experimental science. Isis, 74(1):38–52, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232279.

Knorr:1983:GBM

[808] Wilbur Knorr. The geometry of burning-mirrors in antiquity. Isis, 74(1): 53–73, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232280.

Guerlac:1983:CWD

[809] Henry Guerlac. Can we date Newton’s early optical experiments? Isis, 74(1):74–80, March 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232281.

Whitrow:1983:IHS

[810] Magda Whitrow. Interrelationships in the history of science: Bibliography as a guide to subjects for research. Isis, 74(1):81–85, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232282.

Cohen:1983:EWH

[811] I. Bernard Cohen. Eloges: Willy Hartner, 22 January 1905–16 May 1981. Isis, 74(1):86–87, March 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232283.

Morrell:1983:SFC

[812] J. B. Morrell. Susan Faye Cannon (Walter Faw Cannon) 15 October 1925–6 November 1981. Isis, 74(1):88, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232284.

Hahn:1983:BRB

[813] Roger Hahn. Book review: Science and Polity in France at the End of the Old Regime. Isis, 74(1):89–91, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232285.

Glick:1983:BRB

[814] Thomas F. Glick. Book review: The Origins of Academic Geography in the United States. Isis, 74(1):92–97, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232286.

Postnikov:1983:BRB

[815] A. V. Postnikov. Book review: Tooley’s Dictionary of Mapmakers. Isis, 74(1):97–101, March 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232287.

Leary:1983:BRB

[816] David E. Leary. Book review: Human Nature in American Thought: A History by Merle Curti. Isis, 74(1):102–103, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232288.

Kohler:1983:BRB

[817] Robert E. Kohler. Book review: Science in America: A Documentary History, 1900–1939 by Nathan Reingold; Ida H. Reingold. Isis, 74(1): 103–104, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232289.

Multhauf:1983:BRBa

[818] R. P. Multhauf. Book review: Catalog of the Sidney M. Edelstein Collection of the History of Chemistry, Dyeing, and Technology by Moshe Ron. Isis, 74(1):104–105, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232290.

Burnham:1983:BRB

[819] John C. Burnham. Book review: A Guide to Manuscript Collections in the History of Psychology and Related Areas by Michael Sokal; Patrice Rafail. Isis, 74(1):105, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232291.

Dobbs:1983:BRBa

[820] B. J. T. Dobbs. Book review: The Reenchantment of the World by Morris Berman. Isis, 74(1):105–106, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232292.

Hull:1983:BRB

[821] David L. Hull. Book review: Darwinism Defended: A Guide to the Evolution Controversies by Michael Ruse. Isis, 74(1):106–107, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232293.

Morman:1983:BRB

[822] Edward T. Morman. Book review: Ahead of Its Time: The Engineering Societies Library, 1913–80 by Ellis Mount. Isis, 74(1):107, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232294.

Hickel:1983:BRB

[823] Erika Hickel. Book review: Aus der Fruhzeit der chemischen Konstitutionsforschung: Die Tropanalkaloide Atropin und Cocain in Wissenschaft und Wirtschaft by Brigitte Hoppe. Isis, 74(1):107–108, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232295.

Larson:1983:BRBa

[824] James L. Larson. Book review: “En Fond Som Aldrig Kan Aldras”: Stiftelsen Lars Hiertas Minne. En alterblick 1878–1978 by Wilhelm Odelberg. Isis, 74(1):108–109, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232296.

Gray:1983:BRB

[825] J. J. Gray. Book review: Mathematical Perspectives: Essays on Mathematics and Its Historical Development by Joseph W. Dauben. Isis, 74 (1):109–110, March 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232297.

Pycior:1983:BRBa

[826] Helena Pycior. Book review: History of Binary and Other Nondecimal Numeration by Anton Glaser. Isis, 74(1):110, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232298.

Olesko:1983:BRBa

[827] Kathryn Olesko. Book review: Social History of Nineteenth-Century Mathematics by Herbert Mehrtens; Henk Bos; Ivo Schneider. Isis, 74 (1):110–111, March 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232299.

Tarbell:1983:BRB

[828] D. Stanley Tarbell. Book review: Carolina Chemists: Sketches from Chapel Hill by Maurice M. Bursey. Isis, 74(1):111–112, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232300.

Kopperl:1983:BRB

[829] Sheldon J. Kopperl. Book review: Inorganic Coordination Compounds by George B. Kauffman. Isis, 74(1):112–113, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232301.

Trenn:1983:BRBa

[830] Thaddeus J. Trenn. Book review: Die Geschichte der Radioaktivitat: Unter besonderer Berucksichtigung der Transurane by Cornelius Keller. Isis, 74(1):113–114, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232302.

Brooke:1983:BRB

[831] John Hedley Brooke. Book review: Jacob Berzelius: The Emergence of His Chemical System by Evan M. Melhado. Isis, 74(1):114, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232303.

Hetherington:1983:BRB

[832] Norriss S. Hetherington. Book review: Astronomy for the Enquiring Mind by Eric M. Rogers. Isis, 74(1):114–115, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232304.

Manning:1983:BRBa

[833] Thomas G. Manning. Book review: The Geological Sciences in the Antebellum South by James X. Corgan. Isis, 74(1):115, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232305.

Aldrich:1983:BRB

[834] Michele L. Aldrich. Book review: American Geological Literature, 1669 to 1850 by Robert M. Hazen; Margaret Hindle Hazen. Isis, 74(1):115– 116, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232306.

McGraw:1983:BRB

[835] Donald J. McGraw. Book review: The Fungus Fighters: Two Women Scientists and Their Discovery by Richard S. Baldwin. Isis, 74(1):116– 117, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232307.

Churchill:1983:BRB

[836] Frederick B. Churchill. Book review: Die Illustrationen Ernst Haeckels zur Abstammungs- und Entwicklunsgeschichte: Diskussion im Wissenschaftlichen und Nichtwissenschaftlichen Schrifttum by Reinhard Gursch. Isis, 74(1):117, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232308.

Kottler:1983:BRBa

[837] Malcom Kottler. Book review: Dobzhansky’s Genetics of Natural Populations I-XLIII by R. C. Lewontin; John A. Moore; William B. Provine; Bruce Wallace. Isis, 74(1):117–118, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232309.

Kimmelman:1983:BRB

[838] Barbara A. Kimmelman. Book review: The Maine Agricultural Experiment Station by David C. Smith. Isis, 74(1):118, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232310.

Stocking:1983:BRB

[839] George W. Stocking. Book review: Culture in Comparative and Evolutionary Perspective: E. B. Tylor and the Making of “Primitive Culture” by Joan Leopold. Isis, 74(1):119–120, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232311.

Mark:1983:BRB

[840] Joan Mark. Book review: Beyond Separate Spheres: Intellectual Roots of Modern Feminism by Rosalind Rosenberg. Isis, 74(1):120, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232312.

Tarr:1983:BRB

[841] Joel A. Tarr. Book review: From Streetcar to Superhighway: American City Planners and Urban Transportation, 1900–1940. (Technology and Urban Growth) by Mark S. Foster. Isis, 74(1):120–121, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232313.

Jensen:1983:BRB

[842] J. Vernon Jensen. Book review: Ethics in an Age of Pervasive Technology by Melvin Kranzberg. Isis, 74(1):121–122, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232314.

Cardwell:1983:BRB

[843] Donald Cardwell. Book review: The Selected Papers of Boulton and Watt. Volume I: The Engine Partnership, 1775–1825 by Jennifer Tann. Isis, 74(1):122–123, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232315.

Shinn:1983:BRB

[844] Terry Shinn. Book review: The Making of Technological Man: The Social Origins of French Engineering Education by John Hubbel Weiss. Isis, 74(1):123–124, March 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232316.

Copenhaver:1983:BRB

[845] Brian P. Copenhaver. Book review: Les alchimistes grecs. Volume I, Papyrus de Leyde, papyrus de Stockholm, fragments de recettes by Robert Halleux. Isis, 74(1):124, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232317.

Neuenschwander:1983:BRB

[846] Erwin Neuenschwander. Book review: Philosophy of Mathematics and Deductive Structure in Euclid’s Elements by Ian Mueller. Isis, 74(1):124– 126, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232318.

Longrigg:1983:BRB

[847] James Longrigg. Book review: Empedocles: The Extant Fragments by M. R. Wright. Isis, 74(1):126, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232319.

Berggren:1983:BRBa

[848] J. L. Berggren. Book review: Maxime Planude: Le grand calcun selon les Indiens by Andre Allard. Isis, 74(1):126–127, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232320.

Siraisi:1983:BRB

[849] Nancy G. Siraisi. Book review: Arnaldi de Villanova opera medica omnia. Vol. XVI: Translatio libri Galieni de rigore et tremore et iectigatione et spasmo by Arnald of Villanova; Michael McVaugh. Isis, 74(1):127–128, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232321.

Sylla:1983:BRB

[850] Edith D. Sylla. Book review: De numeris datis by Jordanus de Nemore; Barnabas Bernard Hughes. Isis, 74(1):128–129, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232322.

Thomson:1983:BRBa

[851] Ron B. Thomson. Book review: Infinity and Continuity in Ancient and Medieval Thought by Norman Kretzmann. Isis, 74(1):129–130, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232323.

Grant:1983:BRB

[852] Edward Grant. Book review: On the Threshold of Exact Science: Selected Writings of Anneliese Maier on Late Medieval Natural Philosophy by Anneliese Maier; Steven D. Sargent. Isis, 74(1):130–131, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232324.

Anonymous:1983:CHL

[853] Anonymous. Corrigendum: Histoire litteraire de la France. Volume 41: Suite du quatorzième siècle. Isis, 74(1):131, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232325. See [714].

Albury:1983:BRB

[854] W. R. Albury. Book review: Les monades by Etienne Bonnot de Condillac; Laurence L. Bongie. Isis, 74(1):131–132, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232326.

Schaffer:1983:BRBa

[855] Simon Schaffer. Book review: Universal Natural History and Theory of the Heavens by Immanuel Kant; Stanley L. Jaki. Isis, 74(1):132– 133, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232327.

Shea:1983:BRBa

[856] William R. Shea. Book review: Otto von Guericke by Fritz Krafft. Isis, 74(1):133, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232328.

Appel:1983:BRB

[857] Toby A. Appel. Book review: From Natural History to the History of Nature: Readings from Buffon and His Critics by John Lyon; Phillip Sloan. Isis, 74(1):133–134, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232329.

Kiefer:1983:BRB

[858] Joseph H. Kiefer. Book review: The Letters and Papers of Sir John Hill (1714–1775) by G. S. Rousseau. Isis, 74(1):134–135, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232330.

Ewan:1983:BRB

[859] Joseph Ewan. Book review: The Poison Tree: Selected Writings of Rumphius on the Natural History of the Indies by Georg Eberhard Rumpf; E. M. Beekman. Isis, 74(1):135, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232331.

Chambers:1983:BRB

[860] David Wade Chambers. Book review: Octavio F. Mossotti: En la amanecer de la ciencia argentina by Maximo Baron. Isis, 74(1):135– 136, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232332.

Burke:1983:BRB

[861] Michael E. Burke. Book review: El cientifico español ante su historia: La ciencia en España entre 1750–1850 by Santiago Garma. Isis, 74(1):136– 137, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232333.

Lenoir:1983:BRBa

[862] Timothy Lenoir. Book review: Schelling: Seine Bedeutung für eine Philosophie der Natur und der Geschichte by Ludwig Hasler. Isis, 74 (1):137–138, March 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232334.

Sloan:1983:BRB

[863] Phillip R. Sloan. Book review: The Red Notebook of Charles Darwin by Sandra Herbert. Isis, 74(1):138–139, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232335.

Dunbar:1983:BRB

[864] G. S. Dunbar. Book review: Carl Ritter–Geltung und Deutung by Karl Lenz; Carl Ritter. Isis, 74(1):139–140, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232336.

Middleton:1983:BRBa

[865] W. E. Knowles Middleton. Book review: New Eye for the Navy: The Origin of Radar at the Naval Research Laboratory by David Kite Allison. Isis, 74(1):140, March 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232337.

Sterling:1983:BRBa

[866] Keir B. Sterling. Book review: The Zest for Life, or Waldo Had a Pretty Good Run: The Life of Waldo LaSalle Schmitt by Richard E. Blackwelder. Isis, 74(1):140–141, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232338.

Pyenson:1983:BRB

[867] Lewis Pyenson. Book review: Emmy Noether, 1882–1935 by Auguste Dick; Heidi I. Blocher. Isis, 74(1):141, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232339.

Kuchment:1983:BRBa

[868] Mark Kuchment. Book review: Landau: A Great Physicist and Teacher by Anna Livanova; J. B. Sykes. Isis, 74(1):141–143, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232340.

Reingold:1983:BRB

[869] Nathan Reingold. Book review: The Rise of Robert Millikan: Portrait of a Life in American Science by Robert H. Kargon. Isis, 74(1):143– 144, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232341.

Rabkin:1983:BRB

[870] Yakov M. Rabkin. Book review: Polywater by Felix Franks. Isis, 74(1): 144–146, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232342.

Weingart:1983:BRB

[871] Peter Weingart. Book review: Policymaking in a Nuclear Program: The Case of the West German Breeder Reactor by Otto Keck. Isis, 74(1):146– 147, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232343.

Badash:1983:BRBa

[872] Lawrence Badash. Book review: Kennedy, Khrushchev, and the Test Ban by Glenn T. Seaborg; Benjamin S. Loeb. Isis, 74(1):147–148, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232344.

Stern:1983:BRB

[873] Nancy Stern. Book review: History of Programming Languages by Richard L. Wexelblat. Isis, 74(1):148, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232345.

Anonymous:1983:BMa

[874] Anonymous. Back matter. Isis, 74(1):149–152, March 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232346.

Anonymous:1983:FMb

[875] Anonymous. Front matter. Isis, 74(2):153–157, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233100.

Roe:1983:JTN

[876] Shirley A. Roe. John Turberville Needham and the generation of living organisms. Isis, 74(2):158–184, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233101.

Jenks:1983:AMA

[877] Stuart Jenks. Astrometeorology in the middle ages. Isis, 74(2):185– 210, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233102.

Pycior:1983:AMA

[878] Helena M. Pycior. Augustus De Morgan’s algebraic work: The three stages. Isis, 74(2):211–226, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233103.

Fancher:1983:BOF

[879] Raymond E. Fancher. Biographical origins of Francis Galton’s psychology. Isis, 74(2):227–233, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233104.

Osborne:1983:ADC

[880] Catherine Osborne. Archimedes on the dimensions of the Cosmos. Isis, 74(2):234–242, June 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233105.

Hannaway:1983:AMH

[881] Owen Hannaway, Robert Kargon, and Audrey B. Davis. Annual meeting of the History of Science Society 28–31 October 1982. Isis, 74(2):243– 248, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233106.

Shapin:1983:SHS

[882] Steven Shapin. Symposium on history of science today: Tradition and change Uppsala, Sweden, 18–19 November 1982. Isis, 74(2):248–249, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233107.

Edge:1983:BRB

[883] David Edge. Book review: The Social Basis of Scientific Discoveries. Isis, 74(2):250–256, June 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233108.

Multhauf:1983:BRBb

[884] R. P. Multhauf. Book review: Reference Books for the Historian of Science: A Handlist by S. A. Jayawardene. Isis, 74(2):257, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233109.

Montgomery:1983:BRBa

[885] William Montgomery. Book review: Concise Dictionary of Scientific Biography by James F. Maurer; A Biographical Encyclopedia of Scientists by John Daintith; Sarah Mitchell; Elizabeth Tootil. Isis, 74(2):257– 258, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233110.

Bhaskar:1983:BRB

[886] Roy Bhaskar. Book review: Scientific Progress: A Study concerning the Nature of the Relation between Successive Scientific Theories by Craig Dilworth; Scientific Revolutions by Ian Hacking. Isis, 74(2):258–259, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233111.

Burian:1983:BRB

[887] Richard M. Burian. Book review: Abusing Science: The Case against Creationism by Philip Kitcher. Isis, 74(2):259–260, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233112.

Baird:1983:BRB

[888] Davis Baird. Book review: What? Where? When? Why?: Essays on Induction, Space and Time, Explanation by Robert McLaughlin. Isis, 74(2):260–261, June 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233113.

Huff:1983:BRB

[889] Toby E. Huff. Book review: Karl Popper by Anthony O’Hear. Isis, 74(2): 261–262, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233114.

Schroeder-Gudehus:1983:BRB

[890] Brigitte Schroeder-Gudehus. Book review: Knowledge and Power in a Global Society by William M. Evan. Isis, 74(2):262, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233115.

Vucinich:1983:BRBa

[891] Alexander Vucinich. Book review: Geschichte der Akademie der Wissenschaften der UdSSR by G. D. Komkov; B. V. Levsin; L. K. Semenov; Gerhard Basler; Conrad Grau; Dieter Muhle. Isis, 74(2):262–263, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233116.

Lenoir:1983:BRBb

[892] Timothy Lenoir. Book review: On the Roads to Modernity: Conscience, Science, and Civilizations by Benjamin Nelson; Toby E. Huff. Isis, 74(2): 263–264, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233117.

Whorton:1983:BRB

[893] James Whorton. Book review: Fire in America: A Cultural History of Wildland and Rural Fire by Stephen J. Pyne. Isis, 74(2):264, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233118.

Rosner:1983:BRB

[894] David Rosner. Book review: Madhouses, Mad-Doctors, and Madmen: The Social History of Psychiatry in the Victorian Era by Andrew Scull. Isis, 74(2):264–265, June 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233119.

Paul:1983:BRBa

[895] H. W. Paul. Book review: Darwin & Co. by Pierre Thuillier. Isis, 74(2): 265–266, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233120.

Cowan:1983:BRB

[896] Ruth Schwartz Cowan. Book review: Biology, Medicine, and Society, 1840–1940 by Charles Webster. Isis, 74(2):266–267, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233121.

Hutton:1983:BRB

[897] Sarah Hutton. Book review: Francis Godwins “The Man in the Moone”: Die Entdeckung des Romans als Medium der Auseinandersetzung mit Zeitproblemen by Anke Janssen. Isis, 74(2):267, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233122.

Theerman:1983:BRB

[898] Paul Theerman. Book review: American Literature and the Universe of Force by Ronald E. Martin. Isis, 74(2):267–268, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233123.

Crowe:1983:BRB

[899] Michael J. Crowe. Book review: Plurality of Worlds: The Origins of Extraterrestrial Life Debate from Democritus to Kant by Steven J. Dick. Isis, 74(2):268–270, June 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233124.

Scholz:1983:BRB

[900] E. Scholz. Book review: E. B. Christoffel: The Influence of His Work on Mathematics and the Physical Sciences by P. L. Butzer; F. Feher. Isis, 74(2):270–271, June 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233125.

Wilson:1983:BRBa

[901] Curtis A. Wilson. Book review: Stellar Astronomy: Historical Studies by Michael Hoskin. Isis, 74(2):271–272, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233126.

Schaffer:1983:BRBb

[902] Simon Schaffer. Book review: Uranus and the Outer Planets by Garry Hunt. Isis, 74(2):272, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233127.

Bromberg:1983:BRB

[903] Joan Bromberg. Book review: Transmutation: Natural and Artificial by T. J. Trenn. Isis, 74(2):272–273, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233128.

Kish:1983:BRBa

[904] George Kish. Book review: Early Thematic Mapping in the History of Cartography by Arthur H. Robinson. Isis, 74(2):273, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233129.

Servos:1983:BRB

[905] John W. Servos. Book review: From Medical Chemistry to Biochemistry: The Making of a Biomedical Discipline by Robert E. Kohler. Isis, 74(2): 273–275, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233130.

Farber:1983:BRBa

[906] Paul Lawrence Farber. Book review: Lamarck et son temps–Lamarck et notre temps: Colloque international dans le cadre du Centre d’Etudes et de Recherches interdisciplines de Chantilly. Isis, 74(2):275, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233131.

Reeds:1983:BRB

[907] Karen Reeds. Book review: History of Botanical Science: An Account of the Development of Botany from Ancient Times to the Present Day by A. G. Morton; The Garden of Eden: The Botanic Garden and the Re-Creation of Paradise by John Prest. Isis, 74(2):275–277, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233132.

Liebersohn:1983:BRB

[908] Harry Liebersohn. Book review: Geschichte der Soziologie: Studien zur kognitiven, sozialen, und historischen Identitat einer Disziplin by Wolf Lepenies. Isis, 74(2):277–278, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233133.

Vogel:1983:BRB

[909] Morris J. Vogel. Book review: Medicine in Colonial Massachusetts, 1620–1820: A Conference Held 25 and 26 May 1978 by the Colonial Society of Massachusetts by Philip Cash; Eric H. Christianson; J. Worth Estes. Isis, 74(2):278–279, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233134.

Larson:1983:BRBb

[910] James Larson. Book review: Ars Chirurgica: Kirurgi och medicinsk teori vid Lunds Universitet under 1700-talet by Jan Larsson. Isis, 74(2): 279, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233135.

Decker:1983:BRB

[911] Hannah S. Decker. Book review: A Mobius Strip: Fin-de-Siècle Neuropsychiatry and Paul Mobius by Francis Schiller. Isis, 74(2):279–280, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233136.

Parascandola:1983:BRB

[912] John Parascandola. Book review: The Enchanted Ring: The Untold Story of Penicillin by John C. Sheehan. Isis, 74(2):280–281, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233137.

Pursell:1983:BRB

[913] Carroll Pursell. Book review: The Teaching of History of Technology in USA — a Critical Survey in 1978 by Svante Lindqvist. Isis, 74(2): 281, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233138.

Fries:1983:BRB

[914] Russell I. Fries. Book review: Yankee Enterprise: The Rise of the American System of Manufactures by Otto Mayr; Robert C. Post. Isis, 74(2): 281–283, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233139.

Finn:1983:BRB

[915] Bernard S. Finn. Book review: Technology in America: A History of Individuals and Ideas by Carroll W. Pursell. Isis, 74(2):283, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233140.

Unguru:1983:BRB

[916] Sabetai Unguru. Book review: Bibliographia Euclideana: Die Geisteslinien der Tradition in den Editionen der “Elemente” (ΣT OIXEIA des Euklid) (um 365-300). Handschriften–Inkunabeln–Fruhdrucke (16. Jahrhundert). Textkritische Editionen des 17.-20. Jahrhunderts. Editionen der Opera minora (16.-20. Jahrhundert) by Max Steck. Edited by Menso Folkerts. Isis, 74(2):283–284, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233141.

Pingree:1983:BRB

[917] David Pingree. Book review: Sisyadhivrddhida Tantra of Lalla, with the Commentary of Mallikarjuna Suri by Bina Chatterjee. Isis, 74(2):284– 285, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233142.

Jayawardene:1983:BRB

[918] S. A. Jayawardene. Book review: Practical Mathematics in the Italian Renaissance: A Catalog of Italian Abbacus Manuscripts and Printed Books to 1600 by Warren Van Egmond. Isis, 74(2):285–286, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233143.

Shapin:1983:BRB

[919] Steven Shapin. Book review: The Origins and Nature of the Scottish Enlightenment by R. H. Campbell; Andrew S. Skinner. Isis, 74(2):286, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233144.

Staum:1983:BRB

[920] Martin S. Staum. Book review: La Physiologie des lumières: Empirisme, modeles et theories by Francois Duchesneau. Isis, 74(2):287–288, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233145.

VanHelden:1983:BRB

[921] Albert Van Helden. Book review: Huygens et la France. Isis, 74(2): 288, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233146.

Golinski:1983:BRB

[922] Jan V. Golinski. Book review: Memoir on Heat; Read to the Royal Academy of Sciences June 28, 1783, by Messrs. Lavoisier & De La Place of the Same Academy by Antoine Laurent Lavoisier; Pierre Simon; Marquis de Laplace; Henry Guerlac. Isis, 74(2):288–289, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233147.

Sterling:1983:BRBb

[923] Keir B. Sterling. Book review: Elliott Coues: Naturalist and Frontier Historian by Paul Russell Cutright; Michael J. Brodhead. Isis, 74(2): 290–291, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233148.

Kottler:1983:BRBb

[924] Malcolm Jay Kottler. Book review: Alfred Russel Wallace by Martin Fichman. Isis, 74(2):291, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233149.

Porter:1983:BRBa

[925] Charlotte M. Porter. Book review: Rafinesque: A Sketch of His Life with Bibliography by Thomas Jefferson Fitzpatrick; Charles Boewe. Isis, 74(2):291–292, June 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233150.

Richards:1983:BRA

[926] Joan L. Richards. Book review: Anthony Hyman: Charles Babbage: Pioneer of the Computer. Isis, 74(2):292, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233151.

Ruse:1983:BRB

[927] Michael Ruse. Book review: The Development of Darwin’s Theory: Natural History, Natural Theology, and Natural Selection, 1838–1859 by Dov Ospovat. Isis, 74(2):292–293, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233152.

Kuklick:1983:BRB

[928] Bruce Kuklick. Book review: Writings of Charles S. Peirce: A Chronological Edition. Volume I: 1857–1866 by Charles S. Peirce; Max H. Fisch. Isis, 74(2):293–294, June 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233153.

Carlson:1983:BRB

[929] Elof Axel Carlson. Book review: The Material Basis of Evolution by Richard Goldschmidt; Richard Goldschmidt: Controversial Geneticist and Creative Biologist by Leonie K. Piternick. Isis, 74(2):294–296, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233154.

Constant:1983:BRB

[930] Edward W. Constant. Book review: Bringing Aerodynamics to America by Paul A. Hanle. Isis, 74(2):296, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233155.

Holmes:1983:BRB

[931] Frederic L. Holmes. Book review: Reminiscences and Reflections by Hans Krebs. Isis, 74(2):296–297, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233156.

Priesner:1983:BRB

[932] Claus Priesner. Book review: The Chain Straighteners: Fruitful Innovation: The Discovery of Linear and Stereoregular Synthetic Polymers by Frank G. McMillan; Polymer Science Overview: A Tribute to Herman F. Mark by G. Allan Stahl. Isis, 74(2):297–298, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233157.

vonMeyenn:1983:BRB

[933] Karl von Meyenn. Book review: Les conceptions quantiques de 1911 a 1937 by Girolamo Ramunni. Isis, 74(2):298–300, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233158.

Friedel:1983:BRB

[934] Robert Friedel. Book review: The Great Television Race: A History of the American Television Industry 1925–1941 by Joseph H. Udelson. Isis, 74(2):300, June 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233159.

Anonymous:1983:BMb

[935] Anonymous. Back matter. Isis, 74(2):301–304, June 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233160.

Anonymous:1983:FMc

[936] Anonymous. Front matter. Isis, 74(3):305–309, September 1983. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232591.

Jacyna:1983:ITT

[937] L. S. Jacyna. Immanence or transcendence: Theories of life and organization in Britain, 1790–1835. Isis, 74(3):310–329, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232592.

Goldstein:1983:NVE

[938] Bernard R. Goldstein and Alan C. Bowen. A new view of early Greek astronomy. Isis, 74(3):330–340, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232593.

Hunt:1983:PVT

[939] Bruce J. Hunt. “Practice vs. theory”: The British electrical debate, 1888–1891. Isis, 74(3):341–355, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232594.

Rosenberg:1983:SAS

[940] Charles Rosenberg. Science in American society: A generation of historical debate. Isis, 74(3):356–367, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232595.

Abraham:1983:MMT

[941] Gary A. Abraham. Misunderstanding the Merton thesis: A boundary dispute between history and sociology. Isis, 74(3):368–387, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232596.

Saliba:1983:ONT

[942] George Saliba. An observational notebook of a Thirteenth-Century astronomer. Isis, 74(3):388–401, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232597.

Anonymous:1983:NR

[943] Anonymous. News of record. Isis, 74(3):402–403, September 1983. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232598.

Watanabe:1983:EBI

[944] Masao Watanabe. Eloge: Bun-Ichi Tamamushi, 18 October 1898–26 July 1982. Isis, 74(3):403–404, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232599.

Roger:1983:BRB

[945] Jacques Roger and Michael T. Ghiselin. Book review: The Growth of Biological Thought: Diversity, Evolution, and Inheritance. Isis, 74(3): 405–413, September 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232600.

Allen:1983:BRB

[946] Garland E. Allen. Book review: Genes, Radiation, and Society: The Life and Work of H, by Elof Axel Carlson. J. Muller. Isis, 74(3):413–416, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232601.

Ovitt:1983:BRB

[947] George Ovitt, Jr. Book review: Dawn of Modern Science: From the Arabs to Leonardo da Vinci by Thomas Goldstein. Isis, 74(3):417–418, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232602.

Hoskin:1983:BRB

[948] Michael Hoskin. Book review: The History of Modern Astronomy and Astrophysics: A Selected, Annotated Bibliography by David H. DeVorkin. Isis, 74(3):418, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232603.

Trenn:1983:BRBb

[949] Thaddeus J. Trenn. Book review: Literature on the History of Physics in the 20th Century by J. L. Heilbron; Bruce R. Wheaton; J. G. May; Robin Rider; David Robinson. Isis, 74(3):418–419, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232604.

Elliott:1983:BRB

[950] Clark A. Elliott. Book review: Pure and Applied Science Books, 1876– 1980. Isis, 74(3):419–420, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232605.

Porter:1983:BRBb

[951] Charlotte M. Porter. Book review: Biographical Dictionary of Rocky Mountain Naturalists: A Guide to the Writings and Collections of Botanists, Zoologists, Geologists, Artists, and Photographers, 1682–1932 by Joseph Ewan; Nesta Dunn Ewan. Isis, 74(3):420, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232606.

Edsall:1983:BRB

[952] John T. Edsall. Book review: A Bio-Bibliography for the History of the Biochemical Sciences since 1800 by Joseph S. Fruton. Isis, 74(3):421, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232607.

Airaksinen:1983:BRB

[953] Timo Airaksinen. Book review: Progress and Its Discontents by Gabriel Almond; Marvin Chodorow; Roy H. Pearce. Isis, 74(3):421–422, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232608.

Richards:1983:BRBc

[954] Robert J. Richards. Book review: Is Science Sexist? And Other Problems in the Biomedical Sciences by Michael Ruse. Isis, 74(3):422–423, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232609.

Rynasiewicz:1983:BRB

[955] Robert Rynasiewicz. Book review: Ideas of Matter: From Ancient Times to Bohr and Einstein by Mendel Sachs. Isis, 74(3):423–424, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232610.

Layton:1983:BRB

[956] David Layton. Book review: From Armstrong to Nuffield: Studies in Twentieth–Century Science Education in England and Wales by E. W. Jenkins. Isis, 74(3):424–425, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232611.

Hodge:1983:BRB

[957] M. J. S. Hodge. Book review: Victorian Science and Victorian Values: Literary Perspectives by James Paradis; Thomas Postlewait. Isis, 74(3): 425–427, September 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232612.

Daston:1983:BRB

[958] Lorraine J. Daston. Book review: Penser les mathematiques: Seminaire de philosophie et mathematiques de l’École normale superieure by Francois Guenard; Gilbert Lelievre. Isis, 74(3):427–428, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232613.

Pycior:1983:BRBb

[959] Helena Pycior. Book review: The Boole-De Morgan Correspondence, 1842–1864 by G. C. Smith. Isis, 74(3):428–429, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232614.

Shea:1983:BRBb

[960] William R. Shea. Book review: La Relativita da Faraday a Einstein by Enrico Bellone. Isis, 74(3):429, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232615.

Hackmann:1983:BRB

[961] Willem D. Hackmann. Book review: Elements of Early Modern Physics by J. L. Heilbron. Isis, 74(3):429–430, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232616.

Mulholland:1983:BRB

[962] James Mulholland. Book review: A History of Platinum and Its Allied Metals by Donald McDonald; Leslie B. Hunt. Isis, 74(3):430–431, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232617.

Wilson:1983:BRBb

[963] Curtis A. Wilson. Book review: The 1769 Transit of Venus: The Baja California Observations of Jean-Baptiste Chappe d’Auteroche, Vicente de Doz, and Joaquin Velazquez Cardenas de Leon by Doyce B. Nunis; James Donahue; Maynard J. Geiger; Iris Wilson Engstrand. Isis, 74(3): 431–432, September 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232618.

Ryan:1983:BRB

[964] Michael T. Ryan. Book review: The Great Map of Mankind: Perceptions of New Worlds in the Age of Enlightenment by P. J. Marshall; Glyndwr Williams. Isis, 74(3):432–433, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232619.

Brieger:1983:BRB

[965] Gert H. Brieger. Book review: Bad Blood: The Tuskegee Syphilis Experiment by James H. Jones. Isis, 74(3):433–434, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232620.

Howell:1983:BRB

[966] Joel Howell. Book review: Medical Thinking: A Historical Preface by Lester S. King. Isis, 74(3):434, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232621.

Torrance:1983:BRB

[967] George M. Torrance. Book review: Medicine in Canadian Society: Historical Perspectives by S. E. D. Shortt. Isis, 74(3):435, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232622.

Kranakis:1983:BRB

[968] Eda Fowlks Kranakis. Book review: A History of Theory of Structures in the Nineteenth Century by T. M. Charlton. Isis, 74(3):435–436, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232623.

Cheape:1983:BRB

[969] Charles W. Cheape. Book review: The Port of New York: A History of the Rail and Terminal System from the Grand Central Electrification to the Present by Carl W. Condit. Isis, 74(3):436–437, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232624.

Hughes:1983:BRB

[970] Thomas P. Hughes. Book review: Bottled Energy: Electrical Engineering and the Evolution of Chemical Energy Storage by Richard H. Schallenberg. Isis, 74(3):437–438, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232625.

Hall:1983:BRB

[971] A. Rupert Hall. Book review: The Social Context of Innovation: Bureaucrats, Families, and Heroes in the Early Industrial Revolution, as Foreseen in Bacon’s “New Atlantis.” by Anthony F. C. Wallace. Isis, 74(3):438, September 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232626.

Mauskopf:1983:BRB

[972] Seymour H. Mauskopf. Book review: Science and the Paranormal: Probing the Existence of the Supernatural by George O. Abell; Barry Singer. Isis, 74(3):438–439, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232627.

Pinch:1983:BRB

[973] Trevor Pinch. Book review: Psychical Research: A Guide to Its History, Principles, and Practices by I. Grattan-Guinness. Isis, 74(3):439, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232628.

Lennox:1983:BRB

[974] James G. Lennox. Book review: Aristotle on the Generation of Animals: A Philosophical Study by Johannes Morsink. Isis, 74(3):440–441, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232629.

Berggren:1983:BRBb

[975] J. L. Berggren. Book review: Enseignment et recherche mathematiques dans le Maghreb des XIIIe-XIVe siècles (Etude partielle) by A. Djebbar. Isis, 74(3):441, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232630.

Savage-Smith:1983:BRB

[976] Emilie Savage-Smith. Book review: The Elements of Vision: The Micro-Cosmology of Galenic Visual Theory according to Hunayn ibn Ishaq by Bruce Stansfield Eastwood. Isis, 74(3):441–442, September 1983. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232631.

McCluskey:1983:BRBa

[977] Stephen C. McCluskey. Book review: The Paris Codex: Decoding an Astronomical Ephemeris by Gregory M. Severin. Isis, 74(3):442–443, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232632.

Lindberg:1983:BRB

[978] David C. Lindberg. Book review: The Philosophy of Robert Grosseteste by James McEvoy. Isis, 74(3):443, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232633.

Livesay:1983:BRB

[979] Steven J. Livesay. Book review: Une vision du monde a la fin du XI-IIe siècle: Commentaire du Dialogue de Placides et Timeo by Claude Thomasset. Isis, 74(3):444, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232634.

Cook:1983:BRB

[980] Harold J. Cook. Book review: Obstetrics and Gynaecology in Tudor and Stuart England by Audrey Eccles. Isis, 74(3):444–445, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232635.

Ackerknecht:1983:BRB

[981] Erwin H. Ackerknecht. Book review: On Monsters and Marvels by Ambroise Pare; Janis L. Pallister. Isis, 74(3):445, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232636.

Gaukroger:1983:BRB

[982] Stephen Gaukroger. Book review: Descartes’ Philosophy of Science by Desmond M. Clarke. Isis, 74(3):445–446, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232637.

Osler:1983:BRB

[983] Margaret J. Osler. Book review: The Newtonian Revolution: With Illustrations of the Transformation of Scientific Ideas by I. Bernard Cohen. Isis, 74(3):446–447, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232638.

Larson:1983:BRBc

[984] James Larson. Book review: Nemesis divina by Carl von Linne; Wolf Lepenies; Lars Gustafsson; Ruprecht Volz. Isis, 74(3):447–448, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232639.

Moran:1983:BRB

[985] Bruce T. Moran. Book review: Wissenschaftsgeschichte um Wilhelm Schickard: Vorträge bei dem Symposion der Universitat Tübingen im 500 Jahr ihres Bestehens am 24. und 25. Juni 1977 by Friedrich Seck. Isis, 74(3):448–449, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232640.

Goetzmann:1983:BRB

[986] William H. Goetzmann. Book review: James Graham Cooper: Pioneer Western Naturalist by Eugene Coan. Isis, 74(3):449–450, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232641.

Lenoir:1983:BRBc

[987] Timothy Lenoir. Book review: Two Great Scientists of the Nineteenth Century: Correspondence of Emil Du Bois-reymond and Carl Ludwig by Paul F. Cranefield; Sabie Lichter Ayed. Isis, 74(3):450–451, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232642.

Greene:1983:BRB

[988] Mott T. Greene. Book review: Alfred Wegener by Hans-Gunther Korber. Isis, 74(3):451, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232643.

Melhado:1983:BRB

[989] Evan M. Melhado. Book review: Electrons and Valence: Development of the Theory, 1900–1925 by Anthony N. Stranges. Isis, 74(3):451–452, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232644.

Kuchment:1983:BRBb

[990] Mark Kuchment. Book review: Refusenik: Trapped in the Soviet Union by Mark Ya. Azbel; Grace Pierce Forbes. Isis, 74(3):452–454, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232645.

Wise:1983:BRB

[991] George Wise. Book review: Fusion: Science, Politics, and the Invention of a New Energy Source by Joan Lisa Bromberg. Isis, 74(3):454–455, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232646.

Anonymous:1983:BMc

[992] Anonymous. Back matter. Isis, 74(3):456–458, September 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232647.

Anonymous:1983:FMd

[993] Anonymous. Front matter. Isis, 74(4):459–463, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232206.

Frangsmyr:1983:GEH

[994] Tore Frangsmyr. Guest editorial: History of science in Sweden. Isis, 74 (4):464–468, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232207.

Beller:1983:MTB

[995] Mara Beller. Matrix theory before schrödinger: Philosophy, problems, consequences. Isis, 74(4):469–491, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232208.

Williams:1983:WWA

[996] L. Pearce Williams. What were Ampere’s earliest discoveries in electrodynamics? Isis, 74(4):492–508, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232209.

Lindberg:1983:SEC

[997] David C. Lindberg. Science and the early Christian Church. Isis, 74 (4):509–530, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232210.

King:1983:AM

[998] David A. King. The astronomy of the Mamluks. Isis, 74(4):531–555, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232211.

Fowler:1983:ERO

[999] David H. Fowler and Dennis Rawlins. Eratosthenes’ ratio for the obliquity of the ecliptic. Isis, 74(4):556–562, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232212.

Feingold:1983:NCC

[1000] Mordechai Feingold. “Novita Celesti e Crisi del Sapere” International Conference of Galilean Studies, Pisa, Padua, Venice, Florence, 18–26 March 1983. Isis, 74(4):563–564, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232213.

VanHelden:1983:ESB

[1001] Albert Van Helden. Eloge: Salomon Bochner, 20 August 1899–2 May 1982. Isis, 74(4):565, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232214.

Ringer:1983:BRB

[1002] Fritz K. Ringer. Book review: The Emergence of Modern Universities in France, 1863–1914. Isis, 74(4):566–568, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232215.

Nye:1983:BRB

[1003] Robert A. Nye. Book review: Distorting Mirrors: Visions of the Crowd in Late Nineteenth-Century France. Isis, 74(4):568–573, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232216.

Gelfand:1983:BRB

[1004] Toby Gelfand. Book review: Death Is a Social Disease: Public Health and Political Economy in Early Industrial France. Isis, 74(4):573–576, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232217.

Paul:1983:BRBb

[1005] Diane B. Paul. Book review: Nature, Human Nature, and Society: Marx, Darwin, Biology, and the Human Sciences by Paul Heyer. Isis, 74(4): 577, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232218.

McMullin:1983:BRB

[1006] Ernan McMullin. Book review: Science, Revolution, and Discontinuity by John Krige. Isis, 74(4):577–579, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232219.

Stroup:1983:BRB

[1007] Alice Stroup. Book review: Medicine and Its Technology: An Introduction to the History of Medical Instrumentation by Audrey B. Davis. Isis, 74(4):579–580, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232220.

Turner:1983:BRB

[1008] R. Steven Turner. Book review: Bildungspolitik in Preussen zur Zeit des Kaiserreichs by Peter Baumgart. Isis, 74(4):580–581, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232221.

Middleton:1983:BRBb

[1009] W. E. Knowles Middleton. Book review: Macchine, scuola, industria: Dal mestiere alla professionalita operaia by Comune di Bologna. Isis, 74(4):581, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232222.

Manning:1983:BRBb

[1010] Thomas G. Manning. Book review: The Geological Society of America: Life History of a Learned Society by Edwin B. Eckel. Isis, 74(4):581–582, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232223.

Irgolic:1983:BRB

[1011] Kurt J. Irgolic. Book review: Technik, Ingenieure und Gesellschaft: Geschichte des Vereins Deutscher Ingenieure 1856–1981 by Karl-Heinz Ludwig; Wolfgang Konig. Isis, 74(4):582–583, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232224.

Krohn:1983:BRB

[1012] Roger G. Krohn. Book review: The Cognitive Paradigm: Cognitive Science, a Newly Explored Approach to the Study of Cognition Applied to an Analysis of Science and Scientific Knowledge by Marc De Mey. Isis, 74 (4):583–584, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232225.

Low:1983:BRB

[1013] Reinhard Low. Book review: The Formation of the German Chemical Community (1720–1795) by Karl Hufbauer. Isis, 74(4):584–585, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232226.

Gieryn:1983:BRB

[1014] Thomas F. Gieryn. Book review: Quality in Science by Marcel C. La Follette. Isis, 74(4):585–586, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232227.

Bowler:1983:BRB

[1015] Peter J. Bowler. Book review: Statistics in Britain, 1865–1930: The Social Construction of Scientific Knowledge by Donald A. MacKenzie. Isis, 74(4):586–588, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232228.

Miller:1983:BRB

[1016] David Philip Miller. Book review: The Transformation of Intellectual Life in Victorian England by T. W. Heyck. Isis, 74(4):588–589, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232229.

Olesko:1983:BRBb

[1017] Kathryn Olesko. Book review: Die Natur als arbeitende Maschine: Zur Entstehung des Energiebegriffs in der Physik, 1840–1850 by Herbert Breger. Isis, 74(4):589, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232230.

Hufbauer:1983:BRB

[1018] Karl Hufbauer. Book review: Cosmic Discovery: The Search, Scope, and Heritage of Astronomy by Martin Harwit. Isis, 74(4):589–590, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232231.

Harman:1983:BRB

[1019] P. M. Harman. Book review: Das Selbstverstandnis der Physik im Wandel der Zeit: Vorlesungen zum Historischen Erfahrungsraum physikalischen Erkennens by Fritz Krafft. Isis, 74(4):590, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232232.

Hindle:1983:BRB

[1020] Brooke Hindle. Book review: A Calculating People: The Spread of Numeracy in Early America by Patricia Cline Cohen. Isis, 74(4):590–591, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232233.

Johnson:1983:BRB

[1021] Dale M. Johnson. Book review: Geschichte des Mannigfaltigkeitsbegriffs von Riemann bis Poincaré by Erhard Scholz. Isis, 74(4):591–592, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232234.

Hamlin:1983:BRB

[1022] Christopher Hamlin. Book review: The Road to Jaramillo: Critical Years of the Revolution in Earth Science by William Glen. Isis, 74(4):592–593, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232235.

Davies:1983:BRB

[1023] G. L. Herries Davies. Book review: Geology in the Nineteenth Century: Changing Views of a Changing World by Mott T. Greene. Isis, 74(4): 593–594, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232236.

Schlee:1983:BRB

[1024] Susan B. Schlee. Book review: No Sea Too Deep: The History of Oceanographic Instruments by Anita McConnell; 150 Years of Service on the Seas: A Pictorial History of the U.S. Naval Oceanographic Office from 1830 to 1980. Volume I: 1830–1946 by Marc I. Pinsel. Isis, 74(4):594– 595, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232237.

Davis:1983:BRB

[1025] Charles H. Davis. Book review: An Adventure in Applied Science: A History of the International Rice Research Institute by Robert F. Chandler. Isis, 74(4):595–596, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232238.

Mayr:1983:BRB

[1026] Ernst Mayr. Book review: Allgemeine Betrachtungen über die Triebe der Thiere hauptsachlich über ihre Kunsttriebe by Hermann Samuel Reimarus; Jurgen von Kempski. Isis, 74(4):596–597, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232239.

Vucinich:1983:BRBb

[1027] Alexander Vucinich. Book review: Russian Physicians in the Era of Reform and Revolution, 1856–1905 by Nancy Mandelker Frieden. Isis, 74(4):597–598, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232240.

Scarborough:1983:BRB

[1028] John Scarborough. Book review: Green Pharmacy: A History of Herbal Medicine by Barbara Griggs. Isis, 74(4):598–599, December 1983. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232241.

Risse:1983:BRB

[1029] Guenter B. Risse. Book review: Andreas Roschlaub und die Romantische Medizin: Die philosophischen Grundlagen der modernen Medizin by Nelly Tsouyopoulos. Isis, 74(4):599–600, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232242.

Parrott:1983:BRB

[1030] Bruce Parrott. Book review: Industrial Innovation in the Soviet Union by Ronald Amann; Julian Cooper. Isis, 74(4):600–601, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232243.

Williams:1983:BRB

[1031] Rosalind Williams. Book review: Energy and Transport: Historical Perspective on Policy Issues by George H. Daniels; Mark H. Rose. Isis, 74(4):601, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232244.

Brittain:1983:BRB

[1032] James E. Brittain. Book review: Engineers, Managers, and Politicians: The First Fifteen Years of Nationalised Electricity Supply in Britain by Leslie Hannah. Isis, 74(4):601–602, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232245.

Smith:1983:BRB

[1033] Roger Smith. Book review: The Secularization of the Soul: Psychical Research in Modern Britain by John J. Cerullo. Isis, 74(4):602–603, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232246.

Leventhal:1983:BRB

[1034] Robert S. Leventhal. Book review: From Locke to Saussure: Essays on the Study of Language and Intellectual History by Hans Aarsleff. Isis, 74 (4):603–604, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232247.

Thomson:1983:BRBb

[1035] Ron B. Thomson. Book review: Medieval Mensuration: “Quadrans vetus” and “Geometrie due sunt partes principales. …” by Nan L. Hahn. Isis, 74(4):604–605, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232248.

Thibodeau:1983:BRB

[1036] Sharon Gibbs Thibodeau. Book review: Traite de l’astrolabe by Jean Philopon; A. P. Segonds. Isis, 74(4):605–606, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232249.

McVaugh:1983:BRB

[1037] Michael McVaugh. Book review: El formulario de los hospitales de Ibn Abi l-Bayan by Jose Luis Valverde; Carmen Pena Munoz. Isis, 74(4): 606, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232250.

McCluskey:1983:BRBb

[1038] Stephen C. McCluskey. Book review: When Stars Came Down to Earth: Cosmology of the Skidi Pawnee Indians of North America by Von Del Chamberlain. Isis, 74(4):606–607, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232251.

Chamberlain:1983:BRA

[1039] Von Del Chamberlain. Book review: Archaeoastronomy in the Americas by Ray A. Williamson. Isis, 74(4):607–608, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232252.

Kish:1983:BRBb

[1040] George Kish. Book review: Lazarus Secretarius: The First Hungarian Mapmaker by Lajos Stegena. Isis, 74(4):608–609, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232253.

Dobbs:1983:BRBb

[1041] B. J. T. Dobbs. Book review: Contemporary Newtonian Research by Zev Bechler. Isis, 74(4):609–610, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232254.

McClellan:1983:BRB

[1042] James E. McClellan, III. Book review: Between Orthodoxy and the Enlightenment: Jean-Robert Chouet and the Introduction of Cartesian Science in the Academy of Geneva by Michael Heyd. Isis, 74(4):610–611, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232255.

Debus:1983:BRB

[1043] Allen G. Debus. Book review: Joan Baptista van Helmont: Reformer of Science and Medicine by Walter Pagel. Isis, 74(4):611–613, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232256.

Larson:1983:BRBd

[1044] James Larson. Book review: Charles Bonnet and the Order of the Known by Lorin Anderson. Isis, 74(4):613–614, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232257.

Taylor:1983:BRB

[1045] Kenneth L. Taylor. Book review: Voltaire’s Attitude toward Geology by Marguerite Carozzi. Isis, 74(4):614, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232258.

Schufle:1983:BRB

[1046] Joseph A. Schufle. Book review: Carl Wilhelm Scheele by Heinz Cassebaum. Isis, 74(4):614–615, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232259.

Langins:1983:BRB

[1047] Janis Langins. Book review: Calendar of the Correspondence of Pierre Simon Laplace by Roger Hahn. Isis, 74(4):615–616, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232260.

Burkhardt:1983:BRB

[1048] Richard W. Burkhardt, Jr. Book review: The Letters of Erasmus Darwin by Desmond King-Hele. Isis, 74(4):616–617, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232261.

Farber:1983:BRBb

[1049] Paul Lawrence Farber. Book review: Lamarck the Mythical Precursor: A Study of the Relations between Science and Ideology by Madeleine Barthelemy-Madaule; M. H. Shrank. Isis, 74(4):617, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232262.

Boewe:1983:BRB

[1050] Charles Boewe. Book review: Lucy Audubon: A Biography by Carolyn E. Delatte. Isis, 74(4):617–618, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232263.

Stanton:1983:BRB

[1051] William Stanton. Book review: Looking Far North: The Harriman Expedition to Alaska, 1899 by William H. Goetzmann; Kay Sloan. Isis, 74 (4):618–619, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232264.

Montgomery:1983:BRBb

[1052] William Montgomery. Book review: Charles Darwin, 1809–1882; Anton Dohrn, 1840–1909: Correspondence by Christine Groeben. Isis, 74(4): 619, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232265.

Pyne:1983:BRB

[1053] Stephen J. Pyne. Book review: Joseph LeConte: Gentle Prophet of Evolution by Lester Stephens. Isis, 74(4):619–620, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232266.

Numbers:1983:BRB

[1054] Ronald L. Numbers. Book review: Crusaders for Fitness: The History of American Health Reformers by James C. Whorton. Isis, 74(4):620–621, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232267.

Blackmore:1983:BRB

[1055] John Blackmore. Book review: Vom Positivismus zur “wissenschaftlichen Weltauffassung”–Am Beispiel der Wirkungsgeschichte von Ernst Mach in Osterreich von 1895 bis 1934 by Friedrich Stadler. Isis, 74(4):621–622, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232268.

Badash:1983:BRBb

[1056] Lawrence Badash. Book review: The Cult of the Atom: The Secret Papers of the Atomic Energy Commission by Daniel Ford. Isis, 74(4): 622–623, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232269.

Swenson:1983:BRB

[1057] Loyd S. Swenson, Jr. Book review: JPL and the American Space Program: A History of the Jet Propulsion Laboratory by Clayton R. Koppes.

Isis, 74(4):623–624, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232270.

Nelkin:1983:BRB

[1058] Dorothy Nelkin. Book review: Genetic Alchemy: The Social History of the Recombinant DNA Controversy by Sheldon Krimsky. Isis, 74(4): 624–625, December 1983. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232271.

Anonymous:1983:BMd

[1059] Anonymous. Back matter. Isis, 74(4):626–629, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232272.

Anonymous:1983:VI

[1060] Anonymous. Volume information. Isis, 74(4):630–640, December 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232205.

Anonymous:1983:FMe

[1061] Anonymous. Front matter. Isis, 74(5):1–4, ???? 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232202.

Neu:1983:OHE

[1062] John Neu. One hundred eighth critical bibliography of the history of science and its cultural influences. Isis, 74(5):5–215, ???? 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232203.

Anonymous:1983:BMe

[1063] Anonymous. Back matter. Isis, 74(5):216, ???? 1983. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232204.

Anonymous:1984:FMa

[1064] Anonymous. Front matter. Isis, 75(1):1–5, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232347.

Thackray:1984:ESS

[1065] Arnold Thackray. Editorial: Sarton, science, and history. Isis, 75(1): 6–9, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232348.

Edsall:1984:LJH

[1066] John T. Edsall. Lawrence J. Henderson and George Sarton. Isis, 75(1): 10–13, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232349.

Cohen:1984:HE

[1067] I. Bernard Cohen. A Harvard education. Isis, 75(1):13–21, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232350.

Hall:1984:BC

[1068] A Rupert Hall. Beginnings in Cambridge. Isis, 75(1):22–25, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232351.

Crombie:1984:BO

[1069] A. C. Crombie. Beginnings at Oxford. Isis, 75(1):25–28, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232352.

Kuhn:1984:PRT

[1070] Thomas S. Kuhn. Professionalization recollected in tranquility. Isis, 75 (1):29–32, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232353.

vanderGhinst:1984:SIV

[1071] Irenee van der Ghinst, Hosam Elkhadem, and George Sarton. Sarton and Irenee van der Ghinst. Isis, 75(1):33–38, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232354.

DeMey:1984:SEA

[1072] Marc De Mey and George Sarton. Sarton’s earliest ambitions at the University of Ghent. Isis, 75(1):39–45, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232355.

Dibner:1984:SLB

[1073] Bern Dibner and George Sarton. Sarton letters at the Burndy Library. Isis, 75(1):45–49, March 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232356.

Frangsmyr:1984:SN

[1074] Tore Frangsmyr, Johan Nordström, and George Sarton. Sarton and Nordström. Isis, 75(1):49–55, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232357.

Mikulinsky:1984:SV

[1075] S. R. Mikulinsky, George Sarton, and Vladimir Ivanovich Vernadsky. Sarton and Vernadsky. Isis, 75(1):56–62, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232358.

Hilts:1984:HSU

[1076] Victor L. Hilts. History of science at the University of Wisconsin. Isis, 75(1):63–94, March 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232359.

Rossiter:1984:HPS

[1077] Margaret W. Rossiter. The history and philosophy of science program at the national science foundation. Isis, 75(1):95–104, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232360.

Gilbert:1984:EKP

[1078] G. Nigel Gilbert and Michael Mulkay. Experiments are the key: Participants’ histories and historians’ histories of science. Isis, 75(1):105–125, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232361.

Shapin:1984:THR

[1079] Steven Shapin. Talking history: Reflections on discourse analysis. Isis, 75(1):125–130, March 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232362.

Holmes:1984:LKI

[1080] Frederic L. Holmes. Lavoisier and Krebs: The individual scientist in the near and deeper past. Isis, 75(1):131–142, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232363.

Diamond:1984:BWP

[1081] Solomon Diamond. Buckle, wundt, and Psychology’s use of history. Isis, 75(1):143–152, March 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232364.

Wheaton:1984:ISH

[1082] Bruce R. Wheaton. Inventory of Sources for History of Twentieth– Century Physics (ISHTCP). Isis, 75(1):153–157, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232365.

Elliott:1984:JCA

[1083] Clark A. Elliott. Joint committee on archives of science and technology (jcast): Summary from the final report. Isis, 75(1):158–162, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232366.

Aspray:1984:LIH

[1084] William Aspray. Literature and institutions in the history of computing. Isis, 75(1):162–170, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232367.

White:1984:BRB

[1085] Lynn White, Jr. and Jonathan D. Spence. Book review: Science and Civilisation in China. Isis, 75(1):171–189, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232368.

Sicherman:1984:BRB

[1086] Barbara Sicherman, John Lankford, and Daniel J. Kevles. Book review: Women Scientists in America: Struggles and Strategies to 1940. Isis, 75(1):189–203, March 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232369.

Molland:1984:BRB

[1087] A. George Molland. Book review: Science Deified & Science Defied: The Historical Significance of Science in Western Culture from the Bronze Age to the Beginnings of the Modern Era ca. 3500 B.C. to ca. A.D. 1640 by Richard Olson. Isis, 75(1):204, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232370.

Multhauf:1984:BRB

[1088] R. P. Multhauf. Book review: Dictionary of the History of Science by W. F. Bynum; E. J. Browne; Roy Porter. Isis, 75(1):204–205, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232371.

Hall:1984:BRBa

[1089] A. Rupert Hall. Book review: Editing Texts in the History of Science and Medicine by Trevor H. Levere. Isis, 75(1):205–206, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232372.

Paul:1984:BRBa

[1090] H. W. Paul. Book review: Nature et methode de l’histoire des sciences by Francois Russo. Isis, 75(1):206–207, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232373.

VanEgmond:1984:BRB

[1091] Warren Van Egmond. Book review: A Bibliography of Early Modern Algebra, 1500–1800. by Robin E. Rider. Isis, 75(1):207–208, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232374.

Elliott:1984:BRB

[1092] Clark A. Elliott. Book review: The History of Science and Technology in the United States: A Critical and Selective Bibliography by Marc Rothenberg. Isis, 75(1):208–209, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232375.

Morrell:1984:BRB

[1093] Jack Morrell. Book review: The Manuscript Papers of British Scientists, 1600–1940 by The Royal Commission on Historical Manuscripts. Isis, 75 (1):209–210, March 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232376.

Freudenthal:1984:BRB

[1094] Gad Freudenthal. Book review: Der Wissenschaftler und das Irrationale by Hans Peter Duerr; Die Logik der Weltbilder: Sinnstrukturen im Wandel der Geschichte by Gunter Dux. Isis, 75(1):210–211, March 1984.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232377.

Dunlap:1984:BRB

[1095] Thomas R. Dunlap. Book review: Science in a Free Society by Paul Feyerabend. Isis, 75(1):211–212, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232378.

Dupree:1984:BRB

[1096] A. Hunter Dupree. Book review: A Patron for Pure Science. Volume I: The National Science Foundation’s Formative Years, 1945–1957 by J. Merton England. Isis, 75(1):212–213, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232379.

Kargon:1984:BRB

[1097] Robert Kargon. Book review: Science for Industry: A Short History of the Imperial College of Science and Technology and Its Antecedents by A. Rupert Hall. Isis, 75(1):213, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232380.

Hatch:1984:BRB

[1098] Robert A. Hatch. Book review: The Royal Society and Its Fellows, 1660–1700: The Morphology of an Early Scientific Institution by Michael Hunter. Isis, 75(1):213–215, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232381.

Woolf:1984:BRB

[1099] Patricia Woolf. Book review: Betrayers of the Truth: Fraud and Deceit in the Halls of Science by William Broad; Nicholas Wade. Isis, 75(1): 215, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232382.

Galatzer-Levy:1984:BRB

[1100] Robert M. Galatzer-Levy. Book review: Jewish Origins of the Psychoanalytic Movement by Dennis B. Klein. Isis, 75(1):216, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232383.

Slade:1984:BRB

[1101] Joseph W. Slade. Book review: Bucknell Review by Harry Garvin; Volume 27, Number 2: Science and Literature by James M. Heath. Isis, 75(1):216–217, March 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232384.

Deane:1984:BRB

[1102] Thatcher E. Deane. Book review: Zhongguo da baike quanshu: Tianwenxue (The Greater Encyclopedia of China: Astronomy.) by Zhang Yuzhe. Isis, 75(1):217, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232385.

Dott:1984:BRB

[1103] R. H. Dott, Jr. Book review: Earth Science History: Journal of the History of the Earth Sciences Society by Gerald M. Freeman; Kennard B. Bork. Isis, 75(1):217–218, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232386.

Dunbar:1984:BRBa

[1104] G. S. Dunbar. Book review: Grosse Geographen: Pioniere–Aussenseiter– Gelehrte by Hanno Beck. Isis, 75(1):218–219, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232387.

Burkhardt:1984:BRB

[1105] Richard W. Burkhardt, Jr. Book review: Ethology: Its Nature and Relations with Other Sciences by Robert A. Hinde. Isis, 75(1):219–220, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232388.

Langham:1984:BRB

[1106] Ian Langham. Book review: Missing Links: The Hunt for Earliest Man by John Reader. Isis, 75(1):220–221, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232389.

Matthews:1984:BRB

[1107] Fred Matthews. Book review: American Sociological Theory: A Critical History by Robert Bierstedt. Isis, 75(1):221–222, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232390.

Handler:1984:BRBa

[1108] Richard Handler. Book review: Freud and Anthropology: A History and Reappraisal by Edwin R. Wallace. Isis, 75(1):222–223, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232391.

Reeds:1984:BRBa

[1109] Karen Reeds. Book review: The Discovery of Insulin by Michael Bliss. Isis, 75(1):223–224, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232392.

Stevens:1984:BRB

[1110] Rosemary Stevens. Book review: The Healthiest City: Milwaukee and the Politics of Health Reform by Judith Walzer Leavitt. Isis, 75(1): 224, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232393.

Brieger:1984:BRB

[1111] Gert H. Brieger. Book review: A Celebration of Medical History: The Fiftieth Anniversary of the Johns Hopkins Institute of the History of Medicine and the Welch Medical Library by Lloyd G. Stevenson. Isis, 75(1):224–225, March 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232394.

Hughes:1984:BRB

[1112] Thomas P. Hughes. Book review: The Pursuit of Power: Technology, Armed Force, and Society since A.D. 1000 by William H. McNeill. Isis, 75(1):225–227, March 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232395.

Pursell:1984:BRBa

[1113] Carroll W. Pursell, Jr. Book review: A Short History of Twentieth-Century Technology, c. 1900–c. 1950 by Trevor I. Williams. Isis, 75(1): 227–228, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232396.

Grant:1984:BRB

[1114] Edward Grant. Book review: Aristotle and the Renaissance by Charles B. Schmitt. Isis, 75(1):228–229, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232397.

Edwards:1984:BRB

[1115] William F. Edwards. Book review: Prelude to Galileo: Essays on Medieval and Sixteenth-Century Sources of Galileo’s Thought by William Wallace. Isis, 75(1):229–230, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232398.

Emerson:1984:BRB

[1116] Roger L. Emerson. Book review: The Radical Enlightenment: Pantheists, Freemasons, and Republicans by Margaret Jacob. Isis, 75(1):230– 231, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232399.

Gaukroger:1984:BRB

[1117] Stephen Gaukroger. Book review: L’orientation de la science cartésienne by Geraud Tournadre. Isis, 75(1):231, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232400.

Bowler:1984:BRB

[1118] Peter J. Bowler. Book review: Archetypes and Ancestors: Palaeontology in Victorian London, 1850–1875 by Adrian Desmond. Isis, 75(1):232– 233, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232401.

Olesko:1984:BRB

[1119] Kathryn Olesko. Book review: The Mathematical Manuscripts of Karl Marx by Karl Marx; C. Aronson; M. Meo. Isis, 75(1):233–234, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232402.

Stuewer:1984:BRN

[1120] Roger H. Stuewer. Book review: Niels Bohr; L. Rosenfeld; J. Rud Nielson; Ulrich Hoyer: Collected Works, Volumes 1–4. Isis, 75(1):234–235, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232403.

Erikson:1984:BRB

[1121] G. E. Erikson. Book review: The Seven Ages of a Medical Scientist: An Autobiography by George W. Corner. Isis, 75(1):235–236, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232404.

Anonymous:1984:BMa

[1122] Anonymous. Back matter. Isis, 75(1):237–240, March 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232405.

Anonymous:1984:FMb

[1123] Anonymous. Front matter. Isis, 75(2):241–247, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231823.

Gillespie:1984:BFB

[1124] Richard Gillespie. Ballooning in France and Britain, 1783–1786: Aerostation and adventurism. Isis, 75(2):248–268, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231824.

Wisan:1984:GPS

[1125] Winifred Lovell Wisan. Galileo and the process of scientific creation. Isis, 75(2):269–286, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231825.

Park:1984:BGD

[1126] Katharine Park, Lorraine J. Daston, and Peter L. Galison. Bacon, Galileo, and Descartes on imagination and analogy. Isis, 75(2):287–289, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231826.

Park:1984:BEG

[1127] Katharine Park. Bacon’s “enchanted glass”. Isis, 75(2):290–302, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231827.

Daston:1984:GAI

[1128] Lorraine J. Daston. Galilean analogies: Imagination at the bounds of sense. Isis, 75(2):302–310, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231828.

Galison:1984:DCI

[1129] Peter Galison. Descartes’s comparisons: From the invisible to the visible. Isis, 75(2):311–326, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231829.

Ruscelli:1984:ASG

[1130] Girolamo Ruscelli, William Eamon, and Francoise Paheau. The Accademia Segreta of Girolamo Ruscelli: A Sixteenth-Century Italian scientific society. Isis, 75(2):327–342, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231830.

Perrin:1984:DLR

[1131] C. E. Perrin. Did Lavoisier report to the Academy of Sciences on his own book? Isis, 75(2):343–348, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231831.

Bulmer-Thomas:1984:GTP

[1132] Ivor Bulmer-Thomas. Guldin’s theorem — or Pappus’s? Isis, 75(2):348– 352, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231832.

Gregory:1984:AMH

[1133] Frederick Gregory and Robert A. Hatch. Annual meeting of the History of Science Society, 27–30 October 1983. Isis, 75(2):353–360, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231833.

Reingold:1984:IS

[1134] Nathan Reingold and Ann Hibner Koblitz. International symposia. Isis, 75(2):360–363, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231834.

Davis:1984:EWP

[1135] Audrey B. Davis. Eloge: Walter Pagel, 14 November 1898–1 April 1983. Isis, 75(2):363–365, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231835.

Mahoney:1984:BRB

[1136] Michael S. Mahoney. Book review: The Mathematical Papers of Isaac Newton, by Isaac Newton; D. T. Whiteside. Volume VIII: 1697–1722. Isis, 75(2):366–372, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231836.

McMullin:1984:BRBa

[1137] Ernan McMullin. Book review: Science and Hypothesis: Historical Essays on Scientific Methodology. Isis, 75(2):372–377, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231837.

Klein:1984:BRA

[1138] Martin J. Klein. Book review: Abraham Pais: “Subtle Is the Lord … ”: The Science and the Life of Albert Einstein. Isis, 75(2):377–379, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231838.

Hannah:1984:BRB

[1139] Leslie Hannah. Book review: Networks of Power: Electrification in Western Society, 1880–1930. Isis, 75(2):379–382, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231839.

Steneck:1984:BRB

[1140] Nicholas H. Steneck. Book review: Much Ado about Nothing: Theories of Space and Vacuum from the Middle Ages to the Scientific Revolution by Edward Grant. Isis, 75(2):383–384, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231840.

Williams:1984:BRBa

[1141] L. Pearce Williams. Book review: Geschichte der Naturwissenschaften by Hans Wussing. Isis, 75(2):384–385, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231841.

Rothenberg:1984:BRB

[1142] Marc Rothenberg. Book review: The Papers of Nathaniel Bowditch in the Boston Public Library: Guide to the Microfilm Edition by James W. Montgomery, Jr.; Laura V. Monti. Isis, 75(2):385, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231842.

Schneer:1984:BRB

[1143] Cecil J. Schneer. Book review: The Earth Sciences: An Annotated Bibliography by Roy Porter. Isis, 75(2):385–386, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231843.

Fine:1984:BRB

[1144] Arthur Fine. Book review: How the Laws of Physics Lie by Nancy Cartwright. Isis, 75(2):386–387, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231844.

Siegel:1984:BRB

[1145] Daniel M. Siegel. Book review: Metaphysics and Natural Philosophy: The Problem of Substance in Classical Physics by P. M. Harman; Energy, Force, and Matter: The Conceptual Development of Nineteenth-Century Physics by P. M. Harman. Isis, 75(2):387–389, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231845.

Hacking:1984:BRB

[1146] Ian Hacking. Book review: Scientific Explanation: Papers Based on Herbert Spencer Lectures Given in the Oxford University by A. F. Heath. Isis, 75(2):389–390, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231846.

Cassidy:1984:BRB

[1147] David C. Cassidy. Book review: Scientific Explanation and Atomic Physics by Edward M. MacKinnon. Isis, 75(2):390–391, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231847.

Barker:1984:BRB

[1148] Peter Barker. Book review: Philosophy and the New Physics by Jonathan Powers. Isis, 75(2):391–392, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231848.

Dunbar:1984:BRBb

[1149] G. S. Dunbar. Book review: La formation de l’école française de géographie (1870–1914) by Vincent Berdoulay. Isis, 75(2):392–393, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231849.

McClellan:1984:BRB

[1150] James E. McClellan, III. Book review: The Papers of the Dublin Philosophical Society, 1683–1708 by K. Theodore Hoppen. Isis, 75(2):393, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231850.

Sheets-Pyenson:1984:BRB

[1151] Susan Sheets-Pyenson. Book review: Wiley: One Hundred and Seventy Five Years of Publishing by John Hammond Moore; A. Wayne Anderson. Isis, 75(2):393–394, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231851.

Hannaway:1984:BRB

[1152] Owen Hannaway. Book review: Praelectiones Physicae: Historischkritische Edition by Joachim Jungius; Christopher Meinel. Isis, 75(2): 394–395, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231852.

Rocke:1984:BRBa

[1153] Alan J. Rocke. Book review: Liebigs Experimentalvorlesung: Vorlesungsbuch und Kekules Mitschrift by Otto Paul Kratz; Claus Priesner. Isis, 75(2):395–396, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231853.

Bedini:1984:BRBa

[1154] Silvio A. Bedini. Book review: Los relojes y automatas de Juanelo Turriano by Jose A. Garcia-Diego. Isis, 75(2):396–397, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231854.

Gaston:1984:BRB

[1155] Jerry Gaston. Book review: Science and Society: Studies in the Sociology of Science by Joseph Agassi. Isis, 75(2):397, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231855.

Eastwood:1984:BRBa

[1156] Bruce Eastwood. Book review: The Art of Describing: Dutch Art in the Seventeenth Century by Svetlana Alpers. Isis, 75(2):398, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231856.

VanHelden:1984:BRB

[1157] Albert Van Helden. Book review: The Mathematical Science of Christopher Wren by J. A. Bennett. Isis, 75(2):398–399, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231857.

Hall:1984:BRBb

[1158] Diana Long Hall. Book review: A Comparison of Selected Poetic and Scientific Works of Albrecht von Haller by Dorothy Roller Wiswall; Otto G. Graf. Isis, 75(2):399–400, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231858.

Diamond:1984:BRB

[1159] Harold G. Diamond. Book review: Collected Papers of J. E. Littlewood by J. E. Littlewood; London Mathematical Society. Isis, 75(2):400–401, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231859.

Birkhoff:1984:BRB

[1160] Garrett Birkhoff and William Aspray. Book review: Zermelo’s Axiom of Choice: Its Origins, Development, and Influence by Gregory H. Moore. Isis, 75(2):401–402, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231860.

Heims:1984:BRB

[1161] Steve J. Heims. Book review: Neyman: From Life by Constance Reid. Isis, 75(2):402–403, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231861.

Buchwald:1984:BRB

[1162] Jed Z. Buchwald. Book review: La polarisation rotatoire naturelle: De la structure de la lumière à celle des molécules by Jean Rosmorduc. Isis, 75 (2):403, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231862.

Schultz:1984:BRB

[1163] Susan Schultz. Book review: Frontiers of Geological Exploration of Western North America: A Symposium by Alan E. Leviton; Peter U. Rodda; Ellis L. Yochelson; Michele L. Aldrich. Isis, 75(2):403–404, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231863.

Farley:1984:BRB

[1164] John Farley. Book review: The Eclipse of Darwinism: Anti-Darwinian Evolution Theories in the Decades around 1900 by Peter J. Bowler. Isis, 75(2):404–405, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231864.

Egerton:1984:BRBa

[1165] Frank N. Egerton. Book review: The Secular Ark: Studies in the History of Biogeography by Janet Browne. Isis, 75(2):405–406, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231865.

Casada:1984:BRB

[1166] James A. Casada. Book review: Botanical Exploration of Southern Africa: An Illustrated History of Early Botanical Literature on the Cape Flora [with] Biographical Accounts of the Leading Plant Collectors and Their Activities in Southern Africa from the Days of the East India Company until Modern Times by Mary Gunn; L. E. Codd. Isis, 75(2):406– 407, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231866.

Jones:1984:BRBa

[1167] Greta Jones. Book review: The Idea of Race in Science: Great Britain, 1800–1960 by Nancy Stepan. Isis, 75(2):407–408, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231867.

Meikle:1984:BRBa

[1168] Jeffrey L. Meikle. Book review: American Design Ethic: A History of Industrial Design by Arthur J. Pulos. Isis, 75(2):408, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231868.

Langins:1984:BRB

[1169] Janis Langins. Book review: Stronger Than a Hundred Men: A History of the Vertical Water Wheel by Terry S. Reynolds. Isis, 75(2):408–409, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231869.

Fries:1984:BRB

[1170] Russell I. Fries. Book review: Charles Blacker Vignoles: Romantic Engineer by K. H. Vignoles. Isis, 75(2):409–410, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231870.

Yans-McLaughlin:1984:BRB

[1171] Virginia Yans-McLaughlin. Book review: Margaret Mead and Samoa: The Making and Unmaking of an Anthropological Myth by Derek Freeman. Isis, 75(2):410–411, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231871.

Handler:1984:BRBb

[1172] Richard Handler. Book review: Ruth Benedict: Patterns of a Life by Judith Schachter Modell. Isis, 75(2):411–412, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231872.

Hinsley:1984:BRB

[1173] Curtis M. Hinsley. Book review: The Fall of Natural Man: The American Indian and the Origins of Comparative Ethnology by Anthony Pagden. Isis, 75(2):412–413, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231873.

Toomer:1984:BRB

[1174] G. J. Toomer. Book review: Science and Speculation: Studies in Hellenistic Theory and Practice by Jonathan Barnes; Jacques Brunschwig; Myles Burnyeat; Malcolm Schofield. Isis, 75(2):413–415, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231874.

Mueller:1984:BRB

[1175] Ian Mueller. Book review: L’architecture du divin: Mathematique et philosophie chez Plotin et Proclus by Annick Charles-Saget. Isis, 75(2): 415, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231875.

Unguru:1984:BRB

[1176] Sabetai Unguru. Book review: The Medieval Latin Translation of the Data of Euclid by Shuntaro Ito. Isis, 75(2):415–416, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231876.

Dales:1984:BRBa

[1177] Richard C. Dales. Book review: Roger Bacon’s Philosophy of Nature: A Critical Edition, with English Translation, Introduction, and Notes, of “De multiplicatione specierum” and “De speculis comburentibus” by David C. Lindberg. Isis, 75(2):416–417, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231877.

Siraisi:1984:BRBa

[1178] Nancy G. Siraisi. Book review: Girolamo Cardano, 1501–1576: Physician, Natural Philosopher, Mathematician, Astrologer, and Interpreter of Dreams by Markus Fierz; Helga Niman. Isis, 75(2):417, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231878.

Oppenheimer:1984:BRB

[1179] Jane M. Oppenheimer. Book review: Carteggio (1768–1798) by Leopoldo M. A. Caldani; Lazzaro Spallanzani; Giuseppe Ongaro. Isis, 75(2):417– 418, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231879.

Osler:1984:BRB

[1180] Margaret J. Osler. Book review: Principles of Philosophy by Rene Descartes; Valentine Rodger Miller; Reese P. Miller. Isis, 75(2):418– 419, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231880.

Shea:1984:BRBa

[1181] William R. Shea. Book review: Amici e corrispondenti di Galileo by Antonio Favoro; Paulo Galluzi. Isis, 75(2):419, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231881.

Daumas:1984:BRB

[1182] Maurice Daumas. Book review: The Montgolfier Brothers and the Invention of Aviation, 1783–1784: With a Word on the Importance of Ballooning for the Science of Heat and the Art of Building Railroads by Charles Coulston Gillispie. Isis, 75(2):419–421, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231882.

Shapin:1984:BRB

[1183] Steven Shapin. Book review: Henry Stubbe, Radical Protestantism and the Early Enlightenment by James R. Jacob. Isis, 75(2):421–422, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231883.

Forbes:1984:BRB

[1184] Eric G. Forbes. Book review: Briefwechsel zwischen Alexander von Humboldt und Peter Gustav Lejeune Dirichlet by Kurt-R. Bierman. Isis, 75 (2):422–423, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231884.

Curd:1984:BRB

[1185] Martin Curd. Book review: Ludwig Boltzmann: Man–Physicist– Phylosopher by Engelbert Broda; Larry Gay. Isis, 75(2):423, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231885.

Bernstein:1984:BRB

[1186] Howard Bernstein. Book review: Philosophie als Analysis: Studien zur Entwicklung philosophischer Analysiskonzeptionen unter dem Einfluss mathematischer Methodenmodelle im 17. und fruhen 18. Jahrhundert by Hans-Jurgen Engfer. Isis, 75(2):423–424, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231886.

Blaisdell:1984:BRB

[1187] Muriel L. Blaisdell. Book review: Evolution without Evidence: Charles Darwin and the Origin of Species by Barry G. Gale. Isis, 75(2):424– 425, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231887.

Reeves:1984:BRB

[1188] Barbara J. Reeves. Book review: Pio IX e le macchine ottiche, con reproduzione anastatica del Saggio fisico-matematico proposto da Giovan Maria Mastai Ferretti a conclusione dei suoi studi presso le Scuole Pie di Volterra il 7 settembre 1809 by Pier Luigi Pizzamiglio; Giorgio Tabarroni. Isis, 75(2):425–426, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231888.

Finocchiaro:1984:BRB

[1189] Maurice A. Finocchiaro. Book review: Scienza e filosofia nella cultura positivistica by Antonio Santucci. Isis, 75(2):426–427, June 1984. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231889.

Rudwick:1984:BRB

[1190] Martin Rudwick. Book review: Adam Sedgwick: Geologist and Dalesman, 1785–1873: A Biography in Twelve Themes by Colin Speakman. Isis, 75(2):427, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231890.

Rossiter:1984:BRB

[1191] Margaret W. Rossiter. Book review: George Washington Carver: Scientist and Symbol by Linda O. McMurry. Isis, 75(2):427–428, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231891.

Glick:1984:BRS

[1192] Thomas F. Glick. Book review: Societat Catalana de Ciències Físiques, Químiques i Matemátiques: Centenari de la naixença d’Albert Einstein (Catalan) [Centenary of the Birth of Albert Einstein]. Isis, 75(2):428– 429, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231892.

Goldberg:1984:BRH

[1193] Stanley Goldberg. Book review: Harry Woolf: Some Strangeness in the Proportion: A Centennial Symposium to Celebrate the Achievements of Albert Einstein; Gerald Holton and Yehuda Elkana: Albert Einstein: Historical and Cultural Perspectives. Isis, 75(2):429–430, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231893.

DelSesto:1984:BRBa

[1194] Steven L. Del Sesto. Book review: Rationality and Ritual: The Windscale Inquiry and Nuclear Decisions in Britain by Bryan Wynne. Isis, 75(2): 430–431, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231894.

Blake:1984:BRB

[1195] John B. Blake. Book review: Disease and Its Control: The Shaping of Modern Thought by Robert P. Hudson. Isis, 75(2):432, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231895.

Baatz:1984:BRBa

[1196] Simon Baatz. Book review: A Catalogue of the Manuscripts and Archives of the Library of The College of Physicians of Philadelphia by Rudolf Hirsch. Isis, 75(2):432–433, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231896.

Warner:1984:BRB

[1197] John Harley Warner. Book review: Sources in the History of American Pharmacology by John Parascandola; Elizabeth Keeney. Isis, 75(2):433– 434, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231897.

Whorton:1984:BRB

[1198] James Whorton. Book review: History of Clinical Chemistry by J. Buttner. Isis, 75(2):434, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231898.

Brandt:1984:BRB

[1199] Allan M. Brandt. Book review: Compulsory Health Insurance: The Continuing Debate by Ronald L. Numbers. Isis, 75(2):434–435, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231899.

Carson:1984:BRB

[1200] Ronald A. Carson. Book review: Health/Medicine and the Faith Traditions: An Inquiry into Religion and Medicine by Martin E. Marty; Kenneth L. Vaux. Isis, 75(2):435, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231900.

Absuch:1984:BRB

[1201] Tzvi Absuch. Book review: Die babylonisch-assyrische Medizin in Texten und Untersuchungen. Volumes V + VI: Keilschrifttexte aus Ninive by Franz Kocher. Isis, 75(2):436–437, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231901.

Scarborough:1984:BRBa

[1202] John Scarborough. Book review: Galen’s Commentary on the Hippocratic Treatise “Airs, Waters, Places” in the Hebrew Translation of Solomon ha-Me ati by Galen; Abraham Wasserstein. Isis, 75(2):437, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231902.

Demaitre:1984:BRB

[1203] Luke Demaitre. Book review: Medizin im mittelalterlichen Abendland by Gerhard Baader; Gundolf Keil. Isis, 75(2):437–438, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231903.

Savage-Smith:1984:BRBa

[1204] Emilie Savage-Smith. Book review: Vorlesungen über die Medizin im Islam by Felix Klein-Franke. Isis, 75(2):438–440, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231904.

Eide:1984:BRB

[1205] Elling O. Eide. Book review: An Enumeration of Chinese Materia Medica by Shiu-ying Hu; Y. C. Kong; Paul P. H. But. Isis, 75(2):440–441, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231905.

Gelfand:1984:BRBa

[1206] Toby Gelfand. Book review: William Hunter, 1718–1783: A Memoir by Samuel Fort Simmons; John Hunter; C. H. Brock. Isis, 75(2):441, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231906.

Maulitz:1984:BRB

[1207] Russell C. Maulitz. Book review: Laennec: Catalogue des manuscrits scientifiques by Lydie Boulle; Mirko D. Grmek; Catherine Lupovici; Janine Samion-Contet. Isis, 75(2):442, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231907.

Benson:1984:BRB

[1208] Keith R. Benson. Book review: The Etiology, Concept, and Prophylaxis of Childbed Fever by Ignaz Semmelweis; K. Codell Carter. Isis, 75(2): 442–443, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231908.

Parssinen:1984:BRB

[1209] Terry M. Parssinen. Book review: Endangered Lives: Public Health in Victorian Britain by Anthony S. Wohl. Isis, 75(2):443, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231909.

Mazumdar:1984:BRB

[1210] Pauline M. H. Mazumdar. Book review: Cholera: The American Scientific Experience, 1947–1980 by W. E. van Heyningen; John R. Seal. Isis, 75(2):444, June 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231910.

Marks:1984:BRB

[1211] Harry M. Marks. Book review: Pure Politics and Impure Science: The Swine Flu Affair by Arthur M. Silverstein. Isis, 75(2):444–445, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231911.

Anonymous:1984:BMb

[1212] Anonymous. Back matter. Isis, 75(2):446–??, June 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231912.

Anonymous:1984:FMc

[1213] Anonymous. Front matter. Isis, 75(3):451–455, September 1984. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232937.

Maienschein:1984:WDS

[1214] Jane Maienschein. What determines sex? A study of converging approaches, 1880–1916. Isis, 75(3):456–480, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232938.

Eastwood:1984:DRS

[1215] Bruce Stansfield Eastwood. Descartes on refraction: Scientific versus rhetorical method. Isis, 75(3):481–502, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232939.

Crawford:1984:AAH

[1216] Elisabeth Crawford. Arrhenius, the atomic hypothesis, and the 1908 Nobel prizes in physics and chemistry. Isis, 75(3):503–522, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232940.

Turner:1984:CCN

[1217] Steven Turner, Edward Kerwin, and David Woolwine. Careers and creativity in Nineteenth-Century physiology: Zloczower redux. Isis, 75(3): 523–529, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232941.

Pan:1984:CDC

[1218] Jixing Pan. Charles Darwin’s Chinese sources. Isis, 75(3):530–534, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232942.

Kidwell:1984:WAB

[1219] Peggy Aldrich Kidwell. Women astronomers in Britain, 1780–1930. Isis, 75(3):534–546, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232943.

McReynolds:1984:CTO

[1220] Paul McReynolds and Klaus Ludwig. Christian Thomasius and the origin of psychological rating scales. Isis, 75(3):546–553, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232944.

Numbers:1984:SRA

[1221] Ronald L. Numbers. Science and religion in the American society of Church history. Isis, 75(3):554, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232945.

Gregory:1984:BRB

[1222] Frederick Gregory. Book review: The Strategy of Life: Teleology and Mechanics in Nineteenth-Century German Biology. Isis, 75(3):555–558, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232946.

Reverby:1984:BRB

[1223] Susan Reverby. Book review: The Social Transformation of American Medicine. Isis, 75(3):559–562, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232947.

Vucinich:1984:BRB

[1224] Alexander Vucinich. Book review: The Revolution in Science 1500– 1750 by A. Rupert Hall. Isis, 75(3):563, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232948.

Meikle:1984:BRBb

[1225] Jeffrey L. Meikle. Book review: The Show of Science by Robin E. Rider. Isis, 75(3):564, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232949.

Shank:1984:BRB

[1226] Michael Shank. Book review: Science: Its History and Development among the World’s Cultures by Colin Ronan. Isis, 75(3):564–565, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232950.

McLeod:1984:BRB

[1227] Alexander Dean McLeod. Book review: New Perspective on the History of Science by Seiki Shia; Frederick Y. Coe. Isis, 75(3):565–566, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232951.

Shapere:1984:BRB

[1228] Dudley Shapere. Book review: A Bibliography of the Philosophy of Science, 1945–1981 by Richard J. Blackwell. Isis, 75(3):566–567, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232952.

Crawford:1984:BRB

[1229] Susan Crawford. Book review: Sociology of Sciences: An Annotated Bibliography on Invisible Colleges, 1972–1981 by Daryl E. Chubin. Isis, 75(3):567, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232953.

Thackray:1984:BRB

[1230] Arnold Thackray. Book review: Information Sources in the History of Science and Medicine by Pietro Corsi; Paul Weindling. Isis, 75(3):567– 569, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232954.

Browne:1984:BRB

[1231] Janet Browne. Book review: Bibliography of the Rhinoceros: An Analysis of the Literature on the Recent Rhinoceroses in Culture, History, and Biology by L. C. Rookmaaker. Isis, 75(3):569, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232955.

McVaugh:1984:BRB

[1232] Michael R. McVaugh. Book review: Diccionario historico de la ciencia moderna en España by Jose M. Lopez Pinero; Thomas F. Glick; Victor Navarro Brotons; Eugenio Portela Marcos. Isis, 75(3):569–570, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232956.

Rhees:1984:BRB

[1233] David Rhees. Book review: A Biographical Dictionary of Scientists by Trevor I. Williams. Isis, 75(3):570–571, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232957.

Weingart:1984:BRB

[1234] Peter Weingart. Book review: Between Science and Values by Loren R. Graham. Isis, 75(3):571–572, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232958.

Hall:1984:BRBc

[1235] A. Rupert Hall. Book review: Nature Mathematized: Historical and Philosophical Case Studies in Classical Modern Natural Philosophy. Volume I by William Shea. Isis, 75(3):572–573, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232959.

Donovan:1984:BRB

[1236] Arthur Donovan. Book review: Education and Opportunity in Victorian Scotland: Schools and Universities by R. D. Anderson. Isis, 75(3):573, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232960.

Christianson:1984:BRB

[1237] Eric H. Christianson. Book review: A Lamp in the Forest: Natural Philosophy in Transylvania University, 1799–1859 by Ash Gobar; J. Hill Hamon. Isis, 75(3):573–575, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232961.

OHara:1984:BRB

[1238] James G. O’Hara. Book review: A History of Mathematics Education in England by Geoffrey Howson. Isis, 75(3):575, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232962.

Weisz:1984:BRB

[1239] George Weisz. Book review: The Transformations of Higher Learning 1860–1930: Expansion, Diversification, Social Opening, and Professionalization in England, Germany, Russia, and the United States by Konrad H. Jarausch. Isis, 75(3):575–577, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232963.

Gillespie:1984:BRB

[1240] Richard Gillespie. Book review: Partial Progress: The Politics of Science and Technology by David Albury; Joseph Schwartz. Isis, 75(3):577, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232964.

Inkster:1984:BRB

[1241] Ian Inkster. Book review: Sociology of Scientific Knowledge: A Source Book by H. M. Collins. Isis, 75(3):577–578, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232965.

Epstein:1984:BRB

[1242] Julia L. Epstein. Book review: Women in Science: Portraits from a World in Transition by Vivian Gornick. Isis, 75(3):578–579, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232966.

Cozzens:1984:BRB

[1243] Susan E. Cozzens. Book review: The Social Construction of Science: A Comparative Study of Goal Direction, Research Evolution, and Legitimation by Tom Jagtenberg. Isis, 75(3):579–580, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232967.

Webb:1984:BRBa

[1244] George E. Webb. Book review: Creationism, Science, and the Law by Marcel C. La Follette; The Creation Controversy by Dorothy Nelkin. Isis, 75(3):580–581, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232968.

Kevles:1984:BRB

[1245] Daniel J. Kevles. Book review: Birth Control and the Population Question in England, 1877–1930 by Richard Allen Soloway. Isis, 75(3):581– 582, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232969.

Larson:1984:BRBa

[1246] Ed Larson. Book review: Science and Law: An Essential Alliance by William A. Thomas. Isis, 75(3):582–583, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232970.

Wachhorst:1984:BRB

[1247] Wyn Wachhorst. Book review: The Culture of Time and Space, 1880– 1918 by Stephen Kern. Isis, 75(3):583–584, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232971.

Jacob:1984:BRB

[1248] Margaret C. Jacob. Book review: The Changing of the Gods by Frank E. Manuel. Isis, 75(3):584–585, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232972.

Bozeman:1984:BRB

[1249] T. D. Bozeman. Book review: The Princeton Theology: Scripture, Science and Theological Method from Archibald Alexander to Benjamin Breckinridge Warfield by Mark A. Noll. Isis, 75(3):585–586, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232973.

Jones:1984:BRBb

[1250] Greta Jones. Book review: The Wider Domain of Evolutionary Thought by David Oldroyd; Ian Langham. Isis, 75(3):586–587, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232974.

Theerman:1984:BRB

[1251] Paul Theerman. Book review: Relations of Literature and Science, 1980– 1981 by Walter Schatzberg. Isis, 75(3):587–588, September 1984. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232975.

Attfield:1984:BRB

[1252] Robin Attfield. Book review: Man and the Natural World: A History of the Modern Sensibility by Keith Thomas. Isis, 75(3):588–589, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232976.

Bedini:1984:BRBb

[1253] Silvio A. Bedini. Book review: Il circolo mathematico di Palermo by Aldo Brigaglia; Guido Masotto. Isis, 75(3):589–590, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232977.

Enros:1984:BRB

[1254] Philip C. Enros. Book review: The Great Mental Calculators: The Psychology, Methods, and Lives of Calculating Prodigies, Past and Present by Steven B. Smith. Isis, 75(3):590–591, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232978.

Brown:1984:BRB

[1255] Laurie M. Brown. Book review: Storia della fisica: Un contributo per l’insegnamento della fisica by Fabio Bevilacqua. Isis, 75(3):591, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232979.

Williams:1984:BRBb

[1256] L. Pearce Williams. Book review: Ampere et la creation de l’electrodynamique (1820–1827) by Christine Blondel. Isis, 75(3):591– 592, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232980.

Stuewer:1984:BRBb

[1257] Roger H. Stuewer. Book review: Statistical Physics and the Atomic Theory of Matter from Boyle and Newton to Landau and Onsager by Stephen G. Brush. Isis, 75(3):592–593, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232981.

Smith:1984:BRBa

[1258] Robert W. Smith. Book review: Frame of the Universe: A History of Physical Cosmology by Frank Durham; Robert D. Purrington. Isis, 75(3):593, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232982.

Hackmann:1984:BRB

[1259] W. D. Hackmann. Book review: The Effluvial Theory of Electricity by Roderick Weir Home. Isis, 75(3):593–595, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232983.

Buerger:1984:BRB

[1260] Martin J. Buerger. Book review: Crystallography in North America by Dan McLachlan,; Jenny P. Glusker. Isis, 75(3):595, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232984.

Paul:1984:BRBb

[1261] E. Robert Paul. Book review: The Expanding Universe: Astronomy’s “Great Debate,” 1900–1930 by Robert W. Smith. Isis, 75(3):595–596, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232985.

Carroll:1984:BRB

[1262] P. Thomas Carroll. Book review: Roger Adams: Scientist and Statesman by D. Stanley Tarbell; Ann Tracy Tarbell. Isis, 75(3):596–598, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232986.

Rehbock:1984:BRB

[1263] Philip F. Rehbock. Book review: The Discovery of Fossil Fishes in Scotland up to 1845, with Checklists of Agassiz’s Figured Specimens by S. M. Andrews. Isis, 75(3):598, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232987.

Norton:1984:BRB

[1264] B. Norton. Book review: Evolution from Molecules to Men by D. S. Bendall. Isis, 75(3):598–599, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232988.

Pyne:1984:BRB

[1265] Stephen J. Pyne. Book review: This Land, This South: An Environmental History by Albert E. Cowdrey. Isis, 75(3):599–600, September 1984.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232989.

Egerton:1984:BRBb

[1266] Frank N. Egerton. Book review: Sacred Sands: The Struggle for Community in the Indiana Dunes by J. Ronald Engel. Isis, 75(3):600, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232990.

Glass:1984:BRB

[1267] Bentley Glass. Book review: A Feeling for the Organism: The Life and Work of Barbara McClintock by Evelyn Fox Keller. Isis, 75(3):600–601, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232991.

Mayr:1984:BRB

[1268] Ernst Mayr. Book review: Dear Lord Rothschild: Birds, Butterflies, and History by Miriam Rothschild. Isis, 75(3):601–603, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232992.

Lennox:1984:BRB

[1269] James G. Lennox. Book review: Nature Animated: Historical and Philosophical Case Studies in Greek Medicine, Nineteenth-Century and Recent Biology, Psychiatry, and Psychoanalysis by Michael Ruse. Isis, 75 (3):603–604, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232993.

Rosenberg:1984:BRB

[1270] Charles E. Rosenberg. Book review: The D.O.’s: Osteopathic Medicine in America by Norman Gevitz. Isis, 75(3):604–605, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232994.

Dwyer:1984:BRB

[1271] Ellen Dwyer. Book review: Mental Illness and American Society, 1875– 1940 by Gerald Grob. Isis, 75(3):605–606, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232995.

Grob:1984:BRB

[1272] Gerald N. Grob. Book review: Secret Passions, Secret Remedies: Narcotic Drugs in British Society, 1820–1930 by Terry M. Parssinen. Isis, 75 (3):606–607, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232996.

Burnham:1984:BRB

[1273] John C. Burnham. Book review: The Occult in America: New Historical Perspectives by Howard Kerr; Charles L. Crow. Isis, 75(3):607, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232997.

Preus:1984:BRB

[1274] Anthony Preus. Book review: Zenon d’Elée: Prolégomènes aux doctrines du continue: Étude historique et critique des fragments et témoinages by Maurice Caveing. Isis, 75(3):607–608, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232998.

Eamon:1984:BRB

[1275] William Eamon. Book review: Fachprosa-Studien: Beitrage zur mittelalterlichen Wissenschafts- und Geistesgeschichte by Gundolf Keil. Isis, 75 (3):608–609, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232999.

Savage-Smith:1984:BRBb

[1276] Emilie Savage-Smith. Book review: Kitab al-Manazir: Books I-II-III (On Direct Vision) by Al-Hasan ibn al-Haytham; Abdelhamid I. Sabra. Isis, 75(3):609–611, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233000.

Kennedy:1984:BRB

[1277] E. S. Kennedy. Book review: A Bibliography of the Works of Abu’l-Rayhan al-Biruni by Ahmad Saeed Khan. Isis, 75(3):611, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233001.

Berggren:1984:BRB

[1278] J. L. Berggren. Book review: Geschichte der Zahlentheorie im Orient: im Mittelalter und zu Beginn der Neuzeit unter besonder Berucksichtigung persischer Mathematiker by Alireza Djafari Naini. Isis, 75(3):611–612, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233002.

Rocher:1984:BRB

[1279] Ludo Rocher. Book review: The Indian Response to European Technology and Culture (A.D. 1498–1707) by Ahsan Jan Qaisar. Isis, 75(3):612–613, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233003.

Moran:1984:BRBa

[1280] Bruce T. Moran. Book review: Lay Culture, Learned Culture: Books and Social Change in Strasbourg, 1480–1599 by Miriam Usher Chrisman. Isis, 75(3):613–614, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233004.

Wisan:1984:BRB

[1281] Winifred Lovell Wisan. Book review: Cause, Experiment, and Science: A Galilean Dialogue Incorporating a New English Translation of Galileo’s “Bodies That Stay Atop Water, or Move in It.” by Stillman Drake. Isis, 75(3):614, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233005.

Ballester:1984:BRB

[1282] Luis Garcia Ballester. Book review: Spanish Scientists in the New World: The Eighteenth-Century Expeditions by Iris H. W. Engstrand. Isis, 75 (3):614–616, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233006.

Palter:1984:BRB

[1283] Robert Palter. Book review: Descartes: Philosophy, Mathematics and Physics by Stephen Gaukroger. Isis, 75(3):616–617, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233007.

Larson:1984:BRBb

[1284] James Larson. Book review: Svensk lardomshistoria: Gustavianska tiden by Sten Lindroth; Gunnar Eriksson. Isis, 75(3):617–618, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233008.

Calinger:1984:BRB

[1285] Ronald S. Calinger. Book review: Der junge Leibniz II: Der Ubergang vom Atomismus zu einem mechanistischen Aristotelismus; Der revidierte Anschluss an Pierre Gassendi by Konrad Moll. Isis, 75(3):618–619, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233009.

Baatz:1984:BRBb

[1286] Simon Baatz. Book review: Charles Willson Peale and His World by Edgar P. Richardson; Brooke Hindle; Lillian Miller. Isis, 75(3):619–620, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233010.

Perrin:1984:BRB

[1287] Carl Perrin. Book review: Joseph Black, 1728–1799: A Commemorative Symposium by A. D. C. Simpson. Isis, 75(3):620–621, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233011.

Howard:1984:BRB

[1288] Don A. Howard. Book review: Gesamtausgabe. Volume VIII: Ludwig Boltzmann: Internationale Tagung anlasslich des 75. Jahrestages seines Todes, 1981: Ausgewahlte Abhandlungen by Ludwig Boltzmann; Roman U. Sexl; John Blackmore. Isis, 75(3):621, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233012.

Bjork:1984:BRB

[1289] Daniel W. Bjork. Book review: William James on Exceptional Mental States: The 1896 Lowell Lectures by Eugene Taylor. Isis, 75(3):621–623, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233013.

DelSesto:1984:BRBb

[1290] Steven L. Del Sesto. Book review: Crisis Contained: The Department of Energy at Three Mile Island by Philip L. Cantelon; Robert C. Williams; Global Fission: The Battle over Nuclear Power by Jim Falk. Isis, 75 (3):623–625, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233014.

Aspray:1984:BRB

[1291] William Aspray. Book review: Alan Turing: The Enigma by Andrew Hodges. Isis, 75(3):625–626, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233015.

Roland:1984:BRBa

[1292] Alex Roland. Book review: The First 25 Years in Space: A Symposium by Allan A. Needell. Isis, 75(3):626–627, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233016.

Moyer:1984:BRB

[1293] Albert E. Moyer. Book review: Project Y: The Los Alamos Story by David S. Hawkins; Edith Truslow; Ralph Carlisle Smith. Isis, 75(3): 627–628, September 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233017.

Anonymous:1984:BMc

[1294] Anonymous. Back matter. Isis, 75(3):629–634, September 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233018.

Anonymous:1984:FMd

[1295] Anonymous. Front matter. Isis, 75(4):635–641, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232410.

Stranges:1984:FBR

[1296] Anthony N. Stranges. Friedrich Bergius and the rise of the German synthetic fuel industry. Isis, 75(4):642–667, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232411.

Cantor:1984:BAR

[1297] Geoffrey Cantor. Berkeley’s the analyst revisited. Isis, 75(4):668–683, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232412.

Weiner:1984:CES

[1298] Douglas R. Weiner. Community ecology in Stalin’s Russia: “socialist” and “bourgeois” science. Isis, 75(4):684–696, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232413.

Nye:1984:SDQ

[1299] Mary Jo Nye. Scientific decline: Is quantitative evaluation enough? Isis, 75(4):697–708, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232414.

dAlverny:1984:EML

[1300] M. T. d’Alverny. Eloge: Maria Luisa Righini Bonelli, 11 November 1917–18 December 1981. Isis, 75(4):709–711, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232415.

Kuklick:1984:BRB

[1301] Henrika Kuklick. Book review: Person and Myth: Maurice Leenhardt in the Melanesian World. Isis, 75(4):712–717, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232416.

Daston:1984:BRB

[1302] Lorraine J. Daston. Book review: The Nature of Mathematical Knowledge. Isis, 75(4):717–721, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232417.

Eastwood:1984:BRBb

[1303] Bruce Eastwood. Book review: The Soul of Science: An Audiovisual History of Science from Anaximander to Einstein by Bill Stonebarger; Michael Stonebarger. Isis, 75(4):722–723, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232418.

Tarbell:1984:BRB

[1304] D. Stanley Tarbell. Book review: The History of Chemical Technology: An Annotated Bibliography by Robert Multhauf. Isis, 75(4):723, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232419.

Leplin:1984:BRB

[1305] Jarrett Leplin. Book review: Testing Scientific Theories by John Earman. Isis, 75(4):723–724, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232420.

Lenoir:1984:BRB

[1306] Timothy Lenoir. Book review: Ludwik Fleck: Erfahrung und Tatsache: Gesammelte Aufsatze mit einer Einleitung by Lothar Shafer; Thomas Schnelle. Isis, 75(4):724–725, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232421.

Dales:1984:BRBb

[1307] Richard C. Dales. Book review: Time, Creation and the Continuum: Theories in Antiquity and the Early Middle Ages by Richard Sorabji. Isis, 75(4):725–726, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232422.

Todd:1984:BRB

[1308] Edmund N. Todd. Book review: Wissenschaft, Staat, Mazene: Anfange moderner Wissenschaftspolitik in Grossbritannien 1850–1920 by Peter Alter. Isis, 75(4):726–727, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232423.

Seidel:1984:BRB

[1309] Robert W. Seidel. Book review: Science, Technology, and Society in the Time of Alfred Nobel by Carl Gustaf Bernhard; Elisabeth Crawford; Per Sorbom. Isis, 75(4):727–728, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232424.

Morley:1984:BRB

[1310] Jane Morley. Book review: Books and Society in History by Kenneth E. Carpenter. Isis, 75(4):728–729, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232425.

Reingold:1984:BRB

[1311] Nathan Reingold. Book review: Professions and Professional Ideology in America by Gerald L. Geison. Isis, 75(4):729, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232426.

Schofield:1984:BRB

[1312] Robert E. Schofield. Book review: Metropolis and Province: Science in British Culture, 1780–1850 by Ian Inkster; Jack Morrell. Isis, 75 (4):729–730, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232427.

Segal:1984:BRB

[1313] Daniel Segal. Book review: The Social Relations of Physics, Mysticism, and Mathematics: Studies in Social Structure, Interests and Ideas by Sal Restivo. Isis, 75(4):730–732, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232428.

Catudal:1984:BRB

[1314] Jacques N. Catudal. Book review: Values and Scientists by John A. White. Isis, 75(4):732, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232429.

Ware:1984:BRB

[1315] R. Dean Ware. Book review: Gregorian Reform of the Calendar: Proceedings of the Vatican Conference to Commemorate Its 400th Anniversary, 1582–1982 by G. V. Coyne; M. A. Hoskin; O. Pedersen. Isis, 75 (4):732–733, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232430.

Hetherington:1984:BRB

[1316] Norriss S. Hetherington. Book review: Glimpsing an Invisible Universe: The Emergence of X-Ray Astronomy by Richard F. Hirsh. Isis, 75(4): 733–734, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232431.

Swenson:1984:BRB

[1317] Loyd S. Swenson, Jr. Book review: American Physics in Transition: A History of Conceptual Change in the Late Nineteenth Century by Albert E. Moyer. Isis, 75(4):734–735, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232432.

Saliba:1984:BRB

[1318] George Saliba. Book review: Astronomy and History: Selected Essays by O. Neugebauer. Isis, 75(4):735–736, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232433.

Swerdlow:1984:BRB

[1319] N. M. Swerdlow. Book review: Narratio prima by Georg Joachim Rheticus; Henri Hugonnard-Roche; Jean–Pierre Verdet; Michel-Pierre Lerner; Alain Segonds. Isis, 75(4):736–737, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232434.

Gerstner:1984:BRB

[1320] Patsy Gerstner. Book review: It Began with a Stone: A History of Geology from the Stone Age to the Age of Plate Tectonics by Henry Faul; Carol Faul. Isis, 75(4):738, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232435.

Dunbar:1984:BRBc

[1321] G. S. Dunbar. Book review: Sheets of Many Colours: The Mapping of Ireland’s Rocks, 1750–1890 by Gordon L. Herries Davies. Isis, 75 (4):738–739, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232436.

Kupsch:1984:BRB

[1322] W. O. Kupsch. Book review: Antarctica Observed: Who Discovered the Antarctic Continent? by A. G. E. Jones; Ross in the Antarctic: The Voyages of James Clark Ross in Her Majesty’s Ships “Erebus” and “Terror,” 1839–1843 by M. J. Ross. Isis, 75(4):739, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232437.

Sapp:1984:BRB

[1323] Jan Sapp. Book review: Histoire de la genetique et de l’evolutionisme en France by Denis Buican. Isis, 75(4):739–741, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232438.

Montgomery:1984:BRB

[1324] William Montgomery. Book review: Gametes & Spores: Ideas about Sexual Reproduction, 1750–1914 by John Farley. Isis, 75(4):741–742, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232439.

Morton:1984:BRB

[1325] A. G. Morton. Book review: Landmarks of Botanical History by Edward Lee Greene; Frank N. Egerton; Robert P. McIntosh; Rogers McVaugh.

Isis, 75(4):742–743, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232440.

Outram:1984:BRB

[1326] Dorinda Outram. Book review: Darwin in Italia: Impresa scientifica e frontiere culturali by Giuliano Pancaldi. Isis, 75(4):743–744, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232441.

Beddall:1984:BRB

[1327] Barbara G. Beddall. Book review: Darwin: La aventura de un espiritu by Desiderio Papp. Isis, 75(4):744–745, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232442.

Danzinger:1984:BRB

[1328] Kurt Danzinger. Book review: Kurt Koffka: An Unwitting Self-Portrait by Molly Harrower. Isis, 75(4):745, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232443.

Colley:1984:BRB

[1329] Linda Colley. Book review: Racial Myth in English History: Trojans, Teutons, and Anglo-Saxons by Hugh A. MacDougall. Isis, 75(4):745– 746, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232444.

Marcus:1984:BRB

[1330] George E. Marcus. Book review: Observers Observed: Essays on Ethnographic Fieldwork by George W. Stocking. Isis, 75(4):746, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232445.

Frey:1984:BRB

[1331] Emil F. Frey. Book review: The Physicians of Pharaonic Egypt by Paul Ghalioungui. Isis, 75(4):746–747, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232446.

Glick:1984:BRBb

[1332] Thomas F. Glick. Book review: Medicina e farmacia nella caricature politiche italiane, 1848–1914 by William H. Helfand; Sergio Rocchietta.

Isis, 75(4):747, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232447.

Fee:1984:BRB

[1333] Elizabeth Fee. Book review: Princes and Peasants: Smallpox in History by Donald R. Hopkins. Isis, 75(4):747–748, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232448.

Fye:1984:BRB

[1334] W. Bruce Fye. Book review: The Chief: Doctor William Osler by R. Palmer Howard. Isis, 75(4):748–749, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232449.

Gelfand:1984:BRBb

[1335] Toby Gelfand. Book review: Les maladies et la medicine en Pays basque nord a la fin de l’Ancien Regime (1690–1789) by Pierre L. Thillaud. Isis, 75(4):749, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232450.

Lang:1984:BRB

[1336] Helen S. Lang. Book review: Essays in Ancient Greek Philosophy. Volume II by John P. Anton; Anthony Preus. Isis, 75(4):750, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232451.

Scarborough:1984:BRBb

[1337] John Scarborough. Book review: Science, Folklore and Ideology: Studies in the Life Sciences in Ancient Greece by G. E. R. Lloyd. Isis, 75(4): 750–752, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232452.

Wardy:1984:BRB

[1338] Robert Wardy. Book review: Nature, Change and Agency in Aristotle’s “Physics”: A Philosophical Study by Sarah Waterlow. Isis, 75(4):752– 753, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232453.

Siraisi:1984:BRBb

[1339] Nancy G. Siraisi. Book review: The Black Death: Natural and Human Disaster in Medieval Europe by Robert S. Gottfried. Isis, 75(4):753–754, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232454.

Shea:1984:BRBb

[1340] William R. Shea. Book review: Vom Einfluss der Gestirne auf die Gesundheit und den Charakter des Menschen by Gundolf Keil; Freidrich Lenhardt; Christoph Weisser. Isis, 75(4):754–755, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232455.

Tachau:1984:BRB

[1341] Katherine H. Tachau. Book review: William of Auvergne and Robert Grosseteste: New Ideas of Truth in the Early Thirteenth Century by Steven P. Marrone. Isis, 75(4):755–756, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232456.

Peters:1984:BRB

[1342] Edward Peters. Book review: Les debuts de l’imagerie alchimique (XIVe-XVe siècles) by Barbara Obrist. Isis, 75(4):756–757, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232457.

Bullough:1984:BRB

[1343] Vern L. Bullough. Book review: Corps et chirurgie à l’apogée du Moyen Âge: Savoir et imaginaire du corps chez Henri de Mondeville, chirurgien de Philippe le Bel by Marie-Christine Pouchelle. Isis, 75(4):757–758, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232458.

Gingerich:1984:BRB

[1344] Owen Gingerich. Book review: Studies in the Islamic Exact Sciences by E. S. Kennedy; David King; Mary Helen Kennedy. Isis, 75(4):758, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232459.

DeYoung:1984:BRB

[1345] Gregg De Young. Book review: Agricultural Innovation in the Early Islamic World: The Diffusion of Crops and Farming Techniques, 700–1100 by Andrew M. Watson. Isis, 75(4):758–759, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232460.

Shumaker:1984:BRB

[1346] Wayne Shumaker. Book review: Astrology in the Renaissance: The Zodiac of Life by Eugenio Garin; Carolyn Jackson; June Allen; Clare Robertson. Isis, 75(4):759, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232461.

Henry:1984:BRB

[1347] John Henry. Book review: Thomas Harriot: A Biography by John W. Shirley. Isis, 75(4):759–760, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232462.

Reeds:1984:BRBb

[1348] Karen Reeds. Book review: La bibliothèque et le laboratoire de Guy de la Brosse au Jardin des Plantes a Paris by Rio Howard. Isis, 75(4): 760–761, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232463.

McMullin:1984:BRBb

[1349] Ernan McMullin. Book review: Certain Philosophical Questions: Newton’s Trinity Notebook by J. E. McGuire; Martin Tamny. Isis, 75(4): 761–762, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232464.

Rousseau:1984:BRB

[1350] G. S. Rousseau. Book review: Universal Languages and Scientific Taxonomy in the Seventeenth Century by M. M. Slaughter. Isis, 75(4):762–763, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232465.

Westfall:1984:BRB

[1351] Richard S. Westfall. Book review: From Paracelsus to Newton: Magic and the Making of Modern Science by Charles Webster. Isis, 75(4):763, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232466.

Smolka:1984:BRB

[1352] Josef Smolka. Book review: Zivot a dilo Vaclava Prokopa Divise. (The Life and Work of V. P. Divis.) by Josef Haubelt. Isis, 75(4):763–764, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232467.

Albury:1984:BRB

[1353] W. R. Albury. Book review: Linnaeus: The Man and His Work by Tore Frangsmyr; Michael Srigley; Bernard Vowles. Isis, 75(4):764–765, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232468.

Dougherty:1984:BRB

[1354] F. W. P. Dougherty. Book review: The Correspondence between Albrecht von Haller and Charles Bonnet by Otto Sonntag. Isis, 75(4):765–766, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232469.

Gillispie:1984:BRB

[1355] Charles C. Gillispie. Book review: Patrick Matthew and Natural Selection: Nineteenth Century Gentleman-Farmer, Naturalist and Writer by W. J. Dempster. Isis, 75(4):766–767, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232470.

Rocke:1984:BRBb

[1356] A. J. Rocke. Book review: Amedeo Avogadro: A Scientific Biography by Mario Morselli. Isis, 75(4):767–768, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232471.

Webb:1984:BRBb

[1357] George Webb. Book review: Realm of the Long Eyes: A Brief History of Kitt Peak National Observatory by James E. Kloeppel. Isis, 75(4): 768–769, December 1984. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232472.

Gillmor:1984:BRB

[1358] C. Stewart Gillmor. Book review: Origins of Magnetospheric Physics by James A. Van Allen. Isis, 75(4):769, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232473.

Kingsland:1984:BRB

[1359] Sharon Kingsland. Book review: The Ecological Century: A Personal Appraisal by E. Barton Worthington. Isis, 75(4):769–770, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232474.

Beaver:1984:BRB

[1360] Donald deB. Beaver. Book review: Turing’s Man: Western Culture in the Computer Age by David J. Bolter. Isis, 75(4):770–771, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232475.

Pursell:1984:BRBb

[1361] Carroll Pursell. Book review: The Culture of Technology by Arnold Pacey. Isis, 75(4):771–772, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232476.

Carlson:1984:BRB

[1362] W. Bernard Carlson. Book review: Explaining Technical Change: A Case Study in the Philosophy of Science by Jon Elster. Isis, 75(4):772–773, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232477.

Ihde:1984:BRB

[1363] Don Ihde. Book review: The Nature of Technological Change: Are Models of Scientific Change Relevant? by Rachel Laudan. Isis, 75(4):773–774, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232478.

Smith:1984:BRBb

[1364] Cecil O. Smith, Jr. Book review: Research in Philosophy and Technology. Volume 5, 1982 by Paul Durbin. Isis, 75(4):774–775, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232479.

McGaw:1984:BRB

[1365] Judith A. McGaw. Book review: More Work for Mother: The Ironies of Household Technology from the Open Hearth to the Microwave by Ruth Schwartz Cowan. Isis, 75(4):775–777, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232480.

Lubrano:1984:BRB

[1366] Linda L. Lubrano. Book review: Politics and Technology in the Soviet Union by Bruce Parrott. Isis, 75(4):777–778, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232481.

Constant:1984:BRB

[1367] Edward W. Constant, II. Book review: Inside the Black Box: Technology and Economics by Nathan Rosenberg. Isis, 75(4):778–779, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232482.

Ravetz:1984:BRB

[1368] Alison Ravetz. Book review: Machina ex Dea: Feminist Perspectives on Technology by Joan Rothschild. Isis, 75(4):779–780, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232483.

Kranakis:1984:BRB

[1369] Eda Kranakis. Book review: Dictionnaire des ingénieurs militaires, 1691–1791 by Anne Blanchard; Le Corps des Ponts et Chaussées by A. Brunot; R. Coquand. Isis, 75(4):780–781, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232484.

Lankton:1984:BRB

[1370] Larry D. Lankton. Book review: The Introduction of the Use of Mild Steel into the Shipbuilding and Marine Engine Industries by J. F. Clarke; F. Storr. Isis, 75(4):781–782, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232485.

Basalla:1984:BRB

[1371] George Basalla. Book review: Metropolitan Corridor: Railroads and the American Scene by John R. Stilgoe. Isis, 75(4):782–783, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232486.

Roland:1984:BRBb

[1372] Alex Roland. Book review: The Winged Gospel: America’s Romance with Aviation, 1900–1950 by Joseph Corn; Winged Wonders: The Story of the Flying Wings by E. T. Wooldridge. Isis, 75(4):783–784, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232487.

Trimble:1984:BRB

[1373] William F. Trimble. Book review: The Eagle Aloft: Two Centuries of the Balloon in America by Tom D. Crouch. Isis, 75(4):784–785, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232488.

Reynolds:1984:BRB

[1374] Terry S. Reynolds. Book review: Water Power in Scotland, 1550–1870 by John Shaw. Isis, 75(4):785–786, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232489.

Thomson:1984:BRB

[1375] Ron B. Thomson. Book review: Les temps de l’eau: La cite, l’eau et les techniques, nord de la France, fin IIIe-debut XIXe siècle by Andre Guillerme. Isis, 75(4):786, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232490.

Pershey:1984:BRB

[1376] Edward J. Pershey. Book review: Material Culture of the Wooden Age by Brooke Hindle. Isis, 75(4):787, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232491.

Hall:1984:BRBd

[1377] A. Rupert Hall. Book review: Revolution in Time: Clocks and the Making of the Modern World by David S. Landes. Isis, 75(4):787–788, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232492.

Landels:1984:BRB

[1378] J. G. Landels. Book review: The Earliest Wheeled Transport: From the Atlantic Coast to the Caspian Sea by Stuart Piggott. Isis, 75(4):788–789, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232493.

Knobloch:1984:BRB

[1379] Eberhard Knobloch. Book review: Corpus der italienischen Zeichnungen, 1300–1450. Volume I: Addenda zu Sudund Mittelitalien. Part 4: Katalog 717–719: Mariano Taccola by Bernhard Degenhardt; Annegrit Schmitt; Hans-Joachim Eberhardt. Isis, 75(4):789–790, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232494.

Moran:1984:BRBb

[1380] Bruce T. Moran. Book review: The Printing Revolution in Early Modern Europe by Elizabeth Eisenstein. Isis, 75(4):790–791, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232495.

Lewis:1984:BRB

[1381] W. David Lewis. Book review: Boss Kettering: Wizard of General Motors by Stuart W. Leslie. Isis, 75(4):791–792, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232496.

Anonymous:1984:BMd

[1382] Anonymous. Back matter. Isis, 75(4):793–797, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232497.

Anonymous:1984:VI

[1383] Anonymous. Volume information. Isis, 75(4):798–810, December 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232409.

Anonymous:1984:FMe

[1384] Anonymous. Front matter. Isis, 75(5):1–4, ???? 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232273.

Neu:1984:OHN

[1385] John Neu. One hundred ninth critical bibliography of the history of science and its cultural influences. Isis, 75(5):5–206, ???? 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232274.

Anonymous:1984:BMe

[1386] Anonymous. Back matter. Isis, 75(5):??, ???? 1984. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232275.

Anonymous:1985:FMa

[1387] Anonymous. Front matter. Isis, 76(1):1–5, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232788.

Elkana:1985:GEH

[1388] Yehuda Elkana and Michael Segre. Guest editorial: History of science in Israel. Isis, 76(1):6–10, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232789.

Westfall:1985:SPG

[1389] Richard S. Westfall. Science and patronage: Galileo and the telescope. Isis, 76(1):11–30, March 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232790.

Lederer:1985:HNL

[1390] Susan Eyrich Lederer. Hideyo Noguchi’s luetin experiment and the antivivisectionists. Isis, 76(1):31–48, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232791.

Coleman:1985:CBD

[1391] William Coleman. The cognitive basis of the discipline: Claude Bernard on physiology. Isis, 76(1):49–70, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232792.

Rahman:1985:STI

[1392] A. Rahman. Science and technology in Indian culture. Isis, 76(1):71– 72, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232793.

Gillispie:1985:EMD

[1393] Charles C. Gillispie. Eloge: Maurice Daumas, 1910–1984. Isis, 76(1): 72–74, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232794.

Hall:1985:BRB

[1394] A. Rupert Hall. Book review: Correspondance du P, by Marin Mersenne; Cornelis de Waard; Armand Beaulieu. Marin Mersenne, religieux minime. Isis, 76(1):75–80, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232795.

Lankford:1985:BRB

[1395] John Lankford. Book review: Cecilia Payne-Gaposchkin: An Autobiography and Other Recollections. Isis, 76(1):80–83, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232796.

Gieryn:1985:BRB

[1396] Thomas F. Gieryn. Book review: Functions and Uses of Disciplinary History by Loren Graham; Wolf Lepenies; Peter Weingart. Isis, 76(1): 84–85, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232797.

Nye:1985:BRBa

[1397] Mary Jo Nye. Book review: Sciences et techniques en perspective by Jean Dhombres. Isis, 76(1):85, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232798.

Beaulieu:1985:BRB

[1398] A. Beaulieu. Book review: The Collection Boulliau (BN, FF. 13019– 13059): An Inventory by Ismael Boulliau; Robert A. Hatch. Isis, 76(1): 85–86, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232799.

Kohlstedt:1985:BRBa

[1399] Sally Gregory Kohlstedt. Book review: Guide to the Smithsonian Archives. Isis, 76(1):86–87, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232800.

Lubrano:1985:BRB

[1400] Linda L. Lubrano. Book review: Empire of Knowledge: The Academy of Sciences of the USSR (1917–1970) by Alexander Vucinich. Isis, 76(1): 87–88, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232801.

Engel:1985:BRB

[1401] A. J. Engel. Book review: How the Ph.D. Came to Britain: A Century of Struggle for Post-graduate Education by Renate Simpson. Isis, 76(1): 88–89, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232802.

Duncan:1985:BRB

[1402] A. M. Duncan. Book review: Science and the English Enlightenment by Simon Schaffer. Isis, 76(1):89–90, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232803.

Blackwell:1985:BRB

[1403] Richard J. Blackwell. Book review: Culture and Cultural Entities: Toward a New Unity of Science by Joseph Margolis. Isis, 76(1):90–91, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232804.

Wise:1985:BRB

[1404] M. Norton Wise. Book review: Springs of Scientific Creativity: Essays on Founders of Modern Science by Rutherford Aris; H. Ted Davis; Roger H. Stuewer. Isis, 76(1):91–92, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232805.

Losee:1985:BRB

[1405] John P. Losee. Book review: Great Scientific Experiments: Twenty Experiments That Changed Our View of the World by Rom Harre. Isis, 76(1):92–93, March 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232806.

Jones:1985:BRBa

[1406] Greta Jones. Book review: Darwin’s Plots: Evolutionary Narrative in Darwin, George Eliot and Nineteenth-Century Fiction by Gillian Beer. Isis, 76(1):93–94, March 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232807.

Eamon:1985:BRBa

[1407] William Eamon. Book review: Pseudo-Aristotle: The Secret of Secrets– Sources and Influences by W. F. Ryan; Charles B. Schmitt. Isis, 76(1): 94–95, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232808.

Barber:1985:BRBa

[1408] Bernard Barber. Book review: Opening Pandora’s Box: A Sociological Analysis of Scientists’ Discourse by G. Nigel Gilbert; Michael Mulkay. Isis, 76(1):95–97, March 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232809.

Miller:1985:BRBa

[1409] David Philip Miller. Book review: Science Observed: Perspectives on the Social Study of Science by Karin Knorr-Cetina; Michael Mulkay. Isis, 76(1):97–98, March 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232810.

Paul:1985:BRB

[1410] H. W. Paul. Book review: Pour une prospective de la science: Recherche et technologie: Les enjeux de l’avenir by Pierre Papon. Isis, 76(1):98– 99, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232811.

Grabiner:1985:BRBa

[1411] Judith V. Grabiner. Book review: Classics of Mathematics by Ronald S. Calinger. Isis, 76(1):99–100, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232812.

Siegel:1985:BRB

[1412] Daniel M. Siegel. Book review: Maxwell on Saturn’s Rings by Stephen G. Brush; C. W. F. Everitt; Elizabeth Garber. Isis, 76(1):100, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232813.

Dick:1985:BRB

[1413] Steven J. Dick. Book review: Planetary, Lunar, and Solar Positions: New and Full Moons, A.D. 1650–1805 by Owen Gingerich; Barbara Welther. Isis, 76(1):101, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232814.

Schweber:1985:BRB

[1414] S. S. Schweber. Book review: The Birth of Particle Physics by Laurie M. Brown; Lillian Hoddeson. Isis, 76(1):101–102, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232815.

Brown:1985:BRBa

[1415] Laurie M. Brown and Robert Michaelson. Book review: The Infancy of Atomic Physics: Hercules in His Cradle by Alex Keller; The Question of the Atom from the Karlsruhe Congress to the First Solvay Conference, 1860–1911 by Mary Jo Nye. Isis, 76(1):102–104, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232816.

Stuewer:1985:BRB

[1416] Roger H. Stuewer. Book review: The Tiger and the Shark: Empirical Roots of Wave-Particle Dualism by Bruce R. Wheaton. Isis, 76(1):104– 105, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232817.

Kish:1985:BRBa

[1417] George Kish. Book review: The History of Geography: Translations of Some French and German Essays by Gary S. Dunbar. Isis, 76(1):105– 106, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232818.

Faul:1985:BRB

[1418] Carol Faul. Book review: Memoirs of an Unrepentant Field Geologist: A Candid Profile of Some Geologists and Their Science, 1921–1981 by F. J. Pettijohn. Isis, 76(1):106, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232819.

Taylor:1985:BRB

[1419] Kenneth L. Taylor. Book review: The Great Chain of History: William Buckland and the English School of Geology (1814–1849) by Nicolas A. Rupke. Isis, 76(1):106–108, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232820.

Lyon:1985:BRB

[1420] John Lyon. Book review: The Ethics of Environmental Concern by Robin Attfield. Isis, 76(1):108, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232821.

Landauer:1985:BRB

[1421] Lyndall Landauer. Book review: Marshes of the Ocean Shore: Development of an Ecological Ethic by Joseph V. Siry. Isis, 76(1):108–109, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232822.

Samelson:1985:BRB

[1422] Franz Samelson. Book review: The Assault on Truth: Freud’s Suppression of the Seduction Theory by Jeffrey Moussaieff Masson. Isis, 76(1): 109–110, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232823.

Blustein:1985:BRB

[1423] Bonnie E. Blustein. Book review: Revue Internationale d’Histoire de la Psychiatrie by J. Arveiller. Isis, 76(1):110, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232824.

Castle:1985:BRB

[1424] Peter W. Castle. Book review: In Freud’s Shadow: Adler in Context by Paul E. Stepansky. Isis, 76(1):110–111, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232825.

Davis:1985:BRB

[1425] Audrey Davis. Book review: Hospital with a Heart: Women Doctors and the Paradox of Separatism at the New England Hospital, 1862–1969 by Virginia G. Drachman. Isis, 76(1):111–112, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232826.

Jones:1985:BRBb

[1426] Daniel P. Jones. Book review: Women and Health in America by Judith Walzer Leavitt. Isis, 76(1):112–113, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232827.

Brieger:1985:BRBa

[1427] Gert H. Brieger. Book review: The Structure of American Medical Practice, 1875–1944 by George Rosen; Charles E. Rosenberg. Isis, 76(1): 113, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232828.

Bolter:1985:BRB

[1428] J. David Bolter. Book review: The Computer Comes of Age: The People, the Hardware, and the Software by Rene Moreau; J. Howlett; Engines of the Mind: A History of the Computer by Joel Shurkin. Isis, 76(1):113– 115, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232829.

Ceruzzi:1985:BRB

[1429] Paul E. Ceruzzi. Book review: Memories That Shaped an Industry: Decisions Leading to IBM System/360 by Emerson W. Pugh. Isis, 76 (1):115, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232830.

McKim:1985:BRB

[1430] Richard McKim. Book review: Lucretius and Epicurus by Diskin Clay. Isis, 76(1):115–116, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232831.

Rubin:1985:BRB

[1431] Jody Rubin. Book review: Les maladies à l’aube de la civilisation occidentale: Researches sur la realité pathologique dans le monde grec préhistorique, archaique et classique by Mirko D. Grmek. Isis, 76(1): 116–117, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232832.

Goldstein:1985:BRB

[1432] Bernard R. Goldstein. Book review: Ptolemy’s Almagest by Ptolemy; G. J. Toomer. Isis, 76(1):117–118, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232833.

Scarborough:1985:BRBa

[1433] John Scarborough. Book review: Die Schrift des Rufus von Ephesos über die Gelbsucht in arabischer und lateinischer Ubersetzung by Manfred Ullmann. Isis, 76(1):118, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232834.

Waldron:1985:BRB

[1434] H. A. Waldron. Book review: Lead and Lead Poisoning in Antiquity by Jerome O. Nriagu. Isis, 76(1):118–120, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232835.

Eastwood:1985:BRBa

[1435] Bruce Eastwood. Book review: Hexaemeron by Robert Grosseteste; Richard C. Dales; Servus Gieben. Isis, 76(1):120–121, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232836.

Laird:1985:BRB

[1436] W. R. Laird. Book review: Tempo e spazio nel Medioevo by Massimo Parodi. Isis, 76(1):121, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232837.

Livesey:1985:BRB

[1437] Steven J. Livesey. Book review: The Seven Liberal Arts in the Middle Ages by David L. Wagner. Isis, 76(1):121–123, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232838.

Kealey:1985:BRB

[1438] Edward J. Kealey. Book review: A Latin Technical Phlebotomy and Its Middle English Translation by Linda E. Voights; Michael R. McVaugh. Isis, 76(1):123, March 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232839.

Ragep:1985:BRB

[1439] F. Jamil Ragep. Book review: The Book of the Reasons behind Astronomical Tables (Kitab fi ilal al-zijat): A Facsimile Reproduction of the Unique Arabic Text Contained in the Bodleian MS Arch. Seld. A.11 by Ali ibn Sulayman al-Hashimi; Fuad I. Haddad; E. S. Kennedy. Isis, 76(1):123–124, March 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232840.

Savage-Smith:1985:BRB

[1440] Emilie Savage-Smith. Book review: Islamic Cosmology: A Study of as-Suyuti’s al-Hay a as-saniya fi l-hay a as-sunniya with Critical Edition, Translation, and Commentary by Anton M. Heinen. Isis, 76(1):124– 125, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232841.

Deane:1985:BRB

[1441] Thatcher E. Deane. Book review: Les premières théories planétaires chinoises by Michel Teboul. Isis, 76(1):126, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232842.

Guerrini:1985:BRB

[1442] Anita Guerrini. Book review: Medicina e filosofia tra ’500 e ’600 by Giancarlo Zanier. Isis, 76(1):127, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232843.

Rousseau:1985:BRBa

[1443] G. S. Rousseau. Book review: Boerhaave’s Orations by Hermann Boerhaave; E. Kegel-Brinkgreve; A. M. Luyendijk-Elshout. Isis, 76(1):127– 128, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232844.

Jacob:1985:BRB

[1444] Margaret C. Jacob. Book review: The Apocalyptic Politics of Richard Price and Joseph Priestley: A Study in Late Eighteenth-Century English Republican Millennialism by Jack Fruchtman, Jr. Isis, 76(1):128, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232845.

Warner:1985:BRB

[1445] Deborah Jean Warner. Book review: At the Sign of the Compass and Quadrant: The Life and Times of Anthony Lamb by Silvio A. Bedini. Isis, 76(1):128–129, March 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232846.

Turner:1985:BRBa

[1446] Frank M. Turner. Book review: Science, God, and Nature in Victorian Canada by Carl Berger. Isis, 76(1):129, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232847.

Knight:1985:BRBa

[1447] David Knight. Book review: Chemical Atomism in the Nineteenth Century by Alan J. Rocke. Isis, 76(1):129–130, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232848.

Bowler:1985:BRBa

[1448] Peter J. Bowler. Book review: Hugh Miller: Outrage and Order — a Biography and Selected Writings by George Rosie. Isis, 76(1):130– 131, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232849.

Sokal:1985:BRB

[1449] Michael M. Sokal. Book review: A Matter of Consequences: Part Three of an Autobiography by B. F. Skinner. Isis, 76(1):131–132, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232850.

Trenn:1985:BRB

[1450] Thaddeus J. Trenn. Book review: Rutherford: Simple Genius by David Wilson. Isis, 76(1):132–133, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232851.

Anonymous:1985:BMa

[1451] Anonymous. Back matter. Isis, 76(1):134–138, March 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232852.

Anonymous:1985:FMb

[1452] Anonymous. Front matter. Isis, 76(2):139–143, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231743.

Dear:1985:TVR

[1453] Peter Dear. Totius in Verba: Rhetoric and authority in the early Royal Society. Isis, 76(2):144–161, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231744.

Kohler:1985:INS

[1454] Robert E. Kohler. Innovation in normal science: Bacterial physiology. Isis, 76(2):162–181, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231745.

Owens:1985:PSG

[1455] Larry Owens. Pure and sound government: Laboratories, playing fields, and gymnasia in the Nineteenth-Century search for order. Isis, 76(2): 182–194, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231746.

Melhado:1985:CPC

[1456] Evan M. Melhado. Chemistry, physics, and the chemical revolution. Isis, 76(2):195–211, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231747.

Kitts:1985:AMH

[1457] David B. Kitts, Mary Jo Nye, Audrey B. Davis, and Edith Sylla. Annual meeting of the History of Science Society 27–30 December 1984. Isis, 76(2):212–217, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231748.

Anonymous:1985:DD

[1458] Anonymous. Doctoral dissertations, 1982–1984. Isis, 76(2):218–219, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231749.

Anonymous:1985:DSF

[1459] Anonymous. Donors to and supporters of the fund drive. Isis, 76(2): 219–221, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231750.

Burnham:1985:BRB

[1460] John C. Burnham. Book review: American Science in the Age of Jefferson. Isis, 76(2):222–224, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231751.

Molland:1985:BRB

[1461] A. George Molland. Book review: The Discoverers: A History of Man’s Search to Know His World and Himself. Isis, 76(2):224–227, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231752.

Vucinich:1985:BRB

[1462] Alexander Vucinich. Book review: Povijest egzaktnih znanosti u Hrvata [History of Exact Sciences in Croatia] by Zarko Dadic. Isis, 76(2):228– 229, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231753.

Mauskopf:1985:BRB

[1463] Seymour H. Mauskopf. Book review: Science under Scrutiny: The Place of History and Philosophy of Science by R. W. Home. Isis, 76(2):229– 230, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231754.

Sheets-Pyenson:1985:BRB

[1464] Susan Sheets-Pyenson. Book review: Critical Issues in the History of Canadian Science, Technology, and Medicine by Richard A. Jarrell; Arnold E. Roos. Isis, 76(2):230, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231755.

Knight:1985:BRBb

[1465] David Knight. Book review: Nineteenth Century Short Title Catalogue. Series I. Phase I: 1801–1815 by Gwen Averley. Isis, 76(2):230–231, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231756.

Swerdlow:1985:BRB

[1466] N. M. Swerdlow. Book review: Scientific Writings and Astronomical Tables in Cracow: A Census of Manuscript Sources (XIVth-XVIth Centuries) by Grazyna Rosińska. Isis, 76(2):231, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231757.

Milton:1985:BRB

[1467] J. R. Milton. Book review: Hegel and the Sciences by Robert S. Cohen; Marx W. Wartofsky. Isis, 76(2):232–233, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231758.

Nordmann:1985:BRB

[1468] Alfred Nordmann. Book review: Edgar Zilsel und die Einheit der Erkenntnis by Johann Dvorak. Isis, 76(2):233, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231759.

Heyd:1985:BRB

[1469] Michael Heyd. Book review: Atom und Individuum im Zeitalter Newtons: Zur Genese der mechanistischen Natur-und Sozialphilosophie by Gideon Freudenthal. Isis, 76(2):233–234, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231760.

Phillips:1985:BRB

[1470] Joanne H. Phillips. Book review: Medizinische Instrumente aus Sepulkralfunden der romischen Kaiserzeit by Ernst Kunzl; Franz Josef Hassel; Susanna Kunzl. Isis, 76(2):234–235, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231761.

Hansen:1985:BRBa

[1471] Bert Hansen. Book review: Les Instruments astronomiques du Moyen Age by Emmanuel Poulle. Isis, 76(2):235–236, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231762.

McKnight:1985:BRB

[1472] John L. McKnight. Book review: Nineteenth-Century Scientific Instruments by Gerard L’E. Turner. Isis, 76(2):236–237, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231763.

Miller:1985:BRBb

[1473] David Philip Miller. Book review: From Clergyman to Don: The Rise of the Academic Profession in Nineteenth-Century Oxford by A. J. Engel. Isis, 76(2):237–238, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231764.

Pegg:1985:BRB

[1474] Barry Pegg. Book review: The Mathematicians’ Apprenticeship: Science, Universities, and Society in England, 1561–1640 by Mordechai Feingold. Isis, 76(2):238–239, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231765.

Ackerknecht:1985:BRB

[1475] Erwin H. Ackerknecht. Book review: Das Medizinisch-Chirurgische Institut in Zurich, 1782–1833: Vorlaufer der Medizinischen Fakultat der Universitat Zurich by Moritz Leisibach. Isis, 76(2):239–240, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231766.

Nye:1985:BRBb

[1476] Mary Jo Nye. Book review: Professions and the French State, 1700–1900 by Gerald Geison. Isis, 76(2):240–241, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231767.

Latour:1985:BRB

[1477] Bruno Latour. Book review: Rationality and Relativism by Martin Hollis; Steven Lukes. Isis, 76(2):241–242, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231768.

Schweikart:1985:BRB

[1478] Larry Schweikart. Book review: Science in Government by J. Ronayne. Isis, 76(2):242–243, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231769.

Greenblatt:1985:BRB

[1479] Samuel H. Greenblatt. Book review: The Culture of Biomedicine by D. Heyward Brock. Isis, 76(2):243, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231770.

Shapiro:1985:BRB

[1480] Barbara Shapiro. Book review: Probability and Literary Form: Philosophic Theory and Literary Practice in the Augustan Age by Douglas Lane Patey. Isis, 76(2):243–244, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231771.

Auerbach:1985:BRB

[1481] Nina Auerbach. Book review: George Eliot and Nineteenth-Century Science: The Make-Believe of a Beginning by Sally Shuttleworth. Isis, 76 (2):244–245, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231772.

Stavroudis:1985:BRB

[1482] Orestes N. Stavroudis. Book review: Optics after Newton: Theories of Light in Britain and Ireland, 1704–1840 by G. N. Cantor. Isis, 76(2): 245–246, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231773.

Norberg:1985:BRB

[1483] Arthur L. Norberg. Book review: Mercury’s Perihelion from Le Verrier to Einstein by N. T. Roseveare. Isis, 76(2):246–247, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231774.

DeVorkin:1985:BRB

[1484] David H. DeVorkin. Book review: The Early Years of Radio Astronomy: Reflections Fifty Years After Jansky’s Discovery by W. T. Sullivan III. Isis, 76(2):247–249, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231775.

Grattan-Guinness:1985:BRB

[1485] I. Grattan-Guinness. Book review: Erdgestalt, Kosmologie und Weltanschauung: Die Geschichte der Geodasie als Teil der Kulturgeschichte der Menschheit by Volker Bialas. Isis, 76(2):249, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231776.

Schultz:1985:BRB

[1486] Susan Schultz. Book review: Revolution in the Earth Sciences: Advances in the Past Half-Century by Shelby J. Boardman. Isis, 76(2):249–250, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231777.

Davies:1985:BRBa

[1487] Gordon L. Herries Davies. Book review: Krakatau 1883: The Volcanic Eruption and Its Effects by Tom Simkin; Richard S. Fiske; Sarah Melcher; Elizabeth Nielson. Isis, 76(2):250–251, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231778.

Cohen:1985:BRBa

[1488] Seymour S. Cohen. Book review: L’esprit des proteines: Histoire et philosophie biochimiques by Claude Debru. Isis, 76(2):251–252, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231779.

Winsor:1985:BRB

[1489] Mary P. Winsor. Book review: The Philosophical Naturalists: Themes in Early Nineteenth-Century British Biology by Philip R. Rehbock. Isis, 76(2):252–253, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231780.

Mabee:1985:BRB

[1490] Carleton Mabee and Susan Mabee Newhouse. Book review: With a Daughter’s Eye: A Memoir of Margaret Mead and Gregory Bateson by Mary Catherine Bateson. Isis, 76(2):253–255, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231781.

Danziger:1985:BRBa

[1491] Kurt Danziger. Book review: Explorations in the History of Psychology in the United States by Josef Brožek. Isis, 76(2):255–256, June 1985.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231782.

Stocking:1985:BRB

[1492] George W. Stocking, Jr. Book review: Essays in the History of Linguistic Anthropology by Dell H. Hymes. Isis, 76(2):256–257, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231783.

Avrith:1985:BRB

[1493] Gale Avrith. Book review: David Boyle: From Artisan to Archaeologist by Gerald Killan. Isis, 76(2):257–258, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231784.

Ospovat:1985:BRBa

[1494] Alexander M. Ospovat. Book review: Those Who Came Before: Southwestern Archeology in the National Park System by Robert H. Lister; Florence C. Lister. Isis, 76(2):258, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231785.

Clifford:1985:BRB

[1495] James Clifford. Book review: Histoires de l’anthropologie: XVIe-XIXe siècles by Britta Rupp-Eisenreich. Isis, 76(2):258–259, June 1985. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231786.

Ramsey:1985:BRB

[1496] Matthew Ramsey. Book review: Le chaudron et la lancette: Croyances populaires et medecine preventive, 1798–1830 by Yves-Marie Berce. Isis, 76(2):259, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231787.

LaBerge:1985:BRB

[1497] Ann F. La Berge. Book review: Dr. Richard Bright, 1789–1858 by Pamela Bright. Isis, 76(2):259–260, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231788.

vanKeuren:1985:BRB

[1498] David van Keuren. Book review: African Fever: A Study of British Science, Technology, and Politics in West Africa, 1787–1864 by Dennis G. Carlson. Isis, 76(2):260–261, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231789.

Lugg:1985:BRBa

[1499] Andrew Lugg. Book review: Philosophy and Technology by Paul T. Durbin; Friedrich Rapp. Isis, 76(2):261, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231790.

Jarrell:1985:BRB

[1500] Richard A. Jarrell. Book review: Mechanical Engineering at the National Research Council of Canada, 1929–1951 by W. E. Knowles Middleton. Isis, 76(2):261–262, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231791.

McMahon:1985:BRB

[1501] A. Michal McMahon. Book review: Great Waters: A History of Boston’s Water Supply by Fern L. Nesson. Isis, 76(2):262, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231792.

McCleary:1985:BRB

[1502] Peter McCleary. Book review: Architecture and the Crisis of Modern Science by Alberto Perez-Gomez. Isis, 76(2):262–263, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231793.

Glick:1985:BRBa

[1503] Thomas F. Glick. Book review: Problemi d’acque a Bologna in eta moderna. Isis, 76(2):263–264, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231794.

Robinson:1985:BRB

[1504] Lisa Mae Robinson. Book review: Lammot du Pont and the American Explosives Industry, 1850–1884 by Norman B. Wilkinson. Isis, 76(2): 264–265, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231795.

DeLacy:1985:BRB

[1505] Philip De Lacy. Book review: Galen on Respiration and the Arteries: An Edition with English Translation and Commentary of “De usu respirationis,” “An in arteriis natura sanguis contineatur,” “De usu pulsuum,” and “De causis respirationis.” by Galen; David J. Furley; J. S. Wilkie. Isis, 76(2):265–266, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231796.

Preus:1985:BRB

[1506] Anthony Preus. Book review: Le triangle hippocratique dans le monde greco-romain: Le malade, sa maladie et son medecin by Danielle Gourevitch. Isis, 76(2):266, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231797.

Scarborough:1985:BRBb

[1507] John Scarborough. Book review: A Commentary on the Prognosticon of Hippocrates by Stephanus the Philosopher; John M. Duffy. Isis, 76(2): 266–267, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231798.

Clagett:1985:BRBa

[1508] Marshall Clagett. Book review: The First Translation of Euclid’s Elements, Commonly Ascribed to Adelard of Bath. Books I–VIII and Books X.36–XV.2 by Euclid; H. L. L. Busard. Isis, 76(2):267–268, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231799.

Weisheipl:1985:BRB

[1509] James A. Weisheipl. Book review: Studies in the History of Medieval Optics by David C. Lindberg. Isis, 76(2):268–270, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231800.

Ovitt:1985:BRBa

[1510] George Ovitt, Jr. Book review: Approaches to Nature in the Middle Ages by Lawrence D. Roberts. Isis, 76(2):270–271, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231801.

Cunningham:1985:BRB

[1511] Andrew Cunningham. Book review: Fracastoro’s Syphilis by Girolamo Fracastoro; Geoffrey Eatough. Isis, 76(2):271, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231802.

Lamb:1985:BRB

[1512] Ursula Lamb. Book review: Sir Francis Drake and the Famous Voyage, 1577–1580: Essays Commemorating the Quadricentennial of Drake’s Circumnavigation of the Earth by Norman J. W. Thrower. Isis, 76(2): 271–272, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231803.

Murray:1985:BRBa

[1513] T. J. Murray. Book review: The Neuroanatomy of Leonardo da Vinci by Edwin M. Todd. Isis, 76(2):272–273, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231804.

Hackmann:1985:BRB

[1514] W. D. Hackmann. Book review: Isaac Beeckman (1588–1637) en de Mechanisering van het Wereldbeeld by Klaas van Berkel. Isis, 76(2):273– 274, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231805.

Shapin:1985:BRB

[1515] Steven Shapin. Book review: The Uses of Science in the Age of Newton by John G. Burke. Isis, 76(2):274–275, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231806.

Merchant:1985:BRB

[1516] Carolyn Merchant. Book review: The Principles of the Most Ancient and Modern Philosophy by Anne Conway; Peter Loptson. Isis, 76(2): 275–276, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231807.

Dooley:1985:BRB

[1517] Brendan Dooley. Book review: Scienza, natura, religione: Mondo newtoniano e cultura italiana nel primo settecento by Vincenzo Ferrone. Isis, 76(2):276–277, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231808.

Field:1985:BRB

[1518] J. V. Field. Book review: The Secret of the Universe: Mysterium cosmographicum by Johannes Kepler; A. M. Duncan. Isis, 76(2):277–278, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231809.

OHara:1985:BRB

[1519] J. G. O’Hara. Book review: Die newtonsche Kraft: Metamorphose eines wissenschaftlichen Begriffs by Werner Kutschmann. Isis, 76(2):278–279, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231810.

Harth:1985:BRB

[1520] Phillip Harth. Book review: Probability and Certainty in Seventeenth-Century England: A Study of the Relationships between Natural Science, Religion, History, Law, and Literature by Barbara J. Shapiro. Isis, 76(2): 279–280, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231811.

Jones:1985:BRBc

[1521] Greta Jones. Book review: Benjamin Kidd: Portrait of a Social Darwinist by D. P. Crook. Isis, 76(2):280–281, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231812.

Gooding:1985:BRBa

[1522] David Gooding. Book review: The Demon in the Aether: The Story of James Clerk Maxwell by Martin Goldman. Isis, 76(2):281, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231813.

Kremer:1985:BRB

[1523] Richard Lynn Kremer. Book review: Über die Erhaltung der Kraft by Hermann von Helmholtz; Christa Kirsten. Isis, 76(2):281–282, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231814.

Marcus:1985:BRB

[1524] Alan I. Marcus. Book review: Boussingault: Chemist and Agriculturist by F. W. J. McCosh. Isis, 76(2):282–283, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231815.

Weinstock:1985:BRB

[1525] Robert Weinstock. Book review: Inner Exile: Recollections of a Life with Werner Heisenberg by Elisabeth Heisenberg; S. Cappellari; C. Morris. Isis, 76(2):283, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231816.

Ziman:1985:BRB

[1526] John Ziman. Book review: Cambridge Physics in the Thirties by John Hendry; The National Physical Laboratory: A History by Edward Pyatt. Isis, 76(2):283–284, June 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231817.

McGraw:1985:BRB

[1527] Donald J. McGraw. Book review: Alexander Fleming: The Man and the Myth by Gwyn Macfarlane. Isis, 76(2):284–285, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231818.

Ludmerer:1985:BRB

[1528] Kenneth M. Ludmerer. Book review: Black Apollo of Science: The Life of Ernest Everett Just by Kenneth R. Manning. Isis, 76(2):285–286, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231819.

Wyden:1985:BRB

[1529] Peter Wyden. Book review: The Day the Sun Rose Twice: The Story of the Trinity Site Nuclear Explosion, July 16, 1945 by Ferenc Morton Szasz; The American Atom: A Documentary History of Nuclear Policies from the Discovery of Fission to the Present, 1939–1984 by Robert C. Williams; Philip L. Cantelon. Isis, 76(2):286–288, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231820.

Ihde:1985:BRB

[1530] Aaron J. Ihde. Book review: A Time to Remember: The Autobiography of a Chemist by Alexander Todd. Isis, 76(2):288, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231821.

Anonymous:1985:BMb

[1531] Anonymous. Back matter. Isis, 76(2):289–292, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231822.

Anonymous:1985:FMc

[1532] Anonymous. Front matter. Isis, 76(3):293–298, September 1985. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232853.

Kargon:1985:KCI

[1533] Robert Kargon and Elizabeth Hodes. Karl Compton, Isaiah Bowman, and the politics of science in the Great Depression. Isis, 76(3):300–318, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232854.

Riddle:1985:AMC

[1534] John M. Riddle. Ancient and Medieval chemotherapy for cancer. Isis, 76 (3):319–330, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232855.

Williams:1985:AIN

[1535] Elizabeth A. Williams. Anthropological institutions in Nineteenth-Century France. Isis, 76(3):331–348, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232856.

McGuire:1985:NAA

[1536] J. E. McGuire, Martin Tamny, and Newton. Newton’s astronomical apprenticeship: Notes of 1664/5. Isis, 76(3):349–365, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232857.

Peishan:1985:HMS

[1537] Li Peishan. History of modern science and technology in the People’s Republic of China. Isis, 76(3):366–370, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232858.

Beaver:1985:EDJ

[1538] Donald deB. Beaver. Eloge: Derek John deSolla Price (22 January 1922–3 September 1983). Isis, 76(3):371–374, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232859.

Numbers:1985:BRBa

[1539] Ronald L. Numbers. Book review: Did the Devil Make Darwin Do It? Modern Perspectives on the Creation-Evolution Controversy. Isis, 76(3): 375–377, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232860.

Edge:1985:BRB

[1540] David Edge. Book review: Metaphoric Process: The Creation of Scientific and Religious Understanding by Mary Gerhart; Allan Melvin Russell. Isis, 76(3):377–378, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232861.

Eastwood:1985:BRBb

[1541] Bruce Eastwood. Book review: In Defense of the Earth’s Centrality and Immobility: Scholastic Reaction to Copernicanism in the Seventeenth Century by Edward Grant. Isis, 76(3):378–379, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232862.

Wallace:1985:BRB

[1542] William A. Wallace. Book review: Galileo eretico by Pietro Redondi. Isis, 76(3):379–380, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232863.

Finocchiaro:1985:BRBa

[1543] Maurice A. Finocchiaro. Book review: I documenti del processo di Galileo Galilei by Sergio M. Pagano; Antonio G. Luciani. Isis, 76(3):380–381, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232864.

Moore:1985:BRBa

[1544] James R. Moore. Book review: Victorian Science and Religion: A Bibliography with Emphasis on Evolution, Belief, and Unbelief, Comprised of Works Published from ca. 1900 to 1975 by Sydney Eisen; Bernard V. Lightman. Isis, 76(3):381, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232865.

Wells:1985:BRBa

[1545] Kentwood D. Wells. Book review: The Establishment of Human Antiquity by Donald K. Grayson. Isis, 76(3):382, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232866.

Turner:1985:BRBb

[1546] Frank M. Turner. Book review: Science and Religion in the Nineteenth Century by Tess Cosslett. Isis, 76(3):382–383, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232867.

Kaplan:1985:BRB

[1547] Barbara Beigun Kaplan. Book review: From Aristotle to Darwin and Back Again: A Journey in Final Causality, Species, and Evolution by Etienne Gilson; John Lyon. Isis, 76(3):383, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232868.

Beatty:1985:BRB

[1548] John Beatty. Book review: Evolution: The History of an Idea by Peter J. Bowler. Isis, 76(3):383–384, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232869.

Landes:1985:BRB

[1549] David S. Landes. Book review: From the American System to Mass Production, 1800–1932: The Development of Manufacturing Technology in the United States. Isis, 76(3):385–388, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232870.

Heilbron:1985:BRB

[1550] J. L. Heilbron. Book review: The Historical Development of Quantum Theory. Isis, 76(3):388–393, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232871.

Eamon:1985:BRBb

[1551] William Eamon. Book review: Occult and Scientific Mentalities in the Renaissance. Isis, 76(3):393–396, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232872.

Lindberg:1985:BRB

[1552] David C. Lindberg. Book review: Album of Science: Antiquity and the Middle Ages by John E. Murdoch. Isis, 76(3):397–398, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232873.

Glick:1985:BRBb

[1553] Thomas F. Glick. Book review: The New History: The 1980’s and Beyond by Theodore K. Rabb; Robert I. Rotberg. Isis, 76(3):398–399, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232874.

Sanchez-Ron:1985:BRB

[1554] Jose M. Sanchez-Ron. Book review: The History of Modern Physics: An International Bibliography by Stephen G. Brush; Lanfranco Belloni; An Inventory of Published Letters to and from Physicists, 1900–1950 by Bruce R. Wheaton; J. L. Heilbron. Isis, 76(3):399–400, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232875.

Rechenberg:1985:BRB

[1555] Helmut Rechenberg. Book review: Werner Heisenberg: A Bibliography of His Writings by David Cassidy; Martha Baker. Isis, 76(3):400–401, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232876.

Fee:1985:BRB

[1556] Elizabeth Fee. Book review: Illustrated Catalogue of the Slide Archive of Historical Medical Photographs at Stony Brook by Center for Photographic Images of Medicine and Health Care; Rima D. Apple. Isis, 76(3):401, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232877.

Moravcsik:1985:BRB

[1557] Michael J. Moravcsik. Book review: Essays of an Information Scientist by Eugene Garfield. Isis, 76(3):401–402, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232878.

Pycior:1985:BRB

[1558] Helena Pycior. Book review: Carl Friedrich Gauss: A Bibliography by Uta C. Merzbach. Isis, 76(3):402, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232879.

Beckinsale:1985:BRB

[1559] Robert P. Beckinsale. Book review: Geology Emerging: A Catalog Illustrating the History of Geology (1500–1850) from a Collection in the Library of the University of Illinois at Urbana-Champaign by Dederick C. Ward; Albert V. Carozzi. Isis, 76(3):402–403, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232880.

Lugg:1985:BRBb

[1560] Andrew Lugg. Book review: Science and Values: The Aims of Science and Their Role in Scientific Debate by Larry Laudan. Isis, 76(3):403–404, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232881.

Reed:1985:BRB

[1561] Edward S. Reed. Book review: Reason and the Search for Knowledge: Investigations in the Philosophy of Science by Dudley Shapere. Isis, 76 (3):404–405, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232882.

Kohlstedt:1985:BRBb

[1562] Sally Gregory Kohlstedt. Book review: America’s Castle: The Evolution of the Smithsonian Building and Its Institution, 1840–1878 by Kenneth Hafertepe. Isis, 76(3):405–406, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232883.

Feingold:1985:BRBa

[1563] Mordechai Feingold. Book review: The Aristotelian Tradition and Renaissance Universities by Charles B. Schmitt. Isis, 76(3):406–407, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232884.

Oldroyd:1985:BRB

[1564] David R. Oldroyd. Book review: Social Darwinism in France by Linda L. Clark. Isis, 76(3):407–409, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232885.

Ravetz:1985:BRB

[1565] Jerome R. Ravetz. Book review: Polish Contributions to the Science of Science by Bohdan Walentynowicz. Isis, 76(3):409, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232886.

Hughes:1985:BRB

[1566] Thomas P. Hughes. Book review: Industriekultur: Peter Behrens and the AEG, 1907–1914 by Tilmann Buddensieg; Henning Rogge; Gabriele Heidecker; Karin Wilhelm; Sabine Bohle; Fritz Neumeyer; Iain Boyd Whyte. Isis, 76(3):409–411, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232887.

Crowe:1985:BRBa

[1567] Michael J. Crowe. Book review: The View from Planet Earth: Man Looks at the Cosmos by Vincent Cronin. Isis, 76(3):411, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232888.

Rousseau:1985:BRBb

[1568] G. S. Rousseau. Book review: Poems of Science by John Heath-Stubbs; Phillips Salman. Isis, 76(3):411–412, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232889.

Slade:1985:BRB

[1569] Joseph W. Slade. Book review: The Vital Science: Biology and the Literary Imagination, 1860–1900 by Peter Morton. Isis, 76(3):412–413, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232890.

Sluga:1985:BRB

[1570] Hans Sluga. Book review: Frege: Logical Excavations by G. P. Baker; P. M. S. Hacker; Frege’s Conception of Numbers as Objects by Crispin Wright. Isis, 76(3):413–415, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232891.

Jones:1985:BRBd

[1571] Charles V. Jones. Book review: Mathematics: People, Problems, Results by Douglas M. Campbell; John C. Higgins. Isis, 76(3):415, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232892.

Unguru:1985:BRBa

[1572] Sabetai Unguru. Book review: Rechenstein, Experiment, Sprache: Historische Fallstudien zur Entstehung der exakten Wissenschaften by Peter Damerow; Wolfgang Lefevre. Isis, 76(3):415–416, September 1985.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232893.

Swinbank:1985:BRB

[1573] Peter Swinbank. Book review: The Story of Astronomy in Edinburgh from Its Beginnings until 1975 by Hermann A. Bruck. Isis, 76(3):416– 417, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232894.

Hetherington:1985:BRB

[1574] Norriss S. Hetherington. Book review: The History of Astronomy from Herschel to Hertzsprung by Dieter B. Herrmann; Kevin Krisciunas. Isis, 76(3):417–418, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232895.

Brush:1985:BRBa

[1575] Stephen G. Brush. Book review: Cosmology and Astrophysics: Essays in Honor of Thomas Gold by Yervant Terzian; Elizabeth Bilson. Isis, 76(3):418, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232896.

Ospovat:1985:BRBb

[1576] Alexander Ospovat. Book review: Die Genesis der Geologie als Wissenschaft: Studie zu den kognitiven Prozessen und gesellschaftlichen Bedingungen bei der Herausbildung der Geologie als naturwissenschaftliche Disziplin an der Wende des 18. und 19. Jahrhundert by Martin Guntau. Isis, 76(3):418–419, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232897.

Kish:1985:BRBb

[1577] George Kish. Book review: United States Coastal Charts, 1783–1861 by Peter J. Guthorn. Isis, 76(3):419, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232898.

Montgomery:1985:BRB

[1578] William Montgomery. Book review: Charles Darwin, 1809–1882: A Centennial Commemorative by Roger G. Chapman; Cleveland T. Duval. Isis, 76(3):420–421, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232899.

Sterling:1985:BRB

[1579] Keir B. Sterling. Book review: The Emergence of Ornithology as a Scientific Discipline: 1760–1850 by Paul L. Farber. Isis, 76(3):421–422, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232900.

Egerton:1985:BRB

[1580] Frank N. Egerton. Book review: Zoological Philosophy: An Exposition with Regard to the Natural History of Animals by J. B. Lamarck; Hugh Elliot. Isis, 76(3):422–423, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232901.

Tobey:1985:BRB

[1581] Ronald C. Tobey. Book review: Discoverers of the Lost World: An Account of Some of Those Who Brought Back to Life South American Mammals Long Buried in the Abyss of Time by George Gaylord Simpson. Isis, 76(3):423–424, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232902.

Overfield:1985:BRB

[1582] Richard A. Overfield. Book review: Contributions to the History of North American Natural History by Alwyne Wheeler. Isis, 76(3):424– 425, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232903.

Blackden:1985:BRB

[1583] Stephanie Blackden. Book review: The History of Medicine in Ireland by John F. Fleetwood. Isis, 76(3):425, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232904.

Risse:1985:BRB

[1584] Guenter B. Risse. Book review: Historical Origins of the Concept of Neurosis by Jose M. Lopez Pinero; Doris Ines Berrios. Isis, 76(3):425– 426, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232905.

Kidwell:1985:BRBa

[1585] Clara Sue Kidwell. Book review: The Admirable Secrets of Physick and Chyrurgery by Thomas Palmer; Thomas Rogers Forbes. Isis, 76(3):426– 427, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232906.

Reynolds:1985:BRB

[1586] Terry S. Reynolds. Book review: The Microwave Debate by Nicholas H. Steneck. Isis, 76(3):427–428, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232907.

Smith:1985:BRBa

[1587] Robert W. Smith. Book review: Beyond Velikovsky: The History of a Public Controversy by Henry H. Bauer. Isis, 76(3):428–429, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232908.

Glidden:1985:BRB

[1588] David K. Glidden. Book review: Epicurus’ Scientific Method by Elizabeth Asmis. Isis, 76(3):429–430, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232909.

Pinault:1985:BRB

[1589] Jody Rubin Pinault. Book review: Cinq cents ans de bibliographie hippocratique, 1473–1982 by G. Maloney; R. Savoie. Isis, 76(3):430, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232910.

Thomson:1985:BRB

[1590] Ron B. Thomson. Book review: Archimedes in the Middle Ages. Volume V: Quasi-Archimedean Geometry in the Thirteenth Century by Marshall Clagett. Isis, 76(3):430–431, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232911.

Toomer:1985:BRB

[1591] G. J. Toomer. Book review: Les Tables Alphonsines avec les canons de Jean de Saxe by Emmanuel Poulle. Isis, 76(3):431–433, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232912.

Pingree:1985:BRB

[1592] David Pingree. Book review: Nicephore Gregoras: Calcul de l’eclipse de soleil due 16 juillet 1330 by Nicephorus Gregoras; J. Mogenet; A.

Tihon; R. Royez; A. Berg. Isis, 76(3):433, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232913.

Miller:1985:BRBc

[1593] Timothy S. Miller. Book review: Gesundheitswesen und Wohltatigkeit im Spiegel der byzantinischen Klostertypika by Robert Volk. Isis, 76(3): 433–434, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232914.

Dols:1985:BRB

[1594] Michael W. Dols. Book review: Le Livre de la methode du medicin de Ali b. Ridwan (998–1067) by Ali ibn Ridwan; Jacques Grand’Henry. Isis, 76 (3):434–435, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232915.

King:1985:BRB

[1595] David A. King. Book review: Glossar der arabischen Fachausdrucke in der mittelalterlichen europaischen Astrolabliteratur by Paul Kunitzsch. Isis, 76(3):435, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232916.

Sheehan:1985:BRB

[1596] Helen E. Sheehan. Book review: Western Medicine in India: Epidemics and Other Diseases by O. P. Jaggi; Western Medicine in India: Medical Education and Research; Western Medicine in India: Public Health and Its Administration; Western Medicine in India: Social Impact. Isis, 76 (3):435–437, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232917.

Elman:1985:BRB

[1597] Benjamin Elman. Book review: The Development and Decline of Chinese Cosmology by John Henderson. Isis, 76(3):437–438, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232918.

Pan:1985:BRB

[1598] Jixing Pan. Book review: China and Charles Darwin by James Reeve Pusey. Isis, 76(3):438–439, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232919.

Maddin:1985:BRB

[1599] Robert Maddin. Book review: Kodo Zuroku: Illustrated Book on the Smelting of Copper by Masuda Tsuna; Cyril Stanley Smith; Zenryu Shirakawa. Isis, 76(3):439–440, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232920.

Moran:1985:BRB

[1600] Bruce T. Moran. Book review: Copernicus and the Scientific Revolution by Edward Rosen. Isis, 76(3):440–441, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232921.

Murray:1985:BRBb

[1601] T. J. Murray. Book review: A History of Neurophysiology in the 17th and 18th Centuries: From Concept to Experiment by Mary A. B. Brazier. Isis, 76(3):441, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232922.

Kleinert:1985:BRB

[1602] Andreas Kleinert. Book review: Théories électriques du XVIIIe siècle: Correspondance entre l’Abbé Nollet (1700–1770) et le physicien génèvois Jean Jallabert (1712–1768) by Isaac Benguigui. Isis, 76(3):442, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232923.

Dobbs:1985:BRB

[1603] Betty Jo Teeter Dobbs. Book review: In the Presence of the Creator: Isaac Newton and His Times by Gale E. Christianson. Isis, 76(3):443– 444, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232924.

Cohen:1985:BRBb

[1604] Albert Cohen. Book review: Quantifying Music: The Science of Music at the First Stage of the Scientific Revolution, 1580–1650 by H. Floris Cohen. Isis, 76(3):444–445, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232925.

Browne:1985:BRB

[1605] Janet Browne. Book review: Augerius Clutius and His “De Hemerobio”: An Early Work on Ephemeroptera by Frans P. M. Francissen; Ad W.

M. Mol. Isis, 76(3):445–446, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232926.

Osler:1985:BRB

[1606] Margaret J. Osler. Book review: John Locke and Medicine by Patrick Romanell. Isis, 76(3):446–447, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232927.

Crowe:1985:BRBb

[1607] Michael J. Crowe. Book review: William Rowan Hamilton: Portrait of a Prodigy by Sean O’Donnell. Isis, 76(3):447–448, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232928.

Bowler:1985:BRBb

[1608] Peter J. Bowler. Book review: T. H. Huxley’s Place in Natural Science by Mario A. di Gregorio. Isis, 76(3):448–449, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232929.

Pyenson:1985:BRB

[1609] Lewis Pyenson. Book review: Scienza e colonialismo nel Canada ottocentesco by Clelia Pighetti. Isis, 76(3):449–450, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232930.

Kidwell:1985:BRBb

[1610] Peggy Aldrich Kidwell. Book review: Maclear and Herschel: Letters and Diaries at the Cape of Good Hope, 1834–1838 by Brian Warner; Nancy Warner. Isis, 76(3):450, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232931.

McVittie:1985:BRB

[1611] G. C. McVittie. Book review: Eddington: The Most Distinguished Astrophysicist of His Time by S. Chandrasekhar. Isis, 76(3):451, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232932.

Brown:1985:BRBb

[1612] Laurie M. Brown. Book review: Gesammelte Werke; Collected Works by Werner Heisenberg; W. Blum; H.-P. Durr; H. Rechenberg. Isis, 76 (3):451–452, September 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232933.

Garciadiego:1985:BRB

[1613] Alejandro Garciadiego. Book review: La España posible de la Segunda Republica: La oferta a Einstein de una Catedra extraordinaria en la Universidad Central (Madrid, 1933) by Jose M. Sanchez-Ron; Thomas F. Glick. Isis, 76(3):452–453, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232934.

Drachman:1985:BRB

[1614] Virginia G. Drachman. Book review: Alice Hamilton: A Life in Letters by Barbara Sicherman. Isis, 76(3):453, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232935.

Anonymous:1985:BMc

[1615] Anonymous. Back matter. Isis, 76(3):454–458, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232936.

Anonymous:1985:FMd

[1616] Anonymous. Front matter. Isis, 76(4):459–465, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233020.

Thackray:1985:EB

[1617] Arnold Thackray. An end and a beginning. Isis, 76(4):466–469, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233021.

Merton:1985:GSE

[1618] Robert K. Merton. George Sarton: Episodic recollections by an unruly apprentice. Isis, 76(4):470–486, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233022.

Glick:1985:GSS

[1619] Thomas F. Glick. George Sarton and the Spanish Arabists. Isis, 76 (4):487–499, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233023.

Johnson:1985:ACI

[1620] Jeffrey A. Johnson. Academic chemistry in Imperial Germany. Isis, 76 (4):500–524, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233024.

Rydell:1985:FDS

[1621] Robert W. Rydell. The fan dance of science: American World’s fairs in the great depression. Isis, 76(4):525–542, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233025.

Smith:1985:GPE

[1622] A. Mark Smith. Galileo’s proof for the Earth’s motion from the movement of sunspots. Isis, 76(4):543–551, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233026.

Josephson:1985:SHS

[1623] Paul R. Josephson. Soviet historians and the structure of scientific revolutions. Isis, 76(4):551–559, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233027.

Blondel:1985:APR

[1624] Christine Blondel and L. Pearce Williams. Ampere and the programming of research. Isis, 76(4):559–561, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233028.

Melhado:1985:IS

[1625] Evan M. Melhado, Gad Freudenthal, Zafer Toprak, Selcuk Tozeren, Selim Deringil, Yakov Rabkin, and Ivo Schneider. International symposia. Isis, 76(4):562–566, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233029.

Wallace:1985:EJA

[1626] William A. Wallace. Eloge: James Athanasius Weisheipl, O.P. 3 July 1923–30 December 1984. Isis, 76(4):566–567, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233030.

Ortiz:1985:EJB

[1627] Eduardo L. Ortiz and Lewis Pyenson. Eloge: Jose Babini, 11 May 1897– 18 May 1984. Isis, 76(4):567–569, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233031.

Moore:1985:BRBb

[1628] James R. Moore. Book review: A Calendar of the Correspondence of Charles Darwin, 1821–1882. Isis, 76(4):570–580, December 1985. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233032.

Reingold:1985:BRB

[1629] Nathan Reingold. Book review: Transformation and Tradition in the Sciences: Essays in Honor of I, by Everett Mendelsohn. Bernard Cohen. Isis, 76(4):580–584, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233033.

Smith:1985:BRBb

[1630] Cyril Stanley Smith. Book review: The Scientific Reinterpretation of Form. Isis, 76(4):584–587, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233034.

Coleman:1985:BRB

[1631] William Coleman. Book review: Crime, Madness, and Politics in Modern France: The Medical Concept of National Decline. Isis, 76(4):587– 590, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233035.

McClellan:1985:BRB

[1632] James E. McClellan. Book review: The Classics of Science: A Study of Twelve Enduring Scientific Works by Derek Gjertsen. Isis, 76(4):591, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233036.

Gregory:1985:BRB

[1633] Michael Gregory. Book review: The Atom by Bill Stonebarger; Michael Stonebarger. Isis, 76(4):591–593, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233037.

Galatzer-Levy:1985:BRB

[1634] Robert M. Galatzer-Levy. Book review: Genius, Creativity and Leadership: Historiometric Inquiries by Dean K. Simonton. Isis, 76(4):593–594, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233038.

Mahoney:1985:BRB

[1635] Michael S. Mahoney. Book review: An Annotated Bibliography on the History of Data Processing by James W. Cortada. Isis, 76(4):594–595, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233039.

Fraser:1985:BRB

[1636] Craig G. Fraser. Book review: The History of Mathematics from Antiquity to the Present: A Selective Bibliography by Joseph W. Dauben. Isis, 76(4):595–596, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233040.

Ovitt:1985:BRBb

[1637] George Ovitt, Jr. Book review: Middle English Prose: A Critical Guide to Major Authors and Genres by A. S. G. Edwards. Isis, 76(4):596, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233041.

Harman:1985:BRB

[1638] P. M. Harman. Book review: The History of Classical Physics: A Selected, Annotated Bibliography by R. W. Home; Mark J. Gittins. Isis, 76 (4):596–597, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233042.

Danziger:1985:BRBb

[1639] Kurt Danziger. Book review: A Century of Serial Publications in Psychology, 1850–1950: An International Bibliography by Donald V. Osier; Robert H. Wozniak. Isis, 76(4):597, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233043.

Wells:1985:BRBb

[1640] Ellen B. Wells. Book review: Wie finde ich Literatur zur Geschichte der Naturwissenschaften und Technik? by Burghard Weiss. Isis, 76(4): 597–598, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233044.

Numbers:1985:BRBb

[1641] Ronald L. Numbers. Book review: American Medicine Comes of Age 1840–1920: Essays to Commemorate the Founding of the Journal of The American Medical Association, July 14, 1883 by Lester S. King. Isis, 76 (4):598, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233045.

Miller:1985:BRBd

[1642] David Philip Miller. Book review: Science and Technology in Manchester: Two Hundred Years of the Lit. and Phil. by Chris E. Makepeace. Isis, 76(4):598–599, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233046.

Reeves:1985:BRB

[1643] Barbara J. Reeves. Book review: I Congressi degli Scienziati Italiani nell’eta del positivismo by Giuliano Pancaldi. Isis, 76(4):599–600, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233047.

Feingold:1985:BRBb

[1644] Mordechai Feingold. Book review: The John Tradescants: Gardeners to the Rose and Lily Queen by Prudence Leith-Ross; Tradescant’s Rarities: Essays on the Foundation of the Ashmolean Museum, 1683; with a Catalogue of the Surviving Early Collections by Arthur MacGregor; Elias Ashmole, 1617–1692: A Tercentenary Exhibition by Michael Hunter; The Ashmolean Museum and Oxford Science, 1683–1983 by A. V. Simcock. Isis, 76(4):600–602, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233048.

Noble:1985:BRB

[1645] David W. Noble. Book review: A Strategic Analysis of Science and Technology Policy by Harvey A. Averch. Isis, 76(4):602–603, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233049.

Barber:1985:BRBb

[1646] Bernard Barber. Book review: Scientific Rationality: The Sociological Turn by James Robert Brown. Isis, 76(4):603–604, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233050.

Hoch:1985:BRB

[1647] Paul K. Hoch. Book review: Refugee Scholars in America: Their Impact and Their Experiences by Lewis Coser. Isis, 76(4):604–605, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233051.

Goodstein:1985:BRB

[1648] Judith R. Goodstein. Book review: Understanding Relativity: Origin and Impact of a Scientific Revolution by Stanley Goldberg. Isis, 76(4): 605–606, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233052.

Roland:1985:BRB

[1649] Alex Roland. Book review: Technological Utopianism in American Culture by Howard P. Segal. Isis, 76(4):606–607, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233053.

Olesko:1985:BRB

[1650] Kathryn Olesko. Book review: Zur Entstehung des modernen Systems wissenschaftlicher Disziplinen: Physik in Deutschland, 1740–1890 by Rudolf Stichweh. Isis, 76(4):607–608, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233054.

Reeds:1985:BRB

[1651] Karen Reeds. Book review: Drawn from Nature: The Botanical Art of Joseph Prestele and His Sons by Charles van Ravenswaay. Isis, 76 (4):608–609, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233055.

Bylebyl:1985:BRB

[1652] Jerome J. Bylebyl. Book review: The Paradox of Rembrandt’s Anatomy of Dr. Tulp. by William Schupbach. Isis, 76(4):609–610, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233056.

Gooding:1985:BRBb

[1653] David Gooding. Book review: Cockroft and the Atom by Guy Hartcup; T. E. Allibone. Isis, 76(4):610–611, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233057.

Pickering:1985:BRB

[1654] Andy Pickering. Book review: The Quest for Quarks by Brian McCusker. Isis, 76(4):611–612, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233058.

Unger:1985:BRB

[1655] Richard W. Unger. Book review: Die Theorie des Segelns in ihrer Entwicklung: Geschichte eines Problems der nautischen Mechanik by Ludwig Rank. Isis, 76(4):612, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233059.

Brown:1985:BRBc

[1656] Laurie M. Brown. Book review: Otto Hahn and the Rise of Nuclear Physics by William R. Shea. Isis, 76(4):612–613, December 1985. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233060.

Cantor:1985:BRB

[1657] Geoffrey Cantor. Book review: Brewster and Wheatstone on Vision by Nicholas J. Wade. Isis, 76(4):613–614, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233061.

Davies:1985:BRBb

[1658] Gordon L. Herries Davies. Book review: Contributions to the History of Geological Mapping by E. Dudich. Isis, 76(4):614–615, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233062.

Janik:1985:BRB

[1659] Allan Janik. Book review: Theorie der Geowissenschaft by Wolf von Engelhardt; Jorg Zimmermann. Isis, 76(4):615–616, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233063.

Greene:1985:BRB

[1660] Mott T. Greene. Book review: Die Einfuhrung des Naturselbstdruckes und der Photographie in die erdwissenschaftliche Dokumentation by Franz Kirchheimer. Isis, 76(4):616–617, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233064.

Rossiter:1985:BRB

[1661] Margaret W. Rossiter. Book review: Science, Agriculture, and the Politics of Research by Lawrence Busch; William B. Lacy. Isis, 76(4):617– 618, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233065.

Bell:1985:BRB

[1662] C. Ritchie Bell. Book review: Frank N. Meyer: Plant Hunter in Asia by Isabel Shipley Cunningham. Isis, 76(4):618–619, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233066.

Albrow:1985:BRB

[1663] Martin Albrow. Book review: Max Weber and the Methodology of the Social Sciences by Toby E. Huff. Isis, 76(4):619, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233067.

Peel:1985:BRB

[1664] J. D. Y. Peel. Book review: Functionalism Historicized: Essays on British Social Anthropology by George W. Stocking. Isis, 76(4):619–620, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233068.

Brieger:1985:BRBb

[1665] Gert H. Brieger. Book review: Medicine and the Management of Living: Taming the Last Great Beast by William Ray Arney; Bernard J. Bergen. Isis, 76(4):620–621, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233069.

Rosenberg:1985:BRB

[1666] Charles Rosenberg. Book review: Jelliffe: American Psychoanalyst and Physician, and His Correspondence with Sigmund Freud and C. G. Jung by John C. Burnham; William McGuire. Isis, 76(4):621–622, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233070.

Hannaway:1985:BRB

[1667] Caroline Hannaway. Book review: Science and Medicine in France: The Emergence of Experimental Physiology, 1790–1855 by John E. Lesch. Isis, 76(4):622–623, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233071.

Gelfand:1985:BRB

[1668] Toby Gelfand. Book review: The Clinical Consultations of Giambattista Morgagni: The Edition of Enrico Benassi (1935) by Giambattista Morgagni; Saul Jarcho. Isis, 76(4):623–624, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233072.

Morley:1985:BRB

[1669] Jane Morley. Book review: Publishers and Librarians: A Foundation for Dialogue by Mary Biggs. Isis, 76(4):624–625, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233073.

Lewis:1985:BRB

[1670] W. David Lewis. Book review: Pioneer Plastic: The Making and Selling of Celluloid by Robert Friedel. Isis, 76(4):625–626, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233074.

Kranakis:1985:BRB

[1671] Eda Fowlks Kranakis. Book review: Bridge to the Future: A Centennial Celebration of the Brooklyn Bridge by Margaret Latimer; Brooke Hindle; Melvin Kranzberg. Isis, 76(4):626–627, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233075.

Berggren:1985:BRB

[1672] J. L. Berggren. Book review: Mathematik in Antike und Orient by Helmuth Gericke. Isis, 76(4):627–628, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233076.

Clagett:1985:BRBb

[1673] Marshall Clagett. Book review: Les pretres-ouab de Sekhmet et les conjurateurs de Serket by Frederique von Kanel. Isis, 76(4):628–629, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233077.

Landels:1985:BRB

[1674] J. G. Landels. Book review: Greek and Roman Mechanical Water-Lifting Devices: The History of a Technology by John Peter Oleson. Isis, 76(4): 629–630, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233078.

Hansen:1985:BRBb

[1675] Bert Hansen. Book review: Le “Quaestiones de sensu” attribuite a Oresme e Alberto de Sassonia by Jole Agrimi. Isis, 76(4):630–631, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233079.

deLuca:1985:BRB

[1676] Emilie Patton deLuca. Book review: “Livre des simples medecines,” Codex Bruxellensis IV 1024: A 15th-century French Herbal by Enid Roberts; William T. Stearn; Carmelia Opsomer. Isis, 76(4):631–632, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233080.

Varisco:1985:BRB

[1677] Daniel Martin Varisco. Book review: Über eine anwa -Tradition mit bisher unbekannten Sternnamen by Paul Kunitzsch. Isis, 76(4):632, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233081.

Riddle:1985:BRB

[1678] John M. Riddle. Book review: The Arabic Materia Medica of Dioscorides by Mahmoud M. Sadek. Isis, 76(4):633–634, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233082.

Major:1985:BRB

[1679] John S. Major. Book review: Science and Civilisation in China. Volume VI: Biology and Biological Technology. Part 2: Agriculture by Joseph Needham; Francesca Bray. Isis, 76(4):634–635, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233083.

Unguru:1985:BRBb

[1680] Sabetai Unguru. Book review: Leonardo da Vinci’s Elements of the Science of Man by Kenneth D. Keele. Isis, 76(4):635–636, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233084.

Hatch:1985:BRB

[1681] Robert A. Hatch. Book review: Libro del nuevo cometa; Littera ad Bartholomaeum Reisacherum; Summa del prognostico del cometa by Jeronimo Munoz; Victor Navarro Brotons; Elizabeth Ladd. Isis, 76(4): 636–637, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233085.

Palmer:1985:BRB

[1682] Richard Palmer. Book review: Professori e promotori di medicina nello Studio di Padova dal 1405 al 1509: Repertorio bio-bibliografico by Tiziana Pesenti. Isis, 76(4):637, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233086.

Ormsby-Lennon:1985:BRB

[1683] Hugh Ormsby-Lennon. Book review: John Case and Aristotelianism in Renaissance England by Charles B. Schmitt. Isis, 76(4):637–638, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233087.

Joy:1985:BRB

[1684] Lynn S. Joy. Book review: William Whiston: Honest Newtonian by James E. Force. Isis, 76(4):639–640, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233088.

Finocchiaro:1985:BRBb

[1685] Maurice A. Finocchiaro. Book review: Novita celesti e crisi del sapere: Atti del Convegno internazionale di Studi galileiani by Paolo Galluzzi. Isis, 76(4):640, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233089.

Gingerich:1985:BRB

[1686] Owen Gingerich. Book review: Gesammelte Werke. Volume XI, Part 1: Ephemerides novae motuum coelestium by Johannes Kepler; Volker Bialas. Isis, 76(4):640–641, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233090.

Simcock:1985:BRB

[1687] A. V. Simcock. Book review: Thomas Beddoes, M.D., 1760–1808: Chemist, Physician, Democrat by Dorothy A. Stansfield. Isis, 76(4): 641–642, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233091.

Wisan:1985:BRB

[1688] Winifred Lovell Wisan. Book review: Galileo and His Sources: The Heritage of the Collegio Romano in Galileo’s Science by William A. Wallace. Isis, 76(4):642–643, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233092.

Rocke:1985:BRB

[1689] A. J. Rocke. Book review: Science versus Practice: Chemistry in Victorian Britain by Robert F. Bud; Gerrylynn K. Roberts. Isis, 76(4): 643–644, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233093.

Hilts:1985:BRB

[1690] Victor L. Hilts. Book review: American Medicine and Statistical Thinking, 1800–1860 by James H. Cassedy. Isis, 76(4):644–645, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233094.

Grabiner:1985:BRBb

[1691] Judith V. Grabiner. Book review: A Convergence of Lives: Sofia Kovalevskaia, Scientist, Writer, Revolutionary by Ann Hibner Koblitz. Isis, 76(4):645–646, December 1985. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233095.

Brush:1985:BRBb

[1692] Stephen G. Brush. Book review: James E. Keeler, Pioneer American Astrophysicist, and the Early Development of American Astrophysics by Donald E. Osterbrock. Isis, 76(4):646–647, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233096.

Hirsh:1985:BRB

[1693] Richard F. Hirsh. Book review: On Mars: Exploration of the Red Planet 1958–1978 by Edward Clinton Ezell; Linda Neuman Ezell. Isis, 76(4): 648, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233097.

Vietor:1985:BRB

[1694] Richard H. K. Vietor. Book review: Controlling the Atom: The Beginnings of Nuclear Regulation, 1946–1962 by George Mazuzan; J. Samuel Walker. Isis, 76(4):648–649, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233098.

Anonymous:1985:BMd

[1695] Anonymous. Back matter. Isis, 76(4):650–672, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233099.

Anonymous:1985:VI

[1696] Anonymous. Volume information. Isis, 76(4):655–667, December 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233019.

Anonymous:1985:FMe

[1697] Anonymous. Front matter. Isis, 76(5):1–4, ???? 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231435.

Neu:1985:OHT

[1698] John Neu. One hundred tenth critical bibliography of the history of science and its cultural influences. Isis, 76(5):5–240, ???? 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231436.

Anonymous:1985:BMe

[1699] Anonymous. Back matter. Isis, 76(5):??, ???? 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231437.

Anonymous:1986:FMa

[1700] Anonymous. Front matter. Isis, 77(1):1–6, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232498.

Colp:1986:CMD

[1701] Ralph Colp, Jr. “Confessing a murder”: Darwin’s first revelations about transmutation. Isis, 77(1):8–32, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232499.

Weiss:1986:WSL

[1702] Sheila Faith Weiss. Wilhelm Schallmayer and the logic of German eugenics. Isis, 77(1):33–46, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232500.

Stewart:1986:PLP

[1703] Larry Stewart. Public lectures and private patronage in Newtonian England. Isis, 77(1):47–58, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232501.

Greenberg:1986:MPE

[1704] John L. Greenberg. Mathematical physics in Eighteenth-Century France. Isis, 77(1):59–78, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232502.

Alexandre:1986:CHT

[1705] Josette Alexandre. La comete de Halley a travers les ouvrages et manuscrits de l’observatoire de Paris. Isis, 77(1):79–84, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232503.

McMullin:1986:GBO

[1706] Ernan McMullin. Giordano Bruno at Oxford. Isis, 77(1):85–94, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232504.

Elton:1986:EGR

[1707] Lewis Elton. Einstein, General Relativity, and the German press, 1919– 1920. Isis, 77(1):95–103, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232505.

vanderWaerden:1986:AOS

[1708] B. L. van der Waerden. Aristarchos’s observation of the Summer solstice. Isis, 77(1):103–104, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232506.

Grant:1986:EER

[1709] Edward Grant. Eloge: Edward Rosen, 12 December 1906–28 March 1985. Isis, 77(1):105–106, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232507.

Beller:1986:BRW

[1710] Mara Beller. Book review: Wie es eigentlich gewesen? [How was it actually?] The Creation of Quantum Mechanics and the Bohr–Pauli Dialogue. Isis, 77(1):107–109, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232508.

Westfall:1986:BRB

[1711] Richard S. Westfall. Book review: Revolution in Science by I. Bernard Cohen. Isis, 77(1):109–110, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232509.

MacKinnon:1986:BRBa

[1712] Edward MacKinnon. Book review: Physical Sciences and History of Physics by R. S. Cohen; M. W. Wartofsky. Isis, 77(1):110–111, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232510.

Hacking:1986:BRBa

[1713] Ian Hacking. Book review: Science and Reality: Recent Work in the Philosophy of Science by James T. Cushing; C. F. Delaney; Gary M. Gutting. Isis, 77(1):111–112, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232511.

McCauley:1986:BRB

[1714] Robert N. McCauley. Book review: Science as Cognitive Process: Toward an Empirical Philosophy of Science by Robert A. Rubinstein; Charles D. Laughlin,; John McManus. Isis, 77(1):112–113, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232512.

Rynasiewicz:1986:BRB

[1715] Robert Rynasiewicz. Book review: How Things Are: Studies in Predication and the History of Philosophy and Science by James Bogen; James E. McGuire. Isis, 77(1):113–114, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232513.

Kitcher:1986:BRB

[1716] Patricia Kitcher. Book review: Kant and the Double Government Methodology: Supersensibility and Method in Kant’s Philosophy of Science by Robert E. Butts. Isis, 77(1):114–115, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232514.

Koslow:1986:BRB

[1717] Arnold Koslow. Book review: Thinking Matter: Materialism in Eighteenth-Century Britain by John W. Yolton. Isis, 77(1):115–116, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232515.

Ribe:1986:BRB

[1718] Neil M. Ribe. Book review: The Transcendent Science: Kant’s Conception of Biological Methodology by Clark Zumbach. Isis, 77(1):116, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232516.

MacKinnon:1986:BRBb

[1719] Edward MacKinnon. Book review: Images of Science: Essays on Realism and Empiricism by Paul M. Churchland; Clifford A. Hooker; Bas C. van Fraassen. Isis, 77(1):116–117, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232517.

Howard:1986:BRB

[1720] Don Howard. Book review: The Philosophy of Niels Bohr: The Framework of Complementarity by Henry J. Folse. Isis, 77(1):117–118, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232518.

Galison:1986:BRB

[1721] Peter Galison. Book review: Representing and Intervening: Introductory Topics in the Philosophy of Natural Science by Ian Hacking. Isis, 77(1): 118–120, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232519.

Curd:1986:BRB

[1722] Martin Curd. Book review: Scientific Realism by Jarrett Leplin. Isis, 77(1):120–121, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232520.

Schumacher:1986:BRB

[1723] John A. Schumacher. Book review: The Limits of Science by Nicholas Rescher. Isis, 77(1):121–123, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232521.

Hesse:1986:BRB

[1724] Mary Hesse. Book review: Scientific Explanation and the Causal Structure of the World by Wesley C. Salmon. Isis, 77(1):123–124, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232522.

Nicholas:1986:BRB

[1725] John M. Nicholas. Book review: Science and Scepticism by John Watkins. Isis, 77(1):124–125, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232523.

Hatfield:1986:BRB

[1726] Gary Hatfield. Book review: A Theory of Method by Husain Sarkar. Isis, 77(1):125, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232524.

Shea:1986:BRBa

[1727] William R. Shea. Book review: Imagery in Scientific Thought: Creating Twentieth–Century Physics by Arthur I. Miller. Isis, 77(1):125–126, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232525.

Ihde:1986:BRBa

[1728] Don Ihde. Book review: Data, Instruments, and Theory: A Dialectical Approach to Understanding Science by Robert John Ackermann. Isis, 77(1):126–127, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232526.

Schmaus:1986:BRB

[1729] Warren Schmaus. Book review: Modern Science and Human Values by William W. Lowrance. Isis, 77(1):127–128, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232527.

Hull:1986:BRBa

[1730] David L. Hull. Book review: Evolution at a Crossroads: The New Biology and the New Philosophy of Science by David J. Depew; Bruce H. Weber. Isis, 77(1):128–129, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232528.

Miller:1986:BRBa

[1731] David Philip Miller. Book review: All Scientists Now: The Royal Society in the Nineteenth Century. Isis, 77(1):130–133, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232529.

North:1986:BRB

[1732] J. D. North. Book review: Astrophysics and Twentieth–Century Astronomy to 1950. Isis, 77(1):133–136, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232530.

Swerdlow:1986:BRB

[1733] N. M. Swerdlow. Book review: The Optical Papers of Isaac Newton, by Isaac Newton; Alan E. Shapiro. Volume 1: The Optical Lectures, 1670– 1672. Isis, 77(1):136–140, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232531.

Lyon:1986:BRB

[1734] John Lyon. Book review: Uneasy Genius: The Life and Work of Pierre Duhem by Stanley L. Jaki. Isis, 77(1):141–142, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232532.

Hall:1986:BRBa

[1735] A. Rupert Hall. Book review: On the Shoulders of Giants: A Shandean Postscript by Robert K. Merton. Isis, 77(1):142–143, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232533.

Pickles:1986:BRB

[1736] John Pickles. Book review: The History of Modern Geography: An Annotated Bibliography of Selected Works by Gary S. Dunbar. Isis, 77(1):

143–144, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232534.

Numbers:1986:BRB

[1737] Ronald L. Numbers. Book review: The History of the Health Care Sciences and Health Care, 1700–1980: A Selective Annotated Bibliography by Jonathon Erlen. Isis, 77(1):144–145, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232535.

Reingold:1986:BRB

[1738] Nathan Reingold. Book review: The Beginnings of the Nobel Institution: The Science Prizes, 1901–1915 by Elisabeth Crawford; The Nobel Prize by Peter Wilhelm. Isis, 77(1):145–146, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232536.

Rosenberg:1986:BRB

[1739] Robert Rosenberg. Book review: A Century of Electrical Engineering and Computer Science at MIT, 1882–1982 by Karl L. Wildes; Nilo A. Lindgren. Isis, 77(1):146–147, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232537.

Thackray:1986:BRBa

[1740] Arnold Thackray. Book review: The Lamp of Learning: Taylor & Francis and the Development of Science Publishing by W. M. Brock; A. J. Meadows. Isis, 77(1):147–148, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232538.

MacLeod:1986:BRB

[1741] Roy MacLeod. Book review: Scientists, Society, and State: The Social Relations of Science Movement in Great Britain, 1931–1947 by William McGucken. Isis, 77(1):148–150, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232539.

Pauly:1986:BRB

[1742] Philip J. Pauly. Book review: Expository Science: Forms and Functions of Popularization by Terry Shinn; Richard Whitley. Isis, 77(1):150– 151, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232540.

Mahoney:1986:BRBa

[1743] Michael S. Mahoney. Book review: La correspondance de Blaise Pascal et de Pierre de Fermat: La géometrie du hasard ou le début du calcul des probabilités by Pierre-Jose About; Michel Boy. Isis, 77(1):151–152, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232541.

Rider:1986:BRB

[1744] Robin E. Rider. Book review: The Analytic Art: Nine Studies in Algebra, Geometry, and Trigonometry from the Opus restitutae mathematicae analyseos, seu Algebra nova by Francois Viete; T. Richard Witmer. Isis, 77(1):152–153, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232542.

Calinger:1986:BRB

[1745] Ronald Calinger. Book review: Number Theory: An Approach through History, from Hammurapi to Legendre by Andre Weil. Isis, 77(1):153– 154, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232543.

Ceruzzi:1986:BRB

[1746] Paul Ceruzzi. Book review: Punched Card Methods in Scientific Computation by Wallace J. Eckert; Calculating Machines: Recent and Prospective Developments and Their Impact on Mathematical Physics and Calculating Instruments and Machines by Douglas R. Hartree. Isis, 77(1): 154–156, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232544.

Williams:1986:BRBa

[1747] M. R. Williams. Book review: Mr. Babbage’s Secret: The Tale of a Cypher — and APL by Ole Immanuel Franksen. Isis, 77(1):156, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232545.

Williams:1986:BRBb

[1748] Bernard O. Williams. Book review: The Preparation of Programs for an Electronic Digital Computer, with Special Reference to the “ED-SAC” and the Use of a Library of Subroutines by Maurice V. Wilkes; David J. Wheeler; Stanley Gill. Isis, 77(1):157, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232546.

Smith:1986:BRBa

[1749] A. Mark Smith. Book review: L’ottica dalle origini all’inizio del ’700 by Fabio Bevilacqua; Maria Grazia Ianniello. Isis, 77(1):158, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232547.

Lankford:1986:BRB

[1750] John Lankford. Book review: Rings: Discoveries from Galileo to Voyager by James Elliot; Richard Kerr. Isis, 77(1):158–159, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232548.

Melhado:1986:BRBa

[1751] Evan M. Melhado. Book review: The Central Science: Essays on the Uses of Chemistry by George B. Kauffman; H. Harry Szmant; Chemistry and Modern Society by John Parascandola; James C. Whorton. Isis, 77(1):159–161, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232549.

Gearhart:1986:BRB

[1752] Clayton Gearhart. Book review: Comet Halley: Once in a Lifetime by Mark Littmann; Donald K. Yeomans. Isis, 77(1):161, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232550.

Vucinich:1986:BRBa

[1753] Alexander Vucinich. Book review: Petersburgskaia astronomicheskaia shkola XVIII v. by N. I. Nevskaia; Le développement de l’astronomie en U.R.S.S., 1917–1935 by E. Nicolaidis. Isis, 77(1):161–162, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232551.

Paul:1986:BRB

[1754] H. W. Paul. Book review: Physique et physiciens en France, 1918– 1940 by Dominique Pestre. Isis, 77(1):162–163, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232552.

Rocke:1986:BRBa

[1755] A. J. Rocke. Book review: Recent Developments in the History of Chemistry by C. A. Russell. Isis, 77(1):163–164, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232553.

Beer:1986:BRB

[1756] John J. Beer. Book review: The Tincal Trail: A History of Borax by N. J. Travis; E. J. Cocks. Isis, 77(1):164–165, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232554.

Gingerich:1986:BRBa

[1757] Owen Gingerich. Book review: Measuring the Universe: Cosmic Dimensions from Aristarchus to Halley by Albert Van Helden. Isis, 77(1): 165–166, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232555.

Traynham:1986:BRB

[1758] James G. Traynham. Book review: A History of the Nomenclature of Organic Chemistry by Pieter Eduard Verkade. Isis, 77(1):166–167, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232556.

Allen:1986:BRB

[1759] Garland E. Allen. Book review: Natural Selection, Heredity, and Eugenics; Including Selected Correspondence of R. A. Fisher with Leonard Darwin and Others by J. H. Bennett. Isis, 77(1):168–169, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232557.

Reeds:1986:BRBa

[1760] Karen Reeds. Book review: Botanical Drugs of the Americas in the Old and New Worlds: Invitational Symposium at the Washington Congress, 1983. Amerikanische pflanzliche Arzneien in der Alten und Neuen Welt: Einladungs Symposium anlasslich des Kongresses in Washington, 1983 by Wolfgang-Hagen Hein; K. Dietrich. Isis, 77(1):169, March 1986. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232558.

Kingsland:1986:BRB

[1761] Sharon Kingsland. Book review: The Background of Ecology: Concept and Theory by Robert P. McIntosh. Isis, 77(1):169–170, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232559.

Teigen:1986:BRB

[1762] Philip M. Teigen. Book review: The English Plant Names in “The Grete Herball” (1526): A Contribution to the Historical Study of English Plant-Name Usage by Mats Ryden. Isis, 77(1):171, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232560.

Samelson:1986:BRB

[1763] Franz Samelson. Book review: From Darwinism to Behaviourism: Psychology and the Minds of Animals by Robert Boakes. Isis, 77(1):171–172, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232561.

Gillespie:1986:BRB

[1764] Neal C. Gillespie. Book review: The Taming of Evolution: The Persistence of Nonevolutionary Views in the Study of Humans by Davydd J. Greenwood. Isis, 77(1):172–173, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232562.

Callahan:1986:BRB

[1765] William J. Callahan. Book review: M. Seoane y la introduccion en España del sistema sanitario liberal, 1791–1870 by J. M. Lopez Pinero. Isis, 77(1):173–174, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232563.

Hildreth:1986:BRB

[1766] Martha L. Hildreth. Book review: Medical Heritage by Sharon Romm. Isis, 77(1):174–175, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232564.

Leavitt:1986:BRB

[1767] Judith W. Leavitt. Book review: Sympathy and Science: Women Physicians in American Medicine by Regina Markell Morantz-Sanchez. Isis, 77(1):175–176, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232565.

Warner:1986:BRBa

[1768] John Harley Warner. Book review: A Calculus of Suffering: Pain, Professionalism, and Anesthesia in Nineteenth-Century America by Martin S. Pernick. Isis, 77(1):176–177, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232566.

Thomson:1986:BRB

[1769] Margaret S. Thomson. Book review: A Generous Confidence: Thomas Story Kirkbride and the Art of Asylum Keeping, 1840–1883 by Nancy Tomes. Isis, 77(1):177–178, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232567.

Porter:1986:BRBa

[1770] Roy Porter. Book review: Poison in the Pot: The Legacy of Lead by Richard P. Wedeen. Isis, 77(1):178–179, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232568.

Kline:1986:BRB

[1771] Ronald R. Kline. Book review: The Continuous Wave: Technology and American Radio, 1900–1932 by Hugh G. J. Aitken. Isis, 77(1):179– 180, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232569.

Swenson:1986:BRB

[1772] Loyd S. Swenson, Jr. Book review: Living and Working in Space: A History of Skylab by W. David Compton; Charles D. Benson. Isis, 77 (1):180–181, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232570.

Noble:1986:BRB

[1773] David W. Noble. Book review: Industrial Evolution: Organization, Structure, and Growth of the Pennsylvania Iron Industry, 1750–1860 by Paul F. Paskoff. Isis, 77(1):181–182, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232571.

Eastwood:1986:BRB

[1774] Bruce Eastwood. Book review: Pseudo-Bede: De mundi celestis terrestrisque constitutione –A Treatise on the Universe and the Soul by Charles Burnett. Isis, 77(1):182–183, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232572.

Freudenthal:1986:BRB

[1775] Gad Freudenthal. Book review: Graceful Reason: Essays in Ancient and Medieval Philosophy by Lloyd P. Gerson. Isis, 77(1):183–184, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232573.

Saliba:1986:BRBa

[1776] George Saliba. Book review: Oriens-Occidens: Ausgewahlte Schriften zur Wissenschafts- und Kulturgeschichte. Volume II by Willy Hartner; Y. Maeyama. Isis, 77(1):184–185, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232574.

Shank:1986:BRB

[1777] Michael H. Shank. Book review: Nicolaus Cusanus und die Entstehung der exakten Wissenschaften by Fritz Nagel. Isis, 77(1):185–186, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232575.

Grant:1986:BRB

[1778] Edward Grant. Book review: Nicole Oresme and the Marvels of Nature: A Study of His De causis mirabilium with Critical Edition, Translation, and Commentary by Nicole Oresme; Bert Hansen. Isis, 77(1):186–187, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232576.

Demaitre:1986:BRBa

[1779] Luke Demaitre. Book review: Scholastic Medicine and Philosophy: A Study of Commentaries on Galen’s Tegni (ca. 1300–1450) by Per-Gunnar Ottosson. Isis, 77(1):187–188, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232577.

Bos:1986:BRB

[1780] H. J. M. Bos. Book review: Albertus (Magnus)’ Commentaar op Euclides’ Elementen der Geometrie by P. M. J. E. Tummers. Isis, 77(1): 188, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232578.

Glick:1986:BRBa

[1781] Thomas F. Glick. Book review: Los moriscos y la medicina: Un capitulo de la medicina y la ciencia marginadas en la España del siglo XVI by Luis Garcia Ballester. Isis, 77(1):189, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232579.

Debus:1986:BRBa

[1782] Allen G. Debus. Book review: Los impresos cientificos españoles de los siglos XV y XVI: Inventario, bibliometria, y thesaurus by Jose M. Lopez Pinero; Francesc Bujosa Homar; Victor Navarro Brotons; Eugenio Portela Marco; Maria Luz Lopez Terrada; Jose Pardo Tomas. Isis, 77 (1):189–190, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232580.

Wallace:1986:BRB

[1783] William A. Wallace. Book review: The Louvain Lectures (Lectiones lovanienses) of Bellarmine and the Autograph Copy of His 1616 Declaration to Galileo by Robert Bellarmine; Ugo Baldini; George V. Coyne; Francois de Aguilon, S.J. (1576–1617); Scientist and Architect by August Ziggelaar. Isis, 77(1):190–192, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232581.

Finocchiaro:1986:BRB

[1784] Maurice A. Finocchiaro. Book review: The Galileo Affair: A Meeting of Faith and Science by G. V. Coyne; M. Heller; J. Zycinski. Isis, 77(1): 192, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232582.

Applebaum:1986:BRB

[1785] Wilbur Applebaum. Book review: The Birth of History and Philosophy of Science: Kepler’s A Defence of Tycho against Ursus with Essays on Its Provenance and Significance by N. Jardine. Isis, 77(1):192–194, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232583.

Deason:1986:BRB

[1786] Gary B. Deason. Book review: Francis Bacon’s Natural Philosophy: A New Source. A Transcription of Manuscript Hardwick 72A with Translation and Commentary by Francis Bacon; Graham Rees; Christopher Upton. Isis, 77(1):194–195, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232584.

Ormsby-Lennon:1986:BRB

[1787] Hugh Ormsby-Lennon. Book review: The Works of Thomas Vaughan by Thomas Vaughan; Alan Rudrum; Jennifer Drake-Brockman. Isis, 77(1):195–196, March 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232585.

Crosland:1986:BRB

[1788] Maurice Crosland. Book review: Cauchy: Un mathematicien legitimiste au XIXe siècle by Bruno Belhoste. Isis, 77(1):196–197, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232586.

Gregory:1986:BRBa

[1789] Frederick Gregory. Book review: Utopie und Wissenschaft im fruhen Deutschen Sozialismus. Theoriebildung und Wissenschaftsbegriff bei Wilhelm Weitling by Lothar Knatz. Isis, 77(1):197–198, March 1986. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232587.

Reeves:1986:BRB

[1790] Barbara J. Reeves. Book review: Tecnica e cultura della crisi (1914– 1939) by Michela Nacci. Isis, 77(1):198, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232588.

Hirsh:1986:BRB

[1791] Richard F. Hirsh. Book review: The X-Ray Universe by Wallace Tucker; Riccardo Giacconi. Isis, 77(1):198–199, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232589.

Anonymous:1986:BMa

[1792] Anonymous. Back matter. Isis, 77(1):200–204, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232590.

Anonymous:1986:FMb

[1793] Anonymous. Front matter. Isis, 77(2):205–211, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232648.

Abbri:1986:GEH

[1794] Ferdinando Abbri and Paolo Rossi. Guest editorial: History of science in Italy. Isis, 77(2):212–218, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232649.

Morawski:1986:OKB

[1795] J. G. Morawski. Organizing knowledge and behavior at Yale’s institute of human relations. Isis, 77(2):219–242, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232650.

Leventhal:1986:EPD

[1796] Robert S. Leventhal. The emergence of philological discourse in the German states, 1770–1810. Isis, 77(2):243–260, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232651.

Stepan:1986:RGR

[1797] Nancy Leys Stepan. Race and gender: The role of analogy in science. Isis, 77(2):261–277, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232652.

Kerker:1986:SMR

[1798] Milton Kerker, Svedberg, and Per Stenius. The Svedberg and molecular reality: An autobiographical postscript. Isis, 77(2):278–282, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232653.

Hill:1986:GWN

[1799] David K. Hill. Galileo’s work on 116v: A new analysis. Isis, 77(2):283– 291, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232654.

Webb:1986:SPO

[1800] George E. Webb. Solar physics and the origins of dendrochronology. Isis, 77(2):291–301, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232655.

VanHelden:1986:AMH

[1801] Albert Van Helden and Audrey B. Davis. Annual meeting of the History of Science Society, 31 October–3 November 1985. Isis, 77(2):302–305, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232656.

Anonymous:1986:SM

[1802] Anonymous. Sarton Medals. Isis, 77(2):305–307, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232657.

Anonymous:1986:DSF

[1803] Anonymous. Donors to and supporters of the fund drive. Isis, 77(2): 308–309, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232658.

Anonymous:1986:DSE

[1804] Anonymous. Dedication and symposium at the Edelstein Center Jerusalem, June 1985. Isis, 77(2):309–310, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232659.

Olby:1986:BRB

[1805] Robert Olby, R. C. Lewontin, and Daniel J. Kevles. Book review: In the Name of Eugenics: Genetics and the Uses of Human Heredity by Daniel J. Kevles. Isis, 77(2):311–319, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232660.

Schaffer:1986:BRB

[1806] Simon Schaffer. Book review: The Dark Abyss of Time: The History of the Earth and the History of Nations from Hooke to Vico. Isis, 77(2): 320–323, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232661.

Lowenthal:1986:BRB

[1807] David Lowenthal. Book review: Voyage into Substance: Art, Science, and the Illustrated Travel Account, 1760–1840. Isis, 77(2):324–327, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232662.

Stafford:1986:BRB

[1808] Barbara M. Stafford. Book review: The Art of Captain Cook’s Voyages by Rudiger Joppien; Bernard Smith; European Vision and the South Pacific by Bernard Smith. Isis, 77(2):328, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232663.

Baatz:1986:BRB

[1809] Simon Baatz. Book review: Latrobe’s Views of America, 1795–1820: Selections from the Watercolors and Sketches by Edward C. Carter; John C. Van Horne; Charles E. Brownell; Tina H. Sheller; Stephen F. Lintner; J. Frederick Fausz; Geraldine C. Vickers. Isis, 77(2):328–329, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232664.

Kauffman:1986:BRB

[1810] George B. Kauffman. Book review: The Periodic Table. Isis, 77(2):330– 332, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232665.

Badash:1986:BRB

[1811] Lawrence Badash. Book review: Bird of Passage: Recollections of a Physicist by Rudolf Peierls. Isis, 77(2):332–333, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232666.

Kargon:1986:BRB

[1812] Robert H. Kargon. Book review: The Education of a College President: A Memoir by James R. Killian. Isis, 77(2):333–334, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232667.

Messeri:1986:BRB

[1813] Peter Messeri. Book review: The Intellectual and Social Organization of the Sciences by Richard Whitley. Isis, 77(2):335–336, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232668.

Kren:1986:BRB

[1814] Claudia Kren. Book review: Mapping the Cosmos by Jane Chance; R. O. Wells. Isis, 77(2):336, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232669.

Multhauf:1986:BRB

[1815] R. P. Multhauf. Book review: The Barchas Collection: The Making of Modern Science by Henry E. Lowood. Isis, 77(2):336–337, June 1986.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232670.

Sivin:1986:BRB

[1816] Nathan Sivin. Book review: A Historical Catalogue of Scientists and Scientific Books: From the Earliest Times to the Close of the Nineteenth Century by Robert Mortimer Gascoigne. Isis, 77(2):337, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232671.

Eamon:1986:BRBa

[1817] William Eamon. Book review: Medieval Science and Technology: A Selected, Annotated Bibliography by Claudia Kren. Isis, 77(2):337–338, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232672.

Porter:1986:BRBb

[1818] Charlotte M. Porter. Book review: Natural History in America, 1609– 1860: Printed Works in the Collections of the American Philosophical Society, the Historical Society of Pennsylvania, and the Library Company of Philadelphia by Andrea J. Tucher. Isis, 77(2):338–339, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232673.

Fancher:1986:BRB

[1819] Raymond E. Fancher. Book review: Hommes superieurs, hommes inferieurs? La controverse sur l’heredite de l’intelligence by Gerard Lemaine; Benjamin Matalon; Clemencon Mireille; Gomis Alain; Ramunni Girolamo; Ability, Merit, and Measurement: Mental Testing and English Education, 1880–1940 by Gillian Sutherland; Stephen Sharp. Isis, 77(2):339–341, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232674.

Lewenstein:1986:BRB

[1820] Bruce V. Lewenstein. Book review: Scientists and Journalists: Reporting Science as News by Sharon M. Friedman; Sharon Dunwoody; Carol L. Rogers. Isis, 77(2):341–342, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232675.

Theerman:1986:BRB

[1821] Paul Theerman. Book review: Science and Social Change in Britain and Europe, 1700–1900 by Colin Russell. Isis, 77(2):342–343, June 1986.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232676.

Schlebecker:1986:BRB

[1822] John T. Schlebecker. Book review: “Venerate the Plough”: A History of the Philadelphia Society for Promoting Agriculture, 1785–1985 by Simon Baatz. Isis, 77(2):343–344, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232677.

Outram:1986:BRB

[1823] Dorinda Outram. Book review: Science Reorganized: Scientific Societies in the Eighteenth Century by James E. McClellan. Isis, 77(2):344, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232678.

Rowe:1986:BRB

[1824] David E. Rowe. Book review: Der Briefwechsel David Hilbert-Felix Klein (1886–1918) by Gunther Frei. Isis, 77(2):345, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232679.

Trenn:1986:BRBa

[1825] Thaddeus J. Trenn. Book review: From Protyle to Proton: William Prout and the Nature of Matter, 1785–1985 by W. H. Brock. Isis, 77(2): 345–346, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232680.

Moyer:1986:BRB

[1826] Albert E. Moyer. Book review: The Young Einstein: The Advent of Relativity by Lewis Pyenson. Isis, 77(2):346–347, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232681.

Ariew:1986:BRB

[1827] Roger Ariew and Peter Barker. Book review: Chronology of Eclipses and Comets, A.D. 1–1000 by D. Justin Schove; Alan Fletcher. Isis, 77(2): 347–348, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232682.

Wise:1986:BRB

[1828] George Wise. Book review: Chemistry in America, 1876–1976: Historical Indicators by Arnold Thackray; Jeffrey L. Sturchio; P. Thomas Carroll; Robert Bud. Isis, 77(2):348–349, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232683.

Glen:1986:BRB

[1829] William Glen. Book review: History of Geophysics. Volume I by C. Stewart Gillmor. Isis, 77(2):349–351, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232684.

Melhado:1986:BRBb

[1830] Evan M. Melhado. Book review: Die Entdeckung des Isomorphismus: Eine Fallstudie zur Geschichte der Mineralogie und der Chemie by Hans-Werner Schutt. Isis, 77(2):351–352, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232685.

Thompson:1986:BRB

[1831] Paul Thompson. Book review: Just before the Origin: Alfred Russel Wallace’s Theory of Evolution by John Langdon Brooks. Isis, 77(2): 352–353, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232686.

Manier:1986:BRB

[1832] Edward Manier. Book review: Dimensions of Darwinism: Themes and Counterthemes in Twentieth–Century Evolutionary Theory by Marjorie Grene. Isis, 77(2):353–355, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232687.

Egerton:1986:BRB

[1833] Frank N. Egerton. Book review: Modeling Nature: Episodes in the History of Population Ecology by Sharon E. Kingsland. Isis, 77(2):355, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232688.

Hull:1986:BRBb

[1834] David L. Hull. Book review: Vaulting Ambition: Sociobiology and the Quest for Human Nature by Philip Kitcher. Isis, 77(2):356–357, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232689.

Rossiter:1986:BRB

[1835] Margaret W. Rossiter. Book review: Agricultural Science and the Quest for Legitimacy: Farmers, Agricultural Colleges, and Experiment Stations, 1870–1890 by Alan I. Marcus. Isis, 77(2):357–358, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232690.

Mayr:1986:BRB

[1836] Ernst Mayr. Book review: Evolutionary Theory: The Unfinished Synthesis by Robert G. B. Reid. Isis, 77(2):358–359, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232691.

DeJager-Seerveld:1986:BRB

[1837] Tim DeJager-Seerveld. Book review: The Structure of Biological Science by Alexander Rosenberg. Isis, 77(2):359–360, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232692.

Cocks:1986:BRBa

[1838] Geoffrey Cocks. Book review: Psychologie im Nationalsozialismus by Carl Friedrich Graumann. Isis, 77(2):360, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232693.

Stocking:1986:BRBa

[1839] George W. Stocking, Jr. Book review: “So Much That Is New”: Baldwin Spencer, 1860–1929: A Biography by D. J. Mulvaney; J. H. Calaby. Isis, 77(2):360–361, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232694.

Danziger:1986:BRB

[1840] Kurt Danziger. Book review: Psychiatry in an Anthropological and Biomedical Context: Philosophical Presuppositions and Implications of German Psychiatry, 1820–1870 by Gerlof Verwey. Isis, 77(2):361–362, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232695.

Jones:1986:BRB

[1841] James H. Jones. Book review: No Magic Bullet: A Social History of Venereal Disease in the United States since 1880 by Allan M. Brandt.

Isis, 77(2):362–363, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232696.

Grob:1986:BRB

[1842] Gerald N. Grob. Book review: The Anatomy of Madness: Essays in the History of Psychiatry by W. F. Bynum; Roy Porter; Michael Shepherd. Isis, 77(2):363–364, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232697.

Patterson:1986:BRB

[1843] K. David Patterson. Book review: The Caribbean Slave: A Biological History by Kenneth D. Kiple. Isis, 77(2):364–365, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232698.

Berggren:1986:BRB

[1844] J. L. Berggren. Book review: Ibn al-Haytham’s Completion of the Conics by J. P. Hogendijk. Isis, 77(2):365–367, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232699.

Unguru:1986:BRB

[1845] Sabetai Unguru. Book review: Gli “inizi” della matematica greca: Le congetture kantiano-popperiane di Arpad Szabo by Silvestro Marcucci. Isis, 77(2):367, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232700.

Siraisi:1986:BRB

[1846] Nancy G. Siraisi. Book review: Arnaldi de Villanova opera medica omnia. Volume XV: Commentum supra tractatum Galieni De malicia complexionis diverse by Arnald of Villanova; Luis Garcia Ballester; Eustaquio Sanchez Salor; Doctrina Galieni De Interioribus by Richard J. Durling. Isis, 77(2):367–368, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232701.

Dales:1986:BRB

[1847] Richard C. Dales. Book review: Opera omnia. Volume IV: Schriften zur Naturwissenschaften: Briefe by Dietrich von Freiberg; Maria-Rita Pagnoni-Sturlese; Rudolf Rehn; Loris Sturlese; William A. Wallace. Isis, 77(2):368–369, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232702.

Saliba:1986:BRBb

[1848] George Saliba. Book review: Theory and Observation in Ancient and Medieval Astronomy by Bernard R. Goldstein. Isis, 77(2):369–370, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232703.

Getz:1986:BRB

[1849] Faye Marie Getz. Book review: Hippocrates Latinus: Repertorium of Hippocratic Writings in the Latin Middle Ages by Pearl Kibre. Isis, 77 (2):370, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232704.

McVaugh:1986:BRB

[1850] Michael McVaugh. Book review: Studies in Medieval Science: Alchemy, Astrology, Mathematics and Medicine by Pearl Kibre. Isis, 77(2):370– 371, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232705.

Kennedy:1986:BRB

[1851] E. S. Kennedy. Book review: The Astronomy of Levi ben Gerson (1288– 1344): A Critical Edition of Chapters 1–20 with Translation and Commentary by Levi ben Gerson; Bernard R. Goldstein. Isis, 77(2):371–372, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232706.

Scarborough:1986:BRBa

[1852] John Scarborough. Book review: The Birth of the Hospital in the Byzantine Empire by Timothy S. Miller. Isis, 77(2):372–373, June 1986. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232707.

Demaitre:1986:BRBb

[1853] Luke E. Demaitre. Book review: Doctors and Medicine in Early Renaissance Florence by Katharine Park. Isis, 77(2):373–374, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232708.

Debus:1986:BRBb

[1854] Allen G. Debus. Book review: The Medical Renaissance of the Sixteenth Century by A. Wear; R. K. French; I. M. Lonie. Isis, 77(2):374–375, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232709.

Jacob:1986:BRBa

[1855] James R. Jacob. Book review: The Art of Memory: A Treatise Useful for Such as Are to Speak in Publick by Marius D’Assigny; The Immortality of the Human Soul, Demonstrated by the Light of Nature by Walter Charleton; A Free and Impartial Censure of the Platonick Philosophie: Being a Letter Written to His Much Honoured Friend, M.N.B. by Samuel Parker. Isis, 77(2):375–376, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232710.

Shea:1986:BRBb

[1856] William R. Shea. Book review: Carteggio, 1649–1656 by Paolo Galluzzi; Maurizio Torrini. Isis, 77(2):376–377, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232711.

Donahue:1986:BRB

[1857] William H. Donahue. Book review: Le secret du monde by Jean Kepler; Alain Segonds; Louis-Paul Cousin. Isis, 77(2):377–378, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232712.

Porter:1986:BRBc

[1858] Roy Porter. Book review: Under Newton’s Shadow: Astronomical Practices in the Seventeenth Century by Lesley Murdin. Isis, 77(2):378–379, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232713.

Barfoot:1986:BRB

[1859] Michael Barfoot. Book review: Church and University in the Scottish Enlightenment: The Moderate Literati of Edinburgh by Richard B. Sher. Isis, 77(2):379, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232714.

Glick:1986:BRBb

[1860] Thomas F. Glick. Book review: Historia de la ciencia en Mexico: Estudios y textos. Volume II: Siglo XVII by Elias Trabulse. Isis, 77(2): 380, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232715.

Rocke:1986:BRBb

[1861] A. J. Rocke. Book review: Justus von Liebig und August Wilhelm Hofmann in ihren Briefen (1841–1873) by William Hodson Brock. Isis, 77 (2):380–381, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232716.

Porter:1986:BRBd

[1862] Theodore M. Porter. Book review: The Cultural Meaning of Popular Science: Phrenology and the Organization of Consent in Nineteenth-Century Britain by Roger Cooter. Isis, 77(2):381–383, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232717.

Appel:1986:BRB

[1863] Toby A. Appel. Book review: Georges Cuvier: Vocation, Science and Authority in Post-Revolutionary France by Dorinda Outram. Isis, 77(2): 383–384, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232718.

Dauben:1986:BRBa

[1864] Joseph W. Dauben. Book review: Writings of Charles S. Peirce: A Chronological Edition. Volume II:1867–1871 by Charles S. Peirce; Edward C. Moore; Max H. Fisch; Christian J. W. Kloesel; Don D. Roberts; Lynn A. Zeigler. Isis, 77(2):384–386, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232719.

Kohlstedt:1986:BRB

[1865] Sally Gregory Kohlstedt. Book review: All the World’s a Fair: Visions of Empire at American International Expositions, 1876–1916 by Robert W. Rydell. Isis, 77(2):386–387, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232720.

Rechenberg:1986:BRB

[1866] H. Rechenberg. Book review: Scientific Correspondence with Bohr, Einstein, Heisenberg and Others. Volume II: 1930–1939 by Wolfgang Pauli; Karl von Meyenn; Armin Hermann; Victor F. Weisskopf. Isis, 77(2): 387–388, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232721.

Heims:1986:BRB

[1867] Steve J. Heims. Book review: Niels Bohr: A Centenary Volume by A. P. French; P. J. Kennedy. Isis, 77(2):388–389, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232722.

Pyenson:1986:BRBa

[1868] Lewis Pyenson. Book review: El secreto atomico de Huemul: Cronica del origen de la energia atomica en la Argentina by Mario Mariscotti. Isis, 77(2):389–391, June 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232723.

Terry:1986:BRB

[1869] Robert H. Terry. Book review: The Geography of Peace and War by David Pepper; Alan Jenkins. Isis, 77(2):391, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232724.

Anonymous:1986:BMb

[1870] Anonymous. Back matter. Isis, 77(2):392–398, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232725.

Anonymous:1986:FMc

[1871] Anonymous. Front matter. Isis, 77(3):399–403, September 1986. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231605.

Cornell:1986:NGD

[1872] John F. Cornell. Newton of the grassblade? Darwin and the problem of organic teleology. Isis, 77(3):404–421, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231606.

Rowe:1986:JMG

[1873] David E. Rowe. “Jewish mathematics” at Göttingen in the era of Felix Klein. Isis, 77(3):422–449, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231607.

Kay:1986:WMS

[1874] Lily E. Kay. W. M. Stanley’s crystallization of the tobacco mosaic virus, 1930–1940. Isis, 77(3):450–472, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231608.

Wisan:1986:GGC

[1875] Winifred Lovell Wisan. Galileo and God’s creation. Isis, 77(3):473–486, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231609.

Yunming:1986:ACS

[1876] Zhang Yunming. Ancient Chinese sulfur manufacturing processes. Isis, 77(3):487–497, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231610.

Kertesz:1986:NIO

[1877] G. A. Kertesz. Notes on Isis von Oken, 1817–1848. Isis, 77(3):497–503, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231611.

Hall:1986:EHG

[1878] Marie Boas Hall. Eloge: Henry Guerlac, 10 June 1910–29 May 1985. Isis, 77(3):504–506, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231612.

Multhauf:1986:EWS

[1879] Robert P. Multhauf. Eloge: Wilson L. Scott, 5 August 1909–10 November 1983. Isis, 77(3):507, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231613.

Turner:1986:BRBa

[1880] Frank M. Turner. Book review: The Great Devonian Controversy: The Shaping of Scientific Knowledge among Gentlemanly Specialists. Isis, 77 (3):508–511, September 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231614.

Melhado:1986:BRBc

[1881] Evan M. Melhado. Book review: Lavoisier and the Chemistry of Life: An Exploration of Scientific Creativity. Isis, 77(3):511–514, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231615.

Glick:1986:BRBc

[1882] Thomas F. Glick. Book review: Dynamis: Acta Hispanica ad Medicinae Scientiarumque Historiam Illustrandum. Isis, 77(3):514–517, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231616.

Papatheofanis:1986:BRB

[1883] Frank J. Papatheofanis. Book review: About Science by Barry Barnes. Isis, 77(3):518–519, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231617.

Kremer:1986:BRB

[1884] Richard L. Kremer. Book review: History and Philosophy of Science: Selected Papers by Joseph W. Dauben; Virginia Staudt Sexton. Isis, 77(3):519, September 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231618.

Pycior:1986:BRB

[1885] Helena M. Pycior. Book review: Some People and Places in Irish Science and Technology by Charles Mollan; William Davis; Brendan Finucane. Isis, 77(3):519–520, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231619.

Hall:1986:BRBb

[1886] A. Rupert Hall. Book review: Religion, Science, and World-view: Essays in Honor of Richard S. Westfall by Margaret J. Osler; Paul Lawrence Farber. Isis, 77(3):520–521, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231620.

Gooding:1986:BRB

[1887] David Gooding. Book review: Faraday to Einstein: Constructing Meaning in Scientific Theories by Nancy J. Nersessian. Isis, 77(3):521–523, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231621.

Ihde:1986:BRBb

[1888] Don Ihde. Book review: Philosophy and Sociology of Science: An Introduction by Stewart Richards. Isis, 77(3):523, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231622.

Warner:1986:BRBb

[1889] John Harley Warner. Book review: Learning to Heal: The Development of American Medical Education by Kenneth M. Ludmerer. Isis, 77(3):

523–524, September 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231623.

Green:1986:BRB

[1890] Monica H. Green. Book review: Scienza e filosofia all Universita di Padova nel Quattrocento by Antonino Poppi. Isis, 77(3):524–525, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231624.

Wheaton:1986:BRBa

[1891] Bruce R. Wheaton. Book review: Constructing Quarks: A Sociological History of Particle Physics by Andrew Pickering. Isis, 77(3):525–527, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231625.

Zinberg:1986:BRB

[1892] Dorothy S. Zinberg. Book review: America’s Golden Bough: The Science Advisory Intertwist by Thaddeus J. Trenn. Isis, 77(3):527, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231626.

Brown:1986:BRBa

[1893] Laurie M. Brown. Book review: Gesammelte Werke — Collected Works. Series A/Part 1: Wissenschaftliche Originalarbeiten — Original Scientific Papers by Werner Heisenberg; W. Blum; H.-P. Durr; H. Rechenberg. Isis, 77(3):527–528, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231627.

Turner:1986:BRBb

[1894] R. Steven Turner. Book review: Physics History from AAPT Journals by Melba Newell Phillips. Isis, 77(3):528–529, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231628.

vonMayrhauser:1986:BRB

[1895] Richard T. von Mayrhauser. Book review: The Intelligence Men: Makers of the IQ Controversy by Raymond E. Fancher. Isis, 77(3):529–530, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231629.

Reed:1986:BRB

[1896] Edward S. Reed. Book review: The Mind’s New Science: A History of the Cognitive Revolution by Howard Gardner. Isis, 77(3):530–532, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231630.

Burnham:1986:BRB

[1897] John C. Burnham. Book review: The Origins of Behaviorism: American Psychology, 1870–1920 by John M. O’Donnell. Isis, 77(3):532–533, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231631.

Stocking:1986:BRBb

[1898] George W. Stocking, Jr. Book review: A Century of Controversy: Ethnological Issues from 1860 to 1960 by Elman R. Service. Isis, 77(3):533, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231632.

Thackray:1986:BRBb

[1899] Arnold Thackray. Book review: Behavioral and Social Science: Fifty Years of Discovery by Neil M. Smelser; Dean R. Gerstein. Isis, 77(3): 534, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231633.

Nye:1986:BRBa

[1900] Robert Nye. Book review: Madness, Morality, and Medicine: A Study of the York Retreat, 1796–1914 by Anne Digby. Isis, 77(3):534–536, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231634.

Porter:1986:BRBe

[1901] R. S. Porter. Book review: Surgeons at the Bailey: English Forensic Medicine to 1878 by Thomas Rogers Forbes. Isis, 77(3):536–537, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231635.

Risse:1986:BRB

[1902] Guenter B. Risse. Book review: The Royal Protomedicato: The Regulation of the Medical Profession in the Spanish Empire by John Tate Lanning; John Jay TePaske. Isis, 77(3):537–538, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231636.

McNeill:1986:BRB

[1903] William H. McNeill. Book review: The Impact of the Plague in Tudor and Stuart England by Paul Slack. Isis, 77(3):538–539, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231637.

Scarborough:1986:BRBb

[1904] John Scarborough. Book review: Bibliografia Hipocratica by Blas Bruni Celli. Isis, 77(3):539, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231638.

Barham:1986:BRB

[1905] James Barham. Book review: A History of Ancient Psychiatry by Giuseppe Roccatagliata. Isis, 77(3):539–540, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231639.

Furley:1986:BRB

[1906] David Furley. Book review: Vestigia Democritea: Die Rezeption der Lehre von den Atomen in der antiken Naturwissenschaft und Medizin by Alfred Stuckelberger; Bernhard Wyss; Denis van Berchem; Olof Gigon. Isis, 77(3):540–541, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231640.

Lang:1986:BRB

[1907] Helen S. Lang. Book review: Truth and Scientific Knowledge in the Thought of Henry of Ghent by Steven P. Marrone. Isis, 77(3):541–542, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231641.

Moran:1986:BRB

[1908] Bruce T. Moran. Book review: “Wo zwei zusammenkommen in rechter Ehr… ”: Sozio- und psychogenetische Studien über Eheschliessungsvorgange vom 12. bis 15. Jahrhundert by Michael Schroter. Isis, 77(3): 542–543, September 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231642.

Gingerich:1986:BRBb

[1909] Owen Gingerich. Book review: Gesamtausgabe. Volume II: De revolutionibus: Kritischer Text by Nicolaus Copernicus; Heribert M. Nobis; Bernhard Sticker. Isis, 77(3):543–544, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231643.

Osler:1986:BRBa

[1910] Margaret J. Osler. Book review: Science, Faith, and Politics: Francis Bacon and the Utopian Roots of the Modern Age — a Commentary on Bacon’s Advancement of Learning by Jerry Weinberger. Isis, 77(3):544, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231644.

Jongsma:1986:BRB

[1911] Calvin Jongsma. Book review: George Boole: His Life and Work by Desmond MacHale. Isis, 77(3):544–545, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231645.

Nye:1986:BRBb

[1912] Mary Jo Nye. Book review: From Knowledge to Power: The Rise of the Science Empire in France, 1860–1939 by Harry W. Paul. Isis, 77 (3):545–546, September 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231646.

Porter:1986:BRBf

[1913] Charlotte M. Porter. Book review: The Expeditions of John Charles Fremont. Volume III: Travels from 1848 to 1854 by Mary Lee Spence. Isis, 77(3):546–547, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231647.

Vucinich:1986:BRBb

[1914] Alexander Vucinich. Book review: Kapitza, Rutherford, and the Kremlin by Lawrence Badash. Isis, 77(3):547–548, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231648.

Cocks:1986:BRBb

[1915] Geoffrey Cocks. Book review: Geschichte der deutschen Psychologie im 20. Jahrhundert: Ein Uberblick by Mitchell G. Ash; Ulfried Geuter. Isis, 77(3):548–549, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231649.

Decker:1986:BRB

[1916] Hannah S. Decker. Book review: Psychotherapy in the Third Reich: The Goring Institute by Geoffrey Cocks. Isis, 77(3):549–550, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231650.

Pyenson:1986:BRBb

[1917] Lewis Pyenson. Book review: La ciencia periferica: Ciencia y sociedad en Venezuela by Elena Diaz; Yolanda Texera; Hebe Vessuri; Ciencia academica en la Venezuela moderna: Historia reciente y perspectivas de las disciplinas cientificas by Hebe M. Vessuri. Isis, 77(3):550–551, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231651.

Durbin:1986:BRB

[1918] Paul T. Durbin. Book review: Technology’s Storytellers: Reweaving the Human Fabric by John M. Staudenmaier. Isis, 77(3):551–552, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231652.

Mitchum:1986:BRB

[1919] Carl Mitchum. Book review: A Theory of Technology: Continuity and Change in Human Development by Thomas R. DeGregori; International Technology Transfer: Concepts, Measures, and Comparisons by Nathan Rosenberg; Claudio Frischtak. Isis, 77(3):552–554, September 1986. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231653.

Leslie:1986:BRB

[1920] Stuart W. Leslie. Book review: The Making of a Profession: A Century of Electrical Engineering in America by A. Michal McMahon. Isis, 77 (3):554–556, September 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231654.

Mowery:1986:BRB

[1921] David C. Mowery. Book review: The Telephone Enterprise: The Evolution of the Bell System’s Horizontal Structure, 1876–1909 by Robert W. Garnet; The Anatomy of a Business Strategy: Bell, Western Electric, and the Origins of the American Telephone Industry by George David Smith; From Invention to Innovation: The Case of Long-Distance Telephone Transmission at the Turn of the Century by Neil H. Wasserman. Isis, 77(3):556–557, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231655.

Smith:1986:BRBb

[1922] John K. Smith. Book review: Rohm and Haas: History of a Chemical Company by Sheldon Hochheiser. Isis, 77(3):557–558, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231656.

Layton:1986:BRB

[1923] Edwin T. Layton, Jr. Book review: The Making of American Industrial Research: Science and Business at GE and Bell, 1876–1926 by Leonard S. Reich; Willis R. Whitney, General Electric, and the Origins of U.S. Industrial Research by George Wise. Isis, 77(3):558–560, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231657.

Dick:1986:BRB

[1924] Steven J. Dick. Book review: The Jodrell Bank Telescopes by Bernard Lovell. Isis, 77(3):560–561, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231658.

Friedel:1986:BRB

[1925] Robert Friedel. Book review: Origins of Materials and Processes by John Delmonte. Isis, 77(3):561, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231659.

Morris:1986:BRB

[1926] Peter J. T. Morris. Book review: Synthetic Rubber: The Project That Had to Succeed by Vernon Herbert; Attilio Bisio. Isis, 77(3):561–562, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231660.

Roland:1986:BRB

[1927] Alex Roland. Book review: To Engineer Is Human: The Role of Failure in Successful Design by Henry Petroski. Isis, 77(3):562–563, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231661.

Carlson:1986:BRB

[1928] W. Bernard Carlson. Book review: Technology on Trial: The Introduction of Steam Power into Sweden, 1715–1736 by Svante Lindqvist.

Isis, 77(3):564–565, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231662.

Reynolds:1986:BRB

[1929] Terry S. Reynolds. Book review: A History of Industrial Power in the United States, 1780–1930. Volume II: Steam Power by Louis C. Hunter. Isis, 77(3):565–566, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231663.

Staudenmaier:1986:BRB

[1930] John M. Staudenmaier. Book review: Military Enterprise and Technological Change: Perspectives on the American Experience by Merritt Roe Smith. Isis, 77(3):566–567, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231664.

Smith:1986:BRBc

[1931] Cyril Stanley Smith. Book review: Islamic Metalwork in the Freer Gallery of Art by Esin Atil; W. T. Chase; Paul Jett. Isis, 77(3):567–568, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231665.

Varey:1986:BRB

[1932] J. E. Varey. Book review: The Twenty-One Books of Devices and of Machines by Pseudo-Juanelo Turriano; Jose A. Garcia-Diego. Isis, 77 (3):568–570, September 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231666.

Bedini:1986:BRB

[1933] Silvio A. Bedini. Book review: The Correspondence and Miscellaneous Papers of Benjamin Henry Latrobe. Volume I: 1784–1804 by John C. Van Horne; Lee W. Formwalt. Isis, 77(3):570, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231667.

McGaw:1986:BRB

[1934] Judith A. McGaw. Book review: Mechanical Metamorphosis: Technological Change in Revolutionary America by Neil Longley York. Isis, 77 (3):571–572, September 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231668.

Hendry:1986:BRB

[1935] John Hendry. Book review: Memoirs of a Computer Pioneer by Maurice Wilkes. Isis, 77(3):572–573, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231669.

Anonymous:1986:BMc

[1936] Anonymous. Back matter. Isis, 77(3):574–578, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231670.

Anonymous:1986:FMd

[1937] Anonymous. Front matter. Isis, 77(4):579–583, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233162.

Hannaway:1986:LDA

[1938] Owen Hannaway. Laboratory design and the aim of science: Andreas Libavius versus Tycho Brahe. Isis, 77(4):584–610, December 1986. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233163.

Servos:1986:MPS

[1939] John W. Servos. Mathematics and the physical sciences in America, 1880–1930. Isis, 77(4):611–629, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233164.

Aronson:1986:DRH

[1940] Naomi Aronson. The discovery of resistance historical accounts and scientific careers. Isis, 77(4):630–646, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233165.

Perrin:1986:LTC

[1941] C. E. Perrin and Antoine Laurent Lavoisier. Lavoisier’s thoughts on calcination and combustion, 1772–1773. Isis, 77(4):647–666, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233166.

Mahoney:1986:EKV

[1942] Michael S. Mahoney and Ivo Schneider. Eloge: Kurt Vogel, 30 September 1888–27 October 1985. Isis, 77(4):667–669, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233167.

Forrester:1986:BRB

[1943] John Forrester. Book review: The Foundations of Psychoanalysis: A Philosophical Critique. Isis, 77(4):670–674, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233168.

Cowan:1986:BRB

[1944] Ruth Schwartz Cowan. Book review: Reflections on Gender and Science. Isis, 77(4):674–676, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233169.

Woodward:1986:BRB

[1945] William R. Woodward. Book review: Natur und Erfahrung: Von der mittelalterlichen zur neuzeitlichen Naturwissenschaft by Michael Heidelberger; Sigrun Thiessen. Isis, 77(4):677–678, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233170.

Crowe:1986:BRB

[1946] Michael J. Crowe. Book review: Breakthroughs: A Chronology of Great Achievements in Science and Mathematics, 1200–1930 by Claire L. Parkinson. Isis, 77(4):678–679, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233171.

DeVorkin:1986:BRB

[1947] David DeVorkin. Book review: Astronomy and Astronautics: An Enthusiast’s Guide to Books and Periodicals by Andy Lusis. Isis, 77(4): 679–680, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233172.

Stannard:1986:BRB

[1948] David E. Stannard. Book review: Historiography and Causation in Psychoanalysis: An Essay on Psychoanalytic and Historical Epistemology by Edwin R. Wallace. Isis, 77(4):680–681, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233173.

Cohen:1986:BRB

[1949] I. Bernard Cohen. Book review: Mind from Matter? An Essay on Evolutionary Epistemology by Max Delbruck; Gunther S. Stent; Ernst Peter Fischer; Solomon W. Golomb; David Presti; Hansjakob Seiler. Isis, 77 (4):681–683, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233174.

Blackwell:1986:BRB

[1950] Richard J. Blackwell. Book review: Scientific Knowledge and Philosophic Thought by Harold Himsworth. Isis, 77(4):683, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233175.

Kuklick:1986:BRB

[1951] Bruce Kuklick. Book review: Writings of Charles S. Peirce: A Chronological Edition Volume III:1872–1878 by Charles S. Peirce; Christian J. W. Kloesel. Isis, 77(4):683, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233176.

Stock:1986:BRB

[1952] John T. Stock. Book review: Nineteenth-Century Scientific Instruments and Their Makers by P. R. de Clercq. Isis, 77(4):684–685, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233177.

VanHelden:1986:BRB

[1953] Albert Van Helden. Book review: The Cosmic Inquirers: Modern Telescopes and Their Makers by Wallace Tucker; Karen Tucker. Isis, 77 (4):685–686, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233178.

Brown:1986:BRBb

[1954] Theodore M. Brown. Book review: A System of Scientific Medicine: Philanthropic Foundations in the Flexner Era by Howard S. Berliner. Isis, 77(4):686–687, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233179.

Young:1986:BRB

[1955] James Harvey Young. Book review: Fair Play in the Marketplace: The First Battle for Pure Food and Drugs by Mitchell Okun. Isis, 77(4):

687–688, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233180.

Reeds:1986:BRBb

[1956] Karen Reeds. Book review: Medieval Medical Miniatures by Peter Murray Jones; Ars Medica: Art, Medicine, and the Human Condition by Diane R. Karp. Isis, 77(4):688–690, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233181.

Lind:1986:BRB

[1957] L. R. Lind. Book review: Art and Anatomy in Renaissance Italy by Bernard Schulz. Isis, 77(4):690, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233182.

Eamon:1986:BRBb

[1958] William Eamon. Book review: Nicolas Chuquet, Renaissance Mathematician: A Study with Extensive Translation of Chuquet’s Mathematical Manuscript Completed in 1484 by Graham Flegg; Cynthia Hay; Barbara Moss. Isis, 77(4):690–691, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233183.

Dauben:1986:BRBb

[1959] Joseph W. Dauben. Book review: Collected Works Volume I:Publications, 1929–1936 by Kurt Godel; Solomon Feferman; John W. Dawson; Stephen C. Kleene; Gregory H. Moore; Robert M. Solovay; Jean van Heijenoort. Isis, 77(4):691–692, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233184.

Hacking:1986:BRBb

[1960] Ian Hacking. Book review: Observation, Experiment, and Hypothesis in Modern Physical Science by Peter Achinstein; Owen Hannaway. Isis, 77 (4):692–693, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233185.

Wilson:1986:BRBa

[1961] Curtis Wilson. Book review: Cometary Theory in Fifteenth-Century Europe by Jane L. Jervis. Isis, 77(4):693–694, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233186.

Stuewer:1986:BRB

[1962] Roger H. Stuewer. Book review: Frederick Soddy (1877–1956): Early Pioneer in Radiochemistry by George B. Kauffman. Isis, 77(4):694–695, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233187.

Wheaton:1986:BRBb

[1963] Bruce R. Wheaton. Book review: History of Physics by Spencer R. Weart; Melba Phillips. Isis, 77(4):695–696, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233188.

Newcomb:1986:BRB

[1964] Sally Newcomb. Book review: Geologists and Ideas: A History of North American Geology by E. T. Drake; William Jordan. Isis, 77(4):696–697, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233189.

Kenzer:1986:BRB

[1965] Martin S. Kenzer. Book review: On Geography: And Its History by D. R. Stoddart. Isis, 77(4):697–698, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233190.

Limoges:1986:BRB

[1966] Camille Limoges. Book review: De Darwin au darwinisme: Science et ideologie by Yvette Conry. Isis, 77(4):698–700, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233191.

Haraway:1986:BRB

[1967] Donna Haraway. Book review: The Social Meaning of Modern Biology: From Social Darwinism to Sociobiology by Howard L. Kaye. Isis, 77 (4):700–701, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233192.

Tang:1986:BRB

[1968] Paul C. L. Tang. Book review: Evolution as a Religion: Strange Hopes and Stranger Fears by Mary Midgley. Isis, 77(4):701–702, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233193.

Ash:1986:BRB

[1969] Mitchell G. Ash. Book review: Die Professionalisierung der deutschen Psychologie im Nationalsozialismus by Ulfried Geuter. Isis, 77(4):702– 704, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233194.

Handler:1986:BRB

[1970] Richard Handler. Book review: Anthropology as Cultural Critique: An Experimental Moment in the Human Sciences by George E. Marcus; Michael M. J. Fischer. Isis, 77(4):704–705, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233195.

Feuerwerker:1986:BRB

[1971] Elie Feuerwerker. Book review: On the Causes of Fevers (1839) by William Budd; Dale C. Smith. Isis, 77(4):705, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233196.

Bearman:1986:BRB

[1972] David Bearman. Book review: The History of Scurvy and Vitamin C. by Kenneth J. Carpenter. Isis, 77(4):705–706, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233197.

Fleming:1986:BRB

[1973] Donald Fleming. Book review: Penicillin: Meeting the Challenge by Gladys L. Hobby. Isis, 77(4):706–707, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233198.

Epstein:1986:BRB

[1974] Julia L. Epstein. Book review: The Captured Womb: A History of the Medical Care of Pregnant Women by Ann Oakley. Isis, 77(4):707–708, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233199.

Maulitz:1986:BRB

[1975] Russell Maulitz. Book review: Hematology, the Blossoming of a Science: A Story of Inspiration and Effort by Maxwell M. Wintrobe. Isis, 77 (4):708–709, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233200.

Scarborough:1986:BRBc

[1976] John Scarborough. Book review: Arcana mundi: Magic and the Occult in the Greek and Roman Worlds: A Collection of Ancient Texts by Georg Luck. Isis, 77(4):709–710, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233201.

Sylla:1986:BRB

[1977] Edith D. Sylla. Book review: On Maxima and Minima: Chapter 5 of Rules for Solving Sophismata, with an Anonymous Fourteenth-Century Discussion by William Heytesbury; John Longeway. Isis, 77(4):710–711, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233202.

King:1986:BRB

[1978] David A. King. Book review: Planispheric Astrolabes from the National Museum of American History by Sharon Gibbs; George Saliba. Isis, 77 (4):711–713, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233203.

Mahoney:1986:BRBb

[1979] Michael S. Mahoney. Book review: L’oeuvre algebrique d’al-Khayyam by Al-Khayyam; Roshdi Rashed; Ahmad Djebbar. Isis, 77(4):713, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233204.

Tee:1986:BRB

[1980] G. J. Tee. Book review: The Algebra Geometrica of Paolo Bonasoni, ca. 1575: Being the Only Known Work of This Nearly Forgotten Renaissance Mathematician by Paolo Bonasoni; Robert Schmidt. Isis, 77(4):713–714, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233205.

Aiton:1986:BRB

[1981] E. J. Aiton. Book review: Complete Works. Volume III: Minor Works by Nicholas Copernicus; Pawel Czartoryski; Edward Rosen; Erna Hilfstein. Isis, 77(4):714–715, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233206.

Osler:1986:BRBb

[1982] Margaret J. Osler. Book review: Ideas, Qualities, and Corpuscles: Locke and Boyle on the External World by Peter Alexander. Isis, 77(4):715– 716, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233207.

Hankins:1986:BRB

[1983] Thomas L. Hankins. Book review: Opera omnia. Series IV A: Commercium epistolicum. Volume VI: Correspondance de Leonhard Euler avec P.-L. M. de Maupertuis et Frederic II by Leonhard Euler; Pierre Costabel; Eduard Winter; Asot T. Grigorijan; Adolph P. Juskevic; Emil A. Fellmann. Isis, 77(4):716–717, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233208.

Dobbs:1986:BRB

[1984] B. J. T. Dobbs. Book review: Physics at the Royal Society during Newton’s Presidency by J. L. Heilbron. Isis, 77(4):717–718, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233209.

Fix:1986:BRB

[1985] Andrew Fix. Book review: Religious Origins of Modern Science: Belief in Creation in Seventeenth-Century Thought by Eugene M. Klaaren. Isis, 77(4):718–719, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233210.

Jacob:1986:BRBb

[1986] Margaret C. Jacob. Book review: Leviathan and the Air Pump: Hobbes, Boyle, and the Experimental Life by Steven Shapin; Simon Schaffer. Isis, 77(4):719–720, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233211.

Beddall:1986:BRB

[1987] Barbara G. Beddall. Book review: Sir Francis Galton, padre de la eugenesia by Raquel Alvarez Pelaez. Isis, 77(4):720–721, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233212.

Miller:1986:BRBb

[1988] David Philip Miller. Book review: Faraday Rediscovered: Essays on the Life and Work of Michael Faraday, 1791–1867 by David Gooding; Frank A. J. L. James. Isis, 77(4):721–722, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233213.

Williams:1986:BRBc

[1989] L. Pearce Williams. Book review: Wranglers and Physicists: Studies on Cambridge Mathematical Physics in the Nineteenth Century by P. M. Harman. Isis, 77(4):722–723, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233214.

Gregory:1986:BRBb

[1990] Frederick Gregory. Book review: Samuel Thomas Soemmering und die Gelehrten der Goethezeit: Beitrage eines Symposiums in Mainz vom 19. bis 21. Mai 1983 by Gunter Mann; Franz Dumont. Isis, 77(4):723–724, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233215.

Wilson:1986:BRBb

[1991] Andrew D. Wilson. Book review: Freud’s Discovery of Psychoanalysis: The Politics of Hysteria by William J. McGrath. Isis, 77(4):724–725, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233216.

Sherman:1986:BRB

[1992] Paul D. Sherman. Book review: “Martyr of Science”: Sir David Brewster, 1781–1868 by A. D. Morrison-Low; J. R. R. Christie. Isis, 77(4): 725–726, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233217.

Porter:1986:BRBg

[1993] Charlotte M. Porter. Book review: The Fort Tejon Letters, 1857–1859 by John Xantus; Ann Zwinger. Isis, 77(4):726–727, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233218.

Turner:1986:BRBc

[1994] Frank M. Turner. Book review: Darwin’s Metaphor: Nature’s Place in Victorian Culture by Robert M. Young. Isis, 77(4):727–728, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233219.

Kimmelman:1986:BRB

[1995] Barbara A. Kimmelman. Book review: A Gardener Touched with Genius: The Life of Luther Burbank by Peter Dreyer. Isis, 77(4):728–729, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233220.

Trenn:1986:BRBb

[1996] Thaddeus J. Trenn. Book review: Presidential Management of Science and Technology: The Johnson Presidency by W. Henry Lambright. Isis, 77(4):729–730, December 1986. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233221.

Harwit:1986:BRB

[1997] Martin Harwit. Book review: Early History of Cosmic Ray Studies: Personal Reminiscences with Old Photographs by Yataro Sekido; Harry Elliot. Isis, 77(4):730–731, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233222.

Anonymous:1986:BMd

[1998] Anonymous. Back matter. Isis, 77(4):732–750, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233223.

Anonymous:1986:VI

[1999] Anonymous. Volume information. Isis, 77(4):736–749, December 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233161.

Anonymous:1986:FMe

[2000] Anonymous. Front matter. Isis, 77(5):1–4, ???? 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232406.

Neu:1986:OHE

[2001] John Neu. One hundred eleventh critical bibliography of the history of science and its cultural influences. Isis, 77(5):5–265, ???? 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232407.

Anonymous:1986:BMe

[2002] Anonymous. Back matter. Isis, 77(5):266, ???? 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232408.

Anonymous:1987:FMa

[2003] Anonymous. Front matter. Isis, 78(1):1–5, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232726.

Richards:1987:AMH

[2004] Joan L. Richards. Augustus De Morgan, the history of mathematics, and the foundations of algebra. Isis, 78(1):6–30, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232727.

Rabkin:1987:TIS

[2005] Yakov M. Rabkin. Technological innovation in science: The adoption of infrared spectroscopy by chemists. Isis, 78(1):31–54, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232728.

McMullin:1987:BC

[2006] Ernan McMullin. Bruno and Copernicus. Isis, 78(1):55–74, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232729.

Donovan:1987:SSP

[2007] Arthur Donovan, George W. Stocking, Jr., and David Leary. Symposia and study programs. Isis, 78(1):75–79, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232730.

Webster:1987:ECB

[2008] Charles Webster. Eloge: Charles Bernard Schmitt (1933–1986). Isis, 78 (1):80–81, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232731.

Hannaway:1987:BRB

[2009] Owen Hannaway. Book review: Between the Library and the Laboratory: The Language of Chemistry in Eighteenth-Century France. Isis, 78(1): 82–85, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232732.

Reeds:1987:BRBa

[2010] Karen Reeds. Book review: Dioscorides on Pharmacy and Medicine. Isis, 78(1):85–88, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232733.

Steneck:1987:BRB

[2011] Nicholas Steneck. Book review: The Light of Nature: Essays in the History and Philosophy of Science Presented to A. C. Crombie by J. D. North; J. J. Roche. Isis, 78(1):89, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232734.

Foley:1987:BRBa

[2012] Vernard Foley. Book review: Bronze Age, Greek and Roman Technology: A Select, Annotated Bibliography by John Peter Oleson. Isis, 78(1):89– 90, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232735.

Cutcliffe:1987:BRBa

[2013] Stephen Cutcliffe. Book review: Accounts of European Science, Technology, and Medicine Written by American Travelers Abroad, 1735–1860, in the Collections of the American Philosophical Society by Darwin Stapleton. Isis, 78(1):90–91, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232736.

Morman:1987:BRBa

[2014] Edward T. Morman. Book review: The Truman G. Blocker, Jr., History of Medicine Collections: Books and Manuscripts by Larry J. Wygant. Isis, 78(1):92, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232737.

Porter:1987:BRBa

[2015] Theodore M. Porter. Book review: Who’s Who in Economics: A Biographical Dictionary of Major Economists, 1700–1986 by Mark Blaug. Isis, 78(1):92–93, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232738.

Brooke:1987:BRB

[2016] John H. Brooke. Book review: The Politics and Rhetoric of Scientific Method: Historical Studies by John A. Schuster; Richard R. Yeo. Isis, 78(1):93–94, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232739.

Divine:1987:BRB

[2017] Robert A. Divine. Book review: Ban the Bomb: A History of SANE, the Committee for a Sane Nuclear Policy; 1957–1985 by Milton Katz. Isis, 78(1):94–95, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232740.

Tatarewicz:1987:BRBa

[2018] Joseph N. Tatarewicz. Book review: The University of Arizona’s Lunar and Planetary Laboratory: Its Founding and Early Years by Ewen A. Whitaker. Isis, 78(1):95–96, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232741.

Miller:1987:BRBa

[2019] Jane A. Miller. Book review: Hypatia’s Heritage: A History of Women in Science from Antiquity to the Nineteenth Century by Margaret Alic. Isis, 78(1):96–97, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232742.

Florman:1987:BRB

[2020] Samuel C. Florman. Book review: Social Responses to Technological Change by Augustine Brannigan; Sheldon Goldenberg. Isis, 78(1):97– 98, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232743.

Cantelon:1987:BRB

[2021] Philip L. Cantelon. Book review: Insuring against Disaster: The Nuclear Industry on Trial by John W. Johnson. Isis, 78(1):98–99, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232744.

Lewenstein:1987:BRB

[2022] Bruce Lewenstein. Book review: The Nemesis Affair: A Story of the Death of Dinosaurs and the Ways of Science by David M. Raup. Isis, 78(1):99–100, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232745.

Weininger:1987:BRB

[2023] Stephen J. Weininger. Book review: Languages of Nature: Critical Essays on Science and Literature by Ludmilla Jordanova. Isis, 78(1):100– 101, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232746.

Smith:1987:BRB

[2024] Robert W. Smith. Book review: Cosmological Constants: Papers in Modern Cosmology by Jeremy Bernstein; Gerald Feinberg. Isis, 78(1): 101, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232747.

Hilfstein:1987:BRBa

[2025] Erna Hilfstein. Book review: Las quimeras de los cielos: Aspectos epistemologicos de la revolucion copernicana by Alberto Elena. Isis, 78(1): 101–102, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232748.

Furukawa:1987:BRBa

[2026] Yasu Furukawa. Book review: History of Polyolefins: The World’s Most Widely Used Polymers by Raymond B. Seymour; Tai Cheng. Isis, 78(1): 102–103, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232749.

Dunbar:1987:BRBa

[2027] G. S. Dunbar. Book review: A Short History of Geomorphology by Keith J. Tinkler. Isis, 78(1):103, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232750.

Greene:1987:BRBa

[2028] Mott T. Greene. Book review: The Dark Side of the Earth by Robert Muir Wood. Isis, 78(1):103–105, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232751.

Martin:1987:BRB

[2029] Calvin Martin. Book review: Ecological Imperialism: The Biological Expansion of Europe, 900–1900 by Alfred W. Crosby. Isis, 78(1):105– 106, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232752.

Hagan:1987:BRBa

[2030] William J. Hagan, Jr. Book review: Microcosmos: Four Billion Years of Microbial Evolution by Lynn Margulis; Dorion Sagan; Origins: A Skeptic’s Guide to the Creation of Life on Earth by Robert Shapiro. Isis, 78(1):106–107, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232753.

Woodward:1987:BRBa

[2031] William R. Woodward. Book review: Urmanuskript “Über das Gedachtniss” 1880 by Hermann Ebbinghaus; Experimental-psychologische Apparate und Methoden: Die Austellung bei dem 1. Kongress für experimentelle Psychologie 1904 by Robert Sommer; Geschichte für die Gegenwart: Vortrage und Aufsatze zur Psychologiegeschichte by Werner Traxel. Isis, 78(1):108, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232754.

Gieryn:1987:BRB

[2032] Thomas F. Gieryn. Book review: Journal of Social Issues, Volume 42, Number 1. Fifty Years of Psychology and Social Issues by Benjamin Harris; Rhoda K. Unger; Ross Stagner. Isis, 78(1):109–110, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232755.

Mark:1987:BRBa

[2033] Joan Mark. Book review: History, Evolution, and the Concept of Culture: Selected Papers by Alexander Lesser; Sidney W. Mintz. Isis, 78(1):110– 111, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232756.

Numbers:1987:BRB

[2034] Ronald L. Numbers. Book review: Masters of Madness: Social Origins of the American Psychiatric Profession by Constance M. McGovern. Isis, 78(1):111–112, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232757.

Barry:1987:BRB

[2035] Jonathan Barry. Book review: Medicine and Industrial Society: A History of Hospital Development in Manchester and Its Region, 1752– 1946 by John V. Pickstone. Isis, 78(1):112–113, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232758.

Pickstone:1987:BRB

[2036] J. V. Pickstone. Book review: Hospital Life in Enlightenment Scotland: Care and Teaching at the Royal Infirmary of Edinburgh by Guenter B. Risse. Isis, 78(1):113, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232759.

Porter:1987:BRBb

[2037] Roy Porter. Book review: The Age of Agony: The Art of Healing, c. 1700–1800 by Guy Williams. Isis, 78(1):113–114, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232760.

Mahoney:1987:BRBa

[2038] Michael S. Mahoney. Book review: IBM’s Early Computers by Charles J. Bashe; Lyle R. Johnson; John H. Palmer; Emerson W. Pugh. Isis, 78(1):114–115, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232761.

Wise:1987:BRB

[2039] George Wise. Book review: Imagining Tomorrow: History, Technology, and the American Future by Joseph J. Corn. Isis, 78(1):115–117, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232762.

Israel:1987:BRB

[2040] Paul B. Israel. Book review: The History of N. V. Philips’ Gloelampenfabrieken. Volume I: The Origin of the Dutch Incandescent Lamp Industry by A. Heerding; Derek S. Jordan. Isis, 78(1):117–118, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232763.

Condit:1987:BRB

[2041] Carl W. Condit. Book review: The Railway Station: A Social History by Jeffrey Richards; John M. MacKenzie. Isis, 78(1):118–119, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232764.

Cowan:1987:BRBa

[2042] Ruth Schwartz Cowan. Book review: The Whale and the Reactor: A Search for Limits in an Age of High Technology by Langdon Winner. Isis, 78(1):119–120, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232765.

Jones:1987:BRBa

[2043] Alexander Jones. Book review: Le “Grand commentaire” de Theon d’Alexandrie aux Tables faciles de Ptolemee, Livre I: Histoire du texte, edition critique, traduction by Theon of Alexandria; Joseph Mogenet; Anne Tihon. Isis, 78(1):120–121, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232766.

Tachau:1987:BRB

[2044] Katherine H. Tachau. Book review: The Physics of William of Ockham by Andre Goddu. Isis, 78(1):121–122, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232767.

Hall:1987:BRBa

[2045] Bert S. Hall. Book review: Science and Technology in Medieval Society by Pamela O. Long. Isis, 78(1):122–124, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232768.

Taylor:1987:BRBa

[2046] Jerome Taylor. Book review: The Intellectual Revolution in Twelfth-Century Europe by Tina Stiefel. Isis, 78(1):124–126, March 1987. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232769.

Bowman:1987:BRB

[2047] Inci A. Bowman. Book review: Symposium on Byzantine Medicine by John Scarborough. Isis, 78(1):126, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232770.

Glick:1987:BRB

[2048] Thomas F. Glick. Book review: Western Science in the Arab World: The Impact of Darwinism, 1860–1930 by Adel A. Ziadat. Isis, 78(1): 126–127, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232771.

Scarborough:1987:BRBa

[2049] John Scarborough. Book review: Religious Healing in the Veda by Kenneth G. Zysk. Isis, 78(1):127–128, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232772.

Garber:1987:BRB

[2050] Daniel Garber. Book review: Discours de la methode pour bien conduire sa raison et chercher la verite dans les sciences, plus la Dioptrique, les Meteores, et la Geometrie qui son des essais de cete methode by Rene Descartes; Jean-Robert Armogathe; Vincent Carraud. Isis, 78(1):128– 129, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232773.

Jardine:1987:BRB

[2051] Lisa Jardine. Book review: Francis Bacon: Terminologia e fortuna nel XVII secolo by Marta Fattori. Isis, 78(1):129–130, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232774.

Jacob:1987:BRB

[2052] James R. Jacob. Book review: Guilford Review, Spring 1986, Volume 23, Number 1. Perspectives on the Seventeenth-Century World of Viscountess Anne Conway by Carol Stoneburner; O. Theodor Benfey; Robert Kraus. Isis, 78(1):130–131, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232775.

Osler:1987:BRBa

[2053] Margaret J. Osler. Book review: Causa e spiegazione: La fisica de Pierre Gassendi by Marco Messeri. Isis, 78(1):131, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232776.

Lundgren:1987:BRB

[2054] Anders Lundgren. Book review: Torbern Bergman: A Man Before His Time by J. A. Schufle. Isis, 78(1):131–132, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232777.

Gregory:1987:BRBa

[2055] Frederick Gregory. Book review: The Other Mind of Europe: Goethe as a Scientist by J. P. S. Uberoi. Isis, 78(1):132, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232778.

Melosh:1987:BRBa

[2056] Barbara Melosh. Book review: Florence Nightingale: Saint, Reformer or Rebel? by Raymond G. Hebert. Isis, 78(1):132–133, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232779.

Kremer:1987:BRB

[2057] Richard L. Kremer. Book review: The Papers of Joseph Henry. Volume V: The Princeton Years: January 1841–December 1843 by Joseph Henry; Nathan Reingold; Marc Rothenberg; Kathleen W. Dorman; Paul H. Theerman; Arthur P. Molella; Joan F. Steiner. Isis, 78(1):133–134, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232780.

Munday:1987:BRB

[2058] Pat Munday. Book review: The Letters from Gerrit Jan Mulder to Justus Liebig (1838–1846) by Gerrit Jan Mulder; H. A. M. Snelders. Isis, 78(1): 135, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232781.

Hoopes:1987:BRB

[2059] James Hoopes. Book review: William James: His Life and Thought by Gerald E. Myers. Isis, 78(1):135–137, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232782.

Rigden:1987:BRB

[2060] John S. Rigden. Book review: Reflections of a Physicist by Anatole Abragam. Isis, 78(1):137, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232783.

Swann:1987:BRB

[2061] John P. Swann. Book review: The Life of Ernst Chain: Penicillin and Beyond by Ronald W. Clark. Isis, 78(1):137–138, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232784.

Allison:1987:BRBa

[2062] David K. Allison. Book review: Seek and Strike: Sonar, Anti-Submarine Warfare and the Royal Navy, 1914–1954 by Willem Hackmann. Isis, 78(1):138–139, March 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232785.

Hughes:1987:BRB

[2063] Sally Smith Hughes. Book review: Radiant Science, Dark Politics: A Memoir of the Nuclear Age by Martin D. Kamen. Isis, 78(1):139–140, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232786.

Anonymous:1987:BMa

[2064] Anonymous. Back matter. Isis, 78(1):141–144, March 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232787.

Anonymous:1987:FMb

[2065] Anonymous. Front matter. Isis, 78(2):145–151, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231519.

Grant:1987:COL

[2066] Edward Grant. Celestial orbs in the Latin middle ages. Isis, 78(2):152– 173, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231520.

Schiebinger:1987:MWB

[2067] Londa Schiebinger. Maria winkelmann at the Berlin academy: A turning point for women in science. Isis, 78(2):174–200, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231521.

Petit:1987:CBH

[2068] Annie Petit. Claude bernard and the history of science. Isis, 78(2):201– 219, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231522.

Holmes:1987:SWS

[2069] Frederic L. Holmes. Scientific writing and scientific discovery. Isis, 78(2): 220–235, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231523.

Daston:1987:AMH

[2070] Lorraine J. Daston, Michael S. Mahoney, Edith D. Sylla, and Frederick Gregory. Annual meeting of the History of Science Society 23–26 October 1986. Isis, 78(2):236–239, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231524.

Coleman:1987:SMC

[2071] William Coleman. 1986 Sarton Medal citation. Isis, 78(2):239–241, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231525.

Anonymous:1987:DSF

[2072] Anonymous. Donors to and supporters of the fund drive. Isis, 78(2): 241–243, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231526.

Buchwald:1987:BRB

[2073] Jed Z. Buchwald. Book review: Intellectual Mastery of Nature: Theoretical Physics from Ohm to Einstein. Isis, 78(2):244–249, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231527.

Montgomery:1987:BRBa

[2074] William Montgomery. Book review: The Darwinian Heritage. Isis, 78(2): 249–252, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231528.

Allison:1987:BRBb

[2075] David K. Allison. Book review: Public History: An Introduction by Barbara J. Howe; Emory L. Kemp. Isis, 78(2):253, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231529.

Bearman:1987:BRB

[2076] David Bearman. Book review: Essays of an Information Scientist. Volume VIII: 1985: Ghostwriting and Other Essays by Eugene Garfield. Isis, 78(2):253–254, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231530.

Benfey:1987:BRB

[2077] Otto Theodor Benfey. Book review: Diamond Dealers and Feather Merchants: Tales from the Sciences by Irving M. Klotz. Isis, 78(2):254–255, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231531.

Restivo:1987:BRB

[2078] Sal Restivo. Book review: Dialogues with Scientists and Sages: The Search for Unity by Renee Weber. Isis, 78(2):255–256, June 1987. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231532.

Sharrer:1987:BRB

[2079] G. Terry Sharrer. Book review: Bibliography of Publications by the Faculty, Staff, and Students of the University of California, 1876–1980, on Grapes, Wines, and Related Subjects by Maynard A. Amerine; Herman Phaff. Isis, 78(2):256, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231533.

Koppes:1987:BRB

[2080] Clayton R. Koppes. Book review: The Frank J. Malina Collection at the California Institute of Technology: Guide to a Microfiche Edition by Judith R. Goodstein; Carol H. Buge. Isis, 78(2):256–257, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231534.

Stroup:1987:BRB

[2081] Alice Stroup. Book review: Herbals and Closely Related Medico-Botanical Works, 1472–1753 by Sally Haines Hocker. Isis, 78(2):257, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231535.

Dick:1987:BRBa

[2082] Steven J. Dick. Book review: The Extraterrestrial Life Debate, 1750– 1900: The Idea of a Plurality of Worlds from Kant to Lowell by Michael Crowe. Isis, 78(2):257–259, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231536.

Lyon:1987:BRBa

[2083] John Lyon. Book review: The Identification of Progress in Learning by Torsten Hagerstrand. Isis, 78(2):259, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231537.

Joy:1987:BRB

[2084] Lynn S. Joy. Book review: Descartes’s Gambit by Peter J. Markie. Isis, 78(2):259–260, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231538.

McMullin:1987:BRB

[2085] Ernan McMullin. Book review: Is Science Progressive? by Ilkka Niiniluoto. Isis, 78(2):260–261, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231539.

Miller:1987:BRBb

[2086] David Philip Miller. Book review: The Botanists: A History of the Botanical Society of the British Isles through a Hundred and Fifty Years by David Elliston Allen. Isis, 78(2):261–263, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231540.

Kohler:1987:BRBa

[2087] Robert E. Kohler. Book review: Inventing the NIH: Federal Biomedical Research Policy, 1887–1937 by Victoria A. Harden. Isis, 78(2):263–264, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231541.

Lewis:1987:BRBa

[2088] Jane Lewis. Book review: Health Policies, Health Politics: The British and American Experience, 1911–1965 by Daniel M. Fox. Isis, 78(2):264– 265, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231542.

Kuklick:1987:BRB

[2089] Henrika Kuklick. Book review: Professional Powers: A Study of the Institutionalization of Formal Knowledge by Eliot Freidson. Isis, 78(2): 265–266, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231543.

Sorensen:1987:BRB

[2090] Knut H. Sorensen. Book review: Radical Science Essays by Les Levidow. Isis, 78(2):266–267, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231544.

Durant:1987:BRB

[2091] John R. Durant. Book review: The Evolution of Darwin’s Religious Views by Frank Burch Brown. Isis, 78(2):267–268, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231545.

Goldberg:1987:BRB

[2092] Stanley Goldberg. Book review: Einstein as Myth and Muse by Alan J. Friedman; Carol C. Donley. Isis, 78(2):268–269, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231546.

Turner:1987:BRB

[2093] Frank M. Turner. Book review: God and Nature: Historical Essays on the Encounter between Christianity and Science by David C. Lindberg; Ronald L. Numbers. Isis, 78(2):269–270, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231547.

Moore:1987:BRB

[2094] James R. Moore. Book review: Evolution and Creation by Ernan Mc-Mullin. Isis, 78(2):270–271, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231548.

Hilfstein:1987:BRBb

[2095] Erna Hilfstein. Book review: Mathemata: Festschrift für Helmuth Gericke by Menso Folkerts; Uta Lindgren. Isis, 78(2):271–272, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231549.

Daston:1987:BRB

[2096] Lorraine Daston. Book review: The Rise of Statistical Thinking, 1820– 1900 by Theodore M. Porter. Isis, 78(2):272–274, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231550.

Cartwright:1987:BRB

[2097] Nancy Cartwright. Book review: The Shaky Game: Einstein, Reality, and the Quantum Theory by Arthur Fine. Isis, 78(2):274–275, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231551.

Warner:1987:BRB

[2098] Deborah Jean Warner. Book review: Short Guide to Modern Star Names and Their Derivations by Paul Kunitzsch; Tim Smart. Isis, 78(2):275, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231552.

Friedel:1987:BRB

[2099] Robert Friedel. Book review: Polymer Pioneers: A Popular History of the Science and Technology of Large Molecules by Peter J. T. Morris.

Isis, 78(2):275–276, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231553.

Cahan:1987:BRB

[2100] David Cahan. Book review: Physics for a New Century: Papers Presented at the 1904 St. Louis Congress by Katherine R. Sopka. Isis, 78(2): 276–277, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231554.

Multhauf:1987:BRB

[2101] R. P. Multhauf. Book review: Fuller’s Earth: A History of Calcium Montmorillonite by Robert H. S. Robertson. Isis, 78(2):277–278, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231555.

Reeds:1987:BRBb

[2102] Karen Reeds. Book review: Printmaking in the Service of Botany: Catalogue of an Exhibition by Gavin D. R. Bridson; Donald E. Wendel; James M. White. Isis, 78(2):278, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231556.

Pointon:1987:BRB

[2103] Marcia Pointon. Book review: Catalogue of the Natural History Drawings Commissioned by Joseph Banks on the Endeavour Voyage, 1768–1771, Held in the British Museum (Natural History). Part 3: Zoology by Alwyne Wheeler. Isis, 78(2):278–280, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231557.

Kutzbach:1987:BRB

[2104] Gisela Kutzbach. Book review: Namias Symposium by John O. Roads. Isis, 78(2):280–281, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231558.

Goetzmann:1987:BRB

[2105] William H. Goetzmann. Book review: Magnificent Voyagers: The U.S. Exploring Expedition, 1838–1842 by Herman J. Viola; Carolyn Margolis. Isis, 78(2):281–282, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231559.

Triplet:1987:BRB

[2106] Rodney G. Triplet. Book review: One Hundred Years of Psychological Research in America: G. Stanley Hall and the Johns Hopkins Tradition by Stewart H. Hulse; Bert F. Green. Isis, 78(2):282–283, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231560.

Lomax:1987:BRB

[2107] Elizabeth Lomax. Book review: History and Research in Child Development by Alice Boardman Smuts; John W. Hagen. Isis, 78(2):283, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231561.

Sica:1987:BRB

[2108] Alan Sica. Book review: Robert K. Merton: An Intellectual Portrait by Piotr Sztompka. Isis, 78(2):283–285, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231562.

Fissell:1987:BRBa

[2109] Mary E. Fissell. Book review: No Magic Bullet: A Social History of Venereal Disease in the United States since 1800, with a New Chapter on AIDS by Allan M. Brandt. Isis, 78(2):285, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231563.

Bynum:1987:BRB

[2110] W. F. Bynum. Book review: The Decline of the Old Medical Regime in Stuart London by Harold J. Cook. Isis, 78(2):285–286, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231564.

Verbrugge:1987:BRB

[2111] Martha H. Verbrugge. Book review: “Hydropathic Highway to Health”: Women and Water-Cure in Antebellum America by Jane B. Donegan. Isis, 78(2):286–288, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231565.

Fee:1987:BRB

[2112] Elizabeth Fee. Book review: In the Patient’s Best Interest: Women and the Politics of Medical Decisions by Sue Fisher. Isis, 78(2):288, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231566.

Andrews:1987:BRB

[2113] Margaret W. Andrews. Book review: Private Practice, Public Payment: Canadian Medicine and the Politics of Health Insurance, 1911–1966 by C. David Naylor. Isis, 78(2):288–289, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231567.

Rosner:1987:BRBa

[2114] David Rosner. Book review: The Effects of Arts, Trades, and Professions on Health and Longevity by Charles Turner Thackrah. Isis, 78(2):289– 290, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231568.

Donegan:1987:BRBa

[2115] Jane B. Donegan. Book review: Midwives in History and Society by Jean Towler; Joan Bramall. Isis, 78(2):290–291, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231569.

Todd:1987:BRB

[2116] Edmund N. Todd. Book review: Monopoly’s Moment: The Organization and Regulation of Canadian Utilities, 1830–1930 by Christopher Armstrong; H. V. Nelles. Isis, 78(2):291–292, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231570.

Lewis:1987:BRBb

[2117] W. David Lewis. Book review: The Golden Age of the Great Passenger Airships: Graf Zeppelin and Hindenburg by Harold G. Dick; Douglas H. Robinson. Isis, 78(2):292–293, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231571.

Morrow:1987:BRB

[2118] John H. Morrow, Jr. Book review: The Official History of the Royal Canadian Air Force. Volume II: The Creation of a National Air Force by W. A. B. Douglas. Isis, 78(2):293–294, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231572.

Slade:1987:BRB

[2119] Joseph W. Slade. Book review: The Social History of the Machine Gun by John Ellis. Isis, 78(2):294–295, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231573.

Hall:1987:BRBb

[2120] A. Rupert Hall. Book review: Firearms and Fortifications: Military Architecture and Siege Warfare in Sixteenth-Century Siena by Simon Pepper; Nicholas Adams. Isis, 78(2):295–296, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231574.

Cowan:1987:BRBb

[2121] Ruth Schwartz Cowan. Book review: The Whale and the Reactor: A Search for Limits in an Age of High Technology by Langdon Winner. Isis, 78(2):296–297, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231575.

Scarborough:1987:BRBb

[2122] John Scarborough. Book review: Studi su papiri greci di logica e medicina by W. Cavini; M. C. Donnini Maccio; M. S. Funghi; D. Manetti. Isis, 78(2):297–298, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231576.

Foley:1987:BRBb

[2123] Vernard Foley. Book review: Production and Exchange of Stone Tools: Prehistoric Obsidian in the Aegean by Robin Torrence. Isis, 78(2):298– 299, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231577.

Peters:1987:BRB

[2124] Edward Peters. Book review: Medieval Cosmology: Theories of Infinity, Place, Time, Void, and the Plurality of Worlds by Pierre Duhem; Roger Ariew. Isis, 78(2):299–300, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231578.

Gueriguian:1987:BRB

[2125] John L. Gueriguian. Book review: Bark’ Galianosi: The Greek-Armenian Dictionary to Galen by John A. C. Greppin. Isis, 78(2):300– 301, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231579.

Sylla:1987:BRBa

[2126] Edith Sylla. Book review: Pseudo-Aristoteles Latinus: A Guide to Latin Works Falsely Attributed to Aristotle before 1500 by Charles B. Schmitt; Dilwyn Knox. Isis, 78(2):301, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231580.

Vickers:1987:BRB

[2127] Brian Vickers. Book review: Pic de la Mirandole: Contribution a la connaissance de l’humanisme philosophique renaissant by Jacques Queron. Isis, 78(2):301–302, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231581.

Mahoney:1987:BRBb

[2128] Michael S. Mahoney. Book review: The Pendulum Clock, or Geometrical Demonstration concerning the Motion of Pendula as Applied to Clocks by Christiaan Huygens; Richard J. Blackwell. Isis, 78(2):302–303, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231582.

Shea:1987:BRBa

[2129] William R. Shea. Book review: Correspondance du P. Marin Mersenne, religieux minime. Volume XVI: 1648 by P. Marin Mersenne; Cornelis de Waard; Armand Beaulieu. Isis, 78(2):303–304, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231583.

Osler:1987:BRBb

[2130] Margaret J. Osler. Book review: The Natural Philosophy of Leibniz by Kathleen Okruhlik; James Robert Brown. Isis, 78(2):304–305, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231584.

Elliott:1987:BRBa

[2131] J. H. Elliott. Book review: Six Galleons for the King of Spain: Imperial Defense in the Early Seventeenth Century by Carla Rahn Phillips. Isis, 78(2):305–306, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231585.

Lenoir:1987:BRBa

[2132] Timothy Lenoir. Book review: Dokumente einer Freundschaft: Briefwechsel zwischen Hermann von Helmholtz und Emil du Bois-Reymond, 1846–1894 by Hermann von Helmholtz; Emil du Bois-Reymond; Christa Kirsten. Isis, 78(2):306–307, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231586.

Melosi:1987:BRB

[2133] Martin V. Melosi. Book review: The Politics of Energy Research and Development by John Byrne; Daniel Rich. Isis, 78(2):307–308, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231587.

Lowood:1987:BRBa

[2134] Henry Lowood. Book review: Portraits of Success: Impressions of Silicon Valley Pioneers by Carolyn Caddes. Isis, 78(2):308–309, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231588.

Hagan:1987:BRBb

[2135] William J. Hagan, Jr. Book review: Origins of Life by Freeman Dyson. Isis, 78(2):309–310, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231589.

Jones:1987:BRBb

[2136] Daniel P. Jones. Book review: The Poisonous Cloud: Chemical Warfare in the First World War by L. F. Haber. Isis, 78(2):310–311, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231590.

Harris:1987:BRB

[2137] Ruth R. Harris. Book review: Gathering Rare Ores: The Diplomacy of Uranium Acquisition, 1943–1954 by Jonathan E. Helmreich. Isis, 78(2): 311–312, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231591.

Channell:1987:BRB

[2138] David F. Channell. Book review: One Hundred Years of Science and Technology in Texas: A Sigma Xi Centennial Volume by Leo J. Klosterman; Loyd S. Swenson,; Sylvia Rose. Isis, 78(2):312–313, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231592.

Kevles:1987:BRB

[2139] Daniel J. Kevles. Book review: The Militarization of Space: U.S. Policy, 1945–1984 by Paul B. Stares. Isis, 78(2):313–314, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231593.

Koblitz:1987:BRB

[2140] Ann Hibner Koblitz. Book review: Women Scientists from Antiquity to the Present: An Index by Caroline L. Herzenberg; Women in Science, Antiquity through the Nineteenth Century: A Biographical Dictionary with Annotated Bibliography by Marilyn Bailey Ogilvie. Isis, 78(2):315– 316, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231594.

Wheaton:1987:BRB

[2141] Bruce R. Wheaton. Book review: The Bragg Family in Adelaide: A Pictorial Celebration by John Jenkin. Isis, 78(2):316, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231595.

Richards:1987:BRB

[2142] Joan L. Richards. Book review: Alfred North Whitehead: The Man and His Work. Volume I: 1861–1910 by Victor Lowe. Isis, 78(2):317, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231596.

Hoddeson:1987:BRB

[2143] Lillian Hoddeson. Book review: A Life in Science by Nevill Mott. Isis, 78(2):317–319, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231597.

Tee:1987:BRB

[2144] Garry J. Tee. Book review: Ada: A Life and a Legacy by Dorothy Stein. Isis, 78(2):319–320, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231598.

Ewan:1987:BRB

[2145] Joseph Ewan. Book review: Martha Maxwell: Rocky Mountain Naturalist by Maxine Benson. Isis, 78(2):320–321, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231599.

Porter:1987:BRBc

[2146] Charlotte M. Porter. Book review: A Yankee Botanist in the Carolinas: The Reverend Moses Ashley Curtis, D.D. (1808–1872) by Edmund Berkeley; Dorothy Smith Berkeley. Isis, 78(2):321, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231600.

Browne:1987:BRB

[2147] Janet Browne. Book review: Cambridge and Clare by Harry Godwin. Isis, 78(2):322, June 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231601.

Greene:1987:BRBb

[2148] Mott T. Greene. Book review: The Ocean of Truth: A Personal History of Global Tectonics by H. W. Menard. Isis, 78(2):322–324, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231602.

Marvin:1987:BRB

[2149] Ursula B. Marvin. Book review: Alfred Wegener: The Father of Continental Drift by Martin Schwarzbach; Carla Love. Isis, 78(2):324–325, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231603.

Anonymous:1987:BMb

[2150] Anonymous. Back matter. Isis, 78(2):326–330, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231604.

Anonymous:1987:FMc

[2151] Anonymous. Front matter. Isis, 78(3):331–335, September 1987. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231998.

Churchill:1987:HTV

[2152] Frederick B. Churchill. From heredity theory to vererbung: The transmission problem, 1850–1915. Isis, 78(3):336–364, September 1987. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231999.

Nyhart:1987:DBG

[2153] Lynn Nyhart. The disciplinary breakdown of German morphology, 1870– 1900. Isis, 78(3):365–389, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232000.

Harwood:1987:NSS

[2154] Jonathan Harwood. National styles in science: Genetics in Germany and the United States between the World Wars. Isis, 78(3):390–414, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232001.

Ewan:1987:EFV

[2155] Joseph Ewan. Eloge: Frans Verdoorn, 24 July 1906–18 May 1984. Isis, 78 (3):415–416, September 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232002.

Shapin:1987:BRB

[2156] Steven Shapin. Book review: The Correspondence of Henry Oldenburg. Isis, 78(3):417–424, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232003.

Pickering:1987:BRBa

[2157] Andy Pickering. Book review: The Second Creation: Makers of the Revolution in Twentieth–Century Physics. Isis, 78(3):425–428, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232004.

Robinson:1987:BRB

[2158] Eric H. Robinson. Book review: Engines of Change: The American Industrial Revolution, 1790–1860. Isis, 78(3):429–431, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232005.

Fitzgerald:1987:BRB

[2159] Deborah Fitzgerald. Book review: The Day the Universe Changed by James Burke. Isis, 78(3):432–433, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232006.

Hall:1987:BRBc

[2160] Marie Boas Hall. Book review: Le edizioni dei testi filosofici e scientifici del ’500 e del ’600: Problemi di metodo e prospettive di ricerca by Guido Canziani; Gianni Paganini. Isis, 78(3):433–434, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232007.

Mark:1987:BRBb

[2161] Joan Mark. Book review: “The Proper Study of Mankind”: An Annotated Bibliography of Manuscript Sources on Anthropology and Archaeology in the Library of the American Philosophical Society by David K. van Keuren. Isis, 78(3):434–435, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232008.

Tiffany:1987:BRB

[2162] Paul Tiffany. Book review: A Guide to Iron and Steel: Pictures in the Hagley Museum and Library by Jon M. Williams. Isis, 78(3):435–436, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232009.

Manier:1987:BRB

[2163] Edward Manier. Book review: Neurophilosophy: Toward a Unified Science of the Mind-Brain by Patricia Smith Churchland. Isis, 78(3):436– 437, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/232010.

Lugg:1987:BRB

[2164] Andrew Lugg. Book review: The Arch of Knowledge: An Introductory Study of the History of the Philosophy and Methodology of Science by David Oldroyd. Isis, 78(3):437–438, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232011.

Wynne:1987:BRB

[2165] Brian Wynne. Book review: The Concept of Physical Law by Norman Swartz. Isis, 78(3):438–439, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232012.

Anderson:1987:BRB

[2166] R. G. W. Anderson. Book review: The History and Preservation of Chemical Instrumentation by John T. Stock; Mary Virginia Orna. Isis, 78 (3):439–440, September 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232013.

Forgan:1987:BRB

[2167] Sophie Forgan. Book review: The Royal Institution: An Informal History by Gwendy Caroe. Isis, 78(3):440–441, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232014.

Osterbrock:1987:BRB

[2168] Donald E. Osterbrock. Book review: Big and Bright: A History of the McDonald Observatory by David S. Evans; J. Derral Mulholland. Isis, 78 (3):441–442, September 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232015.

Pauly:1987:BRB

[2169] Philip J. Pauly. Book review: Defining Biology: Lectures from the 1890s by Jane Maienschein. Isis, 78(3):442–443, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232016.

Weisz:1987:BRB

[2170] George Weisz. Book review: L’histoire de l’enseignement, XIXe-XXe siècles: Guide du chercheur by Jean–Pierre Briand; Penelope Caspard-Karydis; Jean-Michel Chapoulie; Therese Charmasson; Serge Chassagne; Alain Choppin; Martine Sonnet. Isis, 78(3):443–444, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232017.

Sylla:1987:BRBb

[2171] Edith Sylla. Book review: Logicae artis compendium by Robert Sanderson; E. J. Ashworth. Isis, 78(3):444, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232018.

Lewis:1987:BRBc

[2172] Jane Lewis. Book review: Medical Care, Medical Costs: The Search for a Health Insurance Policy by Rashi Fein. Isis, 78(3):444–445, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232019.

Melosh:1987:BRBb

[2173] Barbara Melosh. Book review: Unequal Colleagues: The Entrance of Women into the Professions, 1890–1940 by Penina Migdal Glazer; Miriam Slater. Isis, 78(3):446–447, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232020.

Mitman:1987:BRB

[2174] Gregg Mitman. Book review: Science as Politics by Les Levidow. Isis, 78 (3):447–448, September 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232021.

Cuff:1987:BRB

[2175] Robert Cuff. Book review: TVA and the Tellico Dam, 1936–1979: A Bureaucratic Crisis in Post-Industrial America by William Bruce Wheeler; Michael J. McDonald. Isis, 78(3):448–449, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232022.

Elliott:1987:BRBb

[2176] Clark A. Elliott. Book review: Francis Calley Gray and Art Collecting for America by Marjorie B. Cohn. Isis, 78(3):449–450, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232023.

Meikle:1987:BRB

[2177] Jeffrey L. Meikle. Book review: The Mechanic Muse by Hugh Kenner. Isis, 78(3):450, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232024.

Lurie:1987:BRB

[2178] Edward Lurie. Book review: The Rights of Memory: Essays on History, Science, and American Culture by Taylor Littleton. Isis, 78(3):450–451, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232025.

Roberts:1987:BRB

[2179] Jon H. Roberts. Book review: Without God, Without Creed: The Origins of Unbelief in America by James Turner. Isis, 78(3):451–452, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232026.

Forman:1987:BRB

[2180] Paul Forman. Book review: Basic Bethe: Seminal Articles on Nuclear Physics, 1936–1937 by Hans A. Bethe; Robert F. Bacher; M. Stanley Livingston. Isis, 78(3):453, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232027.

Kohler:1987:BRBb

[2181] Robert E. Kohler. Book review: Essays on the History of Organic Chemistry in the United States, 1875–1955 by Dean Stanley Tarbell; Ann Tracy Tarbell. Isis, 78(3):453–454, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232028.

Good:1987:BRB

[2182] Gregory A. Good. Book review: History of Geophysics. Volume II by C. Stewart Gillmor. Isis, 78(3):454–455, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/232029.

Glen:1987:BRB

[2183] William Glen. Book review: Perspectives on a Dynamic Earth by T. R. Paton; J. Clarke. Isis, 78(3):455, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232030.

Gillespie:1987:BRB

[2184] Richard Gillespie. Book review: The Ice: A Journey to Antarctica by Stephen J. Pyne. Isis, 78(3):456–457, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232031.

Fleming:1987:BRB

[2185] James R. Fleming. Book review: Minerals, Lands, and Geology for the Common Defence and General Welfare. Volume III: 1904–1939 by Mary C. Rabbitt. Isis, 78(3):457–458, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232032.

Dunbar:1987:BRBb

[2186] G. S. Dunbar. Book review: Dictionary of Quotations in Geography by James O. Wheeler; Francis M. Sibley. Isis, 78(3):458–459, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232033.

Bowler:1987:BRB

[2187] Peter J. Bowler. Book review: The Piltdown Inquest by Charles Blinderman. Isis, 78(3):459, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232034.

vanKeuren:1987:BRB

[2188] David K. van Keuren. Book review: Theories of Human Evolution: A Century of Debate, 1844–1944 by Peter J. Bowler. Isis, 78(3):459–461, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232035.

Kaye:1987:BRB

[2189] Howard L. Kaye. Book review: Evolution and Social Life by Tim Ingold. Isis, 78(3):461–462, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232036.

Lederberg:1987:BRB

[2190] Joshua Lederberg. Book review: Apprentice to Genius: The Making of a Scientific Dynasty by Robert Kanigel. Isis, 78(3):462–463, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232037.

Montgomery:1987:BRBb

[2191] William Montgomery. Book review: The Eye of Reason: Charles Darwin in Australasia by John Laurent; Margaret Campbell. Isis, 78(3):463, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232038.

Cittadino:1987:BRB

[2192] Eugene Cittadino. Book review: The Ecologists: From Merry Naturalists to Saviours of the Nation by Thomas Soderqvist. Isis, 78(3):463–464, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232039.

Schabas:1987:BRB

[2193] Margaret Schabas. Book review: Lives of the Laureates: Seven Nobel Economists by William Breit; Roger W. Spencer. Isis, 78(3):464–465, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232040.

Buxton:1987:BRB

[2194] William Buxton. Book review: The Nationalization of the Social Sciences by Samuel Z. Klausner; Victor M. Lidz. Isis, 78(3):465–466, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232041.

Matthews:1987:BRB

[2195] Fred Matthews. Book review: Evaluating Chicago Sociology: A Guide to the Literature, with an Annotated Bibliography by Lester R. Kurtz. Isis, 78(3):466–467, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232042.

Hilgard:1987:BRB

[2196] Ernest R. Hilgard. Book review: Behaviorism and Logical Positivism: A Reassessment of the Alliance by Laurence D. Smith. Isis, 78(3):467–468, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232043.

Vinovskis:1987:BRB

[2197] Maris Vinovskis. Book review: Medicine and American Growth, 1800– 1860 by James H. Cassedy. Isis, 78(3):468–469, September 1987. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232044.

Apple:1987:BRB

[2198] Rima D. Apple. Book review: Breasts, Bottles, and Babies: A History of Infant Feeding by Valerie A. Fildes. Isis, 78(3):469–470, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232045.

Lenoir:1987:BRBb

[2199] Timothy Lenoir. Book review: Der Aufstieg der Arzte im 19. Jahrhundert: Vom gelehrten Stand zum professionellen Experten: Das Beispiel Preussens by Claudia Huerkamp. Isis, 78(3):470–472, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232046.

Gottfried:1987:BRB

[2200] Robert S. Gottfried. Book review: Epidemics and Mortality in Early Modern Japan by Ann Bowman Jannetta. Isis, 78(3):472–473, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232047.

Donegan:1987:BRBb

[2201] Jane B. Donegan. Book review: Brought to Bed: Childbearing in America, 1750–1950 by Judith Walzer Leavitt. Isis, 78(3):473–475, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232048.

Melman:1987:BRB

[2202] Seymour Melman. Book review: Military Spending and Industrial Decline: A Study of the American Machine Tool Industry by Anthony Di-Filippo. Isis, 78(3):475–476, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232049.

Bruland:1987:BRB

[2203] Kristine Bruland. Book review: Technology, Innovation and Change by Brian Elliott. Isis, 78(3):476–477, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232050.

Mark:1987:BRBc

[2204] Robert Mark. Book review: Building Construction before Mechanization by John Fitchen. Isis, 78(3):477–478, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232051.

Millard:1987:BRB

[2205] Andre Millard. Book review: Electrical Engineering: The Second Century Begins by Harlow Freitag. Isis, 78(3):478–479, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232052.

Maddin:1987:BRB

[2206] Robert Maddin. Book review: Out of the Fiery Furnace: The Impact of Metals on the History of Mankind by Robert Raymond. Isis, 78(3): 479–480, September 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232053.

Pimlott:1987:BRB

[2207] J. L. Pimlott. Book review: From Waterloo to Balaclava: Tactics, Technology, and the British Army, 1815–1854 by Hew Strachan. Isis, 78(3): 480–481, September 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232054.

Crouch:1987:BRB

[2208] Tom D. Crouch. Book review: The Wright Flyer: An Engineering Perspective by Howard S. Wolko. Isis, 78(3):481, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232055.

Fuller:1987:BRB

[2209] Robert C. Fuller. Book review: Essays in Psychical Research by William James; Fredson Bowers; Ignas K. Skrupskelis. Isis, 78(3):481–482, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232056.

Moran:1987:BRB

[2210] Bruce T. Moran. Book review: Die Alchemie in der europaischen Kulturund Wissenschaftsgeschichte by Christoph Meinel. Isis, 78(3):482–483, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232057.

Westrum:1987:BRBa

[2211] Ron Westrum. Book review: The UFO Verdict: Examining the Evidence by Robert Sheaffer. Isis, 78(3):484, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232058.

Vallance:1987:BRB

[2212] J. T. Vallance. Book review: Aristotle on Nature and Living Things: Philosophical and Historical Studies Presented to David M. Balme on His Seventieth Birthday by Allan Gotthelf. Isis, 78(3):484–485, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232059.

Livesey:1987:BRB

[2213] Steven J. Livesey. Book review: Pseudo-Aristotle in the Middle Ages: The Theology and Other Texts by Jill Kraye; W. F. Ryan; C. B. Schmitt. Isis, 78(3):485–486, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232060.

Ovitt:1987:BRB

[2214] George Ovitt, Jr. Book review: Horses, Oxen, and Technological Innovation: The Use of Draught Animals in English Farming from 1066 to 1500 by John Langdon. Isis, 78(3):486–487, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232061.

Kidwell:1987:BRB

[2215] Clara Sue Kidwell. Book review: Native American Mathematics by Michael P. Closs. Isis, 78(3):487–488, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/232062.

Donovan:1987:BRB

[2216] Arthur Donovan. Book review: A Hotbed of Genius: The Scottish Enlightenment, 1730–1790 by David Daiches; Peter Jones; Jean Jones. Isis, 78(3):488–490, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232063.

Shapiro:1987:BRB

[2217] Alan E. Shapiro. Book review: The Newton Handbook by Derek Gjertsen. Isis, 78(3):490, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

232064.

Goodman:1987:BRB

[2218] Dena Goodman. Book review: Rousseau, Nature, and History by Asher Horowitz. Isis, 78(3):490–491, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232065.

Finocchiaro:1987:BRBa

[2219] Maurice A. Finocchiaro. Book review: La rana ambigua: La controversia sull’elettricita animale tra Galvani e Volta by Marcello Pera. Isis, 78(3): 491–492, September 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232066.

Mendoza:1987:BRB

[2220] Eric Mendoza. Book review: Reflexions on the Motive Power of Fire: A Critical Edition with the Surviving Manuscripts by Sadi Carnot; Robert Fox. Isis, 78(3):492–493, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232067.

Morison:1987:BRB

[2221] Elting E. Morison. Book review: The Correspondence and Miscellaneous Papers of Benjamin Henry Latrobe Volume II:1805–1810 by Benjamin Henry Latrobe; John C. Van Horne; Lee W. Formwalt; Darwin Stapleton; Jeffrey A. Cohen. Isis, 78(3):493–494, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232068.

Hamlin:1987:BRB

[2222] Christopher Hamlin. Book review: Pollution and Control: A Social History of the Thames in the Nineteenth Century by Bill Luckin. Isis, 78 (3):494–495, September 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232069.

Bud:1987:BRB

[2223] Robert Bud. Book review: Lancastrian Chemist: The Early Years of Sir Edward Frankland by Colin A. Russell. Isis, 78(3):495–496, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/232070.

Dick:1987:BRBb

[2224] Steven J. Dick. Book review: The Tucson Meteorites: Their History from Frontier Arizona to the Smithsonian by Richard R. Willey. Isis, 78 (3):496–497, September 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/232071.

Harbutt:1987:BRB

[2225] Fraser J. Harbutt. Book review: Britain’s Nuclear Arms Control Policy in the Context of Anglo–American Relations, 1957–68 by J. P. G. Freeman. Isis, 78(3):497–498, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/232072.

Lowood:1987:BRBb

[2226] Henry Lowood. Book review: The Big Score: The Billion Dollar Story of Silicon Valley by Michael S. Malone. Isis, 78(3):498, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232073.

Braun:1987:BRB

[2227] Hans-Joachim Braun. Book review: Diesel: Technology and Society in Industrial Germany by Donald E. Thomas. Isis, 78(3):498–499, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/232074.

McGraw:1987:BRB

[2228] Donald J. McGraw. Book review: Howard Florey: Penicillin and After by Trevor I. Williams. Isis, 78(3):499–500, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232075.

Anonymous:1987:BMc

[2229] Anonymous. Back matter. Isis, 78(3):501–506, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232076.

Anonymous:1987:FMd

[2230] Anonymous. Front matter. Isis, 78(4):507–513, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231914.

Rosenberg:1987:EIS

[2231] Charles Rosenberg. Editorial: Isis at seventy-five. Isis, 78(4):514–517, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231915.

Lindberg:1987:SHR

[2232] David C. Lindberg. Science as handmaiden: Roger Bacon and the Patristic tradition. Isis, 78(4):518–536, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231916.

Todes:1987:DMM

[2233] Daniel P. Todes. Darwin’s Malthusian metaphor and Russian evolutionary thought, 1859–1917. Isis, 78(4):537–551, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231917.

Heilbron:1987:AHS

[2234] J. L. Heilbron. Applied history of science. Isis, 78(4):552–563, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231918.

Newman:1987:NCS

[2235] William Newman and Issac Newton. Newton’s clavis as Starkey’s key. Isis, 78(4):564–574, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231919.

Freudenthal:1987:HSG

[2236] Gad Freudenthal. The historian of Science’s guide to France. Isis, 78(4): 575, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/231920.

Hawley:1987:BRB

[2237] Ellis W. Hawley, Robert E. Kohler, and Nathan Reingold. Book review: Science in the Federal Government: A History of Policies and Activities. Isis, 78(4):576–589, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231921.

Barber:1987:BRB

[2238] Bernard Barber. Book review: Little Science, Big Science, by Derek J. de Solla Price… and Beyond. Isis, 78(4):589–591, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231922.

Baatz:1987:BRB

[2239] Simon Baatz. Book review: Science and Society in Early America: Essays in Honor of Whitfield J. Bell, Jr. by Randolph S. Klein. Isis, 78 (4):592–593, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231923.

Burkhardt:1987:BRB

[2240] Richard W. Burkhardt, Jr. Book review: A Concordance to Darwin’s The Expression of the Emotions in Man and Animals by Paul Barrett; Donald J. Weinshank; Paul Ruhlen; Stephen J. Ozminski; Barbara N. Berghage. Isis, 78(4):593–594, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231924.

Morman:1987:BRBb

[2241] Edward T. Morman. Book review: Manuscripts of the Dibner Collection in the Dibner Library of the History of Science and Technology of the Smithsonian Institution Libraries. Isis, 78(4):594–595, December 1987.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231925.

Lindqvist:1987:BRB

[2242] Svante Lindqvist. Book review: The Machine in the University: Sample Course Syllabi for the History of Technology and Technology Studies by Terry S. Reynolds. Isis, 78(4):595–596, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231926.

Rothenberg:1987:BRB

[2243] Marc Rothenberg. Book review: A Guide to Documentary Editing by Mary-Jo Kline. Isis, 78(4):596–597, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231927.

Bolter:1987:BRB

[2244] Jay David Bolter. Book review: Machines and Intelligence: A Critique of Arguments against the Possibility of Artificial Intelligence by Stuart Goldkind. Isis, 78(4):597, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231928.

Dear:1987:BRB

[2245] Peter Dear. Book review: Der cartesische Materialismus, Maschine, Gesetz und Simulation: Eine Studie der intensionalen Ontologie der Naturwissenschaft by Gisela Loeck. Isis, 78(4):597–598, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231929.

Brown:1987:BRB

[2246] Theodore M. Brown. Book review: Disease and Discovery: A History of the Johns Hopkins School of Hygiene and Public Health, 1916– 1939 by Elizabeth Fee. Isis, 78(4):598–600, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231930.

Leslie:1987:BRB

[2247] Stuart W. Leslie. Book review: Volume I: The Founding Years: The New Frontier by Weldon B. Gibson; SRI. Volume II: The Take-Off Days. Isis, 78(4):600–602, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231931.

Dauben:1987:BRB

[2248] Joseph W. Dauben. Book review: Abrege d’histoire des mathematiques, 1700–1900 by Jean Dieudonne. Isis, 78(4):602, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231932.

Allison:1987:BRBc

[2249] David K. Allison. Book review: Papers of John von Neumann on Computing and Computer Theory by John von Neumann; William Aspray; Arthur Burks. Isis, 78(4):603, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231933.

Tatarewicz:1987:BRBb

[2250] Joseph N. Tatarewicz. Book review: The Analysis of Starlight: One Hundred and Fifty Years of Astronomical Spectroscopy by J. B. Hearnshaw. Isis, 78(4):603–604, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231934.

Needell:1987:BRB

[2251] Allan A. Needell. Book review: Black-Body Theory and the Quantum Discontinuity, 1894–1912 by Thomas S. Kuhn. Isis, 78(4):604–605, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231935.

Stuewer:1987:BRB

[2252] Roger H. Stuewer. Book review: The Particle Hunters by Yuval Ne’eman; Yoram Kirsh. Isis, 78(4):605, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231936.

Furukawa:1987:BRBb

[2253] Yasu Furukawa. Book review: High Performance Polymers: Their Origin and Development by Raymond B. Seymour; Gerald S. Kirshenbaum. Isis, 78(4):605–606, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231937.

Howard:1987:BRB

[2254] Don Howard. Book review: Mach I, Mach II, Einstein und die Relativitatstheorie: Eine Falschung und ihre Folgen by Gereon Wolters. Isis, 78 (4):606–607, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231938.

Westrum:1987:BRBb

[2255] Ron Westrum. Book review: Cosmic Debris: Meteorites in History by John G. Burke. Isis, 78(4):607–608, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231939.

Taylor:1987:BRBb

[2256] Kenneth L. Taylor. Book review: Time’s Arrow, Time’s Cycle: Myth and Metaphor in the Discovery of Geological Time by Stephen Jay Gould. Isis, 78(4):608–609, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231940.

Gillispie:1987:BRB

[2257] Charles C. Gillispie. Book review: Cartography in France, 1660–1848: Science, Engineering, and Statecraft by Josef Konvitz. Isis, 78(4):609– 611, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231941.

McConnell:1987:BRB

[2258] Anita McConnell. Book review: A Brief History of Geomagnetism and a Catalog of the Collection of the National Museum of American History by Robert P. Multhauf; Gregory Good. Isis, 78(4):611–612, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231942.

Farley:1987:BRB

[2259] John Farley. Book review: Aux sources de la biologie. Volume II: Les théories de la génération après la renaissance: La cytologie et la génétique by Rejane Bernier. Isis, 78(4):612, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231943.

Allen:1987:BRB

[2260] David E. Allen. Book review: A History of Shell Collecting by S. Peter Dance. Isis, 78(4):612–613, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231944.

Yoxen:1987:BRB

[2261] Edward Yoxen. Book review: Biotechnology: The University-Industrial Complex by Martin Kenney. Isis, 78(4):613–614, December 1987. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231945.

Larson:1987:BRBa

[2262] Allan Larson. Book review: The Dialectical Biologist by Richard Levins; Richard Lewontin. Isis, 78(4):614–615, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231946.

Bryder:1987:BRB

[2263] Linda Bryder. Book review: The White Plague: Tuberculosis, Man, and Society by Rene Dubos; Jean Dubos. Isis, 78(4):615–616, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231947.

Quen:1987:BRB

[2264] Jacques M. Quen. Book review: Melancholia and Depression: From Hippocratic Times to Modern Times by Stanley W. Jackson. Isis, 78 (4):617–618, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231948.

Thielman:1987:BRB

[2265] Samuel B. Thielman. Book review: Changing Faces of Madness: Early American Attitudes and Treatment of the Insane by Mary Ann Jimenez. Isis, 78(4):618–619, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231949.

Duffin:1987:BRB

[2266] Jacalyn M. Duffin. Book review: Histoire des medecins et pharmaciens de Marine et des Colonies by Pierre Pluchon. Isis, 78(4):619–620, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231950.

Brunton:1987:BRB

[2267] Deborah C. Brunton. Book review: The Eighteenth-Century Campaign to Avoid Disease by James C. Riley. Isis, 78(4):620–621, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231951.

Ludmerer:1987:BRB

[2268] Kenneth M. Ludmerer. Book review: The Therapeutic Perspective: Medical Practice, Knowledge, and Identity in America, 1820–1885 by John Harley Warner. Isis, 78(4):621–622, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231952.

Collins:1987:BRB

[2269] Martin Collins. Book review: History of Rocketry and Astronautics: Proceedings of the Third through the Sixth History Symposia of the International Academy of Astronautics by R. Cargill Hall. Isis, 78(4):622–623, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231953.

Culick:1987:BRB

[2270] F. E. C. Culick. Book review: Another Icarus: Percy Pilcher and the Quest for Flight by Philip Jarrett. Isis, 78(4):623–624, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231954.

Rosenbloom:1987:BRB

[2271] Richard S. Rosenbloom. Book review: Fast Forward: Hollywood, the Japanese, and the Onslaught of the VCR by James Lardner. Isis, 78 (4):624–625, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231955.

Owens:1987:BRB

[2272] Larry Owens. Book review: A History of Computing Technology by Michael R. Williams. Isis, 78(4):625–626, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231956.

Wik:1987:BRB

[2273] Reynold M. Wik. Book review: Fordson, Farmall, and Poppin’ Johnny: A History of the Farm Tractor and Its Impact on America by Robert C. Williams. Isis, 78(4):626, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231957.

Kemp:1987:BRB

[2274] Emory L. Kemp. Book review: Building Bridges: History, Technology, Construction by Hans Wittfoht; Edward Kluttz. Isis, 78(4):627–628, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231958.

Boyer:1987:BRB

[2275] Marjorie N. Boyer. Book review: Harvesting the Air: Windmill Pioneers in Twelfth-Century England by Edward J. Kealey. Isis, 78(4):628–629, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231959.

Cadden:1987:BRB

[2276] Joan Cadden. Book review: The Ages of Man: Medieval Interpretations of the Life Cycle by Elizabeth Sears. Isis, 78(4):629–630, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231960.

Molland:1987:BRB

[2277] A. George Molland. Book review: Robert Grosseteste: The Growth of an English Mind in Medieval Europe by R. W. Southern. Isis, 78(4): 630–631, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231961.

Webster:1987:BRB

[2278] Charles Webster. Book review: Religion and Neoplatonism in Renaissance Medicine by Walter Pagel; Marianne Winder; From Paracelsus to Van Helmont: Studies in Renaissance Medicine and Science by Walter Pagel; Marianne Winder. Isis, 78(4):631–632, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231962.

Hill:1987:BRB

[2279] Christopher Hill. Book review: Hobbes by Tom Sorell. Isis, 78(4):632– 634, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231963.

Shea:1987:BRBb

[2280] William R. Shea. Book review: La vita e le opere di Giovan Battista Hodierna by Mario Pavone. Isis, 78(4):634, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231964.

Finocchiaro:1987:BRBb

[2281] Maurice A. Finocchiaro. Book review: Galileo Galilei: Toward a Resolution of 350 Years of Debate–1633–1983 by Paul Cardinal Poupard; Pope John Paul; Ian Campbell. Isis, 78(4):634–635, December 1987.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231965.

McClellan:1987:BRB

[2282] James E. McClellan, III. Book review: Madame du Chatelet: Scientist, Philosopher, and Feminist of the Enlightenment by Esther Ehrman. Isis, 78(4):635–636, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231966.

Hahn:1987:BRB

[2283] Roger Hahn. Book review: Science and the Enlightenment by Thomas L. Hankins. Isis, 78(4):636–637, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231967.

Bennett:1987:BRB

[2284] J. A. Bennett. Book review: Benjamin Martin: Author, Instrument-Maker and “Country Showman”: Supplement by John R. Millburn; Retailer of the Sciences: Benjamin Martin’s Scientific Instrument Catalogues, 1756–1782 by John R. Millburn. Isis, 78(4):637–638, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231968.

Gregory:1987:BRBb

[2285] Frederick Gregory. Book review: Goethe and the Sciences: A Reappraisal by Frederick Amrine; Francis J. Zucker; Harvey Wheeler. Isis, 78(4):638– 639, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231969.

Churchill:1987:BRB

[2286] Frederick B. Churchill. Book review: Autobiography of Dr. Karl Ernst von Baer by Karl Ernst von Baer; Jane M. Oppenheimer. Isis, 78(4): 639–640, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231970.

Woodward:1987:BRBb

[2287] William R. Woodward. Book review: Das Spiel der Gegensatze: Friedrich Engels’ Philosophie und die Wissenschaften des 19. Jahrhunderts by Sven-Eric Liedman; Michael Tabukasch. Isis, 78(4):640–641, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231971.

Porter:1987:BRBd

[2288] Roy Porter. Book review: Gentlemen of Science: Early Correspondence of the British Association for the Advancement of Science by Jack Morrell; Arnold Thackray. Isis, 78(4):641–642, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231972.

Blustein:1987:BRB

[2289] Bonnie Ellen Blustein. Book review: The Papers of Frederick Law Olmsted. Volume IV: Defending the Union: The Civil War and the U.S. Sanitary Commission, 1861–1863 by Frederick Law Olmsted; Jane Turner Censer. Isis, 78(4):642–643, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231973.

Pickering:1987:BRBb

[2290] Andy Pickering. Book review: The Life It Brings: One Physicist’s Beginnings by Jeremy Bernstein. Isis, 78(4):643–644, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231974.

Pang:1987:BRB

[2291] Alex Soojung-Kim Pang. Book review: Clipped Wings: The American SST Conflict by Mel Horwitch. Isis, 78(4):644–645, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231975.

Logsdon:1987:BRB

[2292] John M. Logsdon. Book review: The Space Station: A Personal Journey by Hans Mark. Isis, 78(4):645–646, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231976.

Heims:1987:BRB

[2293] Steve J. Heims. Book review: Collected Works with Commentaries. Volume IV: Cybernetics, Science, and Society; Ethics, Aesthetics, and Literary Criticism; Book Reviews and Obituaries by Norbert Wiener; P. Masani. Isis, 78(4):646–647, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231977.

Hollinger:1987:BRB

[2294] David A. Hollinger. Book review: Beyond the Laboratory: Scientists as Political Activists in 1930s America by Peter J. Kuznick. Isis, 78 (4):647–648, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231978.

Stocking:1987:BRB

[2295] S. Holly Stocking. Book review: The Science Critic: A Critical Analysis of the Popular Presentation of Science by Maurice Goldsmith. Isis, 78 (4):648–649, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231979.

Woolf:1987:BRB

[2296] Patricia Woolf. Book review: A Difficult Balance: Editorial Peer Review in Medicine by Stephen Lock. Isis, 78(4):649–650, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231980.

LaFollette:1987:BRB

[2297] Marcel Chotkowski LaFollette. Book review: Selling Science: How the Press Covers Science and Technology by Dorothy Nelkin. Isis, 78(4): 650–652, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231981.

Drucker:1987:BRB

[2298] Thomas L. Drucker. Book review: Women of Mathematics: A Biobibliographic Sourcebook by Louise S. Grinstein; Paul J. Campbell. Isis, 78 (4):652–653, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231982.

Fissell:1987:BRBb

[2299] Mary E. Fissell. Book review: “To Toil the Livelong Day”: America’s Women at Work, 1798–1980 by Carol Groneman; Mary Beth Norton. Isis, 78(4):653, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231983.

Ogilvie:1987:BRB

[2300] Marilyn Ogilvie. Book review: Women of Science, Technology, and Medicine: A Bibliography by Else Hoyrup. Isis, 78(4):654, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231984.

Leavitt:1987:BRB

[2301] Judith Walzer Leavitt. Book review: Women Doctors in Gilded-Age Washington: Race, Gender, and Professionalization by Gloria Moldow. Isis, 78(4):654–656, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231985.

Rosner:1987:BRBb

[2302] David Rosner. Book review: The Hawk’s Nest Incident: America’s Worst Industrial Disaster by Martin Cherniack. Isis, 78(4):656–657, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231986.

Lawrence:1987:BRB

[2303] Christopher Lawrence. Book review: Nervous Juyces and the Feeling Heart: The Growth of Sensibility in the Novels of Tobias Smollett by Susan Bourgeois. Isis, 78(4):657–658, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/231987.

Dumain:1987:BRB

[2304] Ralph Dumain. Book review: Songs from Unsung Worlds: Science in Poetry by Bonnie Bilyeu Gordon. Isis, 78(4):658–659, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231988.

Rousseau:1987:BRB

[2305] G. S. Rousseau. Book review: Erasmus Darwin and the Romantic Poets by Desmond King-Hele. Isis, 78(4):659–660, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231989.

Lyon:1987:BRBb

[2306] John Lyon. Book review: Chesterton, a Seer of Science by Stanley L. Jaki. Isis, 78(4):660–661, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231990.

Douglas:1987:BRB

[2307] Susan J. Douglas. Book review: Shifting Gears: Technology, Literature, Culture in Modernist America by Cecelia Tichi. Isis, 78(4):661–662, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231991.

Larson:1987:BRBb

[2308] Edward J. Larson. Book review: The Wisdom of Science: Its Relevance to Culture and Religion by Hanbury Brown. Isis, 78(4):662–663, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/231992.

Blaisdell:1987:BRB

[2309] Muriel L. Blaisdell. Book review: Darwinism and Divinity: Essays on Evolution and Religious Belief by John Durant. Isis, 78(4):663–664, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/231993.

Oakley:1987:BRB

[2310] Francis Oakley. Book review: Theology and the Scientific Imagination from the Middle Ages to the Seventeenth Century by Amos Funkenstein. Isis, 78(4):664–665, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

231994.

Cuello:1987:BRB

[2311] Cesar Cuello. Book review: Technology: Philosophical and Social Aspects by Joseph Agassi. Isis, 78(4):665–666, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231995.

Cutcliffe:1987:BRBb

[2312] Stephen H. Cutcliffe. Book review: Technological Change and the Transformation of America by Steven E. Goldberg; Charles R. Strain. Isis, 78 (4):666–667, December 1987. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/231996.

Anonymous:1987:BMd

[2313] Anonymous. Back matter. Isis, 78(4):668–690, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231997.

Anonymous:1987:VI

[2314] Anonymous. Volume information. Isis, 78(4):673–685, December 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/231913.

Anonymous:1987:FMe

[2315] Anonymous. Front matter. Isis, 78(5):1–4, ???? 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233417.

Neu:1987:OHT

[2316] John Neu. One hundred twelfth critical bibliography of the history of science and its cultural influences. Isis, 78(5):5–274, ???? 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233418.

Anonymous:1987:BMe

[2317] Anonymous. Back matter. Isis, 78(5):??, ???? 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233419.

Anonymous:1988:FMa

[2318] Anonymous. Front matter. Isis, 79(1):1–5, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234438.

Geison:1988:PPD

[2319] Gerald L. Geison and James A. Secord. Pasteur and the process of discovery: The case of optical isomerism. Isis, 79(1):6–36, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234439.

McCormmach:1988:HCT

[2320] Russell McCormmach. Henry Cavendish on the theory of heat. Isis, 79 (1):37–67, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234440.

Meinel:1988:ESC

[2321] Christoph Meinel. Early Seventeenth-Century atomism: Theory, epistemology, and the insufficiency of experiment. Isis, 79(1):68–103, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234441.

Bowler:1988:BRBa

[2322] Peter J. Bowler. Book review: Victorian Anthropology. Isis, 79(1):104– 107, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234442.

Sica:1988:BRB

[2323] Alan Sica. Book review: New School: A History of the New School for Social Research by Peter M. Rutkoff; William B. Scott. Isis, 79(1):108– 109, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234443.

Porter:1988:BRBa

[2324] Theodore M. Porter. Book review: The Search for a Methodology of Social Science: Durkheim, Weber, and the Nineteenth-Century Problem of Cause, Probability, and Action by Stephen P. Turner. Isis, 79(1):109– 110, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234444.

Popplestone:1988:BRB

[2325] John A. Popplestone. Book review: The Shaping of Modern Psychology: An Historical Introduction by L. S. Hearnshaw. Isis, 79(1):110–111, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234445.

Capshew:1988:BRB

[2326] James H. Capshew. Book review: Psychology in America: A Historical Survey by Ernest R. Hilgard; A History of Psychology: Main Currents in Psychological Thought by Thomas Hardy Leahey. Isis, 79(1):111– 112, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234446.

DeMey:1988:BRB

[2327] Marc De Mey. Book review: Piaget’s Theory of Knowledge: Genetic Epistemology and Scientific Reason by Richard F. Kitchener. Isis, 79(1): 112–114, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234447.

Rosenberg:1988:BRB

[2328] Rosalind Rosenberg. Book review: Untold Lives: The First Generation of American Women Psychologists by Elizabeth Scarborough; Laurel Furumoto. Isis, 79(1):114–115, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234448.

Brick:1988:BRB

[2329] Howard Brick. Book review: Talcott Parsons and the Capitalist Nation-State: Political Sociology as a Strategic Vocation by William Buxton. Isis, 79(1):115–116, March 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234449.

Klingemann:1988:BRB

[2330] Carsten Klingemann. Book review: Deutsche Soziologie, 1933–1945: Die Normalität einer Anpassung by Otthein Rammstedt. Isis, 79(1):116– 117, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234450.

vanKeuren:1988:BRBa

[2331] David van Keuren. Book review: Objects and Others: Essays on Museums and Material Culture by George W. Stocking; Malinowski, Rivers, Benedict, and Others: Essays on Culture and Personality by George W. Stocking. Isis, 79(1):117–118, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234451.

Rosenkrantz:1988:BRB

[2332] Barbara Gutmann Rosenkrantz. Book review: Yellow Fever in the North: The Methods of Early Epidemiology. Isis, 79(1):119–122, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234452.

Pyenson:1988:BRBa

[2333] Lewis Pyenson. Book review: The Dilemmas of an Upright Man: Max Planck as Spokesman for German Science. Isis, 79(1):122–126, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234453.

Shapin:1988:BRB

[2334] Steven Shapin. Book review: The Figural and the Literal: Problems of Language in the History of Science and Philosophy, 1630–1800 by Andrew E. Benjamin; Geoffrey N. Cantor; John R. R. Christie. Isis, 79(1):127–128, March 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234454.

DeVorkin:1988:BRB

[2335] David DeVorkin. Book review: Truth and Beauty: Aesthetics and Motivations in Science by S. Chandrasekhar. Isis, 79(1):128–129, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234455.

Wynne:1988:BRB

[2336] Brian Wynne. Book review: An Introduction to Science Studies: The Philosophical and Social Aspects of Science and Technology by John Ziman. Isis, 79(1):129, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234456.

Lowood:1988:BRBa

[2337] Henry Lowood. Book review: The Tradition of Science: Landmarks of Western Science in the Collections of the Library of Congress by Leonard C. Bruno. Isis, 79(1):130, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234457.

Serio:1988:BRB

[2338] Giorgia Foderà Serio. Book review: Catalogo della corrispondenza degli Astronomi di Brera, 1726–1799 by Agnese Mandrino; Guido Tagliaferri; Pasquale Tucci. Isis, 79(1):130–131, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234458.

Hochheiser:1988:BRB

[2339] Sheldon Hochheiser. Book review: A Guide to Archives and Manuscript Collections in the History of Chemistry and Chemical Technology by George D. Tselos; Colleen Wickey. Isis, 79(1):131–132, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234459.

Hackmann:1988:BRB

[2340] Willem D. Hackmann. Book review: Dutch Pioneers of Science by Leo Beek. Isis, 79(1):132–134, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234460.

Morman:1988:BRBa

[2341] Edward T. Morman. Book review: The Man behind the Syndrome by Peter Beighton; Greta Beighton. Isis, 79(1):134, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234461.

Strimpel:1988:BRB

[2342] Oliver Strimpel. Book review: Portraits in Silicon by Robert Slater. Isis, 79(1):134–135, March 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234462.

Latour:1988:BRB

[2343] Bruno Latour. Book review: The Construction of Reality by Michael A. Arbib; Mary B. Hesse. Isis, 79(1):135–137, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234463.

Scanlan:1988:BRB

[2344] James Scanlan. Book review: Reason and Being by B. G. Kuznetsov; Carolyn R. Fawcett; Robert Cohen. Isis, 79(1):137–138, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234464.

Oldroyd:1988:BRB

[2345] David Oldroyd. Book review: Philosophy of Science and Historical Enquiry by John Losee. Isis, 79(1):138–139, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234465.

McEvoy:1988:BRB

[2346] John G. McEvoy. Book review: The Process of Science: Contemporary Philosophical Approaches to Understanding Scientific Practice by Nancy J. Nersessian. Isis, 79(1):139–140, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234466.

Nakayama:1988:BRB

[2347] Shigeru Nakayama. Book review: The Hall of Heavenly Records: Korean Astronomical Instruments and Clocks, 1380–1780 by Joseph Needham; Lu Gwei-Djen; John H. Combridge; John S. Major. Isis, 79(1):140– 141, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234467.

Landes:1988:BRB

[2348] David S. Landes. Book review: The Time Museum: Catalogue of the Collection by Bruce Chandler; The Time Museum: Catalogue of the Collection. Volume I: Time Measuring Instruments. Part 1: Astrolabes and Related Instruments; The Time Museum: Catalogue of the Collection.

Volume I: Time Measuring Instruments. Part 3: Water-Clocks, Sand-Glasses, and Fire-Clocks by A. J. Turner. Isis, 79(1):141–143, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234468.

Guagnini:1988:BRB

[2349] Anna Guagnini. Book review: Education for the Industrial World: The Écoles d’Arts et Métiers and the Rise of French Industrial Engineering by C. R. Day. Isis, 79(1):143–144, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234469.

Bensaude-Vincent:1988:BRB

[2350] Bernadette Bensaude-Vincent and Thérèse Charmasson. Book review: Histoire de l’École Polytechnique by Ambroise Fourcy; La République avait besoin de savants: Les débuts de l’École Polytechnique: L’École centrale des travaux publics et les cours révolutionnaires de l’An III by Janis Langins. Isis, 79(1):144–145, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234470.

Westrum:1988:BRB

[2351] Ron Westrum. Book review: The Enigma of Loch Ness: Making Sense of a Mystery by Henry H. Bauer. Isis, 79(1):145–146, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234471.

Nye:1988:BRBa

[2352] Robert A. Nye. Book review: The Foul and the Fragrant: Odor and the French Social Imagination by Alain Corbin; Miriam Kochan. Isis, 79(1):146–147, March 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234472.

Ezrahi:1988:BRB

[2353] Yaron Ezrahi. Book review: Governing Science and Technology in a Democracy by Malcolm L. Goggin. Isis, 79(1):147–148, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234473.

Collins:1988:BRB

[2354] H. M. Collins. Book review: Laboratory Life: The Construction of Scientific Facts by Bruno Latour; Steve Woolgar. Isis, 79(1):148–149, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234474.

Gregory:1988:BRB

[2355] Frederick Gregory. Book review: Evolutionismus und Christentum by Robert Spaemann; Reinhard Löw; Peter Koslowski. Isis, 79(1):149–150, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234475.

Warner:1988:BRBa

[2356] Deborah Jean Warner. Book review: Art and Cartography: Six Historical Essays by David Woodward. Isis, 79(1):150–151, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234476.

Rowe:1988:BRBa

[2357] David E. Rowe. Book review: Linear Differential Equations and Group Theory from Riemann to Poincaré by Jeremy Gray. Isis, 79(1):151– 152, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234477.

Wise:1988:BRB

[2358] George Wise. Book review: Silicon and Silicones: About Stone-Age Tools, Antique Pottery, Modern Ceramics, Computers, Space Materials and How They All Got That Way by Eugene G. Rochow. Isis, 79 (1):152–153, March 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234478.

Costa:1988:BRB

[2359] Albert B. Costa. Book review: Essays on the History of Organic Chemistry by James G. Traynham. Isis, 79(1):153–154, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234479.

Dick:1988:BRBa

[2360] Steven J. Dick. Book review: Coon Mountain Controversies: Meteor Crater and the Development of Impact Theory by William Graves Hoyt. Isis, 79(1):154–155, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234480.

Porter:1988:BRBb

[2361] Roy Porter. Book review: From Mineralogy to Geology: The Foundations of a Science, 1650–1830 by Rachel Laudan. Isis, 79(1):155–156, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234481.

Egerton:1988:BRB

[2362] Frank N. Egerton. Book review: Contemporary Classics in Plant, Animal, and Environmental Sciences by James T. Barrett. Isis, 79(1):156– 157, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234482.

Adams:1988:BRBa

[2363] Mark B. Adams. Book review: Lysenkoism in China: Proceedings of the 1956 Qingdao Genetics Symposium by Laurence Schneider; Qin Shizhen. Isis, 79(1):157–158, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234483.

Porter:1988:BRBc

[2364] Charlotte M. Porter. Book review: Charles Wright on the Boundary, 1849–1852, or Plantae Wrightianae Revisited by Elizabeth A. Shaw. Isis, 79(1):158, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234484.

Maulitz:1988:BRBa

[2365] Russell C. Maulitz. Book review: Doctor Dock: Teaching and Learning Medicine at the Turn of the Century by Horace W. Davenport. Isis, 79(1):159–160, March 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234485.

Leslie:1988:BRB

[2366] Stuart W. Leslie. Book review: RCA and the VideoDisc: The Business of Research by Margaret B. W. Graham. Isis, 79(1):160–161, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234486.

Stilgoe:1988:BRB

[2367] John R. Stilgoe. Book review: Railroads and the Character of America, 1820–1887 by James A. Ward. Isis, 79(1):162, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234487.

Fulford:1988:BRB

[2368] Michael Fulford. Book review: The Roman Port and Fishery of Cosa: A Center of Ancient Trade by Anna Marguerite McCann; Joanne Bourgeois; Elaine K. Gazda; John Peter Oleson; Elizabeth Lyding Will. Isis, 79(1):162–163, March 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234488.

Peters:1988:BRB

[2369] Edward Peters. Book review: Philoponus and the Rejection of Aristotelian Science by Richard Sorabji. Isis, 79(1):163–164, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234489.

Multhauf:1988:BRB

[2370] Robert P. Multhauf. Book review: Science and Civilisation in China. Volume V: Chemistry and Chemical Technology. Part I: Paper and Printing by Joseph Needham; Tsien Tsuen-Hsuin. Isis, 79(1):164–165, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234490.

Golinski:1988:BRBa

[2371] Jan Golinski. Book review: Chemistry, Alchemy and the New Philosophy, 1550–1700 by Allen G. Debus. Isis, 79(1):165–166, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234491.

Hall:1988:BRB

[2372] A. Rupert Hall. Book review: Spinoza and the Sciences by Marjorie Grene; Debra Nails. Isis, 79(1):166, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234492.

Debus:1988:BRB

[2373] Allen G. Debus. Book review: William Cooper’s A Catalogue of Chymical Books, 1673–88: A Verified Edition by Stanton J. Linden. Isis, 79(1): 167, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234493.

Eisele:1988:BRB

[2374] Carolyn Eisele. Book review: Peirce, Semeiotic, and Pragmatism by Max H. Fisch; Kenneth Laine Ketner; Christian J. W. Kloesel. Isis, 79(1):167–169, March 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234494.

Faust:1988:BRB

[2375] Drew Gilpin Faust. Book review: Henry William Ravenel, 1814–1887: South Carolina Scientist in the Civil War Era by Tamara Miner Haygood. Isis, 79(1):169–170, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234495.

Ewan:1988:BRB

[2376] Joseph Ewan. Book review: André and François André Michaux by Henry Savage,; Elizabeth J. Savage. Isis, 79(1):170–171, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234496.

Pratt:1988:BRB

[2377] Joseph A. Pratt. Book review: Energy and the Federal Government: Fossil Fuel Policies, 1900–1946 by John G. Clark. Isis, 79(1):171–172, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234497.

Rhees:1988:BRB

[2378] David J. Rhees. Book review: The Problem Solvers: A History of Arthur D. Little, Inc. by E. J. Kahn. Isis, 79(1):172–173, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234498.

Stranges:1988:BRB

[2379] Anthony N. Stranges. Book review: The Griffin: The Greatest Untold Espionage Story of World War II by Arnold Kramish. Isis, 79(1):173– 174, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234499.

Needell:1988:BRB

[2380] Allan A. Needell. Book review: Rabi: Scientist and Citizen by John S. Rigden. Isis, 79(1):174–175, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234500.

Trimble:1988:BRB

[2381] William F. Trimble. Book review: Model Research: The National Advisory Committee for Aeronautics, 1915–1958 by Alex Roland. Isis, 79 (1):175–176, March 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234501.

Anonymous:1988:BMa

[2382] Anonymous. Back matter. Isis, 79(1):177–180, March 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234502.

Anonymous:1988:FMb

[2383] Anonymous. Front matter. Isis, 79(2):181–187, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233603.

Smith:1988:PVP

[2384] A. Mark Smith. The psychology of visual perception in Ptolemy’s optics. Isis, 79(2):188–207, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233604.

Koblitz:1988:SWR

[2385] Ann Hibner Koblitz. Science, women, and the Russian intelligentsia: The generation of the 1860s. Isis, 79(2):208–226, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233605.

Peishan:1988:GCQ

[2386] Li Peishan. Genetics in China: The Qingdao Symposium of 1956. Isis, 79(2):227–236, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233606.

Sokal:1988:AMH

[2387] Michael M. Sokal, John W. Servos, Edith Sylla, and Frederick Gregory. Annual meeting of the History of Science Society 29 October–1 November 1987. Isis, 79(2):237–242, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233607.

Grant:1988:SMC

[2388] Edward Grant. 1987 Sarton Medal citation. Isis, 79(2):243–244, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233608.

Anonymous:1988:DSF

[2389] Anonymous. Donors to and supporters of the fund drive. Isis, 79(2): 244–245, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233609.

Savage-Smith:1988:GAW

[2390] Emilie Savage-Smith. Gleanings from an Arabist’s workshop: Current trends in the study of Medieval Islamic science and medicine. Isis, 79(2): 246–266, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233610.

Saliba:1988:BRBa

[2391] George Saliba. Book review: Les arithmétiques by Diophante; Roshdi Rashed; Books IV to VII of Diophantus’ Arithmetica in the Arabic Translation Attributed to Qusta ibn Luqa by Diophantus; Jacques Sesiano. Isis, 79(2):266–270, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233611.

Newman:1988:BRB

[2392] William Newman. Book review: Jabir ibn H .ayyan: Contribution à l’histoire des idées scientifiques dans l’Islam: Jabir et la science grecque by Paul Kraus. Isis, 79(2):270–271, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233612.

Morelon:1988:BRB

[2393] Régis Morelon. Book review: Der Sternkatalog des Almagest: Die arabischmittelalterliche Tradition. Volume I: Die arabischen Ubersetzungen by Claudius Ptolemaüs; Paul Kunitzsch. Isis, 79(2):271–272, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233613.

Gould:1988:BRB

[2394] Stephen Jay Gould. Book review: Sewall Wright and Evolutionary Biology. Isis, 79(2):273–281, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233614.

Adams:1988:BRBb

[2395] Mark B. Adams. Book review: Factors of Evolution: The Theory of Stabilizing Selection. Isis, 79(2):281–284, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233615.

Ruse:1988:BRB

[2396] Michael Ruse. Book review: Understanding Science through Evolution: A Humanist Approach by Arnold M. Clark; Evolution and the Humanities by David Holbrook. Isis, 79(2):284–285, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233616.

Worster:1988:BRB

[2397] Donald Worster. Book review: The Background of Ecology: Concept and Theory by Robert P. McIntosh. Isis, 79(2):285–286, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233617.

Bradie:1988:BRB

[2398] Michael Bradie. Book review: Taking Darwin Seriously: A Naturalistic Approach to Philosophy by Michael Ruse. Isis, 79(2):286–287, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233618.

Deiss:1988:BRBa

[2399] William A. Deiss. Book review: A Catalogue of Natural Science Collections in North-East England, with Biographical Notes on the Collectors by Peter Davis; Christopher Brewer. Isis, 79(2):287–288, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233619.

Reeds:1988:BRBa

[2400] Karen Reeds. Book review: Herbals, Their Origin and Evolution: A Chapter in the History of Botany, 1470–1670 by Agnes Arber; William T. Stearn. Isis, 79(2):288–289, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233620.

Larson:1988:BRBa

[2401] James Larson. Book review: Book of Birds: A Facsimile of the Original Watercolors of Olof Rudbeck the Younger in the Leufsta Collection in Uppsala University Library by Olof Rudbeck; Björn Löwendahl. Isis, 79 (2):289, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233621.

Dougherty:1988:BRB

[2402] F. W. P. Dougherty. Book review: Science against the Unbelievers: The Correspondence of Bonnet and Needham, 1760–1780 by Renato G.

Mazzolini; Shirley A. Roe. Isis, 79(2):289–290, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233622.

Limoges:1988:BRB

[2403] Camille Limoges. Book review: Le voyage de d’Entrecasteaux à la recherche de Lapérouse by Hélène Richard. Isis, 79(2):290–291, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233623.

Outram:1988:BRB

[2404] Dorinda Outram. Book review: The Cuvier-Geoffroy Debate: French Biology in the Decades before Darwin by Toby A. Appel. Isis, 79(2):291– 292, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233624.

Miller:1988:BRBa

[2405] David Philip Miller. Book review: Jupiter Botanicus: Robert Brown of the British Museum by D. J. Mabberley. Isis, 79(2):292–293, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233625.

Bedini:1988:BRB

[2406] Silvio A. Bedini. Book review: The Eagle’s Nest: Natural History and American Ideas, 1812–1842 by Charlotte M. Porter. Isis, 79(2):293– 294, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233626.

Stuckey:1988:BRB

[2407] Ronald L. Stuckey. Book review: Précis ou abrégé des voyages, travaux, et recherches de C. S. Rafinesque (1833); The Original Version of A Life of Travels (1836) by C. S. Rafinesque; Charles Boewe; Georges Reynaud; Beverly Seaton. Isis, 79(2):294–295, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233627.

Rainger:1988:BRBa

[2408] Ronald Rainger. Book review: Simple Curiosity: Letters from George Gaylord Simpson to His Family, 1921–1970 by Léo F. Laporte. Isis, 79(2):295–296, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233628.

Gingerich:1988:BRB

[2409] Owen Gingerich and Robert S. Westman. Book review: Three Imperial Mathematicians: Kepler Trapped between Tycho Brahe and Ursus. Isis, 79(2):297–299, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233629.

Kulik:1988:BRB

[2410] Gary Kulik. Book review: Most Wonderful Machine: Mechanization and Social Change in Berkshire Paper Making, 1801–1885. Isis, 79(2):299– 301, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233630.

Nyhart:1988:BRB

[2411] Lynn K. Nyhart. Book review: The Search for Solutions by Horace Freeland Judson. Isis, 79(2):302–303, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233631.

Gieryn:1988:BRB

[2412] Thomas F. Gieryn. Book review: Philosophy and Sociology of Science: An Introduction by Stewart Richards. Isis, 79(2):303, June 1988. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233632.

Morman:1988:BRBb

[2413] Edward T. Morman. Book review: Catalogue of the Transylvania University Medical Library. Isis, 79(2):304, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233633.

Wickey:1988:BRB

[2414] Colleen Wickey. Book review: A New Guide to the Collections in the Library of the American Philosophical Society by J. Stephen Catlett. Isis, 79(2):304–305, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233634.

Whiteside:1988:BRB

[2415] D. T. Whiteside. Book review: Biobibliography of British Mathematics and Its Applications. Part II: 1701–1760 by R. V. Wallis; P. J. Wallis. Isis, 79(2):305–307, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233635.

Paul:1988:BRBa

[2416] H. W. Paul. Book review: Premices philosophiques by Pierre Duhem; Stanley L. Jaki. Isis, 79(2):307, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233636.

Lloyd:1988:BRB

[2417] Elisabeth A. Lloyd. Book review: The Science Question in Feminism by Sandra Harding. Isis, 79(2):308–309, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233637.

Leplin:1988:BRBa

[2418] Jarrett Leplin. Book review: Varieties of Realism: A Rationale for the Natural Sciences by Rom Harré. Isis, 79(2):309–310, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233638.

Nordmann:1988:BRB

[2419] Alfred Nordmann. Book review: Decisions and Revisions: Philosophical Essays on Knowledge and Value by Isaac Levi; Hard Choices: Decision Making under Unresolved Conflict by Isaac Levi. Isis, 79(2):310–312, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233639.

Schumacher:1988:BRB

[2420] John Schumacher. Book review: New Ways of Knowing: The Sciences, Society, and Reconstructive Knowledge by Marcus G. Raskin; Herbert J. Bernstein; Susan Buck-Morss; Noam Chomsky; Michael Goldhaber; Edward S. Herman; Joseph Turner. Isis, 79(2):312–313, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233640.

Turner:1988:BRB

[2421] Stephen Turner. Book review: Understanding Social Science: A Philosophical Introduction to the Social Sciences by Roger Trigg. Isis, 79(2): 313, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233641.

Pyenson:1988:BRBb

[2422] Lewis Pyenson. Book review: Technological Development in China, India and Japan: Cross-Cultural Perspectives by Erik Baark; Andrew Jamison.

Isis, 79(2):314, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233642.

Pinch:1988:BRB

[2423] Trevor Pinch. Book review: How Institutions Think by Mary Douglas. Isis, 79(2):314–315, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233643.

Sheets-Pyenson:1988:BRB

[2424] Susan Sheets-Pyenson. Book review: Science in the Provinces: Scientific Communities and Provincial Leadership in France, 1860–1930 by Mary Jo Nye. Isis, 79(2):315–317, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233644.

Worboys:1988:BRB

[2425] Michael Worboys. Book review: Scientific Colonialism: A Cross-Cultural Comparison by Nathan Reingold; Marc Rothenberg. Isis, 79(2):317– 318, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233645.

Pickstone:1988:BRB

[2426] J. V. Pickstone. Book review: Physic and Philanthropy: A History of the Wellcome Trust, 1936–1986 by A. Rupert Hall; B. A. Bembridge. Isis, 79(2):318–319, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233646.

Nichelson:1988:BRB

[2427] Brian J. Nichelson. Book review: Engineer in Charge: A History of the Langley Aeronautical Laboratory, 1917–1958 by James R. Hansen. Isis, 79(2):319–320, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233647.

Starbuck:1988:BRB

[2428] David R. Starbuck. Book review: American Archaeology Past and Future: A Celebration of the Society for American Archaeology 1935–1985 by David J. Meltzer; Donald D. Fowler; Jeremy A. Sabloff. Isis, 79(2): 320–321, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233648.

McClellan:1988:BRB

[2429] James E. McClellan, III. Book review: Royal Funding of the Parisian Académie des Sciences during the 1690s by Alice Stroup. Isis, 79(2):

321–322, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233649.

Rousseau:1988:BRB

[2430] G. S. Rousseau. Book review: The Cosmic Web: Scientific Field Models and Literary Strategies in the Twentieth Century by N. Katherine Hayles. Isis, 79(2):322–323, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233650.

Cassidy:1988:BRB

[2431] David Cassidy. Book review: Einstein and the Humanities by Dennis P. Ryan. Isis, 79(2):323–324, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233651.

Rowe:1988:BRBb

[2432] David E. Rowe. Book review: The History of Mathematics: A Reader by John Fauvel; Jeremy Gray. Isis, 79(2):324–325, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233652.

Parshall:1988:BRB

[2433] Karen Hunger Parshall. Book review: Figures de l’infini: Les mathématiques au miroir des cultures by Tony Lévy. Isis, 79(2):325– 326, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233653.

Porter:1988:BRBd

[2434] Theodore M. Porter. Book review: The History of Statistics: The Measurement of Uncertainty before 1900 by Stephen M. Stigler. Isis, 79(2): 326–327, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233654.

Rothenberg:1988:BRB

[2435] Marc Rothenberg. Book review: New Lands, New Men: America and the Second Great Age of Discovery by William H. Goetzmann. Isis, 79(2): 327–329, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233655.

Taylor:1988:BRB

[2436] Kenneth L. Taylor. Book review: Histoire de la géologie by Gabriel Gohau. Isis, 79(2):329–330, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233656.

Bowler:1988:BRBb

[2437] Peter J. Bowler. Book review: Racial Theories by Michael Banton. Isis, 79(2):330–331, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233657.

vanKeuren:1988:BRBb

[2438] David van Keuren. Book review: Bones of Contention: Controversies in the Search for Human Origins by Roger Lewin. Isis, 79(2):331–332, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233658.

Albury:1988:BRB

[2439] W. R. Albury. Book review: William Hunter and the Eighteenth-Century Medical World by W. F. Bynum; Roy Porter. Isis, 79(2):332–333, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233659.

Donegan:1988:BRB

[2440] Jane B. Donegan. Book review: Wash and Be Healed: The Water-Cure Movement and Women’s Health by Susan E. Cayleff. Isis, 79(2):333– 334, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233660.

Maulitz:1988:BRBb

[2441] Russell C. Maulitz. Book review: Johannes Müller and the Nineteenth-Century Origins of Tumor Cell Theory by L. J. Rather; Patricia Rather; John R. Frerichs. Isis, 79(2):334–335, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233661.

Hirsh:1988:BRB

[2442] Richard F. Hirsh. Book review: Concorde and Dissent: Explaining High Technology Failures in Britain and France by Elliot J. Feldman. Isis, 79 (2):336, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233662.

Constant:1988:BRB

[2443] Edward W. Constant. Book review: Whittle: The True Story by John Golley; Bill Gunston; Frank Whittle. Isis, 79(2):337, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233663.

Reed:1988:BRBa

[2444] Germaine M. Reed. Book review: The Modernization of the Louisiana Sugar Industry, 1830–1910 by John Alfred Heitmann. Isis, 79(2):337– 339, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233664.

Freeland:1988:BRB

[2445] Cynthia A. Freeland. Book review: The Origins of Civilization in Greek and Roman Thought by Sue Blundell; Aristotle’s Classification of Animals: Biology and the Conceptual Unity of the Aristotelian Corpus by Pierre Pellegrin; Anthony Preus. Isis, 79(2):339–340, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233665.

Hankinson:1988:BRB

[2446] R. J. Hankinson. Book review: Science in the Early Roman Empire: Pliny the Elder, His Sources and Influence by Roger French; Frank Greenaway. Isis, 79(2):340–341, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233666.

Berchman:1988:BRB

[2447] Robert M. Berchman. Book review: Middle Platonism and Neoplatonism: The Latin Tradition by Stephen Gersh. Isis, 79(2):341–343, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233667.

Sargent:1988:BRB

[2448] Steven D. Sargent. Book review: Mathematics and Its Applications to Science and Natural Philosophy in the Middle Ages: Essays in Honor of Marshall Clagett by Edward Grant; John E. Murdoch. Isis, 79(2):343– 344, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233668.

Inkster:1988:BRB

[2449] Ian Inkster. Book review: The Rice Economies: Technology and Development in Asian Societies by Francesca Bray. Isis, 79(2):344–345, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233669.

Sivin:1988:BRB

[2450] Nathan Sivin. Book review: The Essential Book of Traditional Chinese Medicine by Liu Yanchi; Fang Tingyu; Chen Laidi. Isis, 79(2):345– 346, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233670.

Dear:1988:BRB

[2451] Peter Dear. Book review: French Higher Education in the Seventeenth and Eighteenth Centuries: A Cultural History by L. W. B. Brockliss. Isis, 79(2):346–347, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233671.

Davis:1988:BRB

[2452] Edward B. Davis. Book review: The Newtonian Revolution by I. Bernard Cohen. Isis, 79(2):347, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233672.

Christensen:1988:BRB

[2453] Thomas Christensen. Book review: Roger North’s Cursory Notes of Musicke (c. 1698–1703): A Physical, Psychological and Critical Theory by Roger North; Mary Chan; Jamie C. Kassler; Roger North’s Writings on Music to c. 1703: A Set of Analytic Indexes by Roger North; Janet D. Hine; Mary Chan; Jamie C. Kassler. Isis, 79(2):347–348, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233673.

Shea:1988:BRBa

[2454] William R. Shea. Book review: Galileo: Heretic by Pietro Redondi; Raymond Rosenthal. Isis, 79(2):348–350, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233674.

Barton:1988:BRBa

[2455] Ruth Barton. Book review: Chemistry and Theology in Mid-Victorian London: The Diary of Herbert McLeod, 1860–1870 by Herbert McLeod; Frank A. J. L. James. Isis, 79(2):350–351, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233675.

Aaserud:1988:BRN

[2456] Finn Aaserud. Book review: Niels Bohr: The Philosophical Writings of Niels Bohr. Isis, 79(2):351–352, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233676.

Reingold:1988:BRB

[2457] Nathan Reingold. Book review: By the Bomb’s Early Light: American Thought and Culture at the Dawn of the Atomic Age by Paul Boyer. Isis, 79(2):352–354, June 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233677.

Fries:1988:BRB

[2458] Sylvia Doughty Fries. Book review: The Johnson Years. Volume II: Vietnam, the Environment, and Science by Robert A. Divine. Isis, 79 (2):354–355, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233678.

Kragh:1988:BRB

[2459] Helge Kragh. Book review: Schrödinger: Centenary Celebration of a Polymath by C. W. Kilmister. Isis, 79(2):355–356, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233679.

Schweber:1988:BRB

[2460] S. S. Schweber. Book review: Behram N. Kurşunoğlu and Eugene P. Wigner: Reminiscences about a Great Physicist: Paul Adrien Maurice Dirac. Isis, 79(2):356–357, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233680.

Anonymous:1988:BMb

[2461] Anonymous. Back matter. Isis, 79(2):358–362, June 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233681.

Anonymous:1988:FMc

[2462] Anonymous. Front matter. Isis, 79(3):363–367, September 1988. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234670.

Sturchio:1988:EAE

[2463] Jeffrey L. Sturchio. Editorial: Artifact and experiment. Isis, 79(3): 368–372, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234671.

Shapin:1988:HES

[2464] Steven Shapin. The house of experiment in Seventeenth-Century England. Isis, 79(3):373–404, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234672.

Kohlstedt:1988:CCN

[2465] Sally Gregory Kohlstedt. Curiosities and cabinets: Natural history museums and education on the antebellum campus. Isis, 79(3):405–426, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234673.

Hacking:1988:TOR

[2466] Ian Hacking. Telepathy: Origins of randomization in experimental design. Isis, 79(3):427–451, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234674.

Daston:1988:BRB

[2467] Lorraine J. Daston. Book review: The Origins of Museums: The Cabinet of Curiosities in Sixteenth- and Seventeenth-Century Europe. Isis, 79(3): 452–467, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234675.

Galison:1988:BRB

[2468] Peter Galison. Book review: The Neglect of Experiment. Isis, 79(3): 467–470, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234676.

Zupko:1988:BRB

[2469] Ronald Edward Zupko. Book review: Nesting Weights, Einsatzgewichte, and Piles à Godet: A Catalog of Nested Cup Weights in the Edward Clark Streeter Collection of Weights and Measures by Ellen Zak Danforth. Isis, 79(3):471, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234677.

Pickering:1988:BRB

[2470] Andy Pickering. Book review: How Experiments End by Peter Galison. Isis, 79(3):472–473, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234678.

Lawrence:1988:BRB

[2471] Christopher Lawrence. Book review: Art and Artefact in Laboratory Science: A Study of Shop Work and Shop Talk in a Research Laboratory by Michael Lynch. Isis, 79(3):473, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234679.

Rigden:1988:BRB

[2472] John S. Rigden. Book review: Models of Scientific Development and the Case of Nuclear Magnetic Resonance by Henk Zandvoort. Isis, 79 (3):473–474, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234680.

Holton:1988:EBD

[2473] Gerald Holton and S. S. Schweber. Eloge: Bern Dibner, 1878–1988. Isis, 79(3):475–477, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234681.

Hall:1988:ELW

[2474] Bert S. Hall. Eloge: Lynn White, Jr., 29 April 1907–30 March 1987. Isis, 79(3):478–481, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234682.

Norberg:1988:BRB

[2475] Arthur L. Norberg. Book review: Thomas A. Edison, by Thomas E. Jeffrey; Thomas A. Edison Papers: Motion Picture Catalogs by American Producers and Distributors, 1894–1908: A Microfilm Edition by Charles Musser; Edison’s Electric Light: Biography of an Invention by Robert Friedel; Paul Israel; Bernard S. Finn. Edison Papers: A Selective Microfilm Edition. Isis, 79(3):482–486, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234683.

Golinski:1988:BRBb

[2476] Jan Golinski. Book review: La méthode philosophique en histoire des sciences: Textes 1914–1939 by Hélène Metzger; Gad Freudenthal. Isis, 79 (3):487–488, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234684.

Barnes:1988:BRB

[2477] Barry Barnes. Book review: Knowledge and Power: Toward a Political Philosophy of Science by Joseph Rouse. Isis, 79(3):488–489, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234685.

Pauly:1988:BRB

[2478] Philip J. Pauly. Book review: Historical Writing on American Science: Perspectives and Prospects by Sally Gregory Kohlstedt; Margaret W. Rossiter. Isis, 79(3):489–491, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234686.

Knight:1988:BRBa

[2479] David Knight. Book review: The Origin and Development of Scholarly Historical Periodicals by Margaret F. Steig. Isis, 79(3):491, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234687.

Guerrini:1988:BRB

[2480] Anita Guerrini. Book review: The Prism of Science. The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science. Volume II by Edna Ullmann-Margalit. Isis, 79(3):492, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234688.

Schmaus:1988:BRB

[2481] Warren Schmaus. Book review: Fact and Method: Explanation, Confirmation, and Reality in the Natural and Social Sciences by Richard W. Miller. Isis, 79(3):492–493, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234689.

White:1988:BRB

[2482] Hayden White. Book review: The Growth of Minds and Cultures: A Unified Theory of the Structures of Human Experience by Willem H. Vanderburg. Isis, 79(3):493–494, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234690.

Weiss:1988:BRB

[2483] Burghard Weiss. Book review: Wissenschaftsgeschichte auf den Versammlungen der Gesellschaft Deutscher Naturforscher und ärzte, 1822– 1972: Bibliographie der Vorträge und allgemeine Übersicht by Dietrich von Engelhardt. Isis, 79(3):494–495, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234691.

McClelland:1988:BRB

[2484] Charles E. McClelland. Book review: The Academic Profession: National, Disciplinary, and Institutional Settings by Burton R. Clark. Isis, 79(3):496–497, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234692.

Jones:1988:BRB

[2485] Paul R. Jones. Book review: American Doctors and German Universities: A Chapter in International Intellectual Relations, 1870–1914 by Thomas Neville Bonner. Isis, 79(3):497–498, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234693.

Kuklick:1988:BRBa

[2486] Henrika Kuklick. Book review: Social Science Research and the Government: Comparative Essays on Britain and the United States by Martin Bulmer. Isis, 79(3):498–499, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234694.

Fissell:1988:BRB

[2487] Mary E. Fissell. Book review: The Making of the Modern Body: Sexuality and Society in the Nineteenth Century by Catherine Gallagher; Thomas Laqueur. Isis, 79(3):499–500, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234695.

Hiddinga:1988:BRB

[2488] Anja Hiddinga. Book review: False Prophets: Fraud and Error in Science and Medicine by Alexander Kohn. Isis, 79(3):500–503, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234696.

Rowe:1988:BRBc

[2489] David E. Rowe. Book review: Die Universität Göttingen unter dem Nationalsozialismus: Das verdrängte Kapitel ihrer 250jährigen Geschichte by Heinrich Becker; Hans-Joachim Dahms; Cornelia Wegeler. Isis, 79 (3):503–505, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234697.

Morris:1988:BRB

[2490] Peter Morris. Book review: Die unheilige Allianz der I. G. Farben: Eine Interessengemeinschaft im Dritten Reich by Joseph Borkin; Industry and Ideology: I.G. Farben in the Nazi Era by Peter Hayes. Isis, 79(3):505– 507, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234698.

Stokes:1988:BRB

[2491] Raymond Stokes. Book review: The Paperclip Conspiracy: The Hunt for the Nazi Scientists by Tom Bower. Isis, 79(3):507–508, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234699.

Hayes:1988:BRB

[2492] Peter Hayes. Book review: Reactionary Modernism: Technology, Culture and Politics in Weimar and the Third Reich by Jeffrey Herf. Isis, 79(3): 508–509, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234700.

Weininger:1988:BRB

[2493] Stephen J. Weininger. Book review: One Culture: Essays in Science and Literature by George Levine; Alan Rauch. Isis, 79(3):509–510, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234701.

Moore:1988:BRB

[2494] James R. Moore. Book review: The Origins of Agnosticism: Victorian Unbelief and the Limits of Knowledge by Bernard Lightman. Isis, 79 (3):510–511, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234702.

Roberts:1988:BRB

[2495] Jon H. Roberts. Book review: The Sciences and Theology in the Twentieth Century by A. R. Peacocke. Isis, 79(3):511–512, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234703.

Grabiner:1988:BRB

[2496] Judith V. Grabiner. Book review: Mathematical People: Profiles and Interviews by Donald J. Albers; G. L. Alexanderson. Isis, 79(3):513–514, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234704.

Barton:1988:BRBb

[2497] Ruth Barton. Book review: Notes towards a History of Teaching Statistics by John Bibby. Isis, 79(3):514, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234705.

Fruton:1988:BRB

[2498] Joseph S. Fruton. Book review: Die Konstitutionserforschung der Alkaloide: Die Pyridin-Piperidin-Gruppe by Jürgen Müller. Isis, 79(3): 514–515, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234706.

Blondel:1988:BRB

[2499] Christine Blondel. Book review: From Falling Bodies to Radio Waves: Classical Physicists and Their Discoveries by Emilio Segré. Isis, 79(3): 515, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234707.

Hunt:1988:BRBa

[2500] Bruce J. Hunt. Book review: A History of the Theories of Aether and Electricity by Edmund Whittaker. Isis, 79(3):515–516, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234708.

Rainger:1988:BRBb

[2501] Ronald Rainger. Book review: A Short History of Vertebrate Palaeontology by Eric Buffetaut. Isis, 79(3):516–517, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234709.

Montgomery:1988:BRB

[2502] William Montgomery. Book review: Das System der Medusen: Textband und Atlasband by Ernst Haeckel. Isis, 79(3):517–518, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234710.

Paul:1988:BRBb

[2503] Diane B. Paul. Book review: Beyond the Gene: Cytoplasmic Inheritance and the Struggle for Authority in Genetics by Jan Sapp. Isis, 79(3):

518–519, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234711.

Tuchman:1988:BRB

[2504] Arleen Tuchman. Book review: Medizin und Krieg: Vom Dilemma der Heilberufe 1865 bis 1985 by Johanna Bleker; Heinz-Peter Schmiedebach. Isis, 79(3):519–520, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234712.

Parisi:1988:BRBa

[2505] Thomas Parisi. Book review: A Godless Jew: Freud, Atheism, and the Making of Psychoanalysis by Peter Gay. Isis, 79(3):520–521, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234713.

Swann:1988:BRB

[2506] John P. Swann. Book review: Medical Science and Medical Industry: The Formation of the American Pharmaceutical Industry by Jonathan Liebenau. Isis, 79(3):521–523, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234714.

Kline:1988:BRB

[2507] Ronald Kline. Book review: The History of Television, 1880 to 1941 by Albert Abramson. Isis, 79(3):523–524, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234715.

Seely:1988:BRBa

[2508] Bruce E. Seely. Book review: Los Angeles and the Automobile: The Making of the Modern City by Scott L. Bottles. Isis, 79(3):524–525, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234716.

Wilde:1988:BRB

[2509] Mark Wilde. Book review: Three Farms: Making Milk, Meat, and Money from the American Soil by Mark Kramer. Isis, 79(3):525–526, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234717.

Flink:1988:BRB

[2510] James J. Flink. Book review: Building the American Highway System: Engineers as Policy Makers by Bruce E. Seely. Isis, 79(3):526–527, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234718.

Misa:1988:BRB

[2511] Thomas J. Misa. Book review: Sheffield Steel and America: A Century of Commercial and Technological Interdependence, 1830–1930 by Geoffrey Tweedale. Isis, 79(3):527–528, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234719.

Todd:1988:BRB

[2512] Edmund N. Todd. Book review: Unternehmensstrategien und technischer Fortschritt: Die deutsche und die britische Stahlindustrie 1865–1895 by Ulrich Wengenroth. Isis, 79(3):528–529, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234720.

Strauss:1988:BRB

[2513] Linda M. Strauss. Book review: Histoire de l’informatique. (Histoire des Sciences.) by Philippe Breton. Isis, 79(3):529–530, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234721.

Pfaffenberger:1988:BRB

[2514] Bryan Pfaffenberger. Book review: High-Tech Society: The Story of the Information Technology Revolution by Tom Forester. Isis, 79(3):530–531, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234722.

Mahoney:1988:BRB

[2515] Michael S. Mahoney. Book review: Beyond the Gutenberg Galaxy: Microcomputers and the Emergence of Post-Typographic Culture by Eugene F. Provenzo. Isis, 79(3):531–532, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234723.

Moran:1988:BRB

[2516] Bruce T. Moran. Book review: Astrology, Science and Society: Historical Essays by Patrick Curry. Isis, 79(3):532–534, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234724.

Mauskopf:1988:BRB

[2517] Seymour H. Mauskopf. Book review: Pseudo-Science and Society in Nineteenth-Century America by Arthur Wrobel. Isis, 79(3):534–535, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234725.

Scarborough:1988:BRB

[2518] John Scarborough. Book review: Galen on Bloodletting: A Study of the Origins, Development and Validity of His Opinions, with a Translation of the Three Works by Peter Brain. Isis, 79(3):535–536, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234726.

Osborne:1988:BRB

[2519] Catherine Osborne. Book review: The Greek Cosmologists. Volume I: The Formation of the Atomic Theory and Its Earliest Critics by David Furley. Isis, 79(3):536–537, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234727.

Schofield:1988:BRB

[2520] Malcolm Schofield. Book review: Rethinking Early Greek Philosophy: Hippolytus of Rome and the Presocratics by Catherine Osborne. Isis, 79 (3):537–538, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234728.

Herlihy:1988:BRB

[2521] David Herlihy. Book review: The Restoration of Perfection: Labor and Technology in Medieval Culture by George Ovitt. Isis, 79(3):538–539, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234729.

Wilson:1988:BRB

[2522] Curtis Wilson. Book review: Kepler’s Geometrical Cosmology by J. V. Field. Isis, 79(3):539–540, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234730.

Vucinich:1988:BRBa

[2523] Alexander Vucinich. Book review: Mikhail Vasilievich Lomonosov: His Life and Work by G. E. Pavlova; A. S. Federov; Arthur Aksenov; Richard Hainsworth. Isis, 79(3):540–541, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234731.

Rees:1988:BRB

[2524] Graham Rees. Book review: Francis Bacon’s Philosophy of Science: An Account and Reappraisal by Peter Urbach. Isis, 79(3):541–542, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234732.

Feingold:1988:BRB

[2525] Mordechai Feingold. Book review: English Science, Bacon to Newton by Brian Vickers. Isis, 79(3):542–543, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234733.

Shea:1988:BRBb

[2526] William R. Shea. Book review: Reinterpreting Galileo by William A. Wallace. Isis, 79(3):543–544, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234734.

Enros:1988:BRB

[2527] Philip C. Enros. Book review: Memoir of the Life and Labours of the Late Charles Babbage Esq. F.R.S., by H. W. Buxton, Anthony Hyman. Isis, 79(3):544, September 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234735.

Miller:1988:BRBb

[2528] Jane A. Miller. Book review: Henry Shaw: His Life and Legacies by William Barnaby Faherty. Isis, 79(3):544–545, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234736.

Lurie:1988:BRB

[2529] Edward Lurie. Book review: Nathaniel Southgate Shaler and the Culture of American Science by David N. Livingstone. Isis, 79(3):545–546, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234737.

Dauben:1988:BRBa

[2530] Joseph W. Dauben. Book review: Historical Perspectives on Peirce’s Logic of Science: A History of Science by Charles S. Peirce; Carolyn Eisele. Isis, 79(3):546–548, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234738.

Cornell:1988:BRB

[2531] Thomas D. Cornell. Book review: Alvarez: Adventures of a Physicist by Luis W. Alvarez; Discovering Alvarez: Selected Works of Luis W. Alvarez, with Commentary by His Students and Colleagues by W. Peter Trower. Isis, 79(3):548–549, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234739.

Coopersmith:1988:BRB

[2532] Jonathan Coopersmith. Book review: Krzhizhanovsky by Vladimir Kartsev; Oleg Glebov. Isis, 79(3):550, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234740.

Wheaton:1988:BRB

[2533] Bruce R. Wheaton. Book review: Erinnerungen eines deutschen Naturforschers by Johannes Stark; Andreas Kleinert. Isis, 79(3):550–551, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234741.

Cuff:1988:BRB

[2534] Robert D. Cuff. Book review: The Fall of the Bell System: A Study in Prices and Politics by Peter Temin; Louis Galambos. Isis, 79(3):551–552, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234742.

Anonymous:1988:BMc

[2535] Anonymous. Back matter. Isis, 79(3):553–558, September 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234743.

Anonymous:1988:FMd

[2536] Anonymous. Front matter. Isis, 79(4):559–563, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234745.

Rosenberg:1988:EWT

[2537] Charles E. Rosenberg. Editorial: Woods or trees? Ideas and actors in the history of science. Isis, 79(4):564–570, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234746.

Cohen:1988:PST

[2538] I. Bernard Cohen. The publication of science, technology and society: Circumstances and consequences. Isis, 79(4):571–582, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234747.

Gieryn:1988:DSR

[2539] Thomas F. Gieryn. Distancing science from religion in Seventeenth-Century England. Isis, 79(4):582–593, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234748.

Shapin:1988:UMT

[2540] Steven Shapin. Understanding the Merton thesis. Isis, 79(4):594–605, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234749.

Merton:1988:MES

[2541] Robert K. Merton. The Matthew Effect in science, II: Cumulative advantage and the symbolism of intellectual property. Isis, 79(4):606–623, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234750.

Numbers:1988:GFW

[2542] Ronald L. Numbers. George Frederick Wright: From Christian Darwinist to fundamentalist. Isis, 79(4):624–645, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234751.

Hill:1988:DTG

[2543] David K. Hill. Dissecting trajectories: Galileo’s early experiments on projectile motion and the law of fall. Isis, 79(4):646–668, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234752.

Lox:1988:ASC

[2544] F. Lox and F. Vandamme. Award of the Sarton Chair to Robert K. Merton, at the University of Ghent. Isis, 79(4):669–670, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234753.

Konvitz:1988:BRB

[2545] Josef W. Konvitz. Book review: The History of Cartography, by J. B. Harley; David Woodward. Volume I: Cartography in Prehistoric, Ancient, and Medieval Europe and the Mediterranean. Isis, 79(4):671–675, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234754.

John:1988:BRB

[2546] Richard R. John, Jr. Book review: The Control Revolution: Technological and Economic Origins of the Information Society. Isis, 79(4): 675–679, December 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234755.

Reed:1988:BRBb

[2547] James Reed. Book review: Paths into American Culture: Psychology, Medicine, and Morals by John C. Burnham. Isis, 79(4):680–681, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234756.

Pyenson:1988:BRBc

[2548] Lewis Pyenson. Book review: An Introduction to the Historiography of Science by Helge Kragh. Isis, 79(4):681–682, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234757.

Vucinich:1988:BRBb

[2549] Alexander Vucinich. Book review: Nauka: Istoriia i istoriografiia XIX-XX vv (Science: History and Historiography in the Nineteenth and Twentieth Centuries) by L. A. Markova. Isis, 79(4):682–683, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234758.

Reeds:1988:BRBb

[2550] Karen Reeds. Book review: Flora Portrayed: Classics of Botanical Art from the Hunt Institute Collections by John V. Brindle; James J. White. Isis, 79(4):683–684, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234759.

Nye:1988:BRBb

[2551] Mary Jo Nye. Book review: The Nobel Population, 1901–1937: A Census of the Nominators and Nominees for the Prizes in Physics and Chemistry by Elisabeth Crawford; J. L. Heilbron; Rebecca Ullrich. Isis, 79(4):684, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234760.

Jayawardene:1988:BRB

[2552] S. A. Jayawardene. Book review: A Chronology of the History of Science, 1450–1900 by Robert Mortimer Gascoigne. Isis, 79(4):685, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234761.

Elliott:1988:BRB

[2553] Clark A. Elliott. Book review: Handbook for Research in American History: A Guide to Bibliographies and Other Reference Works by Francis Paul Prucha. Isis, 79(4):685–686, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234762.

Thackray:1988:BRB

[2554] John C. Thackray. Book review: Geologists and the History of Geology: An International Bibliography from the Origins to 1978: Supplement 1979–1984 and Additions by William A. S. Sarjeant. Isis, 79(4): 686–687, December 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234763.

Deiss:1988:BRBb

[2555] William A. Deiss. Book review: Catalogue of Portraits of Naturalists, Mostly Botanists, in the Collections of the Hunt Institute, the Linnean Society of London, and the Conservatoire et Jardin Botaniques de la Ville de Genéve by Michael T. Stieber; Anita L. Karg; Margot Walker; Gavin D. R. Bridson; Hervé M. Burdet; Marie M. Chautemps; Tina Moruzzi-Bayo; Guide to the Botanical Records and Papers in the Archives of the Hunt Institute, Part 2 by Michael T. Stieber; Anita L. Karg; Catalogue of the Botanical Art Collection at the Hunt Institute by James J. White; Elizabeth R. Smith. Isis, 79(4):687–689, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234764.

Burian:1988:BRB

[2556] Richard Burian. Book review: Explaining Science: A Cognitive Approach by Ronald N. Giere. Isis, 79(4):689–691, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234765.

Leplin:1988:BRBb

[2557] Jarrett Leplin. Book review: Construction and Constraint: The Shaping of Scientific Reality by Ernan McMullin. Isis, 79(4):691–692, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234766.

Hansen:1988:BRB

[2558] James R. Hansen. Book review: Engineering the New South: Georgia Tech, 1885–1985 by Robert C. McMath,; Ronald H. Bayor; James E. Brittain; Lawrence Foster; August W. Giebelhaus; Germaine M. Reed. Isis, 79(4):692–693, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234767.

Plotkin:1988:BRB

[2559] Howard Plotkin. Book review: James Lick’s Monument: The Saga of Captain Richard Floyd and the Building of the Lick Observatory by Helen Wright. Isis, 79(4):693–694, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234768.

Polanco:1988:BRB

[2560] Xavier Polanco. Book review: The Knowledge Society: The Growing Impact of Scientific Knowledge on Social Relations by Gernot Böhme; Nico Stehr. Isis, 79(4):694–695, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234769.

Haber:1988:BRB

[2561] Samuel Haber. Book review: The Professions in American History by Nathan O. Hatch. Isis, 79(4):695–696, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234770.

Kidwell:1988:BRB

[2562] Peggy Kidwell. Book review: The Sound of Our Own Voices: Women’s Study Clubs, 1860–1910 by Theodora Penny Martin. Isis, 79(4):696–697, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234771.

Blume:1988:BRB

[2563] Stuart Blume. Book review: Politique de la science et technologie: Evolution de la politique de recherche: France, Royaume-Uni, Allemagne Federale, Japon, Etats-Unis by V. Thévenin. Isis, 79(4):697–698, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234772.

Cantor:1988:BRB

[2564] Geoffrey Cantor. Book review: The Rhetoric of the Human Sciences: Language and Argument in Scholarship and Public Affairs by John S. Nelson; Allan Megill; Donald N. McCloskey. Isis, 79(4):698–699, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234773.

Livingston:1988:BRB

[2565] James C. Livingston. Book review: Hopkins in the Age of Darwin by Tom Zaniello. Isis, 79(4):699–700, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234774.

Dauben:1988:BRBb

[2566] Joseph W. Dauben. Book review: Georg Cantor, 1845–1918 by Walter Purkert; Hans Joachim Ilgauds. Isis, 79(4):700–702, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234775.

Aspray:1988:BRB

[2567] William Aspray. Book review: Science, Computers, and People: From the Tree of Mathematics by Stanislaw Ulam; Mark C. Reynolds; Gian-Carlo Rota. Isis, 79(4):702–703, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234776.

Dick:1988:BRBb

[2568] Steven J. Dick. Book review: Darkness at Night: A Riddle of the Universe by Edward Harrison. Isis, 79(4):703–704, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234777.

Crowe:1988:BRB

[2569] Michael J. Crowe. Book review: Science and Objectivity: Episodes in the History of Astronomy by Norriss S. Hetherington. Isis, 79(4):704–705, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234778.

Aaboe:1988:BRB

[2570] Asger Aaboe. Book review: Early Astronomy: From Babylonia to Copernicus by W. M. O’Neil. Isis, 79(4):705–706, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234779.

Stuewer:1988:BRB

[2571] Roger H. Stuewer. Book review: La découverte du neutron. (1920–1936) by Jules Six. Isis, 79(4):706–707, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234780.

Slotten:1988:BRB

[2572] Hugh R. Slotten. Book review: Chemical Warfare by Edward M. Spiers. Isis, 79(4):707–708, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234781.

Clarke:1988:BRB

[2573] Edwin Clarke. Book review: A History of Neurophysiology in the Nineteenth Century by Mary A. B. Brazier. Isis, 79(4):708–709, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234782.

Blustein:1988:BRB

[2574] Bonnie Ellen Blustein. Book review: Nineteenth-Century Origins of Neuroscientific Concepts by Edwin Clarke; L. S. Jacyna. Isis, 79(4):709–710, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234783.

Parisi:1988:BRBb

[2575] Thomas Parisi. Book review: Medicine, Mind, and the Double Brain: A Study in Nineteenth-Century Thought by Anne Harrington. Isis, 79 (4):710–712, December 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234784.

Larson:1988:BRBb

[2576] James Larson. Book review: Hendrik Adriaan van Reede tot Drakenstein (1636–1691) and Hortus Malabaricus: A Contribution to the History of Dutch Colonial Botany by J. Heniger. Isis, 79(4):712–713, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234785.

Kuklick:1988:BRBb

[2577] Henrika Kuklick. Book review: Sociology and Scientism: The American Quest for Objectivity, 1880–1940 by Robert C. Bannister. Isis, 79(4): 713–714, December 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234786.

Schabas:1988:BRB

[2578] Margaret Schabas. Book review: Economic History and the History of Economics by Mark Blaug. Isis, 79(4):714–715, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234787.

Campbell-Kelly:1988:BRB

[2579] Martin Campbell-Kelly. Book review: The First Electronic Computer: The Atanasoff Story by Alice R. Burks; Arthur W. Burks. Isis, 79(4): 715–716, December 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234788.

Seely:1988:BRBb

[2580] Bruce E. Seely. Book review: The Automobile Age by James J. Flink. Isis, 79(4):716–717, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234789.

Staudenmaier:1988:BRB

[2581] John M. Staudenmaier. Book review: The Civilized Engineer by Samuel C. Florman. Isis, 79(4):717–718, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234790.

Israel:1988:BRB

[2582] Paul Israel. Book review: The American Telegrapher: A Social History, 1860–1900 by Edwin Gabler. Isis, 79(4):718–720, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234791.

Warner:1988:BRBb

[2583] Deborah Jean Warner. Book review: Mothers of Invention: From the Bra to the Bomb, Forgotten Women and Their Unforgettable Ideas by Ethlie Ann Vare; Greg Ptacek. Isis, 79(4):720, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234792.

Livesey:1988:BRB

[2584] Steven J. Livesey. Book review: Schools and Scholars in Fourteenth-Century England by William J. Courtenay. Isis, 79(4):720–721, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234793.

Gill:1988:BRB

[2585] Mary Louise Gill. Book review: Philosophical Issues in Aristotle’s Biology by Allan Gotthelf; James G. Lennox. Isis, 79(4):721–722, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234794.

Riddle:1988:BRB

[2586] John M. Riddle. Book review: Der Liber servitoris des Abulkasis (936– 1013): Übersetzung, Kommentar und Nachdruck der Textfassung von 1471 by Abu al-Qasim al-Zahrawi; Marianne Engeser. Isis, 79(4):722– 723, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234795.

Saliba:1988:BRBb

[2587] George Saliba. Book review: Islamic Technology: An Illustrated History by Ahmad Y. al-Hassan; Donald R. Hill. Isis, 79(4):723–724, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234796.

Kiger:1988:BRB

[2588] Robert W. Kiger. Book review: Science and Civilisation in China. Volume VI: Biology and Biological Technology. Part I: Botany by Joseph Needham; Lu Gwei-Djen; Huang Hsing-Tsung. Isis, 79(4):724–725, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234797.

Pan:1988:BRB

[2589] Jixing Pan. Book review: Science and Civilisation in China. Volume V: Chemistry and Chemical Technology. Part VII: Military Technology: The Gunpowder Epic by Joseph Needham. Isis, 79(4):725–727, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234798.

Goddu:1988:BRB

[2590] André Goddu. Book review: The Genesis of the Copernican World by Hans Blumenberg; Robert M. Wallace. Isis, 79(4):727–728, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234799.

Phillips:1988:BRB

[2591] Carla Rahn Phillips. Book review: Power and Penury: Government, Technology, and Science in Philip II’s Spain by David C. Goodman. Isis, 79(4):728–729, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234800.

Glick:1988:BRB

[2592] Thomas F. Glick. Book review: Vida y técnica en el renacimiento: Manuscrito que escribió, en el siglo XVI, Francisco Lobato vecino de Medina del Campo by Francisco Lobato; Nicolás García Tapia; José A. García–Diego. Isis, 79(4):729–730, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234801.

Cueto:1988:BRB

[2593] Marcos Cueto. Book review: Alvaro Alonso Barba (1569–1662): Investigaciones sobre su vida y obra by Josep M. Barnadas. Isis, 79(4):730–731, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234802.

Sasaki:1988:BRB

[2594] Chikara Sasaki. Book review: Exercises pour les éléments des solides: Progymnasmata de solidorum elementis: Essai en complement d’Euclide by René Descartes; Pierre Costabel. Isis, 79(4):731–732, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234803.

Dooley:1988:BRB

[2595] Brendan Dooley. Book review: Introduzione al pensiero di Giovanni Battista Hodierna: Filosofo matematico e astronomo dei primi Gattopardi by Mario Pavone. Isis, 79(4):732–733, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234804.

Stocking:1988:BRB

[2596] George W. Stocking, Jr. Book review: Isaac L Peyrère (1596–1676): His Life, Work, and Influence by Richard H. Popkin. Isis, 79(4):733–734, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234805.

Brandt:1988:BRB

[2597] Reinhard Brandt. Book review: Riforma della metafisica e sapere scientifico: Saggio su J. H. Lambert (1728–1777) by Fabio Todesco. Isis, 79(4):734, December 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234806.

Finocchiaro:1988:BRB

[2598] Maurice A. Finocchiaro. Book review: The Idea of Unification in Galileo’s Epistemology by Józef M. ’Zyciński. Isis, 79(4):734–735, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234807.

Hunt:1988:BRBb

[2599] Bruce J. Hunt. Book review: James Clerk Maxwell and the Theory of the Electromagnetic Field by John Hendry. Isis, 79(4):735–736, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234808.

Knight:1988:BRBb

[2600] David Knight. Book review: A Preliminary Discourse on the Study of Natural Philosophy by John F. W. Herschel. Isis, 79(4):736–737, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234809.

Lowood:1988:BRBb

[2601] Henry Lowood. Book review: Il laboratorio borghese: Scienza e politica nella Germania dell’ Ottocento by Pierangelo Schiera. Isis, 79(4):737– 738, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234810.

Nelson:1988:BRB

[2602] Daniel Nelson. Book review: Joseph Wharton: Quaker Industrial Pioneer by W. Ross Yates. Isis, 79(4):738–739, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234811.

Pang:1988:BRB

[2603] Alex Soojung-Kim Pang. Book review: Radar Days by E. G. Bowen. Isis, 79(4):739–740, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234812.

Pestre:1988:BRB

[2604] Dominique Pestre. Book review: Louis de Broglie: Un itinéraire scientifique by Louis de Broglie; Georges Lochak. Isis, 79(4):740–741, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234813.

Bowden:1988:BRB

[2605] Martyn J. Bowden. Book review: Carl O. Sauer: A Tribute by Martin S. Kenzer. Isis, 79(4):741–743, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234814.

Hecht:1988:BRB

[2606] Gabrielle Hecht. Book review: Making and Marketing Arms: The French Experience and Its Implications for the International System by Edward A. Kolodziej. Isis, 79(4):743–744, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234815.

Roland:1988:BRB

[2607] Alex Roland. Book review: Wild Blue Yonder: Money, Politics, and the B-1 Bomber by Nick Kotz. Isis, 79(4):744–745, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234816.

Anonymous:1988:BMd

[2608] Anonymous. Back matter. Isis, 79(4):746–768, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234817.

Anonymous:1988:VI

[2609] Anonymous. Volume information. Isis, 79(4):750–762, December 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234744.

Anonymous:1988:FMe

[2610] Anonymous. Front matter. Isis, 79(5):1–4, ???? 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233804.

Neu:1988:OHT

[2611] John Neu. One hundred thirteenth critical bibliography of the history of science and its cultural influences. Isis, 79(5):5–273, ???? 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233805.

Anonymous:1988:BMe

[2612] Anonymous. Back matter. Isis, 79(5):274, ???? 1988. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233806.

Neu:1989:CBH

[2613] John Neu. Current bibliography of the history of science and its cultural influences, 1989. Isis, 80(??):1–256, ???? 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233415.

Anonymous:1989:BMa

[2614] Anonymous. Back matter. Isis, 80(??):??, ???? 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233416.

Anonymous:1989:FMa

[2615] Anonymous. Front matter. Isis, 80(??):i–xvi, ???? 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233414.

Anonymous:1989:FMb

[2616] Anonymous. Front matter. Isis, 80(1):1–5, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234340.

Numbers:1989:EFI

[2617] Ronald L. Numbers. Editorial: Farewells and introductions. Isis, 80(1): 6–10, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234341.

Golinski:1989:NSP

[2618] J. V. Golinski. A noble spectacle: Phosphorus and the public cultures of science in the early Royal Society. Isis, 80(1):11–39, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234342.

Evans:1989:EMP

[2619] Hughes Evans. European malaria policy in the 1920s and 1930s: The epidemiology of minutiae. Isis, 80(1):40–59, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234343.

Schabas:1989:AMW

[2620] Margaret Schabas. Alfred Marshall, W. Stanley Jevons, and the mathematization of economics. Isis, 80(1):60–73, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234344.

Dales:1989:CWA

[2621] Richard C. Dales. The computistical works ascribed to Robert Grosseteste. Isis, 80(1):74–79, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234345.

Sivin:1989:FIC

[2622] Nathan Sivin. Fifth International Conference on the History of Science in China. Isis, 80(1):80–81, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234346.

Woodward:1989:ISI

[2623] William R. Woodward. International Summer institute in the German democratic Republic. Isis, 80(1):81–82, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234347.

Meadows:1989:EEG

[2624] A. J. Meadows. Eloge: Eric Gray Forbes, (1933–1984). Isis, 80(1):82– 83, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234348.

Brandt:1989:BRBa

[2625] Allan M. Brandt. Book review: Mobilizing against AIDS: The Unfinished Story of a Virus. Isis, 80(1):84–86, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234349.

Brumberg:1989:BRB

[2626] Joan Jacobs Brumberg. Book review: Hypochondria: Woeful Imaginings by Susan Baur. Isis, 80(1):87–88, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234350.

King:1989:BRBa

[2627] Nydia M. King. Book review: Historia general de la farmacia: El medicamento a través del tiempo. Volume I by Guillermo Folch Jou; José María Suñé Arbussá; José Luis Valverde López. Isis, 80(1):88–89, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234351.

Pernick:1989:BRB

[2628] Martin S. Pernick. Book review: No Laughing Matter: Historical Aspects of Anaesthesia by Christopher Lawrence; Ghislaine Lawrence; Huw Geddes; Lorraine Ward. Isis, 80(1):89–90, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234352.

LaBerge:1989:BRBa

[2629] Ann La Berge. Book review: Illness and Self in Society by Claudine Herzlich; Janine Pierret; Elborg Forster. Isis, 80(1):90–91, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234353.

Broman:1989:BRB

[2630] Thomas Broman. Book review: Medizin in Berlin by Rolf Winau. Isis, 80(1):91–92, March 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234354.

Brown:1989:BRB

[2631] Theodore M. Brown. Book review: “For the Welfare of Mankind”: The Commonwealth Fund and American Medicine by A. McGehee Harvey; Susan L. Abrams. Isis, 80(1):92–93, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234355.

Howell:1989:BRBa

[2632] Joel D. Howell. Book review: Making Medical Doctors: Science and Medicine at Vanderbilt since Flexner by Timothy C. Jacobson. Isis, 80 (1):93–94, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234356.

Tomes:1989:BRB

[2633] Nancy Tomes. Book review: Homes for the Mad: Life inside Two Nineteenth-Century Asylums by Ellen Dwyer. Isis, 80(1):94–95, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234357.

Brunton:1989:BRBa

[2634] Deborah Brunton. Book review: Leicestershire’s Lunatics: The Institutional Care of Leicestershire’s Lunatics during the Nineteenth Century by Henry Gilbert Orme; William H. Brock. Isis, 80(1):95, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234358.

Bearn:1989:BRB

[2635] Alexander G. Bearn and Gordon C. F. Bearn. Book review: Use and Abuse of Literary Concepts in Medicine by Joanne Trautmann Banks. Isis, 80(1):95–96, March 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234359.

Rousseau:1989:BRBa

[2636] G. S. Rousseau. Book review: Shelley’s Venomed Melody by Nora Crook; Derek Guiton. Isis, 80(1):96–97, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234360.

Forrester:1989:BRB

[2637] John Forrester. Book review: Psychiatry and the Cinema by Krin Gabbard; Glen O. Gabbard. Isis, 80(1):97–99, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234361.

Cook:1989:BRB

[2638] Harold J. Cook. Book review: Sufferers and Healers: The Experience of Illness in Seventeenth-Century England by Lucinda McCray Beier. Isis, 80(1):99–101, March 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234362.

Weiner:1989:BRB

[2639] Dora B. Weiner. Book review: On Insanity and Its Classification by Vincenzo Chiarugi; George Mora. Isis, 80(1):101–102, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234363.

Higonnet:1989:BRB

[2640] Margaret Higonnet and Patrice Higonnet. Book review: Suicide in Victorian and Edwardian England by Olive Anderson. Isis, 80(1):102–103, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234364.

Sussman:1989:BRB

[2641] George D. Sussman. Book review: Disease and Civilization: The Cholera in Paris, 1832 by François Delaporte; Arthur Goldhammer. Isis, 80(1): 103–104, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234365.

Goldman:1989:BRB

[2642] Joanne Abel Goldman. Book review: Newark: The Nation’s Unhealthiest City, 1832–1895 by Stuart Galishoff. Isis, 80(1):104–105, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234366.

Lawrence:1989:BRBa

[2643] Christopher Lawrence. Book review: Medicine and Society in Wakefield and Huddersfield, 1780–1870 by Hilary Marland. Isis, 80(1):105–106, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234367.

Duffin:1989:BRB

[2644] Jacalyn Duffin. Book review: Morbid Appearances: The Anatomy of Pathology in the Early Nineteenth Century by Russell C. Maulitz. Isis, 80(1):106–107, March 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234368.

Fissell:1989:BRB

[2645] Mary E. Fissell. Book review: Death, Dissection, and the Destitute by Ruth Richardson. Isis, 80(1):107–108, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234369.

Melosh:1989:BRBa

[2646] Barbara Melosh. Book review: Able-Bodied Womanhood: Personal Health and Social Change in Nineteenth-Century Boston by Martha H. Verbrugge. Isis, 80(1):108–109, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234370.

Golden:1989:BRB

[2647] Janet Golden. Book review: Mothers and Medicine: A Social History of Infant Feeding, 1890–1950 by Rima D. Apple. Isis, 80(1):109–110, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234371.

Markowitz:1989:BRB

[2648] Gerald Markowitz. Book review: A History of Neglect: Health Care for Blacks and Mill Workers in the Twentieth–Century South by Edward H. Beardsley. Isis, 80(1):111–112, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234372.

Bates:1989:BRBa

[2649] Barbara Bates. Book review: The Last Crusade: The War on Consumption, 1862–1954 by Mark Caldwell. Isis, 80(1):112–113, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234373.

Cravens:1989:BRB

[2650] Hamilton Cravens. Book review: The Inner World of American Psychiatry, 1890–1940: Selected Correspondence by Gerald N. Grob. Isis, 80(1):113–114, March 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234374.

Weiss:1989:BRB

[2651] Sheila Faith Weiss. Book review: Racial Hygiene: Medicine under the Nazis by Robert N. Proctor. Isis, 80(1):114–115, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234375.

MacKenzie:1989:BRB

[2652] Donald MacKenzie. Book review: The Probabilistic Revolution, by Lorenz Kruger; Lorraine J. Daston; Michael Heidelberger; The Probabilistic Revolution. Volume II: Ideas in Science by Lorenz Kruger; Gerd Gigerenzer; Mary S. Morgan; Classical Probability in the Enlightenment by Lorraine J. Daston. Volume I: Ideas in History. Isis, 80(1):116–124, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234376.

Stanton:1989:BRB

[2653] William Stanton. Book review: The Launching of Modern American Science, 1846–1876. Isis, 80(1):124–128, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234377.

Pyenson:1989:BRJ

[2654] Lewis Pyenson. Book review: Just the facts: The Collected Papers of Albert Einstein, by Albert Einstein; John Stachel; David C. Cassidy; Robert Schulmann; Jürgen Renn; Olga Griminger; Gary Smith; Robert Summerfield; The Collected Papers of Albert Einstein. Volume I: The Early Years, 1879–1902. English Translation by Albert Einstein; Anna Beck; Peter Havas. Volume I: The Early Years, 1879–1902. Isis, 80(1): 129–135, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234378.

Hollinger:1989:BRB

[2655] David A. Hollinger. Book review: The New History and the Old: Critical Essays and Reappraisals by Gertrude Himmelfarb. Isis, 80(1):136, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234379.

Harwit:1989:BRB

[2656] Martin Harwit. Book review: Scientific Discovery: Computational Explorations of the Creative Process by Pat Langley; Herbert A. Simon; Gary L. Bradshaw; Jan M. Zytkow. Isis, 80(1):136–137, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234380.

Feldman:1989:BRB

[2657] Theodore S. Feldman. Book review: La Suède intellectuelle et savante by J. F. Battail. Isis, 80(1):137–138, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234381.

Goldberg:1989:BRB

[2658] Stanley Goldberg. Book review: Einstein’s Legacy: The Unity of Space and Time by Julian Schwinger. Isis, 80(1):138–139, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234382.

Pyne:1989:BRB

[2659] Stephen J. Pyne. Book review: Antarctic Science by D. W. H. Walton. Isis, 80(1):139–141, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234383.

Carlson:1989:BRB

[2660] W. Bernard Carlson. Book review: Inventors at Work by Kenneth A. Brown. Isis, 80(1):141–142, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234384.

Davies:1989:BRB

[2661] Gordon L. Herries Davies. Book review: Nostri plena laboris: An Author Index to the RDS Scientific Journals, 1800–1985 by R. Charles Mollan. Isis, 80(1):142–143, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234385.

Weimerskirch:1989:BRB

[2662] Philip J. Weimerskirch. Book review: Milestones in Science and Technology: The Ready Reference Guide to Discoveries, Inventions, and Facts by Ellis Mount; Barbara A. List. Isis, 80(1):143, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234386.

Churchland:1989:BRB

[2663] Paul M. Churchland. Book review: Observation and Objectivity by Harold I. Brown. Isis, 80(1):143–144, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234387.

Finocchiaro:1989:BRB

[2664] Maurice A. Finocchiaro. Book review: Patterns, Thinking, and Cognition: A Theory of Judgment by Howard Margolis. Isis, 80(1):144–145, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234388.

Koertge:1989:BRB

[2665] Noretta Koertge. Book review: From Knowledge to Wisdom: A Revolution in the Aims and Methods of Science by Nicholas Maxwell. Isis, 80(1):146–147, March 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234389.

Smith:1989:BRB

[2666] Michael L. Smith. Book review: Space Colonization: Technology and the Liberal Arts by Charles H. Holbrow; Allan M. Russell; Gorden F. Sutton. Isis, 80(1):147–148, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234390.

Cozzens:1989:BRB

[2667] Susan E. Cozzens. Book review: Scientific Excellence: Origins and Assessment by Douglas N. Jackson; J. Philippe Rushton. Isis, 80(1):148– 149, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234391.

Reich:1989:BRB

[2668] Leonard S. Reich. Book review: The Essential Alfred Chandler: Essays toward a Historical Theory of Big Business by Alfred Chandler; Thomas K. McCraw. Isis, 80(1):149–150, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234392.

Stine:1989:BRB

[2669] Jeffrey K. Stine. Book review: The New Politics of Science by David Dickson. Isis, 80(1):150–151, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234393.

Secord:1989:BRBa

[2670] J. A. Secord. Book review: Making It Whole: A Victorian Circle and the Shape of Their World by Diana Postlethwaite. Isis, 80(1):151–152, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234394.

Privateer:1989:BRB

[2671] Paul Privateer. Book review: Science and the Imagination. (Annals of Scholarship: Metastudies of the Humanities and Social Sciences, Volume 4, Number 1, Fall 1986) by George S. Rousseau. Isis, 80(1):153– 154, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234395.

Grattan-Guinness:1989:BRBa

[2672] I. Grattan-Guinness. Book review: History and Philosophy of Modern Mathematics by William Aspray; Philip Kitcher. Isis, 80(1):154–155, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234396.

Parshall:1989:BRB

[2673] Karen Hunger Parshall. Book review: Discrete Thoughts: Essays on Mathematics, Science, and Philosophy by Mark Kac; Gian-Carlo Rota; Jacob T. Schwartz; Harry Newman. Isis, 80(1):155–156, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234397.

Rowe:1989:BRB

[2674] David E. Rowe. Book review: A Concise History of Mathematics by Dirk J. Struik. Isis, 80(1):156–157, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234398.

Stairs:1989:BRB

[2675] Allen Stairs. Book review: Speakable and Unspeakable in Quantum Mechanics by J. S. Bell. Isis, 80(1):157–158, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234399.

Caneva:1989:BRB

[2676] Kenneth L. Caneva. Book review: History of Physics: Selected Reprints by Stephen G. Brush. Isis, 80(1):158–159, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234400.

Benfey:1989:BRB

[2677] Theodor Benfey. Book review: Insight in Chemistry by William J. Danaher. Isis, 80(1):159, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234401.

Hiebert:1989:BRB

[2678] Erwin Hiebert. Book review: The Principles of the Theory of Heat, Historically and Critically Elucidated by Ernst Mach; Brian McGuinness. Isis, 80(1):159–161, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234402.

Crosby:1989:BRB

[2679] Alfred W. Crosby. Book review: Western Civilization in Biological Perspective: Patterns in Biohistory by Stephen Boyden. Isis, 80(1):161–162, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234403.

Hagan:1989:BRB

[2680] William J. Hagan, Jr. Book review: The Emergence of Life: Darwinian Evolution from the Inside by Sidney Fox. Isis, 80(1):162–163, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234404.

Magner:1989:BRB

[2681] Lois N. Magner. Book review: Evolution and Human Equality by Stephen Jay Gould. Isis, 80(1):163–164, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234405.

Kingsland:1989:BRB

[2682] Sharon Kingsland. Book review: The American Development of Biology by Ronald Rainger; Keith R. Benson; Jane Maienschein. Isis, 80(1):164– 165, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234406.

Herbert:1989:BRB

[2683] Sandra Herbert. Book review: Darwin and the Emergence of Evolutionary Theories of Mind and Behavior by Robert J. Richards. Isis, 80(1): 165–167, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234407.

Jones:1989:BRBa

[2684] Robert Alun Jones. Book review: Enlightenment and Despair: A History of Sociology by Geoffrey Hawthorn. Isis, 80(1):167–169, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234408.

Ceruzzi:1989:BRB

[2685] Paul Ceruzzi. Book review: Creating the Computer: Government, Industry, and High Technology by Kenneth Flamm. Isis, 80(1):169–170, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234409.

Smulyan:1989:BRB

[2686] Susan Smulyan. Book review: When Old Technologies Were New: Thinking about Communications in the Late Nineteenth Century by Carolyn Marvin. Isis, 80(1):170–171, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234410.

Marx:1989:BRB

[2687] Leo Marx. Book review: New Perspectives on Technology and American Culture by Bruce Sinclair. Isis, 80(1):171–172, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234411.

Dawson:1989:BRB

[2688] Virginia P. Dawson. Book review: The Tender Ship: Government Management of Technological Change by Arthur M. Squires. Isis, 80(1):172– 173, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234412.

Harris:1989:BRB

[2689] J. R. Harris. Book review: The Transfer of Early Industrial Technologies to America by Darwin H. Stapleton. Isis, 80(1):173–174, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234413.

Misa:1989:BRB

[2690] Thomas J. Misa. Book review: The Decline of American Steel: How Management, Labor, and Government Went Wrong by Paul A. Tiffany. Isis, 80(1):174–175, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234414.

Hahm:1989:BRB

[2691] David E. Hahm. Book review: Introduction à la méthode d’Aristote by Jean-Paul Dumont. Isis, 80(1):175–176, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234415.

Drucker:1989:BRB

[2692] Thomas L. Drucker. Book review: The Mathematics of Plato’s Academy: A New Reconstruction by D. H. Fowler. Isis, 80(1):176–177, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234416.

vonStaden:1989:BRBa

[2693] Heinrich von Staden. Book review: Civilization of the Ancient Mediterranean: Greece and Rome by Michael Grant; Rachel Kitzinger. Isis, 80(1):177–178, March 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234417.

Green:1989:BRB

[2694] Monica Green. Book review: A History of Women Philosophers. Volume I: Ancient Women Philosophers, 600 B.C.–500 A.D. by Mary Ellen Waith. Isis, 80(1):178–179, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234418.

Cadden:1989:BRB

[2695] Joan Cadden. Book review: Man and the Beasts: De animalibus (Books 22–26) by Albertus Magnus; James J. Scanlan. Isis, 80(1):179–181, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234419.

Eamon:1989:BRB

[2696] William Eamon. Book review: Spanien und der Basler Buchdruck bis 1600: Ein Querschnitt durch die spanische Geistesgeschichte aus der Sicht einer europäischen Buchdruckerstadt by Carlos Gilly. Isis, 80(1): 181–182, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234420.

Jones:1989:BRBb

[2697] Howard Jones. Book review: Pierre Gassendi: From Aristotelianism to a New Natural Philosophy by Barry Brundell. Isis, 80(1):182–183, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234421.

Henry:1989:BRBa

[2698] John Henry. Book review: The Cultural Meaning of the Scientific Revolution by Margaret C. Jacob. Isis, 80(1):183–184, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234422.

Gillispie:1989:BRB

[2699] Charles C. Gillispie. Book review: De la richesse territoriale du Royaume de France by Antoine-Laurent Lavoisier; Jean-Claude Perrot. Isis, 80(1): 184–185, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234423.

Dear:1989:BRB

[2700] Peter Dear. Book review: Lingua universalis: Kryptologie und Theorie der Universal-sprachen im 16. und 17. Jahrhundert by Gerhard F. Strasser. Isis, 80(1):185–186, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234424.

Hunt:1989:BRBa

[2701] Bruce J. Hunt. Book review: Heinrich Hertz: The Beginning of Microwaves by John H. Bryant. Isis, 80(1):186, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234425.

Rocke:1989:BRB

[2702] A. J. Rocke. Book review: Justus von Liebig und August Wilhelm Hofmann in ihren Briefen: Nachtrage, 1845–1869 by Emil Heuser; Regine Zott; Justus von Liebig und Emil Erlenmeyer in ihren Briefen von 1861–1872 by Emil Heuser. Isis, 80(1):186–187, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234426.

Brittain:1989:BRB

[2703] James E. Brittain. Book review: Hertz and the Maxwellians: A Study and Documentation of the Disvoery of Electromagnetic Wave Radiation, 1873–1894 by J. G. O’Hara; W. Pricha. Isis, 80(1):187–188, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234427.

Gregory:1989:BRB

[2704] Frederick Gregory. Book review: Hegel und die Naturwissenschaften by Michael John Petry. Isis, 80(1):188–189, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234428.

Morris:1989:BRB

[2705] Peter J. T. Morris. Book review: Brazil and the Struggle for Rubber: A Study in Environmental History by Warren Dean. Isis, 80(1):189– 190, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234429.

Decker:1989:BRB

[2706] Hannah S. Decker. Book review: Freud: A Life for Our Time by Peter Gay. Isis, 80(1):190–191, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234430.

Schrepfer:1989:BRB

[2707] Susan R. Schrepfer. Book review: Beauty, Health, and Permanence: Environmental Politics in the United States, 1955–1985 by Samuel P.

Hays; Barbara D. Hays. Isis, 80(1):191–193, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234431.

Mack:1989:BRB

[2708] Pamela Mack. Book review: History of British Space Science by Harrie Massey; M. O. Robins. Isis, 80(1):193, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234432.

Hevly:1989:BRB

[2709] Bruce Hevly. Book review: The Business of Science: Winning and Losing in the High-Tech Age by Simon Ramo. Isis, 80(1):193–194, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234433.

Anonymous:1989:BMb

[2710] Anonymous. Back matter. Isis, 80(1):195–200, March 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234434.

Anonymous:1989:FMc

[2711] Anonymous. Front matter. Isis, 80(2):201–205, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234596.

Segre:1989:VLG

[2712] Michael Segre. Viviani’s life of Galileo. Isis, 80(2):206–231, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234597.

Eagan:1989:DCS

[2713] William E. Eagan. The debate over the Canadian Shield, 1880–1905. Isis, 80(2):232–253, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234598.

Logsdon:1989:MHC

[2714] John M. Logsdon. Missing Halley’s comet: The politics of big science. Isis, 80(2):254–280, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234599.

Schwartz:1989:EWD

[2715] Richard A. Schwartz. Einstein and the war department. Isis, 80(2):281– 284, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234600.

Anonymous:1989:EPH

[2716] Anonymous. Eloge: Paul H. Barrett, 7 November 1917–6 November 1987. Isis, 80(2):285, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234601.

vonStaden:1989:BRBb

[2717] Heinrich von Staden. Book review: The Revolutions of Wisdom: Studies in the Claims and Practice of Ancient Greek Science. Isis, 80(2):286– 288, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234602.

Sturdy:1989:BRB

[2718] Steve Sturdy. Book review: History of the American Physiological Society: The First Century, 1887–1987. Isis, 80(2):289–294, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234603.

Multhauf:1989:BRB

[2719] Robert P. Multhauf. Book review: Frederick E. Brasch and the History of Science by Henry Lowood. Isis, 80(2):295, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234604.

Meinel:1989:BRB

[2720] Christoph Meinel. Book review: Der Ursprung der modernen Wissenschaften: Studien zur Entstehung wissenschaftlicher Disziplinen by Martin Guntau; Hubert Laitko. Isis, 80(2):295–296, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234605.

Porter:1989:BRBa

[2721] Theodore M. Porter. Book review: The American Census: A Social History by Margo J. Anderson. Isis, 80(2):296–298, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234606.

Unguru:1989:BRB

[2722] Sabetai Unguru. Book review: A Mathematical History of Division in Extreme and Mean Ratio by Roger Herz-Fischler. Isis, 80(2):298–299, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234607.

Bedini:1989:BRB

[2723] Silvio A. Bedini. Book review: The Divided Circle: A History of Instruments for Astronomy, Navigation, and Surveying by J. A. Bennett. Isis, 80(2):299–300, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234608.

Fitzgerald:1989:BRB

[2724] Deborah Fitzgerald. Book review: First the Seed: The Political Economy of Plant Biotechnology, 1492–2000 by Jack Ralph Kloppenburg. Isis, 80 (2):300–301, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234609.

Ainley:1989:BRB

[2725] Marianne Gosztonyi Ainley. Book review: Histoire des sciences au Québec by Luc Chartrand; Raymond Duchesne; Yves Gingras. Isis, 80 (2):301–302, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234610.

Barkan:1989:BRB

[2726] Diana Barkan. Book review: Chemiker über Chemiker: Wahlvorschläge zur Aufnahme von Chemikern in die Berliner Akademie 1822–1925, von Eilhard Mitscherlich bis Max Bodenstein by Anneliese Greiner; Fritz Welsh; Wolfgang Girnus. Isis, 80(2):302–303, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234611.

Cabral:1989:BRB

[2727] Regis Cabral. Book review: La psiquiatría Argentina by Antonio Alberto Guerrino. Isis, 80(2):303–304, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234612.

Maulitz:1989:BRB

[2728] Russell C. Maulitz. Book review: Festschrift für Jean Starobinski by Karl Haffter. Isis, 80(2):304–305, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234613.

Toomer:1989:BRB

[2729] G. J. Toomer. Book review: Archimedes by E. J. Dijksterhuis; C. Dikshoorn; Wilbur R. Knorr. Isis, 80(2):305–306, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234614.

Hankinson:1989:BRB

[2730] R. J. Hankinson. Book review: Aristotle: The Desire to Understand by Jonathan Lear. Isis, 80(2):306–307, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234615.

Vallance:1989:BRB

[2731] J. T. Vallance. Book review: The Physical World of the Greeks by S. Sambursky; Merton Dagut; The Physical World of Late Antiquity by S. Sambursky; The Physics of the Stoics by S. Sambursky. Isis, 80(2):307– 308, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234616.

Srinivasan:1989:BRBa

[2732] Raman Srinivasan. Book review: Founders of Sciences in Ancient India by Svami Satya Prakash Sarasvati. Isis, 80(2):308–309, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234617.

Eastwood:1989:BRB

[2733] Bruce S. Eastwood. Book review: A History of Twelfth-Century Western Philosophy by Peter Dronke. Isis, 80(2):309–310, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234618.

Pingree:1989:BRB

[2734] David Pingree. Book review: Islamic Mathematical Astronomy by David A. King. Isis, 80(2):310–311, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234619.

Secord:1989:BRBb

[2735] James A. Secord. Book review: Like Engend’ring like: Heredity and Animal Breeding in Early Modern England by Nicholas Russell. Isis, 80(2):311–312, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234620.

Cueto:1989:BRB

[2736] Marcos Cueto. Book review: Medicine in the New World: New Spain, New France, and New England by Ronald L. Numbers. Isis, 80(2):312, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234621.

Applebaum:1989:BRBa

[2737] Wilbur Applebaum. Book review: Newton’s Principia and Its Legacy by D. G. King-Hele; A. R. Hall. Isis, 80(2):313, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234622.

Calinger:1989:BRBa

[2738] Ronald Calinger. Book review: Beiträge zur Wirkungs- und Rezeptionsgeschichte von Gottfried Wilhelm Leibniz by Albert Heinekamp. Isis, 80(2):313–315, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234623.

Crosland:1989:BRBa

[2739] Maurice Crosland. Book review: Fizikalna rasprava o postanku, naravi i koristi umjetnog zraka by Josip Franjo Domin; Drago Grdenić; Snjezana Pausek-Bazdar. Isis, 80(2):315, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234624.

Turner:1989:BRBa

[2740] G. L’E. Turner. Book review: Josef Anton Nagel: Ein Direktor des physikalischen Kabinettes by Johanna Schönburg–Hartenstein. Isis, 80 (2):315–317, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234625.

Kohlstedt:1989:BRB

[2741] Sally Gregory Kohlstedt. Book review: The Selected Papers of Charles Willson Peale and His Family by Charles Willson Peale; Lillian B. Miller; Sidney Hart. Isis, 80(2):317–318, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234626.

Walton:1989:BRB

[2742] John K. Walton. Book review: Mind-Forg’d Manacles; A History of Madness in England from the Restoration to the Regency by Roy Porter. Isis, 80(2):318–319, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234627.

Williams:1989:BRBa

[2743] L. Pearce Williams. Book review: The Age of Science: The Scientific World-View in the Nineteenth Century by David Knight. Isis, 80(2): 319–320, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234628.

Miller:1989:BRB

[2744] David Philip Miller. Book review: The Amateur and the Professional: Antiquarians, Historians and Archaeologists in Victorian England, 1838– 1886 by Philippa Levine. Isis, 80(2):320–321, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234629.

Munday:1989:BRB

[2745] Pat Munday. Book review: Deutschsprachige Chemielehrbücher (1775– 1850) by Bettina Haupt. Isis, 80(2):321–322, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234630.

Rosenberg:1989:BRB

[2746] Robert Rosenberg. Book review: Oliver Heaviside, Sage in Solitude: The Life, Work, and Times of an Electrical Genius of the Victorian Age by Paul J. Nahin. Isis, 80(2):322–324, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234631.

Rudwick:1989:BRB

[2747] Martin Rudwick. Book review: The Iceland Journal of Henry Holland, 1810 by Henry Holland; Andrew Wawn. Isis, 80(2):324, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234632.

Corsi:1989:BRB

[2748] Pietro Corsi. Book review: Charles Darwin’s Notebooks, 1836–1844: Geology, Transmutation of Species, Metaphysical Enquiries by Charles Darwin; Paul H. Barrett; Peter J. Gautrey; Sandra Herbert; David Kohn; Sydney Smith. Isis, 80(2):324–325, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234633.

McGaw:1989:BRB

[2749] Judith A. McGaw. Book review: To Their Own Soil: Agriculture in the Antebellum North by Jeremy Atack; Fred Bateman. Isis, 80(2):325– 327, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234634.

Rupke:1989:BRB

[2750] Nicholas Rupke. Book review: The Animal Estate: The English and Other Creatures in the Victorian Age by Harriet Ritvo. Isis, 80(2):327– 328, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234635.

Glick:1989:BRB

[2751] Thomas F. Glick. Book review: Las ciencias médicas básicas en la Valencia del siglo XIX by José M. López Piñero; María José Báguena Cervellera; José Luis Barona Vilar; José Luis Fresquet Febrer; María Luz López Terrada; José Pardo Tomás; Vicente L. Salavert Fabiani. Isis, 80(2):328–329, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234636.

Brandt:1989:BRBb

[2752] Allan M. Brandt. Book review: The Cholera Years: The United States in 1832, 1849, and 1866 by Charles E. Rosenberg. Isis, 80(2):329–330, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234637.

Kranakis:1989:BRB

[2753] Eda Kranakis. Book review: Transitions in Engineering: Guillaume Henri Dufour and the Early Nineteenth-Century Cable Suspension Bridge by Tom F. Peters. Isis, 80(2):330–331, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234638.

Paradis:1989:BRB

[2754] James Paradis. Book review: Science and Literature in the Nineteenth Century by J. A. V. Chapple. Isis, 80(2):331–332, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234639.

Morawski:1989:BRB

[2755] Jill Morawski. Book review: Clark University, 1887–1987: A Narrative History by William A. Koelsch. Isis, 80(2):332, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234640.

Lambright:1989:BRB

[2756] W. Henry Lambright. Book review: Science and Technology Advice to the President, Congress, and Judiciary by William T. Golden. Isis, 80(2): 333–334, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234641.

Rip:1989:BRB

[2757] Arie Rip. Book review: Acid Rain and Friendly Neighbors: The Policy Dispute between Canada and the United States by Jurgen Schmandt; Hilliard Roderick. Isis, 80(2):334, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234642.

Stokes:1989:BRB

[2758] Raymond G. Stokes. Book review: Pax Atomica: The Nuclear Defense Debate in West Germany during the Adenauer Era by Mark Cioc. Isis, 80(2):334–335, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234643.

Serwer:1989:BRB

[2759] Daniel Serwer. Book review: The Dragon’s Tail: Radiation Safety in the Manhattan Project, 1942–1946 by Barton C. Hacker. Isis, 80(2):335– 336, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234644.

Wheaton:1989:BRB

[2760] Bruce R. Wheaton. Book review: History of CERN. Volume I: Launching the European Organization for Nuclear Research by Armin Hermann; John Krige; Ulrike Mersits; Dominique Pestre; L. Belloni. Isis, 80(2): 336–338, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234645.

Nye:1989:BRB

[2761] Mary Jo Nye. Book review: Langevin: Science et vigilance by Bernadette Bensaude-Vincent. Isis, 80(2):338–339, June 1989. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234646.

Beller:1989:BRB

[2762] Mara Beller. Book review: The Ghost in the Atom: A Discussion of the Mysteries of Quantum Physics by P. C. W. Davies; J. R. Brown. Isis, 80(2):339–340, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234647.

Elliott:1989:BRBa

[2763] J. Eric Elliott. Book review: Vorgeschichte und Gründung des Kaiser-Wilhelm-Instituts für Kohlenforschung in Mülheim a. d. Ruhr by Manfred Rasch. Isis, 80(2):340–341, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234648.

Rothenberg:1989:BRB

[2764] Marc Rothenberg. Book review: A Man Who Loved the Stars: The Autobiography of John A. Brashear by John A. Brashear. Isis, 80(2): 341, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234649.

Pauly:1989:BRB

[2765] Philip J. Pauly. Book review: The National Geographic Society: One Hundred Years of Adventure and Discovery by C. D. B. Bryan; National Geographic: Behind America’s Lens on the World by Howard S. Abramson. Isis, 80(2):341–343, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234650.

Deacon:1989:BRB

[2766] Margaret B. Deacon. Book review: The Mediterranean Was a Desert: A Voyage of the Glomar Challenger by Kenneth J. Hsü. Isis, 80(2): 343, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234651.

Abir-Am:1989:BRB

[2767] Pnina G. Abir-Am. Book review: Contemporary Classics in the Life Sciences by James T. Barrett. Isis, 80(2):344–345, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234652.

Hacker:1989:BRB

[2768] Barton C. Hacker. Book review: Clouds of Secrecy: The Army’s Germ Warfare Tests over Populated Areas by Leonard A. Cole. Isis, 80(2): 345–346, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234653.

Churchill:1989:BRB

[2769] Frederick B. Churchill. Book review: The Heritage of Experimental Embryology: Hans Spemann and the Organizer by Viktor Hamburger. Isis, 80(2):346–347, June 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234654.

Thom:1989:BRB

[2770] Deborah Thom. Book review: Psychological Testing and American Society 1890–1930 by Michael M. Sokal. Isis, 80(2):347–348, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234655.

Bates:1989:BRBb

[2771] Barbara Bates. Book review: The Tuberculosis Movement: A Public Health Campaign in the Progressive Era by Michael E. Teller. Isis, 80 (2):349–350, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234656.

Staudenmaier:1989:BRB

[2772] John M. Staudenmaier. Book review: Mechanization and Maize: Agriculture and the Politics of Technology Transfer in East Africa by Constance G. Anthony. Isis, 80(2):350–351, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234657.

Frost:1989:BRB

[2773] Robert L. Frost. Book review: Les ingénieurs de la crise: Titre et profession entre les deux guerres by André Grelon. Isis, 80(2):351–352, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234658.

Florman:1989:BRB

[2774] Samuel C. Florman. Book review: Hoover Dam: An American Adventure by Joseph E. Stevens. Isis, 80(2):352–353, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234659.

Ducret:1989:BRB

[2775] Jean-Jacques Ducret. Book review: Constructive Evolution: Origins and Development of Piaget’s Thought by Michael Chapman. Isis, 80(2):353– 354, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234660.

Hoyningen-Huene:1989:BRB

[2776] Paul Hoyningen-Huene. Book review: Constructive Philosophy by Paul Lorenzen; Karl Richard Pavlovic. Isis, 80(2):354–355, June 1989. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234661.

Grattan-Guinness:1989:BRBb

[2777] I. Grattan-Guinness. Book review: Wittgenstein: A Life. Volume I: Young Ludwig, 1889–1921 by Brian McGuinness. Isis, 80(2):355, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234662.

Schmaus:1989:BRB

[2778] Warren Schmaus. Book review: Theories of Explanation by Joseph C. Pitt. Isis, 80(2):356, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234663.

Woolgar:1989:BRB

[2779] Steve Woolgar. Book review: Mundane Reason: Reality in Everyday and Sociological Discourse by Melvin Pollner. Isis, 80(2):356–358, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234664.

Lynaugh:1989:BRB

[2780] Joan E. Lynaugh. Book review: American Nursing: A Biographical Dictionary by Vern L. Bullough; Olga Maranjian Church; Alice P. Stein; Dictionary of American Nursing Biography by Martin Kaufman; Joellen Watson Hawkins; Loretta P. Higgins; Alice Howell Friedman. Isis, 80(2): 358–359, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234665.

Porter:1989:BRBb

[2781] Charlotte M. Porter. Book review: Natural History and the New World, 1524–1770: An Annotated Bibliography of Printed Materials in the Library of the American Philosophical Society by Anita Guerrini. Isis, 80 (2):359–360, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234666.

Freudenthal:1989:BRB

[2782] Gad Freudenthal. Book review: Dictionnaire des philosophes by Denis Huisman. Isis, 80(2):360–361, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234667.

Rousseau:1989:BRBb

[2783] G. S. Rousseau. Book review: The Relations of Literature and Science: An Annotated Bibliography of Scholarship, 1880–1980 by Walter Schatzberg; Ronald A. Waite; Jonathan K. Johnson. Isis, 80(2):361– 363, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234668.

Anonymous:1989:BMc

[2784] Anonymous. Back matter. Isis, 80(2):364–368, June 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234669.

Anonymous:1989:FMd

[2785] Anonymous. Front matter. Isis, 80(3):369–373, September 1989. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234931.

Apple:1989:PUR

[2786] Rima D. Apple. Patenting University research: Harry steenbock and the Wisconsin alumni research foundation. Isis, 80(3):374–394, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234932.

Smith:1989:CNA

[2787] Robert W. Smith. The Cambridge network in action: The discovery of Neptune. Isis, 80(3):395–422, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234933.

Newman:1989:TAD

[2788] William Newman. Technology and alchemical debate in the late middle ages. Isis, 80(3):423–445, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234934.

Fancher:1989:GEU

[2789] Raymond E. Fancher. Galton on examinations: An unpublished step in the invention of correlation. Isis, 80(3):446–455, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234935.

Meyer:1989:GCC

[2790] Eric Meyer. Galileo’s cosmogonical calculations. Isis, 80(3):456–468, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234936.

Richards:1989:AMH

[2791] Joan L. Richards, Shirley A. Roe, Michael M. Sokal, Albert Moyer, and William A. Wallace. Annual meeting of the History of Science Society, 27–30 December 1988. Isis, 80(3):469–478, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234937.

Anonymous:1989:CHS

[2792] Anonymous. Contributors to the History of Science Society. Isis, 80 (3):478–480, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234938.

Hannaway:1989:EWC

[2793] Owen Hannaway. Eloge: William Coleman (2 July 1934–29 April 1988). Isis, 80(3):480–484, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234939.

Masani:1989:LE

[2794] P. Masani, Steve J. Heims, Joel S. Schwartz, Owen Gingerich, and Robert Westman. Letters to the Editor. Isis, 80(3):485–487, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234940.

Shank:1989:EOW

[2795] Michael H. Shank. Editorial: Old wine in new wineskins. Isis, 80(3): 488–490, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234941.

Seaborg:1989:BRB

[2796] Glenn T. Seaborg. Book review: China Builds the Bomb. Isis, 80(3): 491–493, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234942.

Schuster:1989:BRB

[2797] John A. Schuster. Book review: Changing Order: Replication and Induction in Scientific Practice. Isis, 80(3):493–496, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234943.

Barnes:1989:BRB

[2798] Barry Barnes. Book review: Science as Power: Discourse and Ideology in Modern Society by Stanley Aronowitz. Isis, 80(3):497–498, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234944.

Barrett:1989:BRB

[2799] Thomas M. Barrett. Book review: Foucault, Marxism, and History: Mode of Production versus Mode of Information by Mark Poster. Isis, 80 (3):498–499, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234945.

Steffens:1989:BRB

[2800] Henry Steffens. Book review: Science and Culture in the Western Tradition: Sources and Interpretations by John G. Burke. Isis, 80(3):499, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234946.

Eardley:1989:BRB

[2801] Steve Eardley. Book review: Signatura rerum: Segni, magia e conoscenza de Paracelso a Leibniz by Massimo Luigi Bianchi. Isis, 80(3):500, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234947.

Shapiro:1989:BRB

[2802] Alan E. Shapiro. Book review: Great Scientists of Old as Heretics in “The Scientific Method” by C. Truesdell. Isis, 80(3):500–501, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234948.

Lafollette:1989:BRB

[2803] Marcel C. Lafollette. Book review: How Superstition Won and Science Lost: Popularizing Science and Health in the United States by John C. Burnham. Isis, 80(3):501–502, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234949.

Polanco:1989:BRB

[2804] Xavier Polanco. Book review: Histoire des sciences et des savants depuis deux siècles, d’après l’opinion des principales académies ou sociétés scientifiques by Alphonse de Candolle. Isis, 80(3):502–503, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234950.

Osborne:1989:BRB

[2805] Michael A. Osborne. Book review: Measures and Men by Witold Kula; Richard Szreter. Isis, 80(3):504–505, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234951.

Srinivasan:1989:BRBb

[2806] Raman Srinivasan. Book review: Organizing for Science: The Making of an Industrial Research Laboratory by Shiv Visvanathan. Isis, 80(3): 505–506, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234952.

vonEngelhardt:1989:BRB

[2807] Dietrich von Engelhardt. Book review: Wissenschaftsgeschichte heute: Ansprachen und wissenschaftliche Vorträge zum 25-jährigen Bestehen des Instituts für Geschichte der Naturwissenschaften, Mathematik und Technik der Universität Hamburg by Christian Hünemörder. Isis, 80(3): 506, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234953.

Low:1989:BRB

[2808] Reinhard Löw. Book review: Beobachtung, Experiment und Theorie in der Naturwissenschaft und Medizin by Reimar Lüst; Eberhard Buchborn; Klaus Dransfeld; Hans Musso; Rudolf K. Thauer; Jan Tesing; Heinz Gibian. Isis, 80(3):506–507, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234954.

Toth:1989:BRB

[2809] Imre Toth. Book review: L’aventure des parallèles: Préhistoire de la géométrie non euclidienne: Précurseurs et attardés by Jean-Claude Pont. Isis, 80(3):507–510, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234955.

Carozzi:1989:BRB

[2810] Albert V. Carozzi. Book review: Histoire de la géologie. Volume I: Des Anciens à la première moitié du XVIIe siècle by François Ellenberger. Isis, 80(3):510–511, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234956.

Dorst:1989:BRB

[2811] Jean Dorst. Book review: An International History of Mammalogy. Volume I: Eastern Europe and Fennoscandia, I by Keir B. Sterling. Isis, 80 (3):511–512, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234957.

Kuklick:1989:BRB

[2812] Henrika Kuklick. Book review: Between Literature and Science: The Rise of Sociology by Wolf Lepenies. Isis, 80(3):512–513, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234958.

King:1989:BRBb

[2813] Peter King. Book review: Der Begriff der praktischen Vernunft nach Johannes Buridanus by Gerhardt Krieger. Isis, 80(3):513–515, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234959.

Saliba:1989:BRBa

[2814] George Saliba. Book review: Oeuvres mathématiques: Algèbre et géométrie au XIIe siècle by Sharaf al-Din al-T .usi; Roshdi Rashed. Isis, 80 (3):515–517, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234960.

Daston:1989:BRB

[2815] Lorraine Daston. Book review: Capitalism and Arithmetic: The New Math of the Fifteenth Century by Frank J. Swetz. Isis, 80(3):517–518, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234961.

Hilfstein:1989:BRB

[2816] Erna Hilfstein. Book review: Optyka w XV wieku: Miedzy nauka ’sredniowieczna a nowozytna (Fifteenth-Century Optics: Between Medieval and Modern Science) by Grazyna Rosińska. Isis, 80(3):518–519, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234962.

Ellenius:1989:BRB

[2817] Allan Ellenius. Book review: Die Kunst und das Studium der Natur vom 14. zum 16. Jahrhundert by Wolfram Prinz; Andreas Beyer. Isis, 80 (3):519–520, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234963.

Wear:1989:BRB

[2818] Andrew Wear. Book review: Avicenna in Renaissance Italy: The Canon and Medical Teaching in Italian Universities after 1500 by Nancy G. Siraisi. Isis, 80(3):520–521, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234964.

Getz:1989:BRB

[2819] Faye Marie Getz. Book review: Rezeption und Rolle der Heilkunde in landessprachigen handschriftlichen Kompendien des Spatmittelalters by Volker Zimmermann. Isis, 80(3):521–522, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234965.

Dorn:1989:BRB

[2820] Harold Dorn. Book review: Leonardo da Vinci: Engineer and Architect by Paolo Galluzzi. Isis, 80(3):522–524, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234966.

Guerrini:1989:BRB

[2821] Anita Guerrini. Book review: Time, the Greatest Innovator: Timekeeping and Time Consciousness in Early Modern Europe by Rachel Doggett; Susan Jaskot; Robert Rand. Isis, 80(3):524–525, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234967.

Burnett:1989:BRB

[2822] C. S. F. Burnett. Book review: Peter Apian und Azophi: Arabische Sternbilder in Ingolstadt im fruhen 16. Jahrhundert by Paul Kunitzsch.

Isis, 80(3):525–526, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234968.

Ron:1989:BRB

[2823] Moshe Ron. Book review: Jews in the Origins of Modern Science and Bacon’s Scientific Utopia: The Life and Work of Joachim Gaunse, Mining Technologist and First Recorded Jew in English-Speaking North America by Lewis S. Feuer. Isis, 80(3):526, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234969.

Applebaum:1989:BRBb

[2824] Wilbur Applebaum. Book review: La structure poétique du monde: Copernic, Kepler by Fernand Hallyn. Isis, 80(3):526–528, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234970.

Struik:1989:BRB

[2825] Dirk Struik. Book review: Tot Heil van’t Menschdom: Culturele genootschappen in Nederland, 1750–1815 by W. W. Mijnhardt. Isis, 80 (3):528–529, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234971.

Calinger:1989:BRBb

[2826] Ronald Calinger. Book review: Immanuel Kant: His Life and Thought by Arsenij Gulyga; Marijan Despalatović. Isis, 80(3):529–531, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234972.

Kors:1989:BRB

[2827] Alan Charles Kors. Book review: Qeuvres philosophiques by Julien Offray de La Mettrie. Isis, 80(3):531, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234973.

Proctor:1989:BRB

[2828] Robert N. Proctor. Book review: Weber, Irrationality, and Social Order by Alan Sica. Isis, 80(3):532–533, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234974.

Porter:1989:BRBc

[2829] Theodore M. Porter. Book review: L’économiste, la science, et la pouvoir: Le cas Walras by Hervé Dumez. Isis, 80(3):533–534, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234975.

Johnson:1989:BRBa

[2830] Jeffrey A. Johnson. Book review: Technische Bildung zwischen Wissenschaft und Praxis: Die Technische Hochschule Braunschweig 1862– 1914 by Helmuth Albrecht. Isis, 80(3):534–535, September 1989. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234976.

Schneider:1989:BRB

[2831] Ivo Schneider. Book review: Maxwell on Molecules and Gases by Elizabeth Garber; Stephen G. Brush; C. W. F. Everitt. Isis, 80(3):535–536, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234977.

Koblitz:1989:BRB

[2832] Ann Hibner Koblitz. Book review: Aleksandr Porfir’ evich Borodin: A Chemist’s Biography by Nikolai A. Figurovskii; Yurii I. Solov’ev; Charlene Steinberg; George B. Kauffman. Isis, 80(3):536–537, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234978.

Good:1989:BRB

[2833] Gregory A. Good. Book review: U.S. Coast Survey vs. Naval Hydrographic Office: A Nineteenth-Century Rivalry in Science and Politics by Thomas G. Manning. Isis, 80(3):537–538, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234979.

Browne:1989:BRB

[2834] Janet Browne. Book review: The Southern Ark: Zoological Discovery in New Zealand, 1769–1900 by J. R. H. Andrews. Isis, 80(3):538–539, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234980.

Henson:1989:BRB

[2835] Pamela M. Henson. Book review: Asa Fitch and the Emergence of American Entomology: With an Entomological Bibliography and a Catalog of Taxonomic Names and Type Specimens by Jeffrey K. Barnes. Isis, 80 (3):539–540, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234981.

Arens:1989:BRB

[2836] Katherine Arens. Book review: J. E. Purkyne and Psychology: With a Focus on Unpublished Manuscripts by Josef Brožek; Jirí Hoskovec. Isis, 80(3):541–542, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234982.

McGraw:1989:BRB

[2837] Donald J. McGraw. Book review: Essays of Robert Koch by Robert Koch; K. Codell Carter. Isis, 80(3):542–543, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234983.

Hickel:1989:BRB

[2838] Erika Hickel. Book review: Die Pharmazie im Spannungsfeld der Neuorientierung: Philipp Lorenz Geiger (1785–1836), Leben, Werk und Wirken: Eine Biographie by Ulrike Thomas. Isis, 80(3):543, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234984.

Badash:1989:BRB

[2839] Lawrence Badash. Book review: March to Armageddon: The United States and the Nuclear Arms Race, 1939 to the Present by Ronald E. Powaski. Isis, 80(3):543–544, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234985.

Hirsh:1989:BRB

[2840] Richard F. Hirsh. Book review: The Realities of Nuclear Power: International Economic and Regulatory Experience by S. D. Thomas. Isis, 80 (3):544–545, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234986.

Levin:1989:BRBa

[2841] Aleksey E. Levin. Book review: Science between the Superpowers by Yakov M. Rabkin. Isis, 80(3):545–546, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234987.

Crowe:1989:BRB

[2842] Gregory D. Crowe. Book review: The Status of Soviet Civil Science by Craig Sinclair. Isis, 80(3):546, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234988.

Lawrence:1989:BRBb

[2843] Christopher Lawrence. Book review: Horace Darwin’s Shop: A History of the Cambridge Scientific Instrument Company, 1878–1968 by M. J. G. Cattermole; A. F. Wolfe. Isis, 80(3):546–547, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234989.

Stuewer:1989:BRB

[2844] Roger H. Stuewer. Book review: Fifty Years of Neutron Diffraction: The Advent of Neutron Scattering by G. E. Bacon. Isis, 80(3):547, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234990.

Jones:1989:BRBc

[2845] Martin R. Jones. Book review: Quantum Concepts in Space and Time by R. Penrose; C. J. Isham. Isis, 80(3):547–549, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234991.

Josephson:1989:BRB

[2846] Paul Josephson. Book review: Razvitie fundamental’nykh printsipov statisticheskoi fisiki v pervoi polovine XX veka. (The Development of Fundamental Principles of Statistic Physics in the First Half of the Twentieth Century) by N. V. Vdovichenko; Edinye teorii poliia v pervoi treti XX veka. (Unified Field Theories in the First Third of the Twentieth Century) by V. P. Vizgin. Isis, 80(3):549–550, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234992.

Roll-Hansen:1989:BRB

[2847] Nils Roll-Hansen. Book review: Race Hygiene and National Efficiency: The Eugenics of Wilhelm Schallmayer by Sheila Faith Weiss. Isis, 80 (3):550–551, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234993.

Bates:1989:BRBc

[2848] Barbara Bates. Book review: Below the Magic Mountain: A Social History of Tuberculosis in Twentieth–Century Britain by Linda Bryder. Isis, 80(3):551–552, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234994.

Howell:1989:BRBb

[2849] Joel D. Howell. Book review: American Pediatrics: The Social Dynamics of Professionalism, 1880–1980 by Sydney A. Halpern. Isis, 80(3):552– 553, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234995.

Paradowski:1989:BRB

[2850] Robert J. Paradowski. Book review: The Roots of Molecular Medicine: A Tribute to Linus Pauling by Richard P. Huemer. Isis, 80(3):553–554, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234996.

Roland:1989:BRB

[2851] Charles G. Roland. Book review: The Nazi Doctors: Medical Killing and the Psychology of Genocide by Robert Jay Lifton. Isis, 80(3):555–556, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234997.

Pang:1989:BRB

[2852] Alex Soojung-Kim Pang. Book review: Radar in World War II by Henry E. Guerlac. Isis, 80(3):556–557, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234998.

Leary:1989:BRB

[2853] William M. Leary. Book review: The Airway to Everywhere: A History of All American Aviation, 1937–1953 by W. David Lewis; William F. Trimble. Isis, 80(3):557–558, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234999.

Braun:1989:BRB

[2854] Hans-Joachim Braun. Book review: Ingenieure und Naturwissenschaftler in der industriellen Forschung und Entwicklung: Berufliche Situation und Perspektiven by Hugo Kossbiel; Arno Bammé; Bernd Martens.

Isis, 80(3):558–559, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235000.

Numbers:1989:BRB

[2855] Ronald L. Numbers. Book review: American Medical Schools and the Practice of Medicine: A History by William G. Rothstein. Isis, 80(3): 559–560, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235001.

Morantz-Sanchez:1989:BRBa

[2856] Regina Morantz-Sanchez. Book review: Disease and Representation: Images of Illness from Madness to AIDS by Sander L. Gilman. Isis, 80 (3):560–561, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235002.

Hoffman:1989:BRB

[2857] Louise E. Hoffman. Book review: The Complete Letters of Sigmund Freud to Wilhelm Fliess, 1887–1904 by Sigmund Freud; Jeffrey Moussaieff Masson. Isis, 80(3):561–562, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235003.

Pressman:1989:BRB

[2858] Jack D. Pressman. Book review: Split Minds, Split Brains: Historical and Current Perspectives by Jacques M. Quen. Isis, 80(3):562–563, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235004.

Kranzberg:1989:BRB

[2859] Melvin Kranzberg. Book review: The History of Techniques by Bertrand Gille. Isis, 80(3):563–564, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235005.

Lankton:1989:BRB

[2860] Larry Lankton. Book review: Innovation and the Rise of the Tunnelling Industry by Graham West. Isis, 80(3):564–565, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235006.

Hyde:1989:BRB

[2861] Charles K. Hyde. Book review: Mining America: The Industry and the Environment, 1800–1980 by Duane A. Smith. Isis, 80(3):565–566, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235007.

Bugos:1989:BRBa

[2862] Glenn Bugos. Book review: Technology, the Economy, and Society: The American Experience by Joel Colton; Stuart Bruchey. Isis, 80(3):566– 567, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235008.

Johnson:1989:BRBb

[2863] Mark Johnson. Book review: Metaphor by David E. Cooper. Isis, 80 (3):567–568, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235009.

Maienschein:1989:BRB

[2864] Jane Maienschein. Book review: Toward a New Philosophy of Biology: Observations of an Evolutionist by Ernst Mayr. Isis, 80(3):568–569, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235010.

Sawyier:1989:BRB

[2865] Fay Horton Sawyier. Book review: Forbidden Knowledge, and Other Essays in the Philosophy of Cognition by Nicholas Rescher. Isis, 80(3): 569–570, September 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235011.

Fores:1989:BRB

[2866] Michael Fores. Book review: Theories of Industrial Society by Richard J. Badham. Isis, 80(3):570–571, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235012.

Hiddinga:1989:BRB

[2867] Anja Hiddinga. Book review: The Woman in the Body: A Cultural Analysis of Reproduction by Emily Martin. Isis, 80(3):571–572, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235013.

Williams:1989:BRBb

[2868] L. Pearce Williams. Book review: The History of Modern Science: A Guide to the Second Scientific Revolution, 1800–1950 by Stephen G. Brush. Isis, 80(3):572–573, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235014.

Levere:1989:BRB

[2869] Trevor H. Levere. Book review: Science and Technology in Canadian History: A Bibliography of Primary Sources to 1914 by R. A. Richardson; B. H. MacDonald. Isis, 80(3):573, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235015.

Anonymous:1989:Ca

[2870] Anonymous. Collections. Isis, 80(3):574–578, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235016.

Anonymous:1989:BMd

[2871] Anonymous. Back matter. Isis, 80(3):579–584, September 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235017.

Anonymous:1989:FMe

[2872] Anonymous. Front matter. Isis, 80(4):585–591, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234173.

Browne:1989:BGE

[2873] Janet Browne. Botany for gentlemen: Erasmus Darwin and “The Loves of the Plants”. Isis, 80(4):592–621, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234174.

Krause:1989:TTD

[2874] David J. Krause. Testing a tradition: Douglass Houghton and the native copper of Lake Superior. Isis, 80(4):622–639, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234175.

Cueto:1989:ABP

[2875] Marcos Cueto. Andean biology in peru: Scientific styles on the periphery. Isis, 80(4):640–658, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234176.

Freudenthal:1989:EJB

[2876] Gad Freudenthal and J. L. Heilbron. Eloge: Joseph Ben-David, 19 August 1920–12 January 1986. Isis, 80(4):659–663, December 1989. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234177.

Rochow:1989:LE

[2877] Eugene G. Rochow and George Wise. Letters to the Editor. Isis, 80(4): 664, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234178.

Nickles:1989:BRB

[2878] Thomas Nickles. Book review: Scrutinizing Science: Empirical Studies of Scientific Change. Isis, 80(4):665–669, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234179.

Richards:1989:BRB

[2879] Joan L. Richards. Book review: New Directions in the Philosophy of Mathematics: An Anthology. Isis, 80(4):669–672, December 1989. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234180.

Bensaude-Vincent:1989:BRB

[2880] Bernadette Bensaude-Vincent. Book review: D’Archimède à Einstein: Les faces cachées de l’invention scientifique by Pierre Thuillier; Les passions du savoir: Essais sur les dimensions culturelles de la science by Pierre Thuillier. Isis, 80(4):673–674, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234181.

Edge:1989:BRB

[2881] David Edge. Book review: Scientific Controversies: Case Studies in the Resolution and Closure of Disputes in Science and Technology by H. Tristram Engelhardt,; Arthur L. Caplan. Isis, 80(4):675–676, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234182.

Greenaway:1989:BRB

[2882] Frank Greenaway. Book review: Museums of Influence by Kenneth Hudson. Isis, 80(4):676, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234183.

Golinski:1989:BRB

[2883] Jan Golinski. Book review: Shaping Written Knowledge: The Genre and Activity of the Experimental Article in Science by Charles Bazerman. Isis, 80(4):676–678, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234184.

Dobbs:1989:BRB

[2884] B. J. T. Dobbs. Book review: Newton and the New Direction in Science by G. V. Coyne, S.J.; M. Heller; J. Zyciński. Isis, 80(4):678–679, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234185.

Lawrence:1989:BRBc

[2885] Christopher Lawrence. Book review: Essays on the Rise and Decline of Bedside Medicine by Mark D. Altschule. Isis, 80(4):679, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234186.

Morantz-Sanchez:1989:BRBb

[2886] Regina Morantz-Sanchez. Book review: Fasting Girls: The Emergence of Anorexia Nervosa as a Modern Disease by Joan Jacobs Brumberg. Isis, 80(4):679–680, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234187.

Melosh:1989:BRBb

[2887] Barbara Melosh. Book review: Photographing Medicine: Images and Power in Britain and America since 1840 by Daniel M. Fox; Christopher Lawrence. Isis, 80(4):681–682, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234188.

Brunton:1989:BRBb

[2888] Deborah C. Brunton. Book review: Disease, Medicine, and Society in England, 1550–1860 by Roy Porter. Isis, 80(4):682, December 1989.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234189.

Wachhorst:1989:BRB

[2889] Wyn Wachhorst. Book review: Seeking Many Inventions: The Idea of Community in America by Philip Abbott. Isis, 80(4):682–683, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234190.

Anderson:1989:BRB

[2890] R. G. W. Anderson. Book review: The Past Is a Foreign Country by David Lowenthal. Isis, 80(4):683–684, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234191.

Berggren:1989:BRB

[2891] J. L. Berggren. Book review: Mathematics and Measurement by O. A. W. Dilke. Isis, 80(4):684–685, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234192.

Ferngren:1989:BRB

[2892] Gary B. Ferngren. Book review: From Democedes to Harvey: Studies in the History of Medicine by Vivian Nutton. Isis, 80(4):685–686, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234193.

Phillips:1989:BRB

[2893] William D. Phillips, Jr. Book review: Geography, Technology, and War: Studies in the Maritime History of the Mediterranean, 649–1571 by John H. Pryor. Isis, 80(4):686–687, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234194.

Livesey:1989:BRB

[2894] Steven J. Livesey. Book review: The Cambridge History of Later Medieval Philosophy: From the Rediscovery of Aristotle to the Disintegration of Scholasticism, 1100–1600 by Norman Kretzman; Anthony Kenny; Jan Pinborg. Isis, 80(4):687–688, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234195.

Saliba:1989:BRBb

[2895] George Saliba. Book review: Qeuvres d’astronomie by Thabit ibn Qurra; Régis Morelon. Isis, 80(4):688–690, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234196.

Kennedy:1989:BRB

[2896] E. S. Kennedy. Book review: In den Gärten der Wissenschaft by Al-Biruni; Gotthard Strohmaier. Isis, 80(4):690–691, December 1989. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234197.

North:1989:BRB

[2897] John D. North. Book review: Innovationen im Quadrivium des 10. und 11. Jahrhunderts: Studien zur Einführung von Astrolab und Abakus im lateinischen Mittelalter by Werner Bergmann. Isis, 80(4):691, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234198.

Goddu:1989:BRB

[2898] André Goddu. Book review: Tractatus de indivisibilibus by Adam de Wodeham; Rega Wood. Isis, 80(4):691–693, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234199.

Ovitt:1989:BRB

[2899] George Ovitt, Jr. Book review: “Unless You Believe, You Shall Not Understand”: Logic, University, and Society in Late Medieval Vienna by Michael H. Shank. Isis, 80(4):693–694, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234200.

Puig:1989:BRB

[2900] Roser Puig. Book review: El “Lunari” de Bernat de Granollachs: Alguns aspectes de la història de l’astronomia a la Catalunya del quatre-cents by Josep Chabàs; Antoni Roca. Isis, 80(4):694, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234201.

Hannaway:1989:BRB

[2901] Owen Hannaway. Book review: Vom Bergkwerck XII Bücher by Georg Agricola. Isis, 80(4):694–696, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234202.

Hutchison:1989:BRB

[2902] Keith Hutchison. Book review: Hermeticism and the Renaissance: Intellectual History and the Occult in Early Modern Europe by Ingrid Merkel; Allen G. Debus. Isis, 80(4):696–697, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234203.

Levin:1989:BRBb

[2903] Aleksey E. Levin. Book review: Nauchnaya revolutsia XVII veka by V. S. Kirsanov. Isis, 80(4):697–698, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234204.

Feigenbaum:1989:BRB

[2904] Lenore Feigenbaum. Book review: Der Briefwechsel von Johann I Bernoulli. Volume II: Der Briefwechsel mit Pierre Varignon, Erster Teil: 1692–1702 by Johann I Bernoulli; Pierre Costabel; Jeanne Peiffer. Isis, 80(4):698–699, December 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234205.

Yoder:1989:BRB

[2905] Joella Yoder. Book review: Pascal’s Arithmetical Triangle by A. W. F. Edwards. Isis, 80(4):699–700, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234206.

Westfall:1989:BRB

[2906] Richard S. Westfall. Book review: Mathematische Grundlagen der Naturphilosophie by Isaac Newton; Ed Dellian. Isis, 80(4):700–701, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234207.

Henry:1989:BRBb

[2907] John Henry. Book review: Millenarianism and Messianism in English Literature and Thought, 1650–1800 by Richard H. Popkin. Isis, 80(4): 701–702, December 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234208.

Vaughn:1989:BRB

[2908] Karen I. Vaughn. Book review: Before Adam Smith: The Emergence of Political Economy, 1662–1776 by Terence Hutchison. Isis, 80(4):702– 703, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234209.

Vucinich:1989:BRB

[2909] Alexander Vucinich. Book review: Rud– der Bosković by Zarko Dadić. Isis, 80(4):703–704, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234210.

Crosland:1989:BRBb

[2910] Maurice Crosland. Book review: Science, Medicine, and Dissent: Joseph Priestley (1733–1804) by R. G. W. Anderson; Christopher Lawrence. Isis, 80(4):705–706, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234211.

vonStaden:1989:BRBc

[2911] Heinrich von Staden. Book review: Las epidemias de Hippócrates: Con observaciones prácticas de los antiquos y modernos by Andrés Piquer. Isis, 80(4):706–707, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234212.

Wood:1989:BRB

[2912] P. B. Wood. Book review: David Hume: Philosopher of Moral Science by Antony Flew. Isis, 80(4):707–708, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234213.

Calinger:1989:BRBc

[2913] Ronald Calinger. Book review: Kant’s Philosophy of Physical Science: Metaphysische Anfangsgründe der Naturwissenschaft, 1786–1986 by Robert E. Butts. Isis, 80(4):708–709, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234214.

Atchley:1989:BRB

[2914] Alan Aubrey Atchley. Book review: Plantas y luces en México: La Real Expedición Científica a Nueva España (1787–1803) by Xavier Lozoya. Isis, 80(4):709–710, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234215.

Burwick:1989:BRB

[2915] Frederick Burwick. Book review: Goethe contra Newton: Polemics and the Project for a New Science of Color by Dennis L. Sepper. Isis, 80 (4):710–711, December 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234216.

DeVorkin:1989:BRB

[2916] David H. DeVorkin. Book review: Elites in Conflict: The Antebellum Clash over the Dudley Observatory by Mary Ann James. Isis, 80(4): 711–712, December 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234217.

Hunt:1989:BRBb

[2917] Bruce J. Hunt. Book review: Oliver Heaviside, the Man by G. F. C. Searle; Ivor Catt. Isis, 80(4):712, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234218.

Wise:1989:BRB

[2918] M. Norton Wise. Book review: Kelvin and Stokes: A Comparative Study in Victorian Physics by David B. Wilson. Isis, 80(4):712–713, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234219.

Olby:1989:BRB

[2919] Robert Olby. Book review: A Century since Mendel by M. Campbell. Isis, 80(4):713–714, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234220.

Beer:1989:BRB

[2920] Gillian Beer. Book review: Darwin and the Novelists: Patterns of Science in Victorian Fiction by George Levine. Isis, 80(4):714–715, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234221.

Roberts:1989:BRB

[2921] Jon H. Roberts. Book review: Darwin’s Forgotten Defenders: The Encounter between Evangelical Theology and Evolutionary Thought by David N. Livingstone. Isis, 80(4):715–717, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234222.

Furner:1989:BRB

[2922] Mary O. Furner. Book review: Breaking the Academic Mould: Economists and American Higher Learning in the Nineteenth Century by William J. Barber. Isis, 80(4):717–718, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234223.

LaBerge:1989:BRBb

[2923] Ann F. La Berge. Book review: Death in Hamburg: Society and Politics in the Cholera Years, 1830–1910 by Richard J. Evans. Isis, 80(4):718– 719, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234224.

Radder:1989:BRB

[2924] Hans Radder. Book review: H. A. Kramers: Between Tradition and Revolution by M. Dresden. Isis, 80(4):719–720, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234225.

Hoch:1989:BRB

[2925] Paul K. Hoch. Book review: The Making of Physicists by Rajkumari Williamson. Isis, 80(4):720–721, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234226.

Weininger:1989:BRB

[2926] Stephen J. Weininger. Book review: Molecules in Physics, Chemistry, and Biology. Volume I: General Introduction to Molecular Sciences by Jean Maruani. Isis, 80(4):722, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234227.

Kater:1989:BRB

[2927] Michael H. Kater. Book review: Murderous Science: Elimination by Scientific Selection of Jews, Gypsies, and Others, Germany 1933–1945 by Benno Muller-Hill; George Fraser. Isis, 80(4):722–723, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234228.

Kenney:1989:BRB

[2928] Martin Kenney. Book review: Biotechnology: Strategies for Life by Elizabeth Antébi; David Fishlock. Isis, 80(4):723–724, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234229.

Bugos:1989:BRBb

[2929] Glenn E. Bugos. Book review: The Biotechnology Revolution: An International Perspective by Alan M. Russell. Isis, 80(4):724–725, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234230.

Koch:1989:BRB

[2930] Ellen B. Koch. Book review: Worse than the Disease: Pitfalls of Medical Progress by Diana B. Dutton; Thomas A. Preston; Nancy E. Pfund. Isis, 80(4):725–726, December 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234231.

Rogers:1989:BRB

[2931] Naomi Rogers. Book review: Great and Desperate Cures: The Rise and Decline of Psychosurgery and Other Radical Treatments for Mental Illness by Elliot S. Valenstein. Isis, 80(4):726–728, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234232.

Reed:1989:BRB

[2932] James Reed. Book review: Birth Control in Germany, 1871–1933 by James Woycke. Isis, 80(4):728, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234233.

Wright:1989:BRB

[2933] David E. Wright. Book review: Farming for Fuel: The Political Economy of Energy Sources in the United States by Folke Dovring. Isis, 80(4):728– 729, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234234.

Wohlforth:1989:BRB

[2934] William C. Wohlforth. Book review: Innovation and the Arms Race: How the United States and the Soviet Union Develop New Military Technologies by Matthew Evangelista. Isis, 80(4):729–730, December 1989.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234235.

McDougall:1989:BRB

[2935] Walter A. McDougall. Book review: Harnessing the Genie: Science and Technology Forecasting for the Air Force, 1944–1986 by Michael H. Gorn. Isis, 80(4):730–731, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234236.

Jakab:1989:BRB

[2936] Peter L. Jakab. Book review: Test Pilots: The Frontiersmen of Flight by Richard P. Hallion. Isis, 80(4):731–732, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234237.

Bud:1989:BRB

[2937] Robert Bud and W. J. Reader. Book review: Science and Corporate Strategy: Du Pont R & D, 1902–1980 by David A. Hounshell; John Kenly Smith. Isis, 80(4):732–734, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234238.

Lowood:1989:BRB

[2938] Henry Lowood. Book review: Science Parks and the Growth of High Technology Firms by C. S. P. Monck; R. B. Porter; P. R. Quintas; D. J. Storey; P. Wynarczyk. Isis, 80(4):734–735, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234239.

Rosenbrock:1989:BRB

[2939] H. H. Rosenbrock. Book review: Forces of Production: A Social History of Industrial Automation by David F. Noble. Isis, 80(4):735, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234240.

Masani:1989:BRB

[2940] P. R. Masani. Book review: The Human Use of Human Beings: Cybernetics and Society by Norbert Wiener. Isis, 80(4):735–737, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234241.

Sigurdsson:1989:BRB

[2941] Skuli Sigurdsson. Book review: The Fourth Dimension and Non– Euclidean Geometry in Modern Art by Linda Dalrymple Henderson. Isis, 80(4):737–738, December 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234242.

Caron:1989:BRB

[2942] Joseph Caron. Book review: Fondements de l’histoire naturelle: Pour une anthropologie de la science by Scott Atran. Isis, 80(4):738–739, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234243.

Turner:1989:BRBb

[2943] Stephen Turner. Book review: Methodology and Epistemology for Social Science: Selected Papers by Donald T. Campbell; E. Samuel Overman. Isis, 80(4):739–740, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234244.

Rothschild:1989:BRB

[2944] Joan Rothschild. Book review: Technology and Responsibility by Paul T. Durbin. Isis, 80(4):740–741, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234245.

Moser:1989:BRB

[2945] Paul K. Moser. Book review: Space-Perception and the Philosophy of Science by Patrick A. Heelan. Isis, 80(4):741–742, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234246.

Weindling:1989:BRB

[2946] Paul Weindling. Book review: Rassenmythos und Sozialwissenschaften in Deutschland: Ein verdrängtes Kapitel sozialwissenschaftlicher Wirkungsgeschichte by Carsten Klingemann. Isis, 80(4):742–744, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234247.

Porter:1989:BRBd

[2947] Theodore M. Porter. Book review: Government and Expertise: Specialists, Administrators, and Professionals by Roy MacLeod. Isis, 80 (4):744–745, December 1989. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234248.

Laudan:1989:BRB

[2948] Larry Laudan. Book review: Scientific Realism: A Critical Reappraisal by Nicholas Rescher. Isis, 80(4):745–746, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234249.

Elliott:1989:BRBb

[2949] Clark A. Elliott. Book review: The Reader’s Adviser: A Layman’s Guide to Literature. Volume V: The Best in the Literature of Science, Technology, and Medicine by Paul T. Durbin; The Reader’s Adviser: A Layman’s Guide to Literature. Volume VI: Indexes by Barbara A. Chernow; George A. Vallasi. Isis, 80(4):746–747, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234250.

Ellioti:1989:BRB

[2950] Eric Ellioti. Book review: Archiv zur Geschichte der Max-Planck-Gesellschaft by Ulrike Emrich; Michael Globig. Isis, 80(4):747, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234251.

Anonymous:1989:Cb

[2951] Anonymous. Collections. Isis, 80(4):747–758, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234252.

Anonymous:1989:BMe

[2952] Anonymous. Back matter. Isis, 80(4):759–786, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234253.

Anonymous:1989:VI

[2953] Anonymous. Volume information. Isis, 80(4):770–785, December 1989. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234172.

Wheaton:1983:TSE

[2954] Bruce R. Wheaton. The tiger and the shark: empirical roots of waveparticle dualism. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-25098-6. xxiv + 355 pp. LCCN QC476.W38 W45 1983.

URL http://www.loc.gov/catdir/description/cam022/82022069.
html;   http://www.loc.gov/catdir/toc/cam023/82022069.html.